A Complete Bibliography of Publications in the Journal of Cell Biology: 2010–2014

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

26 December 2021 Version 1.01

Title word cross-reference

1 [APV^{+}12, BMFC^{+}11, BMFC^{+}13, FRL^{+}13, GFSR11, PSVRB^{+}11]. 13 [FSA^{+}10b, WBMCSS13]. 2 [OSD^{+}14, WGR^{+}12]. 3 [AKA^{+}13, BCB^{+}14b, BCBG10, BSO^{+}14, DCP^{+}10, ECC^{+}13, HSI^{+}14, HGV^{+}14, KSW^{+}11, LMS^{+}10c, MHAK^{+}12, PGCY12, Sho13b, TB12, WZHV11, YYM^{+}11]. 4 [HMB14, dSJDD^{+}11]. 5 [NPL^{+}10]. 5′ → 3′ [YTM^{+}11]. ^{+} [HA12]. ^{+/}− [LWBH12]. ^{−/}− [BMS^{+}11]. ^{2+} [ADF^{+}12, BNM^{+}14, CZD^{+}13, DAB^{+}11, DSP11, DS10, FCA10, IIN^{+}11, KBC^{+}14, KSLF^{+}11, MSR10, MDW^{+}13, SPC^{+}13, SLH^{+}14, WEK^{+}14, YSM10, vBAK^{+}12]. ^{Cdc20} [COW13]. ^{Cdc55} [KST^{+}11, VCF^{+}13]. ^{Cdh1} [COW13, CRJB^{+}11, EMO12, HZT^{+}12]. ^{glued} [RMF^{+}10]. ^{R} [ZFP^{+}13]. ^{Rts1} [ZDM^{+}14]. ^{Sc} [RKK^{+}14]. ^{Slimb} [BDN^{+}13]. ^{TPR} [WJPD11]. _{1} [EMO12, GRH^{+}12]. _{2} [KPJ^{+}13, Sho12-40, Sho14-45, vGLWB12]. _{3} [DCL^{+}12, DWM^{+}12, VLKI14]. L [bCAH^{+}11, FBR^{+}10, GWR^{+}10]. _{S} [BGC^{+}10]. _{v} [CZD^{+}13, TPZ^{+}14]. α [AiIK^{+}13, BLC^{+}12, BSR^{+}11a, BMRM13, BKY^{+}10, BFG^{+}13, CVR10, COG11, CZD^{+}13, CDAK10a, CDAK10b, DPW^{+}11, DPW^{+}12, FBAO^{+}13, GdBP^{+}14, GRHA^{+}12, HKN^{+}10, JGB^{+}13, JCL^{+}11, KKS^{+}14, KHG^{+}13, KTN^{+}12, KLS^{+}13, KMG^{+}11, KWK^{+}11, LGAC13, MH11, Mit12a, NRM^{+}12, NBC^{+}12, PMB^{+}11, RCM^{+}12, RBH^{+}12, RFK^{+}10, SSV^{+}12, SAG^{+}11, SSB^{+}10, Sho10a, Sho13x, SSK^{+}13, TB12, WCC^{+}10, WHDR^{+}10, YYM^{+}11]. α_{4}β_{7} [YZM^{+}12b]. α_{X} [SYS13]. αβ [BR14]. β [AEC^{+}14, BLC^{+}12, BAY^{+}11, BKAB13, BFG^{+}13, BMFC^{+}11, BMFC^{+}13, CSP^{+}10, DSP11, DCP^{+}10, FRL^{+}13, GBSC^{+}12, GRHA^{+}12, HHJ^{+}11, HAB14, HKN^{+}10, JGB^{+}13, JB12, KYHG12, KTN^{+}12, KIL^{+}12, KCK^{+}14, Les11-34, Les13a, LSW^{+}14, MSR10, MCS^{+}13, NRM^{+}12, NLAS^{+}10, ONH^{+}12, PTST12, PXZ^{+}13, PSF^{+}11, RNS^{+}14, RFL13, RIG^{+}12, RB11, RDB^{+}12, RBH^{+}12, Sho10b, Sho11-27, Sho14-64, VYC^{+}11, WHDR^{+}10, YMU^{+}10, YMU^{+}13, ZGCG^{+}14, ZSZ^{+}13, ZCB^{+}10a, ZCB^{+}10b]. β_{1} [MSR10]. β_{1a} [TPZ^{+}14]. β_{2} [oHXK^{+}12, SYS13]. δ [LMS^{+}10c]. η [BBW^{+}13, DPB^{+}10, GSGL11]. γ [BLO^{+}12, CLS^{+}10, EMO12, ENG^{+}12, GOWM12, GF11, JYRL^{+}13, NEMH^{+}10, Sho10c, Sho12a, Sho12b, UKZ^{+}13]. κ [BLC^{+}12, BWL^{+}13, Les12-29, Les14w, LDN^{+}13, WHA^{+}13]. λ [NSZ^{+}13]. ν [KNH^{+}10]. θ [KNH^{+}10]. ζ [HLN^{+}11].

-2 [NLAS^{+}10]. -AAA [AWB^{+}14]. -actin [GF11]. -Actinin-4 [TB12]. -Actinin-4/ [TB12]. -adaptin [KKS^{+}14]. -adrenergic [oHXK^{+}12]. -Arrestin-dependent [MSR10]. -barrel [KIL^{+}12]. -binding-like [LSW^{+}14]. -Calmodulin [WEK^{+}14, DS10]. -Catenin [Les13a, RFL13, Sho10b, CSP^{+}10, MH11, YMU^{+}10, YMU^{+}13, BAY^{+}11]. -deficient [Sho13b]. -dependent [BMFC^{+}13, ECC^{+}13, WBMCSS13, ZPS^{+}10, BMFC^{+}11, CRJB^{+}11, TB12, VLKI14, vBAK^{+}12]. -dystrobrevin [SAG^{+}11]. -evoked [BNM^{+}14]. -Globin [BKAB13]. -helix-finger [WGR^{+}12]. -independent [LSS^{+}12]. -induced [PTST12]. -kinase [AKA^{+}13, DCP^{+}10, HGV^{+}14, LMS^{+}10c, YYM^{+}11]. -kinases [BCBG10]. -mediated [KLS^{+}13, HZT^{+}12, KST^{+}11, KBC^{+}14, KPJ^{+}13]. -phosphatase [NPL^{+}10]. -phosphate [GFSR11, HMB14, dSJDD^{+}11]. -protofilament [FSA^{+}10b]. -regulated [APV^{+}12, PSVRB^{+}11]. -secretase [JYRL^{+}13, UKZ^{+}13]. -Synuclein [PMB^{+}11, Sho10a, WCC^{+}10, KMG^{+}11]. -Tubulin [EMO12, NEMH^{+}10, Sho10c, Sho12a, Sho12b, BR14, CLS^{+}10, ENG^{+}12].

/calmodulin [KSLF^{+}11, MSR10]. /calmodulin-dependent [KSLF^{+}11]. /Mud [WJPD11]. /Mud-mediated [WJPD11].

1 [BLO^{+}12, BLC^{+}12, BSR^{+}11a, BMRM13, BWBC^{+}14, BBJ^{+}10, COG11, CHL^{+}14, CTL^{+}10, CS13, CWB^{+}14, DGH^{+}14, DGS^{+}10, DSP11, DSD^{+}13, DSW^{+}11, EM11, ECK^{+}12, FTJG13, FS10, GWR^{+}10, GTR^{+}13, HSR^{+}10, HZT^{+}12, JBS^{+}12, JBS^{+}13, JCL^{+}11, KHW^{+}10, KCK^{+}14, Les11f, Les14a, LMC^{+}12, MLG^{+}10, MPM11, NGL^{+}12, NLAS^{+}10, OZ14, OFS^{+}10, PHB^{+}11, PLR^{+}13, PXZ^{+}13, PSF^{+}11, QWL^{+}11, RMG^{+}12, RB11, SSB^{+}10, SRKR10, STI^{+}11, Sho10-61, Sho11w, Sho11-48, Sho13x, Sho14d, SGD^{+}10, TTC^{+}14, WOG13, WMV^{+}14, ZBJL^{+}10]. 1-42 [DSP11]. 1/PPH [KHW^{+}10]. 10 [SCR12]. 11 [KBW^{+}12]. 125b [GSC11]. 129 [SHC^{+}13]. 13 [LWZ^{+}10, LZW^{+}12, LZW^{+}13, Sho12e]. 133a [DWM^{+}12]. 133a1 [ZWL^{+}14]. 14 [RCC^{+}12]. 19 [BEJ10, EBBJ11]. 1B [KF11, SPF11].

2 [GZR^{+}14a, GZR^{+}14b, HHL^{+}11, HBG^{+}11, KHG^{+}13, LMS^{+}10b, SSK^{+}13, VYC^{+}11, WOG13, XYM^{+}10]. 2* [BAB12]. 2/ [BBJ^{+}10]. 2/p37/p47 [KSH^{+}13]. 206 [CTL^{+}10]. 21 [APV^{+}12]. 214 [AAE^{+}14]. 22 [XTH^{+}11]. 24 [ALV^{+}12]. 25 [WWB^{+}10]. 26S [HPB^{+}12]. 29a/b [LSM^{+}11]. 2A [BKBR11]. 2D [MHAK^{+}12].

3 [ASE10, HHL^{+}11, HC10, Les11-42, Les12z, LWZ^{+}10, PCO^{+}10, SSdA^{+}14, SNR^{+}11, Sho10-65, Sho13e, SRU^{+}12, WHA^{+}13]. 30c [BAB12]. 34 [KKL^{+}11]. 34s [Sho12-33, WSZ^{+}12]. 3A [NSB^{+}11, ZPB^{+}12, GR11]. 3E [TSH^{+}14].

4 [BJE^{+}12, DPW^{+}11, DPW^{+}12, NKH11, SJRV14]. 4.1 [KHW^{+}10]. 4.1G [IHG^{+}12, Sho12-35]. 4/ [BDC^{+}14, DPW^{+}11, DPW^{+}12, TB12]. 40 [WLN^{+}14]. 40S [OCF^{+}10]. 42 [DSP11]. 46 [TKB^{+}14]. 48/p97 [KSH^{+}13].

5 [LSGVM14, LSCF11, PTST12, ZKW^{+}13]. 5-bisphosphate [MRPR12]. 5/Eg5 [WVvG^{+}13]. 52 [TKB^{+}14]. 53BP1 [GGSN^{+}13, HBC^{+}11, MFB12]. 53BP1-containing [HBC^{+}11].

6 [KOO^{+}14, MHK^{+}10, WLN^{+}14]. 6.2 [ZIG^{+}12]. 60S [BHA^{+}12, KLZ^{+}12, Oef10, ONNB^{+}14, SBP^{+}10a].

7 [LYH^{+}13].

8 [BMS^{+}11, WOG13]. 84 [CTM^{+}14a].

9 [ASB^{+}11, Sho11-40, TNH^{+}11]. 9/XTP3 [BGC^{+}10]. 99 [KWTR10].

A-kinase [SDS^{+}12b]. a1 [WWHH10, KYP^{+}14]. A2 [ABP^{+}12, GKWG^{+}11, MPRT11, MTT^{+}14, RBF^{+}12]. A2-dependent [GKWG^{+}11]. AAA [AWB^{+}14, KLZ^{+}12]. ab/gd [VWD^{+}13]. Abcg2 [DZT^{+}11]. ABCs [Sed10g]. Aberrant [RKG^{+}12, IMG^{+}12]. aberrations [EIE^{+}14]. ability [GSB^{+}13]. Ablation [CYN^{+}13, TAC^{+}13, PGB^{+}10]. abnormal [GP12]. abrogates [ZNH^{+}11]. abscission [BB10, BV11, CSM^{+}12, EDF^{+}10, GMW^{+}13, KFET11, MMU10a, OKNP13]. absence [DWL^{+}11, GMW^{+}13, KMC^{+}14, Sho12-64, SP11]. absent [SDS^{+}12a]. Acb1 [DAS^{+}10, MALS10, Sho11p]. accelerates [CVJ^{+}11]. accentuate [Les10u]. acceptance [Sho14-33]. acceptor [CSG14]. access [GGR12, JAM^{+}13, LRA^{+}10, Ros10b]. accessory [BAH^{+}12]. accumulation [BWS^{+}10, RKS^{+}10]. Accurate [ZSD^{+}14, LvBG^{+}10, VTO^{+}13]. accurately [GCR^{+}12]. Acetylated [Sho10d, FWM^{+}10b]. acetylation [DT14, PoLC^{+}13, SLK^{+}13, XG12]. acetylproteome [MMB^{+}11]. achieved [LNS^{+}13]. Acid [Sho10e, CZM^{+}14, KYP^{+}14, MP13, MMdCOM^{+}11, PRFF13, Sho10f, Sho11-48, SJM^{+}13, TID^{+}10, YDB^{+}11]. acid-dependent [MP13, PRFF13]. acidic [MDW^{+}13]. acidification [SHB^{+}10, WH13]. acids [RFRV12a, RFRV12b]. acinar [LDN^{+}13]. acinar-to-ductal [LDN^{+}13]. Acinus [NTSK14, Sho14a]. acquisition [GP12]. across [GHGH11, QWL^{+}11, SLM^{+}11, TESA10]. act [Les10f, Les10g, Les13f, MS14, Sho10o, Sho10-30, Sho13-45, Sho14l, SWS^{+}11]. Actin [CZC^{+}11, CP11, GCH^{+}14, HTS11, HH10, Les13b, Les13c, OVW10, RBH^{+}12, SGLV10, Sho10g, VGL^{+}14, WHH^{+}11, AA13, ALV^{+}12, AFM^{+}13, BLO^{+}12, BKY^{+}10, BFG^{+}13, BCBG10, BVM^{+}11, CWL^{+}11b, DAB^{+}11, DWPC^{+}11, DMH^{+}12, GF11, GJP^{+}13, HM10, HSI^{+}11, HTT13, HHY^{+}12, HC10, IHM13, JLVH12, JKS14, KBC^{+}14, KNOM11, Les10-41, Les13-42, LSM^{+}11, LKG^{+}13, LhYL^{+}13, MAE^{+}10, OBD^{+}10, OD10, PKG10, Sho11-32, Sho12h, SMS^{+}13, SCR12, TB12, TB13, TTB^{+}13, TPM^{+}13, TCN14, WHWS12, WDG^{+}13, YZPF12, YRU^{+}13, YZM^{+}12a, ZJP^{+}12, lDSB^{+}10a, lDSB^{+}10b, Sed14m, FEHF12, KNOM11]. actin-anchored [lDSB^{+}10a, lDSB^{+}10b]. Actin-based [OVW10, AA13, YRU^{+}13]. actin-binding [DAB^{+}11, SMS^{+}13]. Actin-capping [Les13b]. acting [GZZ^{+}14, JK10, VFNR11, ZFP^{+}13]. Actinin [TB12]. actinome [RSL^{+}11]. action [BPL^{+}11, NSD^{+}14, Sed12h, Sho11n, Sho14-64, TLTW10, XRO^{+}11]. actions [EKJH13]. activate [BAY^{+}11, KLF^{+}14, Sho11-59]. activated [GTR^{+}13, VBB^{+}10]. activates [ADS^{+}13, CKO^{+}10, DKA^{+}13, GHK10a, GGSN^{+}13, HIM^{+}10, HSS^{+}13, KLP14a, KLZ^{+}12, KYHG12, Lin10, LCK^{+}13, LXTM12, MFB12, MSS^{+}10, RTC^{+}12, SZE^{+}11, YDB^{+}11]. activating [AMH11, AEC^{+}14, BGB^{+}13, LSCF11, NSB^{+}11, HMiY^{+}10, Les14w, PLR^{+}13]. Activation [GP10, ZSH10, ZLH^{+}14, ADB^{+}14, BLO^{+}12, Boe12, CDK^{+}10, CO13, CDAK10a, CDAK10b, CTW^{+}10, DSK^{+}11, DGF^{+}14, DCO^{+}12, DCO^{+}16, EAK13, GdBP^{+}14, GB10, GJP^{+}13, HZE^{+}13, HAKK11, HVDG13, IWS^{+}11, KSSK12, KCK^{+}14, LAR^{+}10, LCBG^{+}11, LCfC11, LSOT10, LS13b, MdFF^{+}14, MMU10a, MFGB10, MSR10, MHC^{+}12, MBVT^{+}13, NSSF10, NDS^{+}11, dJPAA^{+}11, PRFF13, PAB^{+}11, PRM^{+}14, SDN^{+}14a, SDN^{+}14b, SRKR10, Sho10-41, SRP^{+}13, SHS^{+}13, TGB10, VYM^{+}10, VEDBC13, WWM^{+}12, XHB^{+}10, YHG^{+}14, YHT^{+}10, ZEG11]. activator [DSM^{+}11, HZT^{+}12, MLBY^{+}10, ZBBG10]. activators [Les11d]. active [CLD11, KHG^{+}13, SZ12b]. activities [FRS^{+}13, JPT^{+}11, KLP^{+}14b, MC10, NTSK14]. activity [ALS^{+}13, AAE^{+}14, AFM^{+}13, BWL^{+}11, BNM^{+}14, BRF^{+}10, CDD13, CZD^{+}13, CS13, CSG14, CRP^{+}14, CPX11, CWB^{+}14, DPL^{+}12, DSL13, DGS^{+}10, DWDW12, EZT^{+}12, EMT^{+}14, EAK13, FDB^{+}13, GWP^{+}11, GVP^{+}11, HFS10, HSK^{+}10, HTS^{+}10, HPB^{+}12, HLN^{+}10, HLS^{+}14, HHC^{+}11, JGB^{+}13, JCN^{+}14, JYRL^{+}13, KLF^{+}14, KSLF^{+}11, KFS^{+}14, LAR^{+}12, LHW10, LNJ^{+}13, LSGVM14, LR13, LT11, MTG^{+}11, MMdCOM^{+}11, MTT^{+}14, OYYK14, OZ14, OPM^{+}12, PSR^{+}10, PPG11a, PPG11b, PKS^{+}10, RNS^{+}14, RJM^{+}12, RB11, RC12, RPK^{+}11, RFK^{+}10, SNZVK12, SNZVK13, Sho13i, Sho13-55, Sho13t, TOI^{+}13, VLI^{+}14, WAW^{+}11, WKGB^{+}10, WDG^{+}13, XOV^{+}10]. activity- [CWB^{+}14, KSLF^{+}11]. activity-dependent [CS13, EMT^{+}14, MTT^{+}14]. activity-regulated [FDB^{+}13]. Actomyosin [NWD^{+}11, APC^{+}13, CWL^{+}11a, HHY^{+}12, KCF^{+}14, Les11a, MCS^{+}13, NT11, dMSMZ14]. acts [CAB^{+}10, OZ14, QWL^{+}11]. Acute [dTLLB12]. Acyl [KHFV^{+}13]. Acyl-CoA [KHFV^{+}13]. Adam [Sed14a]. ADAM10 [DCO^{+}12, DCO^{+}16]. ADAM10/Kuzbanian [DCO^{+}12, DCO^{+}16]. ADAM12 [DYI^{+}13]. Adaptation [VCF^{+}13, KAS^{+}12, RGF^{+}10, RPK^{+}11]. adapter [BSR^{+}14, IHG^{+}12, KHB^{+}11b, WWS^{+}12]. adapters [PPV^{+}14]. adaptin [KKS^{+}14]. adaptive [KMSR12]. Adaptor [Bon14, ASE10, BKBS12, DHL^{+}12, KBS^{+}10, MBM^{+}10, MLY^{+}10, SBP^{+}10a]. adaptor-specific [MLY^{+}10]. adaptors [KSH^{+}13, LCBG^{+}11]. adapts [CPT^{+}12, CPT^{+}14]. adds [Sho10-65]. Adducin [CHL^{+}14, Les14a]. Adducin-1 [CHL^{+}14, Les14a]. Adenomatous [OBD^{+}10]. adenovirus [SYV14]. ADF [HKN^{+}14]. ADF/ [HKN^{+}14]. adherens [BG11a, CVR10, Les11-41, OT11, SNZVK12, SNZVK13, SME^{+}13, Sho12h, SRZ^{+}11, TIT11, TB12, TB13, lDSB^{+}10a, lDSB^{+}10b]. adhesion [ALV^{+}12, BKY^{+}10, BVM^{+}11, BC11a, BC11b, BSO^{+}14, CBBH11, CRP^{+}14, CWC^{+}13, DC12, DWPC^{+}11, EBBJ11, FPAM13, HDH^{+}10, HKR^{+}14, HKN^{+}10, KBW^{+}12, KLS^{+}13, LS13a, LZR^{+}11, LMC^{+}12, NDS^{+}11, OBSG12, OLB13, RBH^{+}12, SPC^{+}13, SSdA^{+}14, Sed13q, Sho12-58, SW12, SFL12, TTB^{+}13, TCN14, TDV^{+}14, WKN^{+}13, WHDR^{+}10, YZM^{+}12b, ZBJL^{+}10, ZSK12]. adhesions [ATW^{+}10, BPT^{+}14, BFG^{+}13, DHB^{+}14, LLU^{+}12a, LLU^{+}12b, PSR^{+}10, Sed11c, Sho12s, Sho12-57, Sho14-30, TTB^{+}13, WM12]. adhesive [HTT11b, VMNLB^{+}11]. Adieu [Sed13b]. adipocytes [CWZ^{+}12]. adipogenesis [HHL^{+}11]. adipose [Sho14-37]. adjacent [HWB^{+}13, SZE^{+}11]. adjustment [Sho11w]. ADP [BPDB^{+}11, CPT^{+}12, CPT^{+}14, COB^{+}12, Leu14]. ADP-ribose [Leu14]. ADP/ATP [COB^{+}12]. adrenal [LSE^{+}10]. adrenergic [oHXK^{+}12, MSR10]. adult [CVJ^{+}11, KWK^{+}11, OLT11, PCO^{+}10, SLS^{+}10, ZLW^{+}13, ZLH^{+}14]. advances [MS12]. advice [Sho11-44]. AFD [WOG13]. affects [LWBH12, TAC^{+}13, ZNP^{+}13]. affinity [KSR^{+}13b, PBPW^{+}14a, PBPW^{+}14b, SHC^{+}13, ZSD^{+}14]. Aflame [Sed10i]. after [APC^{+}13, JEF^{+}11, LPG^{+}10, Les10s, Les11e, Les11-46, MSR10, RZS^{+}14, dSMSS13]. against [ACO12, Les12k, PW12]. AGC [HDK^{+}13]. age [APV^{+}12, Sed11a, Sed14x]. age-associated [APV^{+}12]. aggregates [Les11-40, SMMB11]. Aggregation [SJRV14, KUH^{+}14, KMG^{+}11, LNT^{+}10, RKS^{+}10]. aggregation-induced [KUH^{+}14]. aging [DD10b, Les14-33, LR11b, PL11, RvD13, Sed10a, Sed11k]. AGO3 [HLS^{+}14]. agonist [BGC^{+}14]. Agrin [SBS^{+}12]. Ags1 [CSM^{+}12]. ahead [Six12, Sho11v]. aid [Sho12-50]. AIF [QTL^{+}12, QTL^{+}13]. aiming [Hal14]. ain’t [Sho14x]. Aip1 [OD10]. airway [BLC^{+}14]. Ajuba [NDS^{+}11]. AKAP9 [TMS^{+}12]. Akhmanova [Sed11b]. Akt [CZD^{+}13, CLZ^{+}14, PSVRB^{+}11, SOW^{+}11, Sho13a, HAKK11, MHC^{+}12, XHB^{+}10]. Akt/ [XHB^{+}10]. AKT1 [NTSK14]. Alarmin [DKM^{+}13]. ALC1 [PVM^{+}12]. Alejandro [Sed11a]. Alex [Sed14b]. Alexey [Sed12a]. alignment [DLBG11, GKA^{+}12, JKA^{+}10]. ALIX [DCL^{+}12]. ALK1 [ZKW^{+}13]. ALK1-Smad1 [ZKW^{+}13]. ALK1-Smad1/5 [ZKW^{+}13]. All-purpose [EEP13]. allosteric [KPE^{+}14]. allows [GB10, Les11v]. alone [Sho13o]. along [LSS^{+}12]. Alpha [Sed13a]. ALPS [PMB^{+}11]. ALT [CSH^{+}12]. altered [BBK^{+}13]. altering [KYHG12]. alternative [BLC^{+}12, IIWS14, KNsMK13, LT11, MVC^{+}11, OBC14, RCBY^{+}12, SBR^{+}11, SNT^{+}12, Sho10n, VBB^{+}10]. alters [DMK^{+}12, HIM^{+}10]. Alvarado [Sed11a]. Always [Hal14]. Alzheimer [DSP11, WTH^{+}11]. AMAP1 [ONH^{+}12]. AMBRA1 [DCN^{+}10]. Ambros [Sed13x]. amendment [Sho11s]. Amiloride [KWH^{+}10a, KWH^{+}10b]. amino [CZM^{+}14, MP13, PRFF13, RFRV12a, RFRV12b, YDB^{+}11]. amoeboid [BMÁG^{+}14, Sho14-65]. Amon [Sho11b]. AMP [CTW^{+}10]. amphipathic [SDS^{+}12b]. AMPK [JCL^{+}11, LSS^{+}12, MSK^{+}13b, Sho11a, YMT^{+}13]. AMPK- [JCL^{+}11]. AMPK-dependent [LSS^{+}12]. amplification [BKBR11, HLS^{+}14, KRS11, LMS^{+}10b]. amplifies [Sho11a]. amplify [BPMK^{+}14]. Amy [Sed14c]. amyloid [DSP11, RKK^{+}14, Sho11-27]. amyotrophy [BBK^{+}13]. Ana2 [SDB^{+}10]. analyses [BCB^{+}14b]. Analysis [LhYL^{+}13, DV10, FBZM^{+}10, GSU^{+}12, KWL^{+}12, RPO^{+}14, SSZ^{+}14, SLC^{+}13, TMG^{+}10, TBV^{+}14, WZHV11, ZKR^{+}11]. anaphase [BCB14a, GSP^{+}14, HSN^{+}11, JDS^{+}10, LWK^{+}13, LS13b, NEMH^{+}10, RDPG14, RHK11, Sho12-51, UG10, UTK^{+}13, WBMCSS13, WDB10, WHL^{+}12]. anaphase-promoting [NEMH^{+}10, WDB10]. anchor [Sho12-29, Sho14-39]. Anchorage [CLL^{+}10, GLG12]. anchored [ARF10, BBW^{+}14, YHF13, lDSB^{+}10a, lDSB^{+}10b, BBD^{+}11, FWJ^{+}11]. anchoring [NSS13, She14, SDS^{+}12b, TSL12, TMS^{+}12, ZGEM12]. anchors [FTJG13, HKN^{+}10, RBH^{+}12]. Anderson [BG11b]. Andy [Sed10a]. anemia [GHK10a]. aneuploid [HRK13, TC10]. Aneuploidy [RvD13, RJvD11]. Angelika [Sho11b]. angiogenesis [ZKW^{+}13]. Angiopoietin [NGL^{+}12]. Angiopoietin-1 [NGL^{+}12]. angle [Les14g]. Anillin [AGL^{+}14, EKJH13, Sho14-54]. animals [WPSA13]. Ankyrin [BLI^{+}10, HAB14, LBD^{+}14, Sho14b, Sed11u]. Ankyrin-G [HAB14]. Ankyrin2 [BRP14]. ankyrinB [RGL^{+}13]. ankyrinB-dependent [RGL^{+}13]. Anna [Sed11b]. Annexin [Les11b, GKWG^{+}11]. annexins [LMS^{+}13]. Answering [Sed11l]. answers [Sed14i]. antagonism [DWM^{+}12, GHC^{+}14, KST^{+}11, QJO10]. antagonistic [CSTBM^{+}10, HS10a]. antagonists [SRS10]. antagonize [MPD^{+}12]. antagonizes [CSP^{+}10, LT11]. anterior [BEJ10, GHGH11]. antiapoptotic [FBR^{+}10]. antigen [RGB^{+}13]. Antisense [RHKB12]. Antisense-based [RHKB12]. AP [ASE10, BBJ^{+}10, BSP11, KF11, SPF11]. AP-1 [BBJ^{+}10]. AP-1-mediated [BSP11]. AP-1B [KF11]. AP-1B-dependent [SPF11]. AP-3 [ASE10]. AP2 [KPJ^{+}13]. APC [ADB^{+}14, COW13, CRJB^{+}11, EMO12, HZT^{+}12, IP12, JK10, Les10a, LS13b, LHD^{+}14, NMB^{+}14, PM13, SBEM13, Sed10g, Sho12c, YZL^{+}13, vZOtR^{+}10, BDR^{+}12]. APC/C [ADB^{+}14, COW13, CRJB^{+}11, EMO12, HZT^{+}12, IP12, LS13b, LHD^{+}14, NMB^{+}14, PM13, Sho12c, vZOtR^{+}10]. apical [APC^{+}13, GSU^{+}12, HKI^{+}13, KCF^{+}14, VOSB12, ZME^{+}14]. apically [Sed13l]. aPKC [GHC^{+}14, HKI^{+}13, IMP^{+}12, NSZ^{+}13]. Apobec2 [ERS10]. Apobec2-related [ERS10]. apoptosis [CTW^{+}10, HLN^{+}11, HDH^{+}10, HVW^{+}10, HRWW^{+}13, JRC^{+}13a, JRC^{+}13b, LSOT10, MBO^{+}14, NSSF10, QTL^{+}12, QTL^{+}13, Sho12c, YSN^{+}11, dSMSS13]. Apoptotic [CTY^{+}12, Sho12d, ER10, FA12, KDIE11, LZY^{+}12, Sho11-55]. APP [mFH13]. apparatus [CDH^{+}14, DS10, OYYK14, Sed12a]. appendages [TYN^{+}13]. approach [CTM^{+}14b, Sed11i, Sho10-56]. approaches [YST^{+}11]. aquaporin [SJRV14]. aquaporin-4 [SJRV14]. arborization [CO13]. archipelago [Sho10-45]. Architects [CF13]. architecture [CTM^{+}14a, FS14, HH14b, JJH^{+}10, KBS^{+}10, LCLW11, Leu14, MHS10]. Arf [HDH^{+}10, Sho10h]. Arf1 [CM12c, Sho11c]. Arf6 [MRPR12, SPF11]. arginine [CC12]. Arginylation [ZSK12]. Arginylation-dependent [ZSK12]. ARH [KF11, SBTF13]. ARHGEF7 [BPB^{+}12]. arise [ABP^{+}14, EIE^{+}14]. Arkadia [PHW^{+}13]. ARL [LWZ^{+}10, LZW^{+}12, LZW^{+}13, Sho12e]. ARL-13 [LWZ^{+}10, LZW^{+}12, LZW^{+}13, Sho12e]. ARL-3 [LWZ^{+}10]. Arl1 [CM12c]. Arl13b [CHK^{+}10a, CHK^{+}10b]. Armitage [HLS^{+}14]. arms [KMSR12, Sho10-47]. Arnaud [Sed13b]. Arp2 [IHM13, Les12z, Sho13b, Sho13e, SRU^{+}12, TB12, WHA^{+}13, YZPF12]. Arp2/ [IHM13, Sho13b, TB12, YZPF12]. Arp2/3 [Les12z, Sho13e, SRU^{+}12, WHA^{+}13]. Arp5 [MH14]. Arpc1b [MLBY^{+}10]. Arpc3 [LSM^{+}11]. arrangement [BYY^{+}12, Sho13-34]. array [LCS^{+}13]. Arrayed [VTO^{+}13]. arrays [GBK^{+}14]. Arrest [OBC14, GSJS10, Sho12-41, WAG^{+}10]. arrested [PTBT10]. Arrestin [MSR10, BVL^{+}12]. arrestin-like [BVL^{+}12]. arrests [LZLG13, TL12]. arsenic [dTLLB12]. art [Sed14d]. artemin [BST^{+}11]. arterial [YMU^{+}10, YMU^{+}13]. articular [NSB^{+}11]. ASF [VBB^{+}10]. ASH1 [CT10, HFB^{+}10]. aspect [VOSB12]. aspects [RDB^{+}12, WHH^{+}11]. assay [BKS^{+}13, JDB^{+}12]. Assemblages [TW14]. assemble [KHS^{+}11]. assembled [HHY^{+}12]. Assembly [HHS^{+}14, LCLW11, AMGC14, BKS^{+}11, BPH^{+}14, BSR^{+}11b, BSR^{+}11c, BCB^{+}14b, BM11, CMS10, CHL^{+}14, CP11, CDB^{+}14, CSKW13, CSEH12, DP10, DE10, EIW^{+}12, EUB^{+}14, FSA^{+}10b, FRS^{+}13, GLB10, GBL^{+}11, GC13, GKR^{+}11a, GKR^{+}11b, GRH^{+}12, GDS^{+}12, HYS11, HAB14, HZE^{+}13, HTT13, HWB^{+}13, HSTF13, HTM^{+}14, IHM13, IMG^{+}12, KKUG11, Kar10, KOO^{+}14, KNOM11, KTN^{+}12, KKL^{+}14, KBS^{+}10, LCD^{+}11, LYH^{+}13, LYB^{+}10, LLK11, LvBG^{+}10, LZLG13, MAS11, MGT^{+}10, MMU10a, MXS10, MHS10, MMFS11, MSS^{+}12, MAD10, NAS^{+}11, NAS^{+}12, NAS^{+}13, NSD^{+}14, OZ14, OBD^{+}10, OL12, OD10, OGD^{+}12, PCC11, PoLC^{+}13, PPD^{+}10, RTC^{+}13a, RTC^{+}13b, RMG^{+}12, RBB^{+}14, RGF^{+}10, SFJ^{+}14, SFK^{+}13, SLM^{+}13, SSZ^{+}14, Sho10g, Sho10-51, Sho14-45, Sho14-52, SBP^{+}10b, SIO10, TH11, TB12, TSL12, TNH^{+}11, TKB^{+}14, UG10, UTK^{+}13, VWC^{+}13, WEK^{+}14, WS10, YOA^{+}11, YCP10, ZNA^{+}14]. associate [HYS11]. associated [APV^{+}12, DHL^{+}12, HBM^{+}11, KTN^{+}12, KPI^{+}10, LP13, MVR^{+}10, MPD^{+}12, RKK^{+}14, RFC14, SSL^{+}14, SHC^{+}10, ZSH10, ZYH^{+}11, ZJP^{+}10]. associates [HOS^{+}12, MSZ^{+}12, WWS^{+}12]. association [FCA10, HM10, HPB10, KK13b, LNJ^{+}13, MJEM10, Oef10, PMF12, RKW^{+}13, SKN^{+}13, SBP^{+}10a, YMT^{+}13]. associations [CAB^{+}13, WLK^{+}11]. assumes [VKMI12]. Aster [TGES12, JG10]. Asters [Les12b, Sho10-34]. asthma [ES14]. astral [RW10]. Astrin [SKH^{+}10, DLBG11]. Astrocyte [KWK^{+}11]. astrocytes [PMP^{+}11a, PMP^{+}11b, SJRV14]. astrocytic [CYN^{+}13, KST^{+}10]. Asymmetric [LN11, MBLD11, BMLB^{+}12a, BMLB^{+}12b, BDvdK13, BGY^{+}13, GYZ^{+}12, JG10, KWTR10]. asymmetrically [MXS10, OB12]. asymmetry [BYY^{+}12, DHL^{+}12]. Asynchronous [DWJ^{+}14, BDC^{+}14]. ATAD5 [yLFAM13, Les13d]. ataxia [LgLM^{+}10, Orr12, Les10b]. Atg14L [MMS^{+}10]. Atg18 [Sho13-41, TOI^{+}13]. Atg30 [NOT^{+}14]. Atg37 [NOT^{+}14]. Atg38 [AKA^{+}13]. Atg8 [KBW^{+}10]. Atg9 [MGR^{+}10, RT12, YKW^{+}12]. Atg9-containing [MGR^{+}10]. atlastin [MLSM^{+}11, Sho11g]. ATM [DPL^{+}12, JPT^{+}11, SZ12b, SDC10, YWJ^{+}12]. ATOH1 [GvEM^{+}11]. Atomic [NRM^{+}12]. ATP [AvCG^{+}11, HDK^{+}13, IHM13, RNS^{+}14, COB^{+}12]. ATP-dependent [AvCG^{+}11]. ATPase [CZC^{+}11, CMW11, CCJ^{+}12, KLZ^{+}12, KPH^{+}12, PGAE^{+}13, Sho10f, VM14, WH13, WEK^{+}14, WWHH10, XSJ^{+}10]. ATR [GSGL11, JPT^{+}11, MFB12, MTM^{+}10, VEDBC13]. ATR-dependent [MFB12]. ATR-mediated [GSGL11]. Atrophin [ZFP^{+}13]. atrophy [CZGG12, CLZ^{+}14, MBK^{+}10, RHKB12, SYK^{+}11]. attached [Sho14-46, Sed13q]. attachment [BKG10, BSR^{+}14, CWPW11, DWDW12, GCR^{+}12, MS14, TIM14, VWC^{+}13]. attachments [CYLMM13, CJNS12, DSL13, MGK^{+}12]. attenuates [DPZ^{+}14, KKY^{+}14]. Attenuation [CCGN11, HSK^{+}10, WPM14]. attraction [TSB^{+}14]. augments [PHD^{+}10]. Augmin [KOK^{+}13, Sho10i, Les13f]. Augmin-dependent [KOK^{+}13]. Augmin-ting [Sho10i]. aurora [BWS^{+}10, PKS^{+}10, TUG^{+}10, XOV^{+}10, BKS14, BGB^{+}13, BCB14a, BDB^{+}14, CDD13, CJNS12, DMK^{+}12, FPAM13, HSN^{+}11, IMG^{+}12, KPE^{+}14, KSH^{+}13, Les11c, Les11-37, Les12c, LVB^{+}10, MMU10a, MS14, MLBY^{+}10, NCT^{+}11, NvCL^{+}13, PAB^{+}11, PPG11a, PPG11b, RTC^{+}12, RTC^{+}13a, RTC^{+}13b, RJvD11, RMF^{+}10, SKH^{+}10, Sho11d, Sho13c, UTK^{+}13, WBL11, WUD^{+}12, WDB10, WHL^{+}12, ZBBG10]. Aurora-dependent [NCT^{+}11]. authors [dWMR10]. autism [PYT^{+}13]. autism-linked [PYT^{+}13]. autocatalytic [SRP^{+}13]. autoimmunity [BHMBS^{+}11, BMS^{+}11]. Autoinhibition [HBS^{+}10]. autolysosomal [SWS^{+}13]. Automated [CG10b, DV10]. Autonomous [TKMK10, KHB^{+}11a, KPC^{+}11]. Autophagic [NSS^{+}10, DGH^{+}14, FMI^{+}13, LYH^{+}13, RKS^{+}10]. Autophagosomal [TNV^{+}13, IKU^{+}11]. autophagosome [BMC^{+}11, IM11, IKU^{+}11, KSR^{+}13a, KBW^{+}10, LLR^{+}12, MALS10, MGR^{+}10, MRPR12, Pfe10, YKW^{+}12]. autophagosome-mediated [BMC^{+}11]. autophagosome-resident [IKU^{+}11]. Autophagosomes [MWH12, DAS^{+}10, Krä13]. Autophagy [LZY^{+}12, MMS^{+}10, MFF^{+}13, RT12, RGB^{+}13, Sho10j, AKA^{+}13, CZM^{+}14, CKU^{+}10, CZ10, CMD^{+}13, DCN^{+}10, DCP^{+}10, Les10n, Les12-27, MJFS10, MMB^{+}11, NTSK14, Oka14, ONNB^{+}14, PRM^{+}14, RKS^{+}10, SA10a, Sed10m, Sho10l, Sho10r, Sho10a, Sho14a, Sho14u, Sho14-35, TTM^{+}14, TKL^{+}10, YSN^{+}10, Sed12z]. autophagy-deficient [RKS^{+}10]. autophagy-related [SA10a]. autophagy-specific [AKA^{+}13]. Autophosphorylation [SDC10, GKA^{+}12]. autoregulating [HLN^{+}10]. auxiliary [VYC^{+}11]. avoid [Car12]. away [Les11-34, Les14-36]. aweigh [Sho14b]. axial [BMÁG^{+}14, LHGT^{+}12]. axially [SSV^{+}12]. axis [EBB13, KKL^{+}11, SSK^{+}14, WCQ^{+}13]. Axon [JKS14, WMB12, Les14-35, LCP13, Sed11p, SFB^{+}12, SLH13, SMK14, WMV^{+}14]. Axonal [EEP13, BLI^{+}10, CO13, CS13, FHD^{+}12, mFH13, LgLM^{+}10, LBD^{+}14, SFB^{+}12, Sho13-54, Sho14b, VYC^{+}11, XWE^{+}10, YOMM^{+}11, ZPB^{+}12]. axonemal [DSB^{+}14, OYYK14, YHK10]. axoneme [BYY^{+}12, KSR^{+}13b]. axonogenesis [DHVK10a, DHVK10b]. axons [ASLS14, Les11v, Les12-31, Sho12-35, Sho14-48, ZSZ^{+}13].

B [LBD^{+}14, GKR^{+}11a, GKR^{+}11b, GRH^{+}12, HHS^{+}14, BLC^{+}12, BWL^{+}13, BKS14, BWS^{+}10, BGB^{+}13, BCB14a, BHMBS^{+}11, BGC^{+}10, CDD13, CJNS12, CSTBM^{+}10, DS12, DMK^{+}12, FPAM13, HSN^{+}11, JGA^{+}11, KPE^{+}14, Les11c, Les12c, Les12-29, Les14w, Lin10, LDN^{+}13, LVB^{+}10, MMU10a, MLH12, MS14, NvCL^{+}13, OMW^{+}14, PAB^{+}11, PKS^{+}10, RJvD11, SKH^{+}10, Sho13c, TKB^{+}14, TUG^{+}10, UTK^{+}13, WBL11, WUD^{+}12, WBMCSS13, WDB10, WHA^{+}13, WHL^{+}12, XOV^{+}10, LSM^{+}11]. B-activating [Les14w]. B-dependent [WHA^{+}13]. B-mediated [MMU10a]. B1 [DCO^{+}13, GP10, KYOY13, MJJ^{+}10, vZOtR^{+}10]. B23 [RTC^{+}12, LKLA12, Les12d]. B23/nucleophosmin [LKLA12]. B55 [OMW^{+}14]. B56 [EUB^{+}14, BCB14a]. BAC [HBC^{+}10]. back [HHS13a, HHS13b, Les14l, Les14-35, Sho11-36, Sho13-37, VMNLB^{+}11]. backstage [GGR12]. bacteria [ATKK11, PBG^{+}13, TLSA14]. Bacterial [PR12, BVM^{+}11, KWO11, Sho10p, Sho13-57]. baculovirus [OVW10]. bad [Les11z]. BAG [MHK^{+}10]. BAG-6 [MHK^{+}10]. BAG1 [DGS^{+}10]. Bak [DSM^{+}11, GWR^{+}10, Les11d, SDN^{+}14a, SDN^{+}14b]. Bakal [Sed13f]. balance [AWB^{+}14, CSTBM^{+}10, RLS^{+}14, Sed11o, SHV^{+}13]. Balanced [KSSK12]. Balancing [CM12b]. band [DGF^{+}14, FS14]. bands [CZGG12]. BAR [CLW^{+}14, KSR^{+}13a, QJO10, ZBJL^{+}10]. barbed [HM10, TB13]. barbed-end [TB13]. BARP [BNM^{+}14]. Barr [PASG^{+}12]. Barral [Sed14x]. barrel [KIL^{+}12]. barrier [BLI^{+}10, BKS^{+}13, HSI^{+}14, HSI^{+}11, JCN^{+}14, MKS^{+}13, OGD^{+}12, RBY^{+}11, Sho10-35, SMT^{+}10, SFL12, YMM^{+}10, ZPS^{+}10]. barriers [BC11a, BC11b, SNSyN13]. Barry [Sed12b]. BARs [FSR11]. Barth [CWS^{+}11]. basal [BLC^{+}14, ETC^{+}12, HKN^{+}10, KD11, NSS13, NTSK14, RH10, TYN^{+}13, WLK^{+}11]. basally [Sed13l]. base [EIE^{+}14]. based [AA13, FBR^{+}10, HKN^{+}10, JTN^{+}13, Kin13, OVW10, RHKB12, SLH13, TSH^{+}14, WM10, WBL11, YRU^{+}13]. Basement [KTN^{+}12, HZM^{+}13, KLHS14, Les12-33, Sho13p]. basic [LCfC11]. basis [CBB12, GLG12, PMK^{+}13]. Batten [KKUG11]. Baum [Sed14d]. Bax [GWR^{+}10, HZT^{+}12]. Bax-deficient [GWR^{+}10]. Bazooka [BGY^{+}13, KBKW10]. BBSome [LBS^{+}13, Sho13d]. bc [GRH^{+}12, EMT^{+}14]. Bcl [bCAH^{+}11, GWR^{+}10]. Bcl-x [bCAH^{+}11, GWR^{+}10]. Bcr [DKA^{+}13]. BDNF [DMD^{+}12]. be [Les10o, Sho12b, TGB10]. bearing [CTM^{+}14a]. beat [Sho11j]. beautiful [Sed13v]. beauty [Sed14s]. before [BLI^{+}10, CZC^{+}11, vZOtR^{+}10]. beginning [Sed12z, Sed13p]. beginnings [GK13]. behavior [CG12b, CRL^{+}14, SZ12a, SW10a, SEV^{+}14]. behaviors [PCCR11]. behind [Les13t]. being [SZ12b]. Bellaïche [Sed13y]. Belmonte [Sed14g]. below [Sho14y]. belt [PJS^{+}11, Sho10-65]. Bem1 [SRP^{+}13]. Bem1-mediated [SRP^{+}13]. Bem1p [LN14, Sho14c]. bend [Sho10-61]. bender [FSR11]. bends [Les11-31, Les14k, Sho10-55]. Bennett [Sed11u]. Benny [Sed11c]. Bergmann [Pow14b]. Bertuzzi [Sed12u]. best [Hal14, Les10-36]. beta [Les10a]. Bettencourt [Sho10-39]. better [Les11-44, Sho11e, Sho11t, Sho13-58]. Bettina [Sed10b]. between [ANT^{+}12, Bez12, DWM^{+}12, ETYS^{+}12, GHC^{+}14, HBI^{+}10, HSKAT11, KdKDP12, dSLPRG11, MDW^{+}13, OYYK14, PYT^{+}13, RKG^{+}10, RSS^{+}13, Sho10-27, Sho11g, SHBC12, SLS^{+}10, TYN^{+}13, VCF^{+}13, VG13, WH13, WJW^{+}11, dSJDD^{+}11]. beyond [BM10, FSK^{+}10, HMB14, LeB10, RFC14, Sho10-35]. BH3 [CTW^{+}10, DSM^{+}11]. BH3-binding [DSM^{+}11]. BH3-only [CTW^{+}10]. Bidirectional [MDW^{+}13, FS10, GHC^{+}14, KUN^{+}13]. bifurcation [AGM^{+}10]. Big [Kik13, Sho12f, Les12t, Sed12m, Sed13e, Sed13d, Sed11s]. BIG1 [CM12c]. BIG2 [CM12c]. Bigger [Sho11e, Les13g]. Biggins [Sed12v]. bilayers [dSJDD^{+}11]. Bilder [Sed11e]. Bim [BMS^{+}11, CTW^{+}10]. bind [BBK^{+}13, BBY^{+}12, HYS11, MS12, MHV12]. Binding [HTT13, AVP^{+}14, BKBS12, CC12, CRP^{+}14, DSM^{+}11, DPV^{+}12, DAB^{+}11, DK10a, DKF^{+}11, ECK^{+}12, EJBW12, GSM^{+}14, HDK^{+}13, HTT^{+}11a, HTS11, HKI^{+}13, HAB14, HWB^{+}13, IM11, LNJ^{+}13, LSW^{+}14, MBR^{+}11, MVP^{+}11, NNO^{+}11, NCML^{+}12, PMB^{+}11, RSM^{+}13, RCC^{+}12, SYH^{+}13, Sho12-39, SMS^{+}13, SMT^{+}10, TOI^{+}13, UKZ^{+}13, WAW^{+}11, YYA^{+}11, ZNH^{+}11, vGLWB12, DSM^{+}11, DPV^{+}12, KKK^{+}11, RCBY^{+}12, YZL^{+}13]. binds [DS10, DCL^{+}12, GSP^{+}14, GRHA^{+}12, LCS^{+}10, MJEM10, NOT^{+}14, PAB^{+}10, SBEM13, SMS^{+}13, WJPD11]. bioavailability [NLAS^{+}10]. biochemical [LZW^{+}10]. biochemistry [Sab10]. Biocompatible [NRK^{+}13]. biofilm [Sho11h, VŠH^{+}11]. Biogenesis [BNH12, BMC^{+}11, PGP14, AHL^{+}11, ARF10, BWL^{+}13, BWK^{+}11, BKT13, Dun11, GSM^{+}14, JBS^{+}12, JBS^{+}13, KHFV^{+}13, KA12, KBW^{+}10, MAS11, MGR^{+}10, Sho14-40, ZCB^{+}10a, ZCB^{+}10b]. biological [SMK14]. biology [BM10, BWM12, CC10a, CR10a, CR10b, DN10, Dev14, DK10b, ES14, Gol12b, GCC12, GM11, HSF12, KN12, LRB13, LMN10, MS12, MRA14, Orr12, Sab10, SA10b, SKM10b, Sed10a, Sed10e, Sed10k, Sed11d, Sed11i, Sed11o, Sed12s, Sed13c, Sed13o, Sed14b, Sed14j, SC10a, WM10, WPSA13]. biomedicine [BWM12]. biorientation [STD^{+}10]. bioriented [SKH^{+}10]. biosensor [EAK13]. biotin [RKRB12]. Biphasic [FLN^{+}10, YWC^{+}13, FLN^{+}16]. bipolar [SW10a]. bipolarity [BKK^{+}10, JVS^{+}14]. birth [BG11b]. bis [VSG^{+}12]. bis-phosphate [VSG^{+}12]. bisphosphate [MRPR12]. bite [Sho11-35]. bites [Les13g]. bladder [WHF^{+}11]. Blame [Les13d, Les11z]. Blanpain [Sed12d]. Blasco [Sed11k]. blastocyst [MC10]. Bleb [ZTBK14]. Bleb-driven [ZTBK14]. blebbing [GPCK12]. BLM [HMBC10]. BLOC [PLR^{+}13]. BLOC-1 [PLR^{+}13]. block [IMG^{+}12, ZDS^{+}12]. blocker [Les10a]. blocks [AMH11, GKWG^{+}11, KKS^{+}14, Oka14, SOW^{+}11]. blood [CBB12, HSI^{+}11, Sed10e]. Bmf [CMD^{+}13]. BMI1 [IAMH10]. BMI1-mediated [IAMH10]. BMP [CRL^{+}14, GDO13, KST^{+}10, NLP^{+}10, NLAS^{+}10, dJPAA^{+}11, ZLH^{+}14]. BMP-induced [KST^{+}10]. BMP-regulated [CRL^{+}14]. Bnip3 [QTL^{+}12, QTL^{+}13]. board [Mis10]. Bob [Sed13c]. bodies [CLSO^{+}12, CAB^{+}13, DGS^{+}11, wFLW^{+}13, NNSH11, NWD^{+}11, RBM^{+}11, SFJ^{+}14, SRBL13, Sho13-49, TYN^{+}13, VSMC11, WBcY^{+}11, dTLLB12, COG11]. body [BLC^{+}14, CMW11, Dun11, ETC^{+}12, KD11, Sed12e, SIO10, TGG^{+}11, TL12, WLK^{+}11, YOA^{+}11, RMS^{+}14]. body/transition [WLK^{+}11]. Boi [Sho10k, HZS^{+}10]. bond [HTT11b, KLvdB^{+}13]. bonds [CSHS^{+}13a, CSHS^{+}13b, Sho11i]. Bone [Les13e, RT10, ASB^{+}11, BBJ^{+}10, DK10b, HSK^{+}10, ISZ^{+}11, KPC^{+}11, KPH^{+}12, Les10-29, NLAS^{+}10, SNR^{+}11, Sho10-71, Sho11-40, Sho12-33, SJZ^{+}10, WCQ^{+}13, ZYH^{+}11]. bones [Sho11-56]. Bootstrapping [Sed11a]. border [CMT14, LKG^{+}13, PCCR11, Sed11f]. borders [Les13n]. Both [BMÁG^{+}14, SML^{+}13, CWPW11, FBR^{+}10, MLBY^{+}10, NSB^{+}11, Sho12p, Sho12-27]. Botulinum [SMT^{+}10]. bound [DDH^{+}12, GL10, oHXK^{+}12, KHB^{+}11b, PPV^{+}14, SYS13, WWB^{+}10, ZBBG10, MLSM^{+}11]. boundaries [BGY^{+}13, Dan14]. boundary [EMO12, GHGH11, Les11-29]. box [BCB^{+}14b, CMW11, KLC^{+}10, LC10, vZOtR^{+}10, LNL11]. Boyer [Sed12o]. brain [DMD^{+}12, HSI^{+}11, MVC^{+}11, Sho13-33, TAC^{+}13]. brake [Les11-35, RHK11]. brakes [Les11-39, Sho10-51, Sho11-42, Sho13-54, SS11]. branch [Sho10-56]. branches [Sho11-32, Sho13-28]. branching [Les14-35, LCP13, MAD^{+}11, WMV^{+}14]. BRCA1 [BAS^{+}14, CGW^{+}11, GGSN^{+}13, KA12, Les12e, Les13u, RSS^{+}13, Sho11f].

BRCA2 [RZS^{+}14]. Breaching [HH14a]. break [BNDB^{+}14, BCJ13, IAMH10, KK13b, PLC^{+}11, Yam13, YTM^{+}11]. breakage [JRC^{+}13a, JRC^{+}13b]. breakdown [Les12r, SWS^{+}13, TCB^{+}14]. Breaking [MS12, Sho13e, Les10q, SRP^{+}13, YRU^{+}13]. breaks [CCJ^{+}12, CWG^{+}11, EIE^{+}14, GBJ10, PLL^{+}12, Sho11-41]. breast [ATU^{+}12, CG12b, CGW^{+}11, GGSN^{+}13]. bridge [AGL^{+}14, MLSM^{+}11]. bridges [GSP^{+}14, ZGEM12]. Bridging [Pri14, Sho11g, Sab10]. Briggs [Sed13m]. bring [Les11-33]. Bringing [TALR11, Sed12q]. brings [Les14y, Sho14-29]. brink [Les12b]. Bro1 [WAW^{+}11]. broad [Sed13x]. Brodsky [Sed13i]. broke [Sho14x]. brown [OSD^{+}14, Sho14-37]. Brr6 [TGG^{+}11]. Bruce [Sed12c]. Bruchpilot [Sho13f, MSK^{+}13a]. Brugge [Sed10j]. brush [CMT14]. Btn1 [KKUG11]. BUB [Sho14d, VTO^{+}13]. BUB-1 [Sho14d]. Bub1 [KWL^{+}12, MKH^{+}14, RJvD11, RJM^{+}12, Sho12-54]. Bub3 [Sho13g, dSMSS13]. BubR1 [GKA^{+}12, PoLC^{+}13, Sho12g]. bud [FHKW11]. bud-directed [FHKW11]. Bud3 [KLP14a]. Budding [Sho11h, Sho14e, AMS^{+}13, CWPW11, HTM^{+}14, KLP14a, KST^{+}11, LCS^{+}13, RY11, SKN^{+}13, WMC14, WVT^{+}13, GBL^{+}11]. Building [Sho12h, Sho14f, Oka14, Sho14k, SC10b, Sed13t]. builds [Sho14g]. bulk [KKS^{+}14]. bulks [Les11a, Sho14o]. bundle [BBK^{+}13, SRKR10]. bundles [CWL^{+}11b]. bundling [CWL^{+}11b, SMM^{+}10, WDG^{+}13, ZJP^{+}12]. Burridge [Sed12n]. buttoned [Sho14-48]. Buzz [Sed14d]. bypass [MHKM11].

C [HTS^{+}10, MMFS11, XYM^{+}10, FHD^{+}12, MLH12, RBM^{+}11, ADB^{+}14, BCB^{+}14b, CNP^{+}12, COW13, CRJB^{+}11, EMO12, HMiY^{+}10, HZT^{+}12, IP12, JLVH12, KWDD10, KBW^{+}12, LS13b, LHD^{+}14, NMB^{+}14, PM13, Sho12c, STG13, WBS11, vZOtR^{+}10, FBR^{+}10]. C-cadherin [KBW^{+}12]. c-Jun [FHD^{+}12, RBM^{+}11]. c-myc [MLH12]. C-terminal [CNP^{+}12, JLVH12, STG13]. C. [AGL^{+}14, BDC^{+}14, CTY^{+}12, EM11, GZR^{+}14a, GZR^{+}14b, GHGH11, HYS11, KHW^{+}10, KHG^{+}13, KWTR10, MMVK^{+}12, OGD^{+}12, UHKS11, WOG13, WRCD12, ZBJL^{+}10, ZC11]. C/D [BCB^{+}14b]. C8ORF38 [ZLJ^{+}13]. Ca [ADF^{+}12, BNM^{+}14, CZD^{+}13, DAB^{+}11, DSP11, DS10, FCA10, IIN^{+}11, KBC^{+}14, KSLF^{+}11, MSR10, MDW^{+}13, SPC^{+}13, SLH^{+}14, TPZ^{+}14, WEK^{+}14, YSM10, vBAK^{+}12]. cab [Sho12-53]. Cab45 [vBAK^{+}12]. cables [HHY^{+}12]. Cadherin [HTT11b, Ari10, BKY^{+}10, BGY^{+}13, HTT13, Les13c, Les13n, MDP^{+}10, NBDB12, OT11, RSB13, SRZ^{+}11, TCN14, TDV^{+}14, BEJ10, EBBJ11, HOS^{+}12, KBW^{+}12, NDS^{+}11, RSB13, SPJ^{+}14, SZE^{+}11, SMT^{+}10, YYA^{+}11, lDSB^{+}10a, lDSB^{+}10b]. cadherin-dependent [NBDB12]. cadherin-mediated [BKY^{+}10, TCN14, TDV^{+}14]. Cadherins [Sho11i, NCML^{+}12, NAS^{+}11, NAS^{+}12, NAS^{+}13, SY10, Sed13a]. Caenorhabditis [CHK^{+}10a, CHK^{+}10b, FTJG13, RZA^{+}13, RKW^{+}13, SQC^{+}12]. Cajal [SIO10]. CAL1 [CDB^{+}14, Sho14g]. calcium [BNM^{+}14, Bez12, BJE^{+}12, CFB^{+}12, DWM^{+}12, JC10, KPH^{+}12, Les14e, Les14r, PPG11a, PPG11b, SB14, Sho11d, VBG^{+}13, YFLH12]. calcium-dependent [BJE^{+}12, SB14]. call [Sho13h]. Calmodulin [VPC^{+}14, WEK^{+}14, DS10, MSR10]. calmodulin-dependent [KSLF^{+}11]. calpain [HVOF^{+}14, SPC^{+}13]. calpain-like [HVOF^{+}14]. calyceal [SDS^{+}12a]. Camargo [Sed12g]. CaMKII [LLT^{+}12, Les12f]. cAMP [GVP^{+}11, LLH13, TMS^{+}12]. cAMP/PKA [LLH13]. cAMPs [Sho11r]. can [HSI^{+}14, KMC^{+}14, Les10o, Les11-38, MHKM11, Sho11-35, TGB10, WAG^{+}10, Gol12b]. canal [HC10, MHCvSW11]. Cancer [Les11e, ATU^{+}12, BT12, BKK^{+}10, CG12b, CGW^{+}11, DSMB13, Gol12a, GGSN^{+}13, HJ14, KKL^{+}11, Les12u, Les14d, Les14f, LWW12, MHAK^{+}12, ONH^{+}12, RIG^{+}12, Sed10c, Sed10d, Sed10j, Sed11k, Sed11t, Sed12w, Sho12i, SGT^{+}13, XTH^{+}11]. Canoe [WJPD11]. Canonical [AKB^{+}13, BPMK^{+}14, NSB^{+}11]. cap [SOW^{+}11, Sed11k, Sho12-44, Sho13i, BDN^{+}13]. cap-dependent [SOW^{+}11]. capacity [FHA10, HKR^{+}10]. capping [GLM^{+}10, Les13b, TB13]. capsid [RMG^{+}12]. capture [FHY^{+}10, SBS^{+}12, YWC^{+}13, YZM^{+}12a, Sed12l]. capture-shrinkage [YWC^{+}13]. carcinoma [GWR^{+}10, IWS^{+}11, VES^{+}11]. carcinomas [KDIE11]. cardiac [CCM^{+}11, DWM^{+}12, SZW^{+}11, TKS^{+}13]. Cardiolipin [COB^{+}12]. cardiomyocyte [RBP^{+}13]. cardiomyopathy [HL11, TPSS12]. cardiovascular [BTC^{+}11]. Career [Les14b]. Cargo [HKN^{+}11, MLY^{+}10, BSR^{+}14, EJBW12, KBC^{+}14, LYH^{+}13, MBCKD13, PHB^{+}11, SAoS14, SDD^{+}13, SP11, ZQA^{+}14, vBAK^{+}12]. Cargo- [MLY^{+}10]. cargoes [EJBW12]. Carl [Sed13d]. Carl-Philipp [Sed13d]. Carla [Sed13e]. Carol [Sed11d]. Carole [Sed14e]. carrier [COB^{+}12]. carriers [CLW^{+}14, HKR^{+}10, LMS^{+}10c]. CARTS [WVvG^{+}13]. carvedilol [oHXK^{+}12]. carvedilol-bound [oHXK^{+}12]. cascade [LKLA12, MJFS10, WOG13, ZMW^{+}13]. cascades [TKS^{+}13]. case [Sho13-47]. Casein [GR11, ZEG11]. CASK [SSK^{+}13]. Casp8p41 [SDN^{+}14a, SDN^{+}14b]. Caspase [FA12, Les12g, CO13, DGH^{+}14, NTSK14, Sho14-35, BMS^{+}11]. Caspase-8 [BMS^{+}11]. cast [Les10h]. Casting [Sho10l]. catalytic [GVP^{+}11]. catalyzing [NOS^{+}14]. catches [Les10f, Les10g, Les14v]. Catching [Les13f]. Catenin [Les13a, RFL13, Sho10b, BKY^{+}10, CSP^{+}10, Les10a, MH11, NCML^{+}12, PMF12, SME^{+}13, YMU^{+}10, YMU^{+}13, BAY^{+}11]. cathepsin [GGSN^{+}13]. cation [SHB^{+}10]. Causal [Bar13]. cause [CWS^{+}11, CFB^{+}12, Les12r, LgLM^{+}10, Sho10h, NGM12]. caused [DPZ^{+}14, RCFH10]. causes [BFG^{+}13, CZC^{+}11, CZGG12, ERS10, PoLC^{+}13, STI^{+}11, SPD^{+}13, XHS^{+}13, YWJ^{+}12, vRJMvD10]. causing [LNT^{+}10]. caveola [JBS^{+}12, JBS^{+}13]. Caveolae [AFRZ^{+}14, KWH14]. caveolin [PH10, HSR^{+}10, Les11f, MLG^{+}10, SKV^{+}11]. Caveolin-1 [HSR^{+}10, Les11f]. Caveolin-1-dependent [MLG^{+}10]. Caveolin-1-eNOS [SKV^{+}11]. caveosome [PH10]. cavernous [GLG12]. cavitation [QTL^{+}12, QTL^{+}13]. cavity [AMGC14]. Cbk1 [KKK^{+}11]. Cbp3 [GKR^{+}11a, GKR^{+}11b, GRH^{+}12]. Cbp6 [GKR^{+}11a, GKR^{+}11b, GRH^{+}12]. CCAN [HSTF13]. CCCTC [ZNH^{+}11]. CCM [Sho13i]. CCM1 [FRL^{+}13].

CD2AP [TB13]. CD4 [SDN^{+}14a, SDN^{+}14b]. CD40 [KDIE11]. CD63 [FCA10]. CD95 [FBR^{+}10]. CDC [KSH^{+}13]. CDC-48 [KSH^{+}13]. CDC-48/p97 [KSH^{+}13]. Cdc12 [CSKW13]. Cdc14 [Les10t, MMV^{+}10, RGF^{+}10, Sho10m]. Cdc14A [MBZ^{+}10]. Cdc14B [MBZ^{+}10]. Cdc20 [IP12, MGT^{+}10, MJJ^{+}10, Sho10-33]. Cdc25 [OHC10]. Cdc28 [VCF^{+}13]. Cdc28/Clbs [VCF^{+}13]. Cdc42 [AYS^{+}13, BT13, EZT^{+}12, JDL^{+}14, KLP14a, KWH^{+}10a, KWH^{+}10b, Les13g, MI13b, NLP^{+}10, OKNP13, OPCEM10, QWL^{+}11, QMHM10, RIG^{+}12, RFVE^{+}10, Sho11j, Sho12i, Sho14e, SRP^{+}13, VLI^{+}14, ZME^{+}14]. Cdc42-selective [NLP^{+}10]. Cdc48 [KBW^{+}10]. Cdc48/ [KBW^{+}10]. Cdc55 [BM11, Les11-32, RY11]. CDC6 [ZMW^{+}13, Les11g, SZE^{+}11, ZSH10]. CDH1 [LS13b]. Cdk [VYC^{+}11, EM11, GZR^{+}14a, GZR^{+}14b, GC13, GSJS10, KST^{+}11, TGB10]. CDK-1 [EM11]. CDK-dependent [GC13]. Cdk-mediated [VYC^{+}11]. Cdk1 [vZOtR^{+}10, DSL13, NMB^{+}14, Sho13j, ADB^{+}14, CLO^{+}11, GP10, Lin10, MGK^{+}12, OMW^{+}14, PLC^{+}11]. CDK2 [ZMW^{+}13, CVJ^{+}11]. CDK5 [Sho14h, TQM^{+}14, HSS^{+}13, WDG^{+}13]. CDK5RAP2 [BKG10, CLS^{+}10]. CDKN3 [NBSE^{+}13a, NBSE^{+}13b]. Cdks [Les10c, Sho11k]. Cdo [RSRK13]. Cdr2 [RBB^{+}14]. CDT2 [JEF^{+}11]. Cédric [Sed12d]. Cell [ABVP11, AMO^{+}11, BM10, BWM12, CM12a, Gil10, HSF12, KMSR12, Orr12, PWP11, PTBT10, PCCR11, RG14, RFAA^{+}12, Sho11l, Six12, WGN^{+}13, ALV^{+}12, ANT^{+}12, AFM^{+}13, APC^{+}13, ABP^{+}12, BLO^{+}12, BMÁG^{+}14, BPT^{+}14, BHMBS^{+}11, BDR^{+}12, BKY^{+}10, BvMD^{+}14, BV11, BEJ10, BFG^{+}13, BC11a, BC11b, BTC^{+}11, BGS13a, BGS13b, BW12, BDvdK13, BLC^{+}14, BKE10, CWL^{+}11a, CPS^{+}13, CDAK10a, CDAK10b, CC10a, CHL12, CR10a, CR10b, CTL^{+}10, CYN^{+}13, CG12b, CGW^{+}11, Coo13, CSM^{+}12, CBB12, DV10, DN10, DWPC^{+}11, Dev14, DGF^{+}14, DZT^{+}11, EAB^{+}14, EBBJ11, ES14, FPAM13, FSK^{+}10, FA12, FW10, GdBP^{+}14, GHC^{+}14, GvEM^{+}11, GSU^{+}12, Gol12b, GCC12, GM11, GFSR11, GNHB11, GCV^{+}11, GS11, HKN^{+}14, HHJ^{+}11, HZS^{+}10, HSJ^{+}13, HCC^{+}10, HDH^{+}10, HYTU^{+}10, HKR^{+}14, HKN^{+}10, HK14, HCG^{+}11, IMP^{+}12, IIWS14, JOR^{+}11, KEJ13, KHB^{+}11a]. cell [KPC^{+}11, KKL^{+}11, KPSL12, KN12, KCF^{+}14, KFS^{+}14, KBW^{+}12, KTB^{+}14, KFL^{+}14, LAR^{+}12, LJW13, LLS^{+}11, LCBG^{+}11, LVK^{+}13, LLU^{+}12a, LLU^{+}12b, LDCF^{+}13, LSGVM14, LGM^{+}12, Les10-27, Les11n, Les11y, Les12d, Les13-27, Les13-43, Les14g, Les14f, Les14s, Les14-29, Les14-32, LNTR14, LM13, LNS^{+}13, LZW^{+}10, LSCF11, LZY^{+}12, LT11, LR11b, LP11, LWB^{+}14, LP13, LMN10, LZLG13, LXTM12, MWH12, MLM^{+}11, MGG^{+}12, MJEM10, MMdCOM^{+}11, MRA14, MBM^{+}10, MHAK^{+}12, MFF^{+}13, MSK^{+}13b, MRLLS12, MBO^{+}14, MVN11, MCS^{+}13, MAD^{+}11, NGL^{+}12, NSZ^{+}13, OVW10, OOKH^{+}12, OLB13, OPCEM10, OLT11, Pal14, PGCY12, PPV^{+}14, PRM^{+}14, QJO10, RLS^{+}14, RMT13, RCG^{+}10, RCG^{+}11, RKE14a, RKE14b, RDC^{+}11, RKK^{+}14, Sab10, SBEM13, SA10b, SRS10, SPC^{+}13, SSdA^{+}14, SRBL13, SKM10b, SZW^{+}11, Sed10e, Sed10k, Sed11f, Sed11i, Sed11o, Sed12i, Sed12o, Sed14d, Sed14j, Sha10]. cell [Sho10-31, Sho10-45, Sho10-61, Sho10c, Sho13-56, Sho14t, Sho14-28, Sho14-39, Sho14-60, SMK14, SL14, SMS^{+}13, SEV^{+}14, SGT^{+}13, SFL12, SC10a, SRU^{+}12, SMZL13, TDV^{+}14, TLSA14, TL12, WVvG^{+}13, WM10, WBS11, WWM^{+}12, WRF^{+}13, WPM14, WPSA13, WHH^{+}11, WtLK^{+}13, WHWS12, WLW11, YHG^{+}14, ZBJL^{+}10, ZDM^{+}14, ZSK12, ZPS^{+}10, ZEG11, vGLWB12, Nan14, Sed13c, CFLDM11, SHS^{+}13]. cell-autonomous [KHB^{+}11a, KPC^{+}11]. cell-intrinsic [BHMBS^{+}11]. cell-killing [Les12d]. cell-surface [RKK^{+}14]. CellGeo [TBV^{+}14]. Cells [Sha10, Sho10n, AMH11, AIJI11, ADAB^{+}12, BT12, BHMBS^{+}11, BMS^{+}11, BDB^{+}14, BKK^{+}10, BSO^{+}14, BKE10, CTM^{+}14a, Car12, CLL^{+}10, CLEZ12, CC10b, CZ10, CSHS^{+}13a, CSHS^{+}13b, CFB^{+}12, CWFL13, CSH^{+}12, DMH^{+}12, ECC^{+}13, ER10, FEHF12, FP10, FHD^{+}12, FSA^{+}10a, FSLM11, GCP^{+}14, GvEM^{+}11, Gol12b, GGSN^{+}13, HBC^{+}11, HRK13, HKI^{+}13, HKR^{+}10, HIB^{+}10, IWS^{+}11, IMG^{+}12, KF11, KNPK^{+}10, KOO^{+}14, KCF^{+}14, KBG12, KRS11, KSS^{+}11, LRH^{+}13, LKLA12, LR13, Les10o, Les10-42, Les10p, Les10-29, Les10-36, Les11e, Les11t, Les11w, Les11-33, Les12g, Les12n, Les12u, Les13b, Les13i, Les13g, Les13n, Les13p, Les13t, Les13-31, Les13-37, Les14b, Les14c, Les14d, Les14r, LMW^{+}11, LZR^{+}11, LSE^{+}10, LKG^{+}13, MBZ^{+}10, MH14, NT11, NSS^{+}10, NSD^{+}14, NBS^{+}11, OS13, OSD^{+}14, OWC^{+}10, OB12, dJPAA^{+}11, PJS^{+}11, PWP11, PTBT10, PPG11a, PPG11b, PCCR11]. cells [QB12, QECC10, RZS^{+}14, RIG^{+}12, RKRB12, SDN^{+}14a, SDN^{+}14b, SHC^{+}10, SW10b, Sed11q, Sed12d, Sed12g, Sed13j, Sed14e, Sed14f, Sed14h, Sho10h, Sho10w, Sho10-37, Sho10-38, Sho10-56, Sho10-69, Sho11-47, Sho11-50, Sho11-55, Sho12i, Sho12-41, Sho12-50, Sho13b, Sho13-55, Sho13p, Sho13o, Sho13n, Sho13-31, Sho13-47, Sho13-52, Sho14-43, Sho14-65, SPF11, SPD^{+}13, SRZ^{+}11, TAGJ11, TID^{+}10, TPZ^{+}14, TAC^{+}13, TC10, TP13, TBV^{+}14, VB12, VMNLB^{+}11, VOSB12, WHF^{+}11, WBBD14, WHH^{+}11, Yam13, YWC^{+}13, YST^{+}11, ZTBK14, ZLW^{+}13, ZSH10, ZLH^{+}14]. Cellular [CLS13, HL11, LCP13, RK13, BS13, CBB12, DD10a, DCO^{+}13, FSA^{+}11, HPB^{+}12, HVDG13, Les10i, Les10-37, Les12z, Les13j, Les14-33, Leu14, LhYL^{+}13, MKS^{+}13, Oka14, PBvdS12, Pow14a, RMG^{+}12, RC11, RH10, SZ12a, SSD^{+}14, Sed14a, TPSS12, TW14, XTH^{+}11]. CENP [LBS11, BKS^{+}11, BSR^{+}11b, BSR^{+}11c, CMS10, CDB^{+}14, CWPW11, GKA^{+}12, HSTF13, HTM^{+}14, MS14, MMFS11]. CENP-A [CWPW11]. CENP-C [MMFS11]. CENP-E-dependent [GKA^{+}12]. CENP-I [MS14]. center [BWM12, Sed13u, YKT^{+}13]. Centered [Sho10-39]. central [BGB^{+}13, BCB14a, BDR^{+}12, CVJ^{+}11, DKM^{+}13, DS10, Les13-27, OYYK14, RTC^{+}13a, RTC^{+}13b, Sho10i, Sho13c, UG10, UTK^{+}13]. centralspindlin [LCD^{+}11, MZP^{+}10]. centrins [DWL^{+}11]. centriolar [BIY^{+}13]. Centriole [KFS^{+}14, BKBR11, FRS^{+}13, GZLG11, HKH^{+}10, JVS^{+}14, KOO^{+}14, KD11, LWL^{+}13, MXS10, SKN^{+}12, Sho11-59, SDB^{+}10]. centrioles [CSAPLBD11a, CSAPLBD11b, Sho10-39, SPD^{+}13, TYN^{+}13, WSUT11]. Centrobin [GZLG11]. centromere [BWS^{+}10, CMS10, CWPW11, HSTF13, HZW^{+}12, JKA^{+}10, KIOY10, LBS11, Sho11x, SHV^{+}11, YTT^{+}10]. centromere-specific [LBS11]. centromeres [HZW^{+}12, LBS11, MMFS11, Sho13-60, Sho14g, WBL11]. centromeric [BKS14, RKE14a, RKE14b, WUD^{+}12]. centrosomal [FCE^{+}12, KSH^{+}13, MLBY^{+}10]. Centrosome [TQS^{+}11, YWC^{+}13, ADAB^{+}12, BIY^{+}13, BKG10, BKK^{+}10, CAB^{+}10, DWL^{+}11, DHVK10a, DHVK10b, FCE^{+}12, GCP^{+}14, GR11, HLN^{+}10, HCG^{+}11, KRS11, KFL^{+}14, LR11a, Les13-39, MJEM10, MS12, RTC^{+}12, RFAA^{+}12, Sho12k, Sho13-43, SC10b, TMS^{+}12, ZEG11]. Centrosomes [Les10d, Sho12j, BBD^{+}11, KFL^{+}14, LR13, Sed13z, WSUT11]. CEP120 [LWL^{+}13, MXS10]. Cep152 [CAB^{+}10, HKH^{+}10, Sho10o]. Cep164 [SKN^{+}12]. CEP290 [CTD^{+}10, KKL^{+}14]. Cep55 [BB10]. cerebral [GLG12, HLL^{+}12]. cerevisiae [AKC^{+}12, GSJS10, HBSD12, CT10, MWZ^{+}11, NSBW10, SMMB11, WMCF10, YCP10]. CFTR [OL12]. cGMP [WOG13]. cGMP-dependent [WOG13]. ch [HWB^{+}13, FCE^{+}12]. chain [DDH^{+}12, DHL^{+}12, ECJB10, Les14t, LHS10, RSD^{+}12, SNZVK12, SNZVK13, SLC^{+}13, VGL^{+}14]. chains [JVS^{+}14]. challenge [LRB13]. Chang [Sho11z]. change [ADF^{+}12, DGF^{+}14, Sho11-34]. changes [AOE^{+}10, AOE^{+}12, CLEZ12, KEJ13, KNOM11, LJPJ11, RZS^{+}14, SHN^{+}11, TBV^{+}14, YON^{+}12]. channel [BNM^{+}14, BLI^{+}10, CZD^{+}13, DSP11, IIN^{+}11, JAM^{+}13, MVC^{+}11, TMG12, VYC^{+}11, VBG^{+}13, YSM10]. channels [HJ14]. chaperone [ARF10, BNH12, CPT^{+}12, CPT^{+}14, DGS^{+}10, Sho12l, ZLJ^{+}13]. Chaperones [Les12h, KPI^{+}10, WTBM12]. Chaperoning [Kar10]. Characterization [KIL^{+}12, SBP^{+}10a, AMS^{+}13, KSP^{+}11, WMB^{+}10, iYGL^{+}10, YST^{+}11]. Charcot [LCL12, Sho12-55]. Charcot-Marie-Tooth [LCL12]. charge [XBC^{+}13]. chart [Sho10p]. chauffeur [Les12p]. CHC22 [ECJB10]. CHD [KLP^{+}14b]. CHD-class [KLP^{+}14b]. CHD4 [LPG^{+}10]. checking [Les14i]. checkpoint [BRL14, BKP11, CKO^{+}10, COW13, CGRS^{+}12, DP10, EUB^{+}14, ECK^{+}12, GKA^{+}12, KSB^{+}13, KFH^{+}12, KWL^{+}12, LWK^{+}13, LDL12, MGT^{+}10, MMU10a, MBZ^{+}10, MCHCC10, MKH^{+}14, PoLC^{+}13, RJM^{+}12, RGF^{+}10, STD^{+}10, SFK^{+}13, SLM^{+}13, Sho13g, Sho13u, Sho14-57, VSMC11, VYM^{+}10, VWC^{+}13, VCF^{+}13, WGC11, XHB^{+}10, YFO12, dSMSS13]. checkpoints [MGS14]. chemical [LC10]. chemistry [PMB^{+}11]. Chemokine [Sho13k]. chemoreceptor [PBPW^{+}14a, PBPW^{+}14b]. chemotactic [WHA^{+}13]. chemotaxis [ADF^{+}12, BMRM13, CDK^{+}10, VLKI14, ZTBK14]. Chen [Sed14f]. chew [Sho11-35]. Chibby [BLC^{+}14, ETC^{+}12, Sho14i]. chicken [KNH^{+}10]. chitinase [Sho14-67]. chitinolytic [HMO^{+}14]. CHK [KHW^{+}10]. CHK-1 [KHW^{+}10]. Chk1 [GB10]. Chk2 [PZ14]. Chlamydia [MBM^{+}10, Sed14u]. Chlamydia-induced [MBM^{+}10]. Chlamydomonas [BYY^{+}12, MBLD11]. chloroplast [TO12]. cholera [GCV^{+}11]. cholesterol [DKF^{+}11, HVDG13, Sho13-52]. chondrocyte [NSB^{+}11]. choose [Les12n]. CHOP [BAAW11]. CHOP-regulated [BAAW11]. choreography [Ewe11, Sed12b]. chosen [Sed12y]. Chris [Sed13f]. Christine [Sho10p]. chromaffin [LSE^{+}10]. chromatid [KMSR12, OYH13, STI^{+}11, Sho11-60]. Chromatin [SRBL13, BKS^{+}11, BG10, BZ12, Car12, CCJ^{+}12, FP10, FMG^{+}11, GL10, IPM^{+}13, KAAM11, KLP^{+}14b, KSS^{+}11, LGAC13, LPG^{+}10, yLFAM13, Les13q, Les13r, LLM^{+}10, LLA^{+}12, MMFS11, PASG^{+}12, RSM^{+}13, RGF^{+}10, SKM10a, SNSyN13, Sho13y, SBP^{+}10b, SWV^{+}10, SSK^{+}14, SLC^{+}13, SHV^{+}11, SHV^{+}13, TCB^{+}14, XSJ^{+}10, ZNA^{+}14]. chromatin-bound [GL10]. chromatin-remodeling [LPG^{+}10, RGF^{+}10, SKM10a]. chromatoid [YOA^{+}11]. chromokinesins [WBS^{+}12]. Chromosomal [HRK13, EIE^{+}14, LCD^{+}11, SPD^{+}13, ZF11]. Chromosome [RHK11, TIM14, AMS^{+}13, BM11, BAS^{+}14, CLO^{+}11, DKY^{+}12, DLBG11, GCR^{+}12, GKA^{+}12, HSN^{+}11, IP12, KNW^{+}14, LGAC13, STD^{+}10, SMS^{+}14, SBP^{+}10b, VTO^{+}13, WBS^{+}12, WUD^{+}12, WHL^{+}12, WRCD12, XOV^{+}10, YSO^{+}11, YRU^{+}13, vRJMvD10]. chromosome-specific [DKY^{+}12]. chromosomes [EHUD14, HKW^{+}13, KPC^{+}10, KIOY10, LH11, Les10-38, Les11r, Les13s, Les13-44, SSV^{+}12, Sho12z, Sho12-28, Sho12-51, Sho13-27, SST^{+}12]. Ci [ZFP^{+}13]. CIIA [HHC^{+}11]. cilia [BKS^{+}13, CSAPLBD11a, CSAPLBD11b, GCR^{+}13, HBM^{+}11, IMG^{+}12, KKL^{+}14, LBS^{+}13, Les11i, LAH^{+}12, PBM^{+}11, Sho12e, Sho12-60, Sho14-66, TLS10]. cilia-associated [HBM^{+}11]. ciliary [BLC^{+}14, CHK^{+}10a, CHK^{+}10b, DS10, FLVP10, FSLM11, GBK^{+}14, GGR12, HBG^{+}11, JYRL^{+}13, Kin13, LZW^{+}12, LZW^{+}13, Omr10, SW10b, WLK^{+}11, YSaY^{+}13]. ciliary/flagellar [Kin13]. ciliogenesis [ETC^{+}12, GGR12, HCG^{+}11, KSR^{+}13b, LWZ^{+}10, PTBT10, SKN^{+}12, WLK^{+}11]. ciliopathy [MGG^{+}12]. cilium [Sho10-36]. CIN [Sho10-44]. cingulin [YMT^{+}13]. CIP2A [PRM^{+}14]. circuit [CF13, MMV^{+}10]. circuitry [FPM^{+}14, Sed12o]. circular [GNHB11]. Circulating [YST^{+}11]. circumferential [NT11, OOKH^{+}12]. Circumventing [Sho14j]. Cis [GZZ^{+}14, BKAB13]. Cis-acting [GZZ^{+}14]. cis-elements [BKAB13]. cisternae [LLK11]. cisternal [RPM^{+}13, XW10]. Citron [BVC^{+}11]. CK1 [GOWM12]. CK1- [GOWM12]. CK2 [BRP14, ILD^{+}10]. Cks [vZOtR^{+}10]. clamp [GL10, WMP^{+}14]. CLAMP/Spef1 [WMP^{+}14]. CLAMPing [Les14c]. Clarin [OZ14]. Clarin-1 [OZ14]. Clasp [SMM^{+}10, FSOL14, NSS13, Sho13p]. Clasp-mediated [SMM^{+}10]. CLASP2 [MGK^{+}12, SBS^{+}12, SME^{+}13]. CLASP2-mediated [SBS^{+}12]. CLASPing [Les12i]. CLASPs [Sho13-39]. Class [BCBG10, YDB^{+}11, vGLWB12, DCP^{+}10, KWDD10, UAH^{+}12, VJK^{+}10a, VJK^{+}10b, YHF13, KLP^{+}14b]. classes [SNR^{+}11]. Classic [Sed12w]. classical [NCML^{+}12]. Clathrin [BVM^{+}11, FCE^{+}12, HKR^{+}10, KPJ^{+}13, Les10e, LHS10, Sho12k, ECJB10, GK13, GHK^{+}10b, GCH^{+}14, LADS10, MBCKD13, MLY^{+}10, NPL^{+}10, PDMBW11, Sed13i, SNT^{+}12, Sho12-44, VGL^{+}14, KKS^{+}14, Sho14k, HWB^{+}13]. clathrin-coated [LADS10, NPL^{+}10, SNT^{+}12]. Clathrin-independent [HKR^{+}10, MBCKD13, PDMBW11]. clathrin-mediated [GK13, GHK^{+}10b, GCH^{+}14, MLY^{+}10, Sho12-44]. Clbs [VCF^{+}13]. clean [Sho11-39]. cleans [Les11-46]. Clearance [ER10, LNS^{+}13]. Clearing [Les14d, KMSR12]. Cleavage [BBW^{+}14, BVC^{+}11, CG12b, FLN^{+}10, FLN^{+}16, KMSR12, KRS11, RZF^{+}11, Sho13o, Sho14a, SML^{+}13, SGT^{+}13]. cleave [AWB^{+}14]. cleaves [BXB^{+}12]. CLIC [Sho10-37]. client [KXN10]. CLIP [Les10f, Les10g]. Clk1 [NKH11]. Clk1/4 [NKH11]. clonogenic [LRH^{+}13]. close [Les13-31, Mit12a, Sho10k, VWC^{+}13]. closed [BMLB^{+}12a, BMLB^{+}12b]. closeness [Les10-37]. closer [Sed13m, Sed14f]. Closing [Les14e]. closure [GEN14, LN11, Les11k, SNT^{+}12, Sho11l, YSN^{+}11]. closures [Les13d]. club [Les14o]. CLUH [GSM^{+}14, Sho14-40]. cluster [HTT13, MAE^{+}10]. Clustering [RDLT11, AMH11, ALSN^{+}11, BKK^{+}10, HZW^{+}12, LADS10, SCR12, WLN^{+}14, YKT^{+}13]. clusters [CWPW11, KKMB10, Les14w, RZS^{+}14, SSD^{+}14]. CNTD1 [HSY^{+}14, Sho14l]. Co [KdKDP12, RMG^{+}12]. Co-operation [KdKDP12]. co-opts [RMG^{+}12]. CoA [KHFV^{+}13, NOT^{+}14]. Coa3 [MVP^{+}10]. coat [FSR11, JKS14, Sho14-48, WS10]. coated [BAH^{+}12, LADS10, NPL^{+}10, Sed13m, SNT^{+}12, CLW^{+}14]. Coatomer [BPDB^{+}11]. coats [Sed14n]. code [Jan14, MPRT11]. coenzyme [Sho12-36]. Cofilin [KBC^{+}14, CP11, HKN^{+}14, HTS11]. COG [LHL11, Sho11m]. Cohen [Sed12r, Sed14a]. Cohesin [Les12j, Sho11n, SHV^{+}11, LH11, LHGT^{+}12]. Cohesinopathies [BG10]. cohesion [JVS^{+}14, Sho10-59, YTT^{+}10]. coil [ŻKC^{+}11]. coiled [ŻKC^{+}11]. coiled-coil [ŻKC^{+}11]. coiling [DGF^{+}14]. coincides [BTL^{+}12]. Coleen [Sed12e]. coli [OBD^{+}10, WHF^{+}11]. collaborate [YFO12]. collaborates [WGC11]. collaborative [GJP^{+}13]. collagen [BBJ^{+}10, ISZ^{+}11, Les11-27, MLM^{+}13, MHAK^{+}12, SYS^{+}14, WPL^{+}11a, WPL^{+}11b]. collapse [JDHS10, Les11e]. Collaring [Sed10o]. colleagues [dWMR10]. collective [KTB^{+}14, LAR^{+}12, LKG^{+}13, NBDB12, PCCR11, Sho14t]. collectively [GHK^{+}10b]. colon [MLH12, Sed13g]. colonies [VŠH^{+}11]. combined [HBC^{+}10, MFB12]. come [Sho14-46]. comes [Sed14p]. commitment [LZR^{+}11, SAoS14, Sed12o]. committed [CC10b]. common [LNL11, ZKR^{+}11]. communication [ISZ^{+}11, LXTM12]. communities [Sed13t]. commute [Les11w]. compacted [SSV^{+}12]. companion [Les11l]. Comparative [RSL^{+}11]. comparisons [RZA^{+}13]. compartment [BMC^{+}11, MGR^{+}10, Omr10, SBTF13]. compartments [CGCP^{+}14, LLH13, OHC10, OLT11]. compass [Les12-34, Sho10-58]. compete [IP12, Sho12-38]. competition [LM13, SYV14, Sho10-54]. Competitive [MVP^{+}11, Sho11-49]. complement [SRKR10]. complementary [XW10]. Complete [GSB^{+}13, ZZS13]. completion [ADS^{+}13]. complex [AIJI11, AKA^{+}13, BDR^{+}10, BTL^{+}12, BAY^{+}11, BGY^{+}13, CHS^{+}10, CLS^{+}10, CPX11, CWS^{+}11, CFB^{+}13, DCN^{+}10, DLBG11, EZT^{+}12, EBBJ11, Ewe11, FRL^{+}13, FMPS^{+}12, GBiY^{+}14, GLG12, GTR^{+}13, GRH^{+}12, GDS^{+}12, HBSD12, HTS^{+}10, HZE^{+}13, HCCS^{+}11, HSKAT11, HS10b, IHM13, JGB^{+}13, KXN10, KA12, KWDD10, KPI^{+}10, KBKW10, KHB^{+}11b, LHW10, LHL11, LVK^{+}13, LCD^{+}11, LAH^{+}12, LvBG^{+}10, MHS10, MMS^{+}10, MMC^{+}10, MI13b, MPM11, MKH^{+}14, NEMH^{+}10, NSD^{+}14, OSD^{+}14, ONH^{+}12, PR12, PLR^{+}13, PPD^{+}10, RNS^{+}14, RMG^{+}12, RPM^{+}13, RGF^{+}10, SKM10a, SSdA^{+}14, SKH^{+}10, SLM^{+}11, Sho12n, Sho13i, SDS^{+}12b, SWV^{+}10, SKFH11, SMB12, SRU^{+}12, SIO10, TKB^{+}14, TESA10, TCN14, TBV^{+}14, UHKS11, WM12, WWM^{+}12, WMP^{+}14, WGR^{+}12, WDB10, WGM^{+}12, XOV^{+}10, XZC^{+}12, YSaY^{+}13, YZPF12, YSN^{+}10, ZYF^{+}11, ZFP^{+}13, ZLJ^{+}13, ZNA^{+}14, CJNS12]. complex-dependent [YZPF12]. complex/cyclosome [NEMH^{+}10]. complexes [CWFL13, DK10a, FP10, GSM^{+}12, KBS^{+}10, LMA^{+}13, Les13w, LYH^{+}13, LP11, LHGT^{+}12, LWB^{+}14, MGT^{+}10, MLW13, MRCC^{+}13, OPM^{+}12, RZF^{+}11, RCG^{+}10, RCG^{+}11, RHK11, SRS10, Sho14f, TUG^{+}10, TLS10]. Complexin [DYS^{+}14]. complexity [MVC^{+}11]. component [GDS^{+}12, TAC^{+}13, TL12]. components [Jan14, WBcY^{+}11, WGM^{+}12]. composite [YSO^{+}11]. composition [HFS10, SSD^{+}14]. Compression [KEJ13]. comprises [JKS14]. compromised [Sho10a]. computational [DV10, LHN10, TBV^{+}14]. Concentration [SYS^{+}14, ADF^{+}12, CWL^{+}11a, KNOM11]. concept [BT12]. concert [KBW^{+}10]. condensation [KNW^{+}14, YSO^{+}11]. Condensin [FP10, Les13h, OYH13, BSR^{+}11b, BSR^{+}11c, BDN^{+}13, KNW^{+}14, Les14u, SSV^{+}12, Sho10-69, Sho11n, SSK^{+}14, SHV^{+}11, YSO^{+}11]. condition [Sho10z]. conductance [SNT^{+}12]. cone [JDHS10, TKMK10, VPC^{+}14]. confer [DBUT13, PBPW^{+}14a, PBPW^{+}14b]. confers [CSM^{+}12, LHW10]. confined [HCP^{+}13]. confinement [EZT^{+}12, RFAA^{+}12]. conformation [AMS^{+}13, MLSM^{+}11]. Conformational [YON^{+}12, Boe12, CLEZ12]. conformations [Sho11g]. Congenital [ZYH^{+}11]. congression [WBS^{+}12]. conjugating [vRJMvD10]. connected [Sed14k, Sho10-64]. Connecting [Sho11o, COW13, HH10, SRZ^{+}11]. connection [Les10p, Les11o, Sho12-35]. connections [Sed13g, SMS^{+}14]. connectivity [CF13, DNB13]. Conquering [Sho11b]. consecutive [KFS^{+}14]. consequence [NGM12]. consequences [RC13]. Conserved [MOZ^{+}13, DBH^{+}11, MSZ^{+}12, NCML^{+}12, RSL^{+}11, RFRV12a, RFRV12b, SDB^{+}10, TDV^{+}14, YSaY^{+}13, YSN^{+}10]. consistent [SMZL13]. conspiracy [Les12g]. constitute [AiIK^{+}13, SRKR10]. constitutes [HDK^{+}13]. constitution [Sho11s]. Constitutive [YSM10]. constrain [TIM14, YZPF12]. constrained [KPC^{+}10]. constrains [BMLB^{+}12a, BMLB^{+}12b]. constraints [BLT^{+}11]. constriction [AVP^{+}14, APC^{+}13, CWL^{+}11a, CP11, LYB^{+}10]. construction [Sho12w]. consumption [Les14p]. contact [GSW^{+}11, IMP^{+}12, PvdLA^{+}14, Pri14, SMM^{+}10]. contacts [FWM^{+}10b]. contain [LBS11]. containing [CLSO^{+}12, JKS14, KLC^{+}10, NvCL^{+}13, RMG^{+}12, RGB^{+}13, HBC^{+}11, MGR^{+}10]. contains [FLVP10, WDG^{+}13]. content [CG10b, CTD^{+}10, VvDV^{+}10]. context [Sho12-45]. Continued [VYM^{+}10]. Contractile [LNTR14, BMÁG^{+}14, BSO^{+}14, CP11, CSKW13, EKJH13, LCD^{+}11, MGFG^{+}10, MMVK^{+}12, VTM14]. contractility [FRL^{+}13, NBDB12, NWD^{+}11, PTBT10, dMSMZ14]. contraction [ELH14, MCS^{+}13]. contracts [Sho10-41]. contraptions [Sed14v]. contrasting [PMB^{+}11]. contribute [LCD^{+}11, WGM^{+}12]. contributes [ECK^{+}12, GBSC^{+}12, JK10, LN14, PHD^{+}10, RCFH10, RMF^{+}10, VYM^{+}10]. contributing [MVC^{+}11]. contribution [BR14]. Control [ASB^{+}11, BW12, ANT^{+}12, AXL10, AFRZ^{+}14, Bab14, BEJ10, BCB^{+}14b, BKE10, CPS^{+}13, CWL^{+}11b, CZM^{+}14, CLD11, CF13, EMT^{+}14, FMG^{+}11, GC13, HOS^{+}12, Kin13, KLS^{+}13, KBAW^{+}11, LJW13, LLS^{+}11, LLU^{+}12a, LLU^{+}12b, Les10m, LJLJ11, LOR^{+}10, LAB14, MGT^{+}10, MWP^{+}12, MFGB10, MJEM10, MMdCOM^{+}11, MRCC^{+}13, MAD^{+}11, OHC10, PL10, QECC10, RJM^{+}12, RBB^{+}14, RHK11, RFC14, RBM^{+}11, SZ12a, SZW^{+}11, Sed12v, SLM^{+}13, Sho10v, Sho12g, SL14, SHS^{+}13, SW12, TSB^{+}14, UTK^{+}13, VWD^{+}13, VLI^{+}14, WP14, WGC11, WHDR^{+}10, YMM^{+}10, ZDM^{+}14, vZOtR^{+}10, ABP^{+}12, WtLK^{+}13]. controlled [KYOY13, TQS^{+}11]. controlling [ALV^{+}12, AFM^{+}13, OPM^{+}12, OMV^{+}11, RGF^{+}10, ZBBG10]. controls [ALS^{+}13, AAE^{+}14, ADF^{+}12, BMRM13, BAAW11, BDR^{+}12, BBJ^{+}10, BSO^{+}14, CJNS12, CSS^{+}14, DKMK^{+}11, FRL^{+}13, GOWM12, GYZ^{+}12, GWR12, HSJ^{+}13, HBSD12, HSK^{+}10, JDL^{+}14, KNW^{+}14, KKUG11, LGAC13, LZR^{+}11, LLcK^{+}11, MdFF^{+}14, MSS^{+}12, MRR^{+}12, NMB^{+}14, NBSE^{+}13a, NBSE^{+}13b, NSS^{+}10, NSZ^{+}13, OYYK14, PBD^{+}13, PGAE^{+}13, PRM^{+}14, RCM^{+}12, RBA^{+}11, RFVE^{+}10, RFAA^{+}12, RY11, SMdP^{+}14, SFJ^{+}14, SKH^{+}10, TKS^{+}13, TOI^{+}13, TESA10, TCN14, UKZ^{+}13, XRO^{+}11, ZFA^{+}13, ZEG11, vdBFS^{+}12, vdVMG^{+}11]. Convergence [YMU^{+}10, YMU^{+}13]. convergent [KBW^{+}12, dMSMZ14]. converging [MMB^{+}11, SOW^{+}11]. conversion [WSUT11, ZWL^{+}14]. convertase [MC10]. conveys [CLL^{+}10]. cooking [Pow14a, Sed13w]. cooperate [BGC^{+}14, CSKW13, EMT^{+}14, LVK^{+}13, MGS14, PCO^{+}10, QTL^{+}12, QTL^{+}13, SGLV10, VŠH^{+}11, WLK^{+}11]. cooperates [KF11]. Cooperation [TUG^{+}10, ETYS^{+}12]. cooperative [HTM^{+}14]. coordinate [AIJI11, FPAM13, HSN^{+}11, LWZ^{+}10, Sho13-60]. Coordinated [AVP^{+}14, APC^{+}13, HCC^{+}10, BPL^{+}11, MCHCC10, RKG^{+}10, SSH^{+}13, TLTW10]. coordinately [GC13]. coordinates [BSR^{+}14, BM11, EJBW12, GRH^{+}12, LPG^{+}10, PSVRB^{+}11, Sho13-51, YZM^{+}12a]. coordinating [mFH13]. Coordination [HWB^{+}13, LAR^{+}10, SHS^{+}12, YOMM^{+}11]. coorientation [YTT^{+}10]. COPII [OTLH10, WS10]. COPs [FSR11]. copy [CG10a, JJH^{+}10, PR12]. cord [LSGVM14, Les13m]. core [DMD^{+}12, HTS^{+}10, HSTF13, LNL11, MMC^{+}10, PMP^{+}11a, PMP^{+}11b, RSL^{+}11, Sed10p, Sho12-48, TKB^{+}14, VWC^{+}13, vdBFS^{+}12]. coregulate [VJK^{+}10a, VJK^{+}10b]. corepressor [ZFP^{+}13]. corneal [ZLH^{+}14]. corners [Les10w]. corpse [LZY^{+}12]. corralling [JTN^{+}13]. correct [BWS^{+}10]. correcting [OL12]. Correction [DCO^{+}16, FLN^{+}16, RJM^{+}12]. Correlated [KSW^{+}11]. Cortactin [MLM^{+}11]. cortex [GS11, HKN^{+}10, KEJ13, KCF^{+}14, NSS13, RDC^{+}11, SMS^{+}13, TSL12, vGLWB12]. Cortical [KBG12, BKP11, BXB^{+}12, CMS11, JTN^{+}13, LOR^{+}10, PJS^{+}11, RBB^{+}14, SCL11, TPM^{+}13, VJK^{+}10a, VJK^{+}10b, WLW11]. Cossart [Sed11m]. cotranscriptional [SBR^{+}11]. cotranslational [ADS^{+}13, LAB14, SAoS14]. counted [Sho12b]. counteract [CSS^{+}12, MJJ^{+}10]. counteracting [WHL^{+}12]. counteracts [PSF^{+}11]. counterbalancing [BSO^{+}14]. counterflux [SHB^{+}10]. couple [MCS^{+}13]. coupled [BGC^{+}14, CZGG12, KYP^{+}14, LJW13, TLL^{+}13, WKN^{+}13, JEF^{+}11]. couples [CTW^{+}10, SNSyN13, SEV^{+}14, TTB^{+}13]. Coupling [HK14, CS13, ELH14, GSM^{+}12, HKW^{+}13, TUG^{+}10, WSUT11]. cover [Les11v, Sho11-50]. COX [LMS^{+}10b]. COX-2 [LMS^{+}10b]. COX1 [MVP^{+}10]. Cox14 [MVP^{+}10]. CP110 [FRS^{+}13, KKL^{+}14]. CP110-interacting [KKL^{+}14]. CPAP [CAB^{+}10, LWL^{+}13]. CPC [NCT^{+}11]. cPKC [EL14]. Craig [Sed10c]. cranial [RSB13, SZ12a]. craniofacial [JB12]. Crawling [Les13i, RSD^{+}12, BSO^{+}14]. Crb3 [HKI^{+}13]. create [WBL11]. created [Les10-28, Les14x]. creates [CGCP^{+}14, SKN^{+}13]. Creating [Les13j]. Creb3-like [FHA10]. CrebA [FHA10]. CrebA/Creb3 [FHA10]. CrebA/Creb3-like [FHA10]. credentials [Mit12b]. crest [BMRM13, LTJN^{+}12, Les13-37, MMdCOM^{+}11, RSB13, Sho13x, SMB12, VLG14]. cristae [PvdLA^{+}14]. Critical [GdAJ^{+}12, BHMBS^{+}11, CDK^{+}10, FBAO^{+}13, HMBC10, HSY^{+}14, HAKK11, KBG12, MBK^{+}10, WOG13, ZQA^{+}14]. CRL4 [JEF^{+}11]. crop [Sho14v]. Cross [SLS^{+}10, KKY^{+}14, Les11-29, RKG^{+}10, TPSS12]. crossing [Les13n]. crossover [HSY^{+}14, Sho14l]. crowd [Les13c, Les13-42]. crucial [BNL^{+}10, JDS^{+}10]. crucibles [LKSG13]. Crumbs [CHL12, PSK11, Les12k]. Cryo [FBAO^{+}13, LRB13]. Cryo-electron [LRB13]. Cryoelectron [PBM^{+}11, FBZM^{+}10, OBM^{+}10, YON^{+}12]. cryptic [WDG^{+}13]. Crystal [NRM^{+}12, TKB^{+}14, Les12w]. Cse4 [LBS11]. Csi1 [HZW^{+}12]. cTAGE5 [SYS^{+}14]. CTID [JAM^{+}13]. CtIP [BNDB^{+}14, KFH^{+}12, PLC^{+}11, RSS^{+}13]. CtIP-dependent [KFH^{+}12, PLC^{+}11]. cues [CAK^{+}14, PCCR11]. CUL4B [Sho13l, ZMW^{+}13]. Cullen [Sed14t]. Cullin [HC10]. Cullin-3 [HC10]. Cullin3 [LLcK^{+}11]. Culling [Pal10]. Cultivating [Sed12n]. cultured [KSLF^{+}11]. CUPS [CGCP^{+}14, Sho11p]. curbs [FUK^{+}14]. curvature [BR14, CLM^{+}10, PMB^{+}11, WZHV11, XBC^{+}13]. curve [Sho11v]. cut [Les14m, Sho10-50, Sho12-58]. cutaneous [YMM^{+}10]. cutbacks [Sho11q]. cuts [Les10-34, Les12s, Les13m, Les14-35, Sho11-43, Sho12-39, Sho14a, Sho14-43]. Cx43 [PDKG14]. cyclase [PBPW^{+}14a, PBPW^{+}14b]. cycle [ALSN^{+}11, BKE10, GdAJ^{+}12, LJW13, LLS^{+}11, LP13, PWP11, PTBT10, RCG^{+}10, RCG^{+}11, SZW^{+}11, Sho10-45, TGG^{+}11, TL12, WBS11, WGN^{+}13, WRF^{+}13]. cycles [BJ12, FDB^{+}13, KFS^{+}14, Les14r]. cyclic [PBD^{+}13]. cyclical [LJPJ11]. Cyclin [KYOY13, Lin10, OMW^{+}14, SWC13, BNL^{+}10, DP10, GP10, KMS10, KSSD11, MJJ^{+}10, MFGB10, YFO12, vZOtR^{+}10, ABP^{+}12, Les13-29, LP13]. Cyclin-dependent [SWC13, BNL^{+}10, KMS10, KSSD11]. Cycling [LBS^{+}13, Sho10c]. cyclosome [NEMH^{+}10]. Cyk3 [OKNP13]. CYK4 [BPB^{+}12]. cylinders [Sho13l]. cyst [QMHM10]. cysteine [CSG14]. cysteine-proximal [CSG14]. cystic [OL12]. cytochrome [GKR^{+}11a, GKR^{+}11b, GRH^{+}12, HHS^{+}14]. cytokine [LMS^{+}10c]. cytokines [LDN^{+}13]. cytokinesis [AYS^{+}13, BT13, BVC^{+}11, BPB^{+}12, CSKW13, DC12, EDF^{+}10, FPAM13, KPE^{+}14, LCLW11, MZP^{+}10, OKNP13, RBB^{+}14, Sho11o, WLZ^{+}14, WVT^{+}13, ZFA^{+}13]. Cytokinetic [ABP^{+}14, CTY^{+}12, CP11, RW10, Sho13s]. cytomatrix [FBZM^{+}10, MSK^{+}13a]. cytoplasm [GP10, Sho10q]. Cytoplasmic [GF11, Les10h, VLG14, COG11, CWFL13, DSB^{+}14, FLVP10, KLZ^{+}12, TLTW10, YHK10]. Cytoskeletal [TMPH^{+}10, AMH11, ABVP11, CB12, HVOF^{+}14, IHG^{+}12, SHS^{+}12, SWF12]. Cytoskeleton [Les12l, WBML11, BFG^{+}13, GPCK12, HW11, HKW^{+}13, HC10, JKS14, LKG^{+}13, LOR^{+}10, MBLD11, RSD^{+}12, SCR12, WG11]. cytosol [LJLJ11]. cytosol/ER [LJLJ11]. cytosolic [PHB^{+}13, ZLJ^{+}13].

D [BCB^{+}14b, BSO^{+}14, HSI^{+}14, KSW^{+}11, LBS^{+}13, MHAK^{+}12, MRPR12, PGCY12, WZHV11, YDB^{+}11, vZOtR^{+}10]. Dam1 [LHW10, LMA^{+}13, TUG^{+}10]. damage [BSR^{+}11a, BKG10, BTL^{+}12, BAS^{+}14, CRJB^{+}11, DPV^{+}12, GP12, GBJ10, GSGL11, HLT12, JPT^{+}11, KZR^{+}12, KHW^{+}10, KFH^{+}12, LPG^{+}10, LLA^{+}12, MFB12, MBZ^{+}10, MKL^{+}13, PHW^{+}13, RZS^{+}14, SKM10a, SWV^{+}10, SDC10, XHB^{+}10, ZGCG^{+}14, ZLFC14]. damage-specific [ZLFC14]. damaged [ASLS14, BDvdK13, MSS^{+}10, Sho10-46]. Dan [Pow14a]. dangerous [Mit12c]. Daniel [Sed13g, Sho10r]. Danuser [Sed12i]. data [LRA^{+}10, dWMR10]. DataViewer [WCM12a]. daughter [BV11, CWFL13, FPAM13, MXS10, Sho13n]. daughters [Les12-37]. David [Sed10d, Sed11e]. Dbl [LCS^{+}10]. Dbl3 [ZME^{+}14]. dBruce [NSS^{+}10]. DCC [WLN^{+}14, HZM^{+}13]. Dcp [DGH^{+}14]. Dcp-1 [DGH^{+}14]. DCP1a [RBM^{+}11]. Dcp2 [YCP10]. DDB2 [LLA^{+}12, PVM^{+}12]. DDR1 [Les14f]. DDX6 [RMG^{+}12]. deacetylates [RNS^{+}14]. Deacetylation [CMD^{+}13]. dead [YWJ^{+}12]. deadenylation [HBI^{+}10]. deal [Sed13e]. deals [Sho10e]. Death [DNB13, AMO^{+}11, CHL12, CPX11, KDIE11, LKLA12, Les11w, Les12g, MFF^{+}13, SRS10, SRBL13, Sho10w, Sho11l]. decades [HHS13a, HHS13b]. decapping [HBI^{+}10]. decay [CMW11, LNS^{+}13, VBB^{+}10, YCP10]. Deciding [Sho10s]. decipher [BCJ13]. decision [Gil10]. decondensation [LLA^{+}12]. Deconstructing [Sho10t]. decoupled [TGB10]. decoy [ZNH^{+}11]. decreases [EAB^{+}14, LCHB13]. deep [Les12x, Sed12j, Sed14r, Sho12y]. defeats [Les11-40]. Defective [DWL^{+}11, SDD^{+}13, MHK^{+}10]. Defects [MMU10a, CZ10, CFB^{+}12, DPZ^{+}14, KHB^{+}11a, PoLC^{+}13, Sho11-47, WPL^{+}11a, WPL^{+}11b]. defense [ATKK11, PW12, Swa13]. deficiencies [ZYH^{+}11]. deficiency [MFB12]. deficient [BHMBS^{+}11, MBZ^{+}10, RKS^{+}10, Sho13b, ZZW^{+}13, CZ10, CFB^{+}12, GWR^{+}10]. defies [Sed13s]. define [FDB^{+}13, GvEM^{+}11, SRZ^{+}11]. defines [APV^{+}12, COB^{+}12, HSKAT11, LC10, LN11, SKFH11, WWB^{+}10]. Defining [BTC^{+}11, Sho14m]. deformation [WtLK^{+}13]. degeneration [GWP^{+}11, SFB^{+}12, WMB12, ZZW^{+}13]. deglycanation [MBR^{+}11]. Degradation [SHC^{+}13, BBW^{+}14, BVR11, CTY^{+}12, FUK^{+}14, FAB^{+}10, GL10, GGSN^{+}13, HWE^{+}12, HSR^{+}10, HZT^{+}12, KLC^{+}10, LZY^{+}12, LHD^{+}14, MLM^{+}13, MFGB10, NSS^{+}10, RKG^{+}12, RFC14, SB14, SHBC12, TNV^{+}13, TCX^{+}10, WM12, WWHH10]. degrade [Sho12s]. degraded [JEF^{+}11]. degrading [BDN^{+}13]. degron [BV11]. dEHBP1 [GYZ^{+}12]. delay [MGS14]. delays [PBG^{+}13]. delegates [Les11-28]. delicate [Sed12b]. deliver [Gol12b]. delivery [Les11-27, Mit12c, Sho13-50, YZM^{+}12a]. delta [GR11, ZEG11, DSK^{+}11, GYZ^{+}12]. Deltex [HSKAT11]. dementia [GKWG^{+}11]. demethylase [MKL^{+}13]. Dendritic [dJSTV12, HH10, LLT^{+}12, LSM^{+}11, MRR^{+}12]. denervation [MBK^{+}10]. denervation-induced [MBK^{+}10]. Denise [Sed11f]. DENN [Sho10-52, iYGL^{+}10]. dense [DMD^{+}12, KHG^{+}13, PMP^{+}11a, PMP^{+}11b, Sho12-48, vdBFS^{+}12]. dense-core [PMP^{+}11a, PMP^{+}11b, Sho12-48, vdBFS^{+}12]. density [PBPW^{+}14a, PBPW^{+}14b]. depend [OPCEM10]. dependent [AvCG^{+}11, BWL^{+}11, BPB^{+}12, BNL^{+}10, BBD^{+}11, BJE^{+}12, BMFC^{+}11, BMFC^{+}13, BDvdK13, CWL^{+}11a, CHL12, CS13, CWB^{+}14, CSTBM^{+}10, CRJB^{+}11, DKM^{+}13, DMH^{+}12, DSW^{+}11, ECC^{+}13, EM11, EMT^{+}14, FRL^{+}13, FCA10, FWJ^{+}11, GBK^{+}14, GLB10, GBL^{+}11, GC13, GKWG^{+}11, GTR^{+}13, GR11, GKA^{+}12, GS11, HHJ^{+}11, HSI^{+}11, HTS11, HSK^{+}10, HYTU^{+}10, HOS^{+}12, KOK^{+}13, KKL^{+}11, KFET11, KSB^{+}13, KSLF^{+}11, KMS10, KFH^{+}12, KSSD11, KTB^{+}14, LNT^{+}10, LSS^{+}12, MdFF^{+}14, MLM^{+}11, MBCKD13, MSR10, MALS10, MLG^{+}10, MHC^{+}12, MP13, MFB12, MRLLS12, MTT^{+}14, NCT^{+}11, NBDB12, OSD^{+}14, PKD^{+}11, PHB^{+}11, PLC^{+}11, PRFF13, PLL^{+}12, PSF^{+}11, RCM^{+}12, RGL^{+}13, RKS^{+}10, RB11, RKW^{+}13, RSRK13, RFK^{+}10, SOW^{+}11, SYV14, SB14, SPF11, SWC13, SHN^{+}11, TIT11, TNV^{+}13, TKS^{+}13, TB12, TC10, VLKI14, WOG13, WBMCSS13, WAJ^{+}12, WHA^{+}13, WRCD12, YHF13, YTM^{+}11, YZPF12, ZSK12]. dependent [ZPS^{+}10, lDSB^{+}10a, lDSB^{+}10b, vBAK^{+}12, OOKH^{+}12]. depends [PDMBW11, RPK^{+}11]. dephosphorylation [GTR^{+}13, RSM^{+}13]. depletion [CTW^{+}10, DPZ^{+}14]. depolarization [MSS^{+}10]. depolymerases [JKA^{+}10]. deposition [BMFC^{+}11, BMFC^{+}13, SMMB11]. depths [BS13, Pow10]. derived [SRZ^{+}11]. dermal [DPW^{+}11, DPW^{+}12]. describes [GCR^{+}12]. deselection [HSY^{+}14]. design [CF13]. desmin [CZGG12]. desmoglein [CNP^{+}12, DKA^{+}13]. desmoglein-1 [DKA^{+}13]. desmoplakin [KLS^{+}13, PDKG14, PDKG14]. Desmosomal [NAS^{+}11, NAS^{+}12, NAS^{+}13]. desmosome [Les11q, SCL11]. desmosomes [NAS^{+}11, NAS^{+}12, NAS^{+}13]. destabilization [JLW^{+}10]. destiny [Sho14-63]. destruction [DP10, Krä13, WMB12]. details [NRM^{+}12]. detection [PBPW^{+}14a, PBPW^{+}14b]. detector [WWS^{+}12]. determinant [BRD^{+}13, JG10, dJSTV12]. determinants [LLM^{+}10, SWS^{+}11]. determination [ZGEM12]. determine [AGM^{+}10, BKAB13, FA12, LHD^{+}14, MFF^{+}13, WRF^{+}13]. determined [SSB^{+}10]. determines [DSL13, DHVK10a, DHVK10b, JAM^{+}13, KHG^{+}13, LSGVM14, MSK^{+}13a, SSD^{+}14, SJRV14, TGES12]. Determining [Sho13m, MSC^{+}10]. detyrosinated [ADAB^{+}12]. Deubiquitination [Sho14n, ZLFC14]. developing [DNB13, LSGVM14, PL10, Sho14-44]. Development [Gol12a, AAE^{+}14, CMH^{+}10, CLD11, CSEH12, EL14, FMG^{+}11, GZR^{+}14a, GZR^{+}14b, HH14a, JB12, KT10, KNsMK13, Les11-28, MI13a, NLP^{+}10, PYT^{+}13, SAG^{+}11, SMSP11, Sed13c, Sed13v, Sho14-54, SL14, VB12, VWD^{+}13, WKN^{+}13, ZPS^{+}10]. developmental [GWP^{+}11, Les12-33]. DExD [CMW11]. DExD/H [CMW11]. DGAT2 [XZC^{+}12]. DGK [Mit12a]. DGK- [Mit12a]. Dhh1 [CMW11]. Diacylglycerol [RCM^{+}12]. Dias [Sho10-39]. dichotomy [OCF^{+}10]. dictate [BLT^{+}11]. Dictyostelium [VKMI12, ZTBK14]. Didier [Sed13h]. Dieckmann [Sed11d]. Diet [HSJ^{+}13]. Different [YFO12, HBG^{+}11, KdKDP12, Les14g, Sho12-30, VFNR11, YHK10]. differential [BKAB13, CLL^{+}10, DCO^{+}13, GWR^{+}10, HIB^{+}10, VWD^{+}13, WHH^{+}11]. differentially [BPT^{+}14, NLAS^{+}10]. differentiated [BKE10]. differentiating [ECC^{+}13]. Differentiation [MPD^{+}12, ALV^{+}12, BWL^{+}13, BTC^{+}11, BBJ^{+}10, BLC^{+}14, CVJ^{+}11, CTL^{+}10, CCM^{+}11, DKA^{+}13, IIN^{+}11, KPH^{+}12, KNsMK13, MRLLS12, Sho12-34, TNH^{+}11, WSZ^{+}12, WAG^{+}10, ZME^{+}14]. Differentiation-associated [MPD^{+}12]. differently [Sed11q]. differing [KLvdB^{+}13]. diffuse [GHGH11]. diffuses [RZS^{+}14]. diffusion [BMLB^{+}12a, BMLB^{+}12b, BLI^{+}10, BKS^{+}13, WBL11]. diffusion-based [WBL11]. digs [Les11q]. dilation [KEJ13]. Dillin [Sed10a]. dimensional [LNTR14]. dimer [KSP^{+}11]. dimeric [BPDB^{+}11]. dimers [DGS^{+}10]. dine [Les10-42]. Direct [KSSD11, SMS^{+}14, TP13, FHA10, LN14, MBLD11, SMdP^{+}14, SDN^{+}14a, SDN^{+}14b, Sho13-40, dMSMZ14]. directed [CDAK10a, CDAK10b, FLN^{+}10, FLN^{+}16, FHKW11, PAB^{+}10, LMT^{+}10]. Directing [Sho12l, UKZ^{+}13]. direction [Sho10-38]. directional [CAK^{+}14, GdBP^{+}14, HCG^{+}11, LBWS10, QJO10, SRU^{+}12]. directionality [mFH13, WAJ^{+}12]. directions [Sho14-68]. Directly [GMD^{+}10, BVR11, GRHA^{+}12, JOR^{+}11, JC10, LHL11, LCS^{+}10, SMT^{+}10]. directs [BWS^{+}10, CM12c, DPB^{+}10, HLH^{+}14, MGT^{+}10, PHB^{+}11, SBTF13, Sho14c, Sho14-45, YZL^{+}13]. DIS [NGM12]. disaggregation [WTBM12]. DisAp [GBK^{+}14, Sho14-66]. DisAp-dependent [GBK^{+}14]. DisAp-pearance [Sho14-66]. disassembling [CSTBM^{+}10]. disassembly [BFG^{+}13, IHM13, KYOY13, MWG^{+}12, WDB10]. discharge [Sho10b]. Discoidin [JDL^{+}14]. Disconnecting [HCG^{+}11]. Discovering [LK12]. discovery [Ste12]. Discrete [YHK10]. discrimination [Les13-28]. Disease [LNT^{+}10, PDKG14, CC10a, CPX11, DD10a, DSP11, ER10, HH14a, HPB^{+}12, JCL^{+}11, KKUG11, LCL12, PL11, RvD13, SMSP11, SFB^{+}12, Sho11a, Sho12-55, WK12, WTH^{+}11, WCC^{+}10, NGM12]. Disease-causing [LNT^{+}10]. disease-linked [LCL12]. Dishevelled [MGG^{+}12]. disjunction [BM11]. disorder [HWS14, Les12-33]. disordered [TL12]. disorders [PBvdS12]. dispensable [IHM13, MTM^{+}10, TTB^{+}13]. displacement [STI^{+}11]. display [SW10a, SSH^{+}13]. disposal [BGC^{+}10]. dispose [Sho12d]. disrupt [KFL^{+}14]. disruption [BTL^{+}12, COG11, NB12, RCG^{+}10, RCG^{+}11]. disrupts [CAK^{+}14, LZLG13, SMT^{+}10]. Dissecting [STD^{+}10]. dissection [FBR^{+}10, RTM13]. dissemination [FPM^{+}14, SPJ^{+}14, Sho14-60]. dissociation [CLEZ12, CLZ^{+}14, PHB^{+}13, TLS10]. distal [ASLS14]. distally [MWH12]. distance [Les11c, WPM14]. Distinct [DSK^{+}11, EDF^{+}10, GvEM^{+}11, HCP^{+}13, LAO^{+}10, MAD^{+}11, OKNP13, TO12, ABVP11, ABP^{+}14, BPDB^{+}11, COW13, FSA^{+}11, wFLW^{+}13, GCR^{+}13, GDS^{+}12, HGV^{+}14, JDB^{+}12, JPT^{+}11, KKMB10, KKL^{+}14, LDCF^{+}13, LLH13, LSE^{+}10, MMB^{+}11, NAS^{+}11, NAS^{+}12, NAS^{+}13, OHC10, OLB13, OLT11, PGCY12, TTC^{+}14, TYN^{+}13, TMS^{+}12, VFNR11, WDB10, ZKR^{+}11]. distribution [CHS^{+}10, CSS^{+}14, CSHS^{+}13a, CSHS^{+}13b, DWPC^{+}11, HARS14, LN11, LWBH12, SHN^{+}11]. Disturbed [TAGJ11, HLN^{+}11]. disulfide [KLvdB^{+}13, ZCB^{+}10a, ZCB^{+}10b]. Disuse [Sho10u]. divergent [MOZ^{+}13]. diverse [EJBW12, KXN10, ZGW^{+}14]. diversification [KNH^{+}10]. diversity [KLHS14, OT11]. divide [Sho11b]. dividing [OB12]. Diving [Sed12a]. division [AIJI11, BDC^{+}14, CPS^{+}13, JG10, KZR^{+}12, LDCF^{+}13, LSGVM14, Les13-27, Les14g, LNTR14, NSZ^{+}13, PJS^{+}11, RG14, RKE14a, RKE14b]. divisions [GYZ^{+}12, KST^{+}11]. DJ [DGS^{+}10]. DJ-1 [DGS^{+}10]. DLC1 [Sho14h, TQM^{+}14]. Dlg1 [MJEM10, SMdP^{+}14]. DLK [GWP^{+}11, HRWW^{+}13, Sho11q, XWE^{+}10]. DM1 [RCBY^{+}12]. DNA [BSR^{+}11a, BKG10, BNDB^{+}14, BDC^{+}14, BTL^{+}12, BBW^{+}13, BCJ13, BAS^{+}14, CRJB^{+}11, CCJ^{+}12, CGRS^{+}12, DPV^{+}12, DPB^{+}10, ECC^{+}13, ETI^{+}10, EIE^{+}14, GP12, GHK10a, GZZ^{+}14, GSP^{+}14, GBJ10, GSGL11, GL10, IAMH10, JRC^{+}13a, JRC^{+}13b, JPT^{+}11, KPC^{+}10, KHW^{+}10, KK13b, KLP^{+}14b, KNH^{+}10, KFH^{+}12, KSSD11, LPG^{+}10, LeB10, yLFAM13, Les10s, Les10z, Les10-38, Les10-39, Les14i, Les14h, LvBG^{+}10, LLA^{+}12, MGS14, MFB12, MHKM11, MHV12, MBZ^{+}10, MKL^{+}13, MFR^{+}14, NSS^{+}10, PL11, PLL^{+}12, PHW^{+}13, RSS^{+}13, RZF^{+}11, RZS^{+}14, RFK^{+}10, SKM10a, SCL^{+}14, Sho12-47, Sho14y, SJM^{+}13, SWV^{+}10, SDC10, TIM14, VEDBC13, WGC11, XHB^{+}10, XSJ^{+}10, YTM^{+}11, ZGCG^{+}14, ZLFC14, ZYH^{+}11, ZNA^{+}14]. Dnm1 [BKBS12, Sho12m]. do [DBUT13, Sed11q, Sho14e]. dock [Sho14i, TCN14]. docking [AMGC14, BLC^{+}14, HHJ^{+}11, SKN^{+}12, TQS^{+}11, WLGC11, WGM^{+}12]. does [Bab14, KRS11]. doesn’t [Les10j, Sho10-42, Sho13g]. doing [BH13, LRB13]. domain [ABD14, BBY^{+}12, DSM^{+}11, FSLM11, FSOL14, HDK^{+}13, HTT^{+}11a, HKW^{+}13, IIWS14, JLVH12, JAM^{+}13, JDL^{+}14, KYHG12, KDIE11, KLC^{+}10, LGAC13, LSW^{+}14, MSZ^{+}12, MLSM^{+}11, NvCL^{+}13, OL12, PGAE^{+}13, SLM^{+}11, Sho14-56, TSL12, iYGL^{+}10, ZBJL^{+}10, ZSZ^{+}13]. domains [HBC^{+}11, HWB^{+}13, TYN^{+}13, TP13, WZHV11, vGCMA^{+}14]. Dominique [Pow14b]. don’t [Les10u, Les12-37, Les13s, Sho14x, Les14-36]. dopamine [CPX11]. Dorma [MZP^{+}10]. dormant [GB10]. dorsal [GNHB11, HTT^{+}11a, LN11]. dosage [KK11]. dots [Sho11o]. Double [Les14i, AMO^{+}11, BNDB^{+}14, BCJ13, CCJ^{+}12, CWG^{+}11, GBJ10, IAMH10, JRC^{+}13a, JRC^{+}13b, KK13b, Les10b, PLC^{+}11, PLL^{+}12, Sho10o, Sho11-41, YTM^{+}11, YSN^{+}10]. Double-checking [Les14i]. double-edged [AMO^{+}11]. double-membrane [YSN^{+}10]. double-strand [BNDB^{+}14, BCJ13, CCJ^{+}12, CWG^{+}11, GBJ10, IAMH10, JRC^{+}13a, JRC^{+}13b, KK13b, PLC^{+}11, PLL^{+}12, Sho11-41, YTM^{+}11]. doubles [Les10-30, Sho11c]. Doubling [Les10i]. Doug [Sed10e]. down [HVW^{+}10, LCBG^{+}11, Les10-30, Les10-34, Les11n, Les14c, LT11, Sho11-43, Sho12c, Sho13a, Sho13e, Sho14a, Sho14-29]. down-regulates [LCBG^{+}11, LT11]. down-regulation [HVW^{+}10]. downhill [Les13-30]. downstream [DJL^{+}12, HTM^{+}14]. Doxsey [Sed13u]. Dpb11 [GSP^{+}14]. drag [Les11-43]. drags [Sho11-46]. Drawing [Sho10p]. draws [Les13c]. Drebrin [Les13k, WDG^{+}13]. Drg1 [KLZ^{+}12]. dRich [Les10j]. drive [ADS^{+}13, BLO^{+}12, BMÁG^{+}14, HAB14, LAO^{+}10, LDN^{+}13, RDC^{+}11, WHH^{+}11, YRU^{+}13]. driven [CZC^{+}11, DGF^{+}14, ETRP12, HZM^{+}13, JGB^{+}13, SLM^{+}11, SST^{+}12, WDB10, YWC^{+}13, ZTBK14]. driver [DWJ^{+}14, HSF12]. drives [AVP^{+}14, HRK13, HTM^{+}14, LNJ^{+}13, MVC^{+}11, MLSM^{+}11, OVW10, RJM^{+}12, SAoS14, Sho10-53, Sho10-61, Sho12-42, TKMK10, TGG^{+}11, ZME^{+}14]. driving [YZPF12]. drop [Sho12-29]. droplet [GSW^{+}11, KHFV^{+}13, PGP14, SWS^{+}13, XZC^{+}12]. droplets [AHL^{+}11, Les14p, Sho11-29, Sho12n]. DROSHA [KA12]. Drosophila [ETC^{+}12, GOWM12, HZS^{+}10, HC10, JPT^{+}11, SRS10, SNZVK12, SNZVK13, SDB^{+}10, VJK^{+}10a, VJK^{+}10b, WBcY^{+}11, ABVP11, ATKK11, AIJI11, BMG14, CG12a, CDB^{+}14, CRL^{+}14, CTM^{+}14b, DV10, DGH^{+}14, DPZ^{+}14, DSW^{+}11, FRS^{+}13, GCP^{+}14, GDO13, HFS10, HSJ^{+}13, HSI^{+}11, JG10, KUN^{+}13, KT10, LN11, Les14z, LCHB13, LMW^{+}11, LMT^{+}12, LgLM^{+}10, LKG^{+}13, MRLLS12, MZP^{+}10, MHCvSW11, NSS^{+}10, NSBW10, OBC14, OFS^{+}10, PHB^{+}11, RNS^{+}14, RLS^{+}14, SYK^{+}11, SMM^{+}10, TNV^{+}13, TGES12, VLI^{+}14, WWHH10, YTT^{+}10]. DRP1 [MRLLS12, OWC^{+}10, Sho13-48]. DRP1-dependent [MRLLS12]. Drs2 [Les13l]. drug [LK12, VŠH^{+}11]. drugs [Gol12a, RBS10]. Dual [CMS10, NMB^{+}14, TLTW10, XRO^{+}11, HBS^{+}10, KPI^{+}10, Sho10v, WWHH10]. Dual-mode [NMB^{+}14, XRO^{+}11]. Duchenne [DWJ^{+}14]. ductal [LDN^{+}13]. due [KHB^{+}11a]. duel [Les14z]. duplication [HKH^{+}10, SDB^{+}10, WSUT11]. duration [RHK11]. during [AVP^{+}14, ATU^{+}12, ABP^{+}14, ATKK11, AYS^{+}13, BLO^{+}12, BNDB^{+}14, BVC^{+}11, BPH^{+}14, BPB^{+}12, BDC^{+}14, BEJ10, BM11, BRF^{+}10, BC11a, BC11b, BTC^{+}11, BGS13a, BGS13b, BJ12, BLC^{+}14, BHA^{+}12, CZM^{+}14, CSKW13, CFB^{+}13, CSEH12, CGCP^{+}14, DGF^{+}14, DKMK^{+}11, DJL^{+}12, EL14, FMG^{+}11, FRS^{+}13, FMI^{+}13, GSJS10, GYZ^{+}12, GPCK12, GBJ10, GCH^{+}14, GNHB11, HKN^{+}14, HBSD12, ISZ^{+}11, KdKDP12, KNsMK13, KFS^{+}14, KWTR10, KSR^{+}13b, LN11, LCLW11, LCS^{+}13, LGM^{+}12, Les11-28, Les11-47, Les12j, Les13-30, Les14-32, LJLJ11, LKG^{+}13, MGT^{+}10, MWH12, MGS14, MGFG^{+}10, MMU^{+}10b, MRLLS12, MAE^{+}10, MSZ^{+}12, NEMH^{+}10, NOT^{+}14, NSS^{+}10, NWD^{+}11, NLAS^{+}10, OCF^{+}10, OKNP13, OYH13, OWC^{+}10, OB12, PSVRB^{+}11, PLC^{+}11, PTST12, QTL^{+}12, QTL^{+}13, QMHM10, QJO10, RLS^{+}14, RFVE^{+}10, RFAA^{+}12, RSB13, RMF^{+}10, RC12, RDC^{+}11, SBEM13, SKM10a, SSdA^{+}14, SKN^{+}12, SZJ^{+}10, SHS^{+}12]. during [dMSMZ14, SML^{+}13, TSL12, TCN14, TCB^{+}14, UG10, UTK^{+}13, VB12, VTM14, VSG^{+}12, VLI^{+}14, WBS^{+}12, WWM^{+}12, WRF^{+}13, WLZ^{+}14, WMCF10, WLK^{+}11, WMV^{+}14, WVT^{+}13, XSJ^{+}10, YKW^{+}12, YTM^{+}11, YSN^{+}10, YSM10, ZBJL^{+}10, ZPS^{+}10, ZZW^{+}10, ZKW^{+}13, ZEG11]. duty [Sho13d]. dx [DSW^{+}11]. dynactin [DSD^{+}13, MSZ^{+}12, Sed13w, ZYF^{+}11, LBD^{+}14]. Dynamic [BGY^{+}13, CSHS^{+}13a, CSHS^{+}13b, FHY^{+}10, MRR^{+}12, SBP^{+}10b, VTM14, DBH^{+}11, DCN^{+}10, HSI^{+}11, KSP^{+}11, LMT^{+}12, LWW12, MAE^{+}10, OZT^{+}13, PHB^{+}11, SGC10, TMG^{+}10]. Dynamics [BG11a, AMS^{+}13, BGB^{+}13, BMÁG^{+}14, BKY^{+}10, BHB^{+}11, BR14, CMW11, CPS^{+}13, CB12, DGH^{+}14, DWPC^{+}11, DYS^{+}14, FBR^{+}10, FWM^{+}10b, GCH^{+}14, HSN^{+}11, HH10, KCF^{+}14, LGAC13, LLH13, LMW^{+}11, LDL12, LWB^{+}14, LhYL^{+}13, MAD^{+}11, NCT^{+}11, NPL^{+}10, Pri14, RMS^{+}14, RMT13, RPM^{+}13, SSdA^{+}14, SME^{+}13, Sho11v, Sho14t, SQC^{+}12, SW12, SWF12, TOI^{+}13, VLI^{+}14, WBS^{+}12, WBL11, WBML11, WHWS12, WHDR^{+}10, ZKR^{+}11]. Dynamin [KKS^{+}14, Les13m, SCN^{+}14, Sho14o, AKC^{+}12, BKBS12, CLW^{+}14, KBS^{+}10, LMS^{+}13, RPO^{+}14, SWC13, SWS^{+}13]. dynamin-like [AKC^{+}12]. dynamin-related [CLW^{+}14, SWC13]. dynamin2 [GCH^{+}14, WHF^{+}11]. Dynamin2- [WHF^{+}11]. dynamins [Les10y]. Dynein [DHL^{+}12, JVS^{+}14, WRCD12, BSR^{+}14, BYY^{+}12, DCN^{+}10, DSD^{+}13, DSB^{+}14, ETRP12, EM11, FS10, mFH13, GS11, Kik13, Kin13, KBG12, MJEM10, OYYK14, RTM13, SYV14, Sho10-47, TSL12, YSaY^{+}13, ZYF^{+}11, ZQA^{+}14, ŻKC^{+}11, Les12m, Sed13s]. dynein-based [Kin13]. Dynein-dependent [WRCD12, EM11, GS11]. dynein-driven [ETRP12]. dyneins [YHK10]. Dysbindin [RBP^{+}13]. dysfunction [CPX11, CFB^{+}12, MFB12, PBvdS12]. Dyskerin [SSK^{+}14]. dystonin [RBF^{+}12]. dystonin-a2 [RBF^{+}12]. dystrobrevin [SAG^{+}11]. Dystroglycan [NSS13]. Dystroglycan-mediated [NSS13]. dystrophic [ERS10, GSB^{+}13]. dystrophin [RGL^{+}13, ZC11]. dystrophy [APV^{+}12, CG10a, DWJ^{+}14, RK13].

E-cadherin [Les13n, NDS^{+}11, SPJ^{+}14, SZE^{+}11, SMT^{+}10, YYA^{+}11, lDSB^{+}10a, lDSB^{+}10b]. E-catenin [BKY^{+}10]. E2 [vRJMvD10]. E2F [BKE10, MPD^{+}12]. E2F-induced [BKE10]. E3 [BAY^{+}11, DSW^{+}11, HLH^{+}14, KLF^{+}14, LNJ^{+}13, ONNB^{+}14]. earliest [BTC^{+}11]. Early [HARS14, YFLH12, BDC^{+}14, BSR^{+}14, CSEH12, FWM^{+}10a, FCE^{+}12, HHL^{+}11, KHW^{+}10, KSR^{+}13b, LSS^{+}12, MGR^{+}10, MBM^{+}10, Nan14, SKN^{+}12, Sho14-36, SMZL13, TH11, YKW^{+}12, YWJ^{+}12, ZYF^{+}11, ZQA^{+}14, Sed13d]. eat [Sho10s]. eaters [Les11-46]. Eaton [Sed13v]. EB1 [BKS14, EAK13, JK10, LMW^{+}11, LMT^{+}12, PDKG14]. EB3 [FPAM13, Sho13n]. EBP50 [GLB10, GB12]. Echard [Sed13b]. Echinoid [LN11]. ECM [CFLDM11, Les11-39, MAD^{+}11, Sho11-28, Sho12s, WtLK^{+}13]. ectodomain [NRM^{+}12]. EDEM2 [NOS^{+}14]. edge [FEHF12, HCC^{+}10, HKR^{+}10, LN11, Les13-34, Mit12a, WS10]. edged [AMO^{+}11]. editing [Les11z]. editorial [Mis10]. effect [KNOM11, WHA^{+}13]. effector [BPB^{+}12, BW12, DGH^{+}14, GVP^{+}11, PAB^{+}10, TSH^{+}14]. effects [DCO^{+}13, MRPR12, NGL^{+}12]. efficiencies [FA12, PASG^{+}12]. efficiency [vdBFS^{+}12]. Efficient [BPL^{+}11, DK10a, FHY^{+}10, GSGL11, KNH^{+}10, LvBG^{+}10]. efficiently [BWS^{+}10]. efflux [VŠH^{+}11]. Eg5 [Les11-43, MTG^{+}11, WVvG^{+}13]. EGF [DYI^{+}13, ONH^{+}12, Sho10v]. EGF-induced [ONH^{+}12]. EGFR [FAB^{+}10, Les10-30]. egg [HH14b, Sed14p, Sho13-35, Sho13-45, WRF^{+}13]. Eggert [Sed11s]. eggshell [OGD^{+}12]. Egr1 [JBS^{+}12, JBS^{+}13]. Ei24 [LSW^{+}14]. eIF4A1 [WKGB^{+}10]. eIF4B [EMT^{+}14]. eight [ZZS13]. Eisosome [KHS^{+}11, MSS^{+}12]. eisosomes [Les11m]. ejection [CYLMM13, CSHS^{+}13a, CSHS^{+}13b]. elastase [PMHZ10]. electron [AiIK^{+}13, FAvdB^{+}12, FBAO^{+}13, KSW^{+}11, LRB13]. elegans [AGL^{+}14, BDC^{+}14, CHK^{+}10a, CHK^{+}10b, CTY^{+}12, EM11, FTJG13, GZR^{+}14a, GZR^{+}14b, GHGH11, HYS11, KHW^{+}10, KHG^{+}13, KWTR10, MMVK^{+}12, OGD^{+}12, RKW^{+}13, SQC^{+}12, UHKS11, WOG13, WRCD12, ZBJL^{+}10, ZC11]. element [LHGT^{+}12, SCL^{+}14, WS10]. elements [BPH^{+}14, BKAB13, OZT^{+}13, WOG13]. Elevated [CYLMM13]. Eliminating [Mis13, Oka14]. elimination [KSLF^{+}11, MHK^{+}10]. Elizabeth [Sed14f]. Elm1 [CKO^{+}10, Les10k, MCHCC10]. Elmo [TCN14]. elongation [GBK^{+}14, GZLG11, KMSR12, KWTR10, LWL^{+}13, MTM^{+}10, RHK11, SGLV10, WBMCSS13]. elucidates [RKT^{+}14]. Ema [KWDD10]. embrace [Sed14g]. embryo [BDC^{+}14, Les11k, RLS^{+}14, TGES12]. embryonic [BTC^{+}11, DPL^{+}12, ECC^{+}13, FP10, KWK^{+}11, Les12b, Les13n, LZW^{+}10, LZR^{+}11, OLT11, QB12, YWJ^{+}12, ZPS^{+}10, ZKW^{+}13]. embryonic-to-adult [KWK^{+}11]. embryos [ERS10, GHGH11, KHW^{+}10, KWTR10, Nan14, RZA^{+}13, Sed13d, SQC^{+}12]. Eme1 [DKMK^{+}11]. emerge [Sho10-52]. emergence [OLT11]. emerging [HWS14, RG11]. emigration [MMdCOM^{+}11]. EMILIN1 [DPW^{+}11, DPW^{+}12]. EMILIN1- [DPW^{+}11, DPW^{+}12]. EMT [LNL11, RSB13, VBB^{+}10]. enable [EAB^{+}14]. enables [BKS14, OLB13]. encoded [GSM^{+}14]. end [BHB^{+}11, CFB^{+}13, CWG^{+}11, DYP14, GLM^{+}10, HM10, LHW10, Les13z, ME13, PAB^{+}10, PH10, Sed12z, Sed14o, Sho13-59, Sho14-39, TB13, YTM^{+}11, YWC^{+}13, vdVMG^{+}11, BV11]. end-on [YWC^{+}13]. endo [CFB^{+}12]. endo-lysosomal [CFB^{+}12]. endocytic [HKR^{+}10, NPL^{+}10, NCML^{+}12, PDMBW11, SBTF13, Sho14z]. Endocytosis [Les10l, Les13o, Sho14p, BVL^{+}12, BG11b, GK13, GHK^{+}10b, GWR12, GCH^{+}14, GTS10, oHXK^{+}12, JTN^{+}13, KWH14, Les12-30, MLG^{+}10, MLY^{+}10, Pow10, QJO10, Sed10o, SRCP14, SCN^{+}14, Sho12-44, Sho14o, TID^{+}10]. endocytosis-based [JTN^{+}13]. endoderm [SRZ^{+}11]. endoderm-derived [SRZ^{+}11]. endodermal [WHWS12]. Endogenous [GWR^{+}10, TKL^{+}10, NLAS^{+}10]. endoglin [RB11]. endoglin-dependent [RB11]. endolysosomal [HWE^{+}12, VJK^{+}10a, VJK^{+}10b, WWHH10, YBN^{+}11]. endolysosomes [HSR^{+}10]. endonuclease [DKMK^{+}11]. endonucleolytic [SML^{+}13]. endopeptidase [HMO^{+}14]. endoplasmic [AiIK^{+}13, CTH^{+}11, DPZ^{+}14, LSOT10, MMS^{+}10, MDW^{+}13, RPK^{+}11, WLW11]. Endorepellin [BFG^{+}13]. Endosomal [MRCC^{+}13, CLW^{+}14, CTM^{+}14b, DKF^{+}11, ECJB10, KWDD10, LCL12, MBCKD13, PLR^{+}13, RBA^{+}11, SSdA^{+}14]. endosome [ALF^{+}13, HARS14, KKUG11, ZYF^{+}11, ZQA^{+}14, LHL11]. endosome-Golgi [KKUG11]. endosomes [BSR^{+}14, FMI^{+}13, JDHS10, LMT^{+}10, Les11-45, Les12-27, LLR^{+}12, MMU^{+}10b, Pal14, Sho14-30, Sho14-36]. endothelial [AFM^{+}13, BFG^{+}13, BDB^{+}14, CVR10, CLL^{+}10, FEHF12, FRL^{+}13, HLN^{+}11, HOS^{+}12, JCN^{+}14, KFL^{+}14, Les13-31, LXTM12, MAD^{+}11, Sho12y, SKV^{+}11, TAGJ11, WHF^{+}11, ZPS^{+}10, ZLH^{+}14]. ends [DLBG11, EHUD14, HKN^{+}10, JK10, MOZ^{+}13]. enduring [RN12]. energetic [MSK^{+}13b]. energetics [bCAH^{+}11]. Energy [SWS^{+}11, CTW^{+}10]. engagement [RKG^{+}12, SF12, Sho14m]. engages [SCL^{+}14]. engineered [RPM^{+}13, SBP^{+}10b]. engraftment [BvMD^{+}14]. engulfing [LZY^{+}12]. enhance [ALSN^{+}11, CT10, ONH^{+}12]. enhanced [GSB^{+}13]. enhances [GKA^{+}12, KWH14, TUG^{+}10, VPC^{+}14, XBC^{+}13]. enhancing [JCN^{+}14]. enigma [RN12]. eNOS [SKV^{+}11]. enough [Sho12f]. enriched [WMC14]. Ensconsin [GCP^{+}14]. Ensconsin/ [GCP^{+}14]. ensure [BB10, DK10a, IP12, LR13, Les14p, LHS10, NCT^{+}11]. ensures [GYC^{+}14, KZR^{+}12, LvBG^{+}10]. Ensuring [Les13p]. entanglements [KPC^{+}10]. enters [Les10q]. entosis [Sed11l]. entrances [TLL^{+}13]. entry [ANT^{+}12, BC11a, BC11b, BKS^{+}13, FSLM11, GM11, LLS^{+}11, RCG^{+}10, RCG^{+}11, RY11, SPC^{+}13, SP11]. envelope [BMLB^{+}12a, BMLB^{+}12b, BBD^{+}11, CTM^{+}14a, CSTBM^{+}10, FS10, GSU^{+}12, HH14a, HZW^{+}12, LLK11, Sho11-39, Sho14q, Sho14-59, TGG^{+}11, TCB^{+}14, ZGW^{+}14, ZKR^{+}11]. environment [UHKS11]. environments [LNTR14]. enzymatic [vGCMA^{+}14]. Enzyme [Les14j, Les10m, Les14o, Sho12n, vRJMvD10]. enzymes [GCSB10, Les10f, Les10g, Les10h, XG12]. Epac1 [GVP^{+}11, Sho11r]. EpCAM [MVR^{+}10]. ependymal [GCR^{+}13]. EPG [LYH^{+}13]. EPG-7 [LYH^{+}13]. Epg5 [ZZW^{+}13]. Eph [JGA^{+}11, NJS^{+}10]. EphA2 [HYTU^{+}10, Sho13o, SGT^{+}13]. EphB2 [SSD^{+}14]. Ephexin4 [HYTU^{+}10]. Epiblast [NSS13, Sho13p]. epidermal [ALV^{+}12, GHK^{+}10b, LN11, NSZ^{+}13, SFL12, TLTW10, YMM^{+}10]. epidermis [FMG^{+}11, PL10, RFL13, SCL11]. Epigenetic [BWS^{+}10, STI^{+}11, BKAB13, SLS^{+}10]. epigenetically [SMB12]. Epithelia [Sho13q, BHB^{+}11, KBKW10, Sed11e, Sed12b, Sho14-31, Sho14-42, WBML11, Sed14g]. Epithelial [EZT^{+}12, GFSR11, Sho10w, AEC^{+}14, BMRM13, BG11a, CPS^{+}13, CHL12, CFLDM11, DHB^{+}14, FSLM11, GHC^{+}14, GBSC^{+}12, GWR12, HKI^{+}13, HKN^{+}10, JK10, KF11, KKL^{+}11, KCF^{+}14, LDCF^{+}13, Les13b, LCHB13, LZR^{+}11, MBR^{+}11, MLH12, MSC^{+}10, MSK^{+}13b, NT11, PTST12, QTL^{+}12, QTL^{+}13, QMHM10, RG14, RFVE^{+}10, RFAA^{+}12, SPJ^{+}14, Sho14-53, SPF11, SMB12, SMT^{+}10, TIT11, VB12, VTM14, VOSB12, WHF^{+}11, WJW^{+}11, ZZW^{+}10]. epithelial-restricted [MBR^{+}11]. epithelial-to-mesenchymal [AEC^{+}14, SMB12]. epithelium [GvEM^{+}11]. EPLIN [TIT11]. EPLIN-dependent [TIT11]. Epstein [PASG^{+}12]. equal [Les10-28, Les14x]. equilibrium [CPS^{+}13, KSP^{+}11]. ER-associated [RFC14]. ERAD [BGC^{+}10, KXN10, NOS^{+}14]. ErbB [SAG^{+}11]. erects [Sho14z]. Erik [Sed11g]. ERK1 [RCG^{+}10, RCG^{+}11]. ERK1/ [RCG^{+}10, RCG^{+}11]. ERK8 [GL10]. ERK’d [Sho10-54]. ERKs [NB10]. Erlins [HVDG13]. ERM [SMS^{+}13]. ERMs [MdFF^{+}14]. ERO1 [ZCB^{+}10a, ZCB^{+}10b, AiIK^{+}13, KSSK12]. ERO1- [ZCB^{+}10a, ZCB^{+}10b, AiIK^{+}13]. error [CLO^{+}11, RJM^{+}12]. error-prone [CLO^{+}11]. errors [BWS^{+}10]. escape [Les14f]. escapes [DP10]. Escherichia [WHF^{+}11]. escort [Les11h]. ESCRT [AVP^{+}14, ALF^{+}13, DCL^{+}12, HSKAT11, LCL12, Les11h, Les14k, SSZ^{+}14, SP11, WAW^{+}11]. ESCRT-III [DCL^{+}12, Les11h, Les14k]. ESCRT-III/MVB [DCL^{+}12]. ESCRTs [FSR11]. esophageal [RSRK13]. essential [BKT13, BKBS12, CLSO^{+}12, CHL^{+}14, CCGN11, CSM^{+}12, CGRS^{+}12, FHKW11, HHJ^{+}11, HMO^{+}14, HKR^{+}14, KBKW10, KSSD11, LHGT^{+}12, MXS10, MGG^{+}12, MHKM11, MVP^{+}10, MHK^{+}10, MMC^{+}10, MAD10, NNSH11, OWC^{+}10, RSS^{+}13, RKE14a, RKE14b, SLM^{+}11, SCL11, VSMC11, VLG14, WSUT11, WGR^{+}12, XW10, YSaY^{+}13, ZSK12, vGLWB12]. establish [KLP14a, KRS11, WLK^{+}11]. establishes [BAS^{+}14, KSSK12]. establishment [BG11a, BLI^{+}10, JTN^{+}13, KST^{+}10, MLH12]. eukaryotic [ARF10, Sed14r, LAB14]. evade [Sho13u]. event [SMZL13]. events [GP10, OMV^{+}11, YHT^{+}10]. everything [Sho11-31]. Evi5 [LAR^{+}12]. evidence [KNPK^{+}10]. evoked [BNM^{+}14]. evolution [ACO12, Hyn12, WG11, WD11]. Evolutionary [RZA^{+}13, RFRV12a, RFRV12b]. evolving [BT12]. Ewald [Sed12f]. exceeds [CWPW11]. Excess [KFL^{+}14, HSY^{+}14]. exchange [CM12c, GSW^{+}11, LCS^{+}10, Lev11, iYGL^{+}10]. exchanges [dSJDD^{+}11]. excision [DWL^{+}11, EIE^{+}14, MHV12, PVM^{+}12]. excitability [KK13a, MRR^{+}12]. excitation [ELH14]. exclusion [GC13]. execution [FA12]. exert [NGL^{+}12]. exhibit [DBH^{+}11, TTC^{+}14, ZZW^{+}13]. exhibits [FRS^{+}13]. Existence [PDMBW11]. existential [Sed11l]. existing [GB10]. exists [OTLH10]. exit [AYS^{+}13, BKP11, FWJ^{+}11, GKR^{+}11a, GKR^{+}11b, KMS10, LLS^{+}11, MMV^{+}10, PZ14, RGF^{+}10, RY11, SYS^{+}14, Sho10n, Sho12i, Sho14-32, VSMC11, vZOtR^{+}10]. exits [HTT11b, Les14-27]. exo [SRCP14]. exo- [SRCP14]. Exo70p [LN14]. Exo84 [LZLG13]. exocyst [LZLG13, MI13b, MRCC^{+}13, RMT13, SYH^{+}13]. Exocytosis [TID^{+}10, AOE^{+}10, AOE^{+}12, BNM^{+}14, CZC^{+}11, DYS^{+}14, GYZ^{+}12, GTS10, JOR^{+}11, KF11, Les13o, MMU^{+}10b, MRCC^{+}13, NWD^{+}11, Pfe10, TPM^{+}13, WWB^{+}10, WH13, WMV^{+}14, XRO^{+}11]. exocytotic [PGAE^{+}13]. exonucleolytic [SML^{+}13]. Exosome [CSP^{+}10, HMiY^{+}10, Sho10-53]. Exosomes [HHS13a, HHS13b, CRL^{+}14, RS13]. expansion [GZZ^{+}14, XZC^{+}12, ZSH10]. expectations [Sed13s]. expels [NWD^{+}11]. explain [CSHS^{+}13a, CSHS^{+}13b]. explanation [Sho12-55]. explore [VvDV^{+}10]. Exploring [HA12, Sho14q, Sed12y, Sho10-31]. export [BVR11, LBS^{+}13, LYB^{+}10, SYS^{+}14, SBP^{+}10a, Sho12-46, Sho13d, SIO10, TLTW10, ZDS^{+}12]. express [Les10-41, Les13-35]. Expression [EAB^{+}14, BAAW11, BG10, BAB12, CMD^{+}13, KAAM11, LMC^{+}12, MHC^{+}12, PMP^{+}11a, PMP^{+}11b, SBTF13, SZE^{+}11, TPZ^{+}14, UAH^{+}12]. extend [UHKS11]. extended [ENG^{+}12, LLK11]. extension [ATW^{+}10, KNOM11, KBW^{+}12, KSR^{+}13b, dMSMZ14, SRU^{+}12]. extra [BKK^{+}10]. Extracellular [LMS^{+}13, MCS^{+}13, RS13, BZ12, CAK^{+}14, CLL^{+}10, Hyn12, KCK^{+}14, LWW12, PMHZ10, TLTW10, VŠH^{+}11]. extracts [HH14b, WRF^{+}13]. extraproteasomal [MHCvSW11]. Extreme [STG13]. extrude [Sed13l]. extrusion [GFSR11, TGG^{+}11]. eye [KWK^{+}11, Les11-38, Sed11d]. ezrin [VOSB12].

F [BLO^{+}12, CWL^{+}11b, HTT13, HHY^{+}12, KBC^{+}14, LNL11, WDG^{+}13, YZM^{+}12a, ZBJL^{+}10]. F-actin [BLO^{+}12, CWL^{+}11b, HTT13, HHY^{+}12, KBC^{+}14, WDG^{+}13]. F-actin-mediated [YZM^{+}12a]. F-box [LNL11]. face [Les12q]. faces [DK10b, RC11]. facilitate [MFR^{+}14, OTLH10, SSK^{+}13, TCB^{+}14]. facilitates [ARF10, CDD13, CCJ^{+}12, DLBG11, FEHF12, GPCK12, HKI^{+}13, PHW^{+}13, RMG^{+}12, WGN^{+}13, XZC^{+}12, ZBJL^{+}10, ZKW^{+}13]. factor [BKS^{+}11, BMRM13, BPDB^{+}11, BAB12, CDB^{+}14, ETRP12, GBSC^{+}12, GHK^{+}10b, KKY^{+}14, LPG^{+}10, Les12-31, LSW^{+}14, LSM^{+}11, MHAK^{+}12, NWD^{+}11, NKH11, OWC^{+}10, RMM^{+}10, SDB^{+}10, TLTW10, YMM^{+}10, ZNH^{+}11, ZDS^{+}12, Mis13]. factor-mediated [ZDS^{+}12]. factories [GB10, yLFAM13, SKN^{+}13]. factors [BPH^{+}14, CM12c, FHA10, GCV^{+}11, HHS^{+}14, LNL11, LCS^{+}10, LTJN^{+}12, Mit12b, RT10, iYGL^{+}10]. factory [GB10, Les13d, Sho13m, TGB10]. fail [MJJ^{+}10]. failed [DWJ^{+}14]. failure [KRS11, ZYH^{+}11]. FAK [CWC^{+}13, JCN^{+}14, LLU^{+}12a, LLU^{+}12b, PCC11, PSR^{+}10, SF12, Sho10t, WM12, ZPS^{+}10]. FAK-mediated [PSR^{+}10]. fall [Les13-32]. Family [iYGL^{+}10, KYP^{+}14, LCS^{+}10, Sho11-58, TSB^{+}14, WAG^{+}10, YFLH12]. Family-wide [iYGL^{+}10]. fan [Les12c]. FANCJ [Les13q, Les13r, SNSyN13]. Fanconi [GHK10a]. fancy [Sed14n]. Farp1 [CB12, Sho12p]. Farquhar [Sed13p]. Fasciclin [GWR12]. fascicular [RSRK13]. Fascin [ZJP^{+}12]. fast [DAB^{+}11, KBAW^{+}11]. faster [Sho13r]. fat [OSD^{+}14]. Fate [MSC^{+}10, Gil10, MBO^{+}14, NSZ^{+}13, Sho12x, Sho13-46, UKZ^{+}13, YMU^{+}10, YMU^{+}13]. Fate-determining [MSC^{+}10]. fates [BPMK^{+}14]. FATP1 [XZC^{+}12]. fatten [Sho11-29, Sho12n]. FBH1 [JRC^{+}13a, JRC^{+}13b]. features [KWL^{+}12, MOZ^{+}13, SKFH11]. feed [Les12-27]. Feedback [LMS^{+}10b, BCB14a, MVP^{+}10]. feel [Les11p, Les12o, Les13i]. feeling [Sho10a]. fellowship [Sho12q, Sho14r]. female [CRL^{+}14]. fences [EEP13]. FERM [LS13a]. Fernando [Sed14g, Sed12g]. Feroz [Sed14h]. Fewer [Sho13r]. FG [SKFH11]. FGF [MMdCOM^{+}11, SYK^{+}11]. FGFR1 [CG12b]. FGFR2 [ALS^{+}13]. FHOD1 [AA13]. FHOD3 [ILD^{+}10]. fiber [DHB^{+}14, GBK^{+}14, OBSG12, OBC14, ZWL^{+}14]. fiber-mediated [DHB^{+}14]. fiber-specific [OBC14]. Fibers [Les11i, BW13, GF11, OZT^{+}13, Sho12-57]. fibrillar [KNPK^{+}10]. Fibrillin [NLAS^{+}10]. Fibrillin-1 [NLAS^{+}10]. Fibroblast [YMM^{+}10, DPW^{+}11, DPW^{+}12, SRU^{+}12]. fibroblasts [WAJ^{+}12, Six12]. fibrocystin [FLVP10]. fibrogenic [APV^{+}12]. fibronectin [BMFC^{+}11, BMFC^{+}13, FRL^{+}13, NRM^{+}12]. fibrosis [LMS^{+}10b, OL12, Sho12r]. fidelity [NMB^{+}14]. fifth [Sho11s]. filament [CP11, FS14, GLM^{+}10, HTS11, NSBW10, PKG10, PLC^{+}11, SSZ^{+}14]. filament-forming [NSBW10]. Filamentous [PBG^{+}13, VSG^{+}12]. filaments [CZGG12, DBH^{+}11, HHJ^{+}11, HTS11, JLVH12, JKS14, MAE^{+}10, SBEM13, SGLV10, WBML11]. filopodia [TKMK10, ZJP^{+}12]. filtering [VPC^{+}14]. finally [FSR11]. find [Les11k, Les13h, ME13, Sho10n, Sho13v, Six12]. finds [Les11l, Les11q]. fine [KPH^{+}12]. fine-tuning [KPH^{+}12]. finger [DPV^{+}12, WGR^{+}12]. fingers [Sho10-71, Sed11h]. Fiore [Pow10]. FIP5 [WJW^{+}11]. firing [GB10, Sho13l, ZNP^{+}13]. firmer [Sho12r]. First [Sho14s, Les10z, LK12, NOS^{+}14, PW12, Sed12u, Sed13z, Sed14p]. Fission [CLW^{+}14, CSM^{+}12, ABP^{+}14, ALF^{+}13, AWB^{+}14, CWPW11, GCR^{+}12, HHY^{+}12, KBS^{+}10, LCLW11, LMS^{+}10c, MHKM11, MRLLS12, OWC^{+}10, RPO^{+}14, Sho14-41, WLZ^{+}14]. fit [Bab14]. five [MOZ^{+}13]. fix [Sho11-47, Sho14x, Sed12r]. Fixing [OL12]. flagella [CSAPLBD11a, CSAPLBD11b, Ish14, PBM^{+}11]. flagellar [CTD^{+}10, EIW^{+}12, GDS^{+}12, Kin13, SDS^{+}12b]. flatworms [Sed11a]. Fletcher [Pow14a]. flex [Les11-35]. Flies [Car12, DCO^{+}12, DCO^{+}16]. flight [OBC14]. FLIP [FBR^{+}10]. flipping [XBC^{+}13, Sho10x]. flips [Les12-28, Les13l, Sho12a]. Flo11p [VŠH^{+}11]. Floating [Sed12x]. floodgates [Les14e]. Flotillin [LOR^{+}10]. flow [APC^{+}13, BLO^{+}12, HLN^{+}11, TAGJ11, TTB^{+}13, YZPF12]. flow-induced [HLN^{+}11]. fluctuations [CPT^{+}12, CPT^{+}14, DCO^{+}13]. fluid [LCK^{+}13, Sho14t]. fluidity [KTB^{+}14]. fluorescence [KSW^{+}11, NRK^{+}13, SHL10]. fluorescent [ACS^{+}13, CTM^{+}14b, NRK^{+}13, SRCP14]. flux [DGH^{+}14, RBF^{+}12, Sho14-35, WBMCSS13]. Fly [Les12n, Les13s, Sho14g]. Fn14 [MBK^{+}10]. Focal [ATW^{+}10, Sho12s, BFG^{+}13, CRP^{+}14, CWC^{+}13, DWPC^{+}11, DHB^{+}14, GTS10, LZR^{+}11, LMC^{+}12, OBSG12, PSR^{+}10, SSdA^{+}14, Sed11c, Sho12-57, Sho14-30, SW12, TTB^{+}13, WM12]. foci [BLM^{+}11, BCJ13, TALR11]. focus [TALR11]. focused [HCCS^{+}11]. focuses [HZM^{+}13]. fold [Les14j]. folding [LJPJ11, Sev10]. follicle [AAE^{+}14, HZS^{+}10, HSJ^{+}13, MRLLS12, OLT11]. folliculin [PRFF13]. Follistatin [PBD^{+}13, WWT^{+}12, SGD^{+}10]. Follistatin-mediated [WWT^{+}12]. follow [Sho11-33]. following [HLL^{+}12, MBO^{+}14, VES^{+}11]. follows [ILD^{+}10, PKD^{+}11, Sed13h]. For3 [CSKW13]. Force [EHUD14, FSA^{+}10a, BMÁG^{+}14, Boe12, BW13, CTM^{+}14a, CLEZ12, CSTBM^{+}10, CSHS^{+}13a, CSHS^{+}13b, HOS^{+}12, RC13, Sed11c, SHV^{+}13, WtLK^{+}13]. force-bearing [CTM^{+}14a]. force-dependent [HOS^{+}12]. force-generating [BW13]. force-regulated [CLEZ12]. forces [CYLMM13, KWTR10, LNTR14, MGFG^{+}10, Sed12c, YRU^{+}13]. fork [MFA^{+}14, MFR^{+}14, SNSyN13, ZNP^{+}13]. forks [Les14l, MFR^{+}14, RZF^{+}11, Sho13r, VYM^{+}10]. form [BKS^{+}11, BCB14a, HKR^{+}10, KMC^{+}14, KHB^{+}11b, OZT^{+}13, RFVE^{+}10, SRS10, SYV14, SRBL13, Sed13h, Sho10o, VŠH^{+}11, WBcY^{+}11, vGCMA^{+}14, Sho11-59]. Formation [KBKW10, LLK11, PHB^{+}13, AVP^{+}14, ASB^{+}11, ADAB^{+}12, BLM^{+}11, BBJ^{+}10, BLC^{+}14, CLSO^{+}12, CLS13, DK10a, DGS^{+}10, DYI^{+}13, EZT^{+}12, ETC^{+}12, GBL^{+}11, HTT^{+}11a, HRK13, IMP^{+}12, IM11, ISZ^{+}11, JDL^{+}14, KOO^{+}14, KPJ^{+}13, KKY^{+}14, KIOY10, KSR^{+}13a, KLvdB^{+}13, KYOY13, LMA^{+}13, LGAC13, LCP13, LHGT^{+}12, LLR^{+}12, LOR^{+}10, LXTM12, MI13a, MALS10, MRPR12, NOT^{+}14, NLAS^{+}10, OOKH^{+}12, OKNP13, ONH^{+}12, PMHZ10, PASG^{+}12, PoLC^{+}13, PLC^{+}11, PTST12, QMHM10, QECC10, RJvD11, RKE14a, RKE14b, RBM^{+}11, SSL^{+}14, SB14, SZJ^{+}10, SWS^{+}11, SJZ^{+}10, SFL12, SLH^{+}14, WJW^{+}11, YYM^{+}11, YKW^{+}12, YSN^{+}10, YKT^{+}13, ZJP^{+}12, ZBBG10, vRJMvD10]. formed [TYN^{+}13, TW14, VBG^{+}13]. Formin [ILD^{+}10, Sho12t, Sho13s, AA13, OBD^{+}10, PDMBW11, VLI^{+}14, vGLWB12, GJP^{+}13]. forming [NSBW10]. formins [CSKW13, MMVK^{+}12]. forms [BAY^{+}11, HSTF13, SDS^{+}12b, VSG^{+}12]. fortunes [Sho13-38]. forward [Les10-36]. fosters [SST^{+}12]. found [Sho10-49]. foundation [LS13a]. Four [RC11]. Foxj1a [JYRL^{+}13]. FoxO [CLZ^{+}14, NB12]. FOXO1 [PXZ^{+}13]. Fra [BBJ^{+}10]. Fra-2 [BBJ^{+}10]. Fra-2/ [BBJ^{+}10]. fragile [BBW^{+}13, GZZ^{+}14, Sho14-50]. fragility [MTM^{+}10]. fragmentation [NSS^{+}10, WTBM12]. Frances [Sed13i]. Fraser [KTN^{+}12]. free [FSA^{+}10b]. freely [GHGH11]. Freeman [Sed14o]. frequent [ELH14]. FRETting [Sho13t]. Friends [dWMR10, RS13]. Frizzled [ASB^{+}11, Sho11-40]. Frizzled-9 [ASB^{+}11, Sho11-40]. Frog [Sho13u]. front [Sho11-36, VMNLB^{+}11]. frontiers [Gol12b]. frontotemporal [GKWG^{+}11]. FSGS1 [TB12]. FSGS3 [TB13]. FSGS3/ [TB13]. FSHD [CG10a]. Fsp27 [GSW^{+}11, Sho11-29]. FtsZ1 [TO12]. FtsZ2 [TO12]. fuel [Sho14-37]. Full [KSP^{+}11, ATKK11, GZZ^{+}14, Sho10r]. full-of-bacteria [ATKK11]. Fumiyo [Sed12h]. function [ABP^{+}12, Bez12, BZ12, CLM^{+}10, CYN^{+}13, CSS^{+}14, CFB^{+}13, DD10a, DJL^{+}12, EBB13, EIW^{+}12, FMPS^{+}12, GRHA^{+}12, HTS11, HFB^{+}10, HH10, ILD^{+}10, JGA^{+}11, JCN^{+}14, JAM^{+}13, KPSL12, KKY^{+}14, KSB^{+}13, KPI^{+}10, LTJN^{+}12, Les11-37, LZY^{+}12, NJS^{+}10, OHC10, QWL^{+}11, RBY^{+}11, RMF^{+}10, SSL^{+}14, Sed10k, Sed13h, SFB^{+}12, Sha10, Sho10f, Sho10t, Sho10-43, Sho10-69, Sho13e, Sho13g, Sho14i, SJRV14, SLC^{+}13, SHV^{+}13, TNV^{+}13, TPM^{+}12, TMG12, VSMC11, VOSB12, WLK^{+}11, WWHH10, WAG^{+}10, YHK10, YZPF12, ZPS^{+}10]. Functional [OCF^{+}10, UG10, VBG^{+}13, WD11, AGM^{+}10, BKS^{+}11, BDR^{+}10, DGS^{+}10, FWM^{+}10a, OT11, RPO^{+}14, SKFH11, SC10b, UTK^{+}13, vGCMA^{+}14, TW14]. Functionally [HGV^{+}14, BGC^{+}14]. functioning [JVS^{+}14]. functions [AvCG^{+}11, BKG10, CMW11, CHK^{+}10a, CHK^{+}10b, CLD11, DWL^{+}11, DMK^{+}12, DSW^{+}11, ENG^{+}12, ECJB10, GCR^{+}13, GR11, HC10, HHC^{+}11, IIN^{+}11, Jan14, JB12, KFET11, LCL12, LZW^{+}10, MALS10, MGR^{+}10, RC13, Sho10-27, Sho12-54, TO12, VJK^{+}10a, VJK^{+}10b, WUD^{+}12, ZLFC14, ZPS^{+}10, ZGW^{+}14]. Fungal [Les12o, SST^{+}12, VSG^{+}12]. Fungi [Sed14c]. furlough [Sho13-48]. furrow [FLN^{+}10, FLN^{+}16, RW10]. Fus [YZL^{+}13]. Fusing [Sed14f]. fusion [AWB^{+}14, AKC^{+}12, BJE^{+}12, DYS^{+}14, DJL^{+}12, ELH14, JC10, LMS^{+}13, Les13-43, LAO^{+}10, OOKH^{+}12, PBD^{+}13, RMT13, RKRB12, SZJ^{+}10, SRKR10, SCN^{+}14, Sho12-42, SHS^{+}13, WLGC11]. future [HHS13a, HHS13b]. Fuz [BW12]. Fuzzy [Sho12u]. FYCO1 [Les10n, PAB^{+}10]. Fyn [Sho12v]. FYVE [Sed11h]. FYVE-fingers [Sed11h]. Fz7 [KBW^{+}12].

G [HAB14, BGC^{+}14, BLI^{+}10, CM12c, EMO12, FEHF12, KNOM11, KYP^{+}14, OPM^{+}12, TKS^{+}13]. G-actin [FEHF12, KNOM11]. G.W [BG11b]. G1 [BNDB^{+}14, CGRS^{+}12, HBC^{+}11, KSB^{+}13, MFGB10, Sho12a, WAG^{+}10]. G1/S [Sho12a]. G2 [LLM^{+}10, XHB^{+}10]. GAAC [Sho14u]. GAD67 [KKMB10]. Gag [RMG^{+}12]. Gain [KSB^{+}13, HRK13, Sho14-33]. Gain-of-function [KSB^{+}13]. Gaining [Sed10f]. gamete [ABD14]. gang [Les14t]. gap [IIN^{+}11, Les12-35, Pri14, Sho11g, BPL^{+}11, IKU^{+}11, LAR^{+}12, NLP^{+}10]. GAPDH [LKLA12, SWS^{+}11]. Gardel [Sed13o]. gastrulation [LBWS10]. gate [WLK^{+}11]. gated [AMGC14, BNM^{+}14]. gatekeeper [Sho12-37]. gatekeepers [Omr10]. gathering [Les13-41]. gating [TMG12]. Gaudenz [Sed12i]. gauze [GBL^{+}11]. GCC185 [BSP11]. GCY [WOG13]. GCY-8 [WOG13]. gd [VWD^{+}13]. GDNF [BST^{+}11]. GDP [YON^{+}12, iYGL^{+}10]. GDP-taxol [YON^{+}12]. gear [Sho10-37]. GEF [DKA^{+}13, GBiY^{+}14, HHC^{+}11, NGL^{+}12, QMHM10, RFVE^{+}10]. GEFs [Sho10-52, Sho13v]. Geiger [Sed11c]. gene [ATKK11, ALSN^{+}11, BG10, CAB^{+}13, HKN^{+}11, HPB10, KAAM11, KNH^{+}10, RKW^{+}13, SZW^{+}11, SBTF13, Sho13w, TPZ^{+}14, UAH^{+}12, Sho11z]. general [CZM^{+}14]. generate [SHV^{+}11, SHN^{+}11, VMNLB^{+}11]. generated [JDHS10, SDN^{+}14a, SDN^{+}14b]. generates [KEJ13, SHS^{+}13]. Generating [KK11, BW13]. generation [SHC^{+}10, UG10, FAvdB^{+}12]. genes [LZY^{+}12, SSK^{+}14]. genesis [DS12]. genetic [HCCS^{+}11, SWS^{+}11, VvDV^{+}10]. genetically [MBZ^{+}10]. genetics [Sed14s]. Genome [WBBD14, AvCG^{+}11, BNL^{+}10, DV10, wFLW^{+}13, GSP^{+}14, MGS14, OMV^{+}11, Sed12k, TDV^{+}14]. genome-wide [TDV^{+}14]. genomic [DKMK^{+}11, GL10, YWJ^{+}12, ZNH^{+}11]. genotoxic [NZHL13]. geometries [TBV^{+}14]. geometry [Sed10d]. germ [HYS11, HKR^{+}14, SCL^{+}14, UHKS11, Sed11q]. germ-line [SCL^{+}14]. germline [AGL^{+}14, GZR^{+}14a, GZR^{+}14b, KT10, Sho14-54, SL14]. get [Les10y, Les11d, Les11m, Les13p, Les14n, Lin10, Sed13j, Sed13z, Sho10-67, Sho11-56, Sho12y, Sho12-28, Sho13-41, Sho13-52, Sho13-50, Sho14-51, Six12]. gets [Les10n, Les11-31, Les11-42, Les13-33, Les13-40, Les14k, Sho10b, Sho10q, Sho10-54, Sho11-36, Sho11-49, Sho12c, Sho13y, Sho14-50, Sho14-64]. Getting [Nan14, FSR11, Sed11e, Sed11g, Sed12m]. giant [Sho11-57, Sho14-29]. gilgamesh [GOWM12]. Gillian [Sed13j]. give [Sho10-47, Sho11l]. gives [Les12-32, Sho10-38, Sho11-44, Sho12p, Sho14-68]. GKAP [MJEM10]. Gladfelter [Sed14c]. glands [CRL^{+}14]. Glc [PTS^{+}10]. Glia [CF13]. glial [BLT^{+}11, CC10b]. gliding [GS11]. Glimcher [Sed10k]. gliosecretion [PMP^{+}11a, PMP^{+}11b]. Global [Sho12w, SLC^{+}13, WPL^{+}11a, WPL^{+}11b, PHD^{+}10]. Globin [BKAB13]. globular [HFB^{+}10]. glucan [MCS^{+}13]. Glucose [OSD^{+}14, SLK^{+}13, BVL^{+}12, ZCB^{+}10a, ZCB^{+}10b]. glucosyltransferase [PTS^{+}10]. GLUT1 [OSD^{+}14]. GLUT4 [CWZ^{+}12, XRO^{+}11]. glutamate [ZIG^{+}12]. glutamine [CZM^{+}14]. glutamylases [GCR^{+}13]. glutathionylation [HIM^{+}10]. glycan [HS10b]. glycolysis [Sho14-34]. glycolytic [CZ10]. glycoprotein [LAO^{+}10, NOS^{+}14, PTS^{+}10]. glycoproteins [CSG14]. glycosphingolipid [HFS10]. Glycosphingolipids [Sho10z]. glycosylase [Sho14-33]. glycosylated [STG13]. glycosylation [GCSB10, Sho10-62, CSG14]. glycylases [GCR^{+}13]. Glypican [LSCF11]. Glypican-5 [LSCF11]. GMP [PBD^{+}13]. GMPCPP [YON^{+}12]. Go [Les10o, Les10r, Les10x, Sed14l, Sho13o]. Goda [Sho10-70]. goes [Les11c, Sho10-33, Sho12-27]. Gohta [Sed14i]. Goldstein [Sed13c]. Golgi [BDR^{+}10, BSP11, CDH^{+}14, CM12c, DT14, GCSB10, HMB14, HCG^{+}11, KKUG11, KYP^{+}14, LMS^{+}10c, LP13, OZT^{+}13, OTLH10, RPM^{+}13, RBF^{+}12, Sed10f, Sho11-51, Sho13-32, SC10b, WVvG^{+}13, XW10, YSN^{+}10, vGCMA^{+}14, vBAK^{+}12]. Golgi-associated [LP13]. gone [Dan14]. Good [Sho11u, Sho10m]. Goshima [Sed14i]. Got [Sed14w]. governing [SSZ^{+}14]. governs [PMP^{+}11a, PMP^{+}11b, SME^{+}13]. GPCR [DCL^{+}12, HSI^{+}11, KSP^{+}11, KBAW^{+}11]. GPCR-dependent [HSI^{+}11]. GPCR-mediated [KBAW^{+}11]. GPI [FWJ^{+}11]. GPI-anchored [FWJ^{+}11]. GPR158 [OPM^{+}12]. GPR158/ [OPM^{+}12]. grabby [Les13-33]. Graca [Sed12j]. gradient [HRK13, KAS^{+}12, PHB^{+}11, Sho14-42, VSG^{+}12, WBL11]. gradients [MMdCOM^{+}11]. gradual [CSM^{+}12, CSEH12]. Graham [Sed10f]. granular [PGAE^{+}13]. granule [BKT13, BXB^{+}12, HYS11, YCP10]. granules [KYOY13, LKSG13, PHB^{+}13, Sho13-45, TPM^{+}12, TPM^{+}13, UHKS11, YZL^{+}13]. GRASP [MALS10]. GRASP55 [XW10]. GRASP65 [XW10]. Grasping [Sed10d]. Grb2 [ALS^{+}13]. Great [Sed11b]. Greatwall [OMW^{+}14, Sho13-36, Sho14j, WGN^{+}13]. Grebe [Sed12q]. Greco [Sed14w]. green [Sho14-49]. Grg4 [LTJN^{+}12]. Griffiths [Sed13j]. Grinstein [Sed12t]. grip [Les10y, Sho10h]. GRK5 [CWL^{+}11b]. groove [DSM^{+}11, Les11d]. ground [Sed10f]. Group [DJL^{+}12]. grow [Dun11, Les13-39, Sho12-64, Sho14n]. growing [Sed12b, Sho11-38]. growth [AIJI11, ANT^{+}12, BMG14, BDB^{+}14, CPS^{+}13, CG12a, GCP^{+}14, GBSC^{+}12, GSJS10, GHK^{+}10b, GSW^{+}11, JDHS10, KLP14a, Les11g, Les11-39, LCP13, LBD^{+}14, LZLG13, MHAK^{+}12, MMVK^{+}12, MHCvSW11, dJPAA^{+}11, PRM^{+}14, RT10, RBA^{+}11, Sed14x, Sho12-40, TKMK10, TTB^{+}13, TLTW10, VSG^{+}12, YMM^{+}10, ZNH^{+}11, vGLWB12, vdVMG^{+}11]. GTP [MLSM^{+}11, NNO^{+}11, iYGL^{+}10]. GTP-bound [MLSM^{+}11]. GTP-tubulin [NNO^{+}11]. GTPase [AA13, AKC^{+}12, BPL^{+}11, HMiY^{+}10, LZW^{+}12, LZW^{+}13, PLR^{+}13, Sho10-41, dMSMZ14, SCR12, ZFA^{+}13]. GTPase-activating [HMiY^{+}10, PLR^{+}13]. GTPases [ADS^{+}13, BT13, Bar13, BHB^{+}11, LCS^{+}10, LWZ^{+}10, MP13, PRFF13, TCN14]. Guangshuo [Sed14j]. guanine [LCS^{+}10]. guanylyl [PBPW^{+}14a, PBPW^{+}14b]. guidance [Les14b, LNS^{+}13, PCCR11, Sed11p, SLH13]. guide [BHB^{+}11, SHL10, Sho12-57]. guided [QJO10]. guidelines [Sho11-33]. guideposts [Sed11p]. guides [HTT13, PJS^{+}11, Sho14-52]. Guilt [PMF12]. guilty [Les13-28]. gut [BDR^{+}12]. guts [Les13e]. GW [CLSO^{+}12]. GW/P [CLSO^{+}12]. GW220 [CLSO^{+}12, Sho12x]. GW220/TNGW1 [CLSO^{+}12].

H [CMW11]. H1 [RSM^{+}13, Sho13y, ZJP^{+}10]. H2 [BDN^{+}13]. H2A.Z [GSS^{+}11]. H3 [WMB^{+}10]. H3.X [WMB^{+}10]. H3.Y [WMB^{+}10]. hail [Sho12-53]. hair [AAE^{+}14, OLT11, Sed14w]. Han [Sed10i]. hand [Sho10-47]. handle [Sho11p]. handles [Sed11n]. Hands [Sed11h]. hang [Sho10-32]. Hansenula [SKVvdK11]. haploinsufficiency [WTH^{+}11]. Harald [Sed11h]. Harvesting [Sho14v]. Harvey [Sho11v]. Haspin [WUD^{+}12, DMK^{+}12]. Hax1 [CBBH11]. HB [DYI^{+}13]. HDAC3 [HSK^{+}10]. HDAC3-dependent [HSK^{+}10]. HDAC4 [BWL^{+}11, KAAM11, OMZK14, Sho11-43]. HDAC6 [DT14, LNT^{+}10]. HDAC6-dependent [LNT^{+}10]. HDM2 [GL10, NB10]. HDM2-mediated [GL10]. headed [KHB^{+}11b]. headpiece [ZZS13]. heads [Les14-27, OVL11, Sho11t]. heal [MKS^{+}13]. healing [PXZ^{+}13, SSW^{+}13, Sho13q]. health [ER10, RvD13]. healthy [Sho11u]. hearing [SKM10b]. heart [QWL^{+}11, Sed12e, Sho11-34, Sho12-64, VLI^{+}14]. Heat [Sho13w, KUH^{+}14]. heavy [ECJB10, LHS10, RSD^{+}12, VGL^{+}14]. Hebbian [VG13]. HECT [LNJ^{+}13]. HECT-like [LNJ^{+}13]. Hectd1 [SZ12a]. Hedgehog [ACO12, Les11j, CTH^{+}11, HZS^{+}10, HSJ^{+}13, HK14, LSCF11, Sho10k, Sho10y, TLS10, ZFP^{+}13]. hefty [Les14h]. Hegde [Sed10n]. Heidi [Sed14k]. height [CM12a]. heirs [Sho11-58]. Heisenberg [Sed13d]. held [Sho10j]. Helder [Sed13k]. helicase [HMBC10, RMG^{+}12]. helices [SDS^{+}12b]. helix [WGR^{+}12]. help [Les10-42, Les10y, Les11m, Les11s, Sho13v, Sho13-33, Sho14-31]. helper [Les12m]. helpful [Les12m]. Helping [Les10p, Les11k]. helps [Les11t, Les12u, Les13g, Les13h, Les13-31, Les13-39, Les14j, Sho11l, Sho11-56, Sho11-54, Sho12i, Sho12k, Sho12n, Sho12-28, Sho13c, Sho13o, Sho13n, Sho13-27, Sho13-31, Sho13-41, Sho13-43, Sho14n, Sho14-37, Sho14-33, Sho14-44, Sho14-51]. hemagglutinin [SMT^{+}10]. hematopoietic [PWP11, ZFA^{+}13]. hemichannel [IIN^{+}11]. hemorrhagic [CYN^{+}13]. hemylation [HHS^{+}14]. Heparan [BST^{+}11]. Heparanase [MBR^{+}11]. hepatocellular [IWS^{+}11, VES^{+}11]. hepatocyte [SWS^{+}11]. hepatocytes [SWS^{+}13]. heptameric [FMPS^{+}12]. her [Sed10b]. HERC2 [DPV^{+}12]. herd [Pal10]. heterochromatic [KLP^{+}14b]. Heterochromatin [ZLW^{+}13, BAS^{+}14, KK13b, RDLT11, STI^{+}11, SMZL13]. heterodimers [SRZ^{+}11]. heterooligomerization [JGA^{+}11]. Hetzer [Sho10-35]. Heuser [Sed12l]. hGAAP [SPC^{+}13]. hibernating [OBM^{+}10]. Hic [PTST12]. Hic-5 [PTST12]. hide [PMF12]. hides [Sho10-44]. Hierarchical [OGD^{+}12, AiIK^{+}13, MBVT^{+}13, WBcY^{+}11]. hierarchy [FSLM11]. HIF [COG11, KWK^{+}11, BMRM13, Sho13x]. HIF- [KWK^{+}11]. HIF-1 [COG11, BMRM13, Sho13x]. High [FSA^{+}11, PBPW^{+}14a, PBPW^{+}14b, YOMM^{+}11, BCJ13, CG10b, FAvdB^{+}12, HKR^{+}10, KSW^{+}11, SHC^{+}13, VvDV^{+}10]. high-content [CG10b, VvDV^{+}10]. high-load [YOMM^{+}11]. High-resolution [FSA^{+}11, YOMM^{+}11, BCJ13]. high-throughput [VvDV^{+}10]. Higher [SMZL13]. Higher-order [SMZL13]. highlight [BCB^{+}14b]. highlights [YSN^{+}11]. Highly [KSS^{+}11]. Hilton [Sed10e]. Hippo [Les13t, HBM^{+}11, HK14, LKG^{+}13]. His58 [CWC^{+}13]. Histone [Les14l, Sho13y, ZJP^{+}10, GSS^{+}11, IAMH10, Les11l, MFA^{+}14, MKL^{+}13, RSM^{+}13, TNH^{+}11, WBcY^{+}11, WMB^{+}10]. histones [Les10-40]. history [Sed14r]. hit [Les10b, Sho10g]. HIV [GdAJ^{+}12, Les12p, Les14m, MPM11, RMG^{+}12, SDN^{+}14a, SDN^{+}14b, Sed13m, STI^{+}11, Sho11w]. HIV-1 [MPM11, RMG^{+}12, STI^{+}11, Sho11w]. HJURP [BKS^{+}11, BSR^{+}11b, BSR^{+}11c, Sho11x]. HMGB1 [DKM^{+}13, Les10-42, TKL^{+}10]. hMis14 [KIOY10]. hnRNP [LCfC11, MPRT11, MTT^{+}14]. hold [Les14l, Sed13a]. holds [Les10w, Les14-32, Sed11u]. holin [HMO^{+}14]. home [Les12-31, Sed10e, Sho10k]. homeostasis [BDR^{+}12, CFB^{+}12, DD10b, GDO13, HIM^{+}10, HVDG13, KSSK12, NSZ^{+}13, VB12, YMM^{+}10, ZCB^{+}10a, ZCB^{+}10b, vGCMA^{+}14]. Homeostatic [HHL^{+}11, RPK^{+}11, RH10, VG13]. homologous [BSR^{+}11a, LH11, PHD^{+}10, TYN^{+}13, WRCD12]. homologue [DSW^{+}11, OFS^{+}10, SYH^{+}13, ZLJ^{+}13]. homology [YTM^{+}11]. homology-dependent [YTM^{+}11]. homotypic [AKC^{+}12, SGT^{+}13]. honorable [Sho10b]. Hook [BSR^{+}14, Sho14-36]. Hook1 [MBCKD13]. HookA [ZQA^{+}14]. Hop [Pal14]. hop-off [Pal14]. Hop-on [Pal14]. HOPS [CLM^{+}10, KWDD10, Sho10-27]. hormone [ZGCG^{+}14, ZWL^{+}14]. hormones [Les14-33]. host [AMO^{+}11, GCV^{+}11, MBM^{+}10, SYV14, HW11].

Hot [Sed13k]. housekeeping [Sho11u]. Howard [Sed14l, Sho11z]. HP1 [BSR^{+}11a, KIOY10, STI^{+}11]. HPat [HBI^{+}10]. HPK1 [Les11n]. HRD1 [BGC^{+}10, KXN10, RKG^{+}12]. Hrr25 [TTM^{+}14]. HSF2 [EAB^{+}14]. Hsk1 [MHKM11]. hsp [RKW^{+}13]. hsp-16 [RKW^{+}13]. Hsp100 [WTBM12]. Hsp110 [MWP^{+}12]. Hsp42 [SMMB11]. Hsp70 [BNH12, HPB10, WTBM12]. Hsp90 [DK10a, SZ12a]. HsSAS [KOO^{+}14]. HsSAS-6 [KOO^{+}14]. HtrA2 [YHG^{+}14]. HtrA2/Omi [YHG^{+}14]. hub [PPD^{+}10]. hubs [TDV^{+}14]. HuD [SHC^{+}13]. Huei [Sed12p]. Human [BBY^{+}12, DKY^{+}12, KUH^{+}14, PLL^{+}12, WBS^{+}12, CBB12, DK10a, GdAJ^{+}12, Gol12b, HHL^{+}11, KOO^{+}14, KIOY10, KBG12, KRS11, LZW^{+}10, MBM^{+}10, SML^{+}13, TC10, TPM^{+}12, VWD^{+}13, WK12, ZNH^{+}11, ZLH^{+}14]. humanized [BBY^{+}12]. humans [ABD14]. Humphries [Sed13q]. huntingtin [ETI^{+}10, XHS^{+}13]. Huntington [HPB^{+}12, JCL^{+}11, Sho11a]. HURP [BKK^{+}10]. hybrid [HS10b]. hydrolysis [IHM13]. hydrophobicity [TMG12]. hyperactivation [RJvD11]. hyperresorptive [XTX^{+}13]. Hypertrophic [TPSS12]. hypertrophy [DWM^{+}12, RMM^{+}10, RBP^{+}13, WWT^{+}12]. hypomorphic [MJJ^{+}10]. Hypoxia [Sho13z, BMRM13, CZ10, DYI^{+}13]. hypoxia-induced [CZ10].

I-band [FS14]. I-BAR [QJO10]. IA [DCP^{+}10]. IAP [SRS10]. Ib [BDR^{+}10]. ICAP [FRL^{+}13]. ICAP- [FRL^{+}13, BMFC^{+}13, BMFC^{+}11]. Id2 [GBSC^{+}12]. ideas [Sed12x]. Identification [GCV^{+}11, NSBW10, WMB^{+}10, ZGW^{+}14]. identifies [ASE10, BAH^{+}12, wFLW^{+}13, JDB^{+}12, RSL^{+}11, RKRB12, TSH^{+}14, TMG^{+}10, TDV^{+}14]. Identifying [CAB^{+}13, Sho10-28]. identity [Bar13, FSK^{+}10, HKI^{+}13, KST^{+}10, Les13h, NBC^{+}12, SPJ^{+}14, Sho11-51]. idle [Les10v, Sho11-40, Sho12-33]. If [Sho14x]. ified [Sho10-67]. IFT [Les10r, LAH^{+}12, Sho12-27]. IFT-A [Sho12-27]. IFT52 [TKB^{+}14]. IFT52/46 [TKB^{+}14]. IFT70 [TKB^{+}14]. IFT70/52 [TKB^{+}14]. IGF [GSC11]. IGF-II [GSC11]. Ignoring [Dan14]. IgSF9b [WKN^{+}13]. II [ADB^{+}14, AFM^{+}13, BSR^{+}11b, BSR^{+}11c, BRF^{+}10, BDN^{+}13, FLN^{+}10, FLN^{+}16, GSC11, HAB14, KNW^{+}14, KPC^{+}10, LGAC13, LCS^{+}10, Les13h, Les14u, Les14-34, MSR10, NBDB12, NWD^{+}11, OYH13, PSR^{+}10, RSD^{+}12, RKW^{+}13, RGB^{+}13, RFK^{+}10, SSV^{+}12, Sho10-41, Sho11-36, SFL12, UAH^{+}12, VJK^{+}10a, VJK^{+}10b, WVT^{+}13, YSO^{+}11, YZPF12, ZNH^{+}11, ZSZ^{+}13, ZJP^{+}10, lDSB^{+}10a, lDSB^{+}10b, vGLWB12]. II-dependent [RKW^{+}13, lDSB^{+}10a, lDSB^{+}10b]. II-mediated [BDN^{+}13]. II-spectrin [HAB14]. IIA [DWPC^{+}11, VMNLB^{+}11]. IIA/IIB [VMNLB^{+}11]. IIB [VMNLB^{+}11]. III [AVP^{+}14, BCBG10, HSKAT11, Les11h, Les14k, MVC^{+}11, SSZ^{+}14, WAW^{+}11, YDB^{+}11]. III/MVB [DCL^{+}12]. Ikeda [Sed12h]. IKK [BLC^{+}12, BWL^{+}13, TTC^{+}14]. IKK/ [BWL^{+}13]. IL-1 [TTC^{+}14]. illuminate [HW11]. illuminates [Sed14b]. image [DV10, LRA^{+}10, TSH^{+}14, TMG^{+}10]. image-based [TSH^{+}14]. Imaging [MC10, Sho11-27, WPSA13, ACS^{+}13, BCJ13, DSP11, DE10, KSP^{+}11, NRK^{+}13, RKT^{+}14, RMT13, RKK^{+}14, SBR^{+}11, SRCP14, TP13, YSN^{+}11, YOMM^{+}11]. imbalance [Les11-38]. immature [FSK^{+}10]. immediately [APC^{+}13]. Immobile [WVT^{+}13]. immune [ATKK11, SA10a, Sed10k]. immunity [BZ12, CLC^{+}11]. immuno [CAB^{+}13, Sho13-49]. immunoglobulin [KNH^{+}10]. Immunogold [Sed13n]. immunological [GTS10, SHS^{+}13, TQS^{+}11]. immunoregulatory [KDIE11]. impact [GWR^{+}10, Mis13, NBC^{+}12, PBvdS12, WK12, Mis13]. impair [KSB^{+}13, LNT^{+}10, LgLM^{+}10]. impaired [MBZ^{+}10]. impairs [CYN^{+}13, ETI^{+}10, MHCvSW11, VSMC11, WCC^{+}10]. Impenetrable [BC11a, BC11b]. implanting [MC10]. implicated [LH11]. implicates [DSP11]. implications [DNB13, ER10, GS11, KBS^{+}10, SBR^{+}11, She14, VBG^{+}13, WCC^{+}10]. import [BWK^{+}11, HHJ^{+}11, HLL^{+}12, JLW^{+}10, LSW^{+}14, PKD^{+}11, SDD^{+}13]. importance [SZ12b]. important [AHL^{+}11, CP11, LMS^{+}13, WVvG^{+}13, YKW^{+}12]. importin [HHJ^{+}11, Les11-34, LSW^{+}14, RDB^{+}12]. importin- [HHJ^{+}11, Les11-34, RDB^{+}12]. imprinting [ZNH^{+}11, ZZW^{+}14]. improved [BvMD^{+}14]. inactivating [EMO12]. inactivation [BMS^{+}11, KSSK12, VES^{+}11]. INCENP [XOV^{+}10, NCT^{+}11]. INCENP-aurora [XOV^{+}10]. incision [OMV^{+}11]. including [RDB^{+}12]. inclusion [SWS^{+}11]. incompatible [VLKI14]. inconsequential [Bar13]. incorporate [HHY^{+}12]. increase [KTB^{+}14]. Increased [DSL13, RCBY^{+}12]. increases [BMS^{+}11, DYI^{+}13, Sho12-59, WTH^{+}11]. incredible [MI13a, Sed14v]. Independence [SRP^{+}13]. independent [AvCG^{+}11, AKB^{+}13, BSP11, DSW^{+}11, DCL^{+}12, HSKAT11, HKR^{+}10, LSS^{+}12, MBCKD13, NSD^{+}14, OMW^{+}14, PDMBW11, RCG^{+}10, RCG^{+}11, ZJP^{+}12, ZPS^{+}10, YHF13]. independently [BKY^{+}10, LCD^{+}11, RDLT11, WWT^{+}12, WAG^{+}10]. Indirect [HPB^{+}12, OBC14, YOMM^{+}11]. Individual [SSH^{+}13]. individuality [Sed14c]. induce [BKBR11, HDH^{+}10, LSOT10, LgLM^{+}10, MMB^{+}11, NZHL13, QTL^{+}12, QTL^{+}13, Sho13-38]. induced [CPX11, CZ10, CTW^{+}10, FUK^{+}14, GJP^{+}13, GDO13, HLN^{+}11, KNOM11, KST^{+}10, KUH^{+}14, LSW^{+}14, LLcK^{+}11, MFGB10, MLG^{+}10, MBM^{+}10, MDP^{+}10, MHAK^{+}12, MBK^{+}10, ONNB^{+}14, RC12, SFJ^{+}14, TCX^{+}10, BKE10, LLA^{+}12, ONH^{+}12, PTST12, SPJ^{+}14, XYM^{+}10]. inducer [RBP^{+}13]. induces [AMH11, CMD^{+}13, FPM^{+}14, GWP^{+}11, HBC^{+}11, HVOF^{+}14, JBS^{+}12, JBS^{+}13, KKS^{+}14, RJvD11, WHA^{+}13, YMU^{+}10, YMU^{+}13, YKT^{+}13, ZGCG^{+}14]. inducible [DPV^{+}12, EAB^{+}14]. inducing [MMO^{+}14, TTC^{+}14, VWD^{+}13, XTH^{+}11]. induction [PBD^{+}13]. inequality [Les13j]. INF2 [ADAB^{+}12]. infection [AMO^{+}11, BC11a, BC11b, FMI^{+}13]. infections [Sho14-46]. inflammation [PSVRB^{+}11, Sed10i]. inflammatory [LMC^{+}12, SJM^{+}13]. influences [PPG11a, PPG11b, RGF^{+}10]. influenza [GSU^{+}12]. ingression [FLN^{+}10, FLN^{+}16]. Inheritance [MLW13, CGRS^{+}12, EJBW12, Les13p, Sho12-38, LeB10]. inhibit [BWL^{+}13, MLH12, PDKG14, WSZ^{+}12]. inhibiting [ALS^{+}13, BKP11, CJNS12, GL10, HSS^{+}13, YYA^{+}11]. Inhibition [AYS^{+}13, JCN^{+}14, ZDS^{+}12, ACO12, CZ10, FAB^{+}10, HDK^{+}13, HPB^{+}12, PSF^{+}11, RB11, dSMSS13]. inhibitions [Sho13-33]. inhibitor [DMK^{+}12, STD^{+}10, Sho10y]. inhibitors [GWR^{+}10, Gol12a, WUD^{+}12]. Inhibitory [AMH11, MGT^{+}10, MSC^{+}10, PYT^{+}13, WKN^{+}13]. inhibits [AKB^{+}13, BPB^{+}12, CNP^{+}12, DPW^{+}11, DPW^{+}12, DT14, EM11, GB10, GVP^{+}11, HDH^{+}10, HZT^{+}12, KWH^{+}10a, KWH^{+}10b, KWTR10, KMG^{+}11, LCS^{+}10, LMS10a, LR13, LWK^{+}13, LSW^{+}14, OMW^{+}14, PGB^{+}10, RC12, VES^{+}11, XHS^{+}13]. initial [BLI^{+}10, JKS14]. initiate [KLZ^{+}12, MWH12, MKH^{+}14]. initiates [AMGC14, DSM^{+}11, LR11a, LJPJ11, MRLLS12, NOS^{+}14, OYH13]. initiating [MHV12]. initiation [ETRP12, GZR^{+}14a, GZR^{+}14b, GCSB10, KFH^{+}12, LADS10, MHKM11, PASG^{+}12, RFAA^{+}12, GK13]. Injured [Les13v]. Injury [GDO13, SMK14, XWE^{+}10]. Injury-induced [GDO13]. Inke [Sed10g]. innate [CLC^{+}11, SA10a]. Inner [KIOY10, bCAH^{+}11, COB^{+}12, GSS^{+}11, HCCS^{+}11, LLH13, Les10-33, Les11u, MAD10, SHN^{+}11]. innovation [HSF12]. Ino80 [Les10s, SKM10a]. iNOS [KLC^{+}10]. inositol [NPL^{+}10]. Inoué [Sed11r]. Ins [Les10t, Sed11m]. insert [TPM^{+}13]. insertase [KIL^{+}12]. insertion [TGG^{+}11]. insertion/extrusion [TGG^{+}11]. insight [She14, TKB^{+}14]. insights [DNB13, GCC12, Ish14, Sev10, SHS^{+}12, XG12]. instability [GSP^{+}14, SPD^{+}13, TMG^{+}10, YWJ^{+}12]. instructs [SCR12]. insufficiency [RCFH10]. insufficient [BWS^{+}10]. insulator [SRBL13]. insulin [CWZ^{+}12, JOR^{+}11, XRO^{+}11, ZNH^{+}11, ZCB^{+}10a, ZCB^{+}10b]. insulin-like [ZNH^{+}11]. insulin-stimulated [CWZ^{+}12]. integrate [YFLH12]. Integrated [LZW^{+}10]. integrates [NTSK14, OMZK14]. Integrating [VvDV^{+}10, HMBC10, TKS^{+}13]. Integration [MGFG^{+}10, KdKDP12, LJLJ11]. integrin [ATW^{+}10, BFG^{+}13, BMFC^{+}13, CLEZ12, DPW^{+}11, DPW^{+}12, FRL^{+}13, GRHA^{+}12, KTN^{+}12, LCfC11, Les12-36, LCHB13, MBVT^{+}13, NRM^{+}12, ONH^{+}12, OLB13, PPV^{+}14, RCM^{+}12, RIG^{+}12, RBH^{+}12, SNR^{+}11, SYS13, SF12, Sho10-30, Sho13i, Sho13-55, TKS^{+}13, WHDR^{+}10, YHT^{+}10, YZM^{+}12b, ZZS13, BMFC^{+}11]. integrin-dependent [FRL^{+}13, TKS^{+}13]. integrin-signaling [PPV^{+}14]. integrin/ICAP [BMFC^{+}13, BMFC^{+}11]. integrin/ICAP- [BMFC^{+}13, BMFC^{+}11]. Integrins [GNHB11, Boe12, KYHG12, Les12-32, MVP^{+}11, Mit12a, SHBC12, WBS11]. Integrity [BHA^{+}12, AKA^{+}13, BNL^{+}10, CHL12, FCE^{+}12, GL10, MGS14, NSS13, OMV^{+}11, RGL^{+}13, Sed12k, Sho12k, SJZ^{+}10, VTM14, ZC11]. interact [BEJ10, JDB^{+}12, LOR^{+}10]. interacting [MVC^{+}11, PSK11, RKRB12, KKL^{+}14, WWS^{+}12, ZGW^{+}14]. Interaction [PYT^{+}13, WJW^{+}11, ALF^{+}13, CJNS12, CHL^{+}14, DPW^{+}11, DPW^{+}12, DCN^{+}10, GKWG^{+}11, GZLG11, HCCS^{+}11, HIB^{+}10, KAAM11, KWL^{+}12, LN14, MKH^{+}14, NCT^{+}11, PPD^{+}10, RSS^{+}13, RBF^{+}12, SMdP^{+}14, SYS^{+}14, TTB^{+}13, VvDV^{+}10, WWM^{+}12, ZYF^{+}11]. Interactions [MTT^{+}14, CNP^{+}12, GB12, HW11, HBC^{+}10, MBM^{+}10, MRR^{+}12, MPRT11, PDKG14, PKS^{+}10, RBY^{+}11, VWD^{+}13, XOV^{+}10, ZSD^{+}14, ZZW^{+}14]. interactome [COB^{+}12]. interacts [CO13, CLC^{+}11, GKR^{+}11a, GKR^{+}11b, HKH^{+}10, KWDD10, LHL11, LWL^{+}13, MI13b, SME^{+}13]. intercalation [WMP^{+}14]. intercellular [AGL^{+}14, KLS^{+}13, Sho12-58, SMT^{+}10]. interchromosomal [UG10]. Interdomain [KPE^{+}14]. interest [Sho14-38]. interface [LMA^{+}13, Oef10, Sed13o, XZC^{+}12]. intermediate [JVS^{+}14, SBEM13, SGLV10, WBML11]. intermediates [NZHL13]. internal [Les12-34, SYS13]. internalization [BVM^{+}11, CNP^{+}12, YSM10]. internodes [IHG^{+}12]. interphase [ABP^{+}14, DE10, EL14, FP10, JG10, LC10, LR13, MGT^{+}10, NEMH^{+}10, RKG^{+}10, TH11, vdVMG^{+}11]. interplay [GCH^{+}14, VCF^{+}13, VG13]. Interruption [ZNH^{+}11]. Intersectin [RFVE^{+}10]. interspersed [CSH^{+}12]. intertwines [TIM14]. intestinal [AIJI11, CMT14, GvEM^{+}11]. intestine [Sho11s]. intoxication [GCV^{+}11]. intracellular [Bez12, BGC^{+}14, DSK^{+}11, MVR^{+}10, SGC10, SZ12a]. intrachromosomal [ZNH^{+}11, ZZW^{+}14]. Intraflagellar [SW10b, BW12, EIW^{+}12, LWZ^{+}10, Sho12u, TKB^{+}14]. intralumenal [HSR^{+}10]. intramolecular [HBS^{+}10, MLSM^{+}11, SHV^{+}11]. intrinsic [BHMBS^{+}11]. intrinsically [TL12]. invaders [Sho13-57]. invading [Sho14-43]. invadopodia [BWBC^{+}14, DYI^{+}13, HZM^{+}13, MRCC^{+}13, OOKH^{+}12, PTST12, SGLV10, Sho13z, YYM^{+}11]. invadopodia-driven [HZM^{+}13]. invadopodial [HKN^{+}14]. invadopodium [SLH^{+}14]. invadosome [JDL^{+}14]. invagination [MGFG^{+}10]. invasion [ABP^{+}12, CBB12, DJL^{+}12, HKN^{+}14, MLM^{+}11, MVN11, ONH^{+}12, PTST12, SGT^{+}13, SLH^{+}14, VFNR11, WHF^{+}11, YHG^{+}14]. Invasive [WM12, FPM^{+}14, RCM^{+}12, SHC^{+}10, SZJ^{+}10, Sho14-61, YZM^{+}12a]. inversely [ZNP^{+}13]. Investigating [Sed11n]. involved [BPH^{+}14, Bon14, DAB^{+}11, RTC^{+}13a, RTC^{+}13b, YSN^{+}10, ZGEM12]. involving [KLS^{+}13]. ion [BLI^{+}10, SNT^{+}12]. IP [DWM^{+}12]. Ipl1 [NCT^{+}11]. Ipl1/ [NCT^{+}11]. IQGAP1 [JGB^{+}13]. IRE1 [CCGN11, RPK^{+}11]. irradiation [JEF^{+}11]. IRSp53 [CFLDM11, Sho11-28]. IRSp53-mediated [CFLDM11]. Isabelle [Sed10h]. ischemia [HLL^{+}12]. isn’t [Sho11e, Sho11-57, Sho12-61]. Isoform [DGS^{+}11, CVR10, ECJB10, ILD^{+}10, STG13, TPM^{+}13]. Isoform-specific [DGS^{+}11]. isoforms [GLM^{+}10, GF11, KWK^{+}11]. isolates [Sho13-57]. isolation [YST^{+}11]. Isotropic [KCF^{+}14]. ISWI [KLP^{+}14b]. ISWI- [KLP^{+}14b]. ITS1 [SML^{+}13]. itself [NB12, Sho14w]. IV [HLT12].

Jacobs [Sho10p]. Jagesh [Sed11i]. JAK [MBVT^{+}13, Sho13-55]. JAM [IMP^{+}12]. JCB [WM11, WCM12a]. Jiahuai [Sed10i]. jigsaw [CD14]. JIP1 [mFH13]. Jiri [Sed12k]. JNK [GWP^{+}11, HRWW^{+}13, NSSF10, SEV^{+}14]. JNK-mediated [HRWW^{+}13]. Joan [Sed10j, Sed11j]. job [Les14-30]. Jody [Sed13l]. John [Sed12l, Sed13m]. Joining [Oef10, CFB^{+}13, CWG^{+}11]. jointly [HSN^{+}11]. Jonathon [Sed14l]. Joubert [CHK^{+}10a, CHK^{+}10b]. journey [MI13a, ME13]. Jun [FHD^{+}12, RBM^{+}11]. JunB [GBSC^{+}12, KHB^{+}11a, RMM^{+}10]. junction [EZT^{+}12, HTT11b, IMP^{+}12, IIN^{+}11, KUN^{+}13, MLG^{+}10, OT11, RBY^{+}11, SAG^{+}11, SBS^{+}12, SNZVK12, SNZVK13, TB12, ZME^{+}14]. junctional [GLG12]. junctions [AFM^{+}13, BG11a, BLT^{+}11, CVR10, HOS^{+}12, Les11x, Les11-41, Les13a, NGL^{+}12, SME^{+}13, Sho10-61, Sho12h, Sho12y, Sho13-39, SEV^{+}14, SRZ^{+}11, TIT11, TB13, YMT^{+}13, lDSB^{+}10a, lDSB^{+}10b]. Jürgen [Sed13n]. just [Les13s, Sho14-42].

K-Ras [MLH12]. Kar3Vik1 [RCC^{+}12]. Karen [Sed12m]. KASH [ETYS^{+}12, HKW^{+}13, MSZ^{+}12, Sho14v, ZGEM12]. KASH5 [Sho13-27]. katanin [GTR^{+}13]. Kcnq1 [ZZW^{+}14]. KDM1A [MKL^{+}13]. keep [Les11i, Les11-32, Sed12v, Sho10z, Sho12-33, Sho12-51]. Keeping [Sho14y, Sed11k]. keeps [Les11b, Les11r, Les13-34, Mit12a, Sho10k, Sho10c, Sho11-28, Sho11-48, Sho12-47, Sho12-60, Sho12-62, Sho13l, Sho13k, Sho14w, Sho14-35, Sho14-48]. Keith [Sed12n]. Kelch [HC10]. Kenneth [Sho10-31]. keratin [RC12, SLK^{+}13, WBML11]. keratinocyte [DPW^{+}11, DPW^{+}12, DKA^{+}13, RC12]. keratinocytes [YMM^{+}10]. Keratins [KLS^{+}13]. key [EMO12, HSF12, Les10-35, Les12l, Les14f, MH14, TSH^{+}14, TMG^{+}10]. KH [IIWS14]. Khodjakov [Sed12a]. Kicking [Les14o]. kidney [CC10a, PPG11a, PPG11b]. KIF13B [KWH14, Sho14z]. KIF14 [ATU^{+}12]. KIF17 [EAK13, HBS^{+}10, JK10]. Kif2A [UTK^{+}13]. KIF3A [GdAJ^{+}12]. KIF4 [SSV^{+}12, Sho12-28]. KIF4A [BGB^{+}13, BCB14a, Sho14-27]. kill [Sho10x]. killer [AMH11, Les11w]. killing [Les12d]. kills [SDN^{+}14a, SDN^{+}14b]. Kin4 [BKP11, CKO^{+}10]. Kinase [YWJ^{+}12, ALS^{+}13, AGM^{+}10, AKA^{+}13, BNDB^{+}14, BVC^{+}11, BKP11, BPL^{+}11, CDD13, CKO^{+}10, CCGN11, CLD11, CFLDM11, CTW^{+}10, Dan14, DPL^{+}12, DCP^{+}10, FWM^{+}10a, FAB^{+}10, GR11, GL10, HDK^{+}13, HGV^{+}14, HLN^{+}10, KPE^{+}14, KDIE11, KMS10, KWO11, KWL^{+}12, KSR^{+}13b, KSSD11, KKK^{+}11, LC10, LS13a, Les11o, LZR^{+}11, LMC^{+}12, LVB^{+}10, LP11, LDL12, LMS^{+}10c, MSR10, MHKM11, MCHCC10, MBO^{+}14, OMW^{+}14, PPG11a, PPG11b, RCM^{+}12, RJvD11, RJM^{+}12, RCG^{+}10, RCG^{+}11, RPK^{+}11, RBM^{+}11, SKH^{+}10, TSB^{+}14, WGN^{+}13, WDB10, XWE^{+}10, XOV^{+}10, YYM^{+}11, YFLH12, ZPS^{+}10, ZEG11, MMS^{+}10, SDS^{+}12b, YDB^{+}11]. kinase-1 [LDL12]. Kinase-dead [YWJ^{+}12]. kinase-independent [RCG^{+}10, RCG^{+}11, ZPS^{+}10]. kinase-mediated [CTW^{+}10]. kinase/phosphatase [FWM^{+}10a]. kinases [BNL^{+}10, BCBG10, MBVT^{+}13, RCG^{+}10, RCG^{+}11, SOW^{+}11, SHC^{+}10, Sho13-58, SWC13, YBN^{+}11]. kinastrin [DLBG11]. kinastrin/SKAP [DLBG11]. Kindlin [SNR^{+}11]. Kindlin-3-mediated [SNR^{+}11]. kinesin [BSR^{+}14, EAK13, mFH13, GdAJ^{+}12, HBS^{+}10, HS10a, JDB^{+}12, LMS10a, Les12p, NNO^{+}11, RCC^{+}12, RDPG14, RHK11, Sho13t, WBMCSS13, WDB10, CS13, FS10, RCC^{+}12, WVvG^{+}13]. kinesin- [WBMCSS13]. kinesin-1 [NNO^{+}11, FS10]. Kinesin-1-syntaphilin [CS13]. kinesin-14 [HS10a, RCC^{+}12]. kinesin-2 [HBS^{+}10]. kinesin-3 [BSR^{+}14]. kinesin-5 [HS10a, RHK11, WVvG^{+}13]. Kinesin-5/Eg5 [WVvG^{+}13]. kinesin-8 [RDPG14, WDB10]. kinesins [NAS^{+}11, NAS^{+}12, NAS^{+}13, NSD^{+}14]. kinetic [SBR^{+}11]. kinetics [DE10, HIB^{+}10]. Kinetochore [JKA^{+}10, Les13w, BKS^{+}11, CDD13, CM12b, CYLMM13, CJNS12, CLO^{+}11, CSHS^{+}13a, CSHS^{+}13b, CD14, DK10a, DSL13, DMK^{+}12, DWDW12, ECK^{+}12, FSOL14, GC13, GCR^{+}12, HSTF13, HTM^{+}14, JJH^{+}10, KIOY10, LHW10, LMA^{+}13, Les13-32, LVB^{+}10, LDL12, MGK^{+}12, MHS10, MS14, MAD10, NvCL^{+}13, PPD^{+}10, PKS^{+}10, RKE14a, RKE14b, Sho10-66, Sho14m, Sho14-52, SMS^{+}14, SJ13, SHN^{+}11, TUG^{+}10, VTO^{+}13, ZSD^{+}14, KWL^{+}12]. kinetochore-microtubule [DSL13]. Kinetochores [Les11p, Sho10-32, BRL14, FSOL14, JDS^{+}10, MOZ^{+}13, MKH^{+}14, PZ14, RC13, RDB^{+}12, SKH^{+}10, Sed12v, Sho11-52, Sho14d, VWC^{+}13, WHL^{+}12, Sed10p]. Kip3 [RDPG14]. Klar [GYC^{+}14]. Klein [Sed11p]. KLHL20 [LLcK^{+}11]. Klionsky [Sho10r]. KMN [CJNS12]. KNL [ECK^{+}12]. KNL-1 [ECK^{+}12]. KNL1 [CDD13, LVB^{+}10, VTO^{+}13, EUB^{+}14, KWL^{+}12, Les12q]. knockout [WPL^{+}11a, WPL^{+}11b]. knockouts [VKMI12]. know [Nan14, Sed11e, Sed13i]. knowledge [Sed12n]. knows [Sho14-27]. Koehler [Sed13e]. Kornbluth [Sed14v]. Koushika [Sed13t]. Kozlov [Sed14q]. Kruppel [KKY^{+}14]. Kruppel-like [KKY^{+}14]. Ku [BCJ13, CWG^{+}11, Sho11-41]. Ku70 [ETI^{+}10]. Ku70-mediated [ETI^{+}10]. Kuzbanian [DCO^{+}12, DCO^{+}16]. Kv [VYC^{+}11]. Kv1 [VYC^{+}11]. Kv1.1 [SHC^{+}13].

L [BGC^{+}10, CZD^{+}13, DCL^{+}12, GGSN^{+}13]. L-mediated [GGSN^{+}13]. L-type [CZD^{+}13]. L166P [DGS^{+}10]. L17 [LJPJ11]. labels [DZT^{+}11]. lability [CWS^{+}11]. Lack [ERS10]. lacking [KHB^{+}11a]. lacks [LLM^{+}10, SLM^{+}13]. lacritin [MBR^{+}11]. lamellipodia [GdBP^{+}14, MKS^{+}13, SRU^{+}12]. lamellipodial [IHM13]. Lamellipodin [LVK^{+}13, Sho13-28]. lamellipodium [KNOM11]. Lamin [DCO^{+}13, HSI^{+}14, MWG^{+}12, RCG^{+}10, RCG^{+}11, Sho14-28, CSTBM^{+}10]. Lamin-A [Sho14-28]. lamina [Les12r]. Laminin [HKN^{+}10, CYN^{+}13]. Laminin-based [HKN^{+}10]. LAMTOR2 [SSdA^{+}14]. LAMTOR2/3 [SSdA^{+}14]. large [FAvdB^{+}12, SBP^{+}10b, SDD^{+}13, ZNA^{+}14]. large-scale [SBP^{+}10b]. larger [Sho12-64]. Lasp [Sho14-29, FS14]. Late [Sho14-30, CTM^{+}14b, Les14-34, NSS^{+}10, SSdA^{+}14, XHB^{+}10]. latent [AEC^{+}14, KCK^{+}14, MSS^{+}10]. later [Les10z]. lateral [BMÁG^{+}14, HAB14, KWTR10, PJS^{+}11]. laterally [SSV^{+}12]. LATS1 [CSS^{+}12]. LATS1/ [CSS^{+}12]. lattice [MMVK^{+}12, OZT^{+}13]. Laura [Sed14m]. Laurie [Sed10k, Sed12o]. layer [KEJ13]. LC3 [CKU^{+}10, IM11, PAB^{+}10]. LC8 [GDS^{+}12]. Lck [TQS^{+}11]. LDLR [KF11]. lead [MMU10a]. leading [FEHF12, HCC^{+}10, HKR^{+}10, LN11]. leads [CYN^{+}13, DPL^{+}12, DHVK10a, DHVK10b, GJP^{+}13]. learn [Sed13f]. learning [Sed12g]. leash [Les11b]. leaves [Les11-45, Sho10a]. ledger [CM12b]. Lehmann [Sed11q]. lend [dWMR10]. length [DWDW12, JYRL^{+}13, MWP^{+}12, RDPG14, TGES12, TIM14, UTK^{+}13]. lengths [PKG10]. lesions [MHV12]. less [Bra13]. lessons [ZF11, OB12, Sed14x]. let [Sho10d, KWTR10]. LET-99 [KWTR10]. lethality [DPL^{+}12, YWJ^{+}12]. lets [Les13b, Les13q, Les13r]. leukemia [CAB^{+}13, dTLLB12]. leukocyte [SYS13, TSB^{+}14]. Leukocytes [Sho12-29]. level [yLFAM13, dSLPRG11, ZNP^{+}13]. levels [FA12, HHL^{+}11, HZS^{+}10, HH14b, JYRL^{+}13, LCHB13, TGB10]. Leyser [Sed14s]. LGN [LDCF^{+}13, MdFF^{+}14, PJS^{+}11, SMdP^{+}14, ZZW^{+}10]. Li [Sed11o, Sed12p]. Li-Huei [Sed12p]. liberates [Les13-37]. licenses [Les11g]. Licensing [Mit12b, SQC^{+}12, ZMW^{+}13]. life [GdAJ^{+}12, MH11, VM14]. lifeline [Les13x]. lifespan [yLFAM13]. Lifetime [ACS^{+}13]. Ligand [GJP^{+}13, CLEZ12, HFS10, HSKAT11, KTN^{+}12, LNS^{+}13, PBPW^{+}14a, PBPW^{+}14b, SYS13, VWD^{+}13]. ligand-bound [SYS13]. ligand-independent [HSKAT11]. Ligand-induced [GJP^{+}13]. ligands [OLB13]. ligase [BAY^{+}11, BDN^{+}13, DSW^{+}11, HLH^{+}14, HZT^{+}12, KLF^{+}14, MMO^{+}14, ONNB^{+}14, PHW^{+}13, PMK^{+}13, RKRB12, RKG^{+}12]. ligase-dependent [DSW^{+}11]. ligation [CFB^{+}13]. light [CGK13, DHL^{+}12, JVS^{+}14, Les11-37, Sed12q, SNZVK12, SNZVK13, Sho14u]. light-triggered [CGK13]. Lighting [Sed11f, Sed11r]. like [AKC^{+}12, BNDB^{+}14, BVL^{+}12, FHA10, HVOF^{+}14, HLN^{+}10, HCCS^{+}11, KKY^{+}14, KBW^{+}10, LNJ^{+}13, LSW^{+}14, LDL12, MLM^{+}13, SZJ^{+}10, ZNH^{+}11]. LIM [SHC^{+}10, Sed12y]. limit [HSI^{+}14, Les10-27]. limited [TC10]. limiting [KSH^{+}13]. limits [BGB^{+}13, BKE10, BAB12, CHL12, HZS^{+}10, HLN^{+}10, NB12, Sed12g, Sho10-60, Sho11d, Sho13-56, Sho14u, Sho14-27, WtLK^{+}13]. LINC [Sho13-27, Sho14q]. line [MMVK^{+}12, PW12, SCL^{+}14, Sho10g, UHKS11]. lineage [LZR^{+}11]. linear [JDL^{+}14]. lines [Sed12h]. Lingo [PSF^{+}11, Sho11-48]. Lingo-1 [PSF^{+}11, Sho11-48]. link [ANT^{+}12, Bez12, GF11, HBI^{+}10, LCS^{+}10, Sho14-53, WH13]. linked [LCL12, PYT^{+}13, RDC^{+}11]. linker [ZQA^{+}14]. Linking [GP12, LH11]. links [CB12, HZW^{+}12, JOR^{+}11, KDIE11, LDCF^{+}13, LSOT10, LYH^{+}13, MMC^{+}10, RMT13, SYK^{+}11, WBS11, vdVMG^{+}11]. Lipid [CRP^{+}14, LAR^{+}10, Les14p, Lev11, SWS^{+}13, Sho11-29, AHL^{+}11, GSW^{+}11, KHFV^{+}13, Les13x, MWG^{+}12, PGP14, Sho12n, TP13, XZC^{+}12, dSJDD^{+}11, XZC^{+}12]. lipid-mediated [MWG^{+}12]. lipidomes [GSU^{+}12]. Lipids [Sho14-31, HAB14, Sed12x, PGP14]. lipin [AHL^{+}11]. Lipins [Les12r]. lipophagy [WMC14]. Liprin [KHG^{+}13, SSK^{+}13]. Liprin- [KHG^{+}13, SSK^{+}13]. LIS1 [YOMM^{+}11, ŻKC^{+}11, ETRP12, Les11q, Les12s, Sho13-29, SCL11]. Lissencephaly [DSD^{+}13]. Lissencephaly-1 [DSD^{+}13]. Listeria [Sed11m]. little [Sho10-33]. Live [DE10, LWB^{+}14, RMT13, RKK^{+}14, YSN^{+}11, Les13u, RZS^{+}14, SQC^{+}12, TP13, WPSA13]. Live-cell [RMT13]. lived [LWBH12]. liver [Sho11u]. living [DMH^{+}12, HIB^{+}10, KSS^{+}11]. Liz [Sed10l]. LKB1 [BDC^{+}14, CAK^{+}14, Dan14, MSK^{+}13b, LSS^{+}12, Sho12-30]. LL5 [HKN^{+}10]. load [CPT^{+}12, CPT^{+}14, YOMM^{+}11]. loading [KSS^{+}11]. Local [CO13, FDB^{+}13, LADS10, SFB^{+}12, VOSB12, EAK13, LLT^{+}12, LhYL^{+}13, SB14, WMV^{+}14]. Localization [SDS^{+}12a, BVC^{+}11, GYC^{+}14, HFB^{+}10, HLS^{+}14, KKK^{+}11, LJW13, MTG^{+}11, MdFF^{+}14, MBLD11, NvCL^{+}13, OPM^{+}12, OPCEM10, RGL^{+}13, SZ12a, Sho14-47, SJRV14, VSMC11, WGN^{+}13, XOV^{+}10, YTT^{+}10]. Localized [AOE^{+}10, AOE^{+}12, wFLW^{+}13, LMC^{+}12, MXS10, WP14, YZL^{+}13]. localizes [CLD11, DLBG11]. locally [ASLS14, BGB^{+}13, TCN14]. Locating [Sho11-30]. Location [Sho11-31, RT10, vdBFS^{+}12]. locus [CAB^{+}13, WBcY^{+}11, ZZW^{+}14]. logic [Sed13v]. logistics [Sho12u]. LOK [VOSB12]. LOK/SLK [VOSB12]. Long [KK13b, RW10, Sho10-34, ZZW^{+}14, KMSR12, LWBH12, MDP^{+}10, Sho10-33, WPM14]. long-distance [WPM14]. long-lived [LWBH12]. Long-range [KK13b]. long-term [MDP^{+}10]. look [Sed13m, Sed13f]. Looking [HHS13a, HHS13b]. loop [BCB14a, SHV^{+}11, ZBBG10]. looping [ZNH^{+}11]. loops [SHV^{+}13]. loosens [Sho11-60]. lose [Mit12b, Sho10h, Sho13-33]. Loss [CPX11, DPL^{+}12, Les14q, MHCvSW11, OD10, PoLC^{+}13, SPD^{+}13, WHA^{+}13, CVJ^{+}11, CAK^{+}14, CZGG12, GGSN^{+}13, Sho10u, Sho14o, XTX^{+}13]. lost [Les11-45, MHCvSW11, Sho10-29]. Love [Sed13r]. Low [Sho13-30]. lowering [KWH^{+}10a, KWH^{+}10b]. LPAR2 [KTB^{+}14]. LPAR2-dependent [KTB^{+}14]. LRP1 [KWH14]. LRP6 [KPJ^{+}13]. LRRTMs [KSLF^{+}11]. LSD1 [MKL^{+}13]. LSD1/KDM1A [MKL^{+}13]. Lte1 [BKP11, GSJS10]. Ltn1 [ONNB^{+}14]. Lukas [Sed12k]. Lulu2 [NT11]. lumen [CFLDM11, LDCF^{+}13, RFVE^{+}10, RFAA^{+}12, WJW^{+}11]. luminal [LRH^{+}13]. Lymphatic [SMSP11, CMH^{+}10, LXTM12, XYM^{+}10]. lymphoma [SOW^{+}11]. lysine [SLK^{+}13]. Lysosomal [BKT13, CFB^{+}12, PBvdS12, Bez12, oHXK^{+}12, SHB^{+}10, SHBC12, TNV^{+}13]. Lysosome [IPM^{+}13, EBB13, HKN^{+}11, Les13m]. Lysosome-mediated [IPM^{+}13]. lysosome-related [EBB13, HKN^{+}11]. Lysosomes [Les10u, FUK^{+}14, FCA10, Krä13, MP13, PRFF13, SYV14].

M [MGT^{+}10, PLC^{+}11]. M-phase [PLC^{+}11]. M1 [SJRV14]. M1- [SJRV14]. M18BP1 [MMFS11]. M2 [MLM^{+}13]. M2-like [MLM^{+}13]. M23 [SJRV14]. ma [Sed13b]. Machesky [Sed14m]. machinery [FMI^{+}13, GMW^{+}13, LCL12, LYH^{+}13, PGAE^{+}13, Sho11m, Sho14-55]. machines [Pow14a]. macroautophagy [WCC^{+}10]. Macrophage [LDN^{+}13, FHY^{+}10, PSVRB^{+}11]. Macrophage-secreted [LDN^{+}13]. Macrophages [DSMB13, Sho14-32, GdAJ^{+}12, MLM^{+}13, PBG^{+}13, Sho10-68, SMM^{+}10]. macropinocytosis [GNHB11, KWH^{+}10a, KWH^{+}10b, VLKI14]. MAD [Sho14d]. Mad1 [BRL14, HTS^{+}10, MS14, MKH^{+}14, SFK^{+}13]. Mad1/Mad2 [SFK^{+}13]. Mad2 [HTS^{+}10, IP12, LMS10a, Les10v, MKH^{+}14, SFK^{+}13]. MAD2L2 [LS13b, Sho13-31]. made [CSTBM^{+}10, Sho10-70]. madness [Sed10c]. MagT1 [Sho14-33]. Maiato [Sed13k]. maintain [BGY^{+}13, CTM^{+}14a, CHL12, JVS^{+}14, MSK^{+}13b, SZW^{+}11, Sho12k]. maintained [HSI^{+}11, ZC11]. maintaining [BNL^{+}10, OD10]. maintains [BVC^{+}11, JYRL^{+}13, QB12, RMM^{+}10, Sho11u, TNV^{+}13, TPM^{+}13, WMCF10]. maintenance [AvCG^{+}11, BG11a, BSP11, CSTBM^{+}10, EIW^{+}12, HK14, ILD^{+}10, KST^{+}10, KFH^{+}12, MMVK^{+}12, NDS^{+}11, OLT11, PHD^{+}10, RG14, SNSyN13, Sed12k, WBBD14]. major [BRD^{+}13, COB^{+}12, FHA10, HVOF^{+}14, MSC^{+}10, TQM^{+}14, dJSTV12]. majority [PZ14]. make [Les10p, Les11s, Sed13d, Sho12i, Sho13r, Sho13-37]. makes [Les10-31, Les11o, Lev11, Mit12c, Sho10-50, Sho11q, Sho11w, Sho12-64, Sho13j, Sho13-44, Sho14d]. Making [Les13y, OVL11, Sed13g]. MAL [DKA^{+}13]. malaria [CBB12]. male [BAS^{+}14, CRL^{+}14]. malformations [GLG12]. Mammalian [CLSO^{+}12, HMBC10, HSY^{+}14, RMS^{+}14, SGD^{+}10, CSS^{+}12, CSEH12, DCN^{+}10, HKW^{+}13, LH11, MTG^{+}11, MSZ^{+}12, NOS^{+}14, OCF^{+}10, OTLH10, OWC^{+}10, RNS^{+}14, RPM^{+}13, RKRB12]. mammals [DBH^{+}11, DCO^{+}12, DCO^{+}16]. mammary [BWBC^{+}14, LRH^{+}13]. Man [Sho12-51, WHL^{+}12]. management [Sho10-63]. mandates [Ros13]. Manifestations [LR11b]. Manipulation [KWO11]. manner [CWB^{+}14, GLB10, GKWG^{+}11, lDSB^{+}10a, lDSB^{+}10b]. mannose [MLM^{+}13, NOS^{+}14]. many [Sed10l, Sed11j, Sho11-32]. many-splendored [Sed11j]. map [HCCS^{+}11, Sho11-30, FSA^{+}10b, Les10w, RCG^{+}10, RCG^{+}11]. MAP1B [RBF^{+}12]. Map205 [KPE^{+}14]. Map7 [GCP^{+}14]. MAPK [FWM^{+}10a, MJFS10, MWZ^{+}11, NMB^{+}14]. MAPK-signaling [MWZ^{+}11]. mapping [FSA^{+}11, FMPS^{+}12, VvDV^{+}10]. maps [FAvdB^{+}12]. marcescens [HMO^{+}14]. march [Sho11j]. MARCH2 [oHXK^{+}12]. Margaret [Sed13o, Sed14n]. margin [ZME^{+}14]. marginal [DGF^{+}14]. María [Sed11k]. Marie [LCL12, Sho12-55]. Marilyn [Sed13p]. MARK4 [KSR^{+}13b]. marks [LRH^{+}13]. Markus [Sed12q]. marrow [ZYH^{+}11]. Martin [Sed13q, Sho10-35, Sed14g]. MarvelD3 [SEV^{+}14]. masks [NCML^{+}12]. Mass [WM10, HSK^{+}10, KPH^{+}12, RMM^{+}10, WCQ^{+}13]. master [Dan14, Sed13n, Sho10-43, Sho14-56, ZDM^{+}14, SWF12]. Mastering [Sed13y]. Mastl [ADB^{+}14]. match [Sho12z]. Math [Sed14b]. Matrix [KPSL12, BPT^{+}14, CAK^{+}14, CLL^{+}10, Dan14, GMD^{+}10, Hyn12, KCK^{+}14, LMS^{+}10b, LWW12, MAS11, RSD^{+}12, RT10, VŠH^{+}11, WM12, WPM14]. matters [LRA^{+}10]. Matthew [Sed14o, Sho11-32]. maturation [ATKK11, BIY^{+}13, BPH^{+}14, BHA^{+}12, DHB^{+}14, HKR^{+}14, HSY^{+}14, IMP^{+}12, IKU^{+}11, KLZ^{+}12, KWDD10, KFS^{+}14, LR11a, NKH11, OBSG12, OFS^{+}10, PTS^{+}10, PLR^{+}13, RPM^{+}13, SJZ^{+}10, TMPH^{+}10]. mature [MWH12, Sho13-32]. maturing [Les14t]. maturity [Sho12-57]. Maxence [Sho10-36]. may [Les11-30]. MCAK [BDB^{+}14, DWDW12]. McBride [Sed14k]. Mcl [GWR^{+}10]. Mcl-1 [GWR^{+}10]. Mclk1 [LWBH12]. Mcm [KSS^{+}11]. McMahon [Sho11v]. MCPH1 [Les11r, YSO^{+}11]. MDC1 [WGC11]. MDCK [GSU^{+}12]. MDGAs [PYT^{+}13, Sho13-33]. mDia1 [OBD^{+}10]. Mdm2 [MMO^{+}14, Sho14-34]. Means [Les13z]. measure [Sho10-42, Sho13c]. Measurements [KNOM11]. mechanical [GMD^{+}10, LZW^{+}10, RFL13]. mechanics [TMPH^{+}10]. mechanism [Bab14, BCB^{+}14b, BKE10, CDH^{+}14, CSP^{+}10, DE10, EBBJ11, FAB^{+}10, HYTU^{+}10, HTM^{+}14, JTN^{+}13, KBAW^{+}11, OMW^{+}14, PKG10, SSW^{+}13, SNT^{+}12, TC10, TLS10, WAJ^{+}12, WWHH10, WMV^{+}14]. Mechanisms [Coo13, KOO^{+}14, LM13, McN13, BCJ13, CLS13, GK13, GHK^{+}10b, HBS^{+}10, HL11, HZE^{+}13, HCP^{+}13, Jan14, KKMB10, LCP13, LR11b, LHD^{+}14, MSC^{+}10, MLY^{+}10, OLB13, RK13, SSZ^{+}14, SLH13, WMB12, WRF^{+}13, ZKR^{+}11, GGR12]. Mechanistic [XG12]. mechanosensing [MAD^{+}11, PPV^{+}14, lDSB^{+}10a, lDSB^{+}10b]. Mechanosensitive [TIT11]. Mechanosignaling [OYYK14]. mechanotransducer [JBS^{+}12, JBS^{+}13]. mechanotransducers [BW13]. mechanotransduction [BPT^{+}14, OZ14, SY10]. mediate [CWZ^{+}12, DSK^{+}11, EMT^{+}14, FS10, GR11, HYS11, HYTU^{+}10, JDHS10, MGK^{+}12, MP13, PAB^{+}10, RBH^{+}12, SML^{+}13, SLK^{+}13, TCX^{+}10]. mediated [BKY^{+}10, BG11b, BMC^{+}11, CDK^{+}10, CCGN11, CTW^{+}10, CRJB^{+}11, DHB^{+}14, DAS^{+}10, FMI^{+}13, GdBP^{+}14, GOWM12, GHK10a, GK13, GHK^{+}10b, GSGL11, GCH^{+}14, GGSN^{+}13, GM11, HZT^{+}12, HRWW^{+}13, IAMH10, IPM^{+}13, KST^{+}11, KBC^{+}14, KHW^{+}10, KPJ^{+}13, KK13b, LvDG^{+}10, MLM^{+}13, MWG^{+}12, MLY^{+}10, PDKG14, RMS^{+}14, RSS^{+}13, SNR^{+}11, SFK^{+}13, Sho12-44, SMM^{+}10, TCN14, TDV^{+}14, VYC^{+}11, VBB^{+}10, VOSB12, WWM^{+}12, WJPD11, WWT^{+}12, YYM^{+}11, YHG^{+}14, YZM^{+}12a, ZDS^{+}12, BRF^{+}10, BJE^{+}12, BSP11, BDN^{+}13, CFLDM11, CCJ^{+}12, ETI^{+}10, GL10, HKN^{+}11, HVW^{+}10, JC10, KLS^{+}13, KBAW^{+}11, LNS^{+}13, MMU10a, NSS13, PSR^{+}10, RBP^{+}13, RZF^{+}11, RKS^{+}10, SBS^{+}12, SHS^{+}12, SRP^{+}13, SLH^{+}14, WMV^{+}14, ZZW^{+}14]. mediates [AKC^{+}12, ABD14, CS13, DCL^{+}12, EBBJ11, EDF^{+}10, GEN14, HLN^{+}11, HWB^{+}13, KFET11, KYP^{+}14, LMT^{+}10, LHS10, MDP^{+}10, OOKH^{+}12, PMB^{+}11, QJO10, RSB13, SKN^{+}12, SMB12, TSL12, WLGC11, WBMCSS13, WPM14, WMC14, XYM^{+}10, YZM^{+}12b]. mediator [DKM^{+}13, FHKW11]. medicine [Gol12b]. Meet [Sho12-31, Les11j, Sho12z]. meets [EIE^{+}14]. MEF2 [BWL^{+}11]. megakaryocyte [MI13a]. megalin [SBTF13]. MEI [GTR^{+}13]. MEI-1 [GTR^{+}13]. meiosis [ADB^{+}14, BM11, BAS^{+}14, DSL13, LH11, Les12j, Les12y, MSZ^{+}12, NMB^{+}14, SCL^{+}14, SLM^{+}13, WRCD12, YRU^{+}13, YSM10]. meiosis-specific [SCL^{+}14]. Meiotic [KST^{+}11, LHGT^{+}12, BKK^{+}10, CSEH12, EM11, HH14b, HMBC10, HSY^{+}14, HKW^{+}13, JPT^{+}11, LHN10, OHC10, Sho10-59, Sho12z, Sho13j, Sho13-27, YTT^{+}10, YKT^{+}13]. MEKK1 [SEV^{+}14]. melanogaster [HFS10, NSS^{+}10, NSBW10, WWHH10, YTT^{+}10]. Melanoma [Les14r, CAK^{+}14, DS12, SLH^{+}14]. Melanosomes [Sed12j]. Melina [Sed14p]. melt [Sho14-59]. member [RCC^{+}12]. members [Mis10]. Membrane [KSR^{+}13a, ZSZ^{+}13, ABVP11, AVP^{+}14, AMR11, AOE^{+}10, AOE^{+}12, ANT^{+}12, AXL10, AFRZ^{+}14, Bab14, Bar13, BPDB^{+}11, BWK^{+}11, BRD^{+}13, BJE^{+}12, Bra13, BKBS12, CLM^{+}10, CDAK10a, CDAK10b, bCAH^{+}11, CWL^{+}11b, CSS^{+}14, COB^{+}12, CTD^{+}10, FSLM11, GSS^{+}11, GKWG^{+}11, GSU^{+}12, GPCK12, HZM^{+}13, HKN^{+}14, HFS10, HAB14, HBG^{+}11, HCCS^{+}11, JLW^{+}10, JC10, KEJ13, KHS^{+}11, KKS^{+}14, KdKDP12, KLHS14, KWDD10, KPH^{+}12, KTN^{+}12, KIL^{+}12, KBKW10, LMT^{+}10, Les10-33, Les11m, Les11u, Les12i, Les12-33, LAO^{+}10, LAH^{+}12, LJPJ11, LJLJ11, LWB^{+}14, LYB^{+}10, LMS^{+}10c, MAS11, ME13, MMU^{+}10b, MMC^{+}10, NBC^{+}12, NSSF10, NBS^{+}11, OPCEM10, PKD^{+}11, PDKG14, PGAE^{+}13, PMB^{+}11, RLS^{+}14, SBS^{+}12, SSL^{+}14, Sed14q, SRKR10, SHS^{+}12, Sho10-35, Sho11v, Sho13p, TID^{+}10, TESA10, WMCF10, WAW^{+}11, WZHV11, WLK^{+}11, WWS^{+}12, XBC^{+}13, YKW^{+}12, YSN^{+}10, Pri14]. membrane-cortex [KEJ13]. membrane-less [Bra13]. membranes [BJ12, CHK^{+}10a, CHK^{+}10b, RKT^{+}14, Sho10j, Sho10-55, TCB^{+}14]. memories [Les11s, Sho10-70]. Mena [GRHA^{+}12]. merge [ABP^{+}14]. Merging [Sed10k]. mesenchymal [AEC^{+}14, BMRM13, GBSC^{+}12, KKL^{+}11, LZR^{+}11, PTST12, SMB12]. mesenchyme [SZ12a]. mesoangioblasts [GSB^{+}13]. message [Les11-42, Sed14e]. messenger [TLTW10]. Metabolic [LLS^{+}11, MFGB10, OS13, XG12]. metabolism [BLC^{+}12, Pri14, Sed14r]. metabolites [YBN^{+}11]. metacaspase [MFF^{+}13]. Metadata [LRA^{+}10]. metalloproteinase [WPM14]. metaphase [BRL14, CMS^{+}14, JKA^{+}10, Les13-44]. metaplasia [LDN^{+}13]. metastable [SYS13, WAJ^{+}12]. metastasis [BT12, BWBC^{+}14, JCN^{+}14, Les10-35, Sed11g]. metavinculin [JLVH12]. metazoa [HBI^{+}10]. metazoan [Hyn12, RSL^{+}11]. method [BCJ13, NSD^{+}14, Sed10c]. methylation [VLG14]. methyltransferase [ECC^{+}13, TNH^{+}11]. methyltransferase- [ECC^{+}13]. Mff [OWC^{+}10]. Mgr2 [GSM^{+}12]. MHC [RGB^{+}13, UAH^{+}12]. Mia3 [WPL^{+}11a, WPL^{+}11b]. Mia3/TANGO1 [WPL^{+}11a, WPL^{+}11b]. Mice [KHB^{+}11a, ZZW^{+}13, ABD14, BBY^{+}12, BMS^{+}11, DPL^{+}12, LWBH12, LHGT^{+}12, MJJ^{+}10, MBK^{+}10, NNSH11, PGB^{+}10, RKS^{+}10, SDS^{+}12a, SJZ^{+}10, YOA^{+}11, YWJ^{+}12, ZYH^{+}11]. Michael [Sed11l, Sed13r, Sed14q]. micro [MKS^{+}13]. micro-wounds [MKS^{+}13]. Microbe [HW11]. microclusters [LCBG^{+}11]. microdomain [WMC14]. microdomains [KHFV^{+}13, LOR^{+}10]. micrometer [TP13]. micrometer-scale [TP13]. micropatterning [LWB^{+}14]. microprocessor [KA12]. MicroRNA [AAE^{+}14, BAB12, QECC10, BAAW11, CTL^{+}10, GSC11, HVW^{+}10, KA12, Les12t, LNS^{+}13, LK12, Sho12-64, SLS^{+}10]. microRNA-1 [CTL^{+}10]. microRNA-125b [GSC11]. microRNA-206 [CTL^{+}10]. MicroRNA-214 [AAE^{+}14]. MicroRNA-30c-2* [BAB12]. microRNA-mediated [HVW^{+}10, LNS^{+}13]. microRNA-targeted [LK12]. microRNAs [Les11s, Les12e, MPD^{+}12, Sed13x]. microscopes [Les11-44]. microscopy [CG10b, FAvdB^{+}12, KSW^{+}11, SHL10, Sed11r, SNT^{+}12, YON^{+}12]. Microtubule [COG11, ECK^{+}12, GTR^{+}13, HS10a, MBCKD13, RBF^{+}12, YKT^{+}13, BGB^{+}13, BHB^{+}11, BDB^{+}14, BR14, CYLMM13, CJNS12, CLS^{+}10, CPX11, CWPW11, DK10a, DSL13, DT14, DWDW12, DLBG11, EHUD14, ENG^{+}12, GCP^{+}14, GCR^{+}12, GS11, HSN^{+}11, HKN^{+}10, JG10, JKA^{+}10, KOK^{+}13, KSR^{+}13b, LvDG^{+}10, LHW10, LMA^{+}13, Les11e, Les11-39, LMW^{+}11, LDL12, MGK^{+}12, MOZ^{+}13, MAD^{+}11, NCT^{+}11, NSS13, NB12, PAB^{+}10, PKS^{+}10, RMS^{+}14, RCC^{+}12, RDPG14, SME^{+}13, SMM^{+}10, SCL11, TMG^{+}10, TUG^{+}10, UG10, UTK^{+}13, VWC^{+}13, WBS^{+}12, WMP^{+}14, YWC^{+}13, ZSD^{+}14, vdVMG^{+}11, FSA^{+}10b]. Microtubule-dependent [MBCKD13]. Microtubule-organizing [YKT^{+}13]. microtubule-stabilizing [WMP^{+}14]. Microtubules [OZT^{+}13, Sho11-33, Sho13-34, ADAB^{+}12, BBK^{+}13, BKS14, BGY^{+}13, CTD^{+}10, ETYS^{+}12, FWM^{+}10b, FSOL14, GMW^{+}13, JK10, LCS^{+}13, LLT^{+}12, Les12i, Les12x, Les14c, Les14y, LMT^{+}12, MJEM10, NNO^{+}11, PHB^{+}11, RW10, SBS^{+}12, SGLV10, Sed11b, Sho10d, Sho11-38, Sho13-39, Sho14-58, SMS^{+}13, SW12, SWC13, SFL12, SHN^{+}11, WHH^{+}11, YON^{+}12, YMT^{+}13]. microvesicles [RS13]. Microvillar [Sho12-32, GLB10]. microvilli [Fis14, GB12]. Mid1 [RBB^{+}14]. MiD51 [Les14o, RPO^{+}14]. midbody [BB10, CTY^{+}12, EKJH13, EDF^{+}10, GMW^{+}13, Sho12d]. middle [DC12]. midgut [GDO13]. midzone [RDPG14]. MIG [SCR12]. MIG-10 [SCR12]. MIG6 [FAB^{+}10]. Migrating [Sho10-37, WAJ^{+}12, BDB^{+}14, DHVK10a, DHVK10b, FEHF12, FSA^{+}10a, HKR^{+}10, Les13t, VMNLB^{+}11, Sed14e]. migration [ALV^{+}12, CAK^{+}14, CDAK10a, CDAK10b, Coo13, DT14, DWPC^{+}11, DHVK10a, DHVK10b, FW10, GdBP^{+}14, GSB^{+}13, GNHB11, HSI^{+}14, HBSD12, HCC^{+}10, HYTU^{+}10, HCP^{+}13, HCG^{+}11, KTB^{+}14, KFL^{+}14, LAR^{+}12, LVK^{+}13, Les11y, LP11, LKG^{+}13, MVP^{+}11, MHAK^{+}12, MVN11, NBDB12, PGCY12, PCCR11, QJO10, RCM^{+}12, RIG^{+}12, RC12, SBEM13, SPC^{+}13, SSdA^{+}14, Sed11f, Sho10-31, Sho13x, Sho13-56, Sho14t, Sho14-28, Sho14-53, Six12, SRU^{+}12, TSB^{+}14, TGES12, VFNR11, VLG14, WWM^{+}12, WtLK^{+}13, WHDR^{+}10, YRU^{+}13]. migrations [FS10]. migratory [Les14c]. Miklós [Sed14r]. MIM [Sho10-38]. Mim1 [BWK^{+}11, PKD^{+}11]. Mim1-dependent [PKD^{+}11]. mind [Sed10b, Sed12e, MHS10]. Mindbomb [BAY^{+}11]. Mindbomb1 [DSK^{+}11]. minds [Les12-35]. mineralization [BMFC^{+}11, BMFC^{+}13, KPC^{+}11]. minerals [Sho11-56]. minimal [SRKR10]. Minimizing [WCM12b]. minus [EHUD14]. miR [ALV^{+}12, APV^{+}12, LSM^{+}11, QWL^{+}11, Sho12-33, SHC^{+}13, WSZ^{+}12, XTH^{+}11, ZWL^{+}14]. miR-1 [QWL^{+}11]. miR-129 [SHC^{+}13]. miR-133a1 [ZWL^{+}14]. miR-21 [APV^{+}12]. miR-22 [XTH^{+}11]. miR-24 [ALV^{+}12]. miR-29a [LSM^{+}11]. miR-29a/b [LSM^{+}11]. miR-34s [Sho12-33, WSZ^{+}12]. miR669a [CCM^{+}11]. miR669q [CCM^{+}11]. miRISC [CLSO^{+}12]. miRNA [DWM^{+}12, Sho12o, SGD^{+}10]. miRNA-1 [SGD^{+}10]. miRNA-133a [DWM^{+}12]. miRNA-34 [KKL^{+}11]. miRNAs [Sho11-34]. MIS12 [PPD^{+}10, DK10a]. Misfolded [Les10x]. MISP [ZSK^{+}13]. missegregation [vRJMvD10]. Misshapen [LCHB13, Sho13-35]. missing [Sho12-36, WH13]. MITF [MP13]. Mitochondria [Les10y, ASLS14, GRH^{+}12, HCCS^{+}11, HLS^{+}14, LLH13, MSS^{+}10, MVP^{+}10, OVL11, Pal10, Sho10-46, Sho13-54, Sed14k]. Mitochondrial [ELH14, JLW^{+}10, KZR^{+}12, PL11, She14, Sho12-34, Sho14-35, AWB^{+}14, BWL^{+}13, BWK^{+}11, BNH12, BKBS12, bCAH^{+}11, CS13, CPX11, COB^{+}12, CGRS^{+}12, DGH^{+}14, FHKW11, GSM^{+}14, GSM^{+}12, GKR^{+}11a, GKR^{+}11b, HIM^{+}10, HHS^{+}14, HCCS^{+}11, KIL^{+}12, KBS^{+}10, LNJ^{+}13, LNT^{+}10, MSS^{+}10, MRLLS12, OWC^{+}10, PKD^{+}11, PvdLA^{+}14, RPO^{+}14, Sho11-30, Sho12-47, Sho14-40, Sho14-41, YHG^{+}14, ZLJ^{+}13, FWM^{+}10b]. mitochondrial-focused [HCCS^{+}11]. mitofusin [DPZ^{+}14]. mitofusins [TCX^{+}10]. mitogen [MBR^{+}11]. Mitophagy [ASLS14, LNT^{+}10, MWZ^{+}11, MSS^{+}10, TCX^{+}10]. Mitosis [Les10z, BBW^{+}13, BV11, BMLB^{+}12a, BMLB^{+}12b, BRF^{+}10, COW13, EAB^{+}14, FCE^{+}12, GP12, GP10, GBJ10, HRK13, HBSD12, HTS^{+}10, KNW^{+}14, KSH^{+}13, LR11a, Les13-30, LHD^{+}14, MGT^{+}10, MGS14, MJJ^{+}10, MAE^{+}10, NBSE^{+}13a, NBSE^{+}13b, OB12, PKS^{+}10, RTM13, RKG^{+}10, RMF^{+}10, RDB^{+}12, RDC^{+}11, Sed12v, Sed13k, Sho14w, SSK^{+}14, WBS^{+}12, YFO12]. mitosis-to-interphase [RKG^{+}10]. Mitotic [KPC^{+}10, Les14s, LZLG13, MWG^{+}12, NGM12, SSV^{+}12, WDB10, ANT^{+}12, AYS^{+}13, BRL14, BWS^{+}10, BKP11, CHL^{+}14, CSS^{+}12, CSEH12, FP10, GSJS10, GKA^{+}12, HS10a, HWB^{+}13, KMS10, LC10, LGAC13, LMS10a, Les11t, Les14a, LHS10, MTG^{+}11, MMV^{+}10, NZHL13, PZ14, PJS^{+}11, PL10, RW10, RJM^{+}12, RFVE^{+}10, RGF^{+}10, RY11, SA10b, Sed10h, Sed12a, Sed14i, Sho12-28, Sho12-54, Sho13-31, SHV^{+}11, SHV^{+}13, VSMC11, WGN^{+}13, ZBBG10, ZZW^{+}10, ZSK^{+}13, vZOtR^{+}10]. mixed [Les11-48]. Miz1 [HDH^{+}10, Sho10h]. Mizushima [Sed10m]. MKlp2 [LMS10a]. MKP- [PSVRB^{+}11]. MKS [WLK^{+}11]. MMP [MRCC^{+}13, SGT^{+}13, YZM^{+}12a]. MMPs [LDN^{+}13]. Mnemonic [Les11s]. MOAP [HZT^{+}12]. MOAP-1 [HZT^{+}12]. Mobile [Les11t]. Mobility [WGR^{+}12, dSLPRG11, RZS^{+}14]. mode [HDK^{+}13, LSGVM14, MFF^{+}13, NMB^{+}14, XRO^{+}11]. Model [FBR^{+}10, ANT^{+}12, CPX11, Dun11, FW10, GCR^{+}12, HPB^{+}12, LHN10, SBR^{+}11, SFB^{+}12]. Model-based [FBR^{+}10]. Modeling [SYK^{+}11, SSZ^{+}14]. models [SSW^{+}13, Sed13t]. modern [BWM12]. modes [PGCY12]. modifications [BKAB13]. modified [GCV^{+}11]. modifiers [SWS^{+}11]. modular [Sho14f]. Modularity [KXN10]. modulate [BLM^{+}11, HH14b, NLAS^{+}10, XOV^{+}10, YBN^{+}11]. modulated [HFS10, JGA^{+}11]. modulates [GRHA^{+}12, NSB^{+}11, OLT11, YCP10]. modulating [AAE^{+}14, CZD^{+}13, CDAK10a, CDAK10b, JB12, yLFAM13, SNZVK12, SNZVK13, TOI^{+}13, TGB10]. Modulation [HVOF^{+}14, KAAM11, GBL^{+}11]. modulator [HAKK11, OZ14, YSO^{+}11]. module [FMPS^{+}12, NvCL^{+}13]. modules [MBVT^{+}13, VTO^{+}13, WLK^{+}11]. moesin [Les11t, RDC^{+}11, SMS^{+}13, BWBC^{+}14]. Mogilner [Sed14b].

Molecular [KBS^{+}10, LMA^{+}13, MHS10, RDC^{+}11, BVL^{+}12, CLS13, CBB12, GDS^{+}12, HHC^{+}11, Les14t, LCP13, MS14, RK13, Sed11a, ZSD^{+}14, Jan14, WMB12]. molecule [DMK^{+}12, KSP^{+}11, NRK^{+}13, STD^{+}10, WWS^{+}12]. molecules [SW10b, Sed10l, Sed11s]. moment [Sed12l]. Mónica [Sho10-39]. monitor [HHS^{+}14]. MoniTORing [KK13a]. monocyte [RBH^{+}12]. monomer [KSP^{+}11, OD10]. monomeric [HM10]. monoubiquitination [HLH^{+}14]. Montell [Sed11f]. Moonlighting [MH11]. moonlights [Les14a]. MOR [RKG^{+}10]. Morg1 [HKI^{+}13]. morphogenesis [CWL^{+}11b, EBB13, GOWM12, GWR12, JK10, MH11, MAD^{+}11, PBG^{+}13, QTL^{+}12, QTL^{+}13, RLS^{+}14, RFVE^{+}10, RFAA^{+}12, RSRK13, SMSP11, Sho12-30, VTM14, VvDV^{+}10, VLI^{+}14, ZBJL^{+}10, ZZW^{+}10]. morphogenetic [Les13e, MVR^{+}10, WCQ^{+}13]. morphology [BMLB^{+}12a, BMLB^{+}12b, CLC^{+}11, HVOF^{+}14, LSM^{+}11, PBG^{+}13, Sed12r, Sed13y, Sho12-34, TAC^{+}13, WRF^{+}13]. Morrison [Sho10-56]. mothball [Les11e]. mother [SKN^{+}12, Sho13-43]. mothers [Les12-37]. motif [BKBS12, CLM^{+}10, DCL^{+}12, PPV^{+}14, TMG12]. motifs [BBK^{+}13, DSK^{+}11, MTT^{+}14, PMB^{+}11]. motility [AA13, BMÁG^{+}14, BvMD^{+}14, CBBH11, CGW^{+}11, DS10, DHB^{+}14, GYC^{+}14, GPCK12, GCR^{+}13, HARS14, Kin13, LLU^{+}12a, LLU^{+}12b, LCHB13, MHAK^{+}12, OVW10, dJPAA^{+}11, SST^{+}12, SMM^{+}10, WHWS12, YSaY^{+}13]. motion [SSH^{+}13, WBML11]. motions [WRCD12]. motoneuron [FHD^{+}12, SFB^{+}12]. Motor [DGF^{+}14, Sho14-36, SST^{+}12, CSTBM^{+}10, DCN^{+}10, FHKW11, HBS^{+}10, HZE^{+}13, JDB^{+}12, Les10v, NNO^{+}11, SCL^{+}14, Sho14f, Sed14l]. motor-dependent [CSTBM^{+}10]. Motor-driven [DGF^{+}14, SST^{+}12]. Motoring [Sed10h]. motors [CT10, mFH13, HS10a, KHB^{+}11b, YOMM^{+}11]. mounts [BW13]. mouse [DSB^{+}14, MC10, OHC10, WSZ^{+}12, WPL^{+}11a, WPL^{+}11b, YSN^{+}11]. Move [Fis14, Sed13d, Sed13r, Sho11f, Sho13-44]. movement [FHKW11, HTT13, LBWS10, MAE^{+}10, ZQA^{+}14]. movements [BEJ10, MVR^{+}10, PJS^{+}11, Sho13-60]. moves [EHUD14]. Moving [Sho13-36, Sho12-51, Sho12-62]. MP1 [SSdA^{+}14]. MPS1 [NvCL^{+}13, STD^{+}10, AGM^{+}10, EUB^{+}14, HTS^{+}10, JDS^{+}10, KNW^{+}14, LC10, Les14u, MGT^{+}10]. Mps3 [CSS^{+}14, GSS^{+}11]. MRCK [GdBP^{+}14]. mRNA [BLM^{+}11, COG11, CMW11, CT10, HFB^{+}10, HZE^{+}13, HIB^{+}10, KYOY13, KKK^{+}11, LCfC11, Les11e, RCBY^{+}12, Sho10v, Sho12-46, SHC^{+}13, VBB^{+}10, YCP10, ZDS^{+}12]. mRNA-binding [KKK^{+}11]. mRNA-silencing [BLM^{+}11]. mRNA-transport [HZE^{+}13]. mRNAs [DSD^{+}13, GSM^{+}14, NKH11, Sho10-63, Sho12x, Sho13-29, WP14, YZL^{+}13]. Msb3 [PLR^{+}13]. MT [JK10]. MT1 [MRCC^{+}13, SGT^{+}13, YZM^{+}12a]. MT1-MMP [MRCC^{+}13, SGT^{+}13, YZM^{+}12a]. Mtm [VJK^{+}10a, VJK^{+}10b]. MTOC [BKK^{+}10]. mTOR [BH13, CZM^{+}14, OSD^{+}14, WWT^{+}12]. mTORC1 [PRM^{+}14, Sho11-44, SHC^{+}13, YDB^{+}11]. mTORC2 [Sho14-37, TAC^{+}13]. Mud [Sho11-46]. Mud-mediated [WJPD11]. Müller [Sed14r]. multiciliated [WHH^{+}11]. multifaceted [LZW^{+}10]. multifunctional [PGP14]. multimeric [SRS10]. multimerization [LJW13]. multiorganelle [LBD^{+}14]. Multiple [CT10, GHK^{+}10b, Les11u, LHD^{+}14, MHKM11, WRF^{+}13, DZT^{+}11, HMB14, HMBC10, RDB^{+}12, SNR^{+}11, SLC^{+}13, Ste12]. multiprotein [JKS14, LvBG^{+}10, NSD^{+}14]. multiscale [FW10, Sed13y]. Multispan [PKD^{+}11]. multispanning [BWK^{+}11]. multitude [Sho10-44]. Multivariate [BAH^{+}12]. multivesicular [DGS^{+}11]. mum [Sho13-43]. Munc13 [BJE^{+}12, vdBFS^{+}12]. Munc13-4 [BJE^{+}12]. Munc18 [BJ12, CWB^{+}14, SRKR10]. Munc18-1 [CWB^{+}14, SRKR10]. Munc18c [JOR^{+}11]. Munro [Sho11-51]. murine [BLT^{+}11]. Murphy [Sed12e]. Mus81 [DKMK^{+}11, RZF^{+}11]. Mus81-Eme1 [DKMK^{+}11]. Mus81-mediated [RZF^{+}11]. muscle [BLC^{+}12, BvMD^{+}14, CTL^{+}10, CYN^{+}13, CZGG12, CLZ^{+}14, CCM^{+}11, DZT^{+}11, ELH14, ERS10, GSB^{+}13, GLM^{+}10, GF11, ILD^{+}10, LGM^{+}12, Les10-36, LT11, LXTM12, MHC^{+}12, MBK^{+}10, MH14, OZT^{+}13, PGB^{+}10, dJPAA^{+}11, QECC10, RMM^{+}10, RCBY^{+}12, RSRK13, SSB^{+}10, Sho10u, Sho12t, Sho12r, Sho14k, TPZ^{+}14, VGL^{+}14, VLI^{+}14, WWT^{+}12, WCQ^{+}13, ZWL^{+}14]. muscle-building [Sho14k]. muscle-specific [ILD^{+}10]. muscles [ETYS^{+}12, MMVK^{+}12, OBC14]. muscular [APV^{+}12, CG10a, DWJ^{+}14, RK13, RHKB12, SYK^{+}11]. mutagenesis [ZLFC14]. Mutant [ETI^{+}10, Sho11-35, DS12, Sho11-45, XHS^{+}13, ZYH^{+}11]. mutants [MVN11, Sho12-63]. mutase [MMO^{+}14]. mutation [GKWG^{+}11, GZZ^{+}14]. mutations [CWS^{+}11, KSB^{+}13, LNT^{+}10, LgLM^{+}10, PL11, PDKG14]. Mutual [DWM^{+}12, KMS10, Sho14-38]. MVB [AVP^{+}14, DCL^{+}12]. MVBs [SP11]. myc [MLH12, AIJI11]. myelin [LCfC11, PSF^{+}11, SMK14]. myelin-dependent [PSF^{+}11]. Myelinated [Sho12-35, IHG^{+}12, ZSZ^{+}13]. myelination [HGV^{+}14, NBS^{+}11]. myeloperoxidase [PMHZ10]. Myo1 [Sho10-40]. Myo1c [FEHF12]. Myo2 [FHKW11]. Myo4 [CT10]. Myo4p [HFB^{+}10, Sho11t]. Myo51 [WLZ^{+}14]. myoblast [DJL^{+}12, LMS^{+}13, PBD^{+}13, SZJ^{+}10, TNH^{+}11]. myoblasts [HVW^{+}10, SCN^{+}14]. myocardin [MH14]. myocytes [KUH^{+}14, SZW^{+}11]. MyoD [HVW^{+}10, ZSH10]. myofibril [ILD^{+}10, TNH^{+}11]. myofibrils [FS14]. myofibroblast [MSC^{+}10]. myogenesis [BWL^{+}13, GSC11, PCO^{+}10, SGD^{+}10]. myogenic [BvMD^{+}14, CZ10, ZSH10]. myogenin [TPZ^{+}14]. Myosin [CWL^{+}11a, LCS^{+}10, PSR^{+}10, Sho10-41, Sho11-36, TPM^{+}13, VMNLB^{+}11, AFM^{+}13, DWPC^{+}11, EJBW12, FLN^{+}10, FLN^{+}16, FHKW11, KHB^{+}11b, PSK11, VTM14, VLI^{+}14, WLZ^{+}14, YZPF12, lDSB^{+}10a, lDSB^{+}10b, SNZVK12, SNZVK13, CHL^{+}14, CWZ^{+}12, NBDB12, RSD^{+}12, WVT^{+}13]. myosin-related [FHKW11]. myosins [SFL12]. myostatin [WWT^{+}12]. MYPT1 [CSS^{+}12]. mystery [Sho12-36]. Myt1 [OHC10].

N [AGM^{+}10, Ari10, Les11v, Les12u, MDP^{+}10, NBS^{+}11, RSB13, SRZ^{+}11, Mit12c, BBY^{+}12, BV11, BEJ10, CSG14, EBBJ11, HS10b, NvCL^{+}13, OLB13, RBM^{+}11, SRKR10, WWB^{+}10, YZM^{+}12a, ŻKC^{+}11]. N-cadherin [Ari10, MDP^{+}10, SRZ^{+}11, BEJ10, EBBJ11]. N-end [BV11]. N-glycan [HS10b]. N-glycosylation [CSG14]. N-peptide [SRKR10]. n-Syb [Mit12c]. N-terminal [AGM^{+}10, BBY^{+}12, NvCL^{+}13, RBM^{+}11, ŻKC^{+}11]. N-terminally [WWB^{+}10]. N-WASP [Les11v, Les12u, NBS^{+}11, YZM^{+}12a]. NAC [KPI^{+}10]. Nachury [Sho10-36]. NAD [HA12]. NADPH [LSOT10]. nanocrystals [Les13-36, NRK^{+}13]. nanodomain [SSL^{+}14]. nanodomain-associated [SSL^{+}14]. nanoparticles [RBS10]. nanoscale [AFRZ^{+}14]. nanoscopy [FAvdB^{+}12]. nanoterritories [SHS^{+}13]. nanotopography [KPSL12]. nascent [DDH^{+}12, KPI^{+}10, LLU^{+}12a, LLU^{+}12b, LJLJ11, RKK^{+}14]. natalizumab [Ste12]. Näthke [Sed10g]. native [FBZM^{+}10]. Natural [Les11w, AMH11, SNSyN13, Sho14p]. Nature [Sed14v]. Nck [Les12v, DMH^{+}12]. Nck-dependent [DMH^{+}12]. NCS [WOG13]. NCS-1 [WOG13]. Ndc1 [CSS^{+}14, Les14v]. Ndc80 [LHW10, LMA^{+}13, TUG^{+}10]. Ndel1 [ŻKC^{+}11]. Ndfip1 [HLL^{+}12]. NDPK [SWS^{+}11]. Nebulin [PKG10, Sho10-42, FS14]. necessary [ADAB^{+}12, CSG14]. neck [AVP^{+}14]. necroptosis [BMS^{+}11]. necrosis [MBO^{+}14, ZFA^{+}13]. need [Les13w, Sho12-35]. needed [CSM^{+}12, GSGL11]. negative [CLC^{+}11, Les10u, MVP^{+}10, PLL^{+}12, XBC^{+}13]. negatively [ATU^{+}12, BB10, GDO13, HBM^{+}11, NB12, RDB^{+}12]. neighbor [Nan14]. neighborhood [Sho13w]. neighboring [SKN^{+}13]. Neighborly [SY10]. Neighbors [Les10-27, Sho12-31]. neither [DHVK10a, DHVK10b]. NEMO [Les14w, TTC^{+}14]. nephronectin [KTN^{+}12]. NER [OMV^{+}11]. nerve [CWB^{+}14]. nerves [IHG^{+}12]. nervous [CVJ^{+}11, Coo13, HGV^{+}14]. nesprin [RMS^{+}14]. nesprin-1 [RMS^{+}14]. Nessun [MZP^{+}10]. net [Sho10l]. Net1 [KST^{+}11]. Netrin [Sho14-39, SCR12, HZM^{+}13, WLN^{+}14, WMV^{+}14]. Netrin-1 [WMV^{+}14]. NETs [Les10h]. network [ARF10, AiIK^{+}13, BPH^{+}14, BBD^{+}11, CMS11, CJNS12, GHK10a, IHM13, KMS10, SB14, Sho13a, SKFH11, WVvG^{+}13, WRF^{+}13, vGCMA^{+}14, vBAK^{+}12]. networks [LhYL^{+}13, RDC^{+}11, YZPF12]. Neural [ZC11, BMRM13, CF13, GCP^{+}14, LTJN^{+}12, Les13-37, MMdCOM^{+}11, RSB13, Sho13x, SMB12, VLG14, YSN^{+}11]. neuralgic [BBK^{+}13]. Neuralized [DSK^{+}11]. Neuregulin [SAG^{+}11]. Neuregulin/ [SAG^{+}11]. Neuregulins [Les11x]. neurite [GR11, LLcK^{+}11, PSF^{+}11, TKMK10]. Neurobeachin [NLJ^{+}13]. neuroblast [CMS11]. neuroblasts [JG10]. neurodegeneration [DD10b, FUK^{+}14, HWE^{+}12, JCL^{+}11, LgLM^{+}10, MVC^{+}11, RCFH10]. neuroepithelial [PJS^{+}11]. neuroepithelium [SMdP^{+}14]. Neurog3 [GvEM^{+}11]. neurogenesis [SLS^{+}10]. neuroligin [WKN^{+}13]. neuroligins [PYT^{+}13, KSLF^{+}11]. Neuroligins/ [KSLF^{+}11]. neurological [XHS^{+}13]. neuromuscular [BLT^{+}11, KUN^{+}13, Les11x, SAG^{+}11, SBS^{+}12]. neuron [CPX11, CF13, TAC^{+}13]. Neuronal [EMT^{+}14, IIWS14, ASLS14, CWL^{+}11b, CLC^{+}11, GWP^{+}11, HWE^{+}12, HLL^{+}12, HRWW^{+}13, KK13a, KNsMK13, MPRT11, Sho11q, TNV^{+}13, WGM^{+}12, ZZW^{+}13, Sed10b]. neurons [DNB13, DMD^{+}12, DHVK10a, DHVK10b, KZR^{+}12, KSLF^{+}11, Les13v, MWH12, MTT^{+}14, She14, Sho13k, WOG13, vdBFS^{+}12]. neuropathology [WTH^{+}11]. neuropeptide [CG12a]. Neuropilin [XYM^{+}10]. Neuropilin-2 [XYM^{+}10]. neurotransmitter [KBAW^{+}11, Les11-35, NLJ^{+}13, VM14, WEK^{+}14]. neurotrophin [BMG14, LLcK^{+}11]. neurotrophin-induced [LLcK^{+}11]. neurturin [BST^{+}11]. neurulation [BEJ10]. neutral [KPC^{+}11]. neutralization [CZC^{+}11]. Neutrophil [BZ12, PMHZ10, CBBH11, LOR^{+}10, TSB^{+}14]. neutrophils [Sho14-32]. newborn [Les13m]. Nexins [Sed14t]. NF [BLC^{+}12, BWL^{+}13, Les12-29, Les14w, LDN^{+}13, WHA^{+}13]. NF- [BLC^{+}12, BWL^{+}13, Les12-29, Les14w, LDN^{+}13, WHA^{+}13]. NHE [BWBC^{+}14]. niche [LWW12]. Nicklas [Sed12c]. nitration [SKV^{+}11]. nitric [PBD^{+}13, WHF^{+}11]. nitrosylation [LKLA12]. Nkx2 [QWL^{+}11]. Nkx2-5 [QWL^{+}11]. Nmd3 [Oef10, SBP^{+}10a]. NMDA [BLM^{+}11, HSS^{+}13]. No [Les12w, Les10-39, Les12c, Les11f, Sho10-60, Les12w, Ros13, Sed12g]. Nobel [ME13]. Noboru [Sed10m]. Nodal [WHWS12]. nodes [ABP^{+}14, LCLW11, RBB^{+}14]. Nogo [JDHS10]. noise [SBR^{+}11]. nomenclature [PvdLA^{+}14]. non [VLI^{+}14]. non-muscle [VLI^{+}14]. nonautonomous [WHA^{+}13]. noncanonical [MTT^{+}14, NSB^{+}11]. noncatalytic [CFB^{+}13]. Noncentrosomal [SFL12]. nonchondrocytic [BPMK^{+}14]. nonciliary [SW10b]. Noncoding [Les11z, ZZW^{+}14]. nonhomologous [CFB^{+}13, CWG^{+}11]. nonlinear [SHV^{+}13]. Nonmedially [HHY^{+}12]. Nonpolarized [PGCY12]. nonproliferation [Sho13-34]. Nonrandom [Yam13, ECC^{+}13]. nonsense [VBB^{+}10]. nonsense-mediated [VBB^{+}10]. noodle [Swa13]. nor [DHVK10a, DHVK10b]. normal [DWL^{+}11, GBiY^{+}14, YSaY^{+}13]. Notch [DYI^{+}13, HFS10, VES^{+}11, CTM^{+}14b, DCO^{+}12, DCO^{+}16, HSKAT11, PCO^{+}10, Sho10z, Sho10-65, Sho13z, UKZ^{+}13, VWD^{+}13, YMU^{+}10, YMU^{+}13]. Notch3 [LRH^{+}13]. notes [Sho12-48]. Notochord [EBB13, Sho13-37]. Novel [BBK^{+}13, ZGEM12, ANT^{+}12, APV^{+}12, ABP^{+}12, BST^{+}11, Bez12, BAH^{+}12, BGS13a, BGS13b, BMC^{+}11, BKBS12, CHS^{+}10, CSP^{+}10, CB12, CGW^{+}11, DPV^{+}12, EZT^{+}12, EBBJ11, ECJB10, FRS^{+}13, HDK^{+}13, HMB14, HTT^{+}11a, IKU^{+}11, JDB^{+}12, KWDD10, KBAW^{+}11, LGAC13, LH11, LCK^{+}13, MZP^{+}10, NSBW10, PDMBW11, RCC^{+}12, SSW^{+}13, SYV14, SRZ^{+}11, TSL12, WLZ^{+}14, WMB^{+}10, WMV^{+}14, YSN^{+}11, YSaY^{+}13, ZQA^{+}14, ZSK^{+}13]. novo [BKS^{+}11, MJJ^{+}10, UG10]. NPC [BBD^{+}11, TH11]. NPC-anchored [BBD^{+}11]. NPF [GJP^{+}13]. NPHP [Omr10, WLK^{+}11]. NPHP4 [HBM^{+}11]. NPLY [PPV^{+}14]. Nrf2 [IWS^{+}11, RKS^{+}10]. Nrf2-mediated [RKS^{+}10]. NRSF [PMP^{+}11a, PMP^{+}11b]. Nsk1 [CLO^{+}11]. Nsp1 [CWFL13]. Nsp1p [MLW13]. nuage [HLS^{+}14]. Nuclear [BMLB^{+}12a, BMLB^{+}12b, DDH^{+}12, HSI^{+}14, JCL^{+}11, LMC^{+}12, Sho11-39, BKAB13, BBD^{+}11, BDN^{+}13, CTM^{+}14a, CHS^{+}10, CSS^{+}14, CAB^{+}13, CG12b, CWFL13, CSH^{+}12, DE10, Dun11, EL14, FMPS^{+}12, FTJG13, wFLW^{+}13, FS10, GSM^{+}14, GSS^{+}11, GC13, HH14a, HBSD12, HZW^{+}12, HLL^{+}12, HPB10, KST^{+}11, LCS^{+}13, Les10-28, Les10-33, Les11u, LSW^{+}14, LLK11, MMU10a, MLW13, MMC^{+}10, MPM11, NNSH11, RG11, RN12, RKW^{+}13, SFJ^{+}14, SRBL13, Sed12r, SBP^{+}10a, Sho10-35, Sho11r, Sho13-34, Sho13-49, Sho14q, Sho14-59, SDD^{+}13, SST^{+}12, SKFH11, TGG^{+}11, TGES12, TESA10, TLTW10, TCB^{+}14, UHKS11, WGN^{+}13, WMCF10, WD11, WtLK^{+}13, ZDS^{+}12, ZGEM12, ZGW^{+}14, ZKR^{+}11]. nuclear-encoded [GSM^{+}14]. Nuclear-localized [LMC^{+}12]. nucleated [OZT^{+}13]. nucleates [OBD^{+}10]. nucleation [CLS^{+}10, ENG^{+}12, GJP^{+}13, KOK^{+}13, LSM^{+}11]. nucleators [Les13-42]. nuclei [FS10, Les12b]. nucleic [SJM^{+}13]. nucleocytoplasmic [SST^{+}12]. nucleolus [BVR11]. Nucleophosmin [RTC^{+}12, LKLA12]. Nucleophosmin/ [RTC^{+}12]. nucleoporin [CHS^{+}10, GVP^{+}11, SKFH11, TESA10]. nucleoporins [KAAM11]. nucleosome [LBS11, XSJ^{+}10]. nucleosomes [CMS10]. Nucleostemin [HLT12, QB12]. Nucleotide [MHV12, DWL^{+}11, GEN14, LCS^{+}10, PVM^{+}12]. nucleus [GP10, LCS^{+}13, Lin10, OVW10, Sho12-62]. nudge [Les12-32]. null [LZR^{+}11]. Num1 [TSL12]. NuMA [LDCF^{+}13, MdFF^{+}14, PJS^{+}11]. Numb [KNsMK13]. number [CG10a, JJH^{+}10]. numbers [CF13, Les12v]. Nup133 [BBD^{+}11]. Nup133-dependent [BBD^{+}11]. Nup188 [TESA10]. Nup358 [HHJ^{+}11]. NuRD [SWV^{+}10]. nurse [NSS^{+}10]. Nutrient [OS13, CGCP^{+}14]. Nutrient-sensing [OS13]. nutritional [Sho11-44].

O [GCSB10, HTS^{+}10, Sho10-62]. O-glycosylation [GCSB10, Sho10-62]. OATL1 [IKU^{+}11]. obligate [PTS^{+}10]. Obscurin [RGL^{+}13]. observed [YON^{+}12]. Observing [CLEZ12, Sed13g]. Occludin [RBY^{+}11, Les10l, MLG^{+}10]. occupies [KFET11]. occur [FWM^{+}10b, WAG^{+}10]. occurs [ABVP11, ASLS14, JTN^{+}13, MDW^{+}13, WM12, WDB10]. octamers [KFET11]. Odf2 [TYN^{+}13]. off [BJ12, DSMB13, Les10r, Les14u, Pal14, Sho11-35, Sho12-39, Sho14-43, Sho14-57, BT13]. Ohsumi [Sed12z]. old [Sed14x, Sho10-43, VM14]. oligodendrocytes [LCfC11]. oligomer [OD10]. oligomeric [RZS^{+}14, YSN^{+}10]. oligomerization [CRP^{+}14, DSM^{+}11, IM11]. oligomers [DBUT13, Sho10g]. OMA1 [AWB^{+}14]. Omi [YHG^{+}14]. on-off [Sho14-57]. ON/OFF [BT13]. oncogene [VBB^{+}10]. Oncogenes [NZHL13, Sho13-38, MLH12]. oncogenesis [LSS^{+}12]. oncoprotein [PRM^{+}14]. One [Les11-27, Les11-44, PR12, Sho11t, Sho13-58, Bab14]. Only [Les12x, CTM^{+}14a, LBS^{+}13, PR12, Sho12-32, CTW^{+}10]. onset [JDS^{+}10, KNW^{+}14, KSSD11, LR11a, LWK^{+}13, LS13b]. onto [KSS^{+}11, Sho13p, Sho13y]. oocyte [OHC10, PHB^{+}11, SLM^{+}13]. oocytes [Sho13u, YRU^{+}13]. oogenesis [MRLLS12, MHCvSW11, NSS^{+}10]. OPA1 [AWB^{+}14, Sho14-41]. Open [PASG^{+}12, Les12u, Sho14-62]. Opening [Sho12-37, ZZS13]. opens [Sed14m, Sho14h, Sho14-54]. operate [Sho11-42]. operated [SPC^{+}13, YSM10]. operation [KdKDP12]. opposes [EUB^{+}14, LVB^{+}10]. Opposing [EKJH13, KLP^{+}14b, BCB14a, NGL^{+}12]. opposite [YOMM^{+}11]. OPT [HBC^{+}11]. optimize [DWPC^{+}11]. opts [RMG^{+}12]. Orai1 [JAM^{+}13, SLH^{+}14, YSM10]. Orai1-mediated [SLH^{+}14]. Orc6 [BV11]. orchestrate [KAS^{+}12, KLP^{+}14b]. orchestrates [SLH^{+}14]. orchestrators [SWF12]. order [LHD^{+}14, RG14, RG11, Sho14w, SMZL13]. ordered [ZNA^{+}14]. orderly [BB10]. orders [Sho13-35]. organ [UKZ^{+}13]. organellar [Sed11d, TOI^{+}13]. Organelle [ETYS^{+}12, OB12, ETRP12, Pri14, Sed12f]. Organelles [Sho12-38, Bra13, EBB13, HKN^{+}11, MDW^{+}13]. Organellophagy [Oka14]. organization [AGL^{+}14, AFM^{+}13, AFRZ^{+}14, BDR^{+}10, BG10, BHB^{+}11, BRP14, CB12, DBH^{+}11, HS10a, HCCS^{+}11, KCF^{+}14, LHN10, LhYL^{+}13, RBF^{+}12, SCL11, TPM^{+}12, VGL^{+}14, ZSZ^{+}13, ZF11]. organizational [Sho10p]. organize [GBK^{+}14, HC10, Les13-41]. organized [Les11m, WZHV11]. organizer [Leu14]. organizes [HARS14, IHG^{+}12, SST^{+}12]. organizing [PvdLA^{+}14, YKT^{+}13]. organogenesis [LSS^{+}12]. orient [WLN^{+}14]. orientation [CMS11, DHL^{+}12, JG10, MdFF^{+}14, NSZ^{+}13, PHB^{+}11, PL10, QMHM10, RFVE^{+}10, SMdP^{+}14, Sho14-66, WJPD11, ZZW^{+}10, ZSK^{+}13, NGM12]. origin [GB10, GZZ^{+}14, Sho14-50, ZNP^{+}13]. origins [CSAPLBD11a, CSAPLBD11b, SZE^{+}11]. Orna [Sed12r]. orthologue [AHL^{+}11, KKUG11]. OS-9 [BGC^{+}10]. OS-9/XTP3-B [BGC^{+}10]. oscillation [SLH^{+}14]. Oscillations [MMV^{+}10, CSHS^{+}13a, CSHS^{+}13b, RZA^{+}13]. oscillatory [WLN^{+}14]. Osh4p [Lev11, dSJDD^{+}11]. oskar [GYC^{+}14]. Osmotic [GEN14, Sho14-42, SRBL13]. OST [STG13]. Osteoblast [BMFC^{+}11, BMFC^{+}13, BBJ^{+}10, IIN^{+}11, ISZ^{+}11, KHB^{+}11a, KKY^{+}14, SJZ^{+}10, WSZ^{+}12]. Osteoblasts [Sho11-40, OMZK14, Sho12-33]. osteoclast [KHB^{+}11a, KPH^{+}12, SNR^{+}11, Sho12-42]. osteoclast-mediated [SNR^{+}11]. osteoclastogenesis [AKB^{+}13, XTX^{+}13]. osteoclasts [KKY^{+}14, SCN^{+}14]. osteopenic [KHB^{+}11a]. osteoporosis [XTX^{+}13]. Osteopotentia [SJZ^{+}10]. osteoprotegerin [AKB^{+}13]. other [MJFS10]. Otoferlin [JC10]. Ottoline [Sed14s]. out-of-equilibrium [CPS^{+}13]. outer [BWK^{+}11, CDD13, KIL^{+}12, LLH13, LVB^{+}10, PKD^{+}11, PPD^{+}10]. Outgrowing [Les12y]. outgrowth [GR11, LLcK^{+}11, PSF^{+}11]. outs [Les10t, Sed11m]. Ovastacin [BXB^{+}12, Sho12-39]. Overcoming [Sed11t]. overduplication [HLN^{+}10]. Overexpression [vRJMvD10, RJvD11]. Overholtzer [Sed11l]. Overlap [EJBW12]. overlapping [KKL^{+}14]. own [HLN^{+}10, Les14p, Lin10, Sed12f, Sho11h, Sho14-27]. oxidase [CHL12, LSOT10, ZCB^{+}10a, ZCB^{+}10b]. oxidase-dependent [CHL12]. Oxidative [SFJ^{+}14, SSB^{+}10, BLC^{+}12, KZR^{+}12, LSOT10, MHV12, PXZ^{+}13, Sev10]. oxide [PBD^{+}13, WHF^{+}11]. Oxidoreductase [CSG14]. oxidoreductases [AiIK^{+}13]. oxygen [Fin11, KLvdB^{+}13, Sho13h]. oxysterol [DKF^{+}11]. oxysterol-binding [DKF^{+}11].

P [KPJ^{+}13, Sho12-40, UHKS11, vGLWB12, BHA^{+}12, COG11, CLSO^{+}12, RMS^{+}14, RBM^{+}11, Sho14-47, VJK^{+}10a, VJK^{+}10b]. P-bodies [COG11]. P-body [RMS^{+}14]. P-site [BHA^{+}12]. p107 [SSB^{+}10]. p110 [CVR10, DCP^{+}10, YYM^{+}11]. p110- [DCP^{+}10]. p114RhoGEF [NT11]. p120 [NCML^{+}12, PMF12, SME^{+}13, Sho13-39]. p120-catenin [NCML^{+}12, PMF12, SME^{+}13]. p120RasGAP [MVP^{+}11]. p125A [OTLH10]. p130 [SZW^{+}11]. p130Cas [WM12]. p14 [SSdA^{+}14]. p150 [RMF^{+}10]. p150CAF [BSR^{+}11a]. p150CAF-1 [BSR^{+}11a]. P150Glued [RTC^{+}13a, RTC^{+}13b]. p190RhoGAP [SKV^{+}11]. p24 [FWJ^{+}11]. p25 [ZYF^{+}11]. p37 [KSH^{+}13]. p38 [PSVRB^{+}11, Sed10i]. p38/MKP [PSVRB^{+}11]. p38/MKP- [PSVRB^{+}11]. p400 [CCJ^{+}12, XSJ^{+}10]. p47 [KSH^{+}13, KBW^{+}10]. p53 [BVR11, CMD^{+}13, DKM^{+}13, HLN^{+}11, KSB^{+}13, Les13y, LSW^{+}14, MFGB10, MVN11, Sed12m, Sho10-44, Sho10-45, Sho14-43, TAGJ11, TC10, YHG^{+}14, KKL^{+}11]. p53-dependent [DKM^{+}13, KSB^{+}13, TC10]. p53-induced [LSW^{+}14, MFGB10]. p53-mediated [YHG^{+}14]. p53/miRNA [KKL^{+}11]. p53/miRNA-34 [KKL^{+}11]. p62 [IWS^{+}11, IM11]. p63 [FMG^{+}11, Les11-28]. p73 [Sho14-44]. p97 [KBW^{+}10, TCX^{+}10, KSH^{+}13]. PAFAH [BDR^{+}10]. Pah1p [AHL^{+}11]. PAI [APV^{+}12]. PAI- [APV^{+}12]. pair [GJP^{+}13, Les12d]. paired [DBH^{+}11]. pairing [WRCD12]. pairs [Ewe11]. PAK [BRF^{+}10]. PAK-mediated [BRF^{+}10]. PAK1 [BPB^{+}12, DWPC^{+}11]. PAKs [DJL^{+}12]. Palade [NWD^{+}11]. PALB2 [MFR^{+}14]. palmitoyl [NOT^{+}14]. Palmitoylated [MMU^{+}10b]. Palmitoylation [FCA10, FDB^{+}13, HAB14, SGC10]. Palmitoylation-dependent [FCA10]. Pamela [Sed12s]. Pan [Les10-29]. pancreas [ZCB^{+}10a, ZCB^{+}10b]. pancreas-specific [ZCB^{+}10a, ZCB^{+}10b]. pancreatic [DSMB13, LDN^{+}13]. Pannexin [IIN^{+}11, VBG^{+}13]. Panta [PM13]. Paolo [Pow10]. Papa [Sed14h]. PAPC [KBW^{+}12]. PAR-1-dependent [PHB^{+}11]. PAR-4 [BDC^{+}14]. PAR-4/ [BDC^{+}14]. PAR1 [DCL^{+}12]. Par1b [CFLDM11, LDCF^{+}13]. Par6 [HKI^{+}13]. paracrine [FHD^{+}12]. parasite [CBB12, Les14q]. Paraspeckles [Les11-30, NNSH11]. Parent [Sed14e]. Parkin [Sho10-46, ASLS14, KLF^{+}14, LNJ^{+}13, LNT^{+}10, MSS^{+}10, TCX^{+}10]. Parkinson [CPX11, WCC^{+}10]. PARL [JLW^{+}10]. PARP [CWG^{+}11, Sho11-41]. PARP1 [PVM^{+}12]. part [OTLH10]. participates [DZT^{+}11]. particle [LAR^{+}10]. partner [GSM^{+}12, MZP^{+}10, SFJ^{+}14]. partnership [SC10b]. party [SF12]. Pascale [Sed11m]. pass [GGR12]. passage [Les10-33, TESA10]. passenger [LCD^{+}11, RHK11, XOV^{+}10]. Passengers [Sho11-42]. Passionate [Pow14b]. past [SNSyN13]. pastoris [MALS10]. patch [TSL12]. Patched [Les11j]. path [SHC^{+}10, Sed12y, Sho14-44]. pathogen [AMO^{+}11, RGB^{+}13]. pathogen-containing [RGB^{+}13]. pathogenesis [LKSG13]. Pathogens [HW11, AMR11, KWO11]. pathology [DSP11]. paths [Sho10-31, Sho12-30]. pathway [AAE^{+}14, APV^{+}12, ASE10, ACO12, CDK^{+}10, CG12a, CZM^{+}14, FWM^{+}10a, GFSR11, HBM^{+}11, JDL^{+}14, LBWS10, LCK^{+}13, LN14, LBD^{+}14, LKG^{+}13, MLM^{+}13, MLM^{+}11, MPD^{+}12, NSSF10, OD10, PKD^{+}11, PBD^{+}13, PDMBW11, RKS^{+}10, RBA^{+}11, Sho14u, SEV^{+}14, VES^{+}11, WMB12, YHF13, YYM^{+}11, YDB^{+}11, ZIG^{+}12]. pathways [AMR11, BPB^{+}12, HGV^{+}14, HMBC10, HK14, KKMB10, KdKDP12, LSS^{+}12, LAB14, MJFS10, MWZ^{+}11, MHKM11, MMB^{+}11, MAD^{+}11, NSB^{+}11, OS13, dJPAA^{+}11, RKG^{+}10, RKT^{+}14, Sed12w, SHBC12, TTM^{+}14, ZDM^{+}14]. PATJ [SNZVK12, SNZVK13]. Patronin [WBMCSS13]. pattern [RCC^{+}12]. Pax3 [HVW^{+}10]. Pax7 [CTL^{+}10]. Paxillin [DT14, Les14y, PSR^{+}10]. Pcdp1 [DS10]. PCNA [GHK10a, GL10, HLH^{+}14, JEF^{+}11, yLFAM13, MFA^{+}14, NB10]. PCNA-coupled [JEF^{+}11]. PCP [GOWM12]. PCP-mediated [GOWM12]. PDE [WOG13]. PDE-2 [WOG13]. PDE4D3 [TMS^{+}12]. PDIs [AiIK^{+}13]. PDK1 [CLD11, GdBP^{+}14]. PDK1-mediated [GdBP^{+}14]. PDZ [GB12, LMT^{+}10, LLcK^{+}11]. PDZ-directed [LMT^{+}10]. PDZ-RhoGEF [LLcK^{+}11]. peace [Les14y]. pearance [Sho14-66]. Pellman [Sed10d]. pellucida [ABD14, BXB^{+}12]. pellucidae [BBY^{+}12]. peptidase [BBW^{+}14]. peptide [BBW^{+}14, DMD^{+}12, SRKR10]. performance [Sho10-59, Sho11-27]. performed [DV10]. peri [Sed13u]. Pericentric [SHV^{+}13, BAS^{+}14, SSK^{+}14]. Pericentrin [DD10a, LR11a]. Pericentromere [CMS^{+}14]. pericentromeres [SSH^{+}13]. pericyte [RB11]. perinuclear [TIM14]. peripheral [HGV^{+}14, IHG^{+}12]. periphery [FTJG13]. PERKs [Sho12o]. Perlecan [KUN^{+}13]. permeability [Les11f, OGD^{+}12, SKV^{+}11, VBG^{+}13]. permit [YFO12]. permits [BKK^{+}10]. Peroxisomal [NOT^{+}14, MALS10]. Peroxisome [MAS11, SKVvdK11]. peroxisomes [YHF13]. Persistent [IWS^{+}11, SMM^{+}10]. persistently [DHVK10a, DHVK10b]. perspective [SMK14]. perspectives [Sed14j]. Peter [Les10-29, Sed11n, Sed14t]. Peterson [Sed13s]. PEX19 [YHF13]. PEX19-dependent [YHF13]. Pex25 [SKVvdK11]. PEX26 [Sho13-40, YHF13]. Pex3 [KMC^{+}14]. pexophagy [MJFS10, NOT^{+}14]. PGC [BLC^{+}12, SSB^{+}10]. PGC-1 [BLC^{+}12, SSB^{+}10]. PGL [HYS11]. pH [CWC^{+}13, KWH^{+}10a, KWH^{+}10b, MLM^{+}11, PHB^{+}13, PGAE^{+}13, SRCP14, Sho13-30]. pH-dependent [MLM^{+}11]. pH-sensitive [SRCP14]. phagocytic [FHY^{+}10]. phagocytosis [MI13b, Sed12t]. phagophore [NOT^{+}14]. phagosome [ATKK11, PBG^{+}13]. phagosomes [BCBG10, RGB^{+}13, Sho10-48, Sho11-35]. Phase [Bra13, BNL^{+}10, COW13, JEF^{+}11, LBS^{+}13, LLM^{+}10, MGT^{+}10, MGS14, OYH13, TW14, WMC14, PLC^{+}11, YFO12]. phases [ABVP11, KLvdB^{+}13]. PHD12 [SMB12]. phenotype [DBUT13, ERS10, NSB^{+}11, RSL^{+}11, SZW^{+}11]. phenotypes [DKM^{+}13, LDCF^{+}13]. Philip [Sab10]. Philipp [Sed13d]. pHocal [LS13a]. phosphatase [ALS^{+}13, FWM^{+}10a, LVB^{+}10, NPL^{+}10, OMW^{+}14, RGF^{+}10, Sho14-68, ZBBG10, BKBR11]. phosphate [GFSR11, HMB14, VSG^{+}12, dSJDD^{+}11]. Phosphatidic [KYP^{+}14]. phosphatidylinositol [AKA^{+}13, DCP^{+}10, HMB14, MRPR12, dSJDD^{+}11]. Phosphatidylserine [HAKK11, XBC^{+}13, FSA^{+}11]. phospho [MGK^{+}12]. phospho-switch [MGK^{+}12]. Phosphocaveolin [JBS^{+}12, JBS^{+}13]. Phosphocaveolin-1 [JBS^{+}12, JBS^{+}13]. phosphocycling [VOSB12]. phosphoglycerate [MMO^{+}14]. Phosphoinositide [LMS^{+}10c, YYM^{+}11, BCBG10, CDH^{+}14, HTT^{+}11a, HAB14, RLS^{+}14, TOI^{+}13, VSG^{+}12]. phosphoinositide-binding [HTT^{+}11a, TOI^{+}13]. phospholipase [BDR^{+}10, KYP^{+}14, LBS^{+}13, MRPR12, YDB^{+}11]. phospholipid [Les13l]. phospholipids [OVL11]. phosphoregulates [RSD^{+}12]. phosphoregulation [TOI^{+}13, ZSD^{+}14]. Phosphorylated [DPB^{+}10, MFR^{+}14, LCBG^{+}11]. phosphorylates [CSS^{+}12, IMP^{+}12, KLF^{+}14, RBM^{+}11]. phosphorylating [TTM^{+}14]. Phosphorylation [CLM^{+}10, GSJS10, HDK^{+}13, PAB^{+}11, Sho13-41, ALS^{+}13, BPT^{+}14, BRF^{+}10, BVM^{+}11, CDD13, CLO^{+}11, CKU^{+}10, EUB^{+}14, FPAM13, GZR^{+}14a, GZR^{+}14b, GLB10, GC13, GSGL11, HSS^{+}13, HRWW^{+}13, ILD^{+}10, JOR^{+}11, KNW^{+}14, KST^{+}11, KLS^{+}13, LR11a, LZLG13, MLM^{+}11, NCT^{+}11, NKH11, PSR^{+}10, RBY^{+}11, RTC^{+}12, RTC^{+}13a, RTC^{+}13b, RMF^{+}10, SAG^{+}11, VYC^{+}11, VTM14, WBL11, WDG^{+}13, XHS^{+}13, YMT^{+}13, YCP10, ZBBG10, ZJP^{+}10]. phosphorylation-dependent [GLB10]. photoinactivation [KKS^{+}14]. photoreceptor [CHL12, MBLD11, SDS^{+}12a]. photoreceptors [VPC^{+}14, WWHH10, YBN^{+}11]. pHuji [SRCP14]. Physical [WtLK^{+}13]. physics [Sed13o, Sed14q]. physiological [YHT^{+}10]. physiology [GLM^{+}10, RH10, SMSP11]. PI [HGV^{+}14, Sho12-40, VJK^{+}10a, VJK^{+}10b, YDB^{+}11, vGLWB12]. PI-3-kinase [YDB^{+}11]. PI3 [MMS^{+}10]. PI3-kinase [MMS^{+}10]. PI3K [CVR10, CLZ^{+}14, VJK^{+}10a, VJK^{+}10b, WAJ^{+}12]. PI3K-dependent [WAJ^{+}12]. PI3P [PAB^{+}10]. PI4KIII [NBC^{+}12]. PIASy [RFK^{+}10]. PIASy-dependent [RFK^{+}10]. Pichia [MALS10]. pick [Sho14-63]. picture [Sed11g, Sed12m]. piece [CD14]. Pier [Pow10]. PIH [Sho10-47, YHK10]. Pih1d3 [DSB^{+}14]. PIK3C3 [LBD^{+}14]. PIM [SOW^{+}11]. Pin1 [RSM^{+}13]. Pincher [JDHS10]. Pincher-generated [JDHS10]. Pin’d [Sho13y]. PINK1 [ASLS14, JLW^{+}10, KLF^{+}14, LNJ^{+}13, MSS^{+}10]. Pins [WJPD11]. Pioneering [Sed12c]. PIP [Sho14-45, VLKI14]. PIPK [CDAK10a, CDAK10b]. PIPK- [CDAK10a, CDAK10b]. PIPs [Sho10-48]. piRNA [HLS^{+}14]. piRNAs [KT10]. pit [Les10-31, LADS10, NPL^{+}10, SNT^{+}12]. PKA [BWL^{+}11, DGS^{+}11, SYV14, Sho11-43, Sho12j, TMS^{+}12, LLH13]. PKA-dependent [BWL^{+}11, SYV14]. PKB [CDK^{+}10, XHB^{+}10]. PKC [CWB^{+}14, EAK13, HLN^{+}11, KLS^{+}13]. PKC- [KLS^{+}13]. PKC-dependent [CWB^{+}14]. PKcs [VEDBC13, ZYH^{+}11]. PKR [LSOT10]. place [Les11-32, Sho11x, Sho13-29]. plakoglobin [CLZ^{+}14]. Plakophilin [Sho10-49, WKGB^{+}10]. planar [BDR^{+}12, BW12, Dev14, LCHB13, MGG^{+}12, PJS^{+}11, SMdP^{+}14, dMSMZ14, WHH^{+}11]. plane [KMSR12]. Plant [Les11-31, Pow14b, Sed12q, Sed14s, Sho12-40, TLL^{+}13, ZGEM12]. plants [MFF^{+}13, WLW11, ZGW^{+}14]. Plasma [KPH^{+}12, RLS^{+}14, AOE^{+}10, AOE^{+}12, AXL10, AFRZ^{+}14, Bab14, CDAK10a, CDAK10b, HBG^{+}11, KBKW10, LMT^{+}10, Les11m, LWB^{+}14, MMU^{+}10b, NBC^{+}12, NSSF10, TID^{+}10, WWS^{+}12]. plasmodesmata [TLL^{+}13]. plastic [Sho13-42]. Plasticity [FW10, BGS13a, BGS13b, LLT^{+}12, MDP^{+}10, SBP^{+}10b, VG13]. plasticity-induced [MDP^{+}10]. plate [JKA^{+}10, Sho10r]. Platelet [SW10a, DGF^{+}14, MI13a, TMPH^{+}10]. platelets [TPM^{+}12]. platform [BGS13a, BGS13b, ENG^{+}12, TBV^{+}14]. play [BHMBS^{+}11, COW13, KKL^{+}14, TH11, XW10]. Playing [Les12z, Sed12r]. plays [BSP11, EMO12, WVT^{+}13]. PLC [BLO^{+}12]. PLD [Sho11-44]. PlexinD1 [TSH^{+}14]. PLK1 [CSS^{+}12, LR11a, BB10, KFS^{+}14, MGK^{+}12, ZSK^{+}13]. Plk4 [BKBR11, CAB^{+}10, HKH^{+}10, Sho10o]. PLP [LR13, Sho13-43]. plug [Sho14-34]. Plumbing [Pow10]. pluripotency [QB12]. pluripotent [Gol12b, WBBD14]. plus [BHB^{+}11, DLBG11, HKN^{+}10, JK10, LHW10, PAB^{+}10, vdVMG^{+}11]. PML [HLT12, SFJ^{+}14, Sho13-49, UAH^{+}12, dTLLB12]. pmn [SFB^{+}12]. podosome [DJL^{+}12, Les10-36, PCC11, QECC10, SZJ^{+}10]. podosome-like [SZJ^{+}10]. podosomes [OOKH^{+}12]. podosomes/invadopodia [OOKH^{+}12]. Point [LBS11, Gil10, GTS10, Sho10-68]. pointed [GLM^{+}10]. pointed-end [GLM^{+}10]. points [Les12f]. Pol [BBW^{+}13]. polar [CYLMM13, CSHS^{+}13a, CSHS^{+}13b]. polarities [BDR^{+}12]. Polarity [BYY^{+}12, CPS^{+}13, Les14z, BW12, CMS11, CFLDM11, Dev14, GHC^{+}14, ISZ^{+}11, JTN^{+}13, KBKW10, LDCF^{+}13, LCHB13, MLH12, MGG^{+}12, MJEM10, MSK^{+}13b, OPCEM10, Pow14b, Sed10b, Sed12q, Sho11-36, TLSA14, VMNLB^{+}11, WLN^{+}14, WHH^{+}11]. polarization [LN14, Nan14]. polarize [BDB^{+}14]. polarized [AOE^{+}10, AOE^{+}12, Bon14, DT14, GSJS10, GHGH11, KF11, NNO^{+}11, QMHM10, SPF11, dMSMZ14, WWM^{+}12, vGLWB12]. polarizes [LKG^{+}13, RSD^{+}12]. polarizing [BHB^{+}11]. pole [BKG10, Les13o, RZA^{+}13, SMS^{+}14, TGG^{+}11, TL12, VSMC11]. poleward [WBMCSS13]. Polo [BNDB^{+}14, HLN^{+}10, LDL12, KPE^{+}14]. Polo-like [BNDB^{+}14, HLN^{+}10, LDL12]. Poly [Leu14]. Polychaetoid [DSW^{+}11]. polycystic [CC10a]. Polycystin [HBG^{+}11]. Polycystin-2 [HBG^{+}11]. Polyglutamylation [Sho10-50, LvDG^{+}10]. polymerase [ALSN^{+}11, GSGL11, HM10, MVC^{+}11, RKW^{+}13, DPB^{+}10]. polymerases [KNH^{+}10, ZJP^{+}10]. polymerization [BLO^{+}12, BCBG10, CZC^{+}11, DMH^{+}12, KFET11, RDPG14, RBH^{+}12]. polymerization-dependent [KFET11]. polymers [HW11]. polymorpha [SKVvdK11]. polypeptide [KPI^{+}10, WGR^{+}12]. polyposis [OBD^{+}10].

Polysaccharide [Les13-27]. polysome [HARS14]. Polysomes [Pal14]. polyspermy [BXB^{+}12]. Polytopic [LJPJ11, LJLJ11]. Pom1 [RBB^{+}14]. POM121 [TH11, MMC^{+}10]. Pom33 [CHS^{+}10]. pombe [RKG^{+}10, TGG^{+}11]. pool [CWFL13, MSK^{+}13a, OD10, VJK^{+}10a, VJK^{+}10b]. pools [FSA^{+}11, HMB14]. populations [JDB^{+}12]. PopZ [LJW13]. pore [AMGC14, CHS^{+}10, CWFL13, DYS^{+}14, FMPS^{+}12, HBSD12, LLK11, MMU10a, MLW13, MMC^{+}10, MPM11, SZJ^{+}10, Sho11w, Sho11-27, Sho13w, SKFH11, TESA10, UHKS11]. pores [DSP11, DE10, Les10-28, Les12o, RKW^{+}13, Sho11r, SST^{+}12]. portal [Sho14-62]. portals [BC11a, BC11b]. position [BKP11, CKO^{+}10, KAAM11, Les13o, MCHCC10, VKMI12, ZME^{+}14, dJSTV12]. positional [RZA^{+}13]. positioned [VWC^{+}13]. positioning [ETYS^{+}12, GS11, KBG12, LDCF^{+}13, MJEM10, McN13, RFAA^{+}12, TSL12, ZEG11]. positions [KFET11]. positive [DCP^{+}10, WCQ^{+}13, LLR^{+}12]. positively [LHL11, LWL^{+}13]. possible [LCS^{+}10]. posterior [GHGH11, KWTR10, PHB^{+}11]. postfusion [MLSM^{+}11]. postmitotic [KZR^{+}12, LLK11, SZW^{+}11]. postnatal [CCM^{+}11]. postsynapse [SSL^{+}14]. postsynaptic [CB12, FDB^{+}13, NLP^{+}10, OFS^{+}10]. posttranscriptionally [JB12]. posttranslocationally [STG13]. POT [FTJG13]. POT-1 [FTJG13]. potassium [HJ14, MVC^{+}11]. potato [Sho13-44]. potent [RBP^{+}13, Ste12]. potential [bCAH^{+}11, FA12, JLW^{+}10, LGM^{+}12, RKS^{+}10]. potentiate [DYI^{+}13, SFL12]. potentiates [AFM^{+}13, CPX11, JCL^{+}11, lDSB^{+}10a, lDSB^{+}10b]. PP1 [HSS^{+}13, Sho13-53]. PP2A [BCB14a, EUB^{+}14, KST^{+}11, OMW^{+}14, RY11, VCF^{+}13, ZDM^{+}14]. PP2A-B56 [EUB^{+}14]. PP6 [Sho10-51]. Ppa [LNL11]. PPFR [GTR^{+}13]. PPFR-1-dependent [GTR^{+}13]. PPH-4.1-mediated [KHW^{+}10]. PPM1D [KSB^{+}13]. PPM1D/Wip1 [KSB^{+}13]. pre [OFS^{+}10, HIB^{+}10, KLZ^{+}12, NKH11, PASG^{+}12, RCBY^{+}12]. pre- [OFS^{+}10]. pre-60S [KLZ^{+}12]. pre-mRNA [HIB^{+}10, RCBY^{+}12]. pre-mRNAs [NKH11]. preassembly [DSB^{+}14, YHK10]. precedes [LLK11]. Precise [BT13, LNS^{+}13]. precision [KSW^{+}11]. precrossover [HSY^{+}14]. precursor [CVJ^{+}11, LCLW11, Sho13-46, UKZ^{+}13]. precursors [SW10a]. predicts [LHN10, MHAK^{+}12]. Preferential [NNO^{+}11]. premature [LS13b, STI^{+}11, ZGCG^{+}14]. Preperoxisomal [KMC^{+}14]. preribosome [BPH^{+}14]. Presenilins [Bez12]. presequence [GSM^{+}12, SLM^{+}11]. preserve [MGS14]. preserves [SPJ^{+}14]. press [Sed14f, Ros13]. pressure [Les13b]. Prestress [KCK^{+}14]. presynapse [CLS13]. Presynaptic [BRP14, CLD11, DMD^{+}12, FBZM^{+}10, KKMB10, KHG^{+}13, OZ14, RBA^{+}11, SSK^{+}13, dJSTV12]. prevent [BXB^{+}12, CHL12, CCM^{+}11, GSP^{+}14, KSLF^{+}11, Sho11-34]. preventing [BBW^{+}13, KWH^{+}10a, KWH^{+}10b, Les12t]. prevention [PXZ^{+}13]. prevents [CLO^{+}11, GSJS10, HWE^{+}12, HLT12, HCG^{+}11, JCN^{+}14, LS13b, PZ14, YHG^{+}14]. Prex1 [WHWS12]. price [Les14h]. pricked [Sho10y]. primary [GdAJ^{+}12, IMG^{+}12, MWH12, Sho10-36, Sho12-60]. primate [WMB^{+}10]. primate-specific [WMB^{+}10]. prime [Sho10-28, WPSA13]. primed [DYS^{+}14, Sho14-64, WWB^{+}10]. primer [VYM^{+}10]. priming [LSE^{+}10, VEDBC13, WLGC11]. Principles [ZF11, ZKR^{+}11]. Prion [Sho12-41, Sho14-46, DBUT13, RCFH10, Sed10n, TL12, WTBM12]. prions [KNPK^{+}10, RKK^{+}14]. prior [LS13b]. Prize [ME13]. PRKD2 [ONH^{+}12]. pro [DMD^{+}12, FBR^{+}10]. pro- [FBR^{+}10]. pro-peptide [DMD^{+}12]. proadaptive [BAB12]. proapoptotic [GWP^{+}11]. probe [HMB14, RKT^{+}14, Sed10a]. probed [BHA^{+}12]. Probing [BS13, Sho14-47, SKFH11, FHY^{+}10, GMD^{+}10]. problem [Les10-39, Les12w, Ros13]. problems [Sho13b]. processes [SDS^{+}12a]. Processing [CTH^{+}11, CMW11, IPM^{+}13, NZHL13, RGB^{+}13, SML^{+}13, XHB^{+}10, YTM^{+}11]. processive [HM10, KHB^{+}11b, WRCD12]. produce [BGY^{+}13]. produces [Sho11-58]. product [ZSK12]. production [BBJ^{+}10]. proficiency [MBZ^{+}10]. profiling [BAH^{+}12]. progenitor [BTC^{+}11, LRH^{+}13, TPZ^{+}14]. progenitors [CCM^{+}11, YMU^{+}10, YMU^{+}13, ZFA^{+}13]. Progeria [GCC12, SDD^{+}13]. Progerin [Les13-28]. prognosis [Sho12-43]. program [FPM^{+}14, Sho13-51, TGB10]. programmed [ZFA^{+}13]. programs [DKY^{+}12]. progress [Sho13r]. progression [ATU^{+}12, CSS^{+}12, CGRS^{+}12, FP10, FSK^{+}10, LWW12, LP13, NMB^{+}14, WBS11, WGN^{+}13, XTH^{+}11, ZNP^{+}13, ZSK^{+}13]. project [Sho12-50]. projections [KHG^{+}13]. proliferating [IMG^{+}12]. Proliferation [TC10, BKE10, CTL^{+}10, DPW^{+}11, DPW^{+}12, DCO^{+}13, HZS^{+}10, HSJ^{+}13, HK14, LSCF11, MPD^{+}12, WPM14, WMCF10, WSZ^{+}12, ZLW^{+}13]. prolonged [RGB^{+}13]. Prometaphase [CSTBM^{+}10, vZOtR^{+}10]. promiscuous [RKRB12]. promote [ALSN^{+}11, AA13, BLC^{+}12, BWL^{+}13, BPDB^{+}11, CTY^{+}12, DHB^{+}14, DWDW12, DCO^{+}12, DCO^{+}16, DJL^{+}12, DKA^{+}13, DHL^{+}12, EKJH13, HHL^{+}11, HLL^{+}12, HRWW^{+}13, IMP^{+}12, IIN^{+}11, KWDD10, KCF^{+}14, LVK^{+}13, LCHB13, LZY^{+}12, MMVK^{+}12, MI13b, MBVT^{+}13, MMFS11, MTT^{+}14, RCM^{+}12, RCG^{+}10, RCG^{+}11, SRS10, SSL^{+}14, VTO^{+}13, WBS^{+}12, WJPD11, WKN^{+}13, WRCD12, ZZW^{+}14]. promoted [CLW^{+}14, OSD^{+}14, YMT^{+}13]. Promoter [RKW^{+}13, HPB10]. Promoter- [RKW^{+}13]. promotes [AEC^{+}14, ALF^{+}13, ADAB^{+}12, BSR^{+}11a, BIY^{+}13, BWBC^{+}14, BWK^{+}11, BBW^{+}13, BKP11, BMLB^{+}12a, BMLB^{+}12b, BRP14, BLC^{+}14, CWL^{+}11b, CRP^{+}14, DPV^{+}12, DGF^{+}14, DSD^{+}13, EUB^{+}14, FHD^{+}12, FCE^{+}12, FPM^{+}14, GCP^{+}14, GBL^{+}11, GSM^{+}12, GZZ^{+}14, GSW^{+}11, GKR^{+}11a, GKR^{+}11b, GDS^{+}12, HWE^{+}12, HKN^{+}14, oHXK^{+}12, HRK13, HK14, IHM13, ILD^{+}10, IAMH10, JGB^{+}13, JRC^{+}13a, JRC^{+}13b, KHFV^{+}13, KBC^{+}14, KNsMK13, KSR^{+}13a, KSR^{+}13b, LAR^{+}12, LLU^{+}12a, LLU^{+}12b, LMW^{+}11, LZW^{+}12, LZW^{+}13, LADS10, LBD^{+}14, LP13, LLA^{+}12, MMO^{+}14, MRPR12, MKL^{+}13, NKH11, ONH^{+}12, PoLC^{+}13, dJPAA^{+}11, PLR^{+}13, PTST12, PVM^{+}12, PXZ^{+}13, QB12, RMM^{+}10, RSM^{+}13, RCBY^{+}12, RIG^{+}12, RB11, RDLT11, SPC^{+}13, SB14, SZJ^{+}10, SKV^{+}11, SJM^{+}13, SSK^{+}13, TID^{+}10, TNH^{+}11, UAH^{+}12, WWM^{+}12, WMV^{+}14, XTX^{+}13, YTM^{+}11, ZJP^{+}12, ZLW^{+}13, ZCB^{+}10a, ZCB^{+}10b, ZMW^{+}13]. promoting [CDD13, CLZ^{+}14, GWR12, HLT12, KOO^{+}14, NEMH^{+}10, WKGB^{+}10, WDB10]. prompt [YFO12]. promyelocytic [CAB^{+}13, dTLLB12]. prone [CLO^{+}11]. propagation [SCL^{+}14]. propel [Sho10-48]. propels [BCBG10]. proper [BVC^{+}11, CHS^{+}10, IP12, KLP14a, KBG12, LR13, NCT^{+}11, ZSK^{+}13]. properties [BKS^{+}13, GMD^{+}10, MBLD11]. prophase [BBD^{+}11]. proplatelet [TMPH^{+}10]. proprotein [MC10]. ProSAP1 [SSL^{+}14]. prosaposin [PTS^{+}10]. prosecretory [MBR^{+}11]. prostaglandin [Sho14-38]. protease [AWB^{+}14, BXB^{+}12, MAD10, SDN^{+}14a, SDN^{+}14b, WM12, YHG^{+}14]. proteasomal [KLC^{+}10, MHK^{+}10]. Proteasome [TCX^{+}10, Gol12a, HPB^{+}12, MFGB10, PHB^{+}13]. proteasomes [Sho13-30]. protected [KUH^{+}14]. Protecting [LAB14]. protects [GL10, LKLA12, ONNB^{+}14, RFL13, SHBC12]. Protein [DD10b, Ish14, Sho14-48, XWE^{+}10, ZBBG10, ADS^{+}13, ASE10, ABD14, ACS^{+}13, BIY^{+}13, BWK^{+}11, BVL^{+}12, BKP11, BGC^{+}14, BMLB^{+}12a, BMLB^{+}12b, BPL^{+}11, BMC^{+}11, BJ12, BAB12, CTM^{+}14a, CZD^{+}13, CHK^{+}10a, CHK^{+}10b, CPT^{+}12, CPT^{+}14, CB12, CLO^{+}11, CTH^{+}11, CGK13, CSS^{+}14, CKU^{+}10, CLW^{+}14, CM12c, COB^{+}12, CTW^{+}10, CTD^{+}10, DS10, DKF^{+}11, FS14, FUK^{+}14, FTJG13, FHKW11, GLB10, GB12, GSS^{+}11, HBM^{+}11, HHJ^{+}11, HMO^{+}14, HKI^{+}13, HVOF^{+}14, HDH^{+}10, HKW^{+}13, HBC^{+}10, IIWS14, IHG^{+}12, JJH^{+}10, KdKDP12, KWDD10, KIOY10, KSR^{+}13a, KHB^{+}11b, KLC^{+}10, KYP^{+}14, KKK^{+}11, LAR^{+}10, LNL11, LCfC11, LBS^{+}13, LCL12, Les11-40, Les11-47, Les13b, Les13e, Les13-38, Les13-34, Les14q, LAH^{+}12, LJPJ11, LYH^{+}13, LVB^{+}10, MAS11, MGG^{+}12, MALS10, MVC^{+}11, MBM^{+}10, MLBY^{+}10, MSZ^{+}12, OBD^{+}10, OMW^{+}14, OPM^{+}12, OMV^{+}11, PKD^{+}11, PR12, PLR^{+}13, PPD^{+}10, QJO10, RBY^{+}11, RCFH10]. protein [RCBY^{+}12, RKS^{+}10, RFRV12a, RFRV12b, RKRB12, RH10, SGC10, SAoS14, Sed10n, SBP^{+}10a, SRCP14, SZ12b, SS11, SDS^{+}12b, SMS^{+}13, SMMB11, SWC13, SP11, TKS^{+}13, TB13, TDV^{+}14, TLS10, VLG14, WK12, WMP^{+}14, WCQ^{+}13, WTBM12, WKN^{+}13, WLW11, YWJ^{+}12, YTT^{+}10, YTM^{+}11, YZL^{+}13, ZLW^{+}13, ZKR^{+}11, dSMSS13, BKBR11, RFC14]. protein-coupled [KYP^{+}14]. Proteins [Les11-32, AGL^{+}14, AXL10, BWK^{+}11, BPMK^{+}14, BBW^{+}14, Bon14, BAH^{+}12, BDvdK13, CDD13, ETYS^{+}12, ERS10, wFLW^{+}13, FSLM11, FWJ^{+}11, GSM^{+}14, GHGH11, HYS11, HMiY^{+}10, JDB^{+}12, KHS^{+}11, KK13b, KTN^{+}12, KIL^{+}12, KKL^{+}14, KUH^{+}14, KSS^{+}11, Les10r, Les10x, Les10-39, Les11-29, Les13-33, Les14z, Les14-31, Les14-36, LJLJ11, MHV12, MMU^{+}10b, NSBW10, OCF^{+}10, Omr10, PGP14, RPM^{+}13, RGB^{+}13, RKRB12, SFJ^{+}14, SDS^{+}12a, SA10a, Sed11n, Sed14l, Sho10-47, Sho12-53, Sho14-36, Sho14-59, SHN^{+}11, TTM^{+}14, TESA10, TP13, TCB^{+}14, VWC^{+}13, WLZ^{+}14, YHK10, ZNH^{+}11, ZGW^{+}14, vdVMG^{+}11]. proteoglycan [BST^{+}11]. proteolysis [BWL^{+}11, CRJB^{+}11, MMO^{+}14, SBTF13, WtLK^{+}13]. proteolytic [FPM^{+}14, JLW^{+}10, ZSK12]. proteome [BGS13a, BGS13b, LAB14, Sho13-42]. Proteomic [BCB^{+}14b, BAH^{+}12]. proteomics [HBC^{+}10, WM10]. proteostasis [SFK^{+}13]. proto [VBB^{+}10]. proto-oncogene [VBB^{+}10]. Protocadherin [BEJ10, EBBJ11]. Protocadherin-19 [BEJ10, EBBJ11]. Protocadherins [Les11-33]. protofilament [FSA^{+}10b]. proton [CFB^{+}12]. Protrusion [Sho12-42, Sho12-43, MHAK^{+}12]. protrusive [VMNLB^{+}11]. provide [Les11-30, Sho11p, Sho14-53, TKB^{+}14]. provides [HBI^{+}10, LAR^{+}10, LBWS10, Sho10-40, Sho12-43, Sho13x, Sho14-28, TMS^{+}12, WWHH10]. proximal [CSG14, RKRB12]. proximity [WWS^{+}12]. PrP [RKK^{+}14]. PSEN [CFB^{+}12]. PSEN-deficient [CFB^{+}12]. pseudopods [VKMI12, YZM^{+}12a]. PTB [LT11]. PtdIns [KPJ^{+}13]. PtdIns4 [Sho14-47]. PtdIns4P [NBC^{+}12]. Pten [HLL^{+}12]. PTH [OMZK14]. PtK1 [CSHS^{+}13a, CSHS^{+}13b]. PTP1B [NJS^{+}10]. pull [Les11p, Les12o]. pulling [KWTR10]. pulls [Les14-30, Sho14-34]. pulses [VTM14]. pump [CFB^{+}12, Les13v, Lin10, VM14]. pumps [VŠH^{+}11]. Purkinje [TAC^{+}13]. puromycylation [DDH^{+}12]. purpose [EEP13]. push [Les12b]. pushing [YRU^{+}13]. put [Les10-36, Sho11k, Sho11-47]. puts [Les11-39, Les13o, Sho10-51, Sho11x, Sho11-55, Sho12-41, Sho13i, Sho13-29, Sho13-30, Sho13-54, Sho13-48, Sho14-39]. Putting [Les13-29, Sho12-44, Sho12-45, SS11]. puzzle [CD14, Sed10n]. pVHL [KWK^{+}11, TMG^{+}10]. PX [KSR^{+}13a]. Pyk2 [CVR10]. pyrimidine [Sho12-46, ZDS^{+}12].

Q [Sed14j]. Quality [AXL10, Bab14, LAB14, RFC14, WP14]. Quantitative [FBZM^{+}10, GSU^{+}12, HBC^{+}10, TMG^{+}10, Sed11i]. Questing [Sed14i]. question [Sed13l]. questions [Sed11l]. quiescence [LLS^{+}11, LCS^{+}13]. quiescent [ZSH10]. quiet [Sho11-48].

R406W [GKWG^{+}11]. RA [PSF^{+}11]. RA-RAR- [PSF^{+}11]. Rab [Bar13, Sho10-52, ZIG^{+}12, BRD^{+}13, BPL^{+}11, LAR^{+}12, iYGL^{+}10]. RAB-6.2 [ZIG^{+}12]. Rab10 [CWZ^{+}12]. Rab11 [LLR^{+}12]. Rab11- [LLR^{+}12]. Rab18 [GBiY^{+}14]. Rab21 [MVP^{+}11]. Rab33B [IKU^{+}11]. Rab35 [HMiY^{+}10, Sho10-53]. RAB4 [FPM^{+}14]. RAB5 [FPM^{+}14, PLR^{+}13]. RAB5/RAB4 [FPM^{+}14]. Rab5c [ONH^{+}12]. Rab7 [PAB^{+}10]. RabGEFs [BRD^{+}13]. Rabs [Sho13v]. RAC [KPI^{+}10, DC12, DJL^{+}12, JGB^{+}13, NDS^{+}11, SCR12, YYA^{+}11]. Rac1 [AA13, BPB^{+}12, BDB^{+}14, CVR10, CDAK10a, CDAK10b, HHC^{+}11, KWH^{+}10a, KWH^{+}10b, LCK^{+}13, WWM^{+}12, WHWS12]. Rac1-dependent [BPB^{+}12]. Rac1-specific [HHC^{+}11]. race [ACO12]. RacGAP1 [JGB^{+}13]. RAD18 [GHK10a, HLH^{+}14, DPB^{+}10, Sho14n, ZLFC14].

RAD18-mediated [GHK10a]. RAD21L [LH11]. RAD23 [BTL^{+}12]. RAD51 [RZS^{+}14, CCJ^{+}12, PLC^{+}11]. Rad51-mediated [CCJ^{+}12]. Rad52 [TALR11]. Rad53 [CGRS^{+}12, Sho12-47]. Rad54 [AvCG^{+}11, Sho11-45]. Radial [WMP^{+}14, OYYK14, PBM^{+}11, Sho11-37, SDS^{+}12b]. Radil [ATU^{+}12]. Raf [DS12, MLH12]. raft [RKK^{+}14]. raft-associated [RKK^{+}14]. rafts [Les13x]. Rag [MP13, PRFF13]. RAGE [SJM^{+}13]. rails [Les10r]. Ral [CMS11]. RALT [FAB^{+}10, Les10-30]. RALT/MIG6 [FAB^{+}10]. Ramanujan [Sed10n]. Ramos [Sed13g]. Ran [HHJ^{+}11, Sho11-46, BRF^{+}10, Les13-30, SDD^{+}13]. Ran-dependent [HHJ^{+}11]. RanBP2 [GVP^{+}11, HHJ^{+}11, Les11-34]. RanBP2/Nup358 [HHJ^{+}11]. RanGAP [ZGEM12]. RANGAP1 [RDB^{+}12]. range [KXN10, KK13b]. RanGTP [HRK13, WJPD11]. RANKL [XTX^{+}13]. Rap [MBVT^{+}13]. Rap1 [AFM^{+}13, CMS11, Les13-31, WWS^{+}12]. Rap1-interacting [WWS^{+}12]. Rap1a [ATU^{+}12]. rapamycin [SGD^{+}10]. Raphael [Sed14u]. rapid [GB12, GEN14, KEJ13, RCG^{+}10, RCG^{+}11]. rapidly [GNHB11]. Raposo [Sed12j]. RAR [PSF^{+}11]. Ras [AFRZ^{+}14, CDK^{+}10, Les10-31, MLH12, MMU^{+}10b, RFRV12a, RFRV12b]. Ras-mediated [CDK^{+}10]. rate [ADF^{+}12, ZNP^{+}13]. rather [DKY^{+}12, LLS^{+}11]. RB [VES^{+}11, CZ10, MPD^{+}12, SZW^{+}11, Sho10-54, WAG^{+}10]. Rb-deficient [CZ10]. Rbfox3 [KNsMK13]. Rbfox3-regulated [KNsMK13]. RBM4 [LT11]. Rbms3 [JB12]. RC [PASG^{+}12]. RCP [JGB^{+}13, RCM^{+}12]. RCP-dependent [RCM^{+}12]. RCP-driven [JGB^{+}13]. Re [WCM12b]. reactivate [BRL14]. reactivation [ADB^{+}14]. reactive [Fin11]. readies [Sho13f]. readily [MSK^{+}13a]. Reading [Sed11p]. readout [Jan14]. ready [Les11-27, WPSA13]. Real [SBR^{+}11, LRA^{+}10, dSLPRG11]. Real-time [SBR^{+}11]. realistic [Ros10b]. rearrangements [LAO^{+}10]. receptor [AMH11, ALS^{+}13, ASB^{+}11, BST^{+}11, BGC^{+}14, BG11b, FAB^{+}10, GHK^{+}10b, GM11, GFSR11, HZM^{+}13, HSS^{+}13, JGA^{+}11, JOR^{+}11, JDL^{+}14, KDIE11, LAR^{+}10, Les11n, LYH^{+}13, MLM^{+}13, MVP^{+}11, MSR10, MHCvSW11, NRM^{+}12, NLJ^{+}13, NJS^{+}10, RPO^{+}14, SLM^{+}11, Sho12-61, SJM^{+}13, TTM^{+}14, VWD^{+}13, WLN^{+}14, ZGCG^{+}14, ZIG^{+}12]. receptor-mediated [BG11b, GM11, MLM^{+}13]. Receptors [Les11-35, BKT13, CSH^{+}12, oHXK^{+}12, JGA^{+}11, KYP^{+}14, LZW^{+}12, LZW^{+}13, LADS10, Sho12e, YMM^{+}10]. recipe [Les10-37, Pow14a]. reciprocally [SLK^{+}13]. Reck [Sed13s]. Recognition [BTL^{+}12, ABD14, CMS10, LAR^{+}10, LvBG^{+}10, MPRT11, SAoS14, vZOtR^{+}10]. recombination [BSR^{+}11a, HMBC10, MTM^{+}10, PHD^{+}10]. Reconstituted [WLGC11]. reconstitutes [BJE^{+}12]. Reconstitution [LMT^{+}12, HZE^{+}13]. reconstruction [Sho10-30]. recovery [GSGL11, MFR^{+}14]. Recreation [YHT^{+}10]. recruit [CSH^{+}12, HTS^{+}10, LTJN^{+}12]. recruited [lDSB^{+}10a, lDSB^{+}10b]. recruiting [KWH14, LMW^{+}11, PLR^{+}13]. Recruitment [BRL14, FMI^{+}13, PRFF13, BSR^{+}11a, BKS14, BB10, BWBC^{+}14, BVM^{+}11, BKBS12, CAB^{+}10, DSD^{+}13, HLT12, KKS^{+}14, KSH^{+}13, LLU^{+}12a, LLU^{+}12b, LOR^{+}10, MP13, MS14, OWC^{+}10, PSR^{+}10, PVM^{+}12, RB11, RDB^{+}12, Sho12m, SSK^{+}13, TLS10, VTO^{+}13, WM12, WBcY^{+}11]. recruits [HSTF13, KBC^{+}14, KIOY10, MSS^{+}10, MMFS11, MFR^{+}14, PPV^{+}14, SYH^{+}13, ŻKC^{+}11, vZOtR^{+}10]. recycle [Sho14s]. Recycling [Les12-27, ALF^{+}13, FPM^{+}14, GYZ^{+}12, HKN^{+}14, JGB^{+}13, LLR^{+}12, MMU^{+}10b, ONH^{+}12, SBTF13, Sho13-51, SHBC12, YSM10, ZIG^{+}12]. Red [Sho14-49, CBB12, SRCP14]. redistributes [CWB^{+}14]. Redox [TSB^{+}14, ACS^{+}13, KSSK12, LCK^{+}13, YFLH12]. reduce [KBW^{+}12]. Reduced [FUK^{+}14]. reduces [CLZ^{+}14, SBR^{+}11]. Reduction [DPZ^{+}14]. Reductionism [SJ13]. redundant [Les10-32]. Reelin [LXTM12]. reepithelialization [SSW^{+}13]. reflects [SDD^{+}13]. regeneration [DWJ^{+}14, DZT^{+}11, FHD^{+}12, KN12, LGM^{+}12, MHC^{+}12, Sho12v, Sho12-50, YFLH12]. regenerative [ZLW^{+}13]. regimen [Sho14k]. region [CNP^{+}12, UG10]. regions [AGM^{+}10, SBP^{+}10b]. regular [Les11i]. regulate [AGL^{+}14, AFRZ^{+}14, BLC^{+}12, BKS14, BDB^{+}14, CTL^{+}10, CSS^{+}12, DT14, DCO^{+}12, DCO^{+}16, FP10, GB12, GHK^{+}10b, GLM^{+}10, HGV^{+}14, HHS^{+}14, HVDG13, HCP^{+}13, JAM^{+}13, KSH^{+}13, KBW^{+}10, LNTR14, LZW^{+}10, LBWS10, MLY^{+}10, MAD^{+}11, NTSK14, NOT^{+}14, OPM^{+}12, PMHZ10, PASG^{+}12, PPV^{+}14, PTBT10, QWL^{+}11, SAG^{+}11, SBTF13, SEV^{+}14, SWC13, WLZ^{+}14, ZIG^{+}12, ZME^{+}14]. Regulated [LVB^{+}10, SHS^{+}13, APV^{+}12, ASE10, CMS^{+}14, CLEZ12, CRL^{+}14, FDB^{+}13, GSC11, IIWS14, ILD^{+}10, JKA^{+}10, KFS^{+}14, LCfC11, NB12, PSVRB^{+}11, PHB^{+}13, PBD^{+}13, PCCR11, TUG^{+}10, VCF^{+}13, WWB^{+}10, WHF^{+}11, WMP^{+}14, WWT^{+}12, WDG^{+}13, ŻKC^{+}11, BAAW11, KNsMK13]. regulates [ATU^{+}12, BWL^{+}13, BMG14, BNDB^{+}14, BB10, BPT^{+}14, BWBC^{+}14, BDR^{+}10, BKY^{+}10, BDC^{+}14, BRF^{+}10, BVR11, CVR10, CMS11, CZD^{+}13, CBBH11, CDAK10a, CDAK10b, CMH^{+}10, bCAH^{+}11, CLC^{+}11, CG12a, CZM^{+}14, CG12b, CWC^{+}13, CFLDM11, CRJB^{+}11, CWG^{+}11, CTD^{+}10, DGH^{+}14, DWPC^{+}11, DYS^{+}14, DCN^{+}10, DS10, DWDW12, DWM^{+}12, EL14, FS14, FPAM13, FMG^{+}11, FSLM11, mFH13, GdBP^{+}14, GHC^{+}14, GSM^{+}14, GLB10, GDO13, HBM^{+}11, HKN^{+}11, HTT^{+}11a, HVW^{+}10, HLL^{+}12, HLS^{+}14, IKU^{+}11, ISZ^{+}11, JLW^{+}10, JC10, KUN^{+}13, KA12, KKL^{+}11, KPH^{+}12, KST^{+}10, KKK^{+}11, LCBG^{+}11, LHL11, LGM^{+}12, yLFAM13, LAH^{+}12, LMC^{+}12, LT11, LWL^{+}13, LSM^{+}11, LDL12, LSS^{+}12, LYB^{+}10, LLR^{+}12, LMS^{+}10c, LXTM12, MTG^{+}11, MLM^{+}11, MVR^{+}10, MVP^{+}11, MJFS10, MBM^{+}10, MFA^{+}14, NLP^{+}10, NLJ^{+}13, NCT^{+}11, NT11, NPL^{+}10, NEMH^{+}10, NSSF10, NBDB12, NJS^{+}10, OFS^{+}10, PKG10, PSR^{+}10, PSK11, PKS^{+}10, RNS^{+}14, RBY^{+}11]. regulates [RLS^{+}14, RBB^{+}14, RDB^{+}12, RFK^{+}10, SZ12a, SSdA^{+}14, SBS^{+}12, SPF11, SWV^{+}10, SJZ^{+}10, SMM^{+}10, SGD^{+}10, TIT11, TKL^{+}10, TPZ^{+}14, VYC^{+}11, VBB^{+}10, VTM14, WEK^{+}14, WAW^{+}11, WJW^{+}11, WHWS12, WHDR^{+}10, XWE^{+}10, XSJ^{+}10, YYM^{+}11, YSO^{+}11, YYA^{+}11, ZLFC14, ZBBG10, ZWL^{+}14, ZZW^{+}10]. Regulating [KD11, BBJ^{+}10, CGW^{+}11, CGRS^{+}12, DKMK^{+}11, JYRL^{+}13, KSR^{+}13b, NGL^{+}12, PCO^{+}10, ZMW^{+}13]. Regulation [CKU^{+}10, GCSB10, HMiY^{+}10, RH10, SA10a, SAoS14, TLSA14, CS13, GOWM12, GWP^{+}11, HVW^{+}10, JBS^{+}12, JBS^{+}13, JPT^{+}11, KPE^{+}14, KMS10, KSSD11, LS13a, MLH12, MLG^{+}10, MC10, MVP^{+}10, NMB^{+}14, NSZ^{+}13, OS13, PWP11, PXZ^{+}13, RT12, RY11, SBR^{+}11, SFK^{+}13, SSB^{+}10, Sho11z, Sho11-32, SLK^{+}13, SLC^{+}13, SLS^{+}10, TMS^{+}12, WGN^{+}13, XG12, ZSK12]. regulator [CLC^{+}11, DCP^{+}10, HS10b, KPSL12, LNL11, MBK^{+}10, MH14, OBC14, PLL^{+}12, TMG^{+}10, TQM^{+}14, WCQ^{+}13, YSaY^{+}13, ZDM^{+}14]. Regulators [BNL^{+}10, CTY^{+}12, FHA10, RTM13, Sho12d, VKMI12]. regulatory [BKBR11, FRS^{+}13]. Reik [LeB10]. reinhardtii [MBLD11]. reintroduction [SKVvdK11]. related [BYY^{+}12, CLW^{+}14, DKF^{+}11, EBB13, FHKW11, HKN^{+}11, SA10a, SWC13, TTM^{+}14, ERS10]. relations [SY10]. relay [Sed14e]. relays [BVL^{+}12]. releasable [MSK^{+}13a]. Release [JDS^{+}10, LCBG^{+}11, MKS^{+}13, Sho12-48, CSP^{+}10, DAB^{+}11, DKM^{+}13, GEN14, HSJ^{+}13, KBAW^{+}11, MMV^{+}10, Sho10-28, Sho11c, Sho11d, Sho13f, TMPH^{+}10, VM14, WEK^{+}14, vdBFS^{+}12]. releases [SHC^{+}13]. relief [Les11-30]. relies [RPM^{+}13]. relocation [GCSB10]. remaining [CC10b]. remember [Les10d, Sed12p]. remodel [Sho14s]. remodeled [FWJ^{+}11]. remodeler [CCJ^{+}12]. Remodeling [CGCP^{+}14, ABVP11, BPH^{+}14, BG11a, CWC^{+}13, DWJ^{+}14, FRL^{+}13, FMG^{+}11, HOS^{+}12, KLP^{+}14b, KSR^{+}13a, LPG^{+}10, MRR^{+}12, RGF^{+}10, RDC^{+}11, SKM10a, SHS^{+}12, SWV^{+}10, TIT11, TKS^{+}13, TO12, TCN14]. remodels [Les14q, MPM11]. remotely [AFRZ^{+}14]. removal [SML^{+}13, TCB^{+}14]. remove [Les11-35]. remyelination [CVJ^{+}11]. renal [DHB^{+}14]. renewal [LGM^{+}12, QB12, Sed14x, ZLW^{+}13]. renovations [Les10-27]. reorganization [BDN^{+}13, SBEM13]. reorganizes [LCS^{+}13]. reorient [WAJ^{+}12]. reorientation [ADAB^{+}12, RSRK13]. Repair [Sho11-47, ABVP11, BSR^{+}11a, BNDB^{+}14, BCJ13, CCJ^{+}12, DWL^{+}11, ETI^{+}10, EIE^{+}14, GSB^{+}13, GHK10a, IAMH10, JPT^{+}11, KK13b, KLP^{+}14b, LPG^{+}10, LGM^{+}12, Les10s, Les10z, Les10-39, Les12-28, Les14h, LvBG^{+}10, MHV12, MBZ^{+}10, PSVRB^{+}11, PLC^{+}11, PVM^{+}12, RSS^{+}13, RC12, SKM10a, Sho10m, SWV^{+}10, TID^{+}10, VB12, XSJ^{+}10, YTM^{+}11, ZYH^{+}11, ZNA^{+}14]. repeat [BBK^{+}13, FS14, GZZ^{+}14, SKFH11]. Repeated [KRS11]. repeats [CSH^{+}12]. repertoire [Sho11-53]. Repetitive [Les10-32, RKE14a, RKE14b, DAB^{+}11]. replenishment [VPC^{+}14]. replicate [DKY^{+}12]. replicated [BBW^{+}13]. Replication [HBC^{+}11, OMV^{+}11, TGB10, TLL^{+}13, YTM^{+}11, BDC^{+}14, DPB^{+}10, DKMK^{+}11, DKY^{+}12, EIE^{+}14, GZR^{+}14a, GZR^{+}14b, GB10, GZZ^{+}14, Gil10, JRC^{+}13a, JRC^{+}13b, KSS^{+}11, KSSD11, yLFAM13, Les10c, Les14l, LLM^{+}10, MGS14, MHKM11, MFA^{+}14, MFR^{+}14, NZHL13, PASG^{+}12, RZF^{+}11, SKN^{+}13, SNSyN13, Sho13l, Sho13m, Sho13-38, Sho14y, Sho14-50, SZE^{+}11, SQC^{+}12, VYM^{+}10, WGC11, ZNP^{+}13, ZMW^{+}13]. replicons [SKN^{+}13]. Repo [Sho12-51, WHL^{+}12]. repositioning [YWC^{+}13]. represses [dJPAA^{+}11, SZE^{+}11, XTH^{+}11, ZFP^{+}13]. repressing [CTL^{+}10]. repression [BWL^{+}11, COG11, CMW11, GBSC^{+}12, PSF^{+}11, SHC^{+}13]. repressive [SMB12]. repressors [LTJN^{+}12]. reproductive [CRL^{+}14]. reprogram [CRL^{+}14]. reprogramming [QB12]. reprograms [ZLH^{+}14]. repulsion [SSD^{+}14, SMM^{+}10, SGT^{+}13, ZPB^{+}12]. require [KST^{+}11, RBF^{+}12]. required [ADB^{+}14, ATKK11, AKA^{+}13, AYS^{+}13, BSR^{+}11b, BSR^{+}11c, BV11, BCBG10, BBW^{+}14, BVM^{+}11, CTM^{+}14a, CMS10, CHS^{+}10, CMH^{+}10, DSB^{+}14, ETC^{+}12, FSK^{+}10, FHY^{+}10, GZR^{+}14a, GZR^{+}14b, GBK^{+}14, GBiY^{+}14, GZLG11, GCV^{+}11, HKH^{+}10, HCC^{+}10, HTS^{+}10, HCCS^{+}11, KPE^{+}14, KPJ^{+}13, KNH^{+}10, KFH^{+}12, KWK^{+}11, MBR^{+}11, MWP^{+}12, MWZ^{+}11, MLG^{+}10, MHC^{+}12, MZP^{+}10, MCS^{+}13, NvCL^{+}13, NDS^{+}11, NBS^{+}11, OBSG12, PCC11, PAB^{+}11, QMHM10, RGL^{+}13, RPO^{+}14, RSRK13, SYS^{+}14, SBEM13, SNR^{+}11, SHC^{+}10, SCN^{+}14, SMS^{+}14, SMMB11, SRU^{+}12, TB12, TDV^{+}14, VGL^{+}14, WLW11, XBC^{+}13, YOA^{+}11, YTT^{+}10, ZYF^{+}11, ZSK^{+}13, ZPB^{+}12, vGCMA^{+}14, vBAK^{+}12]. requirement [BGC^{+}10, KPC^{+}11]. requirements [GvEM^{+}11, KLvdB^{+}13, LMA^{+}13, TTC^{+}14]. requires [ASLS14, BPL^{+}11, BMFC^{+}11, BMFC^{+}13, CWFL13, EZT^{+}12, ETYS^{+}12, GFSR11, HM10, IM11, KIOY10, KTN^{+}12, KFS^{+}14, KSS^{+}11, KTB^{+}14, LBS^{+}13, MLW13, MMS^{+}10, NSS13, PR12, SKVvdK11, SWS^{+}13, SFK^{+}13, SPJ^{+}14, UG10, ZSD^{+}14, ZSZ^{+}13]. Rer1p [JYRL^{+}13]. Rescue [CZ10]. rescues [SFB^{+}12]. research [Gol12a, Sed10a]. resection [KFH^{+}12, PLC^{+}11, RSS^{+}13]. resection-mediated [RSS^{+}13]. reserve [Sho10j]. reservoir [CC10b]. resetting [KAS^{+}12]. reshaping [TIT11]. resident [IKU^{+}11, RPM^{+}13]. resides [Sho13w]. resistance [GWR^{+}10, Sed11t, Sho14-28]. resolution [BCJ13, FAvdB^{+}12, FSA^{+}11, FSA^{+}10b, OYH13, PSVRB^{+}11, RKT^{+}14, SHL10, TIM14, YOMM^{+}11]. resolves [RZF^{+}11]. resorption [SNR^{+}11]. respiration [LWBH12]. respond [BLM^{+}11]. response [BKG10, BAB12, CPT^{+}12, CPT^{+}14, CRJB^{+}11, GBSC^{+}12, GBJ10, JRC^{+}13a, JRC^{+}13b, KHW^{+}10, KMSR12, Les14s, LCK^{+}13, MFB12, MKL^{+}13, PLL^{+}12, PHW^{+}13, RKS^{+}10, RH10, SSD^{+}14, SRBL13, SS11, WK12, XWE^{+}10]. responses [BGS13a, BGS13b, CLL^{+}10, SA10a, SJM^{+}13, YFLH12]. responsible [MLM^{+}13, WOG13]. responsive [NKH11]. REST [KST^{+}10, PMP^{+}11a, PMP^{+}11b]. REST/ [PMP^{+}11a, PMP^{+}11b]. restorative [MKS^{+}13]. restore [Sho12-46]. restored [Sho13q]. restores [DGS^{+}10, SKM10a]. restraining [GYC^{+}14]. restraint [Les13k]. restrict [CO13, HVDG13, MPD^{+}12, VMNLB^{+}11]. restricted [BCB14a, MBR^{+}11]. restrictions [Sho10-46]. restricts [BLI^{+}10, VOSB12]. resulting [BvMD^{+}14]. resumption [OHC10]. resveratrol [MMB^{+}11]. retain [MBZ^{+}10]. retains [Sho11-41]. retention [CWG^{+}11]. reticulum [AiIK^{+}13, CTH^{+}11, DPZ^{+}14, GF11, Les11-36, LSOT10, MMS^{+}10, MDW^{+}13, RPK^{+}11, WLW11]. retina [SW10b]. retinoblastoma [RCG^{+}10, RCG^{+}11]. Retinoic [Sho11-48, MMdCOM^{+}11]. retraction [AMH11, GdBP^{+}14, GPCK12]. retrieval [CZC^{+}11, KKS^{+}14]. Retrograde [BMG14, BLO^{+}12, CLW^{+}14, EIW^{+}12, JDHS10, KKUG11, LHL11, NLP^{+}10, Sho11m, TTB^{+}13, XWE^{+}10, YZPF12, ZIG^{+}12]. retromer [ZIG^{+}12]. retrotransposon [YOA^{+}11]. reveal [KNOM11, MMV^{+}10, RZA^{+}13, WUD^{+}12]. revealed [FWM^{+}10a, SSW^{+}13, SNT^{+}12, SBP^{+}10b]. Revealing [Sho11-51]. reveals [ACS^{+}13, BV11, BGS13a, BGS13b, BKS^{+}13, CDH^{+}14, CTM^{+}14b, DE10, EAK13, FSA^{+}11, FBAO^{+}13, FBR^{+}10, HMB14, HZE^{+}13, HCCS^{+}11, HBC^{+}10, HIB^{+}10, Ish14, KWL^{+}12, LWB^{+}14, dSLPRG11, NSD^{+}14, OD10, PGCY12, RKK^{+}14, SBR^{+}11, SSZ^{+}14, SLC^{+}13, TP13, WZHV11, YOMM^{+}11]. reversal [Sho13-38]. reverse [TSB^{+}14, YFO12]. reverses [ZDS^{+}12]. reversible [LvBG^{+}10]. reversine [STD^{+}10]. Review [WCM12b]. revised [SSW^{+}13]. Revisiting [PH10]. Revolving [MAE^{+}10]. Reya [Sed12w]. Rgl [CMS11]. RGS7 [OPM^{+}12]. rhabdomyosarcoma [LSCF11]. rhei [PM13]. Rho [dMSMZ14, LCS^{+}10, MBVT^{+}13, NGL^{+}12, Sed12n, Sho10w, TCN14]. Rho1 [BT13, NSSF10, OKNP13, PDMBW11, SKVvdK11]. RhoA [ABP^{+}12, BVC^{+}11, CBBH11, DKA^{+}13, HCC^{+}10, JGB^{+}13, RBP^{+}13, VFNR11, ZFA^{+}13, ZLH^{+}14]. RhoA/MAL [DKA^{+}13]. Rhoades [Sed10l]. RhoBTB3 [LP13]. RhoC [VFNR11]. rhodopsin [BAAW11, MBLD11, PSK11]. RhoG [HYTU^{+}10]. RhoG-dependent [HYTU^{+}10]. RhoGEF [LLcK^{+}11, YYA^{+}11]. RhoGEF12 [TKS^{+}13]. rhomboids [Sed14o]. RIAM [WWS^{+}12, WHDR^{+}10]. ribbon [HCG^{+}11, OZ14, VPC^{+}14]. ribonuclease [SB14, Sho14-51]. ribose [Leu14]. ribosomal [BHA^{+}12, GKR^{+}11a, GKR^{+}11b, OCF^{+}10, SBP^{+}10a, SML^{+}13]. ribosome [ARF10, DDH^{+}12, Kar10, LJPJ11, Oef10, Sho12-49]. ribosome-anchored [ARF10]. ribosome-bound [DDH^{+}12]. ribosomes [KPI^{+}10, Les14t, OBM^{+}10, ONNB^{+}14]. ribosylation [BPDB^{+}11, CPT^{+}12, CPT^{+}14]. Rich [NLP^{+}10, NNO^{+}11]. Richard [BG11b]. rictor [TAC^{+}13]. ride [Les10l, Les13s, Sho10-32]. right [BT13, KK11, Sho14-44, TKMK10]. right-screw [TKMK10]. RII [DWM^{+}12]. RIM1 [FBAO^{+}13, SSK^{+}13]. RIM1/CASK [SSK^{+}13]. ring [CWL^{+}11a, CP11, CLS^{+}10, CSKW13, EKJH13, ENG^{+}12, GBL^{+}11, GMW^{+}13, HHY^{+}12, HC10, Les11a, LCD^{+}11, MHCvSW11, MCS^{+}13, Sho10-40, Sho12q, Sho12w, Sho13s, TO12]. rings [Sed10j, MHCvSW11]. RIP1 [KDIE11]. Rise [Les13-32]. rising [CM12a]. rivals [Sho10-66]. Rlp24 [KLZ^{+}12]. RNA [ALSN^{+}11, KYOY13, Les11z, MVC^{+}11, MPRT11, MTT^{+}14, RCBY^{+}12, RKW^{+}13, RKE14a, RKE14b, SB14, Sed11j, SLH13, Sho11-49, Sho13-45, Sho14-52, SML^{+}13, TLTW10, YZL^{+}13, ZZW^{+}14, ZJP^{+}10]. RNA-based [SLH13]. RNA-binding [RCBY^{+}12, YZL^{+}13]. RNA-mediated [ZZW^{+}14]. RNAi [ASE10, CG10b, DV10, RSL^{+}11, TSH^{+}14]. RNAs [EMT^{+}14, Les13-35]. RNF111 [PHW^{+}13]. RNF111/Arkadia [PHW^{+}13]. RNF169 [Les12-28, PLL^{+}12]. RNF8 [DPV^{+}12]. Rng [Sho14r]. Rng8 [WLZ^{+}14]. Rng9 [WLZ^{+}14]. RNP [GYC^{+}14, YZL^{+}13]. RNPs [HYS11]. Robinson [Sed14n]. Robo [CO13]. Robust [JTN^{+}13, PL10, BKK^{+}10, BKE10, CSTBM^{+}10, DWDW12, RBB^{+}14]. robustness [GYC^{+}14, SFK^{+}13]. ROCK [ZLH^{+}14, Sho10-34]. Rockefeller [Ros13]. rod [Sho13-37]. Rod1 [Les12-30]. Role [ZPS^{+}10, ASE10, BHMBS^{+}11, BC11a, BC11b, CGW^{+}11, DKF^{+}11, EMO12, EIW^{+}12, FBAO^{+}13, FBR^{+}10, GdAJ^{+}12, Les11-36, Les12t, Les13-27, MSC^{+}10, MHKM11, OD10, PTS^{+}10, RKS^{+}10, RFRV12a, RFRV12b, STD^{+}10, Sha10, Sho10-57, TH11, WVT^{+}13, YSN^{+}11, YOMM^{+}11, ZJP^{+}12, dSMSS13, MPRT11]. roles [BSP11, COW13, EAK13, EDF^{+}10, KKL^{+}14, LC10, OKNP13, SLC^{+}13, VFNR11, XW10]. roll [Sho13-35]. rolling [YZM^{+}12b]. Rong [Sed11o]. root [Sho10-56]. rooted [Sho11-45]. ROS [Les12k]. Rosenblatt [Sed13l]. rosettes [PCC11]. rotation [EM11, TKMK10]. Roth [Sed13n]. round [Sho10-61]. rounds [Sho11-53]. Route [Krä13, Sho13-40, Sho13-50]. routes [HBG^{+}11, Les11u]. RPA [MFR^{+}14]. Rpd3 [ZFP^{+}13]. Rpgrip1l [MGG^{+}12]. rRNA [ALSN^{+}11, BPH^{+}14]. RSC [RGF^{+}10]. rudder [Les12z]. Rüdiger [Sed11p]. ruffle [HTT^{+}11a]. ruffles [GNHB11]. rule [BV11]. rules [Sho14m]. running [Les11-34, Sed10j]. runs [Sho12-49]. Runx1 [OLT11]. Runx2 [HSK^{+}10]. Ruth [Sed11q]. RYK [BAY^{+}11, LBWS10]. RZZ [MS14]. RZZ/Mad1 [MS14].

S [HIM^{+}10, AKC^{+}12, BNL^{+}10, COW13, CGRS^{+}12, EMO12, GSJS10, HBSD12, JEF^{+}11, LKLA12, MGS14, MFGB10, OYH13, Sho12a, WKN^{+}13, YFO12]. S-glutathionylation [HIM^{+}10]. S-nitrosylation [LKLA12]. S-phase [YFO12]. S2 [DV10]. S408 [RBY^{+}11]. S6 [CLD11]. SAC [Les14j]. Sac1 [CDH^{+}14]. Saccharomyces [CT10, MWZ^{+}11, NSBW10, SMMB11, WMCF10, YCP10]. SADS [Sho13-47]. safe [CSM^{+}12]. safeguards [OMV^{+}11]. Sahai [Sed11g]. Sally [Sed14v]. Salmon [Sed10p]. salt [MLSM^{+}11]. Sam68 [VBB^{+}10]. Samara [Sed13s]. same [IP12, Ros10a, Sho11j]. sampling [KAS^{+}12]. Sánchez [Sed11a]. Sandhya [Sed13t]. Sandra [Sed10o]. Sanpodo [Sho13-46, UKZ^{+}13, CTM^{+}14b]. Sar1 [LYB^{+}10, Sho10-55]. SARAF [JAM^{+}13]. sarcolemma [RGL^{+}13]. sarcomere [VGL^{+}14]. sarcomeres [dSLPRG11, Sho12t]. sarcoplasmic [GF11, Les11-36]. Sarcospan [MHC^{+}12]. Sarcospan-dependent [MHC^{+}12]. Sarm1 [CLC^{+}11]. Satb1 [FMG^{+}11]. SATB2 [WSZ^{+}12]. Satellite [Sho14-52, BIY^{+}13, CTL^{+}10, Les10o, RKE14a, RKE14b, SMZL13, ZSH10]. save [Les12-36]. saves [Les12-31]. Saving [Sed14h]. says [Les11f]. scaffold [CAB^{+}10, HCCS^{+}11, KHS^{+}11, LYH^{+}13, MRR^{+}12, Sho11-30, Sho12-32, Sho14z, SDS^{+}12b, VTO^{+}13, ŻKC^{+}11]. scaffold-like [HCCS^{+}11]. scaffolding [GLB10, GB12, VGL^{+}14, WVT^{+}13]. scaffolds [GMW^{+}13]. scalability [CWL^{+}11a]. scale [SBP^{+}10b, TGES12, TP13]. scales [RDPG14, WCM12a]. scaling [Bra13]. SCAM [WKN^{+}13]. SCAR [Sho12-63, VKMI12, LVK^{+}13]. Scar/WAVE [LVK^{+}13]. scattering [KFL^{+}14]. SCF [BDN^{+}13]. schedule [Les10c]. scheme [Sho12-46]. Schizosaccharomyces [RKG^{+}10, TGG^{+}11]. Schmid [Sed10o]. Schroer [Sed13w]. Schuh [Sed14p]. Schwann [Sho11-50, Sho12-50, FSK^{+}10, FHD^{+}12, NBS^{+}11]. science [Hal14, Sed12u]. scientists [Hal14]. scission [BPDB^{+}11, WAW^{+}11]. sclerosis [AIJI11, Ste12]. scope [Sed13x]. Scoring [GGR12]. screen [ASE10, DV10, TSH^{+}14, TDV^{+}14]. screening [BGS13a, BGS13b, CG10b, FWM^{+}10a, wFLW^{+}13, RSL^{+}11, VvDV^{+}10]. screw [TKMK10]. SDF1 [LNS^{+}13]. SDF1-mediated [LNS^{+}13]. Sds22 [PKS^{+}10, Sho12-51, WHL^{+}12]. seal [Les11-41, Sho10e]. seals [Sho12x, Sho13-46]. Sean [Sho10-56, Sho11-51]. search [Sho10-29]. Sec [Sho12-52]. Sec12 [Les14-27, SYS^{+}14]. Sec13 [OTLH10, WS10]. Sec13/Sec31 [OTLH10]. Sec16 [WS10]. Sec1p [BJ12]. Sec1p/Munc18 [BJ12]. Sec31 [OTLH10]. Sec6p [SYH^{+}13]. SecA [WGR^{+}12]. second [BZ12]. Secret [Les10-33, MH11]. secretase [JYRL^{+}13, UKZ^{+}13]. secreted [LDN^{+}13]. secretion [BMC^{+}11, CGK13, DAS^{+}10, HMO^{+}14, HMiY^{+}10, LMS^{+}10c, MALS10, Pfe10, SZ12a, Sho10-53, Sho11p, Sho14-67, WPL^{+}11a, WPL^{+}11b]. Secretory [Sho12-53, ASE10, BKT13, CGCP^{+}14, FWM^{+}10a, FHA10, GvEM^{+}11, KBC^{+}14, LSE^{+}10, LN14, Sed11n, SYH^{+}13, TPM^{+}13, WGM^{+}12, vBAK^{+}12]. secrets [Sho14-67]. secure [BPMK^{+}14]. SecY [PR12]. SecYEG [ADS^{+}13, Les13-33, WGR^{+}12]. Seed [Dun11]. seeds [Sho12-40]. Seg1 [MSS^{+}12]. segment [BLI^{+}10, BGY^{+}13, JKS14]. segments [LJLJ11]. segregate [TP13]. Segregating [Sho10-57, Sho12-54]. segregation [BMLB^{+}12a, BMLB^{+}12b, BDvdK13, CLO^{+}11, ECC^{+}13, GCR^{+}12, HSN^{+}11, IP12, LR13, OB12, Sho12-56, SMS^{+}14, VTO^{+}13, WSUT11, WUD^{+}12, WHL^{+}12, Yam13]. SEL1L [BGC^{+}10]. selected [BBW^{+}14, TLTW10]. selection [RKS^{+}10]. Selective [BWL^{+}11, GWP^{+}11, LMS^{+}10c, MHK^{+}10, Oka14, ONNB^{+}14, PMB^{+}11, TTM^{+}14, ZZW^{+}13, NLP^{+}10]. self [CMS^{+}14, HYS11, IM11, LNJ^{+}13, LGM^{+}12, Les10m, Les13k, MKS^{+}13, QB12, Sho12g, SBP^{+}10b, WMB12, ZLW^{+}13]. self-assembly [SBP^{+}10b]. self-association [LNJ^{+}13]. self-control [Les10m, Sho12g]. self-destruction [WMB12]. self-oligomerization [IM11]. self-regulated [CMS^{+}14]. self-renewal [LGM^{+}12, QB12, ZLW^{+}13]. self-restorative [MKS^{+}13]. self-restraint [Les13k]. selfish [SCL^{+}14]. Semaphorin [TSH^{+}14, ZPB^{+}12]. Sending [Les14-28]. sends [Les11-48, Les14u]. Senescence [MMO^{+}14, BS13, DCO^{+}13, IPM^{+}13, RC11, SMZL13, XTH^{+}11, ZGCG^{+}14]. Senescence-inducing [MMO^{+}14]. Senescent [Sho13-47, DKM^{+}13]. SENP6 [Les10-34, MAD10]. sense [Sho10-38]. sensing [CLM^{+}10, CWC^{+}13, KAS^{+}12, OS13, QJO10, YDB^{+}11]. sensitive [Sho12j, Sho14-55, KPC^{+}10, SRCP14, WMV^{+}14]. sensitivity [KSW^{+}11, SDD^{+}13]. sensitizes [KCK^{+}14]. sensor [ACS^{+}13, FCA10, HTS11, JC10, PGAE^{+}13]. sensors [PMB^{+}11]. sensory [LZW^{+}12, LZW^{+}13, UKZ^{+}13]. sentence [Les11w]. Sentin [LMW^{+}11, LMT^{+}12]. separate [Les10x, Les14-28]. separation [GCP^{+}14, STI^{+}11, TGES12, TW14]. SEPT9 [EDF^{+}10, KFET11]. septation [MCS^{+}13]. Septin [BHB^{+}11, DBH^{+}11, Ewe11, BBK^{+}13, GBL^{+}11, GPCK12, KFET11, Sho11-33, Sho11-53]. Septins [DHB^{+}14, Sho14-53, EKJH13, EDF^{+}10]. septum [CSM^{+}12, OKNP13]. sequence [FLVP10, GZZ^{+}14, RCFH10, TMG12]. Sequential [YRU^{+}13, ZNA^{+}14, HHS^{+}14, LMS^{+}13]. sequentially [LZY^{+}12]. sequestering [TL12]. Sequestration [LS13b, Sho13-48, HSJ^{+}13, KBW^{+}12, MMV^{+}10, SMMB11]. Ser [RTC^{+}13a, RTC^{+}13b]. Ser285 [IMP^{+}12]. Sergio [Sed12t]. serine [CFLDM11, LCBG^{+}11, PAB^{+}11, RTC^{+}12, SDC10]. serine/threonine [CFLDM11, LCBG^{+}11]. serine/threonine-phosphorylated [LCBG^{+}11]. serotonin [DK10b]. serpentine [ASB^{+}11]. Serratia [HMO^{+}14]. serum [CZM^{+}14]. serum/glutamine [CZM^{+}14]. SesB [DGH^{+}14]. set [Les10c]. Set7 [TNH^{+}11]. Set7/9 [TNH^{+}11]. SET8 [JEF^{+}11]. sets [Les12g, Sho10-60, TMG12]. Setting [Sho10-58, Sho11-52, Sho13-49]. settings [Sho13m]. settles [Sed14u]. several [GDS^{+}12]. severing [CP11, GTR^{+}13, LvDG^{+}10]. severs [JLVH12]. Sey1p [AKC^{+}12]. SF2 [VBB^{+}10]. SF2/ASF [VBB^{+}10]. Sgt1 [DK10a]. SH3BP1 [EZT^{+}12, TSH^{+}14]. SH3YL1 [HTT^{+}11a]. Shah [Sed11i]. shape [BSO^{+}14, DGF^{+}14, KEJ13, Les14q, MSS^{+}12, PTBT10, RLS^{+}14, Sed14d, Sha10, Sho10-55, Sho12-28, Sho13-41, Sho14-51, TBV^{+}14, WMCF10, ZGEM12]. shape-shifting [Les14q]. shapes [Sho11y, Sho12-34, VPC^{+}14]. Shaping [CMT14, YSO^{+}11]. share [Les12-37, Sho12-52, Sho13b]. sharp [DYP14, Sed14o]. SHC1 [MBM^{+}10]. shear [LCK^{+}13]. shedding [DYI^{+}13]. sheet [Sho11-39]. Shelterin [Les13-34, FTJG13]. Shh [LAH^{+}12]. shift [CZ10]. shifting [Les14q]. shifts [Les10v]. shine [Les13-36]. Shining [Les11-37]. Shinya [Sed11r]. SHIP2 [NPL^{+}10]. shipment [Les11-27]. ships [Sho11y]. shock [KUH^{+}14, Sho13w]. shocking [Les14s]. short [CPT^{+}12, CPT^{+}14, Les11b, Les12s, Les13y, MTM^{+}10]. short-term [CPT^{+}12, CPT^{+}14]. shortages [Les14l]. shortcut [Sho14-41]. shortening [EM11]. show [Les11x, Sho14-32]. shows [Les12q, Les13k, RCC^{+}12, Sho12g, ZKR^{+}11]. Shp1 [KBW^{+}10]. Shp1/p47 [KBW^{+}10]. Shp2 [ALS^{+}13]. shrinkage [YWC^{+}13]. Shroom [dMSMZ14]. Shs1 [GBL^{+}11, Sho11-53]. shuts [Sho13a]. shuttle [Les14v]. SIAH1 [LKLA12]. Sicily [ZLJ^{+}13]. sick [Pal10]. side [Sho14-61]. sides [Les10j, Les11-47, Les12n, Sho12p]. Siekevitz [Sab10]. Signal [Fin11, RCFH10, BBW^{+}14, HSKAT11, KBAW^{+}11, LAR^{+}10, NCML^{+}12, PBPW^{+}14a, PBPW^{+}14b, SLM^{+}11, TMG12, TLTW10, VWD^{+}13]. signal-stimulated [TLTW10]. signaling [AMH11, ATU^{+}12, AKB^{+}13, AFRZ^{+}14, ABP^{+}12, BLO^{+}12, BWL^{+}13, BMG14, BVL^{+}12, BAY^{+}11, Bez12, BPMK^{+}14, BPL^{+}11, BAS^{+}14, CDH^{+}14, CO13, CMS11, CSP^{+}10, CHL12, CLL^{+}10, CG12a, CFLDM11, CLZ^{+}14, DKA^{+}13, ETC^{+}12, FWM^{+}10a, FHD^{+}12, FBR^{+}10, GBJ10, GDO13, HCC^{+}10, HCP^{+}13, Ish14, JB12, JOR^{+}11, JDHS10, JYRL^{+}13, KUN^{+}13, KDIE11, KWH^{+}10a, KWH^{+}10b, KWO11, LPG^{+}10, LCBG^{+}11, LBS^{+}13, LLH13, LSCF11, LAH^{+}12, LCK^{+}13, LWB^{+}14, LXTM12, MGT^{+}10, MVR^{+}10, MWG^{+}12, MDW^{+}13, NLP^{+}10, NSSF10, OMZK14, OPM^{+}12, PoLC^{+}13, PGCY12, PPV^{+}14, PBD^{+}13, PPG11a, PPG11b, Pri14, RBA^{+}11, SNR^{+}11, SYK^{+}11, Sho10z, Sho11-28, Sho11-43, Sho14-38, SKV^{+}11, SWV^{+}10, SHS^{+}13, TKS^{+}13, TSH^{+}14, TSB^{+}14, TPM^{+}12, TQS^{+}11, TLS10, VES^{+}11, VMNLB^{+}11, WHWS12, WHA^{+}13, XTX^{+}13, YYM^{+}11, YMU^{+}10, YMU^{+}13, YYA^{+}11, YFLH12, ZFP^{+}13, ZKW^{+}13, ZLH^{+}14, ZME^{+}14, MWZ^{+}11]. signalosomes [KPJ^{+}13]. signals [DWM^{+}12, FSLM11, LLS^{+}11, Les11-48, LZW^{+}10, LBWS10, MKS^{+}13, SOW^{+}11, Sho14-49, TMS^{+}12]. silencing [EUB^{+}14, ECK^{+}12, HKN^{+}11, KT10, KHW^{+}10, KK13b, LDL12, RDLT11, SZW^{+}11, YOA^{+}11, BLM^{+}11]. Silent [Les13-35, Sho12-60]. silico [SSW^{+}13]. Silicon [Les13-36, NRK^{+}13]. Silver [Sed12s]. similar [BKK^{+}10]. SIMPLE [LCL12, Sho12-55]. Simulating [Sho12-56]. SIN [RKG^{+}10]. Single [DSP11, ABD14, BGS13a, BGS13b, CLEZ12, DE10, KSP^{+}11, KHB^{+}11b, LBS11, NRK^{+}13, SP11, SGT^{+}13, Sed10l]. single-cell [BGS13a, BGS13b]. Single-channel [DSP11]. single-headed [KHB^{+}11b]. single-molecule [NRK^{+}13]. Sip1 [LNL11, Les13-37, RSB13]. Sir3 [RDLT11, Sho11-54]. siRNA [HKN^{+}11]. siRNA-mediated [HKN^{+}11]. SIRT [Sho10-67]. SIRT-ified [Sho10-67]. SIRT1 [PHD^{+}10]. SIRT2 [SLK^{+}13, RNS^{+}14]. Sirt2/mammalian [RNS^{+}14]. SIRT3 [RNS^{+}14]. sister [OYH13, STI^{+}11]. sisters [Les13h, Les14-28]. site [BVC^{+}11, BBW^{+}13, DHVK10a, DHVK10b, IM11, IP12, KLP14a, LvBG^{+}10, Oef10, PvdLA^{+}14, BHA^{+}12]. sites [ATW^{+}10, BVM^{+}11, CSG14, CWPW11, DK10a, DPB^{+}10, EJBW12, FWJ^{+}11, GSW^{+}11, HSY^{+}14, LLH13, PGP14, Pri14, SYS^{+}14, STG13, SDC10, SMMB11, VWC^{+}13, lDSB^{+}10a, lDSB^{+}10b]. situ [LRB13, OBM^{+}10]. SIVA1 [HLH^{+}14]. Six1 [LGM^{+}12]. Sixt [Sed13r]. size [ANT^{+}12, BGB^{+}13, Bra13, CWL^{+}11a, EL14, HH14b, KHG^{+}13, Les10-34, Les13-28, MSK^{+}13a, Sho14-29, TAC^{+}13, ZDM^{+}14]. size-dependent [CWL^{+}11a]. Ska [CJNS12, Les12c]. SKAP [SKH^{+}10, DLBG11]. SKAP55 [OLB13]. skeletal [BWL^{+}13, CTL^{+}10, CCM^{+}11, DZT^{+}11, ELH14, GSB^{+}13, GSC11, GLM^{+}10, GF11, LGM^{+}12, MBK^{+}10, PGB^{+}10, RMM^{+}10, RCBY^{+}12, SJZ^{+}10, SGD^{+}10, VGL^{+}14, WWT^{+}12, WCQ^{+}13]. skeletogenesis [BPMK^{+}14]. Skeleton [Les10-35, Les10-29]. skin [AAE^{+}14, Les11-28, Les11-48, Les12x, Les13a, OLT11, Sed12j]. Skp1 [DK10a]. sky-induced [FUK^{+}14]. SLAIN2 [vdVMG^{+}11]. SLD [GZR^{+}14a, GZR^{+}14b]. SLD-2 [GZR^{+}14a, GZR^{+}14b]. sleeping [Sho11-57]. Sli15 [NCT^{+}11]. Sli15/INCENP [NCT^{+}11]. Slicer [HLS^{+}14]. slide [Les13-30, Sho10d]. sliding [FWM^{+}10b]. Slit [CO13]. SLK [MdFF^{+}14, VOSB12]. SLK-dependent [MdFF^{+}14]. SLM1 [IIWS14]. sloppy [Les11-46]. Slow [Les12-29, Sho13-31]. Slug [LNL11]. SM [BJ12]. SMAD1 [LSGVM14]. SMAD1/5 [LSGVM14, ZKW^{+}13]. Smad3 [MSC^{+}10, WWT^{+}12]. small [AA13, CM12c, DMK^{+}12, KUH^{+}14, LWZ^{+}10, LZW^{+}12, LZW^{+}13, STD^{+}10, Sed11s, TPM^{+}13, Sed13e]. small-molecule [DMK^{+}12]. Smaug1 [BLM^{+}11]. SMC1 [YTT^{+}10]. smell [DN10]. SMK [KHW^{+}10]. SMK-1 [KHW^{+}10]. SMK-1/PPH-4.1-mediated [KHW^{+}10]. Smn [SYK^{+}11]. Smooth [Les10-36, LXTM12, RSRK13, CYN^{+}13, MH14, dJPAA^{+}11, QECC10]. snack [Sho14u]. Snail [LNL11]. Snail1 [KKL^{+}11, LZR^{+}11]. Snail1-dependent [KKL^{+}11]. Snail2 [SMB12]. SNAP [WWB^{+}10]. SNARE [BJE^{+}12, HWE^{+}12, JOR^{+}11, JC10, KKUG11, Krä13, SRKR10, WEK^{+}14, WMV^{+}14, WGM^{+}12, ZPB^{+}12]. SNARE-mediated [BJE^{+}12, JC10, WMV^{+}14]. Snc2p [SYH^{+}13]. Snf7 [WAW^{+}11]. SnoN [ZKW^{+}13]. snoRNP [BCB^{+}14b]. SNP’d [Sho14-50]. snRNA [SIO10]. snRNPs [HIB^{+}10]. SNX [CLW^{+}14]. SNX17 [SHBC12]. SNX18 [KSR^{+}13a, WJW^{+}11]. SNX27 [LMT^{+}10]. SOAR [JAM^{+}13]. SOCS [KLC^{+}10]. soeur [Sed13b]. soft [RSD^{+}12]. softness [HSI^{+}14]. solid [Kin13]. solid-state [Kin13]. solidifies [Les13a]. SOLO [Sho10-59, YTT^{+}10]. Soluble [DBUT13, BKS^{+}13, MLSM^{+}11]. soma [MWH12]. somatic [HBG^{+}11, QB12, SPD^{+}13]. some [Sho10q]. sorted [Sed14t]. Sortilins [Sho13-50]. Sorting [FWJ^{+}11, ASE10, Bon14, BKK^{+}10, BKT13, DCL^{+}12, ECJB10, oHXK^{+}12, HKN^{+}11, HKR^{+}10, KBC^{+}14, LMT^{+}10, MBCKD13, SPF11, SHBC12, vBAK^{+}12]. SOS1 [HHC^{+}11]. source [YKW^{+}12]. sources [Les14-29]. sourcing [Sho12w]. sows [Sho12-40]. Sox10 [FSK^{+}10]. SOXC [BPMK^{+}14]. SoxE [LTJN^{+}12, Sho12-45]. space [HA12, Sed14g, Sho11z, WtLK^{+}13]. spaces [HCP^{+}13, Sha10]. spacing [FS14, Sho14q]. SPARC [RB11]. Spare [Les12-30]. sparks [Les10k]. sparse [JKS14]. Spastin [Sho13-51, LvDG^{+}10, ALF^{+}13]. spastin-mediated [LvDG^{+}10]. Spatial [BLT^{+}11, MPRT11, RY11, KSW^{+}11, LAR^{+}10, RBB^{+}14]. spatially [AFM^{+}13, BCB14a, BHB^{+}11, FPAM13, HARS14, WHDR^{+}10]. Spatiotemporal [LJW13, KNOM11, RMT13]. spatiotemporally [LCK^{+}13]. SPCA1 [KBC^{+}14]. specializations [YZM^{+}12b]. species [Fin11, MOZ^{+}13, QWL^{+}11]. Specific [VWD^{+}13, AKA^{+}13, BRD^{+}13, CM12c, DGS^{+}11, DKY^{+}12, FMG^{+}11, HPB10, HS10b, IIWS14, ILD^{+}10, LBS11, LT11, MLY^{+}10, NNSH11, OBC14, SCL^{+}14, SWC13, VEDBC13, WMB^{+}10, ZLFC14, ZCB^{+}10a, ZCB^{+}10b, ALSN^{+}11, HHC^{+}11]. specifically [FSOL14, KIOY10]. specification [BWS^{+}10, BTC^{+}11]. speckles [HPB10]. Spectraplakins [SWF12]. Spectrin [LgLM^{+}10, ZSZ^{+}13, HAB14]. spectrins [EEP13]. spectrometry [WM10]. spectrometry-based [WM10]. spectrum [HWS14]. speed [Les14-33, MFA^{+}14]. Spef1 [WMP^{+}14]. sperm [ADF^{+}12, BBY^{+}12, DSB^{+}14, KAS^{+}12, Les12a, Les12-34, Les14n, Sho12-39, Sho14-44, Sho14-55]. spermatocytes [MZP^{+}10]. Spermidine [MMB^{+}11]. spermiogenesis [YOA^{+}11]. Sphingolipid [Les11-38, Sho11-55]. Sphingomyelin [Sho14-56, vGCMA^{+}14]. Sphingomyelinase [Sho11-56, KPC^{+}11, Sho10e, TID^{+}10]. Sphingosine [YBN^{+}11, GFSR11]. sphingosine- [GFSR11]. spin [Sho11-37]. spinal [LSGVM14, RHKB12, SYK^{+}11]. Spindle [SFK^{+}13, Sho14-58, SHN^{+}11, VWC^{+}13, BKS14, BKG10, BGB^{+}13, BCB14a, BKP11, BM11, BKK^{+}10, CMS11, CKO^{+}10, CMS^{+}14, CHL^{+}14, CSTBM^{+}10, COW13, CSEH12, DP10, DWDW12, DHL^{+}12, EM11, EHUD14, EUB^{+}14, ECK^{+}12, GMD^{+}10, GS11, HH14b, HWB^{+}13, JVS^{+}14, KOK^{+}13, KBG12, KWL^{+}12, KWTR10, Les10e, Les14a, LHS10, LDL12, LHN10, MTG^{+}11, MdFF^{+}14, MGT^{+}10, MGS14, MWP^{+}12, MOZ^{+}13, McN13, MCHCC10, MKH^{+}14, NGM12, PoLC^{+}13, PJS^{+}11, PL10, QMHM10, RTC^{+}13a, RTC^{+}13b, RZA^{+}13, RDPG14, RFVE^{+}10, RGF^{+}10, RHK11, SMdP^{+}14, SA10b, STD^{+}10, Sed10h, Sed12c, Sed14i, SLM^{+}13, Sho10i, Sho10-34, Sho10-51, Sho10-58, Sho13c, Sho13j, Sho14-57, SMS^{+}14, SSK^{+}14, SSH^{+}13, TGG^{+}11, TSL12, TL12, UG10, UTK^{+}13, VSMC11, VCF^{+}13, VvDV^{+}10, WBS^{+}12, WBMCSS13, WJPD11, WDB10, ZBBG10, ZZW^{+}10, ZSK^{+}13, dSMSS13, NCT^{+}11]. spindle-associated [DHL^{+}12]. spindles [LHS10, RW10, Sho11-46]. spine [EBB13, LSM^{+}11, MDP^{+}10]. spines [HH10, MRR^{+}12]. spinocerebellar [LgLM^{+}10, Orr12]. spiral [SSZ^{+}14]. splendored [Sed11j]. splice [DAB^{+}11, JLVH12, SWC13]. splice-specific [SWC13]. splicing [HIB^{+}10, IIWS14, KNsMK13, Les14-31, LT11, MVC^{+}11, NKH11, OBC14, RCBY^{+}12, SBR^{+}11, VBB^{+}10]. split [JDB^{+}12]. SPOC [Les10k]. spoil [Les12-30]. spoke [SDS^{+}12b]. spokes [OYYK14, PBM^{+}11, Sho11-37]. spontaneous [WEK^{+}14]. spot [Sho11-45, Sho12j]. SPP [Les14-30]. spread [Sho13n]. Spreading [Les13-38, CGW^{+}11, DV10, Les14-32, PPV^{+}14, WLW11]. spring [Sho11n, SHV^{+}11, SHV^{+}13]. sprouting [XYM^{+}10]. SPRY [KLC^{+}10]. SPSB2 [KLC^{+}10, Sho10-60]. spurs [Les12-33]. squeeze [Les13i, Sho11-55]. SR [NKH11]. SRC [HHL^{+}11, RC12, TSB^{+}14, YFLH12]. SRC-2 [HHL^{+}11]. SRC-3 [HHL^{+}11]. SREBP [HVDG13]. SRF [RBP^{+}13]. SRF-mediated [RBP^{+}13]. SRGP [Sho10-61, ZBJL^{+}10]. SRGP-1 [Sho10-61, ZBJL^{+}10]. SSB [KPI^{+}10]. Ssd1 [KKK^{+}11]. SSX2IP [BIY^{+}13, Les13-39]. stability [AGL^{+}14, ACS^{+}13, BBW^{+}13, BRP14, CZD^{+}13, DKMK^{+}11, GZLG11, HLN^{+}10, HTT13, LHS10, NB12, RBF^{+}12, SNZVK12, SNZVK13, SCL11, XSJ^{+}10]. stabilization [FCE^{+}12, GCR^{+}13, MGG^{+}12, MDP^{+}10, PKG10, PVM^{+}12]. stabilize [CYLMM13, FSOL14, RGB^{+}13, WHL^{+}12]. stabilized [MSS^{+}10]. stabilizes [BKBR11, CHK^{+}10a, CHK^{+}10b, JK10, MGK^{+}12, RSM^{+}13, RBH^{+}12, SDC10, SMS^{+}13, TB13, WLN^{+}14]. stabilizing [Ari10, bCAH^{+}11, PSK11, UAH^{+}12, WMP^{+}14]. stable [ENG^{+}12, KSS^{+}11, TP13]. stacking [XW10]. stage [FSK^{+}10]. stages [Les13-43, RMT13]. Stainier [Sed13h]. stalled [DPB^{+}10, MFR^{+}14, RZF^{+}11, VYM^{+}10]. stand [Sho14-31]. standard [Sho11-52]. stands [Les12k, Sho12b]. star [Sho11h]. Stardust [KBKW10]. Starting [DSMB13]. starts [Sho11-39]. starvation [CZM^{+}14, CGCP^{+}14, ONNB^{+}14]. starvation-induced [ONNB^{+}14]. STAT3 [SFB^{+}12]. state [GC13, Kin13, PM13, SYS13, Sho14p, SJRV14, WWB^{+}10, ZBBG10]. stationary [OZT^{+}13, WMC14]. status [LLS^{+}11, MS14, SSB^{+}10]. Staufen1 [RCBY^{+}12]. Stay [Les12-31, Les13q, Les13r, Les13-31, Sed12e, Sho13-43]. Stay-at-home [Les12-31]. stays [Sho11-45]. Ste20 [YCP10]. steady [Les12-29, PM13]. steals [Les12d]. steep [HRK13, VSG^{+}12]. steering [Sho14-55]. Stefano [Sed12u]. Steitz [Sed11j]. Stem [Sho13-52, AIJI11, BT12, BvMD^{+}14, BTC^{+}11, BDvdK13, ECC^{+}13, FP10, GCP^{+}14, Gol12b, HZS^{+}10, HSJ^{+}13, HK14, LSGVM14, LGM^{+}12, LR13, Les10p, Les14b, LZW^{+}10, LR11b, NSZ^{+}13, OS13, OLT11, PWP11, QB12, Sed12d, Sed12g, Sho10-56, Sho11-47, SL14, VB12, WPM14, WBBD14, Yam13, ZLW^{+}13, Sed12o]. Stenmark [Sed11h]. step [BTC^{+}11, Dun11, ECJB10, NOS^{+}14, Sed12t, Sho14e]. Stephen [Sed13u]. steps [BPDB^{+}11, GDS^{+}12, Kik13, LMS^{+}13, LSE^{+}10, MGR^{+}10, SKN^{+}12, TH11, YKW^{+}12, ZZS13]. sterol [WMC14]. sterol-enriched [WMC14]. sterols [dSJDD^{+}11]. stick [Sho11-54, Sho14-44]. stickiness [Les14-29]. sticks [Les12j]. Sticky [BVC^{+}11, EKJH13]. Sticky/Citron [BVC^{+}11]. Stiff [Les11-39, RSD^{+}12, Sed11t, Sho14-28]. stiffening [KUH^{+}14, LMS^{+}10b]. stiffness [HSI^{+}14, JKA^{+}10, NBDB12]. STIM1 [HIM^{+}10, JAM^{+}13, SLH^{+}14]. STIM1- [SLH^{+}14]. STIM1L [DAB^{+}11]. stimulated [CWZ^{+}12, GNHB11, TLTW10]. stimulates [BvMD^{+}14, CLS^{+}10, LvDG^{+}10, LSCF11, WKGB^{+}10]. stimulating [HK14]. stimulation [HSS^{+}13, MSR10, MBO^{+}14, SPC^{+}13]. stimulus [KNOM11]. stimulus-induced [KNOM11]. sting [Sho12-63]. Stochastic [LvBG^{+}10, SKN^{+}13, GCR^{+}12]. Stoichiometry [CC12, DMH^{+}12]. stop [Les10-31, Les12d, Sho13g]. stops [Les11-34, Les13n]. storage [CWZ^{+}12, PHB^{+}13, PBvdS12, Sho13-30]. store [SPC^{+}13, YSM10]. store-operated [SPC^{+}13, YSM10]. stored [DMD^{+}12]. storehouse [Les14-31]. stories [Sed10l, Sed12d]. story [Sho10-62, Sho12p]. strand [BNDB^{+}14, BCJ13, CCJ^{+}12, CWG^{+}11, GBJ10, IAMH10, JRC^{+}13a, JRC^{+}13b, KK13b, PLC^{+}11, PLL^{+}12, Sho11-41, YTM^{+}11]. strands [Sho10-45]. strategies [KLHS14]. strategy [Sho12m]. streak [Sed14c]. strength [BvMD^{+}14, CSM^{+}12, DWPC^{+}11, LSGVM14, SSD^{+}14, VWD^{+}13, dJSTV12]. strengthen [Les10e]. strengthens [TB13]. Stress [LKSG13, NKH11, Sho10-63, Sho12-57, BGS13a, BGS13b, CTW^{+}10, DPZ^{+}14, DHB^{+}14, EAB^{+}14, HBC^{+}11, JRC^{+}13a, JRC^{+}13b, Les10i, Les11-30, Les14s, LSOT10, LCK^{+}13, MMO^{+}14, MSK^{+}13b, NZHL13, OBSG12, PW12, PXZ^{+}13, RKS^{+}10, RPK^{+}11, SFJ^{+}14, SRBL13, Sho12o, YCP10, BW13]. stress-induced [CTW^{+}10]. stress-inducible [EAB^{+}14]. Stress-responsive [NKH11]. stresses [RFL13]. stretch [Car12, Les11t]. stretching [SSH^{+}13]. striatal [JCL^{+}11]. striated [GBK^{+}14, MMVK^{+}12]. Stringent [BGC^{+}10]. strings [RKK^{+}14, Sho14-46]. stroke [CYN^{+}13]. stromal [SHC^{+}10]. Structural [GLG12, KWL^{+}12, OT11, PMK^{+}13, RPO^{+}14, SHS^{+}12, SSZ^{+}14, YZM^{+}12b, HSTF13, LAO^{+}10, LRB13, SDS^{+}12b, SKFH11]. Structure [FMPS^{+}12, OBM^{+}10, WS10, BCB^{+}14b, BSP11, CDH^{+}14, CMS^{+}14, DT14, FP10, GBiY^{+}14, HFB^{+}10, Les12w, LAH^{+}12, NRM^{+}12, SKM10a, SNSyN13, SZJ^{+}10, SBP^{+}10b, TO12, VEDBC13, WZHV11, WD11]. structure-specific [VEDBC13]. structures [BYY^{+}12, CPS^{+}13, CWL^{+}11b, DWL^{+}11, HSI^{+}11, KNPK^{+}10, PASG^{+}12, SRBL13, TTC^{+}14, TKB^{+}14]. strung [Les11r]. STT3B [Les13-40, STG13]. Stuck [DC12]. studied [OBM^{+}10]. studies [Sed12c]. study [Sed11c]. style [Ros10a]. subapoptotic [NTSK14]. subcomplex [MLW13, OTLH10]. subcomplexes [MMC^{+}10]. subdomains [FDB^{+}13]. Subgroup [BRF^{+}10]. submembranous [JKS14, KWH^{+}10a, KWH^{+}10b]. submitochondrial [LWBH12]. subpopulation [NNSH11]. subpopulation-specific [NNSH11]. subprocesses [WDB10]. subsets [YHK10]. substance [Ros10a]. Substrate [AMGC14, NBDB12, LHD^{+}14, MLBY^{+}10, PTS^{+}10, RKS^{+}10, RKG^{+}12, ZSK^{+}13]. Substrate-gated [AMGC14]. substrates [BGC^{+}10, MHK^{+}10, Sho14-62, WTBM12]. Subunit [GBL^{+}11, AMGC14, BKBR11, BHA^{+}12, CLM^{+}10, DCP^{+}10, KWL^{+}12, LH11, SBP^{+}10a, TPZ^{+}14, VYC^{+}11, WEK^{+}14, ZYF^{+}11, ZLJ^{+}13]. Subunit-dependent [GBL^{+}11]. subunits [ALSN^{+}11, OCF^{+}10, Oef10]. Subversion [AMR11]. successfully [Sho11-38]. Sue [Sed12v]. sufficient [BRL14, BKS^{+}11, DBUT13, OBSG12, SP11]. SuFu [TLS10]. sugary [DYP14]. suggests [ANT^{+}12]. sulfate [BST^{+}11]. SUMO [DPV^{+}12, Les11-40, MAD10, PHW^{+}13, Sho12e, Sho12-45, SLC^{+}13]. SUMO-binding [DPV^{+}12]. SUMO-targeted [PHW^{+}13]. SUMOs [Les10-34]. SUMOylated [HLT12, LTJN^{+}12]. Sumoylation [KMG^{+}11, SFJ^{+}14, DPV^{+}12, HLN^{+}11, LZW^{+}12, LZW^{+}13, RFK^{+}10, TAGJ11]. SUN [CTM^{+}14a, CSS^{+}14, FTJG13, GSS^{+}11, Sho14-59, TCB^{+}14, ZGEM12, ZGW^{+}14]. SUN-1 [FTJG13]. SUN-interacting [ZGW^{+}14]. SUN1 [MSZ^{+}12, TH11]. Sup35 [KNPK^{+}10]. super [RKT^{+}14, SHL10]. super-resolution [RKT^{+}14, SHL10]. superfamily [RCC^{+}12, RFRV12a, RFRV12b]. supply [MFA^{+}14]. support [Les11x, Sho13-34]. supporting [Les11-36]. supports [ELH14, LLT^{+}12, PRFF13, SNZVK12, SNZVK13, SHB^{+}10]. suppress [HZS^{+}10, HDH^{+}10, ZLFC14]. suppresses [BGB^{+}13, BNM^{+}14, BMS^{+}11, BDN^{+}13, HRWW^{+}13, JGB^{+}13, MTM^{+}10, PYT^{+}13, TPZ^{+}14, XHB^{+}10]. suppressing [CMD^{+}13, KZR^{+}12]. suppression [FAB^{+}10, LMS^{+}10b, RSS^{+}13, RJM^{+}12, RDPG14]. suppressor [BDR^{+}12, HDH^{+}10, NBSE^{+}13a, NBSE^{+}13b, TQM^{+}14, dSMSS13]. supramolecular [TTC^{+}14]. surface [OVW10, RKK^{+}14, WVvG^{+}13]. surfaces [Les14w]. surprising [ACS^{+}13, Sho14-60]. surveillance [GEN14, SIO10]. survival [AMO^{+}11, CCGN11, EAB^{+}14, FHD^{+}12, HSI^{+}14, HLL^{+}12, KZR^{+}12, KPH^{+}12, MBM^{+}10, SOW^{+}11, SEV^{+}14, ZPS^{+}10]. suspect [Sho11-61]. suspects [Sho10-28]. sustain [Sho14-58]. Sustained [HTS^{+}10, BLO^{+}12, XTX^{+}13]. Suzanne [Sed13v]. Sweet [Les10-37]. switch [BWL^{+}13, BVL^{+}12, HHC^{+}11, KK11, Les12-28, Les13l, MGK^{+}12, MS14, RZA^{+}13, RSB13, Sho10x, Sho12a, Sho14-57, WBMCSS13]. switches [BT13, CLM^{+}10, Les11-47, MBO^{+}14]. Switching [Les11-41, Sho13-53, HTT11b, SYV14]. sword [AMO^{+}11]. Syb [Mit12c]. SYD [KHG^{+}13, OFS^{+}10]. Syd-1 [OFS^{+}10]. SYD-2 [KHG^{+}13]. symmetry [SRP^{+}13, Yam13, YRU^{+}13]. sympathetic [OMZK14]. symphony [Sed12i]. symptoms [XHS^{+}13]. synapse [BLM^{+}11, BRP14, CLD11, GTS10, KK13a, KSLF^{+}11, PYT^{+}13, SHS^{+}13, TQS^{+}11, WKN^{+}13]. synapses [Ari10, LLT^{+}12, NLJ^{+}13]. synapsin [XHS^{+}13]. synapsin-1 [XHS^{+}13]. Synaptic [HSS^{+}13, Sho10-64, XHS^{+}13, BMG14, CG12a, FUK^{+}14, FBAO^{+}13, HWE^{+}12, MSK^{+}13a, MRR^{+}12, RBA^{+}11, SBS^{+}12, Sho13f, Sho13-53, SSK^{+}13, SCR12, VPC^{+}14, VG13]. Synaptobrevin [WWB^{+}10, SYH^{+}13, HWE^{+}12, ZPB^{+}12]. synaptogenic [CB12]. synaptotagmin [FCA10, WLGC11]. synchronizes [DYS^{+}14, GP10]. syncytial [AGL^{+}14, TGES12]. syndapin [SSL^{+}14]. syndecan [BST^{+}11, CLC^{+}11, MBR^{+}11, Les11-42, PCO^{+}10, Sho10-65]. syndecan-1 [MBR^{+}11]. syndecan-2 [CLC^{+}11]. syndecan-3 [BST^{+}11, Les11-42, PCO^{+}10, Sho10-65]. syndrome [CHK^{+}10a, CHK^{+}10b, CWS^{+}11, KTN^{+}12, Sho14-50]. synergizes [OBD^{+}10]. Synergy [HSKAT11]. Syntaphilin [Sho13-54, CS13]. syntaxin [SRKR10, LHL11, WWB^{+}10]. Syntaxin17 [TNV^{+}13]. Syntaxin17-dependent [TNV^{+}13]. synthase [BGC^{+}14, RNS^{+}14, WHF^{+}11]. synthesis [BBW^{+}13, GKR^{+}11a, GKR^{+}11b, GRH^{+}12, HS10b, Les14i, MJJ^{+}10, NBC^{+}12, OCF^{+}10, VYM^{+}10, ZDS^{+}12]. Synthesizing [Sed12s]. synthetase [KHFV^{+}13]. Synuclein [PMB^{+}11, Sho10a, WCC^{+}10, KMG^{+}11]. System [ZKR^{+}11, BSO^{+}14, CVJ^{+}11, CC12, CGK13, Coo13, HGV^{+}14, HKR^{+}10, Kin13, MFGB10, MBK^{+}10, NT11, PvdLA^{+}14]. Systematic [AMS^{+}13, RTM13]. Systems [MRA14, SA10b, Sho12-49]. Syx [NGL^{+}12]. Syx’d [Sho12y].

T [Sho13-55, TPM^{+}12, ADAB^{+}12, BLO^{+}12, BMS^{+}11, HCC^{+}10, LCBG^{+}11, Les11n, OLB13, SDN^{+}14a, SDN^{+}14b, Sed13j, SHS^{+}13, YWC^{+}13, ZBBG10, ZEG11]. T-cell [SHS^{+}13]. T-loop [ZBBG10]. TACC3 [HWB^{+}13, LHS10]. tafazzin [CWS^{+}11]. tag [Les14h]. tagging [CTM^{+}14b, Ish14]. Tail [YHF13, BBW^{+}14, CNP^{+}12, FLN^{+}10, FLN^{+}16, FLVP10, HFB^{+}10, JLVH12, Sho10-40, Sho13-59]. Tail-anchored [YHF13, BBW^{+}14]. tails [OVL11]. TAK1 [MBO^{+}14]. take [Les10j, Les11-38, Les13b, Les13g, Pal14, Sho13-31]. takes [HBG^{+}11, Les10l, Les10-41, Les11-36, Sho10v, Sho12-30, Sho12-63, Sho13z, Sho13-40]. taking [SF12, Sho10-35]. Talin [BWBC^{+}14, KYHG12, Les12-32, Les14-32, PPV^{+}14, Sho12-58, LLU^{+}12a, LLU^{+}12b, SF12, WBS11, WWM^{+}12, ZSK12, Sho14-61]. Talin-bound [PPV^{+}14]. talin-mediated [WWM^{+}12]. talk [KKY^{+}14, RKG^{+}10, SLS^{+}10, TPSS12]. tall [Sho14-31]. Talpid3 [KKL^{+}14]. tandem [dJPAA^{+}11]. Tangled [Les10-38]. TANGO1 [WPL^{+}11a, WPL^{+}11b]. Tannishtha [Sed12w]. Tao [GWR12, CM12a]. TAp73 [HKR^{+}14]. Tapping [CC10b]. Tara [YYA^{+}11]. target [DK10a, KKMB10, LvBG^{+}10, Sed13j, Sho13v, SGD^{+}10]. Targeted [PGB^{+}10, HSR^{+}10, PHW^{+}13, LK12]. Targeting [GSS^{+}11, HJ14, LSM^{+}11, RBS10, SOW^{+}11, SW12, ADS^{+}13, BKAB13, BRD^{+}13, BRP14, CDAK10a, CDAK10b, CM12c, DGS^{+}11, FLN^{+}10, FLN^{+}16, FLVP10, IM11, KdKDP12, LAR^{+}10, LZW^{+}12, LZW^{+}13, LHS10, LJLJ11, LVB^{+}10, MMS^{+}10, MPRT11, PDKG14, SAoS14, SKH^{+}10, Sho11-49, SWC13, VYC^{+}11, WSZ^{+}12, WLW11, vGLWB12, DS12]. targets [COG11, CWL^{+}11b, FHY^{+}10, FCA10, FSOL14, KLC^{+}10, LP13, MKH^{+}14, SHC^{+}13, VFNR11, WTBM12, YHF13]. Tarp [MBM^{+}10]. taste [CR10a, CR10b]. Tat [AMGC14, Sho12-52, Sho14-62]. TatC [AMGC14]. tau [GKWG^{+}11, Les11b]. taxol [YON^{+}12]. TBC1D10A [HMiY^{+}10]. TBC1D14 [LLR^{+}12]. TBC1D24 [FUK^{+}14]. TBC1D24/sky [FUK^{+}14]. TBC1D24/sky-induced [FUK^{+}14]. Tbx1 [CMH^{+}10]. TCR [OLB13, VWD^{+}13]. TCR-ab [VWD^{+}13]. TCR-ab/gd [VWD^{+}13]. TDRD5 [YOA^{+}11]. team [Les12h, Sho10-66]. teams [Les10e]. teamwork [Les13w]. Ted [Sed10p]. tell [Sed10l, Sed12d]. Telocentrosomes [Les13-41]. telomerase [MTM^{+}10]. telomere [HLT12, Les13-41, MFB12, MTM^{+}10, PHD^{+}10, RDLT11, YKT^{+}13]. Telomeres [Sho10-67, CSH^{+}12, DKY^{+}12, FTJG13, Les13-34, MTM^{+}10, MSZ^{+}12, SCL^{+}14, Sho11-54, YKT^{+}13]. Telomeropathies [HWS14]. Tem1 [VSMC11]. Template [FSA^{+}10b, ECC^{+}13, OBSG12, WS10, Yam13]. Template-free [FSA^{+}10b]. Temporal [KAS^{+}12]. temporally [KYOY13]. temporary [Sho12-32]. Tenascin [AEC^{+}14, Sho14-63]. Tenascin-X [AEC^{+}14, Sho14-63]. tensile [NT11]. Tension [OBSG12, BW13, CMS^{+}14, HTS11, MKS^{+}13, Sho14-58, SHN^{+}11]. tension-dependent [HTS11, SHN^{+}11]. term [CPT^{+}12, CPT^{+}14, MDP^{+}10]. terminal [AGM^{+}10, KFET11, YHT^{+}10, BBY^{+}12, CNP^{+}12, JLVH12, NvCL^{+}13, RBM^{+}11, STG13, ŻKC^{+}11]. terminally [BKE10, WWB^{+}10]. terminals [CWB^{+}14]. terminate [CDH^{+}14]. termination [BPL^{+}11]. terminus [OLB13, WBS11]. territories [BLT^{+}11]. territory [Sho11-50]. test [Sho10f, Sho12-49, Sho13-52]. Testing [Sho13-56]. testis [HKR^{+}14, HS10b]. testis-specific [HS10b]. tether [LGAC13, SSK^{+}14]. tethering [BSP11, CLM^{+}10, FBAO^{+}13, Sho10-27, TPM^{+}13]. tethers [BBD^{+}11, CTD^{+}10, GVP^{+}11]. Tetrahymena [BKT13]. tetrameric [KHB^{+}11b]. tetraspanins [DCO^{+}12, DCO^{+}16]. TFEB [MP13]. TGF [AEC^{+}14, JB12, KCK^{+}14, NLAS^{+}10, PTST12, PXZ^{+}13, RB11, Sho14-64]. TGF- [AEC^{+}14, JB12, KCK^{+}14, NLAS^{+}10, PTST12, PXZ^{+}13, RB11, Sho14-64]. TGN [LHL11, Sho12-53]. Tha4 [AMGC14]. thanks [Les12x]. their [Car12, LR13, Les10x, Les10-36, Les11x, Les13h, Les13p, Les14p, Les14-28, ME13, MC10, Mit12b, SBEM13, Sho10h, Sho10-52, Sho11h, Sho11-50, Sho11-56, Sho12i, Sho12z, Sho12-38, Sho13b, Sho13n, Sho13v, Sho13-29, Sho13-60, YBN^{+}11]. theme [CG10a]. Themes [LM13]. themselves [Les13-35, Sed14h]. therapeutic [HSF12, HA12, Ste12]. therapeutics [TPSS12]. therapy [RHKB12]. there [BH13, Sed13z, WP14]. thereby [EUB^{+}14]. thermal [GYC^{+}14]. thermotaxis [WOG13]. thermotransduction [WOG13]. thin [CZGG12, GLM^{+}10]. thing [Sed11j]. things [DSMB13, Sed11q, Sed11s, Sed13u]. think [Sed13i]. thiol [ACS^{+}13]. Thompson [Sed10c, Sed12b]. those [Les14-36]. threads [Sho10q]. three [ABVP11, HHS13a, HHS13b, LNTR14, WGM^{+}12, Les13-42]. three-dimensional [LNTR14]. threonine [CFLDM11]. threonine-phosphorylated [LCBG^{+}11]. threshold [TMG12]. throughout [CSH^{+}12]. throughput [VvDV^{+}10]. throw [Les14-36]. throws [Les13t]. Thyroid [Les14-33, ZWL^{+}14, ZGCG^{+}14]. Tiam1 [Sho12-59, WWM^{+}12]. tidies [Les10s]. tiered [FAB^{+}10]. ties [MS12, MS14, Sho11-60]. tight [IMP^{+}12, Les13a, Les14-32, MLG^{+}10, RBY^{+}11, Sha10, SEV^{+}14, YMT^{+}13]. tightens [Les10-38]. tightly [PBD^{+}13]. TIM23 [SLM^{+}11]. Tim50 [SLM^{+}11]. time [dSLPRG11, SBR^{+}11, Sho11z, WPSA13]. timely [ADB^{+}14, JDS^{+}10]. timers [Sho13-45]. Timing [BT13, DSL13, Gil10, LLM^{+}10, MGT^{+}10, MMdCOM^{+}11, PBG^{+}13, Sho10-29, TGB10]. TIMs [Sed13e]. ting [Sho10i]. Tinman [QWL^{+}11, Sho11j]. Tinman/ [QWL^{+}11]. tip [Les13-40]. Tipping [Sed11o]. tips [DWDW12, Les13s, Sed11b, Sed13k, Sho11-38]. TIRFM [AOE^{+}10, AOE^{+}12]. tissue [FMG^{+}11, KTB^{+}14, MGFG^{+}10, PSVRB^{+}11, RC12, Sho14-37, VTM14, YFLH12]. tissue-specific [FMG^{+}11]. Titin [dSLPRG11, Sho11-57, KUH^{+}14]. Tks5 [OOKH^{+}12]. Tks5-dependent [OOKH^{+}12]. TLR9 [TPM^{+}12]. TNF [MLG^{+}10, MBO^{+}14, TTC^{+}14]. TNF-induced [MLG^{+}10]. TNGW1 [CLSO^{+}12]. Tobias [Sed12x]. TOG [HWB^{+}13, FCE^{+}12]. together [Les11-33, Sed11u, Sed13a, Sho11-54, SHV^{+}11, XYM^{+}10]. TOGL [FSOL14]. tolerant [Sho14n]. Tollo [BMG14]. tomography [FBZM^{+}10, FBAO^{+}13, LRB13, OBM^{+}10, PBM^{+}11]. tones [Les11n]. tonic [Sho14-42]. tool [LC10]. tools [Gol12a, WM11]. Tooth [LCL12, Sho12-55]. top [Sho10z]. TopBP1 [GSP^{+}14, Sho11k, WGC11]. TopBP1/ [GSP^{+}14]. Topoisomerase [Les14-34, KPC^{+}10, LGAC13, RZF^{+}11, RFK^{+}10, SSV^{+}12]. topological [AOE^{+}10, AOE^{+}12]. topologically [FSA^{+}11]. topology [KYHG12]. TOR [HDK^{+}13, LP11, Sho13a]. TORC2 [CDK^{+}10]. touch [LMN10]. touches [Les12e]. tour [Pal14]. toxicity [KMG^{+}11]. toxin [GCV^{+}11]. TPR [NvCL^{+}13, SFK^{+}13, SDD^{+}13]. Tpr-mediated [SFK^{+}13]. TPX2 [HH14b, Les11-43, MTG^{+}11, ZBBG10]. Tracing [CSAPLBD11a, CSAPLBD11b, VB12]. track [DSMB13, Les10n, Sho13k]. Tracking [Sho14-65, LHW10, NRK^{+}13, vdVMG^{+}11]. traction [BMÁG^{+}14, Sho14-65, WtLK^{+}13]. trade [Sho11i]. TRAF6 [PGB^{+}10]. traffic [GNHB11, Lev11, MVP^{+}11, ME13, MMU^{+}10b, OPCEM10, RT12, Sho14c, Sho14-49]. trafficking [CTM^{+}14b, DCO^{+}12, DCO^{+}16, DKF^{+}11, GOWM12, LCL12, LAH^{+}12, NLJ^{+}13, NJS^{+}10, PH10, RCM^{+}12, RKT^{+}14, RBA^{+}11, She14, Sho12-61, TLL^{+}13, UKZ^{+}13, WMCF10, YBN^{+}11]. TRAIL [GWR^{+}10]. trans [CM12c, WVvG^{+}13, vGCMA^{+}14, vBAK^{+}12]. trans-Golgi [WVvG^{+}13, vGCMA^{+}14, vBAK^{+}12]. trans-Golgi-specific [CM12c]. transactivator [UAH^{+}12]. transcription [ALSN^{+}11, EAB^{+}14, FHA10, Les12-31, RMM^{+}10, SZE^{+}11, YYA^{+}11, ZJP^{+}10]. transcriptional [JBS^{+}12, JBS^{+}13, LTJN^{+}12]. transduction [Fin11, KBAW^{+}11]. transendothelial [HCC^{+}10, RIG^{+}12]. transfer [AiIK^{+}13, GSW^{+}11]. transferrin [LADS10]. transforming [GBSC^{+}12]. TransgeneOmics [HBC^{+}10]. transgenic [BBY^{+}12, YSN^{+}11]. Transient [DSM^{+}11]. transit [KYP^{+}14, OVW10]. transition [AEC^{+}14, BMRM13, CSEH12, CTD^{+}10, EKJH13, GBSC^{+}12, KKL^{+}11, KD11, KWK^{+}11, Les11-45, MFGB10, MSC^{+}10, PTST12, RKG^{+}10, SMB12, WLK^{+}11]. transitions [Bra13, Gil10, HVOF^{+}14, PSVRB^{+}11]. Translating [TPSS12]. Translation [LCfC11, DDH^{+}12, HHS^{+}14, KYOY13, Les11-47, MVP^{+}10, RN12, SOW^{+}11, Sho10-49, TLTW10, WKGB^{+}10, YZL^{+}13]. Translational [SL14, COG11, CMW11, EMT^{+}14, GCC12]. translocase [AMGC14, GSM^{+}12]. Translocation [LLT^{+}12, CC12, CWZ^{+}12, CG12b, JCL^{+}11, OSD^{+}14, PR12, Sed10n, TMG12, WGR^{+}12]. translocon [LJPJ11, RKG^{+}12]. Transmembrane [LJLJ11, CHS^{+}10, KYHG12, RCFH10]. transmigration [HZM^{+}13]. transmission [CWFL13, DBUT13, FSA^{+}10a, SSK^{+}13, VPC^{+}14]. transport [AMR11, BW12, BJ12, CC12, CHK^{+}10a, CHK^{+}10b, CS13, CLW^{+}14, CT10, ETRP12, EIW^{+}12, FEHF12, mFH13, KKUG11, LHL11, LWZ^{+}10, LgLM^{+}10, LBD^{+}14, MWH12, MTT^{+}14, NNO^{+}11, OTLH10, PAB^{+}10, PHB^{+}11, PSK11, RBF^{+}12, SLM^{+}11, SW10b, Sho11m, Sho12u, Sho14b, SDD^{+}13, SW12, SST^{+}12, TKB^{+}14, WVvG^{+}13, XBC^{+}13, YOMM^{+}11, HZE^{+}13]. transported [DSD^{+}13]. transporter [BVL^{+}12, Les12-30]. transporting [Sho10-36]. transposon [KT10]. transsynaptic [CB12, NLP^{+}10]. trap [Car12, CAB^{+}13, Sho13-49]. traps [BZ12, PMHZ10]. trash [Les13-29]. travel [Les14r]. traveling [Les11l]. travels [DYP14]. Traversing [KLHS14]. TRC40 [YHF13]. TRC40-independent [YHF13]. treatment [RHKB12]. tree [RFRV12a, RFRV12b]. Treslin [KSSD11, Sho11k]. TRF1 [HLT12]. Trichoplein [IMG^{+}12, Sho12-60]. trident [KIOY10]. triggered [CGK13, MWG^{+}12]. triggering [BHMBS^{+}11]. triggers [ALV^{+}12, RKG^{+}12, SGT^{+}13, TTM^{+}14, ZLW^{+}13]. Trim32 [CZGG12, CLZ^{+}14]. Trim39 [HZT^{+}12]. TRIM9 [Les14-35]. Trim’d [Sho12c]. trimming [NOS^{+}14]. Trina [Sed13w]. Trio [YYA^{+}11]. tRNA [SSK^{+}14]. Tropomodulin [GLM^{+}10, GF11]. true [Les13q, Les13r, Sho11-51]. trunk [MMdCOM^{+}11]. trypanosome [HVOF^{+}14]. Tsai [Sed12p]. TspanC8 [DCO^{+}12, DCO^{+}16]. Tuba [QMHM10, JDL^{+}14]. tube [YSN^{+}11]. Tuberous [AIJI11]. tubules [ALF^{+}13, SWS^{+}13]. Tubulin [EMO12, GCR^{+}13, LvDG^{+}10, NEMH^{+}10, Sho10c, Sho12a, Sho12b, BR14, CLS^{+}10, ENG^{+}12, GZLG11, Jan14, PMK^{+}13, Sho13-59, YON^{+}12, NNO^{+}11].

Tudor [Sho11-58]. tuft [GvEM^{+}11]. tumor [ACO12, BWBC^{+}14, BDR^{+}12, FPM^{+}14, HDH^{+}10, JCN^{+}14, Les11g, MVR^{+}10, MVN11, NBSE^{+}13a, NBSE^{+}13b, PoLC^{+}13, RSS^{+}13, RJvD11, RJM^{+}12, SHC^{+}10, Sho13o, TQM^{+}14, YST^{+}11, dSMSS13, vRJMvD10]. tumor-associated [MVR^{+}10, SHC^{+}10]. tumorigenesis [Sho10-57]. tumors [Les12t, RBS10]. tuning [FW10, KPH^{+}12, WtLK^{+}13]. tunnel [GKR^{+}11a, GKR^{+}11b, LJPJ11]. turning [Sho10-68, TKMK10]. turnover [CTH^{+}11, GB12, Sho12-59, Sho13k, SSK^{+}13, TAGJ11, XWE^{+}10]. turns [Les12a]. TWEAK [MBK^{+}10]. TWEAKs [Sho10u]. twice [Les12h]. twin [CC12]. Twins [Sho11-59, BKBR11]. twist [Sed14q, LNL11, Les11y, Sho13x, Sho14-60]. Twist1 [SPJ^{+}14]. Twist1-induced [SPJ^{+}14]. Two [KKMB10, KLvdB^{+}13, KHB^{+}11b, Les11-44, LSE^{+}10, MWZ^{+}11, TYN^{+}13, ABP^{+}14, DK10b, Dun11, ETYS^{+}12, FAB^{+}10, KFS^{+}14, Les13-43, MMC^{+}10, PGCY12, Sho11t, Sho12f, Sho13-58, Sho14e, SDS^{+}12b, WMB^{+}10]. two-step [Dun11, Sho14e]. two-tiered [FAB^{+}10]. Type [CDAK10a, CDAK10b, ISZ^{+}11, BDvdK13, CZD^{+}13, GvEM^{+}11, IIWS14, JGA^{+}11, LgLM^{+}10, SFL12, ZWL^{+}14]. types [ABP^{+}14, DZT^{+}11, YFO12]. tyrosination [PMK^{+}13]. tyrosine [MBVT^{+}13, PMK^{+}13].

U [SIO10]. UbcH10 [Sho10-57, vRJMvD10]. ubiquinone [LWBH12]. ubiquitin [BAY^{+}11, BDN^{+}13, DSW^{+}11, DCL^{+}12, FMI^{+}13, HLH^{+}14, HZT^{+}12, KLF^{+}14, KBW^{+}10, MFGB10, MMO^{+}14, MHCvSW11, PLL^{+}12, PHW^{+}13, RKG^{+}12, SP11, TTC^{+}14, vRJMvD10, RKS^{+}10, Sed12h, Sho12-61, Sho13-57]. ubiquitin-dependent [PLL^{+}12]. ubiquitin-independent [DCL^{+}12]. ubiquitin-like [KBW^{+}10]. ubiquitinated [HSR^{+}10]. ubiquitination [GHK10a, HRWW^{+}13, LNT^{+}10, LLcK^{+}11, SP11, XSJ^{+}10]. ubiquitous [Sho12-61]. Ubiquitylation [CZGG12, HBSD12, ONNB^{+}14, Sho12-62, DSK^{+}11, IAMH10, JEF^{+}11, LP13]. UBXN [KSH^{+}13]. UBXN-2 [KSH^{+}13]. UBXN-2/p37/p47 [KSH^{+}13]. UDP [PTS^{+}10]. Ugo1 [PKD^{+}11]. ULK1 [LLR^{+}12]. ULK1-positive [LLR^{+}12]. Ulrike [Sed11s]. ultra [FAvdB^{+}12]. ultra-large [FAvdB^{+}12]. ultrasensitive [PBPW^{+}14a, PBPW^{+}14b]. unattached [MKH^{+}14, PZ14]. UNC [CTM^{+}14a, WLN^{+}14]. UNC-40 [WLN^{+}14]. UNC-6 [WLN^{+}14]. UNC-84 [CTM^{+}14a]. uncertain [GK13]. uncommitted [CC10b]. unconfined [HCP^{+}13]. unconstrained [ZSD^{+}14]. Unconventional [DAS^{+}10, MALS10, Pfe10, BMC^{+}11]. uncouple [RW10]. uncouples [PLC^{+}11]. uncovers [Les12z]. undamaged [ZNA^{+}14]. under-replicated [BBW^{+}13]. underlies [CWL^{+}11a, KXN10, NNO^{+}11]. underlying [CLS13, LCP13, MMV^{+}10, RK13, SLH13]. undermine [Les10-29]. Understanding [Sho14-66, Kik13, Sho12r]. unexpected [Lev11, dSLPRG11]. Unfaithful [Les10-39]. Unfinished [Les12-33]. unfolded [AXL10, BAB12, CPT^{+}12, CPT^{+}14, RH10, SS11, WK12]. unfolding [SMZL13]. unified [SMK14]. Uniform [PvdLA^{+}14]. unique [BKS^{+}13, CNP^{+}12, DE10, PKD^{+}11, ZGW^{+}14]. united [Les12k]. units [TW14]. universal [DKY^{+}12]. University [Ros13]. unkind [Les14m]. unload [KSS^{+}11]. unloading [MFA^{+}14, SAoS14]. Unlocking [Les12-34, Sho14-67]. Unpacking [Sho10-69]. untransformed [KRS11]. unusual [NZHL13]. Unwrapping [Les10-40]. up-regulates [YYA^{+}11]. up-regulating [ZMW^{+}13]. up-regulation [MLH12, PXZ^{+}13]. Updating [Ros10b]. upon [AOE^{+}10, AOE^{+}12]. UPR [CCGN11]. uproot [Sho14-30]. upstream [LNJ^{+}13, QWL^{+}11, VKMI12]. uptake [OSD^{+}14]. uropathogenic [WHF^{+}11]. uropod [HCC^{+}10, LOR^{+}10]. use [GCV^{+}11]. Usher [SDS^{+}12a]. Using [Boe12, CAB^{+}13, DV10, DKY^{+}12, TYN^{+}13, Sed10a]. USP1 [CRJB^{+}11]. utilize [NAS^{+}11, NAS^{+}12, NAS^{+}13]. utrophin [MHC^{+}12]. UV [CRJB^{+}11, GSGL11, JEF^{+}11, LLA^{+}12, SKM10a]. UV-induced [LLA^{+}12]. UV-mediated [CRJB^{+}11].

v [WHDR^{+}10, CZC^{+}11, EJBW12, KNH^{+}10, KHB^{+}11b, PSK11, PGAE^{+}13, RNS^{+}14, VM14, WLZ^{+}14, Sho10f, WEK^{+}14]. v-ATPase [WEK^{+}14]. V0 [WWHH10]. V0-ATPase [WWHH10]. V0a1 [WEK^{+}14]. Va [CWZ^{+}12]. vaccinia [AA13]. vacillations [Sho14-39]. vacuolar [AMR11, BPL^{+}11, WMC14]. vacuole [Sho13-41]. vacuoles [EBB13, Sho13-37]. Valdivia [Sed14u]. Valentina [Sed14w]. Valerie [Sed11t]. Vann [Sed11u]. Variant [CSH^{+}12, DAB^{+}11, GSS^{+}11, JLVH12]. variants [WMB^{+}10]. variations [CG10a, LM13]. various [RMT13]. vascular [CYN^{+}13, KWK^{+}11, QECC10, SMSP11, YMU^{+}10, YMU^{+}13]. VASP [HM10, WHDR^{+}10]. VCAM [LMC^{+}12]. VCAM-1 [LMC^{+}12]. VE [HOS^{+}12]. VE-cadherin [HOS^{+}12]. VEGF [CLL^{+}10, NGL^{+}12, XYM^{+}10]. VEGFR3 [XYM^{+}10, CMH^{+}10]. ventral [MKS^{+}13]. Vernos [Sed10h]. versatile [GDS^{+}12]. versus [HCP^{+}13]. Vertebrate [JJH^{+}10, MBZ^{+}10, BV11, BW12, DWL^{+}11, Sho13-37, SPD^{+}13, SJ13, TLS10]. very [Sed12f]. Vesicle [WMCF10, AVP^{+}14, BSP11, BJ12, BLC^{+}14, CZC^{+}11, CWZ^{+}12, DYS^{+}14, FUK^{+}14, FBAO^{+}13, GOWM12, HWE^{+}12, JDB^{+}12, LSE^{+}10, NNO^{+}11, PAB^{+}10, PMP^{+}11a, PMP^{+}11b, PMB^{+}11, RMT13, Sho10-28, SHS^{+}13, SCR12, WWB^{+}10, WLGC11, WS10, WGM^{+}12, XRO^{+}11, YSN^{+}10, vdBFS^{+}12]. vesicles [BAH^{+}12, DMD^{+}12, HSR^{+}10, KMC^{+}14, MSK^{+}13a, RS13, Sed13m, SYH^{+}13, Sho10-64, Sho11c, Sho12-48, Sho13f, YKW^{+}12, ME13, Sed14n]. vesicular [KKMB10, SKN^{+}12, Sho14c, WH13, XBC^{+}13, ZPB^{+}12]. vesiculation [SWS^{+}13]. vessel [CMH^{+}10, LXTM12]. VI [TPM^{+}13]. via [ABP^{+}12, BWL^{+}13, CVR10, CNP^{+}12, CFLDM11, DPV^{+}12, DGH^{+}14, DKA^{+}13, FHD^{+}12, FAB^{+}10, GWP^{+}11, GNHB11, HBS^{+}10, HSJ^{+}13, HLN^{+}11, JGB^{+}13, JEF^{+}11, JTN^{+}13, JBS^{+}12, JBS^{+}13, JDL^{+}14, KKUG11, KA12, KKK^{+}11, KFL^{+}14, LLR^{+}12, MMS^{+}10, MMO^{+}14, MRPR12, NSSF10, Oka14, OMW^{+}14, OLB13, PSF^{+}11, QWL^{+}11, RB11, RHK11, SYS^{+}14, SPC^{+}13, SGT^{+}13, WM12, WEK^{+}14, WGR^{+}12, WHWS12, WKN^{+}13, WDB10, WWS^{+}12, XTX^{+}13, YHF13, ZWL^{+}14]. viability [HHJ^{+}11]. Victor [Sed13x]. view [Sho10-36, Sho13-32]. views [Sed12w]. VII [FCA10]. vimentin [SGLV10]. Vinculin [BPT^{+}14, HOS^{+}12, Les12-35, TTB^{+}13, lDSB^{+}10a, lDSB^{+}10b, CRP^{+}14, JLVH12, PSR^{+}10, Sho14-45]. vinculins [Les14x]. Viral [WLW11, ZDS^{+}12]. Virtual [FAvdB^{+}12]. virulence [ZDS^{+}12]. Virus [Les10-41, AA13, GSU^{+}12, GM11, Les10q, Les11-31, PASG^{+}12, SYV14, Sho13-44, TLL^{+}13]. vision [SC10a]. Visualization [GS11, dSLPRG11]. visualize [Boe12]. visualized [AOE^{+}10, AOE^{+}12, DDH^{+}12, FSA^{+}10b]. Visualizing [Sed12k, Sed14a]. vitro [BKS^{+}13, HZE^{+}13, OBM^{+}10, QECC10, SSW^{+}13]. vivo [BKE10, DHVK10a, DHVK10b, DJL^{+}12, HZM^{+}13, HKN^{+}14, HLL^{+}12, HBC^{+}10, KNPK^{+}10, KLHS14, KTB^{+}14, LRH^{+}13, LVK^{+}13, MLG^{+}10, MHV12, OD10, PR12, QECC10, SMM^{+}10, ZQA^{+}14]. voltage [BNM^{+}14, Sed14a]. voltage-gated [BNM^{+}14]. volume [Les13v]. Vousden [Sed12m]. Vpr [Sho11-60, STI^{+}11]. Vps [KWDD10]. Vps1 [CLW^{+}14]. VPS35 [Les11-45, WTH^{+}11, XTX^{+}13]. Vps4 [AVP^{+}14]. Vps41 [CLM^{+}10, Sho10-27]. Vps45p [BJ12]. Vps74 [CDH^{+}14, Sho14-68]. VSV [LAO^{+}10]. vulnerability [ZZW^{+}13].

Wagner [Sho10p]. Wait [Les12-36, Les14-36, Sho13j]. walkabout [Sed14l]. wall [MCS^{+}13]. Wallenda [XWE^{+}10]. Wallenda/DLK [XWE^{+}10]. walls [KOK^{+}13]. Walter [Sed11n]. Walther [Sed12x]. wander [Les13-44]. Warren [Sed10f]. WARTS [CSS^{+}12]. WASH [CZC^{+}11, Les11-46, MRCC^{+}13]. WASP [Les11v, Les12u, NBS^{+}11, Sho12-63, VKMI12, YZM^{+}12a, BHMBS^{+}11]. WASp-deficient [BHMBS^{+}11]. wasting [PGB^{+}10]. watch [Sho12-47]. WAVE [LVK^{+}13]. waves [APC^{+}13, Sho13q]. way [Les10d, Les12f, Les14d, Sed11f, Sed11r, Sed13c, Sed13r, Sed14m, Sho10-33, Sho14h, Sho14-54, Six12, Yam13]. ways [Les10x, Les14-28, Sho12-27]. WD40 [HKI^{+}13]. wear [Sed14n]. Weaver [Sed11t]. webs [RKK^{+}14]. Wee1 [DKMK^{+}11, LWK^{+}13]. Wee1B [OHC10]. Weibel [NWD^{+}11, Sed12f]. weight [Les13t]. Welch [Sho11-32]. well [Sed12p, Sed14k, Sho10-64]. Wendell [Sed12y]. Wg [ETC^{+}12]. WHAMM [SHS^{+}12]. WHAMM-mediated [SHS^{+}12]. Where [BH13]. which [Les10d, SDS^{+}12a]. Who [SF12]. Whole [wFLW^{+}13, DV10, Sed11g]. Whole-genome [wFLW^{+}13]. whom [SF12]. whose [PMB^{+}11]. wide [TDV^{+}14, iYGL^{+}10]. Willebrand [NWD^{+}11]. Winckler [Sed10b]. wing [Les12n]. Wingless [DYP14]. wings [Sho13n]. Wip1 [KSB^{+}13]. within [JKA^{+}10, dSLPRG11, lDSB^{+}10a, lDSB^{+}10b]. without [Les10q, Les13u, OBSG12, Sho11-40]. Wnt [Les11-48, BPMK^{+}14, NSB^{+}11, AAE^{+}14, AKB^{+}13, BAY^{+}11, CSP^{+}10, GR11, KUN^{+}13, KK11, KBW^{+}12, Les13-38, dJPAA^{+}11]. Wnt-11 [KBW^{+}12]. WNT-3A [NSB^{+}11]. Wnt-3a-dependent [GR11]. Wnt/ [BAY^{+}11]. Wnt5b [LBWS10]. Wnt7a [BvMD^{+}14]. Wolf [LeB10]. work [Les12s, Les13y, Les14t, Les14u, Les14-34]. Working [Sho13-59]. workload [Sho12-52]. works [Sho10-71]. world [LRA^{+}10]. Wound [SSW^{+}13, ABVP11, GEN14, PXZ^{+}13, RC12, TSB^{+}14, YFLH12]. wound-induced [RC12]. wounded [Sho14-42, TID^{+}10]. wounding [APC^{+}13]. wounds [Les11-41, MKS^{+}13]. wrapping [NBS^{+}11]. wrestles [Sho12e]. wrong [DSMB13, Sho11-61].

X [AEC^{+}14, Sho14-63, BAS^{+}14, CHL^{+}14, GZZ^{+}14, Sho14-50, bCAH^{+}11, GWR^{+}10]. XBP1 [BAB12]. XendoU [SB14]. Xenopus [BSR^{+}11b, BSR^{+}11c, SLM^{+}13, EL14, HH14b, LHN10, WRF^{+}13]. XII [AMS^{+}13, ISZ^{+}11]. XMAP215 [LMT^{+}12]. XPC [BTL^{+}12]. XTP3-B [BGC^{+}10].

Yamada [Sho10-31]. Yamashita [Sed13z]. Yap [Sed13a]. Yeast [Sho13-60, AHL^{+}11, ABP^{+}14, AMS^{+}13, BGS13a, BGS13b, BKBS12, CCGN11, CWPW11, CSM^{+}12, DBH^{+}11, DBUT13, FHKW11, FSOL14, GBL^{+}11, GCR^{+}12, GKR^{+}11a, GKR^{+}11b, GRH^{+}12, HTM^{+}14, HHY^{+}12, KKUG11, KLP14a, KNPK^{+}10, KST^{+}11, KKK^{+}11, LCLW11, LCS^{+}13, Les12-37, Les13p, MLW13, MJFS10, MHS10, MHKM11, OKNP13, PDMBW11, RY11, RDLT11, SKN^{+}13, Sed14x, Sho11h, Sho12-41, Sho14e, SSH^{+}13, TL12, VŠH^{+}11, VvDV^{+}10, WLZ^{+}14, WMC14, WAW^{+}11, WZHV11, WVT^{+}13, ZF11]. YME1L [AWB^{+}14]. Yohanns [Sed13y]. Yoshinori [Sed12z]. young [Sed12e]. YPX [DCL^{+}12]. Yukiko [Sed13z, Sho10-70]. Yurt [GHC^{+}14]. Yves [Sed14x].

Z [MMVK^{+}12, CZGG12, Sho12q, TO12]. Z-bands [CZGG12]. Z-line [MMVK^{+}12]. Z-ring [TO12]. Zds1 [RY11]. Zds1/Zds2 [RY11]. Zds2 [RY11]. zebrafish [ERS10, YFLH12]. Zelboraf [DS12]. Zfp521 [HSK^{+}10, Sho10-71]. zinc [Sho10-71, DPV^{+}12]. ZO [DSW^{+}11]. ZO-1 [DSW^{+}11]. zona [ABD14, BXB^{+}12, Les14n]. zonae [BBY^{+}12]. zone [CLD11, CTD^{+}10, WLK^{+}11]. zones [KHG^{+}13]. ZP2 [ABD14, BBY^{+}12, BXB^{+}12].

References

Alvarez:2013:FFS

[AA13] Diego E. Alvarez and Hervé Agaisse. The formin FHOD1 and the small GTPase Rac1 promote vaccinia virus actin-based motility. Journal of Cell Biology, 202(7):1075–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

              (electronic). URL http://jcb.rupress.org/content/202/
              7/1075.

Ahmed:2014:MCS

[AAE^{+}14] Mohammed I. Ahmed, Majid Alam, Vladimir U. Emelianov, Krzysztof Poterlowicz, Ankit Patel, Andrey A. Sharov, Andrei N. Mardaryev, and Natalia V. Botchkareva. MicroRNA-214 controls skin and hair follicle development by modulating the activity of the Wnt pathway. Journal of Cell Biology, 207(4):549–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/4/549.

Avella:2014:SDZ

[ABD14] Matteo A. Avella, Boris Baibakov, and Jurrien Dean. A single domain of the ZP2 zona pellucida protein mediates gamete recognition in mice and humans. Journal of Cell Biology, 205 (6):801–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/6/801.

Arsic:2012:NFC

[ABP^{+}12] Nikola Arsic, Nawal Bendris, Marion Peter, Christina Begon-Pescia, Cosette Rebouissou, Gilles Gadéa, Nathalie Bouquier, Frédéric Bibeau, Bénédicte Lemmers, and Jean Marie Blanchard. A novel function for Cyclin A2: Control of cell invasion via RhoA signaling. Journal of Cell Biology, 196(1): 147–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/1/147.

Akamatsu:2014:CNF

[ABP^{+}14] Matthew Akamatsu, Julien Berro, Kai-Ming Pu, Irene R. Tebbs, and Thomas D. Pollard. Cytokinetic nodes in fission yeast arise from two distinct types of nodes that merge during interphase. Journal of Cell Biology, 204(6):977–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
6/977.

Abreu-Blanco:2011:CWR

[ABVP11] Maria Teresa Abreu-Blanco, Jeffrey M. Verboon, and Susan M. Parkhurst. Cell wound repair in Drosophila occurs through three distinct phases of membrane and cytoskeletal remodeling. Journal of Cell Biology, 193(3):455–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
3/455.

Atwood:2012:HPI

[ACO12] Scott X. Atwood, Anne Lynn S. Chang, and Anthony E. Oro. Hedgehog pathway inhibition and the race against tumor evolution. Journal of Cell Biology, 199(2):193–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
2/193.

Avezov:2013:LIF

[ACS^{+}13] Edward Avezov, Benedict C. S. Cross, Gabriele S. Kaminski Schierle, Mikael Winters, Heather P. Harding, Eduardo Pinho Melo, Clemens F. Kaminski, and David Ron. Lifetime imaging of a fluorescent protein sensor reveals surprising stability of ER thiol redox. Journal of Cell Biology, 201(2):337–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
2/337.

Andres-Delgado:2012:IPF

[ADAB^{+}12] Laura Andrés-Delgado, Olga M. Antón, Francesca Bartolini, Ana Ruiz-Sáenz, Isabel Correas, Gregg G. Gundersen, and Miguel A. Alonso. INF2 promotes the formation of detyrosinated microtubules necessary for centrosome reorientation in T cells. Journal of Cell Biology, 198(6):1025–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
6/1025.

Adhikari:2014:MRT

[ADB^{+}14] Deepak Adhikari, M. Kasim Diril, Kiran Busayavalasa, Sanjiv Risal, Shoma Nakagawa, Rebecca Lindkvist, Yan Shen, Vincenzo Coppola, Lino Tessarollo, Nobuaki R. Kudo, Philipp Kaldis, and Kui Liu. Mastl is required for timely activation of APC/C in meiosis I and Cdk1 reactivation in meiosis II. Journal of Cell Biology, 206(7):843–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/7/843.

Alvarez:2012:RCC

[ADF^{+}12] Luis Alvarez, Luru Dai, Benjamin M. Friedrich, Nachiket D. Kashikar, Ingo Gregor, René Pascal, and U. Benjamin Kaupp. The rate of change in Ca^{2+} concentration controls sperm chemotaxis. Journal of Cell Biology, 196(5):653–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
5/653.

Akopian:2013:SAG

[ADS^{+}13] David Akopian, Kush Dalal, Kuang Shen, Franck Duong, and Shu ou Shan. SecYEG activates GTPases to drive the completion of cotranslational protein targeting. Journal of Cell Biology, 200(4):397–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/4/397.

Alcaraz:2014:TXP

[AEC^{+}14] Lindsay B. Alcaraz, Jean-Yves Exposito, Nicolas Chuvin, Roxane M. Pommier, Caroline Cluzel, Sylvie Martel, Stéphanie Sentis, Laurent Bartholin, Claire Lethias, and Ulrich Valcourt. Tenascin-x promotes epithelial-to-mesenchymal transition by activating latent TGF-β. Journal of Cell Biology, 205(3):409–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/3/409.

Ando:2013:RPE

[AFM^{+}13] Koji Ando, Shigetomo Fukuhara, Takahiro Moriya, Yutaro Obara, Norimichi Nakahata, and Naoki Mochizuki. Rap1 potentiates endothelial cell junctions by spatially controlling myosin II activity and actin organization. Journal of Cell Biology, 202(6):901–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/6/901.

Ariotti:2014:CRN

[AFRZ^{+}14] Nicholas Ariotti, Manuel A. Fernández-Rojo, Yong Zhou, Michelle M. Hill, Travis L. Rodkey, Kerry L. Inder, Lukas B. Tanner, Markus R. Wenk, John F. Hancock, and Robert G. Parton. Caveolae regulate the nanoscale organization of the plasma membrane to remotely control Ras signaling. Journal of Cell Biology, 204(5):777–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/5/777.

Amini:2014:CEA

[AGL^{+}14] Rana Amini, Eugénie Goupil, Sara Labella, Monique Zetka, Amy S. Maddox, Jean-Claude Labbé, and Nicolas T. Chartier. C. elegans Anillin proteins regulate intercellular bridge stability and germline syncytial organization. Journal of Cell Biology, 206(1):129–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/1/129.

Araki:2010:TRM

[AGM^{+}10] Yasuhiro Araki, Linda Gombos, Suellen P. S. Migueleti, Lavanya Sivashanmugam, Claude Antony, and Elmar Schiebel. N-terminal regions of Mps1 kinase determine functional bifurcation. Journal of Cell Biology, 189(1):41–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/189/1/41.

Adeyo:2011:YLO

[AHL^{+}11] Oludotun Adeyo, Patrick J. Horn, SungKyung Lee, Derk D. Binns, Anita Chandrahas, Kent D. Chapman, and Joel M. Goodman. The yeast lipin orthologue Pah1p is important for biogenesis of lipid droplets. Journal of Cell Biology, 192(6): 1043–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/6/1043.

Araki:2013:EPC

[AiIK^{+}13] Kazutaka Araki, Shun ichiro Iemura, Yukiko Kamiya, David Ron, Koichi Kato, Tohru Natsume, and Kazuhiro Nagata. Ero1-α and PDIs constitute a hierarchical electron transfer network of endoplasmic reticulum oxidoreductases. Journal of Cell Biology, 202(6):861–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/6/861.

Amcheslavsky:2011:TSC

[AIJI11] Alla Amcheslavsky, Naoto Ito, Jin Jiang, and Y. Tony Ip. Tuberous sclerosis complex and Myc coordinate the growth and division of Drosophila intestinal stem cells. Journal of Cell Biology, 193(4):695–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/4/695.

Araki:2013:ARA

[AKA^{+}13] Yasuhiro Araki, Wei-Chi Ku, Manami Akioka, Alexander I. May, Yu Hayashi, Fumio Arisaka, Yasushi Ishihama, and Yoshinori Ohsumi. Atg38 is required for autophagy-specific phosphatidylinositol 3-kinase complex integrity. Journal of Cell Biology, 203(2):299–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/2/299.

Albers:2013:CWS

[AKB^{+}13] Joachim Albers, Johannes Keller, Anke Baranowsky, Frank Timo Beil, Philip Catala-Lehnen, Jochen Schulze, Michael Amling, and Thorsten Schinke. Canonical Wnt signaling inhibits osteoclastogenesis independent of osteoprotegerin. Journal of Cell Biology, 200(4):537–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/4/537.

Anwar:2012:DLG

[AKC^{+}12] Kamran Anwar, Robin W. Klemm, Amanda Condon, Katharina N. Severin, Miao Zhang, Rodolfo Ghirlando, Junjie Hu, Tom A. Rapoport, and William A. Prinz. The dynamin-like GTPase Sey1p mediates homotypic ER fusion in S. cerevisiae. Journal of Cell Biology, 197(2):209–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/2/209.

Allison:2013:ESI

[ALF^{+}13] Rachel Allison, Jennifer H. Lumb, Coralie Fassier, James W. Connell, Daniel Ten Martin, Matthew N. J. Seaman, Jamilé Hazan, and Evan Reid. An ESCRT–spastin interaction promotes fission of recycling tubules from the endosome. Journal of Cell Biology, 202(3):527–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/3/527.

Ahmed:2013:GCP

[ALS^{+}13] Zamal Ahmed, Chi-Chuan Lin, Kin M. Suen, Fernando A. Melo, James A. Levitt, Klaus Suhling, and John E. Ladbury.

Grb2 controls phosphorylation of FGFR2 by inhibiting receptor kinase and Shp2 phosphatase activity. Journal of Cell Biology, 200(4):493–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/4/493.

Albert:2011:RPS

[ALSN^{+}11] Benjamin Albert, Isabelle Léger-Silvestre, Christophe Normand, Martin K. Ostermaier, Jorge Pérez-Fernández, Kostya I. Panov, Joost C. B. M. Zomerdijk, Patrick Schultz, and Olivier Gadal. RNA polymerase I–specific subunits promote polymerase clustering to enhance the rRNA gene transcription cycle. Journal of Cell Biology, 192(2):277–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/192/2/
277.

Amelio:2012:MTE

[ALV^{+}12] Ivano Amelio, Anna Maria Lena, Giuditta Viticchiè, Ruby Shalom-Feuerstein, Alessandro Terrinoni, David Dinsdale, Giandomenico Russo, Claudia Fortunato, Elena Bonanno, Luigi Giusto Spagnoli, Daniel Aberdam, Richard Austen Knight, Eleonora Candi, and Gerry Melino. miR-24 triggers epidermal differentiation by controlling actin adhesion and cell migration. Journal of Cell Biology, 199(2):347–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
2/347.

Aldridge:2014:SGD

[AMGC14] Cassie Aldridge, Xianyue Ma, Fabien Gerard, and Kenneth Cline. Substrate-gated docking of pore subunit Tha4 in the TatC cavity initiates Tat translocase assembly. Journal of Cell Biology, 205(1):51–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/1/51.

Abeyweera:2011:ISB

[AMH11] Thushara P. Abeyweera, Ernesto Merino, and Morgan Huse. Inhibitory signaling blocks activating receptor clustering and induces cytoskeletal retraction in natural killer cells. Journal of Cell Biology, 192(4):675–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/4/675.

Ashida:2011:CDI

[AMO^{+}11] Hiroshi Ashida, Hitomi Mimuro, Michinaga Ogawa, Taira Kobayashi, Takahito Sanada, Minsoo Kim, and Chihiro Sasakawa. Cell death and infection: a double-edged sword for host and pathogen survival. Journal of Cell Biology, 195(6): 931–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/6/931.

Alix:2011:SMT

[AMR11] Eric Alix, Shaeri Mukherjee, and Craig R. Roy. Subversion of membrane transport pathways by vacuolar pathogens. Journal of Cell Biology, 195(6):943–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/6/943.

Albert:2013:SCC

[AMS^{+}13] Benjamin Albert, Julien Mathon, Ashutosh Shukla, Hicham Saad, Christophe Normand, Isabelle Léger-Silvestre, David Villa, Alain Kamgoue, Julien Mozziconacci, Hua Wong, Christophe Zimmer, Purnima Bhargava, Aurélien Bancaud, and Olivier Gadal. Systematic characterization of the conformation and dynamics of budding yeast chromosome XII. Journal of Cell Biology, 202(2):201–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/2/201.

Anastasia:2012:LBM

[ANT^{+}12] Steph D. Anastasia, Duy Linh Nguyen, Vu Thai, Melissa Meloy, Tracy MacDonough, and Douglas R. Kellogg. A link between mitotic entry and membrane growth suggests a novel model for cell size control. Journal of Cell Biology, 197 (1):89–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/1/89.

Anantharam:2010:LTC

[AOE^{+}10] Arun Anantharam, Bibiana Onoa, Robert H. Edwards, Ronald W. Holz, and Daniel Axelrod. Localized topological changes of the plasma membrane upon exocytosis visualized by polarized TIRFM. Journal of Cell Biology, 188(3): 415–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/3/415.

Anantharam:2012:LTC

[AOE^{+}12] Arun Anantharam, Bibiana Onoa, Robert H. Edwards, Ronald W. Holz, and Daniel Axelrod. Localized topological changes of the plasma membrane upon exocytosis visualized by polarized TIRFM. Journal of Cell Biology, 197(3):457–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
197/3/457.

Antunes:2013:CWA

[APC^{+}13] Marco Antunes, Telmo Pereira, João V. Cordeiro, Luis Almeida, and Antonio Jacinto. Coordinated waves of actomyosin flow and apical cell constriction immediately after wounding. Journal of Cell Biology, 202(2):365–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
2/365.

Ardite:2012:PRM

[APV^{+}12] Esther Ardite, Eusebio Perdiguero, Berta Vidal, Susana Gutarra, Antonio L. Serrano, and Pura Muñoz-Cánoves. PAI-1-regulated miR-21 defines a novel age-associated fibrogenic pathway in muscular dystrophy. Journal of Cell Biology, 196 (1):163–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/1/163.

Albanese:2010:RAC

[ARF10] Véronique Albanèse, Stefanie Reissmann, and Judith Frydman. A ribosome-anchored chaperone network that facilitates eukaryotic ribosome biogenesis. Journal of Cell Biology, 189 (1):69–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/1/69.

Arikkath:2010:CSS

[Ari10] Jyothi Arikkath. N-cadherin: stabilizing synapses. Journal of Cell Biology, 189(3):397–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/3/397.

Albers:2011:CBF

[ASB^{+}11] Joachim Albers, Jochen Schulze, F. Timo Beil, Matthias Gebauer, Anke Baranowsky, Johannes Keller, Robert P. Marshall, Kristofer Wintges, Felix W. Friedrich, Matthias Priemel, Arndt F. Schilling, Johannes M. Rueger, Kerstin Cornils, Boris Fehse, Thomas Streichert, Guido Sauter, Franz Jakob, Karl L. Insogna, Barbara Pober, Klaus-Peter Knobeloch, Uta Francke, Michael Amling, and Thorsten Schinke. Control of bone formation by the serpentine receptor Frizzled-9. Journal of Cell Biology, 192(6):1057–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/6/1057.

Asensio:2010:RSI

[ASE10] Cédric S. Asensio, Daniel W. Sirkis, and Robert H. Edwards. RNAi screen identifies a role for adaptor protein AP-3 in sorting to the regulated secretory pathway. Journal of Cell Biology, 191(6):1173–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/6/1173.

Ashrafi:2014:MDM

[ASLS14] Ghazaleh Ashrafi, Julia S. Schlehe, Matthew J. LaVoie, and Thomas L. Schwarz. Mitophagy of damaged mitochondria occurs locally in distal neuronal axons and requires PINK1 and Parkin. Journal of Cell Biology, 206(5):655–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
5/655.

Akbar:2011:FBG

[ATKK11] Mohammed Ali Akbar, Charles Tracy, Walter H. A. Kahr, and Helmut Krämer. The full-of-bacteria gene is required for phagosome maturation during immune defense in Drosophila. Journal of Cell Biology, 192(3):383–??, February 2011. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/192/3/
383.

Ahmed:2012:KNR

[ATU^{+}12] Syed M. Ahmed, Brigitte L. Thériault, Maruti Uppalapati, Catherine W. N. Chiu, Brenda L. Gallie, Sachdev S. Sidhu, and Stéphane Angers. KIF14 negatively regulates Rap1a–Radil signaling during breast cancer progression. Journal of Cell Biology, 199(6):951–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/6/951.

Askari:2010:FAS

[ATW^{+}10] Janet A. Askari, Christopher J. Tynan, Stephen E. D. Webb, Marisa L. Martin-Fernandez, Christoph Ballestrem, and Martin J. Humphries. Focal adhesions are sites of integrin extension. Journal of Cell Biology, 188(6):891–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
6/891.

Agarwal:2011:ADI

[AvCG^{+}11] Sheba Agarwal, Wiggert A. van Cappellen, Aude Guénolé, Berina Eppink, Sam E. V. Linsen, Erik Meijering, Adriaan Houtsmuller, Roland Kanaar, and Jeroen Essers. ATPdependent and independent functions of Rad54 in genome maintenance. Journal of Cell Biology, 192(5):735–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
5/735.

Adell:2014:CBV

[AVP^{+}14] Manuel Alonso Y. Adell, Georg F. Vogel, Mehrshad Pakdel, Martin Müller, Herbert Lindner, Michael W. Hess, and David Teis. Coordinated binding of Vps4 to ESCRT–III drives membrane neck constriction during MVB vesicle formation. Journal of Cell Biology, 205(1):33–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/205/1/33.

Anand:2014:APY

[AWB^{+}14] Ruchika Anand, Timothy Wai, Michael J. Baker, Nikolay Kladt, Astrid C. Schauss, Elena Rugarli, and Thomas Langer. The i -AAA protease YME1L and OMA1 cleave OPA1 to balance mitochondrial fusion and fission. Journal of Cell Biology, 204(6):919–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/6/919.

Apaja:2010:QCU

[AXL10] Pirjo M. Apaja, Haijin Xu, and Gergely L. Lukacs. Quality control for unfolded proteins at the plasma membrane. Journal of Cell Biology, 191(3):553–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/3/553.

Atkins:2013:ICD

[AYS^{+}13] Benjamin D. Atkins, Satoshi Yoshida, Koji Saito, Chi-Fang Wu, Daniel J. Lew, and David Pellman. Inhibition of Cdc42 during mitotic exit is required for cytokinesis. Journal of Cell Biology, 202(2):231–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/2/231.

Behrman:2011:CRM

[BAAW11] Shannon Behrman, Diego Acosta-Alvear, and Peter Walter. A CHOP-regulated microRNA controls rhodopsin expression. Journal of Cell Biology, 192(6):919–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/6/919.

Byrd:2012:MLE

[BAB12] Andrew E. Byrd, Ileana V. Aragon, and Joseph W. Brewer. MicroRNA-30c-2* limits expression of proadaptive factor XBP1 in the unfolded protein response. Journal of Cell Biology, 196(6):689–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/6/689.

Babst:2014:QCP

[Bab14] Markus Babst. Quality control at the plasma membrane: One mechanism does not fit all. Journal of Cell Biology, 205 (1):11–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/1/11.

Borner:2012:MPP

[BAH^{+}12] Georg H. H. Borner, Robin Antrobus, Jennifer Hirst, Gary S. Bhumbra, Patrycja Kozik, Lauren P. Jackson, Daniela A. Sahlender, and Margaret S. Robinson. Multivariate proteomic profiling identifies novel accessory proteins of coated vesicles.

Journal of Cell Biology, 197(1):141–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/1/141.

Barr:2013:RGM

[Bar13] Francis A. Barr. Rab GTPases and membrane identity: Causal or inconsequential? Journal of Cell Biology, 202(2):191–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
202/2/191.

Broering:2014:BED

[BAS^{+}14] Tyler J. Broering, Kris G. Alavattam, Ruslan I. Sadreyev, Yosuke Ichijima, Yasuko Kato, Kazuteru Hasegawa, R. Daniel Camerini-Otero, Jeannie T. Lee, Paul R. Andreassen, and Satoshi H. Namekawa. BRCA1 establishes DNA damage signaling and pericentric heterochromatin of the X chromosome in male meiosis. Journal of Cell Biology, 205(5):663–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/205/
5/663.

Berndt:2011:MEU

[BAY^{+}11] Jason D. Berndt, Atsushi Aoyagi, Peitzu Yang, Jamie N. Anastas, Lan Tang, and Randall T. Moon. Mindbomb 1, an E3 ubiquitin ligase, forms a complex with RYK to activate Wnt/β-catenin signaling. Journal of Cell Biology, 194(5): 737–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/5/737.

Bastos:2010:PNR

[BB10] Ricardo Nunes Bastos and Francis A. Barr. Plk1 negatively regulates Cep55 recruitment to the midbody to ensure orderly abscission. Journal of Cell Biology, 191(4):751–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
4/751.

Bolhy:2011:NDN

[BBD^{+}11] Stéphanie Bolhy, Imène Bouhlel, Elisa Dultz, Tania Nayak, Michela Zuccolo, Xavier Gatti, Richard Vallee, Jan Ellenberg, and Valérie Doye. A Nup133-dependent NPC-anchored network tethers centrosomes to the nuclear envelope in prophase. Journal of Cell Biology, 192(5):855–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/5/855.

Bozec:2010:FAC

[BBJ^{+}10] Aline Bozec, Latifa Bakiri, Maria Jimenez, Thorsten Schinke, Michael Amling, and Erwin F. Wagner. Fra-2/ AP-1 controls bone formation by regulating osteoblast differentiation and collagen production. Journal of Cell Biology, 190(6): 1093–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/6/1093.

Bai:2013:NSR

[BBK^{+}13] Xiaobo Bai, Jonathan R. Bowen, Tara K. Knox, Kaifeng Zhou, Manuela Pendziwiat, Gregor Kuhlenbäumer, Charles V. Sindelar, and Elias T. Spiliotis. Novel septin 9 repeat motifs altered in neuralgic amyotrophy bind and bundle microtubules. Journal of Cell Biology, 203(6):895–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/6/895.

Bergoglio:2013:DSP

[BBW^{+}13] Valérie Bergoglio, Anne-Sophie Boyer, Erin Walsh, Valeria Naim, Gaëlle Legube, Marietta Y. W. T. Lee, Laurie Rey, Filippo Rosselli, Christophe Cazaux, Kristin A. Eckert, and Jean-Sébastien Hoffmann. DNA synthesis by Pol η promotes fragile site stability by preventing under-replicated DNA in mitosis. Journal of Cell Biology, 201(3):395–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/3/395.

Boname:2014:CSP

[BBW^{+}14] Jessica M. Boname, Stuart Bloor, Michal P. Wandel, James A. Nathan, Robin Antrobus, Kevin S. Dingwell, Teresa L. Thurston, Duncan L. Smith, James C. Smith, Felix Randow, and Paul J. Lehner. Cleavage by signal peptide peptidase is required for the degradation of selected tail-anchored proteins. Journal of Cell Biology, 205(6):847–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/205/6/847.

Baibakov:2012:HSB

[BBY^{+}12] Boris Baibakov, Nathan A. Boggs, Belinda Yauger, Galina Baibakov, and Jurrien Dean. Human sperm bind to the Nterminal domain of ZP2 in humanized zonae pellucidae in transgenic mice. Journal of Cell Biology, 197(7):897–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
7/897.

Bonazzi:2011:IBEa

[BC11a] Matteo Bonazzi and Pascale Cossart. Impenetrable barriers or entry portals? The role of cell–cell adhesion during infection. Journal of Cell Biology, 195(3):349–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/3/349.

Bonazzi:2011:IBEb

[BC11b] Matteo Bonazzi and Pascale Cossart. Impenetrable barriers or entry portals? The role of cell–cell adhesion during infection. Journal of Cell Biology, 195(7):1205–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
7/1205.

Chen:2011:BXR

[bCAH^{+}11] Ying bei Chen, Miguel A. Aon, Yi-Te Hsu, Lucian Soane, Xinchen Teng, J. Michael McCaffery, Wen-Chih Cheng, Bing Qi, Hongmei Li, Kambiz N. Alavian, Margaret Dayhoff-Brannigan, Shifa Zou, Fernando J. Pineda, Brian O’Rourke, Young H. Ko, Peter L. Pedersen, Leonard K. Kaczmarek, Elizabeth A. Jonas, and J. Marie Hardwick. Bcl-x_{L} regulates mitochondrial energetics by stabilizing the inner membrane potential. Journal of Cell Biology, 195(2):263–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
2/263.

Bastos:2014:KPB

[BCB14a] Ricardo Nunes Bastos, Michael J. Cundell, and Francis A. Barr. KIF4A and PP2A–B56 form a spatially restricted feedback loop opposing Aurora B at the anaphase central spindle. Journal of Cell Biology, 207(6):683–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
6/683.

Bizarro:2014:PSA

[BCB^{+}14b] Jonathan Bizarro, Christophe Charron, Séverine Boulon, Belinda Westman, Bérengère Pradet-Balade, Franck Vandermoere, Marie-Eve Chagot, Marie Hallais, Yasmeen Ahmad, Heinrich Leonhardt, Angus Lamond, Xavier Manival, Christiane Branlant, Bruno Charpentier, Céline Verheggen, and Edouard Bertrand. Proteomic and 3D structure analyses highlight the C/D box snoRNP assembly mechanism and its control. Journal of Cell Biology, 207(4):463–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/207/4/463.

Bohdanowicz:2010:CCI

[BCBG10] Michal Bohdanowicz, Gabriela Cosío, Jonathan M. Backer, and Sergio Grinstein. Class I and class III phosphoinositide 3-kinases are required for actin polymerization that propels phagosomes. Journal of Cell Biology, 191(5):999–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
191/5/999.

Britton:2013:NMH

[BCJ13] Sébastien Britton, Julia Coates, and Stephen P. Jackson. A new method for high-resolution imaging of Ku foci to decipher mechanisms of DNA double-strand break repair. Journal of Cell Biology, 202(3):579–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/202/3/579.

Braun:2014:RAR

[BDB^{+}14] Alexander Braun, Kyvan Dang, Felinah Buslig, Michelle A. Baird, Michael W. Davidson, Clare M. Waterman, and Kenneth A. Myers. Rac1 and Aurora A regulate MCAK to polarize microtubule growth in migrating endothelial cells. Journal of Cell Biology, 206(1):97–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/1/97.

Benkemoun:2014:PLR

[BDC^{+}14] Laura Benkemoun, Catherine Descoteaux, Nicolas T. Chartier, Lionel Pintard, and Jean-Claude Labbé. PAR-4/ LKB1 regulates DNA replication during asynchronous division of the early C. elegans embryo. Journal of Cell Biology, 205(4): 447–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/4/447.

Buster:2013:SUL

[BDN^{+}13] Daniel W. Buster, Scott G. Daniel, Huy Q. Nguyen, Sarah L. Windler, Lara C. Skwarek, Maureen Peterson, Meredith Roberts, Joy H. Meserve, Tom Hartl, Joseph E. Klebba, David Bilder, Giovanni Bosco, and Gregory C. Rogers. SCF^{Slimb} ubiquitin ligase suppresses condensin II-mediated nuclear reorganization by degrading Cap–H2. Journal of Cell Biology, 201(1):49–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/1/49.

Bechler:2010:PCP

[BDR^{+}10] Marie E. Bechler, Anne M. Doody, Esther Racoosin, Lin Lin, Kelvin H. Lee, and William J. Brown. The phospholipase complex PAFAH Ib regulates the functional organization of the Golgi complex. Journal of Cell Biology, 190(1):45–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
1/45.

Bellis:2012:TSA

[BDR^{+}12] Julien Bellis, Isabelle Duluc, Béatrice Romagnolo, Christine Perret, Maree C. Faux, Denis Dujardin, Caroline Formstone, Sally Lightowler, Robert G. Ramsay, Jean-Noël Freund, and Jan R. De Mey. The tumor suppressor Apc controls planar cell polarities central to gut homeostasis. Journal of Cell Biology, 198(3):331–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/3/331.

Bufalino:2013:ASD

[BDvdK13] Mary Rose Bufalino, Brian DeVeale, and Derek van der Kooy. The asymmetric segregation of damaged proteins is stem cell– type dependent. Journal of Cell Biology, 201(4):523–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
4/523.

Biswas:2010:PCI

[BEJ10] Sayantanee Biswas, Michelle R. Emond, and James D. Jontes. Protocadherin-19 and N-cadherin interact to control cell movements during anterior neurulation. Journal of Cell Biology, 191(5):1029–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/5/1029.

Bezprozvanny:2012:PNL

[Bez12] Ilya Bezprozvanny. Presenilins: a novel link between intracellular calcium signaling and lysosomal function? Journal of Cell Biology, 198(1):7–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/1/7.

Bix:2013:ECE

[BFG^{+}13] Gregory Bix, Jian Fu, Eva M. Gonzalez, Laura Macro, Amy Barker, Shelly Campbell, Mary M. Zutter, Samuel A. Santoro, Jiyeun K. Kim, Magnus Höök, Charles C. Reed, and Renato V. Iozzo. Endorepellin causes endothelial cell disassembly of actin cytoskeleton and focal adhesions through α2β1 integrin. Journal of Cell Biology, 201(4):641–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
4/641.

Bose:2010:CGE

[BG10] Tania Bose and Jennifer L. Gerton. Cohesinopathies, gene expression, and chromatin organization. Journal of Cell Biology, 189(2):201–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/2/201.

Baum:2011:DAJ

[BG11a] Buzz Baum and Marios Georgiou. Dynamics of adherens junctions in epithelial establishment, maintenance, and remodeling. Journal of Cell Biology, 192(6):907–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
6/907.

Brown:2011:RGW

[BG11b] Michael S. Brown and Joseph L. Goldstein. Richard G.W. Anderson (1940–2011) and the birth of receptor-mediated endocytosis. Journal of Cell Biology, 193(4):601–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
4/601.

Bastos:2013:ABS

[BGB^{+}13] Ricardo Nunes Bastos, Sapan R. Gandhi, Ryan D. Baron, Ulrike Gruneberg, Erich A. Nigg, and Francis A. Barr. Aurora B suppresses microtubule dynamics and limits central spindle size by locally activating KIF4A. Journal of Cell Biology, 202 (4):605–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/4/605.

Bernasconi:2010:SRH

[BGC^{+}10] Riccardo Bernasconi, Carmela Galli, Verena Calanca, Toshihiro Nakajima, and Maurizio Molinari. Stringent requirement for HRD1, SEL1L, and OS-9/XTP3-B for disposal of ERAD– L_{S} substrates. Journal of Cell Biology, 188(2):223–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
2/223.

Binda:2014:GPC

[BGC^{+}14] Chantal Binda, Samuel Génier, Andréane Cartier, Jean-François Larrivée, Jana Stankova, Jason C. Young, and Jean-Luc Parent. A G protein–coupled receptor and the intracellular synthase of its agonist functionally cooperate. Journal of Cell Biology, 204(3):377–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/3/377.

Breker:2013:NSCa

[BGS13a] Michal Breker, Melissa Gymrek, and Maya Schuldiner. A novel single-cell screening platform reveals proteome plasticity during yeast stress responses. Journal of Cell Biology, 200 (6):839–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/200/6/839.

Breker:2013:NSCb

[BGS13b] Michal Breker, Melissa Gymrek, and Maya Schuldiner. A novel single-cell screening platform reveals proteome plasticity during yeast stress responses. Journal of Cell Biology, 201 (2):353–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/2/353.

Bulgakova:2013:DMP

[BGY^{+}13] Natalia A. Bulgakova, Ilya Grigoriev, Alpha S. Yap, Anna Akhmanova, and Nicholas H. Brown. Dynamic microtubules produce an asymmetric E–cadherin–Bazooka complex to maintain segment boundaries. Journal of Cell Biology, 201(6): 887–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/6/887.

Betz:2013:WMW

[BH13] Charles Betz and Michael N. Hall. Where is mTOR and what is it doing there? Journal of Cell Biology, 203(4):563–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
203/4/563.

Bussiere:2012:IPS

[BHA^{+}12] Cyril Bussiere, Yaser Hashem, Sucheta Arora, Joachim Frank, and Arlen W. Johnson. Integrity of the P-site is probed during maturation of the 60S ribosomal subunit. Journal of Cell Biology, 197(6):747–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/6/747.

Bowen:2011:SGS

[BHB^{+}11] Jonathan R. Bowen, Daniel Hwang, Xiaobo Bai, Dheeraj Roy, and Elias T. Spiliotis. Septin GTPases spatially guide microtubule organization and plus end dynamics in polarizing epithelia. Journal of Cell Biology, 194(2):187–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
2/187.

Becker-Herman:2011:WDB

[BHMBS^{+}11] Shirly Becker-Herman, Almut Meyer-Bahlburg, Marc A. Schwartz, Shaun W. Jackson, Kelly L. Hudkins, Chaohong Liu, Blythe D. Sather, Socheath Khim, Denny Liggitt, Wenxia Song, Gregg J. Silverman, Charles E. Alpers, and David J. Rawlings. WASp-deficient B cells play a critical, cell-intrinsic role in triggering autoimmunity. Journal of Cell Biology, 194 (5):??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/5/i10.

Barenz:2013:CSP

[BIY^{+}13] Felix Bärenz, Daigo Inoue, Hideki Yokoyama, Justus Tegha-Dunghu, Stephanie Freiss, Stefanie Draeger, Dmytro Mayilo, Ivana Cado, Sabine Merker, Maren Klinger, Burkhard Hoeckendorf, Sahra Pilz, Kerstin Hupfeld, Herbert Steinbeisser, Holger Lorenz, Thomas Ruppert, Joachim Wittbrodt, and Oliver J. Gruss. The centriolar satellite protein SSX2IP promotes centrosome maturation. Journal of Cell Biology, 202 (1):81–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/1/81.

Bryant:2012:SMS

[BJ12] Nia J. Bryant and David E. James. The Sec1p/Munc18 (SM) protein, Vps45p, cycles on and off membranes during vesicle transport. Journal of Cell Biology, 199(5):863–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
5/863.

Boswell:2012:MRC

[BJE^{+}12] Kristin L. Boswell, Declan J. James, Joseph M. Esquibel, Stephen Bruinsma, Ryutaro Shirakawa, Hisanori Horiuchi, and Thomas F. J. Martin. Munc13-4 reconstitutes calciumdependent SNARE-mediated membrane fusion. Journal of Cell Biology, 197(2):301–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/197/2/301.

Bian:2013:GCE

[BKAB13] Qian Bian, Nimish Khanna, Jurgis Alvikas, and Andrew S. Belmont. β-globin cis-elements determine differential nuclear targeting through epigenetic modifications. Journal of Cell Biology, 203(5):767–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/5/767.

Brownlee:2011:PPR

[BKBR11] Christopher W. Brownlee, Joey E. Klebba, Daniel W. Buster, and Gregory C. Rogers. The Protein Phosphatase 2A regulatory subunit Twins stabilizes Plk4 to induce centriole amplification. Journal of Cell Biology, 195(2):231–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
2/231.

Bui:2012:NMY

[BKBS12] Huyen T. Bui, Mary A. Karren, Debjani Bhar, and Janet M. Shaw. A novel motif in the yeast mitochondrial dynamin Dnm1 is essential for adaptor binding and membrane recruitment. Journal of Cell Biology, 199(4):613–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
4/613.

Buttitta:2010:RCC

[BKE10] Laura A. Buttitta, Alexia J. Katzaroff, and Bruce A. Edgar. A robust cell cycle control mechanism limits E2F-induced proliferation of terminally differentiated cells in vivo. Journal of Cell Biology, 189(6):981–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/6/981.

Barr:2010:CFC

[BKG10] Alexis R. Barr, John V. Kilmartin, and Fanni Gergely. CDK5RAP2 functions in centrosome to spindle pole attachment and DNA damage response. Journal of Cell Biology, 189(1):23–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/1/23.

Breuer:2010:HPM

[BKK^{+}10] Manuel Breuer, Agnieszka Kolano, Mijung Kwon, Chao-Chin Li, Ting-Fen Tsai, David Pellman, Stéphane Brunet, and Marie-Hélène Verlhac. HURP permits MTOC sorting for robust meiotic spindle bipolarity, similar to extra centrosome clustering in cancer cells. Journal of Cell Biology, 191(7): 1251–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/7/1251.

Bertazzi:2011:CPL

[BKP11] Daniela Trinca Bertazzi, Bahtiyar Kurtulmus, and Gislene Pereira. The cortical protein Lte1 promotes mitotic exit by inhibiting the spindle position checkpoint kinase Kin4. Journal of Cell Biology, 193(6):1033–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/6/1033.

Barnhart:2011:HCC

[BKS^{+}11] Meghan C. Barnhart, P. Henning J. L. Kuich, Madison E. Stellfox, Jared A. Ward, Emily A. Bassett, Ben E. Black, and Daniel R. Foltz. HJURP is a CENP–A chromatin assembly factor sufficient to form a functional de novo kinetochore. Journal of Cell Biology, 194(2):229–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/2/229.

Breslow:2013:VAE

[BKS^{+}13] David K. Breslow, Elena F. Koslover, Federica Seydel, Andrew J. Spakowitz, and Maxence V. Nachury. An in vitro assay for entry into cilia reveals unique properties of the soluble diffusion barrier. Journal of Cell Biology, 203(1):129–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
1/129.

Banerjee:2014:EES

[BKS14] Budhaditya Banerjee, Cortney A. Kestner, and P. Todd Stukenberg. EB1 enables spindle microtubules to regulate centromeric recruitment of Aurora B. Journal of Cell Biology, 204 (6):947–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/6/947.

Briguglio:2013:LSR

[BKT13] Joseph S. Briguglio, Santosh Kumar, and Aaron P. Turkewitz. Lysosomal sorting receptors are essential for secretory granule biogenesis in Tetrahymena. Journal of Cell Biology, 203(3): 537–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/3/537.

Benjamin:2010:CRA

[BKY^{+}10] Jacqueline M. Benjamin, Adam V. Kwiatkowski, Changsong Yang, Farida Korobova, Sabine Pokutta, Tatyana Svitkina, William I. Weis, and W. James Nelson. αe-catenin regulates actin dynamics independently of cadherin-mediated cell– cell adhesion. Journal of Cell Biology, 189(2):339–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
2/339.

Bakkar:2012:IAN

[BLC^{+}12] Nadine Bakkar, Katherine Ladner, Benjamin D. Canan, Sandya Liyanarachchi, Naresh C. Bal, Meghna Pant, Muthu Periasamy, Qiutang Li, Paul M. L. Janssen, and Denis C. Guttridge. IKKα and alternative NF-κB regulate PGC-1β to promote oxidative muscle metabolism. Journal of Cell Biology, 196(4):497–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/4/497.

Burke:2014:CPC

[BLC^{+}14] Michael C. Burke, Feng-Qian Li, Benjamin Cyge, Takeshi Arashiro, Heather M. Brechbuhl, Xingwang Chen, Saul S. Siller, Matthew A. Weiss, Christopher B. O’Connell, Damon Love, Christopher J. Westlake, Susan D. Reynolds, Ryoko Kuriyama, and Ken-Ichi Takemaru. Chibby promotes ciliary vesicle formation and basal body docking during airway cell differentiation. Journal of Cell Biology, 207(1):123–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
207/1/123.

Brachet:2010:AGR

[BLI^{+}10] Anna Brachet, Christophe Leterrier, Marie Irondelle, Marie-Pierre Fache, Victor Racine, Jean-Baptiste Sibarita, Daniel Choquet, and Bénédicte Dargent. Ankyrin G restricts ion channel diffusion at the axonal initial segment before the establishment of the diffusion barrier. Journal of Cell Biology, 191 (2):383–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/2/383.

Baez:2011:SMS

[BLM^{+}11] María Verónica Baez, Luciana Luchelli, Darío Maschi, Martín Habif, Malena Pascual, María Gabriela Thomas, and Graciela Lidia Boccaccio. Smaug1 mRNA-silencing foci respond to NMDA and modulate synapse formation. Journal of Cell Biology, 195(7):1141–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/7/1141.

Babich:2012:FAP

[BLO^{+}12] Alexander Babich, Shuixing Li, Roddy S. O’Connor, Michael C. Milone, Bruce D. Freedman, and Janis K. Burkhardt. F-actin polymerization and retrograde flow drive sustained PLCγ1 signaling during T cell activation. Journal of Cell Biology, 197 (6):775–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/6/775.

Brill:2011:SCD

[BLT^{+}11] Monika S. Brill, Jeff W. Lichtman, Wesley Thompson, Yi Zuo, and Thomas Misgeld. Spatial constraints dictate glial territories at murine neuromuscular junctions. Journal of Cell Biology, 195(2):293–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/2/293.

Budde:2010:CBB

[BM10] Priya Prakash Budde and Tom Misteli. Cell biology beyond the cell. Journal of Cell Biology, 190(1):7–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/1/7.

Bizzari:2011:CCS

[BM11] Farid Bizzari and Adele L. Marston. Cdc55 coordinates spindle assembly and chromosome disjunction during meiosis. Journal of Cell Biology, 193(7):1213–??, June 2011. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/7/1213.

Bastounis:2014:BCA

[BMÁG^{+}14] Effie Bastounis, Ruedi Meili, Begoña Álvarez-González, Joshua Francois, Juan C. del Álamo, Richard A. Firtel, and Juan C. Lasheras. Both contractile axial and lateral traction force dynamics drive amoeboid cell motility. Journal of Cell Biology, 204(6):1045–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/6/1045.

Bruns:2011:BNC

[BMC^{+}11] Caroline Bruns, J. Michael McCaffery, Amy J. Curwin, Juan M. Duran, and Vivek Malhotra. Biogenesis of a novel compartment for autophagosome-mediated unconventional protein secretion. Journal of Cell Biology, 195(6): 979–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/6/979.

Brunner:2011:OMR

[BMFC^{+}11] Molly Brunner, Angélique Millon-Frémillon, Genevieve Chevalier, Inaam A. Nakchbandi, Deane Mosher, Marc R. Block, Corinne Albigès-Rizo, and Daniel Bouvard. Osteoblast mineralization requires β1 integrin/ICAP-1-dependent fibronectin deposition. Journal of Cell Biology, 194(2):307–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
2/307.

Brunner:2013:OMR

[BMFC^{+}13] Molly Brunner, Angélique Millon-Frémillon, Genevieve Chevalier, Inaam A. Nakchbandi, Deane Mosher, Marc R. Block, Corinne Albigès-Rizo, and Daniel Bouvard. Osteoblast mineralization requires β1 integrin/ICAP-1-dependent fibronectin deposition. Journal of Cell Biology, 201(4):643–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
4/643.

Ballard:2014:RNS

[BMG14] Shannon L. Ballard, Daniel L. Miller, and Barry Ganetzky. Retrograde neurotrophin signaling through Tollo regulates synaptic growth in Drosophila. Journal of Cell Biology, 204 (7):1157–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/7/1157.

Boettcher:2012:NEMa

[BMLB^{+}12a] Barbara Boettcher, Tatiana T. Marquez-Lago, Mathias Bayer, Eric L. Weiss, and Yves Barral. Nuclear envelope morphology constrains diffusion and promotes asymmetric protein segregation in closed mitosis. Journal of Cell Biology, 197(7):921–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
197/7/921.

Boettcher:2012:NEMb

[BMLB^{+}12b] Barbara Boettcher, Tatiana T. Marquez-Lago, Mathias Bayer, Eric L. Weiss, and Yves Barral. Nuclear envelope morphology constrains diffusion and promotes asymmetric protein segregation in closed mitosis. Journal of Cell Biology, 198(1):143–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
198/1/143.

Barriga:2013:HFH

[BMRM13] Elias H. Barriga, Patrick H. Maxwell, Ariel E. Reyes, and Roberto Mayor. The hypoxia factor Hif-1α controls neural crest chemotaxis and epithelial to mesenchymal transition. Journal of Cell Biology, 201(5):759–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/5/759.

Bohgaki:2011:CIC

[BMS^{+}11] Toshiyuki Bohgaki, Julien Mozo, Leonardo Salmena, Elzbieta Matysiak-Zablocki, Miyuki Bohgaki, Otto Sanchez, Andreas Strasser, Anne Hakem, and Razqallah Hakem. Caspase-8 inactivation in T cells increases necroptosis and suppresses autoimmunity in Bim^{−/}− mice. Journal of Cell Biology, 195(2): 277–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/2/277.

Barton:2014:PLK

[BNDB^{+}14] Olivia Barton, Steffen C. Naumann, Ronja Diemer-Biehs, Julia Künzel, Monika Steinlage, Sandro Conrad, Nodar Makharashvili, Jiadong Wang, Lin Feng, Bernard S. Lopez, Tanya T. Paull, Junjie Chen, Penny A. Jeggo, and Markus Löbrich. Polo-like kinase 3 regulates CtIP during DNA double-strand break repair in G1. Journal of Cell Biology, 206(7):877– ??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/7/877.

Blamowska:2012:BMH

[BNH12] Marta Blamowska, Walter Neupert, and Kai Hell. Biogenesis of the mitochondrial Hsp70 chaperone. Journal of Cell Biology, 199(1):125–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/1/125.

Beck:2010:RCD

[BNL^{+}10] Halfdan Beck, Viola Nähse, Marie Sofie Yoo Larsen, Petra Groth, Trevor Clancy, Michael Lees, Mette Jørgensen, Thomas Helleday, Randi G. Syljuåsen, and Claus Storgaard Sørensen. Regulators of cyclin-dependent kinases are crucial for maintaining genome integrity in S phase. Journal of Cell Biology, 188(5):629–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/5/629.

Beguin:2014:BSV

[BNM^{+}14] Pascal Béguin, Kazuaki Nagashima, Ramasubbu N. Mahalakshmi, Réjan Vigot, Atsuko Matsunaga, Takafumi Miki, Mei Yong Ng, Yu Jin Alvin Ng, Chiaw Hwee Lim, Hock Soon Tay, Le-Ann Hwang, Dmitri Firsov, Bor Luen Tang, Nobuya Inagaki, Yasuo Mori, Susumu Seino, Thomas Launey, and Walter Hunziker. BARP suppresses voltage-gated calcium channel activity and Ca^{2+}-evoked exocytosis. Journal of Cell Biology, 205(2):233–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/2/233.

Boettiger:2012:UFV

[Boe12] David Boettiger. Using force to visualize conformational activation of integrins. Journal of Cell Biology, 199(3):423–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
199/3/423.

Bonifacino:2014:API

[Bon14] Juan S. Bonifacino. Adaptor proteins involved in polarized sorting. Journal of Cell Biology, 204(1):7–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/204/1/7.

Bastos:2012:CIR

[BPB^{+}12] Ricardo Nunes Bastos, Xenia Penate, Michelle Bates, Dean Hammond, and Francis A. Barr. CYK4 inhibits Rac1dependent PAK1 and ARHGEF7 effector pathways during cytokinesis. Journal of Cell Biology, 198(5):865–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
5/865.

Beck:2011:CDA

[BPDB^{+}11] Rainer Beck, Simone Prinz, Petra Diestelkötter-Bachert, Simone Röhling, Frank Adolf, Kathrin Hoehner, Sonja Welsch, Paolo Ronchi, Britta Brügger, John A. G. Briggs, and Felix Wieland. Coatomer and dimeric ADP ribosylation factor 1 promote distinct steps in membrane scission. Journal of Cell Biology, 194(5):765–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/5/765.

Bassler:2014:NAF

[BPH^{+}14] Jochen Baßler, Helge Paternoga, Iris Holdermann, Matthias Thoms, Sander Granneman, Clara Barrio-Garcia, Afua Nyarko, Woonghee Lee, Gunter Stier, Sarah A. Clark, Daniel Schraivogel, Martina Kallas, Roland Beckmann, David Tollervey, Elisar Barbar, Irmi Sinning, and Ed Hurt. A network of assembly factors is involved in remodeling rRNA elements during preribosome maturation. Journal of Cell Biology, 207(4): 481–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/4/481.

Brett:2011:ETV

[BPL^{+}11] Christopher L. Brett, Rachael L. Plemel, Braden T. Lobingier, Marissa Vignali, Stanley Fields, and Alexey J. Merz. Efficient termination of vacuolar Rab GTPase signaling requires coordinated action by a GAP and a protein kinase. Journal of Cell Biology, 195(6):1061–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/6/1061.

Bhattaram:2014:SPA

[BPMK^{+}14] Pallavi Bhattaram, Alfredo Penzo-Méndez, Kenji Kato, Kaustav Bandyopadhyay, Abhilash Gadi, Makoto M. Taketo, and Véronique Lefebvre. SOXC proteins amplify canonical WNT signaling to secure nonchondrocytic fates in skeletogenesis. Journal of Cell Biology, 207(5):657–??, December 2014. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/207/5/
657.

Bays:2014:VPD

[BPT^{+}14] Jennifer L. Bays, Xiao Peng, Catlin E. Tolbert, Christophe Guilluy, Ashley E. Angell, Yuan Pan, Richard Superfine, Keith Burridge, and Kris A. DeMali. Vinculin phosphorylation differentially regulates mechanotransduction at cell–cell and cell– matrix adhesions. Journal of Cell Biology, 205(2):251–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
205/2/251.

Brouhard:2014:CTC

[BR14] Gary J. Brouhard and Luke M. Rice. The contribution of αβtubulin curvature to microtubule dynamics. Journal of Cell Biology, 207(3):323–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/3/323.

Brangwynne:2013:PTS

[Bra13] Clifford P. Brangwynne. Phase transitions and size scaling of membrane-less organelles. Journal of Cell Biology, 203(6): 875–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/6/875.

Blumer:2013:RMD

[BRD^{+}13] Julia Blümer, Juliana Rey, Leif Dehmelt, Tomáš Mazel, Yao-Wen Wu, Philippe Bastiaens, Roger S. Goody, and Aymelt Itzen. RabGEFs are a major determinant for specific Rab membrane targeting. Journal of Cell Biology, 200(3):287– ??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/200/3/287.

Bompard:2010:SIP

[BRF^{+}10] Guillaume Bompard, Gabriel Rabeharivelo, Marie Frank, Julien Cau, Claude Delsert, and Nathalie Morin. Subgroup II PAK-mediated phosphorylation regulates Ran activity during mitosis. Journal of Cell Biology, 190(5):807–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
5/807.

Ballister:2014:RMM

[BRL14] Edward R. Ballister, Michelle Riegman, and Michael A. Lampson. Recruitment of Mad1 to metaphase kinetochores is sufficient to reactivate the mitotic checkpoint. Journal of Cell Biology, 204(6):901–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/6/901.

Bulat:2014:PCP

[BRP14] Victoria Bulat, Melanie Rast, and Jan Pielage. Presynaptic CK2 promotes synapse organization and stability by targeting Ankyrin2. Journal of Cell Biology, 204(1):77–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
1/77.

Baker:2013:PDC

[BS13] Darren J. Baker and John M. Sedivy. Probing the depths of cellular senescence. Journal of Cell Biology, 202(1):11–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
1/11.

Burnette:2014:CCA

[BSO^{+}14] Dylan T. Burnette, Lin Shao, Carolyn Ott, Ana M. Pasapera, Robert S. Fischer, Michelle A. Baird, Christelle Der Loughian, Helene Delanoe-Ayari, Matthew J. Paszek, Michael W. Davidson, Eric Betzig, and Jennifer Lippincott-Schwartz. A contractile and counterbalancing adhesion system controls the 3D shape of crawling cells. Journal of Cell Biology, 205(1):83–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
205/1/83.

Brown:2011:GPI

[BSP11] Frank C. Brown, Carmel H. Schindelhaim, and Suzanne R. Pfeffer. GCC185 plays independent roles in Golgi structure maintenance and AP-1-mediated vesicle tethering. Journal of Cell Biology, 194(5):779–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/5/779.

Baldeyron:2011:HRD

[BSR^{+}11a] Céline Baldeyron, Gaston Soria, Danièle Roche, Adam J. L. Cook, and Geneviève Almouzni. HP1α recruitment to DNA damage by p150CAF-1 promotes homologous recombination repair. Journal of Cell Biology, 193(1):81–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/193/1/81.

Bernad:2011:XHCa

[BSR^{+}11b] Rafael Bernad, Patricia Sánchez, Teresa Rivera, Miriam Rodríguez-Corsino, Ekaterina Boyarchuk, Isabelle Vassias, Dominique Ray-Gallet, Alexei Arnaoutov, Mary Dasso, Geneviève Almouzni, and Ana Losada. Xenopus HJURP and condensin II are required for CENP–A assembly. Journal of Cell Biology, 192(4):569–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/4/569.

Bernad:2011:XHCb

[BSR^{+}11c] Rafael Bernad, Patricia Sánchez, Teresa Rivera, Miriam Rodríguez-Corsino, Ekaterina Boyarchuk, Isabelle Vassias, Dominique Ray-Gallet, Alexei Arnaoutov, Mary Dasso, Geneviève Almouzni, and Ana Losada. Xenopus HJURP and condensin II are required for CENP–A assembly. Journal of Cell Biology, 192(5):899–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/5/899.

Bielska:2014:HAC

[BSR^{+}14] Ewa Bielska, Martin Schuster, Yvonne Roger, Adokiye Berepiki, Darren M. Soanes, Nicholas J. Talbot, and Gero Steinberg. Hook is an adapter that coordinates kinesin-3 and dynein cargo attachment on early endosomes. Journal of Cell Biology, 204(6):989–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/6/989.

Bespalov:2011:HSP

[BST^{+}11] Maxim M. Bespalov, Yulia A. Sidorova, Sarka Tumova, Anni Ahonen-Bishopp, Ana Cathia Magalhães, Evgeny Kulesskiy, Mikhail Paveliev, Claudio Rivera, Heikki Rauvala, and Mart Saarma. Heparan sulfate proteoglycan syndecan-3 is a novel receptor for GDNF, neurturin, and artemin. Journal of Cell Biology, 192(1):153–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/1/153.

Baccelli:2012:ECC

[BT12] Irène Baccelli and Andreas Trumpp. The evolving concept of cancer and metastasis stem cells. Journal of Cell Biology, 198 (3):281–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/3/281.

Balasubramanian:2013:TIR

[BT13] Mohan K. Balasubramanian and Evelyn Yaqiong Tao. Timing it right: Precise ON/OFF switches for Rho1 and Cdc42 GT-Pases in cytokinesis. Journal of Cell Biology, 202(2):187–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
202/2/187.

Bondue:2011:DES

[BTC^{+}11] Antoine Bondue, Simon Tännler, Giuseppe Chiapparo, Samira Chabab, Mirana Ramialison, Catherine Paulissen, Benjamin Beck, Richard Harvey, and Cédric Blanpain. Defining the earliest step of cardiovascular progenitor specification during embryonic stem cell differentiation. Journal of Cell Biology, 192 (5):751–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/5/751.

Bergink:2012:RDD

[BTL^{+}12] Steven Bergink, Wendy Toussaint, Martijn S. Luijsterburg, Christoffel Dinant, Sergey Alekseev, Jan H. J. Hoeijmakers, Nico P. Dantuma, Adriaan B. Houtsmuller, and Wim Vermeulen. Recognition of DNA damage by XPC coincides with disruption of the XPC–RAD23 complex. Journal of Cell Biology, 196(6):681–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/6/681.

Bernal:2011:VER

[BV11] Juan A. Bernal and Ashok R. Venkitaraman. A vertebrate N-end rule degron reveals that Orc6 is required in mitosis for daughter cell abscission. Journal of Cell Biology, 192(6): 969–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/6/969.

Bassi:2011:SCK

[BVC^{+}11] Zuni I. Bassi, Koen J. Verbrugghe, Luisa Capalbo, Stephen Gregory, Emilie Montembault, David M. Glover, and Pier Paolo D’Avino. Sticky/Citron kinase maintains proper RhoA localization at the cleavage site during cytokinesis. Journal of Cell Biology, 195(4):595–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/4/595.

Becuwe:2012:MSA

[BVL^{+}12] Michel Becuwe, Neide Vieira, David Lara, Jéssica Gomes-Rezende, Carina Soares-Cunha, Margarida Casal, Rosine Haguenauer-Tsapis, Olivier Vincent, Sandra Paiva, and Sébastien Léon. A molecular switch on an arrestin-like protein relays glucose signaling to transporter endocytosis. Journal of Cell Biology, 196(2):247–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/2/247.

Bonazzi:2011:CPR

[BVM^{+}11] Matteo Bonazzi, Lavanya Vasudevan, Adeline Mallet, Martin Sachse, Anna Sartori, Marie-Christine Prevost, Allison Roberts, Sabrina B. Taner, Jeremy D. Wilbur, Frances M. Brodsky, and Pascale Cossart. Clathrin phosphorylation is required for actin recruitment at sites of bacterial adhesion and internalization. Journal of Cell Biology, 195(3):525–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
195/3/525.

Bentzinger:2014:WSM

[BvMD^{+}14] C. Florian Bentzinger, Julia von Maltzahn, Nicolas A. Dumont, Danny A. Stark, Yu Xin Wang, Kevin Nhan, Jérôme Frenette, D. D. W. Cornelison, and Michael A. Rudnicki. Wnt7a stimulates myogenic stem cell motility and engraftment resulting in improved muscle strength. Journal of Cell Biology, 205(1):97–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/1/97.

Boyd:2011:NDR

[BVR11] Mark T. Boyd, Nikolina Vlatković, and Carlos P. Rubbi. The nucleolus directly regulates p53 export and degradation. Journal of Cell Biology, 194(5):689–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/5/689.

Brooks:2012:CVI

[BW12] Eric R. Brooks and John B. Wallingford. Control of vertebrate intraflagellar transport by the planar cell polarity effector Fuz. Journal of Cell Biology, 198(1):37–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/1/37.

Burridge:2013:TMS

[BW13] Keith Burridge and Erika S. Wittchen. The tension mounts: Stress fibers as force-generating mechanotransducers. Journal of Cell Biology, 200(1):9–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/1/9.

Beaty:2014:TRM

[BWBC^{+}14] Brian T. Beaty, Yarong Wang, Jose Javier Bravo-Cordero, Ved P. Sharma, Veronika Miskolci, Louis Hodgson, and John Condeelis. Talin regulates moesin–NHE–1 recruitment to invadopodia and promotes mammary tumor metastasis. Journal of Cell Biology, 205(5):737–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/205/5/737.

Becker:2011:MIP

[BWK^{+}11] Thomas Becker, Lena-Sophie Wenz, Vivien Krüger, Waltraut Lehmann, Judith M. Müller, Luise Goroncy, Nicole Zufall, Trevor Lithgow, Bernard Guiard, Agnieszka Chacinska, Richard Wagner, Chris Meisinger, and Nikolaus Pfanner. The mitochondrial import protein Mim1 promotes biogenesis of multispanning outer membrane proteins. Journal of Cell Biology, 194(3):387–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/3/387.

Backs:2011:SRM

[BWL^{+}11] Johannes Backs, Barbara C. Worst, Lorenz H. Lehmann, David M. Patrick, Zegeye Jebessa, Michael M. Kreusser, Qiang Sun, Lan Chen, Claudia Heft, Hugo A. Katus, and Eric N. Olson. Selective repression of MEF2 activity by PKAdependent proteolysis of HDAC4. Journal of Cell Biology, 195 (3):403–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/3/403.

Bakkar:2013:INR

[BWL^{+}13] Nadine Bakkar, Jingxin Wang, Katherine J. Ladner, Huating Wang, Jason M. Dahlman, Micheal Carathers, Swarnali Acharyya, Michael A. Rudnicki, Andrew D. Hollenbach, and Denis C. Guttridge. IKK/ NF-κB regulates skeletal myogenesis via a signaling switch to inhibit differentiation and promote mitochondrial biogenesis. Journal of Cell Biology, 202(5): 825–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/5/825.

Budde:2012:CBC

[BWM12] Priya Prakash Budde, Elizabeth H. Williams, and Tom Misteli. Cell biology: At the center of modern biomedicine. Journal of Cell Biology, 199(1):7–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/1/7.

Bassett:2010:ECS

[BWS^{+}10] Emily A. Bassett, Stacey Wood, Kevan J. Salimian, Sandya Ajith, Daniel R. Foltz, and Ben E. Black. Epigenetic centromere specification directs aurora B accumulation but is insufficient to efficiently correct mitotic errors. Journal of Cell Biology, 190(2):177–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/2/177.

Burkart:2012:OCG

[BXB^{+}12] Anna D. Burkart, Bo Xiong, Boris Baibakov, Maria Jiménez-Movilla, and Jurrien Dean. Ovastacin, a cortical granule protease, cleaves ZP2 in the zona pellucida to prevent polyspermy. Journal of Cell Biology, 197(1):37–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/1/37.

Bui:2012:PAA

[BYY^{+}12] Khanh Huy Bui, Toshiki Yagi, Ryosuke Yamamoto, Ritsu Kamiya, and Takashi Ishikawa. Polarity and asymmetry in the arrangement of dynein and related structures in the Chlamydomonas axoneme. Journal of Cell Biology, 198(5): 913–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/5/913.

Brinkmann:2012:NET

[BZ12] Volker Brinkmann and Arturo Zychlinsky. Neutrophil extracellular traps: Is immunity the second function of chromatin? Journal of Cell Biology, 198(5):773–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
5/773.

Cizmecioglu:2010:CAS

[CAB^{+}10] Onur Cizmecioglu, Marc Arnold, Ramona Bahtz, Florian Settele, Lena Ehret, Uta Haselmann-Weiß, Claude Antony, and Ingrid Hoffmann. Cep152 acts as a scaffold for recruitment of Plk4 and CPAP to the centrosome. Journal of Cell Biology, 191(4):731–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/4/731.

Ching:2013:IGL

[CAB^{+}13] Reagan W. Ching, Kashif Ahmed, Paul C. Boutros, Linda Z. Penn, and David P. Bazett-Jones. Identifying gene locus associations with promyelocytic leukemia nuclear bodies using immuno–TRAP. Journal of Cell Biology, 201(2):325–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
2/325.

Chan:2014:LLM

[CAK^{+}14] Keefe T. Chan, Sreeja B. Asokan, Samantha J. King, Tao Bo, Evan S. Dubose, Wenjin Liu, Matthew E. Berginski, Jeremy M. Simon, Ian J. Davis, Shawn M. Gomez, Norman E. Sharpless, and James E. Bear. LKB1 loss in melanoma disrupts directional migration toward extracellular matrix cues. Journal of Cell Biology, 207(2):299–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/207/2/299.

Carmena:2012:FST

[Car12] Mar Carmena. Flies stretch their cells to avoid a chromatin trap. Journal of Cell Biology, 199(5):719–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
5/719.

Cheadle:2012:NSP

[CB12] Lucas Cheadle and Thomas Biederer. The novel synaptogenic protein Farp1 links postsynaptic cytoskeletal dynamics and transsynaptic organization. Journal of Cell Biology, 199(6): 985–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/6/985.

Cowman:2012:CMB

[CBB12] Alan F. Cowman, Drew Berry, and Jake Baum. The cellular and molecular basis for malaria parasite invasion of the human red blood cell. Journal of Cell Biology, 198(6):961–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
198/6/961.

Cavnar:2011:HRN

[CBBH11] Peter J. Cavnar, Erwin Berthier, David J. Beebe, and Anna Huttenlocher. Hax1 regulates neutrophil adhesion and motility through RhoA. Journal of Cell Biology, 193(3):465–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
3/465.

Chapin:2010:CBP

[CC10a] Hannah C. Chapin and Michael J. Caplan. The cell biology of polycystic kidney disease. Journal of Cell Biology, 191(4): 701–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/4/701.

Chong:2010:TGR

[CC10b] S. Y. Christin Chong and Jonah R. Chan. Tapping into the glial reservoir: cells committed to remaining uncommitted. Journal of Cell Biology, 188(3):305–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
3/305.

Celedon:2012:SBT

[CC12] Jose M. Celedon and Kenneth Cline. Stoichiometry for binding and transport by the twin arginine translocation system. Journal of Cell Biology, 197(4):523–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/4/523.

Chawla:2011:AYU

[CCGN11] Aditi Chawla, Sutapa Chakrabarti, Gourisankar Ghosh, and Maho Niwa. Attenuation of yeast UPR is essential for survival and is mediated by IRE1 kinase. Journal of Cell Biology, 193 (1):41–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/1/41.

Courilleau:2012:CRP

[CCJ^{+}12] Céline Courilleau, Catherine Chailleux, Alain Jauneau, Fanny Grimal, Sébastien Briois, Elisa Boutet-Robinet, François Boudsocq, Didier Trouche, and Yvan Canitrot. The chromatin remodeler p400 ATPase facilitates Rad51-mediated repair of DNA double-strand breaks. Journal of Cell Biology, 199(7): 1067–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/7/1067.

Crippa:2011:MMP

[CCM^{+}11] Stefania Crippa, Marco Cassano, Graziella Messina, Daniela Galli, Beatriz G. Galvez, Tomaz Curk, Claudia Altomare, Flavio Ronzoni, Jaan Toelen, Rik Gijsbers, Zeger Debyser, Stefan Janssens, Blaz Zupan, Antonio Zaza, Giulio Cossu, and Maurilio Sampaolesi. miR669a and miR669q prevent skeletal muscle differentiation in postnatal cardiac progenitors. Journal of Cell Biology, 193(7):1197–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/7/1197.

Corbett:2014:NPK

[CD14] Kevin D. Corbett and Arshad Desai. A new piece in the kinetochore jigsaw puzzle. Journal of Cell Biology, 206(4):457–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
206/4/457.

Chao:2010:TPRa

[CDAK10a] Wei-Ting Chao, Alexes C. Daquinag, Felicity Ashcroft, and Jeannette Kunz. Type I PIPK-α regulates directed cell migration by modulating Rac1 plasma membrane targeting and activation. Journal of Cell Biology, 190(2):247–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
2/247.

Chao:2010:TPRb

[CDAK10b] Wei-Ting Chao, Alexes C. Daquinag, Felicity Ashcroft, and Jeannette Kunz. Type I PIPK-α regulates directed cell migration by modulating Rac1 plasma membrane targeting and activation. Journal of Cell Biology, 190(3):479–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
3/479.

Chen:2014:CDC

[CDB^{+}14] Chin-Chi Chen, Mekonnen Lemma Dechassa, Emily Bettini, Mary B. Ledoux, Christian Belisario, Patrick Heun, Karolin Luger, and Barbara G. Mellone. CAL1 is the Drosophila CENP–A assembly factor. Journal of Cell Biology, 204(3): 313–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/3/313.

Caldas:2013:KFP

[CDD13] Gina V. Caldas, Keith F. DeLuca, and Jennifer G. DeLuca. KNL1 facilitates phosphorylation of outer kinetochore proteins by promoting Aurora B kinase activity. Journal of Cell Biology, 203(6):957–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/6/957.

Cai:2014:SVS

[CDH^{+}14] Yiying Cai, Yongqiang Deng, Florian Horenkamp, Karin M. Reinisch, and Christopher G. Burd. Sac1–Vps74 structure reveals a mechanism to terminate phosphoinositide signaling in the Golgi apparatus. Journal of Cell Biology, 206(4):485–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
206/4/485.

Cai:2010:RMA

[CDK^{+}10] Huaqing Cai, Satarupa Das, Yoichiro Kamimura, Yu Long, Carole A. Parent, and Peter N. Devreotes. Ras-mediated activation of the TORC2–PKB pathway is critical for chemotaxis. Journal of Cell Biology, 190(2):233–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/2/233.

Corty:2013:ANC

[CF13] Megan M. Corty and Marc R. Freeman. Architects in neural circuit design: Glia control neuron numbers and connectivity. Journal of Cell Biology, 203(3):395–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
3/395.

Coen:2012:LCH

[CFB^{+}12] Katrijn Coen, Ronald S. Flannagan, Szilvia Baron, Luciene R. Carraro-Lacroix, Dong Wang, Wendy Vermeire, Christine Michiels, Sebastian Munck, Veerle Baert, Shuzo Sugita, Frank Wuytack, Peter Robin Hiesinger, Sergio Grinstein, and Wim Annaert. Lysosomal calcium homeostasis defects, not proton pump defects, cause endo-lysosomal dysfunction in PSENdeficient cells. Journal of Cell Biology, 198(1):23–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
1/23.

Cottarel:2013:NFL

[CFB^{+}13] Jessica Cottarel, Philippe Frit, Oriane Bombarde, Bernard Salles, Aurélie Négrel, Stéphanie Bernard, Penny A. Jeggo, Michael R. Lieber, Mauro Modesti, and Patrick Calsou. A noncatalytic function of the ligation complex during nonhomologous end joining. Journal of Cell Biology, 200(2):173–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
200/2/173.

Cohen:2011:STK

[CFLDM11] David Cohen, Dawn Fernandez, Francisco Lázaro-Diéguez, and Anne Müsch. The serine/threonine kinase Par1b regulates epithelial lumen polarity via IRSp53-mediated cell–ECM signaling. Journal of Cell Biology, 192(3):525–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
3/525.

Cabianca:2010:FCN

[CG10a] Daphne Selvaggia Cabianca and Davide Gabellini. FSHD: copy number variations on the theme of muscular dystrophy. Journal of Cell Biology, 191(6):1049–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/6/1049.

Conrad:2010:AMH

[CG10b] Christian Conrad and Daniel W. Gerlich. Automated microscopy for high-content RNAi screening. Journal of Cell Biology, 188(4):453–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/4/453.

Chen:2012:NSP

[CG12a] Xu Chen and Barry Ganetzky. A neuropeptide signaling pathway regulates synaptic growth in Drosophila. Journal of Cell Biology, 196(4):529–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/196/4/529.

Chioni:2012:FCN

[CG12b] Athina-Myrto Chioni and Richard Grose. FGFR1 cleavage and nuclear translocation regulates breast cancer cell behavior. Journal of Cell Biology, 197(6):801–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/6/801.

Cruz-Garcia:2014:RSC

[CGCP^{+}14] David Cruz-Garcia, Amy J. Curwin, Jean-François Popoff, Caroline Bruns, Juan M. Duran, and Vivek Malhotra. Remodeling of secretory compartments creates CUPS during nutrient starvation. Journal of Cell Biology, 207(6):695–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
6/695.

Chen:2013:LTP

[CGK13] Daniel Chen, Emily S. Gibson, and Matthew J. Kennedy. A light-triggered protein secretion system. Journal of Cell Biology, 201(4):631–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/4/631.

Crider:2012:REM

[CGRS^{+}12] David G. Crider, Luis J. García-Rodríguez, Pallavi Srivastava, Leonardo Peraza-Reyes, Krishna Upadhyaya, Istvan R.

Boldogh, and Liza A. Pon. Rad53 is essential for a mitochondrial DNA inheritance checkpoint regulating G1 to S progression. Journal of Cell Biology, 198(5):793–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
5/793.

Coene:2011:NRB

[CGW^{+}11] Elisabeth D. Coene, Catarina Gadelha, Nicholas White, Ashraf Malhas, Benjamin Thomas, Michael Shaw, and David J. Vaux. A novel role for BRCA1 in regulating breast cancer cell spreading and motility. Journal of Cell Biology, 192 (3):497–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/3/497.

Cevik:2010:JSAa

[CHK^{+}10a] Sebiha Cevik, Yuji Hori, Oktay I. Kaplan, Katarzyna Kida, Tiina Toivenon, Christian Foley-Fisher, David Cottell, Toshiaki Katada, Kenji Kontani, and Oliver E. Blacque. Joubert syndrome Arl13b functions at ciliary membranes and stabilizes protein transport in Caenorhabditis elegans. Journal of Cell Biology, 188(6):953–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/6/953.

Cevik:2010:JSAb

[CHK^{+}10b] Sebiha Cevik, Yuji Hori, Oktay I. Kaplan, Katarzyna Kida, Tiina Toivenon, Christian Foley-Fisher, David Cottell, Toshiaki Katada, Kenji Kontani, and Oliver E. Blacque. Joubert syndrome Arl13b functions at ciliary membranes and stabilizes protein transport in Caenorhabditis elegans. Journal of Cell Biology, 189(1):187–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/1/187.

Chartier:2012:CLO

[CHL12] François J.-M. Chartier, Émilie J.-L. Hardy, and Patrick Laprise. Crumbs limits oxidase-dependent signaling to maintain epithelial integrity and prevent photoreceptor cell death. Journal of Cell Biology, 198(6):991–??, September 2012. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/198/6/
991.

Chan:2014:AEM

[CHL^{+}14] Po-Chao Chan, Rosaline Y. C. Hsu, Chih-Wei Liu, Chien-Chen Lai, and Hong-Chen Chen. Adducin-1 is essential for mitotic spindle assembly through its interaction with myosin–X. Journal of Cell Biology, 204(1):19–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/1/19.

Chadrin:2010:PNT

[CHS^{+}10] Anne Chadrin, Barbara Hess, Mabel San Roman, Xavier Gatti, Bérangère Lombard, Damarys Loew, Yves Barral, Benoit Palancade, and Valérie Doye. Pom33, a novel transmembrane nucleoporin required for proper nuclear pore complex distribution. Journal of Cell Biology, 189(5):795–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
5/795.

Chan:2012:ABC

[CJNS12] Ying Wai Chan, A. Arockia Jeyaprakash, Erich A. Nigg, and Anna Santamaria. Aurora B controls kinetochore–microtubule attachments by inhibiting Ska complex–KMN network interaction. Journal of Cell Biology, 196(5):563–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
5/563.

Caydasi:2010:EKA

[CKO^{+}10] Ayse Koca Caydasi, Bahtiyar Kurtulmus, Maria I. L. Orrico, Astrid Hofmann, Bashar Ibrahim, and Gislene Pereira. Elm1 kinase activates the spindle position checkpoint kinase Kin4. Journal of Cell Biology, 190(6):975–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
6/975.

Cherra:2010:RAP

[CKU^{+}10] Salvatore J. Cherra, Scott M. Kulich, Guy Uechi, Manimalha Balasubramani, John Mountzouris, Billy W. Day, and Charleen T. Chu. Regulation of the autophagy protein LC3 by phosphorylation. Journal of Cell Biology, 190(4):533–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
190/4/533.

Chen:2011:SNR

[CLC^{+}11] Chiung-Ya Chen, Chia-Wen Lin, Chiung-Ying Chang, Si-Tse Jiang, and Yi-Ping Hsueh. Sarm1, a negative regulator of innate immunity, interacts with syndecan-2 and regulates neuronal morphology. Journal of Cell Biology, 193(4):769–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
4/769.

Cheng:2011:SKL

[CLD11] Ling Cheng, Cody Locke, and Graeme W. Davis. S6 kinase localizes to the presynaptic active zone and functions with PDK1 to control synapse development. Journal of Cell Biology, 194(6):921–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/6/921.

Chen:2012:OFR

[CLEZ12] Wei Chen, Jizhong Lou, Evan A. Evans, and Cheng Zhu. Observing force-regulated conformational changes and ligand dissociation from a single integrin on cells. Journal of Cell Biology, 199(3):497–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/3/497.

Chen:2010:AVE

[CLL^{+}10] Tom T. Chen, Alfonso Luque, Sunyoung Lee, Sean M. Anderson, Tatiana Segura, and M. Luisa Iruela-Arispe. Anchorage of VEGF to the extracellular matrix conveys differential signaling responses to endothelial cells. Journal of Cell Biology, 188(4): 595–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/4/595.

Cabrera:2010:PMC

[CLM^{+}10] Margarita Cabrera, Lars Langemeyer, Muriel Mari, Ralf Rethmeier, Ioan Orban, Angela Perz, Cornelia Bröcker, Janice Griffith, Daniel Klose, Heinz-Jürgen Steinhoff, Fulvio Reggiori, Siegfried Engelbrecht-Vandré, and Christian Ungermann. Phosphorylation of a membrane curvature–sensing motif switches function of the HOPS subunit Vps41 in membrane tethering. Journal of Cell Biology, 191(4):845–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
4/845.

Chen:2011:CPK

[CLO^{+}11] Jun-Song Chen, Lucy X. Lu, Melanie D. Ohi, Kevin M. Creamer, Chauca English, Janet F. Partridge, Ryoma Ohi, and Kathleen L. Gould. Cdk1 phosphorylation of the kinetochore protein Nsk1 prevents error-prone chromosome segregation. Journal of Cell Biology, 195(4):583–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
4/583.

Choi:2010:CSM

[CLS^{+}10] Yuk-Kwan Choi, Pengfei Liu, Siu Kwan Sze, Chao Dai, and Robert Z. Qi. CDK5RAP2 stimulates microtubule nucleation by the γ-tubulin ring complex. Journal of Cell Biology, 191(6): 1089–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/6/1089.

Chia:2013:CMM

[CLS13] Poh Hui Chia, Pengpeng Li, and Kang Shen. Cellular and molecular mechanisms underlying presynapse formation. Journal of Cell Biology, 203(1):11–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/1/11.

Castilla-Llorente:2012:MGT

[CLSO^{+}12] Virginia Castilla-Llorente, Lee Spraggon, Miwako Okamura, Saif Naseeruddin, Matthew Adamow, Sarah Qamar, and Jidong Liu. Mammalian GW220/TNGW1 is essential for the formation of GW/P bodies containing miRISC. Journal of Cell Biology, 198(4):529–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/4/529.

Chi:2014:FSB

[CLW^{+}14] Richard J. Chi, Jingxuan Liu, Matthew West, Jing Wang, Greg Odorizzi, and Christopher G. Burd. Fission of SNX– BAR–coated endosomal retrograde transport carriers is promoted by the dynamin-related protein Vps1. Journal of Cell Biology, 204(5):793–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/5/793.

Cohen:2014:TRP

[CLZ^{+}14] Shenhav Cohen, Donghoon Lee, Bo Zhai, Steven P. Gygi, and Alfred L. Goldberg. Trim32 reduces PI3K–Akt–FoxO signaling in muscle atrophy by promoting plakoglobin–PI3K dissociation. Journal of Cell Biology, 204(5):747–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
5/747.

Cai:2012:CHT

[CM12a] Liang Cai and Keith E. Mostov. Cell height: Tao rising. Journal of Cell Biology, 199(7):1023–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/7/1023.

Cane:2012:BKL

[CM12b] Stuart Cane and Thomas J. Maresca. Balancing the kinetochore ledger. Journal of Cell Biology, 198(4):477–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
4/477.

Christis:2012:SGP

[CM12c] Chantal Christis and Sean Munro. The small G protein Arl1 directs the trans-Golgi-specific targeting of the Arf1 exchange factors BIG1 and BIG2. Journal of Cell Biology, 196(3): 327–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/3/327.

Contreras:2013:DPI

[CMD^{+}13] Amelia U. Contreras, Yohannes Mebratu, Monica Delgado, Gilbert Montano, Chien an A. Hu, Stefan W. Ryter, Augustine M. K. Choi, Yuting Lin, Jialing Xiang, Hitendra Chand, and Yohannes Tesfaigzi. Deacetylation of p53 induces autophagy by suppressing Bmf expression. Journal of Cell Biology, 201 (3):427–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/3/427.

Chen:2010:TRV

[CMH^{+}10] Li Chen, Annalisa Mupo, Tuong Huynh, Sara Cioffi, Matthew Woods, Chengliu Jin, Wallace McKeehan, LuAnn Thompson-Snipes, Antonio Baldini, and Elizabeth Illingworth. Tbx1 regulates Vegfr3 and is required for lymphatic vessel development. Journal of Cell Biology, 189(3):417–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/3/417.

Carroll:2010:DRC

[CMS10] Christopher W. Carroll, Kirstin J. Milks, and Aaron F. Straight. Dual recognition of CENP–A nucleosomes is required for centromere assembly. Journal of Cell Biology, 189 (7):1143–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/7/1143.

Carmena:2011:RRR

[CMS11] Ana Carmena, Aljona Makarova, and Stephan Speicher. The Rap1–Rgl–Ral signaling network regulates neuroblast cortical polarity and spindle orientation. Journal of Cell Biology, 195(4):553–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/4/553.

Chacon:2014:PTS

[CMS^{+}14] Jeremy M. Chacón, Soumya Mukherjee, Breanna M. Schuster, Duncan J. Clarke, and Melissa K. Gardner. Pericentromere tension is self-regulated by spindle structure in metaphase. Journal of Cell Biology, 205(3):313–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/205/3/313.

Crawley:2014:SIB

[CMT14] Scott W. Crawley, Mark S. Mooseker, and Matthew J. Tyska. Shaping the intestinal brush border. Journal of Cell Biology, 207(4):441–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/4/441.

Carroll:2011:DHB

[CMW11] Johanna S. Carroll, Sarah E. Munchel, and Karsten Weis. The DExD/H box ATPase Dhh1 functions in translational repression, mRNA decay, and processing body dynamics. Journal of Cell Biology, 194(4):527–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/4/527.

Chen:2012:CTU

[CNP^{+}12] Jing Chen, Oxana E. Nekrasova, Dipal M. Patel, Jodi L. Klessner, Lisa M. Godsel, Jennifer L. Koetsier, Evangeline V. Amargo, Bhushan V. Desai, and Kathleen J. Green. The Cterminal unique region of desmoglein 2 inhibits its internalization via tail–tail interactions. Journal of Cell Biology, 199(4): 699–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/4/699.

Campbell:2013:LCA

[CO13] Douglas S. Campbell and Hitoshi Okamoto. Local caspase activation interacts with Slit–Robo signaling to restrict axonal arborization. Journal of Cell Biology, 203(4):657–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
4/657.

Claypool:2012:CDI

[COB^{+}12] Steven M. Claypool, Yavuz Oktay, Pinmanee Boontheung, Joseph A. Loo, and Carla M. Koehler. Cardiolipin defines the interactome of the major ADP/ATP carrier protein of the mitochondrial inner membrane. Journal of Cell Biology, 197 (7):1029–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/7/1029.

Carbonaro:2011:MDT

[COG11] Marisa Carbonaro, Aurora O’Brate, and Paraskevi Giannakakou. Microtubule disruption targets HIF-1α mRNA to cytoplasmic P-bodies for translational repression. Journal of Cell Biology, 192(1):83–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/1/83.

Cooper:2013:MCM

[Coo13] Jonathan A. Cooper. Mechanisms of cell migration in the nervous system. Journal of Cell Biology, 202(5):725–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
202/5/725.

Clijsters:2013:SCA

[COW13] Linda Clijsters, Janneke Ogink, and Rob Wolthuis. The spindle checkpoint, APC/C^{Cdc20}, and APC/C^{Cdh1} play distinct roles in connecting mitosis to S phase. Journal of Cell Biology, 201(7):1013–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/7/1013.

Chen:2011:AFS

[CP11] Qian Chen and Thomas D. Pollard. Actin filament severing by cofilin is more important for assembly than constriction of the cytokinetic contractile ring. Journal of Cell Biology, 195 (3):485–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/3/485.

Cerruti:2013:PCD

[CPS^{+}13] Benedetta Cerruti, Alberto Puliafito, Annette M. Shewan, Wei Yu, Alexander N. Combes, Melissa H. Little, Federica Chianale, Luca Primo, Guido Serini, Keith E. Mostov, Antonio Celani, and Andrea Gamba. Polarity, cell division, and out-of-equilibrium dynamics control the growth of epithelial structures. Journal of Cell Biology, 203(2):359–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
2/359.

Chambers:2012:ARA

[CPT^{+}12] Joseph E. Chambers, Kseniya Petrova, Giulia Tomba, Michele Vendruscolo, and David Ron. ADP ribosylation adapts an ER chaperone response to short-term fluctuations in unfolded protein load. Journal of Cell Biology, 198(3):371–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
3/371.

Chambers:2014:ARA

[CPT^{+}14] Joseph E. Chambers, Kseniya Petrova, Giulia Tomba, Michele Vendruscolo, and David Ron. ADP ribosylation adapts an ER chaperone response to short-term fluctuations in unfolded protein load. Journal of Cell Biology, 207(4):569–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
4/569.

Choi:2011:LMC

[CPX11] Won-Seok Choi, Richard D. Palmiter, and Zhengui Xia. Loss of mitochondrial complex I activity potentiates dopamine neuron death induced by microtubule dysfunction in a Parkinson’s disease model. Journal of Cell Biology, 192(5):873–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
5/873.

Chaudhari:2010:CBTa

[CR10a] Nirupa Chaudhari and Stephen D. Roper. The cell biology of taste. Journal of Cell Biology, 190(3):285–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
3/285.

Chaudhari:2010:CBTb

[CR10b] Nirupa Chaudhari and Stephen D. Roper. The cell biology of taste. Journal of Cell Biology, 191(2):429–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
2/429.

Cotto-Rios:2011:ACD

[CRJB^{+}11] Xiomaris M. Cotto-Rios, Mathew J. K. Jones, Luca Busino, Michele Pagano, and Tony T. Huang. APC/C^{Cdh1}-dependent proteolysis of USP1 regulates the response to UV-mediated DNA damage. Journal of Cell Biology, 194(2):177–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
2/177.

Corrigan:2014:BRE

[CRL^{+}14] Laura Corrigan, Siamak Redhai, Aaron Leiblich, Shih-Jung Fan, Sumeth M. W. Perera, Rachel Patel, Carina Gandy, S. Mark Wainwright, John F. Morris, Freddie Hamdy, Deborah C. I. Goberdhan, and Clive Wilson. BMP-regulated exosomes from Drosophila male reproductive glands reprogram female behavior. Journal of Cell Biology, 206(5):671–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
5/671.

Chinthalapudi:2014:LBP

[CRP^{+}14] Krishna Chinthalapudi, Erumbi S. Rangarajan, Dipak N. Patil, Eric M. George, David T. Brown, and Tina Izard. Lipid binding promotes oligomerization and focal adhesion activity of vinculin. Journal of Cell Biology, 207(5):643–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
5/643.

Chen:2013:KSC

[CS13] Yanmin Chen and Zu-Hang Sheng. Kinesin-1-syntaphilin coupling mediates activity-dependent regulation of axonal mitochondrial transport. Journal of Cell Biology, 202(2):351–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
202/2/351.

Carvalho-Santos:2011:TOCa

[CSAPLBD11a] Zita Carvalho-Santos, Juliette Azimzadeh, José. B. Pereira-Leal, and Mónica Bettencourt-Dias. Tracing the origins of centrioles, cilia, and flagella. Journal of Cell Biology, 194 (2):165–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/2/165.

Carvalho-Santos:2011:TOCb

[CSAPLBD11b] Zita Carvalho-Santos, Juliette Azimzadeh, José. B. Pereira-Leal, and Mónica Bettencourt-Dias. Tracing the origins of centrioles, cilia, and flagella. Journal of Cell Biology, 195(2): 341–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/2/341.

Courtois:2012:TMM

[CSEH12] Aurélien Courtois, Melina Schuh, Jan Ellenberg, and Takashi Hiiragi. The transition from meiotic to mitotic spindle assembly is gradual during early mammalian development. Journal of Cell Biology, 198(3):357–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/3/357.

Cherepanova:2014:OAN

[CSG14] Natalia A. Cherepanova, Shiteshu Shrimal, and Reid Gilmore. Oxidoreductase activity is necessary for N-glycosylation of cysteine-proximal acceptor sites in glycoproteins. Journal of Cell Biology, 206(4):525–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/206/4/525.

Conomos:2012:VRI

[CSH^{+}12] Dimitri Conomos, Michael D. Stutz, Mark Hills, Axel A. Neumann, Tracy M. Bryan, Roger R. Reddel, and Hilda A. Pickett. Variant repeats are interspersed throughout the telomeres and recruit nuclear receptors in ALT cells. Journal of Cell Biology, 199(6):893–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/6/893.

Civelekoglu-Scholey:2013:DBPa

[CSHS^{+}13a] Gul Civelekoglu-Scholey, Bin He, Muyao Shen, Xiaohu Wan, Emanuele Roscioli, Brent Bowden, and Daniela Cimini. Dynamic bonds and polar ejection force distribution explain kinetochore oscillations in PtK1 cells. Journal of Cell Biology, 201 (4):577–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/4/577.

Civelekoglu-Scholey:2013:DBPb

[CSHS^{+}13b] Gul Civelekoglu-Scholey, Bin He, Muyao Shen, Xiaohu Wan, Emanuele Roscioli, Brent Bowden, and Daniela Cimini. Dynamic bonds and polar ejection force distribution explain kinetochore oscillations in PtK1 cells. Journal of Cell Biology, 202 (3):597–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/3/597.

Coffman:2013:FCF

[CSKW13] Valerie C. Coffman, Jennifer A. Sees, David R. Kovar, and Jian-Qiu Wu. The formins Cdc12 and For3 cooperate during contractile ring assembly in cytokinesis. Journal of Cell Biology, 203(1):101–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/1/101.

Cortes:2012:FYA

[CSM^{+}12] Juan Carlos G. Cortés, Mamiko Sato, Javier Muñoz, M. Belén Moreno, Jose Angel Clemente-Ramos, Mariona Ramos, Hitoshi Okada, Masako Osumi, Angel Durán, and Juan Carlos Ribas. Fission yeast Ags1 confers the essential septum strength needed for safe gradual cell abscission. Journal of Cell Biology, 198(4):637–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/4/637.

Chairoungdua:2010:ERC

[CSP^{+}10] Arthit Chairoungdua, Danielle L. Smith, Pierre Pochard, Michael Hull, and Michael J. Caplan. Exosome release of β-catenin: a novel mechanism that antagonizes Wnt signaling. Journal of Cell Biology, 190(6):1079–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
6/1079.

Chiyoda:2012:LWP

[CSS^{+}12] Tatsuyuki Chiyoda, Naoyuki Sugiyama, Takatsune Shimizu, Hideaki Naoe, Yusuke Kobayashi, Jo Ishizawa, Yoshimi Arima, Hiroshi Tsuda, Masaaki Ito, Kozo Kaibuchi, Daisuke Aoki, Yasushi Ishihama, Hideyuki Saya, and Shinji Kuninaka. LATS1/ WARTS phosphorylates MYPT1 to counteract PLK1 and regulate mammalian mitotic progression. Journal of Cell Biology, 197(5):625–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/5/625.

Chen:2014:SPM

[CSS^{+}14] Jingjing Chen, Christine J. Smoyer, Brian D. Slaughter, Jay R. Unruh, and Sue L. Jaspersen. The SUN protein Mps3 controls Ndc1 distribution and function on the nuclear membrane. Journal of Cell Biology, 204(4):523–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
4/523.

Civelekoglu-Scholey:2010:PSM

[CSTBM^{+}10] Gul Civelekoglu-Scholey, Li Tao, Ingrid Brust-Mascher, Roy Wollman, and Jonathan M. Scholey. Prometaphase spindle maintenance by an antagonistic motor-dependent force balance made robust by a disassembling lamin–B envelope. Journal of Cell Biology, 188(1):49–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/188/1/49.

Chung:2010:MMM

[CT10] Sunglan Chung and Peter A. Takizawa. Multiple Myo4 motors enhance ASH1 mRNA transport in Saccharomyces cerevisiae. Journal of Cell Biology, 189(4):755–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/4/755.

Craige:2010:CTF

[CTD^{+}10] Branch Craige, Che-Chia Tsao, Dennis R. Diener, Yuqing Hou, Karl-Ferdinand Lechtreck, Joel L. Rosenbaum, and George B. Witman. CEP290 tethers flagellar transition zone microtubules to the membrane and regulates flagellar protein content. Journal of Cell Biology, 190(5):927–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
5/927.

Chen:2011:PTH

[CTH^{+}11] Xin Chen, Hanna Tukachinsky, Chih-Hsiang Huang, Cindy Jao, Yue-Ru Chu, Hsiang-Yun Tang, Britta Mueller, Sol Schulman, Tom A. Rapoport, and Adrian Salic. Processing and turnover of the Hedgehog protein in the endoplasmic reticulum. Journal of Cell Biology, 192(5):825–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
5/825.

Chen:2010:MMR

[CTL^{+}10] Jian-Fu Chen, Yazhong Tao, Juan Li, Zhongliang Deng, Zhen Yan, Xiao Xiao, and Da-Zhi Wang. microRNA-1 and microRNA-206 regulate skeletal muscle satellite cell proliferation and differentiation by repressing Pax7. Journal of Cell Biology, 190(5):867–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/190/5/867.

Cain:2014:SPU

[CTM^{+}14a] Natalie E. Cain, Erin C. Tapley, Kent L. McDonald, Benjamin M. Cain, and Daniel A. Starr. The SUN protein UNC-84 is required only in force-bearing cells to maintain nuclear envelope architecture. Journal of Cell Biology, 206(2):163–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
2/163.

Couturier:2014:FTA

[CTM^{+}14b] Lydie Couturier, Mateusz Trylinski, Khallil Mazouni, Léa Darnet, and François Schweisguth. A fluorescent tagging approach in Drosophila reveals late endosomal trafficking of Notch and Sanpodo. Journal of Cell Biology, 207(3):351–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
3/351.

Concannon:2010:AKM

[CTW^{+}10] Caoimhín G. Concannon, Liam P. Tuffy, Petronela Weisová, Helena P. Bonner, David Dávila, Caroline Bonner, Marc C. Devocelle, Andreas Strasser, Manus W. Ward, and Jochen H. M. Prehn. AMP kinase-mediated activation of the BH3only protein Bim couples energy depletion to stress-induced apoptosis. Journal of Cell Biology, 189(1):83–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/189/1/83.

Chai:2012:ARP

[CTY^{+}12] Yongping Chai, Dong Tian, Yihong Yang, Guoxin Feng, Ze Cheng, Wei Li, and Guangshuo Ou. Apoptotic regulators promote cytokinetic midbody degradation in C. elegans. Journal of Cell Biology, 199(7):1047–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/7/1047.

Caillava:2011:CLA

[CVJ^{+}11] Céline Caillava, Renaud Vandenbosch, Beata Jablonska, Cyrille Deboux, Giulia Spigoni, Vittorio Gallo, Brigitte Malgrange, and Anne Baron-Van Evercooren. Cdk2 loss accelerates precursor differentiation and remyelination in the adult central nervous system. Journal of Cell Biology, 193(2):397–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
193/2/397.

Cain:2010:PPI

[CVR10] Robert J. Cain, Bart Vanhaesebroeck, and Anne J. Ridley. The PI3K p110α isoform regulates endothelial adherens junctions via Pyk2 and Rac1. Journal of Cell Biology, 188(6):863–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
188/6/863.

Cijsouw:2014:MRN

[CWB^{+}14] Tony Cijsouw, Jens P. Weber, Jurjen H. Broeke, Jantine A. C. Broek, Desiree Schut, Tim Kroon, Ingrid Saarloos, Matthijs Verhage, and Ruud F. Toonen. Munc18-1 redistributes in nerve terminals in an activity- and PKC-dependent manner. Journal of Cell Biology, 204(5):759–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/5/759.

Choi:2013:PSF

[CWC^{+}13] Chang-Hoon Choi, Bradley A. Webb, Michael S. Chimenti, Matthew P. Jacobson, and Diane L. Barber. pH sensing by FAK–His58 regulates focal adhesion remodeling. Journal of Cell Biology, 202(6):849–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/6/849.

Colombi:2013:TNP

[CWFL13] Paolo Colombi, Brant M. Webster, Florian Fröhlich, and C. Patrick Lusk. The transmission of nuclear pore complexes to daughter cells requires a cytoplasmic pool of Nsp1. Journal of Cell Biology, 203(2):215–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/2/215.

Couto:2011:PRN

[CWG^{+}11] C. Anne-Marie Couto, Hong-Yu Wang, Joanna C. A. Green, Rhian Kiely, Robert Siddaway, Christine Borer, Catherine J. Pears, and Nicholas D. Lakin. PARP regulates nonhomologous end joining through retention of Ku at double-strand breaks. Journal of Cell Biology, 194(3):367–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/3/367.

Calvert:2011:MCU

[CWL^{+}11a] Meredith E. K. Calvert, Graham D. Wright, Fong Yew Leong, Keng-Hwee Chiam, Yinxiao Chen, Gregory Jedd, and Mohan K. Balasubramanian. Myosin concentration underlies cell size-dependent scalability of actomyosin ring constriction. Journal of Cell Biology, 195(5):799–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
5/799.

Chen:2011:GPF

[CWL^{+}11b] Yuejun Chen, Feifei Wang, Hui Long, Ying Chen, Ziyan Wu, and Lan Ma. GRK5 promotes F-actin bundling and targets bundles to membrane structures to control neuronal morphogenesis. Journal of Cell Biology, 194(6):905–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
6/905.

Coffman:2011:CEM

[CWPW11] Valerie C. Coffman, Pengcheng Wu, Mark R. Parthun, and Jian-Qiu Wu. CENP-A exceeds microtubule attachment sites in centromere clusters of both budding and fission yeast. Journal of Cell Biology, 195(4):563–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/4/563.

Claypool:2011:BSM

[CWS^{+}11] Steven M. Claypool, Kevin Whited, Santi Srijumnong, Xianlin Han, and Carla M. Koehler. Barth syndrome mutations that cause tafazzin complex lability. Journal of Cell Biology, 192 (3):447–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/3/447.

Chen:2012:RMV

[CWZ^{+}12] Yu Chen, Yan Wang, Jinzhong Zhang, Yongqiang Deng, Li Jiang, Eli Song, Xufeng S. Wu, John A. Hammer, Tao Xu, and Jennifer Lippincott-Schwartz. Rab10 and myosin–Va mediate insulin-stimulated GLUT4 storage vesicle translocation in adipocytes. Journal of Cell Biology, 198(4):545–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
4/545.

Cane:2013:EPE

[CYLMM13] Stuart Cane, Anna A. Ye, Sasha J. Luks-Morgan, and Thomas J. Maresca. Elevated polar ejection forces stabilize kinetochore–microtubule attachments. Journal of Cell Biology, 200(2):203–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/2/203.

Chen:2013:AAL

[CYN^{+}13] Zu-Lin Chen, Yao Yao, Erin H. Norris, Anna Kruyer, Odella Jno-Charles, Akbarshakh Akhmerov, and Sidney Strickland. Ablation of astrocytic laminin impairs vascular smooth muscle cell function and leads to hemorrhagic stroke. Journal of Cell Biology, 202(2):381–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/2/381.

Ciavarra:2010:RMD

[CZ10] Giovanni Ciavarra and Eldad Zacksenhaus. Rescue of myogenic defects in Rb-deficient cells by inhibition of autophagy or by hypoxia-induced glycolytic shift. Journal of Cell Biology, 191(2):291–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/2/291.

Carnell:2011:APD

[CZC^{+}11] Michael Carnell, Tobias Zech, Simon D. Calaminus, Seiji Ura, Monica Hagedorn, Simon A. Johnston, Robin C. May, Thierry Soldati, Laura M. Machesky, and Robert H. Insall. Actin polymerization driven by WASH causes V–ATPase retrieval and vesicle neutralization before exocytosis. Journal of Cell Biology, 193(5):831–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/5/831.

Catalucci:2013:ART

[CZD^{+}13] Daniele Catalucci, Deng-Hong Zhang, Jaime DeSantiago, Franck Aimond, Guillaume Barbara, Jean Chemin, Désiré Bonci, Eckard Picht, Francesca Rusconi, Nancy D. Dalton, Kirk L. Peterson, Sylvain Richard, Donald M. Bers, Joan Heller Brown, and Gianluigi Condorelli. Akt regulates L-type Ca^{2+} channel activity by modulating Ca_{v} α1 protein stability. Journal of Cell Biology, 200(6):851–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/200/
6/851.

Cohen:2012:UTC

[CZGG12] Shenhav Cohen, Bo Zhai, Steven P. Gygi, and Alfred L. Goldberg. Ubiquitylation by Trim32 causes coupled loss of desmin, Z-bands, and thin filaments in muscle atrophy. Journal of Cell Biology, 198(4):575–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/4/575.

Chen:2014:GAA

[CZM^{+}14] Rui Chen, Yilong Zou, Dongxue Mao, Daxiao Sun, Guanguang Gao, Jingwen Shi, Xiaoqing Liu, Chen Zhu, Mingyu Yang, Wanlu Ye, Qianqian Hao, Ruiqiang Li, and Li Yu. The general amino acid control pathway regulates mTOR and autophagy during serum/glutamine starvation. Journal of Cell Biology, 206(2):173–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/2/173.

Darbellay:2011:SNA

[DAB^{+}11] Basile Darbellay, Serge Arnaudeau, Charles R. Bader, Stephane Konig, and Laurent Bernheim. STIM1L is a new actin-binding splice variant involved in fast repetitive Ca^{2+} release. Journal of Cell Biology, 194(2):335–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/194/2/

Danen:2014:IMB

[Dan14] Erik H. J. Danen. Ignoring matrix boundaries when the LKB1 master kinase is gone. Journal of Cell Biology, 207 (2):167–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/2/167.

Duran:2010:USA

[DAS^{+}10] Juan M. Duran, Christophe Anjard, Chris Stefan, William F. Loomis, and Vivek Malhotra. Unconventional secretion of Acb1 is mediated by autophagosomes. Journal of Cell Biology, 188(4):527–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/4/527.

DeMay:2011:SFE

[DBH^{+}11] Bradley S. DeMay, Xiaobo Bai, Louisa Howard, Patricia Occhipinti, Rebecca A. Meseroll, Elias T. Spiliotis, Rudolf Oldenbourg, and Amy S. Gladfelter. Septin filaments exhibit a dynamic, paired organization that is conserved from yeast to mammals. Journal of Cell Biology, 193(6):1065–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
6/1065.

Dulle:2013:SOS

[DBUT13] Jennifer E. Dulle, Rachel E. Bouttenot, Lisa A. Underwood, and Heather L. True. Soluble oligomers are sufficient for transmission of a yeast prion but do not confer phenotype. Journal of Cell Biology, 203(2):197–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/2/197.

Davies:2012:SMR

[DC12] Tim Davies and Julie C. Canman. Stuck in the middle: Rac, adhesion, and cytokinesis. Journal of Cell Biology, 198(5): 769–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/5/769.

Dores:2012:ABY

[DCL^{+}12] Michael R. Dores, Buxin Chen, Huilan Lin, Unice J. K. Soh, May M. Paing, William A. Montagne, Timo Meerloo, and JoAnn Trejo. ALIX binds a YPX_{3} L motif of the GPCR PAR1 and mediates ubiquitin-independent ESCRT-III/MVB sorting. Journal of Cell Biology, 197(3):407–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/3/407.

DiBartolomeo:2010:DIA

[DCN^{+}10] Sabrina Di Bartolomeo, Marco Corazzari, Francesca Nazio, Serafina Oliverio, Gaia Lisi, Manuela Antonioli, Vittoria Pagliarini, Silvia Matteoni, Claudia Fuoco, Luigi Giunta, Marcello D’Amelio, Roberta Nardacci, Alessandra Romagnoli, Mauro Piacentini, Francesco Cecconi, and Gian Maria Fimia. The dynamic interaction of AMBRA1 with the dynein motor complex regulates mammalian autophagy. Journal of Cell Biology, 191(1):155–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/1/155.

Dornier:2012:TTR

[DCO^{+}12] Emmanuel Dornier, Franck Coumailleau, Jean-François Ottavi, Julien Moretti, Claude Boucheix, Philippe Mauduit, François Schweisguth, and Eric Rubinstein. TspanC8 tetraspanins regulate ADAM10/Kuzbanian trafficking and promote Notch activation in flies and mammals. Journal of Cell Biology, 199(3):481–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/3/481. See correc-

tion [DCO^{+}16].

Dreesen:2013:LBF

[DCO^{+}13] Oliver Dreesen, Alexandre Chojnowski, Peh Fern Ong, Tian Yun Zhao, John E. Common, Declan Lunny, E. Birgitte Lane, Shu Jin Lee, Leah A. Vardy, Colin L. Stewart, and Alan Colman. Lamin B1 fluctuations have differential effects on cellular proliferation and senescence. Journal of Cell Biology, 200(5):605–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/5/605.

Dornier:2016:CTT

[DCO^{+}16] Emmanuel Dornier, Franck Coumailleau, Jean-François Ottavi, Julien Moretti, Claude Boucheix, Philippe Mauduit, François Schweisguth, and Eric Rubinstein. Correction:

TspanC8 tetraspanins regulate ADAM10/Kuzbanian trafficking and promote Notch activation in flies and mammals. Journal of Cell Biology, 213(4):495–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/213/4/495. See

[DCO^{+}12].

Dou:2010:CIP

[DCP^{+}10] Zhixun Dou, Mohar Chattopadhyay, Ji-An Pan, Jennifer L. Guerriero, Ya-Ping Jiang, Lisa M. Ballou, Zhenyu Yue, Richard Z. Lin, and Wei-Xing Zong. The class IA phosphatidylinositol 3-kinase p110-β subunit is a positive regulator of autophagy. Journal of Cell Biology, 191(4):827–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
191/4/827.

Delaval:2010:PCF

[DD10a] Benedicte Delaval and Stephen J. Doxsey. Pericentrin in cellular function and disease. Journal of Cell Biology, 188(2): 181–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/2/181.

Douglas:2010:PHA

[DD10b] Peter M. Douglas and Andrew Dillin. Protein homeostasis and aging in neurodegeneration. Journal of Cell Biology, 190(5): 719–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/5/719.

David:2012:NTV

[DDH^{+}12] Alexandre David, Brian P. Dolan, Heather D. Hickman, Jonathan J. Knowlton, Giovanna Clavarino, Philippe Pierre, Jack R. Bennink, and Jonathan W. Yewdell. Nuclear translation visualized by ribosome-bound nascent chain puromycylation. Journal of Cell Biology, 197(1):45–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/197/1/45.

Dultz:2010:LIS

[DE10] Elisa Dultz and Jan Ellenberg. Live imaging of single nuclear pores reveals unique assembly kinetics and mechanism in interphase. Journal of Cell Biology, 191(1):15–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/191/1/15.

Devenport:2014:CBP

[Dev14] Danelle Devenport. The cell biology of planar cell polarity. Journal of Cell Biology, 207(2):171–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/207/2/171.

Diagouraga:2014:MDM

[DGF^{+}14] Boubou Diagouraga, Alexei Grichine, Arnold Fertin, Jin Wang, Saadi Khochbin, and Karin Sadoul. Motor-driven marginal band coiling promotes cell shape change during platelet activation. Journal of Cell Biology, 204(2):177– ??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/2/177.

DeVorkin:2014:DEC

[DGH^{+}14] Lindsay DeVorkin, Nancy Erro Go, Ying-Chen Claire Hou, Annie Moradian, Gregg B. Morin, and Sharon M. Gorski. The Drosophila effector caspase Dcp-1 regulates mitochondrial dynamics and autophagic flux via SesB. Journal of Cell Biology, 205(4):477–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/4/477.

Deeg:2010:BRF

[DGS^{+}10] Sebastian Deeg, Mathias Gralle, Kamila Sroka, Mathias Bähr, Fred Silvester Wouters, and Pawel Kermer. BAG1 restores formation of functional DJ-1 L166P dimers and DJ-1 chaperone activity. Journal of Cell Biology, 188(4):505–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
4/505.

Day:2011:IST

[DGS^{+}11] Michele E. Day, Guido M. Gaietta, Mira Sastri, Antonius Koller, Mason R. Mackey, John D. Scott, Guy A. Perkins, Mark H. Ellisman, and Susan S. Taylor. Isoform-specific targeting of PKA to multivesicular bodies. Journal of Cell Biology, 193(2):347–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/2/347.

Dolat:2014:SPS

[DHB^{+}14] Lee Dolat, John L. Hunyara, Jonathan R. Bowen, Eva Pauline Karasmanis, Maha Elgawly, Vitold E. Galkin, and Elias T. Spiliotis. Septins promote stress fiber-mediated maturation of focal adhesions and renal epithelial motility. Journal of Cell Biology, 207(2):225–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/2/225.

Dunsch:2012:DLC

[DHL^{+}12] Anja K. Dunsch, Dean Hammond, Jennifer Lloyd, Lothar Schermelleh, Ulrike Gruneberg, and Francis A. Barr. Dynein light chain 1 and a spindle-associated adaptor promote dynein asymmetry and spindle orientation. Journal of Cell Biology, 198(6):1039–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/6/1039.

Distel:2010:CNPa

[DHVK10a] Martin Distel, Jennifer C. Hocking, Katrin Volkmann, and Reinhard W. Köster. The centrosome neither persistently leads migration nor determines the site of axonogenesis in migrating neurons in vivo. Journal of Cell Biology, 191(4): 875–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/4/875.

Distel:2010:CNPb

[DHVK10b] Martin Distel, Jennifer C. Hocking, Katrin Volkmann, and Reinhard W. Köster. The centrosome neither persistently leads migration nor determines the site of axonogenesis in migrating neurons in vivo. Journal of Cell Biology, 191(7):1413–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
191/7/1413.

Duan:2012:GPF

[DJL^{+}12] Rui Duan, Peng Jin, Fengbao Luo, Guofeng Zhang, Nathan Anderson, and Elizabeth H. Chen. Group I PAKs function downstream of Rac to promote podosome invasion during myoblast fusion in vivo. Journal of Cell Biology, 199(1): 169–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/1/169.

Perez:2011:BPM

[dJPAA^{+}11] Vinicio A. de Jesus Perez, Ziad Ali, Tero-Pekka Alastalo, Fumiaki Ikeno, Hirofumi Sawada, Ying-Ju Lai, Thomas Kleisli, Edda Spiekerkoetter, Xiumei Qu, Laura H. Rubinos, Euan Ashley, Manuel Amieva, Shoukat Dedhar, and Marlene Rabinovitch. BMP promotes motility and represses growth of smooth muscle cells by activation of tandem Wnt pathways. Journal of Cell Biology, 192(1):171–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/1/171.

deJong:2012:DPM

[dJSTV12] Arthur P. H. de Jong, Sabine K. Schmitz, Ruud F. G. Toonen, and Matthijs Verhage. Dendritic position is a major determinant of presynaptic strength. Journal of Cell Biology, 197 (2):327–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/2/327.

Davies:2010:HSS

[DK10a] Alexander E. Davies and Kenneth B. Kaplan. Hsp90–Sgt1 and Skp1 target human Mis12 complexes to ensure efficient formation of kinetochore–microtubule binding sites. Journal of Cell Biology, 189(2):261–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/2/261.

Ducy:2010:TFS

[DK10b] Patricia Ducy and Gerard Karsenty. The two faces of serotonin in bone biology. Journal of Cell Biology, 191(1):7–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
1/7.

Dubash:2013:GBA

[DKA^{+}13] Adi D. Dubash, Jennifer L. Koetsier, Evangeline V. Amargo, Nicole A. Najor, Robert M. Harmon, and Kathleen J. Green.

The GEF Bcr activates RhoA/MAL signaling to promote keratinocyte differentiation via desmoglein-1. Journal of Cell Biology, 202(4):653–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/4/653.

Du:2011:ROB

[DKF^{+}11] Ximing Du, Jaspal Kumar, Charles Ferguson, Timothy A. Schulz, Yan Shan Ong, Wanjin Hong, William A. Prinz, Robert G. Parton, Andrew J. Brown, and Hongyuan Yang. A role for oxysterol-binding protein–related protein 5 in endosomal cholesterol trafficking. Journal of Cell Biology, 192 (1):121–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/1/121.

Davalos:2013:PDR

[DKM^{+}13] Albert R. Davalos, Misako Kawahara, Gautam K. Malhotra, Nicholas Schaum, Jiahao Huang, Urvi Ved, Christian M. Beausejour, Jean-Philippe Coppe, Francis Rodier, and Judith Campisi. p53-dependent release of Alarmin HMGB1 is a central mediator of senescent phenotypes. Journal of Cell Biology, 201(4):613–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/4/613.

Dominguez-Kelly:2011:WCG

[DKMK^{+}11] Raquel Domínguez-Kelly, Yusé Martín, Stephane Koundrioukoff, Marvin E. Tanenbaum, Veronique A. J. Smits, René H. Medema, Michelle Debatisse, and Raimundo Freire. Wee1 controls genomic stability during replication by regulating the Mus81-Eme1 endonuclease. Journal of Cell Biology, 194(4): 567–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/4/567.

Drosopoulos:2012:HTR

[DKY^{+}12] William C. Drosopoulos, Settapong T. Kosiyatrakul, Zi Yan, Simone G. Calderano, and Carl L. Schildkraut. Human telomeres replicate using chromosome-specific, rather than universal, replication programs. Journal of Cell Biology, 197(2): 253–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/2/253.

Dunsch:2011:AKS

[DLBG11] Anja K. Dunsch, Emily Linnane, Francis A. Barr, and Ulrike Gruneberg. The astrin–kinastrin/SKAP complex localizes to microtubule plus ends and facilitates chromosome alignment. Journal of Cell Biology, 192(6):959–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/6/959.

Dieni:2012:BPP

[DMD^{+}12] Sandra Dieni, Tomoya Matsumoto, Martijn Dekkers, Stefanie Rauskolb, Mihai S. Ionescu, Ruben Deogracias, Eckart D. Gundelfinger, Masami Kojima, Sigrun Nestel, Michael Frotscher, and Yves-Alain Barde. BDNF and its pro-peptide are stored in presynaptic dense core vesicles in brain neurons. Journal of Cell Biology, 196(6):775–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/196/6/775.

Ditlev:2012:SND

[DMH^{+}12] Jonathon A. Ditlev, Paul J. Michalski, Greg Huber, Gonzalo M. Rivera, William A. Mohler, Leslie M. Loew, and Bruce J. Mayer. Stoichiometry of Nck-dependent actin polymerization in living cells. Journal of Cell Biology, 197(5): 643–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/5/643.

DeAntoni:2012:SMI

[DMK^{+}12] Anna De Antoni, Stefano Maffini, Stefan Knapp, Andrea Musacchio, and Stefano Santaguida. A small-molecule inhibitor of Haspin alters the kinetochore functions of Aurora B. Journal of Cell Biology, 199(2):269–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/2/269.

Simoes:2014:RGS

[dMSMZ14] Sérgio de Matos Simões, Avantika Mainieri, and Jennifer A. Zallen. Rho GTPase and Shroom direct planar polarized actomyosin contractility during convergent extension. Journal of Cell Biology, 204(4):575–??, February 2014. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/4/575.

DeMaria:2010:CBS

[DN10] Shannon DeMaria and John Ngai. The cell biology of smell. Journal of Cell Biology, 191(3):443–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
3/443.

Dekkers:2013:DDN

[DNB13] Martijn P. J. Dekkers, Vassiliki Nikoletopoulou, and Yves-Alain Barde. Death of developing neurons: New insights and implications for connectivity. Journal of Cell Biology, 203(3): 385–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/3/385.

DiFiore:2010:HCD

[DP10] Barbara Di Fiore and Jonathon Pines. How cyclin A destruction escapes the spindle assembly checkpoint. Journal of Cell Biology, 190(4):501–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/4/501.

Day:2010:PRD

[DPB^{+}10] Tovah A. Day, Komariah Palle, Laura R. Barkley, Naoko Kakusho, Ying Zou, Satoshi Tateishi, Alain Verreault, Hisao Masai, and Cyrus Vaziri. Phosphorylated Rad18 directs DNA Polymerase η to sites of stalled replication. Journal of Cell Biology, 191(5):953–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/5/953.

Daniel:2012:LAK

[DPL^{+}12] Jeremy A. Daniel, Manuela Pellegrini, Baeck-Seung Lee, Zhi Guo, Darius Filsuf, Natalya V. Belkina, Zhongsheng You, Tanya T. Paull, Barry P. Sleckman, Lionel Feigenbaum, and André Nussenzweig. Loss of ATM kinase activity leads to embryonic lethality in mice. Journal of Cell Biology, 198(3): 295–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/3/295.

Danielsen:2012:DDI

[DPV^{+}12] Jannie Rendtlew Danielsen, Lou Klitgaard Povlsen, Bine Hare Villumsen, Werner Streicher, Jakob Nilsson, Mats Wikström, Simon Bekker-Jensen, and Niels Mailand. DNA damage– inducible SUMOylation of HERC2 promotes RNF8 binding via a novel SUMO-binding Zinc finger. Journal of Cell Biology, 197(2):179–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/2/179.

Danussi:2011:EII

[DPW^{+}11] Carla Danussi, Alessandra Petrucco, Bruna Wassermann, Eliana Pivetta, Teresa Maria Elisa Modica, Lisa Del Bel Belluz, Alfonso Colombatti, and Paola Spessotto. EMILIN1α4/ α9 integrin interaction inhibits dermal fibroblast and keratinocyte proliferation. Journal of Cell Biology, 195(1): 131–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/1/131.

Danussi:2012:EII

[DPW^{+}12] Carla Danussi, Alessandra Petrucco, Bruna Wassermann, Eliana Pivetta, Teresa Maria Elisa Modica, Lisa Del Bel Belluz, Alfonso Colombatti, and Paola Spessotto. EMILIN1α4/ α9 integrin interaction inhibits dermal fibroblast and keratinocyte proliferation. Journal of Cell Biology, 196(1): 177–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/1/177.

Debattisti:2014:RER

[DPZ^{+}14] Valentina Debattisti, Diana Pendin, Elena Ziviani, Andrea Daga, and Luca Scorrano. Reduction of endoplasmic reticulum stress attenuates the defects caused by Drosophila mitofusin depletion. Journal of Cell Biology, 204(3):303–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
3/303.

DiPetrillo:2010:PCA

[DS10] Christen G. DiPetrillo and Elizabeth F. Smith. Pcdp1 is a central apparatus protein that binds Ca^{2+}-calmodulin and regulates ciliary motility. Journal of Cell Biology, 189(3):601–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
189/3/601.

Davis:2012:GZT

[DS12] Matthew J. Davis and Joseph Schlessinger. The genesis of Zelboraf: Targeting mutant B–Raf in melanoma. Journal of Cell Biology, 199(1):15–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/1/15.

Dong:2014:PRC

[DSB^{+}14] Fenglan Dong, Kyosuke Shinohara, Yanick Botilde, Ryo Nabeshima, Yasuko Asai, Akemi Fukumoto, Toshiaki Hasegawa, Moe Matsuo, Hiroyuki Takeda, Hidetaka Shiratori, Tetsuya Nakamura, and Hiroshi Hamada. Pih1d3 is required for cytoplasmic preassembly of axonemal dynein in mouse sperm. Journal of Cell Biology, 204(2):203–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/2/203.

Dix:2013:LPR

[DSD^{+}13] Carly I. Dix, Harish Chandra Soundararajan, Nikola S. Dzhindzhev, Farida Begum, Beat Suter, Hiroyuki Ohkura, Elaine Stephens, and Simon L. Bullock. Lissencephaly-1 promotes the recruitment of dynein and dynactin to transported mRNAs. Journal of Cell Biology, 202(3):479–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
3/479.

deSaint-Jean:2011:OES

[dSJDD^{+}11] Maud de Saint-Jean, Vanessa Delfosse, Dominique Douguet, Gaëtan Chicanne, Bernard Payrastre, William Bourguet, Bruno Antonny, and Guillaume Drin. Osh4p exchanges sterols for phosphatidylinositol 4-phosphate between lipid bilayers. Journal of Cell Biology, 195(6):965–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
6/965.

Daskalaki:2011:DIM

[DSK^{+}11] Aikaterini Daskalaki, Nevine A. Shalaby, Kristina Kux, Giorgos Tsoumpekos, George D. Tsibidis, Marc A. T. Muskavitch, and Christos Delidakis. Distinct intracellular motifs of Delta mediate its ubiquitylation and activation by Mindbomb1 and Neuralized. Journal of Cell Biology, 195(6):1017–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
6/1017.

Davydenko:2013:ICA

[DSL13] Olga Davydenko, Richard M. Schultz, and Michael A. Lampson. Increased CDK1 activity determines the timing of kinetochore-microtubule attachments in meiosis I. Journal of Cell Biology, 202(2):221–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/202/2/221.

Lopes:2011:TVR

[dSLPRG11] Katharina da Silva Lopes, Agnieszka Pietas, Michael H. Radke, and Michael Gotthardt. Titin visualization in real time reveals an unexpected level of mobility within and between sarcomeres. Journal of Cell Biology, 193(4):785–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
4/785.

Dai:2011:TBA

[DSM^{+}11] Haiming Dai, Alyson Smith, X. Wei Meng, Paula A. Schneider, Yuan-Ping Pang, and Scott H. Kaufmann. Transient binding of an activator BH3 domain to the Bak BH3-binding groove initiates Bak oligomerization. Journal of Cell Biology, 194 (1):39–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/1/39.

Deschenes-Simard:2013:MPC

[DSMB13] Xavier Deschênes-Simard, Yusuke Mizukami, and Nabeel Bardeesy. Macrophages in pancreatic cancer: Starting things off on the wrong track. Journal of Cell Biology, 202(3):403–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
202/3/403.

daSilva:2013:TSR

[dSMSS13] Sara Morais da Silva, Tatiana Moutinho-Santos, and Claudio E. Sunkel. A tumor suppressor role of the Bub3 spindle checkpoint protein after apoptosis inhibition. Journal of Cell Biology, 201(3):385–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/3/385.

Demuro:2011:SCC

[DSP11] Angelo Demuro, Martin Smith, and Ian Parker. Single-channel Ca^{2+} imaging implicates Aβ1-42 amyloid pores in Alzheimer’s disease pathology. Journal of Cell Biology, 195(3):515–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
195/3/515.

Djiane:2011:SDE

[DSW^{+}11] Alexandre Djiane, Hideyuki Shimizu, Marian Wilkin, Sabine Mazleyrat, Martin D. Jennings, Johanna Avis, Sarah Bray, and Martin Baron. Su(dx) E3 ubiquitin ligase-dependent and –independent functions of Polychaetoid, the Drosophila ZO-1 homologue. Journal of Cell Biology, 192(1):189–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
1/189.

Deakin:2014:PIH

[DT14] Nicholas O. Deakin and Christopher E. Turner. Paxillin inhibits HDAC6 to regulate microtubule acetylation, Golgi structure, and polarized migration. Journal of Cell Biology, 206(3):395–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/3/395.

deThe:2012:APL

[dTLLB12] Hugues de Thé, Morgane Le Bras, and Valérie Lallemand-Breitenbach. Acute promyelocytic leukemia, arsenic, and PML bodies. Journal of Cell Biology, 198(1):11–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/198/1/11.

Dundr:2011:SGT

[Dun11] Miroslav Dundr. Seed and grow: a two-step model for nuclear body biogenesis. Journal of Cell Biology, 193(4):605–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
4/605.

DAmbrosio:2010:WGR

[DV10] Michael V. D’Ambrosio and Ronald D. Vale. A whole genome RNAi screen of Drosophila S2 cell spreading performed using automated computational image analysis. Journal of Cell Biology, 191(3):471–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/3/471.

Domnitz:2012:MAM

[DWDW12] Sarah B. Domnitz, Michael Wagenbach, Justin Decarreau, and Linda Wordeman. MCAK activity at microtubule tips regulates spindle microtubule length to promote robust kinetochore attachment. Journal of Cell Biology, 197(2):231–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
2/231.

Dadgar:2014:ARD

[DWJ^{+}14] Sherry Dadgar, Zuyi Wang, Helen Johnston, Akanchha Kesari, Kanneboyina Nagaraju, Yi-Wen Chen, D. Ashley Hill, Terence A. Partridge, Mamta Giri, Robert J. Freishtat, Javad Nazarian, Jianhua Xuan, Yue Wang, and Eric P. Hoffman. Asynchronous remodeling is a driver of failed regeneration in Duchenne muscular dystrophy. Journal of Cell Biology, 207 (1):139–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/1/139.

Dantas:2011:DNE

[DWL^{+}11] Tiago J. Dantas, Yifan Wang, Pierce Lalor, Peter Dockery, and Ciaran G. Morrison. Defective nucleotide excision repair with normal centrosome structures and functions in the absence of all vertebrate centrins. Journal of Cell Biology, 193(2):307–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
193/2/307.

Drawnel:2012:MAB

[DWM^{+}12] Faye M. Drawnel, Dagmar Wachten, Jeffery D. Molkentin, Marjorie Maillet, Jan Magnus Aronsen, Fredrik Swift, Ivar Sjaastad, Ning Liu, Daniele Catalucci, Katsuhiko Mikoshiba, Chihiro Hisatsune, Hanneke Okkenhaug, Simon R. Andrews, Martin D. Bootman, and H. Llewelyn Roderick. Mutual antagonism between IP_{3} RII and miRNA-133a regulates calcium signals and cardiac hypertrophy. Journal of Cell Biology, 199(5): 783–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/5/783.

deCathelineau:2010:FCA

[dWMR10] Aimee deCathelineau, Elizabeth H. Williams, Tom Misteli, and Mike Rossner. Friends, colleagues, authors, lend us your data. Journal of Cell Biology, 191(2):231–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
2/231.

Delorme-Walker:2011:PRF

[DWPC^{+}11] Violaine D. Delorme-Walker, Jeffrey R. Peterson, Jonathan Chernoff, Clare M. Waterman, Gaudenz Danuser, Céline Der-Mardirossian, and Gary M. Bokoch. Pak1 regulates focal adhesion strength, myosin IIA distribution, and actin dynamics to optimize cell migration. Journal of Cell Biology, 193 (7):1289–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/7/1289.

Diaz:2013:NIS

[DYI^{+}13] Begoña Díaz, Angela Yuen, Shinji Iizuka, Shigeki Higashiyama, and Sara A. Courtneidge. Notch increases the shedding of HB–EGF by ADAM12 to potentiate invadopodia formation in hypoxia. Journal of Cell Biology, 201(2):279–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
201/2/279.

Droujinine:2014:SES

[DYP14] Ilia A. Droujinine, Dong Yan, and Norbert Perrimon. A sharp end to sugary Wingless travels. Journal of Cell Biology, 206(7):

819–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/7/819.

Dhara:2014:CSP

[DYS^{+}14] Madhurima Dhara, Antonio Yarzagaray, Yvonne Schwarz, Soumyajit Dutta, Chad Grabner, Paanteha K. Moghadam, Anneka Bost, Claudia Schirra, Jens Rettig, Kerstin Reim, Nils Brose, Ralf Mohrmann, and Dieter Bruns. Complexin synchronizes primed vesicle exocytosis and regulates fusion pore dynamics. Journal of Cell Biology, 204(7):1123–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
7/1123.

Doyle:2011:ALM

[DZT^{+}11] Michelle J. Doyle, Sheng Zhou, Kathleen Kelly Tanaka, Addolorata Pisconti, Nicholas H. Farina, Brian P. Sorrentino, and Bradley B. Olwin. Abcg2 labels multiple cell types in skeletal muscle and participates in muscle regeneration. Journal of Cell Biology, 195(1):147–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/1/147.

Elsing:2014:EHD

[EAB^{+}14] Alexandra N. Elsing, Camilla Aspelin, Johanna K. Björk, Heidi A. Bergman, Samu V. Himanen, Marko J. Kallio, Pia Roos-Mattjus, and Lea Sistonen. Expression of HSF2 decreases in mitosis to enable stress-inducible transcription and cell survival. Journal of Cell Biology, 206(6):735–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
206/6/735.

Espenel:2013:BLK

[EAK13] Cedric Espenel, Bipul R. Acharya, and Geri Kreitzer. A biosensor of local kinesin activity reveals roles of PKC and EB1 in KIF17 activation. Journal of Cell Biology, 203(3): 445–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/3/445.

Ellis:2013:NVL

[EBB13] Kathryn Ellis, Jennifer Bagwell, and Michel Bagnat. Notochord vacuoles are lysosome-related organelles that function in axis and spine morphogenesis. Journal of Cell Biology, 200 (5):667–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/200/5/667.

Emond:2011:CPC

[EBBJ11] Michelle R. Emond, Sayantanee Biswas, Cheasequah J. Blevins, and James D. Jontes. A complex of Protocadherin-19 and N-cadherin mediates a novel mechanism of cell adhesion. Journal of Cell Biology, 195(7):1115–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
7/1115.

Elabd:2013:DMD

[ECC^{+}13] Christian Elabd, Wendy Cousin, Robert Y. Chen, Marc S. Chooljian, Joey T. Pham, Irina M. Conboy, and Michael J. Conboy. DNA methyltransferase-3-dependent nonrandom template segregation in differentiating embryonic stem cells. Journal of Cell Biology, 203(1):73–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/1/73.

Esk:2010:CHC

[ECJB10] Christopher Esk, Chih-Ying Chen, Ludger Johannes, and Frances M. Brodsky. The clathrin heavy chain isoform CHC22 functions in a novel endosomal sorting step. Journal of Cell Biology, 188(1):131–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/1/131.

Espeut:2012:MBK

[ECK^{+}12] Julien Espeut, Dhanya K. Cheerambathur, Lenno Krenning, Karen Oegema, and Arshad Desai. Microtubule binding by KNL-1 contributes to spindle checkpoint silencing at the kinetochore. Journal of Cell Biology, 196(4):469–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
4/469.

Estey:2010:DRS

[EDF^{+}10] Mathew P. Estey, Caterina Di Ciano-Oliveira, Carol D. Froese, Margaret T. Bejide, and William S. Trimble. Distinct roles of septins in cytokinesis: SEPT9 mediates midbody abscission. Journal of Cell Biology, 191(4):741–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
4/741.

Eshed-Eisenbach:2013:ASA

[EEP13] Yael Eshed-Eisenbach and Elior Peles. Axonal spectrins: All-purpose fences. Journal of Cell Biology, 203(3):381– ??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/3/381.

Elting:2014:FSM

[EHUD14] Mary Williard Elting, Christina L. Hueschen, Dylan B. Udy, and Sophie Dumont. Force on spindle microtubule minus ends moves chromosomes. Journal of Cell Biology, 206(2):245–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
206/2/245.

Ensminger:2014:DBC

[EIE^{+}14] Michael Ensminger, Lucie Iloff, Christian Ebel, Teodora Nikolova, Bernd Kaina, and Markus Löbrich. DNA breaks and chromosomal aberrations arise when replication meets base excision repair. Journal of Cell Biology, 206(1):29–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/206/1/29.

Engel:2012:RRI

[EIW^{+}12] Benjamin D. Engel, Hiroaki Ishikawa, Kimberly A. Wemmer, Stefan Geimer, Ken ichi Wakabayashi, Masafumi Hirono, Branch Craige, Gregory J. Pazour, George B. Witman, Ritsu Kamiya, and Wallace F. Marshall. The role of retrograde intraflagellar transport in flagellar assembly, maintenance, and function. Journal of Cell Biology, 199(1):151–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
1/151.

Eves:2012:OCB

[EJBW12] P. Taylor Eves, Yui Jin, Matthew Brunner, and Lois S. Weisman. Overlap of cargo binding sites on myosin V coordinates the inheritance of diverse cargoes. Journal of Cell Biology, 198(1):69–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/1/69.

ElAmine:2013:OAS

[EKJH13] Nour El Amine, Amel Kechad, Silvana Jananji, and Gilles R. X. Hickson. Opposing actions of septins and Sticky on Anillin promote the transition from contractile to midbody ring. Journal of Cell Biology, 203(3):487–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
3/487.

Edens:2014:CRI

[EL14] Lisa J. Edens and Daniel L. Levy. cPKC regulates interphase nuclear size during Xenopus development. Journal of Cell Biology, 206(4):473–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/4/473.

Eisner:2014:MFF

[ELH14] Verónica Eisner, Guy Lenaers, and György Hajnóczky. Mitochondrial fusion is frequent in skeletal muscle and supports excitation–contraction coupling. Journal of Cell Biology, 205 (2):179–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/2/179.

Ellefson:2011:CIM

[EM11] Marina L. Ellefson and Francis J. McNally. CDK-1 inhibits meiotic spindle shortening and dynein-dependent spindle rotation in C. elegans. Journal of Cell Biology, 193(7):1229–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
193/7/1229.

Edgerton-Morgan:2012:TPK

[EMO12] Heather Edgerton-Morgan and Berl R. Oakley. γ-tubulin plays a key role in inactivating APC/C^{Cdh1} at the G_{1}–S boundary. Journal of Cell Biology, 198(5):785–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
5/785.

Eom:2014:NBR

[EMT^{+}14] Taesun Eom, Ilham A. Muslimov, Panayiotis Tsokas, Valerio Berardi, Jun Zhong, Todd C. Sacktor, and Henri Tiedge. Neuronal BC RNAs cooperate with eIF4B to mediate activitydependent translational control. Journal of Cell Biology, 207 (2):237–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/2/237.

Erlemann:2012:ETR

[ENG^{+}12] Sarah Erlemann, Annett Neuner, Linda Gombos, Romain Gibeaux, Claude Antony, and Elmar Schiebel. An extended γ-tubulin ring functions as a stable platform in microtubule nucleation. Journal of Cell Biology, 197(1):59–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/197/1/59.

Elliott:2010:CAC

[ER10] Michael R. Elliott and Kodi S. Ravichandran. Clearance of apoptotic cells: implications in health and disease. Journal of Cell Biology, 189(7):1059–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/7/1059.

Etard:2010:LAR

[ERS10] Christelle Etard, Urmas Roostalu, and Uwe Strähle. Lack of Apobec2-related proteins causes a dystrophic muscle phenotype in zebrafish embryos. Journal of Cell Biology, 189 (3):527–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/3/527.

Erle:2014:CBA

[ES14] David J. Erle and Dean Sheppard. The cell biology of asthma. Journal of Cell Biology, 205(5):621–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/205/5/621.

Enjolras:2012:DCR

[ETC^{+}12] Camille Enjolras, Joëlle Thomas, Brigitte Chhin, Elisabeth Cortier, Jean-Luc Duteyrat, Fabien Soulavie, Maurice J. Kernan, Anne Laurençon, and Bénédicte Durand. Drosophila chibby is required for basal body formation and ciliogenesis but not for Wg signaling. Journal of Cell Biology, 197(2):313–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
197/2/313.

Enokido:2010:MHI

[ETI^{+}10] Yasushi Enokido, Takuya Tamura, Hikaru Ito, Anup Arumughan, Akihiko Komuro, Hiroki Shiwaku, Masaki Sone, Raphaele Foulle, Hirohide Sawada, Hiroshi Ishiguro, Tetsuya Ono, Miho Murata, Ichiro Kanazawa, Nikolai Tomilin, Kazuhiko Tagawa, Erich E. Wanker, and Hitoshi Okazawa. Mutant huntingtin impairs Ku70-mediated DNA repair. Journal of Cell Biology, 189(3):425–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/3/425.

Egan:2012:LIF

[ETRP12] Martin J. Egan, Kaeling Tan, and Samara L. Reck-Peterson. Lis1 is an initiation factor for dynein-driven organelle transport. Journal of Cell Biology, 197(7):971–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/197/7/
971.

Elhanany-Tamir:2012:OPM

[ETYS^{+}12] Hadas Elhanany-Tamir, Yanxun V. Yu, Miri Shnayder, Ankit Jain, Michael Welte, and Talila Volk. Organelle positioning in muscles requires cooperation between two KASH proteins and microtubules. Journal of Cell Biology, 198(5):833–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
198/5/833.

Espert:2014:PBO

[EUB^{+}14] Antonio Espert, Pelin Uluocak, Ricardo Nunes Bastos, Davinderpreet Mangat, Philipp Graab, and Ulrike Gruneberg. PP2A-b56 opposes Mps1 phosphorylation of Knl1 and thereby promotes spindle assembly checkpoint silencing. Journal of Cell Biology, 206(7):833–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/7/833.

Ewers:2011:SPC

[Ewe11] Helge Ewers. Septin pairs, a complex choreography. Journal of Cell Biology, 193(6):959–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/6/959.

Elbediwy:2012:EJF

[EZT^{+}12] Ahmed Elbediwy, Ceniz Zihni, Stephen J. Terry, Peter Clark, Karl Matter, and Maria S. Balda. Epithelial junction formation requires confinement of Cdc42 activity by a novel SH3BP1 complex. Journal of Cell Biology, 198(4):677–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
198/4/677.

Florentin:2012:CLE

[FA12] Anat Florentin and Eli Arama. Caspase levels and execution efficiencies determine the apoptotic potential of the cell. Journal of Cell Biology, 196(4):513–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/196/4/513.

Frosi:2010:TTM

[FAB^{+}10] Yuri Frosi, Sergio Anastasi, Costanza Ballarò, Giulia Varsano, Loriana Castellani, Elena Maspero, Simona Polo, Stefano Alemà, and Oreste Segatto. A two-tiered mechanism of EGFR inhibition by RALT/MIG6 via kinase suppression and receptor degradation. Journal of Cell Biology, 189(3):557–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
3/557.

Faas:2012:VNG

[FAvdB^{+}12] Frank G. A. Faas, M. Cristina Avramut, Bernard M. van den Berg, A. Mieke Mommaas, Abraham J. Koster, and Raimond B. G. Ravelli. Virtual nanoscopy: Generation of ultra-large high resolution electron microscopy maps. Journal of Cell Biology, 198(3):457–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/3/457.

Fernandez-Busnadiego:2013:CET

[FBAO^{+}13] Rubén Fernández-Busnadiego, Shoh Asano, Ana-Maria Oprisoreanu, Eri Sakata, Michael Doengi, Zdravko Kochovski, Magdalena Zürner, Valentin Stein, Susanne Schoch, Wolfgang Baumeister, and Vladan Lučić. Cryo–electron tomography reveals a critical role of RIM1α in synaptic vesicle tethering. Journal of Cell Biology, 201(5):725–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/5/725.

Fricker:2010:MBD

[FBR^{+}10] Nicolai Fricker, Joel Beaudouin, Petra Richter, Roland Eils, Peter H. Krammer, and Inna N. Lavrik. Model-based dissection of CD95 signaling dynamics reveals both a pro- and antiapoptotic role of c–FLIP_{L}. Journal of Cell Biology, 190 (3):377–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/3/377.

Fernandez-Busnadiego:2010:QAN

[FBZM^{+}10] Rubén Fernández-Busnadiego, Benoît Zuber, Ulrike Elisabeth Maurer, Marek Cyrklaff, Wolfgang Baumeister, and Vladan Lučić. Quantitative analysis of the native presynaptic cytomatrix by cryoelectron tomography. Journal of Cell Biology, 188 (1):145–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/1/145.

Flannery:2010:PDA

[FCA10] Andrew R. Flannery, Cecilia Czibener, and Norma W. Andrews. Palmitoylation-dependent association with CD63 targets the Ca^{2+} sensor synaptotagmin VII to lysosomes. Journal of Cell Biology, 191(3):599–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/3/599.

Foraker:2012:CPC

[FCE^{+}12] Amy B. Foraker, Stéphane M. Camus, Timothy M. Evans, Sophia R. Majeed, Chih-Ying Chen, Sabrina B. Taner, Ivan R. Corrêa, Stephen J. Doxsey, and Frances M. Brodsky. Clathrin promotes centrosome integrity in early mitosis through stabilization of centrosomal ch–TOG. Journal of Cell Biology, 198 (4):591–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/4/591.

Fukata:2013:LPC

[FDB^{+}13] Yuko Fukata, Ariane Dimitrov, Gaelle Boncompain, Ole Vielemeyer, Franck Perez, and Masaki Fukata. Local palmitoylation cycles define activity-regulated postsynaptic subdomains. Journal of Cell Biology, 202(1):145–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/1/145.

Fan:2012:MFG

[FEHF12] Yi Fan, Sandeepa M. Eswarappa, Masahiro Hitomi, and Paul L. Fox. Myo1c facilitates G-actin transport to the leading edge of migrating endothelial cells. Journal of Cell Biology, 198(1):47–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/1/47.

Fox:2010:CCL

[FHA10] Rebecca M. Fox, Caitlin D. Hanlon, and Deborah J. Andrew. The CrebA/Creb3-like transcription factors are major and direct regulators of secretory capacity. Journal of Cell Biology, 191(3):479–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/3/479.

Fontana:2012:CJS

[FHD^{+}12] Xavier Fontana, Mariya Hristova, Clive Da Costa, Smriti Patodia, Laura Thei, Milan Makwana, Bradley Spencer-Dene, Morwena Latouche, Rhona Mirsky, Kristjan R. Jessen, Rüdiger Klein, Gennadij Raivich, and Axel Behrens. c-jun in Schwann cells promotes axonal regeneration and motoneuron survival via paracrine signaling. Journal of Cell Biology, 198(1):127–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
198/1/127.

Fortsch:2011:MRM

[FHKW11] Johannes Förtsch, Eric Hummel, Melanie Krist, and Benedikt Westermann. The myosin-related motor protein Myo2 is an essential mediator of bud-directed mitochondrial movement in yeast. Journal of Cell Biology, 194(3):473–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
3/473.

Flannagan:2010:DMP

[FHY^{+}10] Ronald S. Flannagan, Rene E. Harrison, Christopher M. Yip, Khuloud Jaqaman, and Sergio Grinstein. Dynamic macrophage “probing” is required for the efficient capture of phagocytic targets. Journal of Cell Biology, 191(6):1205– ??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/6/1205.

Finkel:2011:STR

[Fin11] Toren Finkel. Signal transduction by reactive oxygen species. Journal of Cell Biology, 194(1):7–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/1/7.

Fischer:2014:MYM

[Fis14] Robert S. Fischer. Move your microvilli. Journal of Cell Biology, 207(1):9–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/1/9.

Fang:2010:BTC

[FLN^{+}10] Xiaodong Fang, Jianying Luo, Ryuichi Nishihama, Carsten Wloka, Christopher Dravis, Mirko Travaglia, Masayuki Iwase, Elizabeth A. Vallen, and Erfei Bi. Biphasic targeting and cleavage furrow ingression directed by the tail of a myosin II. Journal of Cell Biology, 191(7):1333–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/191/7/

1333. See correction [FLN^{+}16].

Fang:2016:CBT

[FLN^{+}16] Xiaodong Fang, Jianying Luo, Ryuichi Nishihama, Carsten Wloka, Christopher Dravis, Mirko Travaglia, Masayuki Iwase, Elizabeth A. Vallen, and Erfei Bi. Correction: Biphasic targeting and cleavage furrow ingression directed by the tail of a myosin II. Journal of Cell Biology, 215(4):591–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/

4/591. See [FLN^{+}10].

Follit:2010:CTF

[FLVP10] John A. Follit, Lixia Li, Yvonne Vucica, and Gregory J. Pazour. The cytoplasmic tail of fibrocystin contains a ciliary targeting sequence. Journal of Cell Biology, 188(1):21–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
188/1/21.

Fessing:2011:PRS

[FMG^{+}11] Michael Y. Fessing, Andrei N. Mardaryev, Michal R. Gdula, Andrey A. Sharov, Tatyana Y. Sharova, Valentina Rapisarda, Konstantin B. Gordon, Anna D. Smorodchenko, Krzysztof Poterlowicz, Giustina Ferone, Yoshinori Kohwi, Caterina Missero, Terumi Kohwi-Shigematsu, and Vladimir A. Botchkarev. p63 regulates Satb1 to control tissue-specific chromatin remodeling during development of the epidermis. Journal of Cell Biology, 194(6):825–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/6/825.

Fujita:2013:RAM

[FMI^{+}13] Naonobu Fujita, Eiji Morita, Takashi Itoh, Atsushi Tanaka, Megumi Nakaoka, Yuki Osada, Tetsuo Umemoto, Tatsuya Saitoh, Hitoshi Nakatogawa, Shouhei Kobayashi, Tokuko Haraguchi, Jun-Lin Guan, Kazuhiro Iwai, Fuminori Tokunaga, Kazunobu Saito, Koutaro Ishibashi, Shizuo Akira, Mitsunori Fukuda, Takeshi Noda, and Tamotsu Yoshimori. Recruitment of the autophagic machinery to endosomes during infection is mediated by ubiquitin. Journal of Cell Biology, 203(1): 115–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/1/115.

Fernandez-Martinez:2012:SFM

[FMPS^{+}12] Javier Fernandez-Martinez, Jeremy Phillips, Matthew D. Sekedat, Ruben Diaz-Avalos, Javier Velazquez-Muriel, Josef D.

Franke, Rosemary Williams, David L. Stokes, Brian T. Chait, Andrej Sali, and Michael P. Rout. Structure–function mapping of a heptameric module in the nuclear pore complex. Journal of Cell Biology, 196(4):419–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/4/419.

Fazzio:2010:CCR

[FP10] Thomas G. Fazzio and Barbara Panning. Condensin complexes regulate mitotic progression and interphase chromatin structure in embryonic stem cells. Journal of Cell Biology, 188(4): 491–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/4/491.

Ferreira:2013:ABS

[FPAM13] Jorge G. Ferreira, António J. Pereira, Anna Akhmanova, and Helder Maiato. Aurora B spatially regulates EB3 phosphorylation to coordinate daughter cell adhesion with cytokinesis. Journal of Cell Biology, 201(5):709–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/5/709.

Frittoli:2014:RRR

[FPM^{+}14] Emanuela Frittoli, Andrea Palamidessi, Paola Marighetti, Stefano Confalonieri, Fabrizio Bianchi, Chiara Malinverno, Giovanni Mazzarol, Giuseppe Viale, Ines Martin-Padura, Massimilliano Garré, Dario Parazzoli, Valentina Mattei, Salvatore Cortellino, Giovanni Bertalot, Pier Paolo Di Fiore, and Giorgio Scita. A RAB5/RAB4 recycling circuitry induces a proteolytic invasive program and promotes tumor dissemination. Journal of Cell Biology, 206(2):307–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/206/2/307.

Faurobert:2013:CIC

[FRL^{+}13] Eva Faurobert, Claire Rome, Justyna Lisowska, Sandra Manet-Dupé, Gwénola Boulday, Marilyne Malbouyres, Martial Balland, Anne-Pascale Bouin, Michelle Kéramidas, Daniel Bouvard, Jean-Luc Coll, Florence Ruggiero, Elisabeth Tournier-Lasserve, and Corinne Albiges-Rizo. CCM1–ICAP-1 complex controls β1 integrin-dependent endothelial contractility and fibronectin remodeling. Journal of Cell Biology, 202 (3):545–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/3/545.

Franz:2013:CEN

[FRS^{+}13] Anna Franz, Hélio Roque, Saroj Saurya, Jeroen Dobbelaere, and Jordan W. Raff. CP110 exhibits novel regulatory activities during centriole assembly in Drosophila. Journal of Cell Biology, 203(5):785–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/5/785.

Fridolfsson:2010:KDN

[FS10] Heidi N. Fridolfsson and Daniel A. Starr. Kinesin-1 and dynein at the nuclear envelope mediate the bidirectional migrations of nuclei. Journal of Cell Biology, 191(1):115–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
1/115.

Fernandes:2014:NRP

[FS14] Isabelle Fernandes and Frieder Schöck. The nebulin repeat protein Lasp regulates I-band architecture and filament spacing in myofibrils. Journal of Cell Biology, 206(4):559–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
4/559.

Fournier:2010:FTM

[FSA^{+}10a] Maxime F. Fournier, Roger Sauser, Davide Ambrosi, Jean-Jacques Meister, and Alexander B. Verkhovsky. Force transmission in migrating cells. Journal of Cell Biology, 188(2): 287–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/2/287.

Fourniol:2010:TFP

[FSA^{+}10b] Franck J. Fourniol, Charles V. Sindelar, Béatrice Amigues, Daniel K. Clare, Geraint Thomas, Mylène Perderiset, Fiona Francis, Anne Houdusse, and Carolyn A. Moores. Templatefree 13-protofilament microtubule–MAP assembly visualized at 8 å resolution. Journal of Cell Biology, 191(3):463– ??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/3/463.

Fairn:2011:HRM

[FSA^{+}11] Gregory D. Fairn, Nicole L. Schieber, Nicholas Ariotti, Samantha Murphy, Lars Kuerschner, Richard I. Webb, Sergio Grinstein, and Robert G. Parton. High-resolution mapping reveals topologically distinct cellular pools of phosphatidylserine. Journal of Cell Biology, 194(2):257–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/2/257.

Finzsch:2010:SRS

[FSK^{+}10] Markus Finzsch, Silke Schreiner, Tatjana Kichko, Peter Reeh, Ernst R. Tamm, Michael R. Bösl, Dies Meijer, and Michael Wegner. Sox10 is required for Schwann cell identity and progression beyond the immature Schwann cell stage. Journal of Cell Biology, 189(4):701–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/4/701.

Francis:2011:HSR

[FSLM11] Stephen S. Francis, Jeff Sfakianos, Bryan Lo, and Ira Mellman. A hierarchy of signals regulates entry of membrane proteins into the ciliary membrane domain in epithelial cells. Journal of Cell Biology, 193(1):219–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/1/219.

Funk:2014:TDS

[FSOL14] Caroline Funk, Verena Schmeiser, Jennifer Ortiz, and Johannes Lechner. A TOGL domain specifically targets yeast CLASP to kinetochores to stabilize kinetochore microtubules. Journal of Cell Biology, 205(4):555–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/205/4/555.

Field:2011:BBE

[FSR11] Mark C. Field, Andrej Sali, and Michael P. Rout. On a bender — BARs, ESCRTs, COPs, and finally getting your coat. Journal of Cell Biology, 193(6):963–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/6/963.

Ferreira:2013:SPP

[FTJG13] Helder C. Ferreira, Benjamin D. Towbin, Thibaud Jegou, and Susan M. Gasser. The shelterin protein POT-1 anchors Caenorhabditis elegans telomeres through SUN-1 at the nuclear periphery. Journal of Cell Biology, 203(5):727–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
203/5/727.

Fernandes:2014:RSV

[FUK^{+}14] Ana Clara Fernandes, Valerie Uytterhoeven, Sabine Kuenen, Yu-Chun Wang, Jan R. Slabbaert, Jef Swerts, Jaroslaw Kasprowicz, Stein Aerts, and Patrik Verstreken. Reduced synaptic vesicle protein degradation at lysosomes curbs TBC1D24/sky-induced neurodegeneration. Journal of Cell Biology, 207(4):453–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/4/453.

Friedl:2010:PCM

[FW10] Peter Friedl and Katarina Wolf. Plasticity of cell migration: a multiscale tuning model. Journal of Cell Biology, 188(1): 11–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/1/11.

Fujita:2011:SGA

[FWJ^{+}11] Morihisa Fujita, Reika Watanabe, Nina Jaensch, Maria Romanova-Michaelides, Tadashi Satoh, Masaki Kato, Howard Riezman, Yoshiki Yamaguchi, Yusuke Maeda, and Taroh Kinoshita. Sorting of GPI-anchored proteins into ER exit sites by p24 proteins is dependent on remodeled GPI. Journal of Cell Biology, 194(1):61–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/1/61.

Farhan:2010:MSE

[FWM^{+}10a] Hesso Farhan, Markus W. Wendeler, Sandra Mitrovic, Eugenio Fava, Yael Silberberg, Roded Sharan, Marino Zerial, and Hans-Peter Hauri. MAPK signaling to the early secretory pathway revealed by kinase/phosphatase functional screening. Journal of Cell Biology, 189(6):997–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/6/997.

Friedman:2010:SDM

[FWM^{+}10b] Jonathan R. Friedman, Brant M. Webster, David N. Mastronarde, Kristen J. Verhey, and Gia K. Voeltz. ER sliding dynamics and ER–mitochondrial contacts occur on acetylated microtubules. Journal of Cell Biology, 190(3):363–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
3/363.

Ge:2010:CIR

[GB10] Xin Quan Ge and J. Julian Blow. Chk1 inhibits replication factory activation but allows dormant origin firing in existing factories. Journal of Cell Biology, 191(7):1285–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
7/1285.

Garbett:2012:PIR

[GB12] Damien Garbett and Anthony Bretscher. PDZ interactions regulate rapid turnover of the scaffolding protein EBP50 in microvilli. Journal of Cell Biology, 198(2):195–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
2/195.

Gerondopoulos:2014:RRG

[GBiY^{+}14] Andreas Gerondopoulos, Ricardo Nunes Bastos, Shin ichiro Yoshimura, Rachel Anderson, Sarah Carpanini, Irene Aligianis, Mark T. Handley, and Francis A. Barr. Rab18 and a Rab18 GEF complex are required for normal ER structure. Journal of Cell Biology, 205(5):707–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/205/5/707.

Giunta:2010:DDS

[GBJ10] Simona Giunta, Rimma Belotserkovskaya, and Stephen P. Jackson. DNA damage signaling in response to double-strand breaks during mitosis. Journal of Cell Biology, 190(2):197–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
190/2/197.

Galati:2014:DDS

[GBK^{+}14] Domenico F. Galati, Stephanie Bonney, Zev Kronenberg, Christina Clarissa, Mark Yandell, Nels C. Elde, Maria Jerka-Dziadosz, Thomas H. Giddings, Joseph Frankel, and Chad G. Pearson. DisAp-dependent striated fiber elongation is required to organize ciliary arrays. Journal of Cell Biology, 207(6): 705–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/6/705.

Garcia:2011:SDM

[GBL^{+}11] Galo Garcia, Aurelie Bertin, Zhu Li, Yi Song, Michael A. McMurray, Jeremy Thorner, and Eva Nogales. Subunitdependent modulation of septin assembly: Budding yeast septin Shs1 promotes ring and gauze formation. Journal of Cell Biology, 195(6):993–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/6/993.

Gervasi:2012:JCI

[GBSC^{+}12] Megan Gervasi, Anna Bianchi-Smiraglia, Michael Cummings, Qiao Zheng, Dan Wang, Song Liu, and Andrei V. Bakin. JunB contributes to Id2 repression and the epithelial–mesenchymal transition in response to transforming growth factor–β. Journal of Cell Biology, 196(5):589–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/196/5/589.

Gascoigne:2013:CDP

[GC13] Karen E. Gascoigne and Iain M. Cheeseman. CDK-dependent phosphorylation and nuclear exclusion coordinately control kinetochore assembly state. Journal of Cell Biology, 201 (1):23–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/1/23.

Gordon:2012:PTI

[GCC12] Leslie B. Gordon, Kan Cao, and Francis S. Collins. Progeria: Translational insights from cell biology. Journal of Cell Biology, 199(1):9–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/1/9.

Grassart:2014:ADD

[GCH^{+}14] Alexandre Grassart, Aaron T. Cheng, Sun Hae Hong, Fan Zhang, Nathan Zenzer, Yongmei Feng, David M. Briner, Gregory D. Davis, Dmitry Malkov, and David G. Drubin. Actin and dynamin2 dynamics and interplay during clathrinmediated endocytosis. Journal of Cell Biology, 205(5):721–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
205/5/721.

Gallaud:2014:EMP

[GCP^{+}14] Emmanuel Gallaud, Renaud Caous, Aude Pascal, Franck Bazile, Jean-Philippe Gagné, Sébastien Huet, Guy G. Poirier, Denis Chrétien, Laurent Richard-Parpaillon, and Régis Giet. Ensconsin/ map7 promotes microtubule growth and centrosome separation in Drosophila neural stem cells. Journal of Cell Biology, 204(7):1111–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/7/1111.

Gay:2012:SMK

[GCR^{+}12] Guillaume Gay, Thibault Courtheoux, Céline Reyes, Sylvie Tournier, and Yannick Gachet. A stochastic model of kinetochore–microtubule attachment accurately describes fission yeast chromosome segregation. Journal of Cell Biology, 196(6):757–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/6/757.

Grau:2013:TGG

[GCR^{+}13] Montserrat Bosch Grau, Gloria Gonzalez Curto, Cecilia Rocha, Maria M. Magiera, Patricia Marques Sousa, Tiziana Giordano, Nathalie Spassky, and Carsten Janke. Tubulin glycylases and glutamylases have distinct functions in stabilization and motility of ependymal cilia. Journal of Cell Biology, 202(3):441–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/3/441.

Gill:2010:RGT

[GCSB10] David J. Gill, Joanne Chia, Jamie Senewiratne, and Frederic Bard. Regulation of O-glycosylation through Golgi–to–ER relocation of initiation enzymes. Journal of Cell Biology, 189 (5):843–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/5/843.

Guimaraes:2011:IHC

[GCV^{+}11] Carla P. Guimaraes, Jan E. Carette, Malini Varadarajan, John Antos, Maximilian W. Popp, Eric Spooner, Thijn R. Brummelkamp, and Hidde L. Ploegh. Identification of host cell factors required for intoxication through use of modified cholera toxin. Journal of Cell Biology, 195(5):751–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
195/5/751.

Gaudin:2012:CRK

[GdAJ^{+}12] Raphaël Gaudin, Bruna Cunha de Alencar, Mabel Jouve, Stefano Bèrre, Emmanuel Le Bouder, Michael Schindler, Aditi Varthaman, François-Xavier Gobert, and Philippe Benaroch. Critical role for the kinesin KIF3A in the HIV life cycle in primary human macrophages. Journal of Cell Biology, 199(3): 467–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/3/467.

Gagliardi:2014:PMA

[GdBP^{+}14] Paolo Armando Gagliardi, Laura di Blasio, Alberto Puliafito, Giorgio Seano, Roberto Sessa, Federica Chianale, Thomas Leung, Federico Bussolino, and Luca Primo. PDK1-mediated activation of MRCKα regulates directional cell migration and lamellipodia retraction. Journal of Cell Biology, 206(3): 415–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/3/415.

Guo:2013:IIB

[GDO13] Zheng Guo, Ian Driver, and Benjamin Ohlstein. Injuryinduced BMP signaling negatively regulates Drosophila midgut homeostasis. Journal of Cell Biology, 201(6):945–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
201/6/945.

Gupta:2012:VMC

[GDS^{+}12] Anjali Gupta, Dennis R. Diener, Priyanka Sivadas, Joel L. Rosenbaum, and Pinfen Yang. The versatile molecular complex component LC8 promotes several distinct steps of flagellar assembly. Journal of Cell Biology, 198(1):115–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
1/115.

Gault:2014:OSM

[GEN14] William J. Gault, Balázs Enyedi, and Philipp Niethammer. Osmotic surveillance mediates rapid wound closure through nucleotide release. Journal of Cell Biology, 207(6):767–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
207/6/767.

Gokhin:2011:CAT

[GF11] David S. Gokhin and Velia M. Fowler. Cytoplasmic γ-actin and tropomodulin isoforms link to the sarcoplasmic reticulum in skeletal muscle fibers. Journal of Cell Biology, 194(1):105–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
194/1/105.

Gu:2011:ECE

[GFSR11] Yapeng Gu, Tetyana Forostyan, Roger Sabbadini, and Jody Rosenblatt. Epithelial cell extrusion requires the sphingosine-1-phosphate receptor 2 pathway. Journal of Cell Biology, 193 (4):667–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/4/667.

Garcia-Gonzalo:2012:SBP

[GGR12] Francesc R. Garcia-Gonzalo and Jeremy F. Reiter. Scoring a backstage pass: Mechanisms of ciliogenesis and ciliary access. Journal of Cell Biology, 197(6):697–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/6/697.

Grotsky:2013:BLA

[GGSN^{+}13] David A. Grotsky, Ignacio Gonzalez-Suarez, Anna Novell, Martin A. Neumann, Sree C. Yaddanapudi, Monica Croke, Montserrat Martinez-Alonso, Abena B. Redwood, Sylvia Ortega-Martinez, Zhihui Feng, Enrique Lerma, Teresa Ramon y Cajal, Junran Zhang, Xavier Matias-Guiu, Adriana Dusso, and Susana Gonzalo. BRCA1 loss activates cathepsin L-mediated degradation of 53BP1 in breast cancer cells. Journal of Cell Biology, 200(2):187–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/2/187.

Gamblin:2014:BAB

[GHC^{+}14] Clémence L. Gamblin, Émilie J.-L. Hardy, François J.-M. Chartier, Nicolas Bisson, and Patrick Laprise. A bidirectional antagonism between aPKC and Yurt regulates epithelial cell polarity. Journal of Cell Biology, 204(4):487–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
4/487.

Goehring:2011:PPD

[GHGH11] Nathan W. Goehring, Carsten Hoege, Stephan W. Grill, and Anthony A. Hyman. PAR proteins diffuse freely across the anterior–posterior boundary in polarized C. elegans embryos. Journal of Cell Biology, 193(3):583–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/3/583.

Geng:2010:RMU

[GHK10a] Liyi Geng, Catherine J. Huntoon, and Larry M. Karnitz. RAD18-mediated ubiquitination of PCNA activates the Fanconi anemia DNA repair network. Journal of Cell Biology, 191 (2):249–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/2/249.

Goh:2010:MMC

[GHK^{+}10b] Lai Kuan Goh, Fangtian Huang, Woong Kim, Steven Gygi, and Alexander Sorkin. Multiple mechanisms collectively regulate clathrin-mediated endocytosis of the epidermal growth factor receptor. Journal of Cell Biology, 189(5):871–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
5/871.

Gilbert:2010:CFT

[Gil10] David M. Gilbert. Cell fate transitions and the replication timing decision point. Journal of Cell Biology, 191(5):899– ??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/5/899.

Graziano:2013:LIA

[GJP^{+}13] Brian R. Graziano, Erin M. Jonasson, Jessica G. Pullen, Christopher J. Gould, and Bruce L. Goode. Ligand-induced activation of a formin–NPF pair leads to collaborative actin nucleation. Journal of Cell Biology, 201(4):595–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
4/595.

Godlee:2013:UBI

[GK13] Camilla Godlee and Marko Kaksonen. From uncertain beginnings: Initiation mechanisms of clathrin-mediated endocytosis. Journal of Cell Biology, 203(5):717–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
5/717.

Guo:2012:CDB

[GKA^{+}12] Yige Guo, Christine Kim, Sana Ahmad, Jiayin Zhang, and Yinghui Mao. CENP-E-dependent BubR1 autophosphorylation enhances chromosome alignment and the mitotic checkpoint. Journal of Cell Biology, 198(2):205–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/198/2/
205.

Gruschke:2011:CCIa

[GKR^{+}11a] Steffi Gruschke, Kirsten Kehrein, Katharina Römpler, Kerstin Gröne, Lars Israel, Axel Imhof, Johannes M. Herrmann, and Martin Ott. Cbp3–Cbp6 interacts with the yeast mitochondrial ribosomal tunnel exit and promotes cytochrome b synthesis and assembly. Journal of Cell Biology, 193(6):1101–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
193/6/1101.

Gruschke:2011:CCIb

[GKR^{+}11b] Steffi Gruschke, Kirsten Kehrein, Katharina Römpler, Kerstin Gröne, Lars Israel, Axel Imhof, Johannes M. Herrmann, and Martin Ott. Cbp3–Cbp6 interacts with the yeast mitochondrial ribosomal tunnel exit and promotes cytochrome b synthesis and assembly. Journal of Cell Biology, 194(1):155–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
194/1/155.

Gauthier-Kemper:2011:FDM

[GKWG^{+}11] Anne Gauthier-Kemper, Carina Weissmann, Nataliya Golovyashkina, Zsofia Sebö-Lemke, Gerard Drewes, Volker Gerke, Jürgen J. Heinisch, and Roland Brandt. The frontotemporal dementia mutation R406W blocks tau’s interaction with the membrane in an annexin A2-dependent manner. Journal of Cell Biology, 192(4):647–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/4/647.

Groehler:2010:CBK

[GL10] Angela L. Groehler and Deborah A. Lannigan. A chromatinbound kinase, ERK8, protects genomic integrity by inhibiting HDM2-mediated degradation of the DNA clamp PCNA. Journal of Cell Biology, 190(4):575–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/4/575.

Garbett:2010:SPE

[GLB10] Damien Garbett, David P. LaLonde, and Anthony Bretscher. The scaffolding protein EBP50 regulates microvillar assembly in a phosphorylation-dependent manner. Journal of Cell Biology, 191(2):397–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/2/397.

Gingras:2012:SBJ

[GLG12] Alexandre R. Gingras, Jian J. Liu, and Mark H. Ginsberg. Structural basis of the junctional anchorage of the cerebral cavernous malformations complex. Journal of Cell Biology, 199 (1):39–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/1/39.

Gokhin:2010:TIR

[GLM^{+}10] David S. Gokhin, Raymond A. Lewis, Caroline R. McKeown, Roberta B. Nowak, Nancy E. Kim, Ryan S. Littlefield, Richard L. Lieber, and Velia M. Fowler. Tropomodulin isoforms regulate thin filament pointed-end capping and skeletal muscle physiology. Journal of Cell Biology, 189(1):95–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
1/95.

Grove:2011:CBR

[GM11] Joe Grove and Mark Marsh. The cell biology of receptormediated virus entry. Journal of Cell Biology, 195(7):1071– ??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/7/1071.

Gatlin:2010:DPM

[GMD^{+}10] Jesse C. Gatlin, Alexandre Matov, Gaudenz Danuser, Timothy J. Mitchison, and Edward D. Salmon. Directly probing the mechanical properties of the spindle and its matrix. Journal of Cell Biology, 188(4):481–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/4/481.

Green:2013:MRS

[GMW^{+}13] Rebecca A. Green, Jonathan R. Mayers, Shaohe Wang, Lindsay Lewellyn, Arshad Desai, Anjon Audhya, and Karen Oegema. The midbody ring scaffolds the abscission machinery in the absence of midbody microtubules. Journal of Cell Biology, 203(3):505–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/3/505.

Gu:2011:ITR

[GNHB11] Zhizhan Gu, Erika H. Noss, Victor W. Hsu, and Michael B. Brenner. Integrins traffic rapidly via circular dorsal ruffles and macropinocytosis during stimulated cell migration. Journal of Cell Biology, 193(1):61–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/1/61.

Goldberg:2012:DPI

[Gol12a] Alfred L. Goldberg. Development of proteasome inhibitors as research tools and cancer drugs. Journal of Cell Biology, 199(4):583–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/4/583.

Goldstein:2012:NFH

[Gol12b] Lawrence S. B. Goldstein. New frontiers in human cell biology and medicine: Can pluripotent stem cells deliver? Journal of Cell Biology, 199(4):577–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/4/577.

Gault:2012:DCG

[GOWM12] William J. Gault, Patricio Olguin, Ursula Weber, and Marek Mlodzik. Drosophila CK1-γ, gilgamesh, controls PCPmediated morphogenesis through regulation of vesicle trafficking. Journal of Cell Biology, 196(5):605–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/196/5/
605.

Gavet:2010:ACB

[GP10] Olivier Gavet and Jonathon Pines. Activation of cyclin B1– Cdk1 synchronizes events in the nucleus and the cytoplasm at mitosis. Journal of Cell Biology, 189(2):247–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
2/247.

Ganem:2012:LAM

[GP12] Neil J. Ganem and David Pellman. Linking abnormal mitosis to the acquisition of DNA damage. Journal of Cell Biology, 199(6):871–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/6/871.

Gilden:2012:SCF

[GPCK12] Julia K. Gilden, Sebastian Peck, Yi-Chun M. Chen, and Matthew F. Krummel. The septin cytoskeleton facilitates membrane retraction during motility and blebbing. Journal of Cell Biology, 196(1):103–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/1/103.

Greer:2011:CKD

[GR11] Yoshimi Endo Greer and Jeffrey S. Rubin. Casein kinase 1 delta functions at the centrosome to mediate Wnt-3a-dependent neurite outgrowth. Journal of Cell Biology, 192 (6):993–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/6/993.

Gruschke:2012:CCC

[GRH^{+}12] Steffi Gruschke, Katharina Römpler, Markus Hildenbeutel, Kirsten Kehrein, Inge Kühl, Nathalie Bonnefoy, and Martin Ott. The Cbp3–Cbp6 complex coordinates cytochrome b synthesis with bc_{1} complex assembly in yeast mitochondria. Journal of Cell Biology, 199(1):137–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/1/137.

Gupton:2012:MBI

[GRHA^{+}12] Stephanie L. Gupton, Daisy Riquelme, Shannon K. Hughes-Alford, Jenny Tadros, Shireen S. Rudina, Richard O. Hynes, Douglas Lauffenburger, and Frank B. Gertler. Mena binds α5 integrin directly and modulates α5β1 function. Journal of Cell Biology, 198(4):657–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/4/657.

Gusnowski:2011:VDD

[GS11] Eva M. Gusnowski and Martin Srayko. Visualization of dyneindependent microtubule gliding at the cell cortex: implications for spindle positioning. Journal of Cell Biology, 194(3):377–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
194/3/377.

Galvez:2013:CRD

[GSB^{+}13] Beatriz G. Galvez, Maurilio Sampaolesi, Silvia Brunelli, Diego Covarello, Manuela Gavina, Barbara Rossi, Gabriela Constantin, Yvan Torrente, and Giulio Cossu. Complete repair of dystrophic skeletal muscle by mesoangioblasts with enhanced migration ability. Journal of Cell Biology, 202(4):715–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
202/4/715.

Ge:2011:IIR

[GSC11] Yejing Ge, Yuting Sun, and Jie Chen. IGF-II is regulated by microRNA-125b in skeletal myogenesis. Journal of Cell Biology, 192(1):69–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/1/69.

Gohler:2011:AMP

[GSGL11] Thomas Göhler, Simone Sabbioneda, Catherine M. Green, and Alan R. Lehmann. ATR-mediated phosphorylation of DNA polymerase η is needed for efficient recovery from UV damage. Journal of Cell Biology, 192(2):219–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/2/219.

Geymonat:2010:PLC

[GSJS10] Marco Geymonat, Adonis Spanos, Sanne Jensen, and Steven G. Sedgwick. Phosphorylation of Lte1 by Cdk prevents polarized growth during mitotic arrest in S. cerevisiae. Journal of Cell Biology, 191(6):1097–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/6/1097.

Gebert:2012:MPC

[GSM^{+}12] Michael Gebert, Sandra G. Schrempp, Carola S. Mehnert, Anna K. Heißwolf, Silke Oeljeklaus, Raffaele Ieva, Maria Bohnert, Karina von der Malsburg, Sebastian Wiese, Thomas Kleinschroth, Carola Hunte, Helmut E. Meyer, Ilka Haferkamp, Bernard Guiard, Bettina Warscheid, Nikolaus Pfanner, and Martin van der Laan. Mgr2 promotes coupling of the mitochondrial presequence translocase to partner complexes. Journal of Cell Biology, 197(5):595–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/5/595.

Gao:2014:CRM

[GSM^{+}14] Jie Gao, Désirée Schatton, Paola Martinelli, Henriette Hansen, David Pla-Martin, Esther Barth, Christian Becker, Janine Altmueller, Peter Frommolt, Marco Sardiello, and Elena I. Rugarli. CLUH regulates mitochondrial biogenesis by binding mRNAs of nuclear-encoded mitochondrial proteins. Journal of Cell Biology, 207(2):213–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/2/213.

Germann:2014:TDB

[GSP^{+}14] Susanne M. Germann, Vera Schramke, Rune Troelsgaard Pedersen, Irene Gallina, Nadine Eckert-Boulet, Vibe H. Oestergaard, and Michael Lisby. TopBP1/ dpb11 binds DNA anaphase bridges to prevent genome instability. Journal of Cell Biology, 204(1):45–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/1/45.

Gardner:2011:TSP

[GSS^{+}11] Jennifer M. Gardner, Christine J. Smoyer, Elizabeth S. Stensrud, Richard Alexander, Madelaine Gogol, Winfried Wiegraebe, and Sue L. Jaspersen. Targeting of the SUN protein Mps3 to the inner nuclear membrane by the histone variant H2A.Z. Journal of Cell Biology, 193(3):489–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
3/489.

Gerl:2012:QAL

[GSU^{+}12] Mathias J. Gerl, Julio L. Sampaio, Severino Urban, Lucie Kalvodova, Jean-Marc Verbavatz, Beth Binnington, Dirk Lindemann, Clifford A. Lingwood, Andrej Shevchenko, Cornelia Schroeder, and Kai Simons. Quantitative analysis of the lipidomes of the influenza virus envelope and MDCK cell apical membrane. Journal of Cell Biology, 196(2):213–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
2/213.

Gong:2011:FPL

[GSW^{+}11] Jingyi Gong, Zhiqi Sun, Lizhen Wu, Wenyi Xu, Nicole Schieber, Dijin Xu, Guanghou Shui, Hongyuan Yang, Robert G. Parton, and Peng Li. Fsp27 promotes lipid droplet growth by lipid exchange and transfer at lipid droplet contact sites. Journal of Cell Biology, 195(6):953–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
6/953.

Gomes:2013:MSK

[GTR^{+}13] José-Eduardo Gomes, Nicolas Tavernier, Bénédicte Richaudeau, Etienne Formstecher, Thomas Boulin, Paul E. Mains, Julien Dumont, and Lionel Pintard. Microtubule severing by the katanin complex is activated by PPFR-1-dependent MEI-1 dephosphorylation. Journal of Cell Biology, 202(3):431–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
202/3/431.

Griffiths:2010:ISF

[GTS10] Gillian M. Griffiths, Andy Tsun, and Jane C. Stinchcombe. The immunological synapse: a focal point for endocytosis and exocytosis. Journal of Cell Biology, 189(3):399–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
3/399.

Gerbe:2011:DAN

[GvEM^{+}11] François Gerbe, Johan H. van Es, Leila Makrini, Bénédicte Brulin, Georg Mellitzer, Sylvie Robine, Béatrice Romagnolo, Noah F. Shroyer, Jean-François Bourgaux, Christine Pignodel, Hans Clevers, and Philippe Jay. Distinct ATOH1 and Neurog3 requirements define tuft cells as a new secretory cell type in the intestinal epithelium. Journal of Cell Biology, 192(5):767–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
192/5/767.

Gloerich:2011:NRT

[GVP^{+}11] Martijn Gloerich, Marjolein J. Vliem, Esther Prummel, Lars A. T. Meijer, Marije G. A. Rensen, Holger Rehmann, and Johannes L. Bos. The nucleoporin RanBP2 tethers the cAMP effector Epac1 and inhibits its catalytic activity. Journal of Cell Biology, 193(6):1009–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/6/1009.

Ghosh:2011:DID

[GWP^{+}11] Arundhati Sengupta Ghosh, Bei Wang, Christine D. Pozniak, Mark Chen, Ryan J. Watts, and Joseph W. Lewcock. DLK induces developmental neuronal degeneration via selective regulation of proapoptotic JNK activity. Journal of Cell Biology, 194(5):751–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/5/751.

Gillissen:2010:EBI

[GWR^{+}10] Bernhard Gillissen, Jana Wendt, Antje Richter, Anja Richter, Annika Müer, Tim Overkamp, Nina Gebhardt, Robert Preissner, Claus Belka, Bernd Dörken, and Peter T. Daniel. Endogenous Bak inhibitors Mcl-1 and Bcl-x_{L}: differential impact on TRAIL resistance in Bax-deficient carcinoma. Journal of Cell Biology, 188(6):851–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/6/851.

Gomez:2012:TCE

[GWR12] Juan Manuel Gomez, Ying Wang, and Veit Riechmann. Tao controls epithelial morphogenesis by promoting Fasciclin 2 endocytosis. Journal of Cell Biology, 199(7):1131–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
7/1131.

Gaspar:2014:KET

[GYC^{+}14] Imre Gaspar, Yanxun V. Yu, Sean L. Cotton, Dae-Hwan Kim, Anne Ephrussi, and Michael A. Welte. Klar ensures thermal robustness of oskar localization by restraining RNP motility. Journal of Cell Biology, 206(2):199–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/2/199.

Giagtzoglou:2012:DCE

[GYZ^{+}12] Nikolaos Giagtzoglou, Shinya Yamamoto, Diana Zitserman, Hillary K. Graves, Karen L. Schulze, Hao Wang, Hayley Klein, Fabrice Roegiers, and Hugo J. Bellen. dEHBP1 controls exocytosis and recycling of Delta during asymmetric divisions. Journal of Cell Biology, 196(1):65–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/196/1/65.

Gudi:2011:CTI

[GZLG11] Radhika Gudi, Chaozhong Zou, Jun Li, and Qingshen Gao. Centrobin–tubulin interaction is required for centriole elongation and stability. Journal of Cell Biology, 193(4):711–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
4/711.

Gaggioli:2014:CPSa

[GZR^{+}14a] Vincent Gaggioli, Eva Zeiser, David Rivers, Charles R. Bradshaw, Julie Ahringer, and Philip Zegerman. CDK phosphorylation of SLD-2 is required for replication initiation and germline development in C. elegans. Journal of Cell Biology, 204(4):507–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/4/507.

Gaggioli:2014:CPSb

[GZR^{+}14b] Vincent Gaggioli, Eva Zeiser, David Rivers, Charles R. Bradshaw, Julie Ahringer, and Philip Zegerman. CDK phosphorylation of SLD-2 is required for replication initiation and germline development in C. elegans. Journal of Cell Biology, 204(6):1075–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/6/1075.

Gerhardt:2014:CAD

[GZZ^{+}14] Jeannine Gerhardt, Nikica Zaninovic, Qiansheng Zhan, Advaitha Madireddy, Sarah L. Nolin, Nicole Ersalesi, Zi Yan, Zev Rosenwaks, and Carl L. Schildkraut. Cis-acting DNA sequence at a replication origin promotes repeat expansion to fragile X full mutation. Journal of Cell Biology, 206(5):599–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
206/5/599.

Houtkooper:2012:ETS

[HA12] Riekelt H. Houtkooper and Johan Auwerx. Exploring the therapeutic space around NAD^{+}. Journal of Cell Biology, 199(2): 205–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/2/205.

He:2014:AGP

[HAB14] Meng He, Khadar M. Abdi, and Vann Bennett. Ankyrin-g palmitoylation and βII-spectrin binding to phosphoinositide lipids drive lateral membrane assembly. Journal of Cell Biology, 206(2):273–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/2/273.

Huang:2011:PCM

[HAKK11] Bill X. Huang, Mohammed Akbar, Karl Kevala, and Hee-Yong Kim. Phosphatidylserine is a critical modulator for Akt activation. Journal of Cell Biology, 192(6):979–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
6/979.

Hall:2014:AAB

[Hal14] Alan Hall. Always aiming for the best: For scientists and for science. Journal of Cell Biology, 207(1):7–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/207/1/7.

Higuchi:2014:EEM

[HARS14] Yujiro Higuchi, Peter Ashwin, Yvonne Roger, and Gero Steinberg. Early endosome motility spatially organizes polysome distribution. Journal of Cell Biology, 204(3):343–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
3/343.

Hubner:2010:QPC

[HBC^{+}10] Nina C. Hubner, Alexander W. Bird, Jürgen Cox, Bianca Splettstoesser, Peter Bandilla, Ina Poser, Anthony Hyman, and Matthias Mann. Quantitative proteomics combined with BAC TransgeneOmics reveals in vivo protein interactions. Journal of Cell Biology, 189(4):739–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/4/739.

Harrigan:2011:RSI

[HBC^{+}11] Jeanine A. Harrigan, Rimma Belotserkovskaya, Julia Coates, Daniela S. Dimitrova, Sophie E. Polo, Charles R. Bradshaw, Peter Fraser, and Stephen P. Jackson. Replication stress induces 53BP1-containing OPT domains in G1 cells. Journal of Cell Biology, 193(1):97–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/1/97.

Hoffmeister:2011:PTD

[HBG^{+}11] Helen Hoffmeister, Karin Babinger, Sonja Gürster, Anna Cedzich, Christine Meese, Karin Schadendorf, Larissa Osten, Uwe de Vries, Anne Rascle, and Ralph Witzgall. Polycystin-2 takes different routes to the somatic and ciliary plasma membrane. Journal of Cell Biology, 192(4):631–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
4/631.

Haas:2010:HPL

[HBI^{+}10] Gabrielle Haas, Joerg E. Braun, Cátia Igreja, Felix Tritschler, Tadashi Nishihara, and Elisa Izaurralde. HPat provides a link between deadenylation and decapping in metazoa. Journal of Cell Biology, 189(2):289–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/2/289.

Habbig:2011:NCA

[HBM^{+}11] Sandra Habbig, Malte P. Bartram, Roman U. Müller, Ricarda Schwarz, Nikolaos Andriopoulos, Shuhua Chen, Josef G. Sägmüller, Martin Hoehne, Volker Burst, Max C. Liebau, H. Christian Reinhardt, Thomas Benzing, and Bernhard Schermer. NPHP4, a cilia-associated protein, negatively regulates the Hippo pathway. Journal of Cell Biology, 193 (4):633–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/4/633.

Hammond:2010:AKM

[HBS^{+}10] Jennetta W. Hammond, T. Lynne Blasius, Virupakshi Soppina, Dawen Cai, and Kristen J. Verhey. Autoinhibition of the kinesin-2 motor KIF17 via dual intramolecular mechanisms. Journal of Cell Biology, 189(6):1013–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/6/1013.

Hayakawa:2012:UNP

[HBSD12] Akira Hayakawa, Anna Babour, Lucie Sengmanivong, and Catherine Dargemont. Ubiquitylation of the nuclear pore complex controls nuclear migration during mitosis in S. cerevisiae. Journal of Cell Biology, 196(1):19–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/196/1/19.

Hudson:2010:DKF

[HC10] Andrew M. Hudson and Lynn Cooley. Drosophila Kelch functions with Cullin-3 to organize the ring canal actin cytoskeleton. Journal of Cell Biology, 188(1):29–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/188/1/29.

Heasman:2010:CRS

[HCC^{+}10] Sarah J. Heasman, Leo M. Carlin, Susan Cox, Tony Ng, and Anne J. Ridley. Coordinated RhoA signaling at the leading edge and uropod is required for T cell transendothelial migration. Journal of Cell Biology, 190(4):553–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
4/553.

Hoppins:2011:MFG

[HCCS^{+}11] Suzanne Hoppins, Sean R. Collins, Ann Cassidy-Stone, Eric Hummel, Rachel M. DeVay, Laura L. Lackner, Benedikt Westermann, Maya Schuldiner, Jonathan S. Weissman, and Jodi Nunnari. A mitochondrial-focused genetic interaction map reveals a scaffold-like complex required for inner membrane organization in mitochondria. Journal of Cell Biology, 195(2): 323–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/2/323.

Hurtado:2011:DGR

[HCG^{+}11] Lidia Hurtado, Cristina Caballero, Maria P. Gavilan, Jesus Cardenas, Michel Bornens, and Rosa M. Rios. Disconnecting the Golgi ribbon from the centrosome prevents directional cell migration and ciliogenesis. Journal of Cell Biology, 193 (5):917–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/5/917.

Hung:2013:DSM

[HCP^{+}13] Wei-Chien Hung, Shih-Hsun Chen, Colin D. Paul, Kimberly M. Stroka, Ying-Chun Lo, Joy T. Yang, and Konstantinos Konstantopoulos. Distinct signaling mechanisms regulate migration in unconfined versus confined spaces. Journal of Cell Biology, 202(5):807–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/202/5/807.

Herkert:2010:ATS

[HDH^{+}10] Barbara Herkert, Anne Dwertmann, Steffi Herold, Mona Abed, Jean-Francois Naud, Florian Finkernagel, Gregory S. Harms, Amir Orian, Michael Wanzel, and Martin Eilers. The Arf tumor suppressor protein inhibits Miz1 to suppress cell adhesion and induce apoptosis. Journal of Cell Biology, 188(6):905–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
188/6/905.

Halova:2013:PTA

[HDK^{+}13] Lenka Hálová, Wei Du, Sara Kirkham, Duncan L. Smith, and Janni Petersen. Phosphorylation of the TOR ATP binding domain by AGC kinase constitutes a novel mode of TOR inhibition. Journal of Cell Biology, 203(4):595–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
4/595.

Heuck:2010:SMG

[HFB^{+}10] Alexander Heuck, Ingrid Fetka, Daniel N. Brewer, Daniela Hüls, Mary Munson, Ralf-Peter Jansen, and Dierk Niessing. The structure of the Myo4p globular tail and its function in ASH1 mRNA localization. Journal of Cell Biology, 189 (3):497–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/3/497.

Hamel:2010:NLA

[HFS10] Sophie Hamel, Jacques Fantini, and François Schweisguth. Notch ligand activity is modulated by glycosphingolipid membrane composition in Drosophila melanogaster. Journal of Cell Biology, 188(4):581–??, February 2010. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/188/4/581.

Heller:2014:FDP

[HGV^{+}14] Bradley A. Heller, Monica Ghidinelli, Jakob Voelkl, Steven Einheber, Ryan Smith, Ethan Grund, Grant Morahan, David Chandler, Luba Kalaydjieva, Filippo Giancotti, Rosalind H. King, Aniko Naray Fejes-Toth, Gerard Fejes-Toth, Maria Laura Feltri, Florian Lang, and James L. Salzer. Functionally distinct PI 3-kinase pathways regulate myelination in the peripheral nervous system. Journal of Cell Biology, 204 (7):1219–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/7/1219.

Hotulainen:2010:ADS

[HH10] Pirta Hotulainen and Casper C. Hoogenraad. Actin in dendritic spines: connecting dynamics to function. Journal of Cell Biology, 189(4):619–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/4/619.

Hatch:2014:BNE

[HH14a] Emily Hatch and Martin Hetzer. Breaching the nuclear envelope in development and disease. Journal of Cell Biology, 205 (2):133–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/2/133.

Helmke:2014:TLM

[HH14b] Kara J. Helmke and Rebecca Heald. TPX2 levels modulate meiotic spindle size and architecture in Xenopus egg extracts. Journal of Cell Biology, 206(3):385–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/3/385.

Hwang:2011:CFM

[HHC^{+}11] Hyun Sub Hwang, Sang Gil Hwang, Jun-Ho Cho, Ji Soo Chae, Kyoung Wan Yoon, Ssang-Goo Cho, and Eui-Ju Choi. CIIA functions as a molecular switch for the Rac1-specific GEF activity of SOS1. Journal of Cell Biology, 195(3):377–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
3/377.

Hamada:2011:RDD

[HHJ^{+}11] Masakazu Hamada, Anna Haeger, Karthik B. Jeganathan, Janine H. van Ree, Liviu Malureanu, Sarah Wälde, Jomon Joseph, Ralph H. Kehlenbach, and Jan M. van Deursen. Randependent docking of importin-β to RanBP2/Nup358 filaments is essential for protein import and cell viability. Journal of Cell Biology, 194(4):597–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/4/597.

Hartig:2011:HLS

[HHL^{+}11] Sean M. Hartig, Bin He, Weiwen Long, Benjamin M. Buehrer, and Michael A. Mancini. Homeostatic levels of SRC-2 and SRC-3 promote early human adipogenesis. Journal of Cell Biology, 192(1):55–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/1/55.

Harding:2013:ELBa

[HHS13a] Clifford V. Harding, John E. Heuser, and Philip D. Stahl. Exosomes: Looking back three decades and into the future. Journal of Cell Biology, 200(4):367–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/4/367.

Harding:2013:ELBb

[HHS13b] Clifford V. Harding, John E. Heuser, and Philip D. Stahl. Exosomes: Looking back three decades and into the future. Journal of Cell Biology, 201(3):485–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/3/485.

Hildenbeutel:2014:AFM

[HHS^{+}14] Markus Hildenbeutel, Eric L. Hegg, Katharina Stephan, Steffi Gruschke, Brigitte Meunier, and Martin Ott. Assembly factors monitor sequential hemylation of cytochrome b to regulate mitochondrial translation. Journal of Cell Biology, 205 (4):511–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/4/511.

Huang:2012:NAF

[HHY^{+}12] Junqi Huang, Yinyi Huang, Haochen Yu, Dhivya Subramanian, Anup Padmanabhan, Rahul Thadani, Yaqiong Tao, Xie Tang, Roland Wedlich-Soldner, and Mohan K. Balasubramanian. Nonmedially assembled F-actin cables incorporate into the actomyosin ring in fission yeast. Journal of Cell Biology, 199(5):831–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/5/831.

Huranova:2010:DIS

[HIB^{+}10] Martina Huranová, Ivan Ivani, Aleš Benda, Ina Poser, Yehuda Brody, Martin Hof, Yaron Shav-Tal, Karla M. Neugebauer, and David Staněk. The differential interaction of snRNPs with pre-mRNA reveals splicing kinetics in living cells. Journal of Cell Biology, 191(1):75–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/1/75.

Hawkins:2010:GAS

[HIM^{+}10] Brian J. Hawkins, Krishna M. Irrinki, Karthik Mallilankaraman, Yu-Chin Lien, Youjun Wang, Cunnigaiper D. Bhanumathy, Ramasamy Subbiah, Michael F. Ritchie, Jonathan Soboloff, Yoshihiro Baba, Tomohiro Kurosaki, Suresh K. Joseph, Donald L. Gill, and Muniswamy Madesh. Sglutathionylation activates STIM1 and alters mitochondrial homeostasis. Journal of Cell Biology, 190(3):391–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
3/391.

Huang:2014:TPC

[HJ14] Xi Huang and Lily Yeh Jan. Targeting potassium channels in cancer. Journal of Cell Biology, 206(2):151–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
2/151.

Huang:2014:CHH

[HK14] Jianhua Huang and Daniel Kalderon. Coupling of Hedgehog and Hippo pathways promotes stem cell maintenance by stimulating proliferation. Journal of Cell Biology, 205(3):325–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
205/3/325.

Hatch:2010:CIP

[HKH^{+}10] Emily M. Hatch, Anita Kulukian, Andrew J. Holland, Don W. Cleveland, and Tim Stearns. Cep152 interacts with Plk4 and is required for centriole duplication. Journal of Cell Biology, 191(4):721–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/4/721.

Hayase:2013:WPM

[HKI^{+}13] Junya Hayase, Sachiko Kamakura, Yuko Iwakiri, Yoshihiro Yamaguchi, Tomoko Izaki, Takashi Ito, and Hideki Sumimoto. The WD40 protein Morg1 facilitates Par6–aPKC binding to Crb3 for apical identity in epithelial cells. Journal of Cell Biology, 200(5):635–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/5/635.

Hotta:2010:LBC

[HKN^{+}10] Azusa Hotta, Tomomi Kawakatsu, Tomoya Nakatani, Toshitaka Sato, Chiyuki Matsui, Taiko Sukezane, Tsuyoshi Akagi, Tomoko Hamaji, Ilya Grigoriev, Anna Akhmanova, Yoshimi Takai, and Yuko Mimori-Kiyosue. Laminin-based cell adhesion anchors microtubule plus ends to the epithelial cell basal cortex through LL5α /β. Journal of Cell Biology, 189(5):901–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
189/5/901.

Harris:2011:CSL

[HKN^{+}11] Dinari A. Harris, Kevin Kim, Kenji Nakahara, Constanza Vásquez-Doorman, and Richard W. Carthew. Cargo sorting to lysosome-related organelles regulates siRNA-mediated gene silencing. Journal of Cell Biology, 194(1):77–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/194/1/77.

Hagedorn:2014:ACP

[HKN^{+}14] Elliott J. Hagedorn, Laura C. Kelley, Kaleb M. Naegeli, Zheng Wang, Qiuyi Chi, and David R. Sherwood. ADF/ cofilin promotes invadopodial membrane recycling during cell invasion in vivo. Journal of Cell Biology, 204(7):1209–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
7/1209.

Howes:2010:CIC

[HKR^{+}10] Mark T. Howes, Matthew Kirkham, James Riches, Katia Cortese, Piers J. Walser, Fiona Simpson, Michelle M. Hill, Alun Jones, Richard Lundmark, Margaret R. Lindsay, Delia J. Hernandez-Deviez, Gordana Hadzic, Adam McCluskey, Rumasia Bashir, Libin Liu, Paul Pilch, Harvey McMahon, Phillip J. Robinson, John F. Hancock, Satyajit Mayor, and Robert G. Parton. Clathrin-independent carriers form a high capacity endocytic sorting system at the leading edge of migrating cells. Journal of Cell Biology, 190(4):675–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/4/675.

Holembowski:2014:TEG

[HKR^{+}14] Lena Holembowski, Daniela Kramer, Dietmar Riedel, Raffaella Sordella, Alice Nemajerova, Matthias Dobbelstein, and Ute M. Moll. TAp73 is essential for germ cell adhesion and maturation in testis. Journal of Cell Biology, 204(7):1173–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
7/1173.

Horn:2013:MKD

[HKW^{+}13] Henning F. Horn, Dae In Kim, Graham D. Wright, Esther Sook Miin Wong, Colin L. Stewart, Brian Burke, and Kyle J. Roux. A mammalian KASH domain protein coupling meiotic chromosomes to the cytoskeleton. Journal of Cell Biology, 202(7):1023–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/7/1023.

Harvey:2011:CMC

[HL11] Pamela A. Harvey and Leslie A. Leinwand. Cellular mechanisms of cardiomyopathy. Journal of Cell Biology, 194(3): 355–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/3/355.

Han:2014:SDE

[HLH^{+}14] Jinhua Han, Ting Liu, Michael S. Y. Huen, Lin Hu, Zhiqiu Chen, and Jun Huang. SIVA1 directs the E3 ubiquitin ligase RAD18 for PCNA monoubiquitination. Journal of Cell Biology, 205(6):811–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/6/811.

Howitt:2012:NRN

[HLL^{+}12] Jason Howitt, Jenny Lackovic, Ley-Hian Low, Adam Naguib, Alison Macintyre, Choo-Peng Goh, Jennifer K. Callaway, Vicki Hammond, Tim Thomas, Matthew Dixon, Ulrich Putz, John Silke, Perry Bartlett, Baoli Yang, Sharad Kumar, Lloyd C. Trotman, and Seong-Seng Tan. Ndfip1 regulates nuclear Pten import in vivo to promote neuronal survival following cerebral ischemia. Journal of Cell Biology, 196(1):29–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/196/1/29.

Holland:2010:PLK

[HLN^{+}10] Andrew J. Holland, Weijie Lan, Sherry Niessen, Heather Hoover, and Don W. Cleveland. Polo-like kinase 4 kinase activity limits centrosome overduplication by autoregulating its own stability. Journal of Cell Biology, 188(2):191–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
2/191.

Heo:2011:PMD

[HLN^{+}11] Kyung-Sun Heo, Hakjoo Lee, Patrizia Nigro, Tamlyn Thomas, Nhat-Tu Le, Eugene Chang, Carolyn McClain, Cynthia A. Reinhart-King, Michael R. King, Bradford C. Berk, Keigi Fujiwara, Chang-Hoon Woo, and Jun ichi Abe. PKCζ mediates disturbed flow-induced endothelial apoptosis via p53 SUMOylation. Journal of Cell Biology, 193(5):867–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
5/867.

Huang:2014:ASA

[HLS^{+}14] Haidong Huang, Yujing Li, Keith E. Szulwach, Guoqiang Zhang, Peng Jin, and Dahua Chen. AGO3 Slicer activity regulates mitochondria–nuage localization of Armitage and piRNA amplification. Journal of Cell Biology, 206(2):217–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
2/217.

Hsu:2012:NPT

[HLT12] Joseph K. Hsu, Tao Lin, and Robert Y. L. Tsai. Nucleostemin prevents telomere damage by promoting PML–IV recruitment to SUMOylated TRF1. Journal of Cell Biology, 197(5):613–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
197/5/613.

Hansen:2010:VPA

[HM10] Scott D. Hansen and R. Dyche Mullins. VASP is a processive actin polymerase that requires monomeric actin for barbed end association. Journal of Cell Biology, 191(3):571–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
3/571.

Hammond:2014:NPP

[HMB14] Gerald R. V. Hammond, Matthias P. Machner, and Tamas Balla. A novel probe for phosphatidylinositol 4-phosphate reveals multiple pools beyond the Golgi. Journal of Cell Biology, 205(1):113–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/1/113.

Holloway:2010:MBH

[HMBC10] J. Kim Holloway, Meisha A. Morelli, Peter L. Borst, and Paula E. Cohen. Mammalian BLM helicase is critical for integrating multiple pathways of meiotic recombination. Journal of Cell Biology, 188(6):779–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/6/779.

Hsu:2010:RES

[HMiY^{+}10] Chieh Hsu, Yuichi Morohashi, Shin ichiro Yoshimura, Natalia Manrique-Hoyos, SangYong Jung, Marcel A. Lauterbach, Mostafa Bakhti, Mads Grønborg, Wiebke Möbius, JeongSeop Rhee, Francis A. Barr, and Mikael Simons. Regulation of exosome secretion by Rab35 and its GTPase-activating proteins TBC1D10A–C. Journal of Cell Biology, 189(2):223–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
2/223.

Hamilton:2014:HEE

[HMO^{+}14] Jaeger J. Hamilton, Victoria L. Marlow, Richard A. Owen, Marília de Assis Alcoforado Costa, Manman Guo, Grant Buchanan, Govind Chandra, Matthias Trost, Sarah J. Coulthurst, Tracy Palmer, Nicola R. Stanley-Wall, and Frank Sargent. A holin and an endopeptidase are essential for chitinolytic protein secretion in Serratia marcescens. Journal of Cell Biology, 207(5):615–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/207/5/615.

Huveneers:2012:VAE

[HOS^{+}12] Stephan Huveneers, Joppe Oldenburg, Emma Spanjaard, Gerard van der Krogt, Ilya Grigoriev, Anna Akhmanova, Holger Rehmann, and Johan de Rooij. Vinculin associates with endothelial VE-cadherin junctions to control force-dependent remodeling. Journal of Cell Biology, 196(5):641–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
5/641.

Hu:2010:HGA

[HPB10] Yan Hu, Matt Plutz, and Andrew S. Belmont. Hsp70 gene association with nuclear speckles is Hsp70 promoter specific. Journal of Cell Biology, 191(4):711–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
4/711.

Hipp:2012:IIP

[HPB^{+}12] Mark S. Hipp, Chetan N. Patel, Kirill Bersuker, Brigit E. Riley, Stephen E. Kaiser, Thomas A. Shaler, Michael Brandeis, and Ron R. Kopito. Indirect inhibition of 26S proteasome activity in a cellular model of Huntington’s disease. Journal of Cell Biology, 196(5):573–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/5/573.

Hasegawa:2013:CGP

[HRK13] Keisuke Hasegawa, Sung Jin Ryu, and Petr Kaláb. Chromosomal gain promotes formation of a steep RanGTP gradient that drives mitosis in aneuploid cells. Journal of Cell Biology, 200(2):151–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/2/151.

Huntwork-Rodriguez:2013:JMP

[HRWW^{+}13] Sarah Huntwork-Rodriguez, Bei Wang, Trent Watkins, Arundhati Sengupta Ghosh, Christine D. Pozniak, Daisy Bustos, Kim Newton, Donald S. Kirkpatrick, and Joseph W. Lewcock. JNK-mediated phosphorylation of DLK suppresses its ubiquitination to promote neuronal apoptosis. Journal of Cell Biology, 202(5):747–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/202/5/747.

Hentrich:2010:MOA

[HS10a] Christian Hentrich and Thomas Surrey. Microtubule organization by the antagonistic mitotic motors kinesin-5 and kinesin-14. Journal of Cell Biology, 189(3):465–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/3/465.

Huang:2010:TSR

[HS10b] Hung-Hsiang Huang and Pamela Stanley. A testis-specific regulator of complex and hybrid N-glycan synthesis. Journal of Cell Biology, 190(5):893–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/5/893.

Hantschel:2012:CBK

[HSF12] Oliver Hantschel and Giulio Superti-Furga. Cell biology: a key driver of therapeutic innovation. Journal of Cell Biology, 199(4):571–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/4/571.

Hatan:2011:DBB

[HSI^{+}11] Meital Hatan, Vera Shinder, David Israeli, Frank Schnorrer, and Talila Volk. The Drosophila blood brain barrier is maintained by GPCR-dependent dynamic actin structures. Journal of Cell Biology, 192(2):307–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/2/307.

Harada:2014:NLS

[HSI^{+}14] Takamasa Harada, Joe Swift, Jerome Irianto, Jae-Won Shin, Kyle R. Spinler, Avathamsa Athirasala, Rocky Diegmiller, P. C. Dave P. Dingal, Irena L. Ivanovska, and Dennis E. Discher. Nuclear lamin stiffness is a barrier to 3D migration, but softness can limit survival. Journal of Cell Biology, 204 (5):669–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/5/669.

Hartman:2013:DCD

[HSJ^{+}13] Tiffiney R. Hartman, Todd I. Strochlic, Yingbiao Ji, Daniel Zinshteyn, and Alana M. O’Reilly. Diet controls Drosophila follicle stem cell proliferation via Hedgehog sequestration and release. Journal of Cell Biology, 201(5):741–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
5/741.

Hesse:2010:ZCB

[HSK^{+}10] Eric Hesse, Hiroaki Saito, Riku Kiviranta, Diego Correa, Kei Yamana, Lynn Neff, Daniel Toben, Georg Duda, Azeddine Atfi, Valérie Geoffroy, William C. Horne, and Roland Baron. Zfp521 controls bone mass by HDAC3-dependent attenuation of Runx2 activity. Journal of Cell Biology, 191(7):1271–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
191/7/1271.

Hori:2011:SBE

[HSKAT11] Kazuya Hori, Anindya Sen, Tom Kirchhausen, and Spyros Artavanis-Tsakonas. Synergy between the ESCRT–III complex and Deltex defines a ligand-independent Notch signal. Journal of Cell Biology, 195(6):1005–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
6/1005.

Hegarat:2011:AAB

[HSN^{+}11] Nadia Hégarat, Ewan Smith, Gowri Nayak, Shunichi Takeda, Patrick A. Eyers, and Helfrid Hochegger. Aurora A and Aurora B jointly coordinate chromosome segregation and anaphase microtubule dynamics. Journal of Cell Biology, 195(7):1103– ??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/7/1103.

Hayer:2010:CUT

[HSR^{+}10] Arnold Hayer, Miriam Stoeber, Danilo Ritz, Sabrina Engel, Hemmo H. Meyer, and Ari Helenius. Caveolin-1 is ubiquitinated and targeted to intralumenal vesicles in endolysosomes for degradation. Journal of Cell Biology, 191(3):615– ??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/3/615.

Hou:2013:SNR

[HSS^{+}13] Hailong Hou, Lu Sun, Benjamin A. Siddoway, Ronald S. Petralia, Hongtian Yang, Hua Gu, Angus C. Nairn, and Houhui Xia. Synaptic NMDA receptor stimulation activates PP1 by inhibiting its phosphorylation by Cdk5. Journal of Cell Biology, 203(3):521–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/3/521.

Hori:2013:CRC

[HSTF13] Tetsuya Hori, Wei-Hao Shang, Kozo Takeuchi, and Tatsuo Fukagawa. The CCAN recruits CENP–A to the centromere and forms the structural core for kinetochore assembly. Journal of Cell Biology, 200(1):45–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/1/45.

Holloway:2014:MCC

[HSY^{+}14] J. Kim Holloway, Xianfei Sun, Rayka Yokoo, Anne M. Villeneuve, and Paula E. Cohen. Mammalian CNTD1 is critical for meiotic crossover maturation and deselection of excess precrossover sites. Journal of Cell Biology, 205(5):633–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/205/
5/633.

Hornung:2014:CMD

[HTM^{+}14] Peter Hornung, Paulina Troc, Francesca Malvezzi, Michael Maier, Zuzana Demianova, Tomasz Zimniak, Gabriele Litos, Fabienne Lampert, Alexander Schleiffer, Matthias Brunner, Karl Mechtler, Franz Herzog, Thomas C. Marlovits, and Stefan Westermann. A cooperative mechanism drives budding yeast kinetochore assembly downstream of CENP–A. Journal of Cell Biology, 206(4):509–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/4/509.

Hewitt:2010:SMA

[HTS^{+}10] Laura Hewitt, Anthony Tighe, Stefano Santaguida, Anne M. White, Clifford D. Jones, Andrea Musacchio, Stephen Green, and Stephen S. Taylor. Sustained Mps1 activity is required in mitosis to recruit O–Mad2 to the Mad1–C–Mad2 core complex. Journal of Cell Biology, 190(1):25–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/1/25.

Hayakawa:2011:AFF

[HTS11] Kimihide Hayakawa, Hitoshi Tatsumi, and Masahiro Sokabe. Actin filaments function as a tension sensor by tensiondependent binding of cofilin to the filament. Journal of Cell Biology, 195(5):721–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/5/721.

Hasegawa:2011:SRD

[HTT^{+}11a] Junya Hasegawa, Emi Tokuda, Takeshi Tenno, Kazuya Tsujita, Haruko Sawai, Hidekazu Hiroaki, Tadaomi Takenawa, and Toshiki Itoh. SH3YL1 regulates dorsal ruffle formation by a novel phosphoinositide-binding domain. Journal of Cell Biology, 193(5):901–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/5/901.

Hong:2011:CEJ

[HTT11b] Soonjin Hong, Regina B. Troyanovsky, and Sergey M. Troyanovsky. Cadherin exits the junction by switching its adhesive bond. Journal of Cell Biology, 192(6):1073–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
6/1073.

Hong:2013:BFA

[HTT13] Soonjin Hong, Regina B. Troyanovsky, and Sergey M. Troyanovsky. Binding to F-actin guides cadherin cluster assembly, stability, and movement. Journal of Cell Biology, 201 (1):131–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/1/131.

Huber:2013:ERS

[HVDG13] Michael D. Huber, Paul W. Vesely, Kaustuv Datta, and Larry Gerace. Erlins restrict SREBP activation in the ER and regulate cellular cholesterol homeostasis. Journal of Cell Biology, 203(3):427–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/3/427.

Hayes:2014:MCC

[HVOF^{+}14] Polly Hayes, Vladimir Varga, Sofia Olego-Fernandez, Jack Sunter, Michael L. Ginger, and Keith Gull. Modulation of a cytoskeletal calpain-like protein induces major transitions in trypanosome morphology. Journal of Cell Biology, 206(3): 377–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/3/377.

Hirai:2010:MRA

[HVW^{+}10] Hiroyuki Hirai, Mayank Verma, Shuichi Watanabe, Christopher Tastad, Yoko Asakura, and Atsushi Asakura. MyoD regulates apoptosis of myoblasts through microRNA-mediated down-regulation of Pax3. Journal of Cell Biology, 191(2): 347–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/2/347.

Haglund:2011:PPM

[HW11] Cat M. Haglund and Matthew D. Welch. Pathogens and polymers: Microbe–host interactions illuminate the cytoskeleton. Journal of Cell Biology, 195(1):7–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/1/7.

Hood:2013:CAD

[HWB^{+}13] Fiona E. Hood, Samantha J. Williams, Selena G. Burgess, Mark W. Richards, Daniel Roth, Anne Straube, Mark Pfuhl, Richard Bayliss, and Stephen J. Royle. Coordination of adjacent domains mediates TACC3–ch–TOG–clathrin assembly and mitotic spindle binding. Journal of Cell Biology, 202(3): 463–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/3/463.

Haberman:2012:SVS

[HWE^{+}12] Adam Haberman, W. Ryan Williamson, Daniel Epstein, Dong Wang, Srisha Rina, Ian A. Meinertzhagen, and P. Robin Hiesinger. The synaptic vesicle SNARE neuronal Synaptobrevin promotes endolysosomal degradation and prevents neurodegeneration. Journal of Cell Biology, 196(2):261–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
196/2/261.

Holohan:2014:TES

[HWS14] Brody Holohan, Woodring E. Wright, and Jerry W. Shay. Telomeropathies: an emerging spectrum disorder. Journal of Cell Biology, 205(3):289–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/205/3/289.

Hynes:2012:EME

[Hyn12] Richard O. Hynes. The evolution of metazoan extracellular matrix. Journal of Cell Biology, 196(6):671–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
6/671.

Hanazawa:2011:PPS

[HYS11] Momoyo Hanazawa, Masafumi Yonetani, and Asako Sugimoto. PGL proteins self associate and bind RNPs to mediate germ granule assembly in C. elegans. Journal of Cell Biology, 192 (6):929–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/6/929.

Hiramoto-Yamaki:2010:EEM

[HYTU^{+}10] Nao Hiramoto-Yamaki, Shingo Takeuchi, Shuhei Ueda, Kohei Harada, Satoshi Fujimoto, Manabu Negishi, and Hironori Katoh. Ephexin4 and EphA2 mediate cell migration through a RhoG-dependent mechanism. Journal of Cell Biology, 190 (3):461–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/3/461.

Heym:2013:VRM

[HZE^{+}13] Roland G. Heym, Dennis Zimmermann, Franziska T. Edelmann, Lars Israel, Zeynep Ökten, David R. Kovar, and Dierk Niessing. In vitro reconstitution of an mRNA-transport complex reveals mechanisms of assembly and motor activation. Journal of Cell Biology, 203(6):971–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
6/971.

Hagedorn:2013:NRD

[HZM^{+}13] Elliott J. Hagedorn, Joshua W. Ziel, Meghan A. Morrissey, Lara M. Linden, Zheng Wang, Qiuyi Chi, Sam A. Johnson, and David R. Sherwood. The netrin receptor DCC focuses invadopodia-driven basement membrane transmigration in vivo. Journal of Cell Biology, 201(6):903–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
6/903.

Hartman:2010:DBL

[HZS^{+}10] Tiffiney R. Hartman, Daniel Zinshteyn, Heather K. Schofield, Emmanuelle Nicolas, Ami Okada, and Alana M. O’Reilly. Drosophila Boi limits Hedgehog levels to suppress follicle stem cell proliferation. Journal of Cell Biology, 191(5):943– ??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/5/943.

Huang:2012:TUL

[HZT^{+}12] Nai-Jia Huang, Liguo Zhang, Wanli Tang, Chen Chen, Chih-Sheng Yang, and Sally Kornbluth. The Trim39 ubiquitin ligase inhibits APC/C^{Cdh1}-mediated degradation of the Bax activator MOAP-1. Journal of Cell Biology, 197(3):361–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
3/361.

Hou:2012:CLC

[HZW^{+}12] Haitong Hou, Zhou Zhou, Yu Wang, Jiyong Wang, Scott P. Kallgren, Tatiana Kurchuk, Elizabeth A. Miller, Fred Chang, and Songtao Jia. Csi1 links centromeres to the nuclear envelope for centromere clustering. Journal of Cell Biology, 199(5): 735–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/5/735.

Ismail:2010:BMH

[IAMH10] Ismail Hassan Ismail, Christi Andrin, Darin McDonald, and Michael J. Hendzel. BMI1-mediated histone ubiquitylation promotes DNA double-strand break repair. Journal of Cell Biology, 191(1):45–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/1/45.

Ivanovic:2012:CAP

[IHG^{+}12] Aleksandra Ivanovic, Ido Horresh, Neev Golan, Ivo Spiegel, Helena Sabanay, Shahar Frechter, Shinichi Ohno, Nobuo Terada, Wiebke Möbius, Jack Rosenbluth, Nils Brose, and Elior Peles. The cytoskeletal adapter protein 4.1G organizes the internodes in peripheral myelinated nerves. Journal of Cell Biology, 196(3):337–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/3/337.

Ingerman:2013:ACA

[IHM13] Elena Ingerman, Jennifer Ying Hsiao, and R. Dyche Mullins. Arp2/ 3 complex ATP hydrolysis promotes lamellipodial actin network disassembly but is dispensable for assembly. Journal of Cell Biology, 200(5):619–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/5/619.

Ishikawa:2011:PFC

[IIN^{+}11] Masaki Ishikawa, Tsutomu Iwamoto, Takashi Nakamura, Andrew Doyle, Satoshi Fukumoto, and Yoshihiko Yamada. Pannexin 3 functions as an ER Ca^{2+} channel, hemichannel, and gap junction to promote osteoblast differentiation. Journal of Cell Biology, 193(7):1257–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/7/1257.

Iijima:2014:NCT

[IIWS14] Takatoshi Iijima, Yoko Iijima, Harald Witte, and Peter Scheiffele. Neuronal cell type–specific alternative splicing is regulated by the KH domain protein SLM1. Journal of Cell Biology, 204(3):331–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/3/331.

Itoh:2011:ONA

[IKU^{+}11] Takashi Itoh, Eiko Kanno, Takefumi Uemura, Satoshi Waguri, and Mitsunori Fukuda. OATL1, a novel autophagosomeresident Rab33B–GAP, regulates autophagosomal maturation. Journal of Cell Biology, 192(5):839–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/5/839.

Iskratsch:2010:FFF

[ILD^{+}10] Thomas Iskratsch, Stephan Lange, Joseph Dwyer, Ay Lin Kho, Cris dos Remedios, and Elisabeth Ehler. Formin follows function: a muscle-specific isoform of FHOD3 is regulated by CK2 phosphorylation and promotes myofibril maintenance. Journal of Cell Biology, 191(6):1159–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/6/1159.

Itakura:2011:PTA

[IM11] Eisuke Itakura and Noboru Mizushima. p62 targeting to the autophagosome formation site requires self-oligomerization but not LC3 binding. Journal of Cell Biology, 192(1):17–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
192/1/17.

Inoko:2012:TAB

[IMG^{+}12] Akihito Inoko, Makoto Matsuyama, Hidemasa Goto, Yuki Ohmuro-Matsuyama, Yuko Hayashi, Masato Enomoto, Miho Ibi, Takeshi Urano, Shigenobu Yonemura, Tohru Kiyono, Ichiro Izawa, and Masaki Inagaki. Trichoplein and Aurora A block aberrant primary cilia assembly in proliferating cells. Journal of Cell Biology, 197(3):391–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/3/391.

Iden:2012:APJ

[IMP^{+}12] Sandra Iden, Steve Misselwitz, Swetha S. D. Peddibhotla, Hüseyin Tuncay, Daniela Rehder, Volker Gerke, Horst Robenek, Atsushi Suzuki, and Klaus Ebnet. aPKC phosphorylates JAM–A at Ser285 to promote cell contact maturation and tight junction formation. Journal of Cell Biology, 196 (5):623–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/5/623.

Izawa:2012:MAC

[IP12] Daisuke Izawa and Jonathon Pines. Mad2 and the APC/C compete for the same site on Cdc20 to ensure proper chromosome segregation. Journal of Cell Biology, 199(1):27–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
199/1/27.

Ivanov:2013:LMP

[IPM^{+}13] Andre Ivanov, Jeff Pawlikowski, Indrani Manoharan, John van Tuyn, David M. Nelson, Taranjit Singh Rai, Parisha P. Shah, Graeme Hewitt, Viktor I. Korolchuk, Joao F. Passos, Hong Wu, Shelley L. Berger, and Peter D. Adams. Lysosomemediated processing of chromatin in senescence. Journal of Cell Biology, 202(1):129–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/202/1/129.

Ishikawa:2014:PTR

[Ish14] Takashi Ishikawa. Protein tagging reveals new insights into signaling in flagella. Journal of Cell Biology, 204(5):631–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
204/5/631.

Izu:2011:TXC

[ISZ^{+}11] Yayoi Izu, Mei Sun, Daniela Zwolanek, Guido Veit, Valerie Williams, Byeong Cha, Karl J. Jepsen, Manuel Koch, and David E. Birk. Type XII collagen regulates osteoblast polarity and communication during bone formation. Journal of Cell Biology, 193(6):1115–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/6/1115.

Inami:2011:PAN

[IWS^{+}11] Yoshihiro Inami, Satoshi Waguri, Ayako Sakamoto, Tsuguka Kouno, Kazuto Nakada, Okio Hino, Sumio Watanabe, Jin Ando, Manabu Iwadate, Masayuki Yamamoto, Myung-Shik Lee, Keiji Tanaka, and Masaaki Komatsu. Persistent activation of Nrf2 through p62 in hepatocellular carcinoma cells. Journal of Cell Biology, 193(2):275–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/2/275.

Yoshimura:2010:FWC

[iYGL^{+}10] Shin ichiro Yoshimura, Andreas Gerondopoulos, Andrea Linford, Daniel J. Rigden, and Francis A. Barr. Family-wide characterization of the DENN domain Rab GDP–GTP exchange factors. Journal of Cell Biology, 191(2):367–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
191/2/367.

Jha:2013:SCD

[JAM^{+}13] Archana Jha, Malini Ahuja, József Maléth, Claudia M. Moreno, Joseph P. Yuan, Min Seuk Kim, and Shmuel Muallem. The STIM1 CTID domain determines access of SARAF to SOAR to regulate Orai1 channel function. Journal of Cell Biology, 202(1):71–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/1/71.

Janke:2014:TCM

[Jan14] Carsten Janke. The tubulin code: Molecular components, readout mechanisms, and functions. Journal of Cell Biology, 206(4):461–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/4/461.

Jayasena:2012:RFC

[JB12] Chathurani S. Jayasena and Marianne E. Bronner. Rbms3 functions in craniofacial development by posttranscriptionally modulating TGF-β signaling. Journal of Cell Biology, 199(3): 453–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/3/453.

Joshi:2012:PMI

[JBS^{+}12] Bharat Joshi, Michele Bastiani, Scott S. Strugnell, Cecile Boscher, Robert G. Parton, and Ivan R. Nabi. Phosphocaveolin-1 is a mechanotransducer that induces caveola biogenesis via Egr1 transcriptional regulation. Journal of Cell Biology, 199 (3):425–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/3/425.

Joshi:2013:PMI

[JBS^{+}13] Bharat Joshi, Michele Bastiani, Scott S. Strugnell, Cecile Boscher, Robert G. Parton, and Ivan R. Nabi. Phosphocaveolin-1 is a mechanotransducer that induces caveola biogenesis via Egr1 transcriptional regulation. Journal of Cell Biology, 200 (5):681–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/200/5/681.

Johnson:2010:OCS

[JC10] Colin P. Johnson and Edwin R. Chapman. Otoferlin is a calcium sensor that directly regulates SNARE-mediated membrane fusion. Journal of Cell Biology, 191(1):187–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
1/187.

Ju:2011:NTA

[JCL^{+}11] Tz-Chuen Ju, Hui-Mei Chen, Jiun-Tsai Lin, Ching-Pang Chang, Wei-Cheng Chang, Jheng-Jie Kang, Cheng-Pu Sun, Mi-Hua Tao, Pang-Hsien Tu, Chen Chang, Dennis W. Dickson, and Yijuang Chern. Nuclear translocation of AMPK-α1 potentiates striatal neurodegeneration in Huntington’s disease. Journal of Cell Biology, 194(2):209–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/2/209.

Jean:2014:IEF

[JCN^{+}14] Christine Jean, Xiao Lei Chen, Ju-Ock Nam, Isabelle Tancioni, Sean Uryu, Christine Lawson, Kristy K. Ward, Colin T. Walsh, Nichol L. G. Miller, Majid Ghassemian, Patric Turowski, Elisabetta Dejana, Sara Weis, David A. Cheresh, and David D. Schlaepfer. Inhibition of endothelial FAK activity prevents tumor metastasis by enhancing barrier function. Journal of Cell Biology, 204(2):247–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/2/247.

Jenkins:2012:NSK

[JDB^{+}12] Brian Jenkins, Helena Decker, Marvin Bentley, Julie Luisi, and Gary Banker. A novel split kinesin assay identifies motor proteins that interact with distinct vesicle populations. Journal of Cell Biology, 198(4):749–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/4/749.

Joset:2010:PGN

[JDHS10] Armela Joset, Dana A. Dodd, Simon Halegoua, and Martin E. Schwab. Pincher-generated Nogo–A endosomes mediate growth cone collapse and retrograde signaling. Journal of Cell Biology, 188(2):271–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/2/271.

Juin:2014:DDR

[JDL^{+}14] Amélie Juin, Julie Di Martino, Birgit Leitinger, Elodie Henriet, Anne-Sophie Gary, Lisa Paysan, Jeremy Bomo, Georges Baffet, Cécile Gauthier-Rouvière, Jean Rosenbaum, Violaine Moreau, and Frédéric Saltel. Discoidin domain receptor 1 controls linear invadosome formation via a Cdc42–Tuba pathway. Journal of Cell Biology, 207(4):517–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
4/517.

Jelluma:2010:RMK

[JDS^{+}10] Nannette Jelluma, Tobias B. Dansen, Tale Sliedrecht, Nicholas P. Kwiatkowski, and Geert J. P. L. Kops. Release of Mps1 from kinetochores is crucial for timely anaphase onset. Journal of Cell Biology, 191(2):281–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/2/281.

Jorgensen:2011:SDP

[JEF^{+}11] Stine Jørgensen, Morten Eskildsen, Kasper Fugger, Lisbeth Hansen, Marie Sofie Yoo Larsen, Arne Nedergaard Kousholt, Randi G. Syljuåsen, Morten Beck Trelle, Ole Nørregaard Jensen, Kristian Helin, and Claus Storgaard Sørensen. SET8 is degraded via PCNA-coupled CRL4(CDT2) ubiquitylation in S phase and after UV irradiation. Journal of Cell Biology, 192 (1):43–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/1/43.

Januschke:2010:IMA

[JG10] Jens Januschke and Cayetano Gonzalez. The interphase microtubule aster is a determinant of asymmetric division orientation in Drosophila neuroblasts. Journal of Cell Biology, 188 (5):693–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/5/693.

Janes:2011:ERF

[JGA^{+}11] Peter W. Janes, Bettina Griesshaber, Lakmali Atapattu, Eva Nievergall, Linda L. Hii, Anneloes Mensinga, Chanly Chheang, Bryan W. Day, Andrew W. Boyd, Philippe I. Bastiaens, Claus Jørgensen, Tony Pawson, and Martin Lackmann. Eph receptor function is modulated by heterooligomerization of A and B type Eph receptors. Journal of Cell Biology, 195(6): 1033–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/6/1033.

Jacquemet:2013:RDR

[JGB^{+}13] Guillaume Jacquemet, David M. Green, Rebecca E. Bridgewater, Alexander von Kriegsheim, Martin J. Humphries, Jim C. Norman, and Patrick T. Caswell. RCP-driven α5β1 recycling suppresses Rac and promotes RhoA activity via the RacGAP1–IQGAP1 complex. Journal of Cell Biology, 202(6): 917–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/6/917.

Johnston:2010:VKP

[JJH^{+}10] Katherine Johnston, Ajit Joglekar, Tetsuya Hori, Aussie Suzuki, Tatsuo Fukagawa, and E. D. Salmon. Vertebrate kinetochore protein architecture: protein copy number. Journal of Cell Biology, 189(6):937–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/6/937.

Jaulin:2010:KSM

[JK10] Fanny Jaulin and Geri Kreitzer. KIF17 stabilizes microtubules and contributes to epithelial morphogenesis by acting at MT plus ends with EB1 and APC. Journal of Cell Biology, 190 (3):443–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/3/443.

Jaqaman:2010:KAW

[JKA^{+}10] Khuloud Jaqaman, Emma M. King, Ana C. Amaro, Jennifer R. Winter, Jonas F. Dorn, Hunter L. Elliott, Nunu Mchedlishvili, Sarah E. McClelland, Iain M. Porter, Markus Posch, Alberto Toso, Gaudenz Danuser, Andrew D. McAinsh, Patrick Meraldi, and Jason R. Swedlow. Kinetochore alignment within the metaphase plate is regulated by centromere stiffness and microtubule depolymerases. Journal of Cell Biology, 188(5):665–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/5/665.

Jones:2014:AIS

[JKS14] Steven L. Jones, Farida Korobova, and Tatyana Svitkina. Axon initial segment cytoskeleton comprises a multiprotein submembranous coat containing sparse actin filaments. Journal of Cell Biology, 205(1):67–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/205/1/67.

Janssen:2012:CTT

[JLVH12] Mandy E. W. Janssen, HongJun Liu, Niels Volkmann, and Dorit Hanein. The C-terminal tail domain of metavinculin, vinculin’s splice variant, severs actin filaments. Journal of Cell Biology, 197(5):585–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/5/585.

Jin:2010:MMP

[JLW^{+}10] Seok Min Jin, Michael Lazarou, Chunxin Wang, Lesley A. Kane, Derek P. Narendra, and Richard J. Youle. Mitochondrial membrane potential regulates PINK1 import and proteolytic destabilization by PARL. Journal of Cell Biology, 191(5): 933–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/5/933.

Jewell:2011:MPI

[JOR^{+}11] Jenna L. Jewell, Eunjin Oh, Latha Ramalingam, Michael A. Kalwat, Vincent S. Tagliabracci, Lixuan Tackett, Jeffrey S. Elmendorf, and Debbie C. Thurmond. Munc18c phosphorylation by the insulin receptor links cell signaling directly to SNARE exocytosis. Journal of Cell Biology, 193(1):185–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
193/1/185.

Joyce:2011:DAA

[JPT^{+}11] Eric F. Joyce, Michael Pedersen, Stanley Tiong, Sanese K. White-Brown, Anshu Paul, Shelagh D. Campbell, and Kim S. McKim. Drosophila ATM and ATR have distinct activities in the regulation of meiotic DNA damage and repair. Journal of Cell Biology, 195(3):359–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/3/359.

Jeong:2013:FPDa

[JRC^{+}13a] Yeon-Tae Jeong, Mario Rossi, Lukas Cermak, Anita Saraf, Laurence Florens, Michael P. Washburn, Patrick Sung, Carl L. Schildkraut, and Michele Pagano. FBH1 promotes DNA double-strand breakage and apoptosis in response to DNA replication stress. Journal of Cell Biology, 200(2):141–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
200/2/141.

Jeong:2013:FPDb

[JRC^{+}13b] Yeon-Tae Jeong, Mario Rossi, Lukas Cermak, Anita Saraf, Laurence Florens, Michael P. Washburn, Patrick Sung, Carl L. Schildkraut, and Michele Pagano. FBH1 promotes DNA double-strand breakage and apoptosis in response to DNA replication stress. Journal of Cell Biology, 201(7):1087–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
201/7/1087.

Jose:2013:RPE

[JTN^{+}13] Mini Jose, Sylvain Tollis, Deepak Nair, Jean-Baptiste Sibarita, and Derek McCusker. Robust polarity establishment occurs via an endocytosis-based cortical corralling mechanism. Journal of Cell Biology, 200(4):407–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/4/407.

Jones:2014:DLI

[JVS^{+}14] Laura A. Jones, Cécile Villemant, Toby Starborg, Anna Salter, Georgina Goddard, Peter Ruane, Philip G. Woodman, Nancy Papalopulu, Sarah Woolner, and Victoria J. Allan. Dynein light intermediate chains maintain spindle bipolarity by functioning in centriole cohesion. Journal of Cell Biology, 207(4): 499–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/4/499.

Jurisch-Yaksi:2013:RMC

[JYRL^{+}13] Nathalie Jurisch-Yaksi, Applonia J. Rose, Huiqi Lu, Tim Raemaekers, Sebastian Munck, Pieter Baatsen, Veerle Baert, Wendy Vermeire, Suzie J. Scales, Daphne Verleyen, Roel Vandepoel, Przemko Tylzanowski, Emre Yaksi, Thomy de Ravel, H. Joseph Yost, Guy Froyen, Cammon B. Arrington, and Wim Annaert. Rer1p maintains ciliary length and signaling by regulating γ-secretase activity and Foxj1a levels. Journal of Cell Biology, 200(6):709–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/6/709.

Kawai:2012:BRM

[KA12] Shinji Kawai and Atsuo Amano. BRCA1 regulates microRNA biogenesis via the DROSHA microprocessor complex. Journal of Cell Biology, 197(2):201–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/197/2/201.

Kehat:2011:MCP

[KAAM11] Izhak Kehat, Federica Accornero, Bruce J. Aronow, and Jeffery D. Molkentin. Modulation of chromatin position and gene expression by HDAC4 interaction with nucleoporins. Journal of Cell Biology, 193(1):21–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/1/21.

Karbstein:2010:CRA

[Kar10] Katrin Karbstein. Chaperoning ribosome assembly. Journal of Cell Biology, 189(1):11–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/1/11.

Kashikar:2012:TSR

[KAS^{+}12] Nachiket D. Kashikar, Luis Alvarez, Reinhard Seifert, Ingo Gregor, Oliver Jäckle, Michael Beyermann, Eberhard Krause, and U. Benjamin Kaupp. Temporal sampling, resetting, and adaptation orchestrate gradient sensing in sperm. Journal of Cell Biology, 198(6):1075–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/6/1075.

Kupchik:2011:NFM

[KBAW^{+}11] Yonatan M. Kupchik, Ofra Barchad-Avitzur, Jürgen Wess, Yair Ben-Chaim, Itzchak Parnas, and Hanna Parnas. A novel fast mechanism for GPCR-mediated signal transduction — control of neurotransmitter release. Journal of Cell Biology, 192(1):137–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/1/137.

Kienzle:2014:CRF

[KBC^{+}14] Christine Kienzle, Nirakar Basnet, Alvaro H. Crevenna, Gisela Beck, Bianca Habermann, Naoko Mizuno, and Julia von Blume. Cofilin recruits F-actin to SPCA1 and promotes Ca^{2+}- mediated secretory cargo sorting. Journal of Cell Biology, 206(5):635–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/5/635.

Kotak:2012:CDC

[KBG12] Sachin Kotak, Coralie Busso, and Pierre Gönczy. Cortical dynein is critical for proper spindle positioning in human cells. Journal of Cell Biology, 199(1):97–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/1/97.

Krahn:2010:FBS

[KBKW10] Michael P. Krahn, Johanna Bückers, Lars Kastrup, and Andreas Wodarz. Formation of a Bazooka–Stardust complex is essential for plasma membrane polarity in epithelia. Journal of Cell Biology, 190(5):751–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/5/751.

Koirala:2010:MAD

[KBS^{+}10] Sajjan Koirala, Huyen T. Bui, Heidi L. Schubert, Debra M. Eckert, Christopher P. Hill, Michael S. Kay, and Janet M. Shaw. Molecular architecture of a dynamin adaptor: implications for assembly of mitochondrial fission complexes. Journal of Cell Biology, 191(6):1127–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/6/1127.

Krick:2010:CPS

[KBW^{+}10] Roswitha Krick, Sebastian Bremer, Evelyn Welter, Petra Schlotterhose, Yvonne Muehe, Eeva-Liisa Eskelinen, and Michael Thumm. Cdc48/ p97 and Shp1/p47 regulate autophagosome biogenesis in concert with ubiquitin-like Atg8. Journal of Cell Biology, 190(6):965–??, September 2010. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/190/6/
965.

Kraft:2012:WFR

[KBW^{+}12] Bianca Kraft, Corinna D. Berger, Veronika Wallkamm, Herbert Steinbeisser, and Doris Wedlich. Wnt-11 and Fz7 reduce cell adhesion in convergent extension by sequestration of PAPC and C-cadherin. Journal of Cell Biology, 198(4):695–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
198/4/695.

Klingner:2014:IAD

[KCF^{+}14] Christoph Klingner, Anoop V. Cherian, Johannes Fels, Philipp M. Diesinger, Roland Aufschnaiter, Nicola Maghelli, Thomas Keil, Gisela Beck, Iva M. Tolić-Nørrelykke, Mark Bathe, and Roland Wedlich-Soldner. Isotropic actomyosin dynamics promote organization of the apical cell cortex in epithelial cells. Journal of Cell Biology, 207(1):107–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
1/107.

Klingberg:2014:PEM

[KCK^{+}14] Franco Klingberg, Melissa L. Chow, Anne Koehler, Stellar Boo, Lara Buscemi, Thomas M. Quinn, Mercedes Costell, Benjamin A. Alman, Elisabeth Genot, and Boris Hinz. Prestress in the extracellular matrix sensitizes latent TGF-β1 for activation. Journal of Cell Biology, 207(2):283–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
2/283.

Kobayashi:2011:RTC

[KD11] Tetsuo Kobayashi and Brian D. Dynlacht. Regulating the transition from centriole to basal body. Journal of Cell Biology, 193(3):435–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/3/435.

Knox:2011:DDK

[KDIE11] Pauline G. Knox, Clare C. Davies, Marina Ioannou, and Aristides G. Eliopoulos. The death domain kinase RIP1 links the immunoregulatory CD40 receptor to apoptotic signaling in carcinomas. Journal of Cell Biology, 192(3):391–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
3/391.

Keller:2012:COB

[KdKDP12] Rebecca Keller, Jeanine de Keyzer, Arnold J. M. Driessen, and Tracy Palmer. Co-operation between different targeting pathways during integration of a membrane protein. Journal of Cell Biology, 199(2):303–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/2/303.

Kapustina:2013:CDM

[KEJ13] Maryna Kapustina, Timothy C. Elston, and Ken Jacobson. Compression and dilation of the membrane-cortex layer generates rapid changes in cell shape. Journal of Cell Biology, 200 (1):95–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/200/1/95.

Kang:2011:ACA

[KF11] Richard S. Kang and Heike Fölsch. ARH cooperates with AP-1B in the exocytosis of LDLR in polarized epithelial cells. Journal of Cell Biology, 193(1):51–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/1/51.

Kim:2011:SOT

[KFET11] Moshe S. Kim, Carol D. Froese, Mathew P. Estey, and William S. Trimble. SEPT9 occupies the terminal positions in septin octamers and mediates polymerization-dependent functions in abscission. Journal of Cell Biology, 195(5):815– ??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/5/815.

Kousholt:2012:CDD

[KFH^{+}12] Arne Nedergaard Kousholt, Kasper Fugger, Saskia Hoffmann, Brian D. Larsen, Tobias Menzel, Alessandro A. Sartori, and Claus Storgaard Sørensen. CtIP-dependent DNA resection is required for DNA damage checkpoint maintenance but not initiation. Journal of Cell Biology, 197(7):869–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
7/869.

Kushner:2014:ECD

[KFL^{+}14] Erich J. Kushner, Luke S. Ferro, Jie-Yu Liu, Jessica R. Durrant, Stephen L. Rogers, Andrew C. Dudley, and Victoria L. Bautch. Excess centrosomes disrupt endothelial cell migration via centrosome scattering. Journal of Cell Biology, 206 (2):257–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/2/257.

Kong:2014:CMR

[KFS^{+}14] Dong Kong, Veronica Farmer, Anil Shukla, Jana James, Richard Gruskin, Shigeo Kiriyama, and Jadranka Loncarek. Centriole maturation requires regulated Plk1 activity during two consecutive cell cycles. Journal of Cell Biology, 206(7): 855–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/7/855.

Kenner:2011:MLJ

[KHB^{+}11a] Lukas Kenner, Astrid Hoebertz, F. Timo Beil, Niamh Keon, Florian Karreth, Robert Eferl, Harald Scheuch, Agnieszka Szremska, Michael Amling, Marina Schorpp-Kistner, Peter Angel, and Erwin F. Wagner. Mice lacking JunB are osteopenic due to cell-autonomous osteoblast and osteoclast defects. Journal of Cell Biology, 195(6):1063–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
6/1063.

Krementsova:2011:TSH

[KHB^{+}11b] Elena B. Krementsova, Alex R. Hodges, Carol S. Bookwalter, Thomas E. Sladewski, Mirko Travaglia, H. Lee Sweeney, and Kathleen M. Trybus. Two single-headed myosin V motors bound to a tetrameric adapter protein form a processive complex. Journal of Cell Biology, 195(4):631–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
4/631.

Kassan:2013:ACS

[KHFV^{+}13] Adam Kassan, Albert Herms, Andrea Fernández-Vidal, Marta Bosch, Nicole L. Schieber, Babu J. N. Reddy, Alba Fajardo, Mariona Gelabert-Baldrich, Francesc Tebar, Carlos Enrich, Steven P. Gross, Robert G. Parton, and Albert Pol. Acyl-CoA synthetase 3 promotes lipid droplet biogenesis in ER microdomains. Journal of Cell Biology, 203(6):985–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
6/985.

Kittelmann:2013:LSD

[KHG^{+}13] Maike Kittelmann, Jan Hegermann, Alexandr Goncharov, Hidenori Taru, Mark H. Ellisman, Janet E. Richmond, Yishi Jin, and Stefan Eimer. Liprin-α / SYD-2 determines the size of dense projections in presynaptic active zones in C. elegans. Journal of Cell Biology, 203(5):849–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
5/849.

Karotki:2011:EPA

[KHS^{+}11] Lena Karotki, Juha T. Huiskonen, Christopher J. Stefan, Natasza E. Ziółkowska, Robyn Roth, Michal A. Surma, Nevan J. Krogan, Scott D. Emr, John Heuser, Kay Grünewald, and Tobias C. Walther. Eisosome proteins assemble into a membrane scaffold. Journal of Cell Biology, 195(5):889– ??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/5/889.

Kim:2010:SPM

[KHW^{+}10] Seung-Hwan Kim, Antonia H. Holway, Suzanne Wolff, Andrew Dillin, and W. Matthew Michael. SMK-1/PPH-4.1-mediated silencing of the CHK-1 response to DNA damage in early C. elegans embryos. Journal of Cell Biology, 189(7):1187–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
7/1187.

Kikkawa:2013:BST

[Kik13] Masahide Kikkawa. Big steps toward understanding dynein. Journal of Cell Biology, 202(1):15–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/1/15.

Klein:2012:CIB

[KIL^{+}12] Astrid Klein, Lars Israel, Sebastian W. K. Lackey, Frank E. Nargang, Axel Imhof, Wolfgang Baumeister, Walter Neupert, and Dennis R. Thomas. Characterization of the insertase for β-barrel proteins of the outer mitochondrial membrane. Journal of Cell Biology, 199(4):599–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/4/599.

King:2013:SSC

[Kin13] Stephen M. King. A solid-state control system for dyneinbased ciliary/flagellar motility. Journal of Cell Biology, 201 (2):173–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/2/173.

Kiyomitsu:2010:ICF

[KIOY10] Tomomi Kiyomitsu, Osamu Iwasaki, Chikashi Obuse, and Mitsuhiro Yanagida. Inner centromere formation requires hMis14, a trident kinetochore protein that specifically recruits HP1 to human chromosomes. Journal of Cell Biology, 188(6):791–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
188/6/791.

Kestler:2011:GWS

[KK11] Hans A. Kestler and Michael Kühl. Generating a Wnt switch: it’s all about the right dosage. Journal of Cell Biology, 193 (3):431–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/3/431.

Kepert:2013:MNE

[KK13a] Inge Kepert and Michael A. Kiebler. MoniTORing neuronal excitability at the synapse. Journal of Cell Biology, 202 (1):7–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/1/7.

Kirkland:2013:LRH

[KK13b] Jacob G. Kirkland and Rohinton T. Kamakaka. Long-range heterochromatin association is mediated by silencing and double-strand DNA break repair proteins. Journal of Cell Biology, 201(6):809–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/6/809.

Kurischko:2011:YCK

[KKK^{+}11] Cornelia Kurischko, Hong Kyung Kim, Venkata K. Kuravi, Juliane Pratzka, and Francis C. Luca. The yeast Cbk1 kinase regulates mRNA localization via the mRNA-binding protein Ssd1. Journal of Cell Biology, 192(4):583–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
4/583.

Kim:2011:PMA

[KKL^{+}11] Nam Hee Kim, Hyun Sil Kim, Xiao-Yan Li, Inhan Lee, Hyung-Seok Choi, Shi Eun Kang, So Young Cha, Joo Kyung Ryu, Dojun Yoon, Eric R. Fearon, R. Grant Rowe, Sanghyuk Lee, Christopher A. Maher, Stephen J. Weiss, and Jong In Yook. A p53/miRNA-34 axis regulates Snail1-dependent cancer cell epithelial–mesenchymal transition. Journal of Cell Biology, 195(3):417–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/3/417.

Kobayashi:2014:CIP

[KKL^{+}14] Tetsuo Kobayashi, Sehyun Kim, Yu-Chun Lin, Takanari Inoue, and Brian David Dynlacht. The CP110-interacting proteins Talpid3 and Cep290 play overlapping and distinct roles in cilia assembly. Journal of Cell Biology, 204(2):215–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
2/215.

Kanaani:2010:TDM

[KKMB10] Jamil Kanaani, Julia Kolibachuk, Hugo Martinez, and Steinunn Baekkeskov. Two distinct mechanisms target GAD67 to vesicular pathways and presynaptic clusters. Journal of Cell Biology, 190(5):911–??, September 2010. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/190/5/911.

Kasprowicz:2014:DPB

[KKS^{+}14] Jaroslaw Kasprowicz, Sabine Kuenen, Jef Swerts, Katarzyna Miskiewicz, and Patrik Verstreken. Dynamin photoinactivation blocks Clathrin and α-adaptin recruitment and induces bulk membrane retrieval. Journal of Cell Biology, 204(7): 1141–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/7/1141.

Kama:2011:YBD

[KKUG11] Rachel Kama, Vydehi Kanneganti, Christian Ungermann, and Jeffrey E. Gerst. The yeast Batten disease orthologue Btn1 controls endosome-Golgi retrograde transport via SNARE assembly. Journal of Cell Biology, 195(2):203–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
2/203.

Kim:2014:KLF

[KKY^{+}14] Jung Ha Kim, Kabsun Kim, Bang Ung Youn, Jongwon Lee, Inyoung Kim, Hong-In Shin, Haruhiko Akiyama, Yongwon Choi, and Nacksung Kim. Kruppel-like factor 4 attenuates osteoblast formation, function, and cross talk with osteoclasts. Journal of Cell Biology, 204(6):1063–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/6/1063.

Kuang:2010:SDC

[KLC^{+}10] Zhihe Kuang, Rowena S. Lewis, Joan M. Curtis, Yifan Zhan, Bernadette M. Saunders, Jeffrey J. Babon, Tatiana B. Kolesnik, Andrew Low, Seth L. Masters, Tracy A. Willson, Lukasz Kedzierski, Shenggen Yao, Emanuela Handman, Raymond S. Norton, and Sandra E. Nicholson. The SPRY domain– containing SOCS box protein SPSB2 targets iNOS for proteasomal degradation. Journal of Cell Biology, 190(1):129–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
190/1/129.

Kane:2014:PPU

[KLF^{+}14] Lesley A. Kane, Michael Lazarou, Adam I. Fogel, Yan Li, Koji Yamano, Shireen A. Sarraf, Soojay Banerjee, and Richard J. Youle. PINK1 phosphorylates ubiquitin to activate Parkin E3 ubiquitin ligase activity. Journal of Cell Biology, 205 (2):143–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/2/143.

Kelley:2014:TBM

[KLHS14] Laura C. Kelley, Lauren L. Lohmer, Elliott J. Hagedorn, and David R. Sherwood. Traversing the basement membrane in vivo: a diversity of strategies. Journal of Cell Biology, 204(3): 291–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/3/291.

Kang:2014:BAC

[KLP14a] Pil Jung Kang, Mid Eum Lee, and Hay-Oak Park. Bud3 activates Cdc42 to establish a proper growth site in budding yeast. Journal of Cell Biology, 206(1):19–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/1/19.

Klement:2014:OIC

[KLP^{+}14b] Karolin Klement, Martijn S. Luijsterburg, Jordan B. Pinder, Chad S. Cena, Victor Del Nero, Christopher M. Wintersinger, Graham Dellaire, Haico van Attikum, and Aaron A. Goodarzi. Opposing ISWI- and CHD-class chromatin remodeling activities orchestrate heterochromatic DNA repair. Journal of Cell Biology, 207(6):717–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/6/717.

Kroger:2013:KCI

[KLS^{+}13] Cornelia Kröger, Fanny Loschke, Nicole Schwarz, Reinhard Windoffer, Rudolf E. Leube, and Thomas M. Magin. Keratins control intercellular adhesion involving PKC-α-mediated desmoplakin phosphorylation. Journal of Cell Biology, 201 (5):681–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/5/681.

Koritzinsky:2013:TPD

[KLvdB^{+}13] Marianne Koritzinsky, Fiana Levitin, Twan van den Beucken, Ryan A. Rumantir, Nicholas J. Harding, Kenneth C. Chu, Paul C. Boutros, Ineke Braakman, and Bradly G. Wouters. Two phases of disulfide bond formation have differing requirements for oxygen. Journal of Cell Biology, 203(4):615– ??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/4/615.

Kappel:2012:RAA

[KLZ^{+}12] Lisa Kappel, Mathias Loibl, Gertrude Zisser, Isabella Klein, Gernot Fruhmann, Christof Gruber, Stefan Unterweger, Gerald Rechberger, Brigitte Pertschy, and Helmut Bergler. Rlp24 activates the AAA–ATPase Drg1 to initiate cytoplasmic pre-60S maturation. Journal of Cell Biology, 199(5):771– ??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/5/771.

Knoops:2014:PVC

[KMC^{+}14] Kèvin Knoops, Selvambigai Manivannan, Małgorzata N. Cepińska, Arjen M. Krikken, Anita M. Kram, Marten Veenhuis, and Ida J. van der Klei. Preperoxisomal vesicles can form in the absence of Pex3. Journal of Cell Biology, 204 (5):659–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/5/659.

Krumova:2011:SIS

[KMG^{+}11] Petranka Krumova, Erik Meulmeester, Manuel Garrido, Marilyn Tirard, He-Hsuan Hsiao, Guillaume Bossis, Henning Urlaub, Markus Zweckstetter, Sebastian Kügler, Frauke Melchior, Mathias Bähr, and Jochen H. Weishaupt. Sumoylation inhibits α-synuclein aggregation and toxicity. Journal of Cell Biology, 194(1):49–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/1/49.

Konig:2010:MRC

[KMS10] Cornelia König, Hiromi Maekawa, and Elmar Schiebel. Mutual regulation of cyclin-dependent kinase and the mitotic exit network. Journal of Cell Biology, 188(3):351–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
3/351.

Kotadia:2012:CEA

[KMSR12] Shaila Kotadia, Emilie Montembault, William Sullivan, and Anne Royou. Cell elongation is an adaptive response for clearing long chromatid arms from the cleavage plane. Journal of Cell Biology, 199(5):745–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/5/745.

King:2012:CBR

[KN12] Ryan S. King and Phillip A. Newmark. The cell biology of regeneration. Journal of Cell Biology, 196(5):553–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
5/553.

Kohzaki:2010:DPR

[KNH^{+}10] Masaoki Kohzaki, Kana Nishihara, Kouji Hirota, Eiichiro Sonoda, Michio Yoshimura, Shigeo Ekino, John E. Butler, Masami Watanabe, Thanos D. Halazonetis, and Shunichi Takeda. DNA polymerases ν and θ are required for efficient immunoglobulin V gene diversification in chicken. Journal of Cell Biology, 189 (7):1117–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/7/1117.

Kiuchi:2011:MSC

[KNOM11] Tai Kiuchi, Tomoaki Nagai, Kazumasa Ohashi, and Kensaku Mizuno. Measurements of spatiotemporal changes in G-actin concentration reveal its effect on stimulus-induced actin assembly and lamellipodium extension. Journal of Cell Biology, 193 (2):365–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/2/365.

Kawai-Noma:2010:VEF

[KNPK^{+}10] Shigeko Kawai-Noma, Chan-Gi Pack, Tomoko Kojidani, Haruhiko Asakawa, Yasushi Hiraoka, Masataka Kinjo, Tokuko Haraguchi, Hideki Taguchi, and Aiko Hirata. In vivo evidence for the fibrillar structures of Sup35 prions in yeast cells. Journal of Cell Biology, 190(2):223–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/2/223.

Kim:2013:RRA

[KNsMK13] Kee K. Kim, Joseph Nam, Yoh suke Mukouyama, and Sachiyo Kawamoto. Rbfox3-regulated alternative splicing of Numb promotes neuronal differentiation during development. Journal of Cell Biology, 200(4):443–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/4/443.

Kagami:2014:MPC

[KNW^{+}14] Yuya Kagami, Keishi Nihira, Shota Wada, Masaya Ono, Mariko Honda, and Kiyotsugu Yoshida. Mps1 phosphorylation of condensin II controls chromosome condensation at the onset of mitosis. Journal of Cell Biology, 205(6):781–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/205/
6/781.

Kamasaki:2013:ADM

[KOK^{+}13] Tomoko Kamasaki, Eileen O’Toole, Shigeo Kita, Masako Osumi, Jiro Usukura, J. Richard McIntosh, and Gohta Goshima. Augmin-dependent microtubule nucleation at microtubule walls in the spindle. Journal of Cell Biology, 202(1):25–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
202/1/25.

Keller:2014:MHA

[KOO^{+}14] Debora Keller, Meritxell Orpinell, Nicolas Olivier, Malte Wachsmuth, Robert Mahen, Romain Wyss, Virginie Hachet, Jan Ellenberg, Suliana Manley, and Pierre Gönczy. Mechanisms of HsSAS-6 assembly promoting centriole formation in human cells. Journal of Cell Biology, 204(5):697–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
5/697.

Kawamura:2010:MCC

[KPC^{+}10] Ryo Kawamura, Lisa H. Pope, Morten O. Christensen, Mingxuan Sun, Ksenia Terekhova, Fritz Boege, Christian Mielke, Anni H. Andersen, and John F. Marko. Mitotic chromosomes are constrained by topoisomerase II–sensitive DNA entanglements. Journal of Cell Biology, 188(5):653–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
5/653.

Khavandgar:2011:CAR

[KPC^{+}11] Zohreh Khavandgar, Christophe Poirier, Christopher J. Clarke, Jingjing Li, Nicholas Wang, Marc D. McKee, Yusuf A. Hannun, and Monzur Murshed. A cell-autonomous requirement for neutral sphingomyelinase 2 in bone mineralization. Journal of Cell Biology, 194(2):277–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/2/277.

Kachaner:2014:IAR

[KPE^{+}14] David Kachaner, Xavier Pinson, Khaled Ben El Kadhi, Karine Normandin, Lama Talje, Hugo Lavoie, Guillaume Lépine, Sébastien Carréno, Benjamin H. Kwok, Gilles R. Hickson, and Vincent Archambault. Interdomain allosteric regulation of Polo kinase by Aurora B and Map205 is required for cytokinesis. Journal of Cell Biology, 207(2):201–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
2/201.

Kim:2012:PMC

[KPH^{+}12] Hyung Joon Kim, Vikram Prasad, Seok-Won Hyung, Zang Hee Lee, Sang-Won Lee, Aditi Bhargava, David Pearce, Youngkyun Lee, and Hong-Hee Kim. Plasma membrane calcium ATPase regulates bone mass by fine-tuning osteoclast differentiation and survival. Journal of Cell Biology, 199(7): 1145–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/7/1145.

Koplin:2010:DFC

[KPI^{+}10] Ansgar Koplin, Steffen Preissler, Yulia Ilina, Miriam Koch, Annika Scior, Marc Erhardt, and Elke Deuerling. A dual function for chaperones SSB–RAC and the NAC nascent polypeptide–associated complex on ribosomes. Journal of Cell Biology, 189(1):57–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/1/57.

Kim:2013:CAR

[KPJ^{+}13] Ingyu Kim, Weijun Pan, Sara A. Jones, Youxin Zhang, Xiaowei Zhuang, and Dianqing Wu. Clathrin and AP2 are required for PtdIns(4,5)P_{2}-mediated formation of LRP6 signalosomes. Journal of Cell Biology, 200(4):419–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/200/
4/419.

Kim:2012:MNR

[KPSL12] Deok-Ho Kim, Paolo P. Provenzano, Chris L. Smith, and Andre Levchenko. Matrix nanotopography as a regulator of cell function. Journal of Cell Biology, 197(3):351–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
3/351.

Kramer:2013:RDA

[Krä13] Helmut Krämer. Route to destruction: Autophagosomes SNARE lysosomes. Journal of Cell Biology, 201(4):495–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
201/4/495.

Krzywicka-Racka:2011:RCF

[KRS11] Anna Krzywicka-Racka and Greenfield Sluder. Repeated cleavage failure does not establish centrosome amplification in untransformed human cells. Journal of Cell Biology, 194 (2):199–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/2/199.

Kleiblova:2013:GFM

[KSB^{+}13] Petra Kleiblova, Indra A. Shaltiel, Jan Benada, Jan evčík, Soňa Pecháčková, Petr Pohlreich, Emile E. Voest, Pavel Dundr, Jiri Bartek, Zdenek Kleibl, René H. Medema, and Libor Macurek. Gain-of-function mutations of PPM1D/Wip1 impair the p53-dependent G1 checkpoint. Journal of Cell Biology, 201(4):511–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/4/511.

Kress:2013:UPP

[KSH^{+}13] Elsa Kress, Françoise Schwager, René Holtackers, Jonas Seiler, François Prodon, Esther Zanin, Annika Eiteneuer, Mika Toya, Asako Sugimoto, Hemmo Meyer, Patrick Meraldi, and Monica Gotta. The UBXN-2/p37/p47 adaptors of CDC-48/p97 regulate mitosis by limiting the centrosomal recruitment of Aurora A. Journal of Cell Biology, 201(4):559–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/4/559.

Ko:2011:NLP

[KSLF^{+}11] Jaewon Ko, Gilberto J. Soler-Llavina, Marc V. Fuccillo, Robert C. Malenka, and Thomas C. Südhof. Neuroligins/ LRRTMs prevent activity- and Ca^{2+} /calmodulin-dependent synapse elimination in cultured neurons. Journal of Cell Biology, 194(2):323–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/2/323.

Kasai:2011:FCG

[KSP^{+}11] Rinshi S. Kasai, Kenichi G. N. Suzuki, Eric R. Prossnitz, Ikuko Koyama-Honda, Chieko Nakada, Takahiro K. Fujiwara, and Akihiro Kusumi. Full characterization of GPCR monomer– dimer dynamic equilibrium by single molecule imaging. Journal of Cell Biology, 192(3):463–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/3/463.

Knaevelsrud:2013:MRP

[KSR^{+}13a] Helene Knævelsrud, Kristiane Søreng, Camilla Raiborg, Karin Håberg, Fredrik Rasmuson, Andreas Brech, Knut Liestøl, Tor Erik Rusten, Harald Stenmark, Thomas P. Neufeld, Sven R. Carlsson, and Anne Simonsen. Membrane remodeling by the PX–BAR protein SNX18 promotes autophagosome formation. Journal of Cell Biology, 202(2):331–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
2/331.

Kuhns:2013:MAR

[KSR^{+}13b] Stefanie Kuhns, Kerstin N. Schmidt, Jürgen Reymann, Daniel F. Gilbert, Annett Neuner, Birgit Hub, Ricardo Carvalho, Philipp Wiedemann, Hanswalter Zentgraf, Holger Erfle, Ursula Klingmüller, Michael Boutros, and Gislene Pereira. The microtubule affinity regulating kinase MARK4 promotes axoneme extension during early ciliogenesis. Journal of Cell Biology, 200(4):505–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/4/505.

Kuipers:2011:HSL

[KSS^{+}11] Marjorie A. Kuipers, Timothy J. Stasevich, Takayo Sasaki, Korey A. Wilson, Kristin L. Hazelwood, James G. McNally, Michael W. Davidson, and David M. Gilbert. Highly stable loading of Mcm proteins onto chromatin in living cells requires replication to unload. Journal of Cell Biology, 192 (1):29–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/1/29.

Kumagai:2011:DRT

[KSSD11] Akiko Kumagai, Anna Shevchenko, Andrej Shevchenko, and William G. Dunphy. Direct regulation of Treslin by cyclindependent kinase is essential for the onset of DNA replication. Journal of Cell Biology, 193(6):995–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/6/995.

Kim:2012:BEA

[KSSK12] Sunghwan Kim, Dionisia P. Sideris, Carolyn S. Sevier, and Chris A. Kaiser. Balanced Ero1 activation and inactivation establishes ER redox homeostasis. Journal of Cell Biology, 196 (6):713–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/6/713.

Kohyama:2010:BIR

[KST^{+}10] Jun Kohyama, Tsukasa Sanosaka, Akinori Tokunaga, Eriko Takatsuka, Keita Tsujimura, Hideyuki Okano, and Kinichi Nakashima. BMP-induced REST regulates the establishment and maintenance of astrocytic identity. Journal of Cell Biology, 189(1):159–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/1/159.

Kerr:2011:MND

[KST^{+}11] Gary W. Kerr, Sourav Sarkar, Katherine L. Tibbles, Mark Petronczki, Jonathan B. A. Millar, and Prakash Arumugam. Meiotic nuclear divisions in budding yeast require PP2A^{Cdc55}- mediated antagonism of Net1 phosphorylation by Cdk. Journal of Cell Biology, 193(7):1157–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/7/1157.

Kukulski:2011:CFE

[KSW^{+}11] Wanda Kukulski, Martin Schorb, Sonja Welsch, Andrea Picco, Marko Kaksonen, and John A. G. Briggs. Correlated fluorescence and 3D electron microscopy with high sensitivity and spatial precision. Journal of Cell Biology, 192(1):111–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
192/1/111.

Khurana:2010:PTS

[KT10] Jaspreet S. Khurana and William Theurkauf. piRNAs, transposon silencing, and Drosophila germline development. Journal of Cell Biology, 191(5):905–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/5/905.

Kuriyama:2014:VCC

[KTB^{+}14] Sei Kuriyama, Eric Theveneau, Alexandre Benedetto, Maddy Parsons, Masamitsu Tanaka, Guillaume Charras, Alexandre Kabla, and Roberto Mayor. In vivo collective cell migration requires an LPAR2-dependent increase in tissue fluidity. Journal of Cell Biology, 206(1):113–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/206/1/113.

Kiyozumi:2012:BMA

[KTN^{+}12] Daiji Kiyozumi, Makiko Takeichi, Itsuko Nakano, Yuya Sato, Tomohiko Fukuda, and Kiyotoshi Sekiguchi. Basement membrane assembly of the integrin α8β1 ligand nephronectin requires Fraser syndrome–associated proteins. Journal of Cell Biology, 197(5):677–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/5/677.

Kotter:2014:HMP

[KUH^{+}14] Sebastian Kötter, Andreas Unger, Nazha Hamdani, Patrick Lang, Matthias Vorgerd, Luitgard Nagel-Steger, and Wolfgang A. Linke. Human myocytes are protected from titin aggregation-induced stiffening by small heat shock proteins. Journal of Cell Biology, 204(2):187–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/2/187.

Kamimura:2013:PRB

[KUN^{+}13] Keisuke Kamimura, Kohei Ueno, Jun Nakagawa, Rie Hamada, Minoru Saitoe, and Nobuaki Maeda. Perlecan regulates bidirectional Wnt signaling at the Drosophila neuromuscular junction. Journal of Cell Biology, 200(2):219–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/200/
2/219.

Kim:2010:NEM

[KWDD10] Sungsu Kim, Yogesh P. Wairkar, Richard W. Daniels, and Aaron DiAntonio. The novel endosomal membrane protein Ema interacts with the class C Vps–HOPS complex to promote endosomal maturation. Journal of Cell Biology, 188(5):717–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
188/5/717.

Koivusalo:2010:AIMa

[KWH^{+}10a] Mirkka Koivusalo, Christopher Welch, Hisayoshi Hayashi, Cameron C. Scott, Moshe Kim, Todd Alexander, Nicolas Touret, Klaus M. Hahn, and Sergio Grinstein. Amiloride inhibits macropinocytosis by lowering submembranous pH and preventing Rac1 and Cdc42 signaling. Journal of Cell Biology, 188(4):547–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/4/547.

Koivusalo:2010:AIMb

[KWH^{+}10b] Mirkka Koivusalo, Christopher Welch, Hisayoshi Hayashi, Cameron C. Scott, Moshe Kim, Todd Alexander, Nicolas Touret, Klaus M. Hahn, and Sergio Grinstein. Amiloride inhibits macropinocytosis by lowering submembranous pH and preventing Rac1 and Cdc42 signaling. Journal of Cell Biology, 189(2):385–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/2/385.

Kanai:2014:KEE

[KWH14] Yoshimitsu Kanai, Daliang Wang, and Nobutaka Hirokawa. KIF13B enhances the endocytosis of LRP1 by recruiting LRP1 to caveolae. Journal of Cell Biology, 204(3):395–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
3/395.

Kurihara:2011:APH

[KWK^{+}11] Toshihide Kurihara, Peter D. Westenskow, Tim U. Krohne, Edith Aguilar, Randall S. Johnson, and Martin Friedlander. Astrocyte pVHL and HIF-α isoforms are required for embryonic-to-adult vascular transition in the eye. Journal of Cell Biology, 195(4):689–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/4/689.

Krenn:2012:SAR

[KWL^{+}12] Veronica Krenn, Annemarie Wehenkel, Xiaozheng Li, Stefano Santaguida, and Andrea Musacchio. Structural analysis reveals features of the spindle checkpoint kinase Bub1– kinetochore subunit Knl1 interaction. Journal of Cell Biology, 196(4):451–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/4/451.

Krachler:2011:MKS

[KWO11] Anne Marie Krachler, Andrew R. Woolery, and Kim Orth. Manipulation of kinase signaling by bacterial pathogens. Journal of Cell Biology, 195(7):1083–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/7/1083.

Krueger:2010:LIL

[KWTR10] Lori E. Krueger, Jui-Ching Wu, Meng-Fu Bryan Tsou, and Lesilee S. Rose. LET-99 inhibits lateral posterior pulling forces during asymmetric spindle elongation in C. elegans embryos. Journal of Cell Biology, 189(3):481–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/3/481.

Kanehara:2010:MHE

[KXN10] Kazue Kanehara, Wei Xie, and Davis T. W. Ng. Modularity of the Hrd1 ERAD complex underlies its diverse client range. Journal of Cell Biology, 188(5):707–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/188/5/707.

Kim:2012:TAI

[KYHG12] Chungho Kim, Feng Ye, Xiaohui Hu, and Mark H. Ginsberg. Talin activates integrins by altering the topology of the β transmembrane domain. Journal of Cell Biology, 197 (5):605–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/5/605.

Kotani:2013:CBM

[KYOY13] Tomoya Kotani, Kyota Yasuda, Ryoma Ota, and Masakane Yamashita. Cyclin B1 mRNA translation is temporally controlled through formation and disassembly of RNA granules. Journal of Cell Biology, 202(7):1041–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
7/1041.

Kunduri:2014:PAP

[KYP^{+}14] Govind Kunduri, Changqing Yuan, Velayoudame Parthibane, Katherine M. Nyswaner, Ritu Kanwar, Kunio Nagashima, Steven G. Britt, Nickita Mehta, Varshika Kotu, Mindy Porterfield, Michael Tiemeyer, Patrick J. Dolph, Usha Acharya, and Jairaj K. Acharya. Phosphatidic acid phospholipase A1 mediates ER–Golgi transit of a family of G protein-coupled receptors. Journal of Cell Biology, 206(1):79–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/1/79.

Kageyama:2012:MDE

[KZR^{+}12] Yusuke Kageyama, Zhongyan Zhang, Ricardo Roda, Masahiro Fukaya, Junko Wakabayashi, Nobunao Wakabayashi, Thomas W. Kensler, P. Hemachandra Reddy, Miho Iijima, and Hiromi Sesaki. Mitochondrial division ensures the survival of postmitotic neurons by suppressing oxidative damage. Journal of Cell Biology, 197(4):535–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/4/535.

Lykke-Andersen:2014:PPE

[LAB14] Jens Lykke-Andersen and Eric J. Bennett. Protecting the proteome: Eukaryotic cotranslational quality control pathways. Journal of Cell Biology, 204(4):467–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
4/467.

Liu:2010:LCT

[LADS10] Allen P. Liu, François Aguet, Gaudenz Danuser, and Sandra L. Schmid. Local clustering of transferrin receptors promotes clathrin-coated pit initiation. Journal of Cell Biology, 191(7): 1381–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/7/1381.

Liem:2012:ICR

[LAH^{+}12] Karel F. Liem, Alyson Ashe, Mu He, Peter Satir, Jennifer Moran, David Beier, Carol Wicking, and Kathryn V. Anderson. The IFT–A complex regulates Shh signaling through cilia structure and membrane protein trafficking. Journal of Cell Biology, 197(6):789–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/6/789.

Libersou:2010:DSR

[LAO^{+}10] Sonia Libersou, Aurélie A. V. Albertini, Malika Ouldali, Virginie Maury, Christine Maheu, Hélène Raux, Felix de Haas, Stéphane Roche, Yves Gaudin, and Jean Lepault. Distinct structural rearrangements of the VSV glycoprotein drive membrane fusion. Journal of Cell Biology, 191(1):199–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
1/199.

Lam:2010:LAS

[LAR^{+}10] Vinh Q. Lam, David Akopian, Michael Rome, Doug Henningsen, and Shu ou Shan. Lipid activation of the signal recognition particle receptor provides spatial coordination of protein targeting. Journal of Cell Biology, 190(4):623–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
4/623.

Laflamme:2012:EPC

[LAR^{+}12] Carl Laflamme, Gloria Assaker, Damien Ramel, Jonas F. Dorn, Desmond She, Paul S. Maddox, and Gregory Emery. Evi5 promotes collective cell migration through its Rab–GAP activity. Journal of Cell Biology, 198(1):57–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/198/1/57.

Lorenzo:2014:PAB

[LBD^{+}14] Damaris Nadia Lorenzo, Alexandra Badea, Jonathan Davis, Janell Hostettler, Jiang He, Guisheng Zhong, Xiaowei Zhuang, and Vann Bennett. A PIK3C3–Ankyrin–B–Dynactin pathway promotes axonal growth and multiorganelle transport. Journal of Cell Biology, 207(6):735–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/207/6/735.

Lawrimore:2011:PCC

[LBS11] Josh Lawrimore, Kerry S. Bloom, and E. D. Salmon. Point centromeres contain more than a single centromere-specific Cse4 (CENP–A) nucleosome. Journal of Cell Biology, 195(4): 573–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/4/573.

Lechtreck:2013:CSP

[LBS^{+}13] Karl F. Lechtreck, Jason M. Brown, Julio L. Sampaio, Julie M. Craft, Andrej Shevchenko, James E. Evans, and George B. Witman. Cycling of the signaling protein phospholipase D through cilia requires the BBSome only for the export phase. Journal of Cell Biology, 201(2):249–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/2/249.

Lin:2010:WRP

[LBWS10] Shengda Lin, Lisa M. Baye, Trudi A. Westfall, and Diane C. Slusarski. Wnt5b–Ryk pathway provides directional signals to regulate gastrulation movement. Journal of Cell Biology, 190(2):263–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/2/263.

Lan:2010:CTB

[LC10] Weijie Lan and Don W. Cleveland. A chemical tool box defines mitotic and interphase roles for Mps1 kinase. Journal of Cell Biology, 190(1):21–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/1/21.

Lasserre:2011:RST

[LCBG^{+}11] Rémi Lasserre, Céline Cuche, Ronnie Blecher-Gonen, Evgeny Libman, Elise Biquand, Anne Danckaert, Deborah Yablonski, Andrés Alcover, and Vincenzo Di Bartolo. Release of serine/ threonine-phosphorylated adaptors from signaling microclusters down-regulates T cell activation. Journal of Cell Biology, 195(5):839–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/5/839.

Lewellyn:2011:CPC

[LCD^{+}11] Lindsay Lewellyn, Ana Carvalho, Arshad Desai, Amy S. Maddox, and Karen Oegema. The chromosomal passenger complex and centralspindlin independently contribute to contractile ring assembly. Journal of Cell Biology, 193(1):155–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
193/1/155.

Laursen:2011:TMB

[LCfC11] Lisbeth S. Laursen, Colin W. Chan, and Charles ffrench Constant. Translation of myelin basic protein mRNA in oligodendrocytes is regulated by integrin activation and hnRNP–K. Journal of Cell Biology, 192(5):797–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/5/797.

Lewellyn:2013:MDI

[LCHB13] Lindsay Lewellyn, Maureen Cetera, and Sally Horne-Badovinac. Misshapen decreases integrin levels to promote epithelial motility and planar polarity in Drosophila. Journal of Cell Biology, 200(6):721–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/6/721.

Liu:2013:NPS

[LCK^{+}13] Yunhao Liu, Caitlin Collins, William B. Kiosses, Ann M. Murray, Monika Joshi, Tyson R. Shepherd, Ernesto J. Fuentes, and Ellie Tzima. A novel pathway spatiotemporally activates Rac1 and redox signaling in response to fluid shear stress. Journal of Cell Biology, 201(6):863–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/201/6/863.

Lee:2012:CMT

[LCL12] Samuel M. Lee, Lih-Shen Chin, and Lian Li. Charcot-marie-tooth disease-linked protein SIMPLE functions with the ES-CRT machinery in endosomal trafficking. Journal of Cell Biology, 199(5):799–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/5/799.

Laporte:2011:AAP

[LCLW11] Damien Laporte, Valerie C. Coffman, I-Ju Lee, and Jian-Qiu Wu. Assembly and architecture of precursor nodes during fission yeast cytokinesis. Journal of Cell Biology, 192(6):1005–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
192/6/1005.

Lewis:2013:CMM

[LCP13] Tommy L. Lewis, Julien Courchet, and Franck Polleux. Cellular and molecular mechanisms underlying axon formation, growth, and branching. Journal of Cell Biology, 202(6):837– ??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/6/837.

Lee:2010:MID

[LCS^{+}10] Chan-Soo Lee, Chang-Ki Choi, Eun-Young Shin, Martin Alexander Schwartz, and Eung-Gook Kim. Myosin II directly binds and inhibits Dbl family guanine nucleotide exchange factors: a possible link to Rho family GTPases. Journal of Cell Biology, 190(4):663–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/4/663.

Laporte:2013:ANM

[LCS^{+}13] Damien Laporte, Fabien Courtout, Bénédicte Salin, Johanna Ceschin, and Isabelle Sagot. An array of nuclear microtubules reorganizes the budding yeast nucleus during quiescence. Journal of Cell Biology, 203(4):585–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/4/585.

Lazaro-Dieguez:2013:PLL

[LDCF^{+}13] Francisco Lázaro-Diéguez, David Cohen, Dawn Fernandez, Louis Hodgson, Sven C. D. van IJzendoorn, and Anne Müsch. Par1b links lumen polarity with LGN–NuMA positioning for distinct epithelial cell division phenotypes. Journal of Cell Biology, 203(2):251–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/2/251.

Liu:2012:PLK

[LDL12] Dan Liu, Olga Davydenko, and Michael A. Lampson. Pololike kinase-1 regulates kinetochore–microtubule dynamics and spindle checkpoint silencing. Journal of Cell Biology, 198(4): 491–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/4/491.

Liou:2013:MSC

[LDN^{+}13] Geou-Yarh Liou, Heike Döppler, Brian Necela, Murli Krishna, Howard C. Crawford, Massimo Raimondo, and Peter Storz. Macrophage-secreted cytokines drive pancreatic acinarto-ductal metaplasia through NF-κB and MMPs. Journal of Cell Biology, 202(3):563–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/202/3/563.

leDuc:2010:VPCa

[lDSB^{+}10a] Quint le Duc, Quanming Shi, Iris Blonk, Arnoud Sonnenberg, Ning Wang, Deborah Leckband, and Johan de Rooij. Vinculin potentiates E-cadherin mechanosensing and is recruited to actin-anchored sites within adherens junctions in a myosin II-dependent manner. Journal of Cell Biology, 189(7):1107–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
189/7/1107.

leDuc:2010:VPCb

[lDSB^{+}10b] Quint le Duc, Quanming Shi, Iris Blonk, Arnoud Sonnenberg, Ning Wang, Deborah Leckband, and Johan de Rooij. Vinculin potentiates E-cadherin mechanosensing and is recruited to actin-anchored sites within adherens junctions in a myosin II-dependent manner. Journal of Cell Biology, 191(4): 891–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/4/891.

LeBrasseur:2010:WRI

[LeB10] Nicole LeBrasseur. Wolf Reik: Inheritance beyond DNA. Journal of Cell Biology, 188(3):302–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/188/3/302.

Leslie:2010:AMT

[Les10a] Mitch Leslie. APC: More than a beta (catenin) blocker. Journal of Cell Biology, 189(7):1055–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/7/1055.

Leslie:2010:ADH

[Les10b] Mitch Leslie. Ataxia’s double hit. Journal of Cell Biology, 189(1):??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/1/6.2.

Leslie:2010:CSR

[Les10c] Mitch Leslie. Cdks set the replication schedule. Journal of Cell Biology, 188(2):??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/2/176.1.

Leslie:2010:CRW

[Les10d] Mitch Leslie. Centrosomes remember which way is up. Journal of Cell Biology, 188(5):615–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/5/615.

Leslie:2010:CTS

[Les10e] Mitch Leslie. Clathrin teams up to strengthen the spindle. Journal of Cell Biology, 189(7):??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/7/1054.1.

Leslie:2010:CCEa

[Les10f] Mitch Leslie. CLIP catches enzymes in the act. Journal of Cell Biology, 191(1):??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/1/2.2.

Leslie:2010:CCEb

[Les10g] Mitch Leslie. CLIP catches enzymes in the act. Journal of Cell Biology, 191(2):431–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/2/431.

Leslie:2010:CEC

[Les10h] Mitch Leslie. Cytoplasmic enzymes cast NETs. Journal of Cell Biology, 191(3):??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/3/434.3.

Leslie:2010:DCS

[Les10i] Mitch Leslie. Doubling up on cellular stress. Journal of Cell Biology, 191(6):??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/6/1044.3.

Leslie:2010:DDT

[Les10j] Mitch Leslie. dRich doesn’t take sides. Journal of Cell Biology, 191(3):??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/3/434.2.

Leslie:2010:ESS

[Les10k] Mitch Leslie. Elm1 sparks the SPOC. Journal of Cell Biology, 190(6):??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/6/944.1.

Leslie:2010:ETO

[Les10l] Mitch Leslie. Endocytosis takes occludin for a ride. Journal of Cell Biology, 189(1):7–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/1/7.

Leslie:2010:ESC

[Les10m] Mitch Leslie. An enzyme with self-control. Journal of Cell Biology, 188(2):177–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/2/177.

Leslie:2010:FGA

[Les10n] Mitch Leslie. FYCO1 gets autophagy on track. Journal of Cell Biology, 188(2):??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/2/176.3.

Leslie:2010:GSC

[Les10o] Mitch Leslie. Go on, satellite cells, be all that you can be. Journal of Cell Biology, 190(5):??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/5/708.2.

Leslie:2010:HSC

[Les10p] Mitch Leslie. Helping stem cells make a connection. Journal of Cell Biology, 191(3):435–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/3/435.

Leslie:2010:HVE

[Les10q] Mitch Leslie. How a virus enters without breaking. Journal of Cell Biology, 191(1):??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/1/2.3.

Leslie:2010:IPG

[Les10r] Mitch Leslie. IFT proteins go off the rails. Journal of Cell Biology, 189(1):??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/1/6.3.

Leslie:2010:ITA

[Les10s] Mitch Leslie. Ino80 tidies up after DNA repair. Journal of Cell Biology, 191(6):??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/6/1044.1.

Leslie:2010:IOC

[Les10t] Mitch Leslie. Ins and outs of Cdc14. Journal of Cell Biology, 190(2):??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/2/160.3.

Leslie:2010:LDA

[Les10u] Mitch Leslie. Lysosomes don’t accentuate the negative. Journal of Cell Biology, 189(7):??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/7/1054.3.

Leslie:2010:MSM

[Les10v] Mitch Leslie. Mad2 shifts a motor into idle. Journal of Cell Biology, 191(6):??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/6/1044.2.

Leslie:2010:MHC

[Les10w] Mitch Leslie. A MAP holds the corners. Journal of Cell Biology, 191(3):??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/3/434.1.

Leslie:2010:MPG

[Les10x] Mitch Leslie. Misfolded proteins go their separate ways. Journal of Cell Biology, 188(2):??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/188/2/176.2.

Leslie:2010:MHD

[Les10y] Mitch Leslie. Mitochondria help dynamins get a grip. Journal of Cell Biology, 191(6):1045–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/6/1045.

Leslie:2010:MFD

[Les10z] Mitch Leslie. Mitosis first, DNA repair later. Journal of Cell Biology, 190(2):??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/2/160.2.

Leslie:2010:NLC

[Les10-27] Mitch Leslie. Neighbors limit cell renovations. Journal of Cell Biology, 188(5):??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/5/614.3.

Leslie:2010:ANP

[Les10-28] Mitch Leslie. Not all nuclear pores created equal. Journal of Cell Biology, 191(1):3–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/1/3.

Leslie:2010:PPB

[Les10-29] Mitch Leslie. Peter Pan bone cells undermine skeleton. Journal of Cell Biology, 189(3):389–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/3/389.

Leslie:2010:RDE

[Les10-30] Mitch Leslie. RALT doubles down on EGFR. Journal of Cell Biology, 189(3):??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/3/388.3.

Leslie:2010:RMP

[Les10-31] Mitch Leslie. Ras makes a pit stop. Journal of Cell Biology, 191(1):??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/1/2.1.

Leslie:2010:RR

[Les10-32] Mitch Leslie. Repetitive but not redundant. Journal of Cell Biology, 190(2):161–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/2/161.

Leslie:2010:SPI

[Les10-33] Mitch Leslie. Secret passage to the inner nuclear membrane. Journal of Cell Biology, 189(7):??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/7/1054.2.

Leslie:2010:SCS

[Les10-34] Mitch Leslie. SENP6 cuts SUMOs down to size. Journal of Cell Biology, 188(5):??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/5/614.2.

Leslie:2010:SKM

[Les10-35] Mitch Leslie. Skeleton key for metastasis. Journal of Cell Biology, 189(3):??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/3/388.2.

Leslie:2010:SMC

[Les10-36] Mitch Leslie. Smooth muscle cells put their best podosome forward. Journal of Cell Biology, 189(1):??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/1/6.1.

Leslie:2010:SRC

[Les10-37] Mitch Leslie. Sweet recipe for cellular closeness. Journal of Cell Biology, 190(5):709–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/190/5/709.

Leslie:2010:TDT

[Les10-38] Mitch Leslie. Tangled DNA tightens chromosomes. Journal of Cell Biology, 188(5):??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/5/614.1.

Leslie:2010:UPN

[Les10-39] Mitch Leslie. Unfaithful proteins no problem for DNA repair. Journal of Cell Biology, 189(3):??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/3/388.1.

Leslie:2010:UNH

[Les10-40] Mitch Leslie. Unwrapping new histones. Journal of Cell Biology, 190(5):??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/5/708.1.

Leslie:2010:VTA

[Les10-41] Mitch Leslie. Virus takes the actin express. Journal of Cell Biology, 190(2):??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/2/160.1.

Leslie:2010:HHC

[Les10-42] Mitch Leslie. With HMGB1’s help, cells dine in. Journal of Cell Biology, 190(5):??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/190/5/708.3.

Leslie:2011:ARB

[Les11a] Mitch Leslie. The actomyosin ring bulks up. Journal of Cell Biology, 195(5):??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/5/704.2.

Leslie:2011:AKT

[Les11b] Mitch Leslie. Annexin keeps tau on a short leash. Journal of Cell Biology, 192(4):??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/4/542.2.

Leslie:2011:ABG

[Les11c] Mitch Leslie. Aurora B goes the distance. Journal of Cell Biology, 194(4):509–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/4/509.

Leslie:2011:BAG

[Les11d] Mitch Leslie. Bak activators get in the groove. Journal of Cell Biology, 194(1):??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/1/2.2.

Leslie:2011:CCM

[Les11e] Mitch Leslie. Cancer cells mothball mRNA after microtubule collapse. Journal of Cell Biology, 192(1):??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/192/1/2.
2.

Leslie:2011:CSN

[Les11f] Mitch Leslie. Caveolin-1 says NO to permeability. Journal of Cell Biology, 193(5):??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/5/800.2.

Leslie:2011:CLT

[Les11g] Mitch Leslie. Cdc6 licenses tumor growth. Journal of Cell Biology, 195(7):??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/7/1066.3.

Leslie:2011:EIE

[Les11h] Mitch Leslie. ESCRT-III has an escort. Journal of Cell Biology, 192(2):??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/2/204.1.

Leslie:2011:FKC

[Les11i] Mitch Leslie. Fibers keep cilia regular. Journal of Cell Biology, 195(1):??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/1/2.1.

Leslie:2011:HMP

[Les11j] Mitch Leslie. Hedgehog, meet Patched. Journal of Cell Biology, 192(4):543–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/4/543.

Leslie:2011:HEF

[Les11k] Mitch Leslie. Helping the embryo find closure. Journal of Cell Biology, 192(2):205–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/2/205.

Leslie:2011:HFT

[Les11l] Mitch Leslie. A histone finds a traveling companion. Journal of Cell Biology, 193(3):??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/3/426.2.

Leslie:2011:HEH

[Les11m] Mitch Leslie. How eisosomes help the plasma membrane get organized. Journal of Cell Biology, 195(5):705–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
5/705.

Leslie:2011:HTC

[Les11n] Mitch Leslie. HPK1 tones down the T cell receptor. Journal of Cell Biology, 195(5):??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/5/704.3.

Leslie:2011:KMC

[Les11o] Mitch Leslie. A kinase makes a connection. Journal of Cell Biology, 194(6):??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/6/808.3.

Leslie:2011:KFP

[Les11p] Mitch Leslie. Kinetochores feel the pull. Journal of Cell Biology, 193(1):3–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/1/3.

Leslie:2011:LFN

[Les11q] Mitch Leslie. Lis1 finds new digs on the desmosome. Journal of Cell Biology, 194(4):??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/4/508.2.

Leslie:2011:MKC

[Les11r] Mitch Leslie. MCPH1 keeps chromosomes strung out. Journal of Cell Biology, 194(6):??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/6/808.2.

Leslie:2011:MMHa

[Les11s] Mitch Leslie. Mnemonic microRNAs help make memories. Journal of Cell Biology, 194(6):809–??, September 2011. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/194/6/
809.

Leslie:2011:MMHb

[Les11t] Mitch Leslie. Mobile moesin helps mitotic cells stretch out. Journal of Cell Biology, 195(1):??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/1/2.3.

Leslie:2011:MRI

[Les11u] Mitch Leslie. Multiple routes to the inner nuclear membrane. Journal of Cell Biology, 193(1):??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/1/2.2.

Leslie:2011:WAA

[Les11v] Mitch Leslie. N-WASP allows axons to cover up. Journal of Cell Biology, 192(2):??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/2/204.3.

Leslie:2011:NKC

[Les11w] Mitch Leslie. Natural killer cells commute death sentence. Journal of Cell Biology, 192(4):??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/4/542.3.

Leslie:2011:NST

[Les11x] Mitch Leslie. Neuregulins show their support for neuromuscular junctions. Journal of Cell Biology, 195(7):1067–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
195/7/1067.

Leslie:2011:NTC

[Les11y] Mitch Leslie. New Twist on cell migration. Journal of Cell Biology, 194(1):??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/1/2.1.

Leslie:2011:NRB

[Les11z] Mitch Leslie. Noncoding RNA to blame for bad editing. Journal of Cell Biology, 193(5):??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/5/800.3.

Leslie:2011:OCS

[Les11-27] Mitch Leslie. One collagen shipment, ready for delivery. Journal of Cell Biology, 193(5):801–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/5/801.

Leslie:2011:PDD

[Les11-28] Mitch Leslie. p63 delegates during skin development. Journal of Cell Biology, 194(6):??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/6/808.1.

Leslie:2011:PPC

[Les11-29] Mitch Leslie. PAR proteins cross the boundary. Journal of Cell Biology, 193(3):427–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/3/427.

Leslie:2011:PMP

[Les11-30] Mitch Leslie. Paraspeckles may provide stress relief. Journal of Cell Biology, 193(1):??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/1/2.1.

Leslie:2011:PVG

[Les11-31] Mitch Leslie. Plant virus gets the bends. Journal of Cell Biology, 193(3):??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/3/426.3.

Leslie:2011:PKC

[Les11-32] Mitch Leslie. Proteins keep Cdc55 in its place. Journal of Cell Biology, 193(3):??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/3/426.1.

Leslie:2011:PBC

[Les11-33] Mitch Leslie. Protocadherins bring cells together. Journal of Cell Biology, 195(7):??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/7/1066.2.

Leslie:2011:RSI

[Les11-34] Mitch Leslie. RanBP2 stops importin-β from running away. Journal of Cell Biology, 194(4):??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/4/508.1.

Leslie:2011:RFR

[Les11-35] Mitch Leslie. Receptors flex to remove neurotransmitter brake. Journal of Cell Biology, 192(1):??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/1/2.3.

Leslie:2011:SRT

[Les11-36] Mitch Leslie. The sarcoplasmic reticulum takes on a supporting role. Journal of Cell Biology, 194(1):3–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/194/1/3.

Leslie:2011:SNL

[Les11-37] Mitch Leslie. Shining new light on the function of Aurora A. Journal of Cell Biology, 195(7):??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/7/1066.1.

Leslie:2011:SIC

[Les11-38] Mitch Leslie. Sphingolipid imbalance can take out an eye. Journal of Cell Biology, 192(4):??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/4/542.1.

Leslie:2011:SEP

[Les11-39] Mitch Leslie. Stiff ECM puts the brakes on microtubule growth. Journal of Cell Biology, 192(2):??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/192/2/
204.2.

Leslie:2011:SDP

[Les11-40] Mitch Leslie. SUMO defeats protein aggregates. Journal of Cell Biology, 194(1):??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/1/2.3.

Leslie:2011:SAJ

[Les11-41] Mitch Leslie. Switching adherens junctions to seal wounds. Journal of Cell Biology, 194(4):??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/4/508.3.

Leslie:2011:SGM

[Les11-42] Mitch Leslie. Syndecan-3 gets the message. Journal of Cell Biology, 192(1):??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/1/2.1.

Leslie:2011:TDE

[Les11-43] Mitch Leslie. TPX2 is a drag on Eg5. Journal of Cell Biology, 195(1):??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/1/2.2.

Leslie:2011:TMB

[Les11-44] Mitch Leslie. Two microscopes are better than one. Journal of Cell Biology, 192(1):3–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/1/3.

Leslie:2011:VLE

[Les11-45] Mitch Leslie. VPS35 leaves endosomes lost in transition. Journal of Cell Biology, 195(5):??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/5/704.1.

Leslie:2011:WCA

[Les11-46] Mitch Leslie. WASH cleans up after sloppy eaters. Journal of Cell Biology, 193(5):??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/5/800.1.

Leslie:2011:WPS

[Les11-47] Mitch Leslie. Why a protein switches sides during translation. Journal of Cell Biology, 195(1):3–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/1/3.

Leslie:2011:WSM

[Les11-48] Mitch Leslie. Wnt sends mixed signals in the skin. Journal of Cell Biology, 193(1):??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/1/2.3.

Leslie:2012:ST

[Les12a] Mitch Leslie. As the sperm turns. Journal of Cell Biology, 196(5):??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/5/548.3.

Leslie:2012:APE

[Les12b] Mitch Leslie. Asters push embryonic nuclei to the brink. Journal of Cell Biology, 197(7):853–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/7/853.

Leslie:2012:ABN

[Les12c] Mitch Leslie. Aurora B is no Ska fan. Journal of Cell Biology, 196(5):??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/5/548.1.

Leslie:2012:BSS

[Les12d] Mitch Leslie. B23 steals to stop cell-killing pair. Journal of Cell Biology, 199(1):??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/1/2.1.

Leslie:2012:BTM

[Les12e] Mitch Leslie. BRCA1 touches up microRNAs. Journal of Cell Biology, 197(2):??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/2/162.2.

Leslie:2012:CPW

[Les12f] Mitch Leslie. CaMKII points the way. Journal of Cell Biology, 198(6):??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/6/954.3.

Leslie:2012:CCS

[Les12g] Mitch Leslie. Caspase conspiracy sets up cells for death. Journal of Cell Biology, 196(4):??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/196/4/390.3.

Leslie:2012:CTT

[Les12h] Mitch Leslie. Chaperones team up twice. Journal of Cell Biology, 198(3):??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/3/266.2.

Leslie:2012:CMM

[Les12i] Mitch Leslie. CLASPing microtubules to the membrane. Journal of Cell Biology, 198(3):??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/3/266.3.

Leslie:2012:CSI

[Les12j] Mitch Leslie. Cohesin sticks to it during meiosis. Journal of Cell Biology, 197(7):??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/7/852.1.

Leslie:2012:CSU

[Les12k] Mitch Leslie. Crumbs stands united against ROS. Journal of Cell Biology, 198(6):??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/6/954.2.

Leslie:2012:CK

[Les12l] Mitch Leslie. Cytoskeleton key. Journal of Cell Biology, 199 (1):3–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/1/3.

Leslie:2012:DHH

[Les12m] Mitch Leslie. Dynein’s helper not so helpful. Journal of Cell Biology, 199(1):??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/1/2.3.

Leslie:2012:FWC

[Les12n] Mitch Leslie. Fly wing cells choose sides. Journal of Cell Biology, 196(5):549–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/5/549.

Leslie:2012:FPF

[Les12o] Mitch Leslie. Fungal pores feel the pull. Journal of Cell Biology, 198(3):??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/3/266.1.

Leslie:2012:HKC

[Les12p] Mitch Leslie. HIV’s kinesin chauffeur. Journal of Cell Biology, 199(3):??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/3/402.1.

Leslie:2012:KSA

[Les12q] Mitch Leslie. Knl1 shows another face. Journal of Cell Biology, 196(4):391–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/4/391.

Leslie:2012:LCL

[Les12r] Mitch Leslie. Lipins cause a lamina breakdown. Journal of Cell Biology, 198(6):??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/6/954.1.

Leslie:2012:LCW

[Les12s] Mitch Leslie. Lis1 cuts its work short. Journal of Cell Biology, 197(7):??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/7/852.3.

Leslie:2012:MBR

[Les12t] Mitch Leslie. microRNA’s big role in preventing tumors. Journal of Cell Biology, 199(1):??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/1/2.2.

Leslie:2012:WHC

[Les12u] Mitch Leslie. N-WASP helps cancer cells open up. Journal of Cell Biology, 199(3):??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/3/402.3.

Leslie:2012:NN

[Les12v] Mitch Leslie. Nck by the numbers. Journal of Cell Biology, 197(5):571–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/5/571.

Leslie:2012:NCS

[Les12w] Mitch Leslie. No crystal structure? No problem. Journal of Cell Biology, 196(4):??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/4/390.1.

Leslie:2012:OSD

[Les12x] Mitch Leslie. Only skin deep, thanks to microtubules. Journal of Cell Biology, 199(3):??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/3/402.2.

Leslie:2012:OM

[Les12y] Mitch Leslie. Outgrowing meiosis. Journal of Cell Biology, 198 (3):267–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/3/267.

Leslie:2012:PAU

[Les12z] Mitch Leslie. Playing with Arp2/3 uncovers cellular rudder. Journal of Cell Biology, 197(2):163–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/2/163.

Leslie:2012:REF

[Les12-27] Mitch Leslie. Recycling endosomes feed autophagy. Journal of Cell Biology, 197(5):??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/5/570.2.

Leslie:2012:RFR

[Les12-28] Mitch Leslie. RNF169 flips the repair switch. Journal of Cell Biology, 197(2):??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/2/162.1.

Leslie:2012:SSN

[Les12-29] Mitch Leslie. Slow but steady for NF-κB. Journal of Cell Biology, 196(4):??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/4/390.2.

Leslie:2012:SRS

[Les12-30] Mitch Leslie. Spare the Rod1, spoil transporter endocytosis. Journal of Cell Biology, 196(2):181–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/196/2/181.

Leslie:2012:SHT

[Les12-31] Mitch Leslie. Stay-at-home transcription factor saves axons. Journal of Cell Biology, 199(3):403–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/3/403.

Leslie:2012:TGI

[Les12-32] Mitch Leslie. Talin gives integrins a nudge. Journal of Cell Biology, 197(5):??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/5/570.1.

Leslie:2012:UBM

[Les12-33] Mitch Leslie. Unfinished basement membrane spurs developmental disorder. Journal of Cell Biology, 197(5):??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
5/570.3.

Leslie:2012:USI

[Les12-34] Mitch Leslie. Unlocking the sperm’s internal compass. Journal of Cell Biology, 198(6):955–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/6/955.

Leslie:2012:VMG

[Les12-35] Mitch Leslie. Vinculin minds the gap. Journal of Cell Biology, 196(5):??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/5/548.2.

Leslie:2012:WSI

[Les12-36] Mitch Leslie. Wait, save that integrin! Journal of Cell Biology, 197(2):??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/2/162.3.

Leslie:2012:WYM

[Les12-37] Mitch Leslie. Why yeast mothers and daughters don’t share. Journal of Cell Biology, 197(7):??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/7/852.2.

Leslie:2013:CST

[Les13a] Mitch Leslie. β-catenin solidifies tight junctions in the skin. Journal of Cell Biology, 202(1):??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/1/2.2.

Leslie:2013:ACP

[Les13b] Mitch Leslie. Actin-capping protein lets epithelial cells take the pressure. Journal of Cell Biology, 203(5):??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
5/712.2.

Leslie:2013:ADC

[Les13c] Mitch Leslie. Actin draws a cadherin crowd. Journal of Cell Biology, 201(1):??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/1/2.3.

Leslie:2013:BAF

[Les13d] Mitch Leslie. Blame ATAD5 for factory closures. Journal of Cell Biology, 200(1):??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/1/2.2.

Leslie:2013:BMP

[Les13e] Mitch Leslie. Bone morphogenetic protein guts it out. Journal of Cell Biology, 201(6):??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/6/780.3.

Leslie:2013:CAA

[Les13f] Mitch Leslie. Catching augmin in the act. Journal of Cell Biology, 202(1):??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/1/2.1.

Leslie:2013:CHC

[Les13g] Mitch Leslie. Cdc42 helps cells take bigger bites. Journal of Cell Biology, 200(1):??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/1/2.3.

Leslie:2013:CIH

[Les13h] Mitch Leslie. Condensin II helps sisters find their identity. Journal of Cell Biology, 200(4):??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/4/362.3.

Leslie:2013:CCF

[Les13i] Mitch Leslie. Crawling cells feel the squeeze. Journal of Cell Biology, 202(5):721–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/202/5/721.

Leslie:2013:CCI

[Les13j] Mitch Leslie. Creating cellular inequality. Journal of Cell Biology, 202(6):829–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/202/6/829.

Leslie:2013:DSS

[Les13k] Mitch Leslie. Drebrin shows self-restraint. Journal of Cell Biology, 202(5):??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/5/720.2.

Leslie:2013:DFP

[Les13l] Mitch Leslie. Drs2 flips the phospholipid switch. Journal of Cell Biology, 202(6):??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/202/6/828.1.

Leslie:2013:DCC

[Les13m] Mitch Leslie. Dynamin 2 cuts the cord for newborn lysosome. Journal of Cell Biology, 203(2):??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/2/166.3.

Leslie:2013:CSE

[Les13n] Mitch Leslie. E-cadherin stops embryonic cells from crossing borders. Journal of Cell Biology, 201(6):??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/201/6/
780.2.

Leslie:2013:EPE

[Les13o] Mitch Leslie. Endocytosis puts exocytosis in pole position. Journal of Cell Biology, 200(4):363–??, February 2013. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/200/4/
363.

Leslie:2013:EYC

[Les13p] Mitch Leslie. Ensuring that yeast cells get their inheritance. Journal of Cell Biology, 203(2):167–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/2/167.

Leslie:2013:FLCa

[Les13q] Mitch Leslie. FANCJ lets chromatin stay true. Journal of Cell Biology, 201(1):??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/1/2.1.

Leslie:2013:FLCb

[Les13r] Mitch Leslie. FANCJ lets chromatin stay true. Journal of Cell Biology, 201(2):351–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/2/351.

Leslie:2013:FCD

[Les13s] Mitch Leslie. Fly chromosomes don’t just ride the tips. Journal of Cell Biology, 200(2):??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/2/126.2.

Leslie:2013:HTW

[Les13t] Mitch Leslie. Hippo throws its weight behind migrating cells. Journal of Cell Biology, 201(6):??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/6/780.1.

Leslie:2013:HLB

[Les13u] Mitch Leslie. How to live without BRCA1. Journal of Cell Biology, 200(2):127–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/2/127.

Leslie:2013:INP

[Les13v] Mitch Leslie. Injured neurons pump up the volume. Journal of Cell Biology, 202(5):??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/202/5/720.1.

Leslie:2013:KCN

[Les13w] Mitch Leslie. Kinetochore complexes need teamwork. Journal of Cell Biology, 200(1):??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/1/2.1.

Leslie:2013:LLR

[Les13x] Mitch Leslie. A lifeline for lipid rafts? Journal of Cell Biology, 202(1):3–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/1/3.

Leslie:2013:MSW

[Les13y] Mitch Leslie. Making short work of p53. Journal of Cell Biology, 201(4):491–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/4/491.

Leslie:2013:ME

[Les13z] Mitch Leslie. Means to an end. Journal of Cell Biology, 201 (6):781–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/6/781.

Leslie:2013:PCR

[Les13-27] Mitch Leslie. Polysaccharide’s central role in cell division. Journal of Cell Biology, 203(2):??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/2/166.2.

Leslie:2013:PGS

[Les13-28] Mitch Leslie. Progerin guilty of size discrimination. Journal of Cell Biology, 201(4):??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/4/490.1.

Leslie:2013:PCT

[Les13-29] Mitch Leslie. Putting Cyclin E in the trash. Journal of Cell Biology, 203(2):??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/2/166.1.

Leslie:2013:RDS

[Les13-30] Mitch Leslie. Ran’s downhill slide during mitosis. Journal of Cell Biology, 200(2):??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/2/126.1.

Leslie:2013:RHE

[Les13-31] Mitch Leslie. Rap1 helps endothelial cells stay close. Journal of Cell Biology, 202(6):??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/202/6/828.2.

Leslie:2013:RFK

[Les13-32] Mitch Leslie. Rise and fall of the kinetochore. Journal of Cell Biology, 201(1):3–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/1/3.

Leslie:2013:SGG

[Les13-33] Mitch Leslie. SecYEG gets grabby with new proteins. Journal of Cell Biology, 200(4):??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/4/362.2.

Leslie:2013:SPK

[Les13-34] Mitch Leslie. Shelterin protein keeps telomeres on edge. Journal of Cell Biology, 203(5):??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/5/712.1.

Leslie:2013:SRE

[Les13-35] Mitch Leslie. Silent RNAs express themselves. Journal of Cell Biology, 203(5):713–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/5/713.

Leslie:2013:SNS

[Les13-36] Mitch Leslie. Silicon nanocrystals shine. Journal of Cell Biology, 202(6):??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/6/828.3.

Leslie:2013:SLN

[Les13-37] Mitch Leslie. Sip1 liberates neural crest cells. Journal of Cell Biology, 203(5):??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/5/712.3.

Leslie:2013:SAW

[Les13-38] Mitch Leslie. Spreading around a Wnt protein. Journal of Cell Biology, 200(2):??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/2/126.3.

Leslie:2013:SHC

[Les13-39] Mitch Leslie. SSX2IP helps the centrosome grow up. Journal of Cell Biology, 202(1):??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/1/2.3.

Leslie:2013:SGT

[Les13-40] Mitch Leslie. STT3B gets the tip. Journal of Cell Biology, 201(1):??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/1/2.2.

Leslie:2013:TOT

[Les13-41] Mitch Leslie. Telocentrosomes organize telomere gathering. Journal of Cell Biology, 200(4):??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/4/362.1.

Leslie:2013:TCA

[Les13-42] Mitch Leslie. Three’s not a crowd for actin nucleators. Journal of Cell Biology, 201(4):??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/4/490.3.

Leslie:2013:TSC

[Les13-43] Mitch Leslie. The two stages of cell fusion. Journal of Cell Biology, 200(1):3–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/1/3.

Leslie:2013:WMC

[Les13-44] Mitch Leslie. Why metaphase chromosomes wander. Journal of Cell Biology, 201(4):??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/4/490.2.

Leslie:2014:AMM

[Les14a] Mitch Leslie. Adducin-1 moonlights at the mitotic spindle. Journal of Cell Biology, 204(1):??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/1/2.1.

Leslie:2014:CGS

[Les14b] Mitch Leslie. Career guidance for stem cells. Journal of Cell Biology, 204(4):463–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/204/4/463.

Leslie:2014:CMM

[Les14c] Mitch Leslie. CLAMPing down on microtubules in migratory cells. Journal of Cell Biology, 206(3):331–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
3/331.

Leslie:2014:CWC

[Les14d] Mitch Leslie. Clearing the way for cancer cells. Journal of Cell Biology, 204(2):149–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/2/149.

Leslie:2014:CCF

[Les14e] Mitch Leslie. Closing the calcium floodgates. Journal of Cell Biology, 205(2):??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/2/128.2.

Leslie:2014:DKC

[Les14f] Mitch Leslie. DDR1 key for cancer cell escape. Journal of Cell Biology, 207(4):??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/4/434.2.

Leslie:2014:DAC

[Les14g] Mitch Leslie. A different angle on cell division. Journal of Cell Biology, 205(6):755–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/6/755.

Leslie:2014:DRH

[Les14h] Mitch Leslie. DNA repair with a hefty price tag. Journal of Cell Biology, 206(7):??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/206/7/814.3.

Leslie:2014:DCD

[Les14i] Mitch Leslie. Double-checking DNA synthesis. Journal of Cell Biology, 204(2):??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/2/148.1.

Leslie:2014:EHF

[Les14j] Mitch Leslie. Enzyme helps fold the SAC. Journal of Cell Biology, 206(7):??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/7/814.1.

Leslie:2014:EIG

[Les14k] Mitch Leslie. ESCRT-III gets the bends. Journal of Cell Biology, 206(6):??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/6/690.3.

Leslie:2014:HSH

[Les14l] Mitch Leslie. Histone shortages hold back replication forks. Journal of Cell Biology, 204(1):??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/1/2.2.

Leslie:2014:HUC

[Les14m] Mitch Leslie. HIV’s unkind cut. Journal of Cell Biology, 206 (7):??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/7/814.2.

Leslie:2014:HSG

[Les14n] Mitch Leslie. How sperm get into the zona. Journal of Cell Biology, 205(6):??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/6/754.2.

Leslie:2014:KME

[Les14o] Mitch Leslie. Kicking MiD51 out of the enzyme club. Journal of Cell Biology, 204(4):??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/4/462.1.

Leslie:2014:LDE

[Les14p] Mitch Leslie. Lipid droplets ensure their own consumption. Journal of Cell Biology, 206(3):??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/3/330.1.

Leslie:2014:LPR

[Les14q] Mitch Leslie. Loss of protein remodels shape-shifting parasite. Journal of Cell Biology, 206(3):??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/3/330.2.

Leslie:2014:MCT

[Les14r] Mitch Leslie. Melanoma cells travel on calcium cycles. Journal of Cell Biology, 207(4):??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/4/434.3.

Leslie:2014:MCS

[Les14s] Mitch Leslie. Mitotic cells’ shocking response to stress. Journal of Cell Biology, 206(6):??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/206/6/690.1.

Leslie:2014:MCG

[Les14t] Mitch Leslie. A molecular chain gang at work in maturing ribosomes. Journal of Cell Biology, 207(4):435–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
4/435.

Leslie:2014:MSC

[Les14u] Mitch Leslie. Mps1 sends condensin II off to work. Journal of Cell Biology, 205(6):??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/6/754.1.

Leslie:2014:NCS

[Les14v] Mitch Leslie. Ndc1 catches the shuttle. Journal of Cell Biology, 204(4):??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/4/462.3.

Leslie:2014:NSN

[Les14w] Mitch Leslie. NEMO surfaces in NF-κB-activating clusters. Journal of Cell Biology, 204(2):??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/2/148.2.

Leslie:2014:AVC

[Les14x] Mitch Leslie. Not all vinculins are created equal. Journal of Cell Biology, 205(2):129–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/205/2/129.

Leslie:2014:PBP

[Les14y] Mitch Leslie. Paxillin brings peace to microtubules. Journal of Cell Biology, 206(3):??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/206/3/330.3.

Leslie:2014:PPD

[Les14z] Mitch Leslie. Polarity proteins duel in Drosophila. Journal of Cell Biology, 204(4):??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/4/462.2.

Leslie:2014:SHE

[Les14-27] Mitch Leslie. Sec12 heads for the exits. Journal of Cell Biology, 206(6):??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/6/690.2.

Leslie:2014:SST

[Les14-28] Mitch Leslie. Sending sisters their separate ways. Journal of Cell Biology, 204(1):3–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/1/3.

Leslie:2014:SCS

[Les14-29] Mitch Leslie. The sources of cell stickiness. Journal of Cell Biology, 204(2):??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/2/148.3.

Leslie:2014:SPI

[Les14-30] Mitch Leslie. SPP pulls an inside job. Journal of Cell Biology, 205(6):??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/6/754.3.

Leslie:2014:SSP

[Les14-31] Mitch Leslie. A storehouse for splicing proteins? Journal of Cell Biology, 206(7):815–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/7/815.

Leslie:2014:THT

[Les14-32] Mitch Leslie. Talin holds tight during cell spreading. Journal of Cell Biology, 205(2):??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/2/128.3.

Leslie:2014:THS

[Les14-33] Mitch Leslie. Thyroid hormones speed cellular aging. Journal of Cell Biology, 204(1):??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/1/2.3.

Leslie:2014:TIW

[Les14-34] Mitch Leslie. Topoisomerase II has to work late. Journal of Cell Biology, 206(6):691–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/6/691.

Leslie:2014:TCB

[Les14-35] Mitch Leslie. TRIM9 cuts back on axon branching. Journal of Cell Biology, 205(2):??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/2/128.1.

Leslie:2014:WDT

[Les14-36] Mitch Leslie. Wait! Don’t throw away those proteins! Journal of Cell Biology, 207(4):??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/4/434.1.

Leung:2014:PAR

[Leu14] Anthony K. L. Leung. Poly(ADP-ribose): an organizer of cellular architecture. Journal of Cell Biology, 205(5):613–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/205/
5/613.

Levine:2011:LTO

[Lev11] Tim P. Levine. Lipid traffic: Osh4p makes an unexpected exchange. Journal of Cell Biology, 195(6):927–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
6/927.

Lane:2013:NCT

[LGAC13] Andrew B. Lane, Juan F. Giménez-Abián, and Duncan J. Clarke. A novel chromatin tether domain controls topoisomerase IIα dynamics and mitotic chromosome formation. Journal of Cell Biology, 203(3):471–??, November 2013. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/203/3/
471.

Lorenzo:2010:SMC

[LgLM^{+}10] Damaris N. Lorenzo, Min gang Li, Sarah E. Mische, Karen R. Armbrust, Laura P. W. Ranum, and Thomas S. Hays. Spectrin mutations that cause spinocerebellar ataxia type 5 impair axonal transport and induce neurodegeneration in Drosophila. Journal of Cell Biology, 189(1):143–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/1/143.

LeGrand:2012:SRS

[LGM^{+}12] Fabien Le Grand, Raphaëlle Grifone, Philippos Mourikis, Christophe Houbron, Carine Gigaud, Julien Pujol, Marjorie Maillet, Gilles Pagès, Michael Rudnicki, Shahragim Tajbakhsh, and Pascal Maire. Six1 regulates stem cell repair potential and self-renewal during skeletal muscle regeneration. Journal of Cell Biology, 198(5):815–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
5/815.

Lee:2011:RNC

[LH11] Jibak Lee and Tatsuya Hirano. RAD21L, a novel cohesin subunit implicated in linking homologous chromosomes in mammalian meiosis. Journal of Cell Biology, 192(2):263–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
192/2/263.

Lu:2014:MMD

[LHD^{+}14] Dan Lu, Jennifer Y. Hsiao, Norman E. Davey, Vanessa A. Van Voorhis, Scott A. Foster, Chao Tang, and David O. Morgan. Multiple mechanisms determine the order of APC/C substrate degradation in mitosis. Journal of Cell Biology, 207 (1):23–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/1/23.

Llano:2012:MCC

[LHGT^{+}12] Elena Llano, Yurema Herrán, Ignacio García-Tuñón, Cristina Gutiérrez-Caballero, Enrique de Álava, José Luis Barbero, John Schimenti, Dirk G. de Rooij, Manuel Sánchez-Martín, and Alberto M. Pendás. Meiotic cohesin complexes are essential for the formation of the axial element in mice. Journal of Cell Biology, 197(7):877–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/197/7/877.

Laufman:2011:CCI

[LHL11] Orly Laufman, WanJin Hong, and Sima Lev. The COG complex interacts directly with Syntaxin 6 and positively regulates endosome–to–TGN retrograde transport. Journal of Cell Biology, 194(3):459–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/3/459.

Loughlin:2010:CMP

[LHN10] Rose Loughlin, Rebecca Heald, and François Nédélec. A computational model predicts Xenopus meiotic spindle organization. Journal of Cell Biology, 191(7):1239–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
7/1239.

Lin:2010:CHC

[LHS10] Chiou-Hong Lin, Chi-Kuo Hu, and Hsiu-Ming Shih. Clathrin heavy chain mediates TACC3 targeting to mitotic spindles to ensure spindle stability. Journal of Cell Biology, 189(7): 1097–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/7/1097.

Lampert:2010:DCC

[LHW10] Fabienne Lampert, Peter Hornung, and Stefan Westermann. The Dam1 complex confers microtubule plus end–tracking activity to the Ndc80 kinetochore complex. Journal of Cell Biology, 189(4):641–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/4/641.

Luo:2013:ALO

[LhYL^{+}13] Weiwei Luo, Cheng han Yu, Zi Zhao Lieu, Jun Allard, Alex Mogilner, Michael P. Sheetz, and Alexander D. Bershadsky. Analysis of the local organization and dynamics of cellular actin networks. Journal of Cell Biology, 202(7):1057– ??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/7/1057.

Lindqvist:2010:CBC

[Lin10] Arne Lindqvist. Cyclin B–Cdk1 activates its own pump to get into the nucleus. Journal of Cell Biology, 189(2):197–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
2/197.

Lin:2011:TSN

[LJLJ11] Pen-Jen Lin, Candice G. Jongsma, Shuren Liao, and Arthur E. Johnson. Transmembrane segments of nascent polytopic membrane proteins control cytosol/ER targeting during membrane integration. Journal of Cell Biology, 195(1):41–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
1/41.

Lin:2011:PMP

[LJPJ11] Pen-Jen Lin, Candice G. Jongsma, Martin R. Pool, and Arthur E. Johnson. Polytopic membrane protein folding at L17 in the ribosome tunnel initiates cyclical changes at the translocon. Journal of Cell Biology, 195(1):55–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
1/55.

Laloux:2013:SCP

[LJW13] Géraldine Laloux and Christine Jacobs-Wagner. Spatiotemporal control of PopZ localization through cell cycle–coupled multimerization. Journal of Cell Biology, 201(6):827–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
6/827.

Lindow:2012:DFM

[LK12] Morten Lindow and Sakari Kauppinen. Discovering the first microRNA-targeted drug. Journal of Cell Biology, 199(3): 407–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/3/407.

Lucas:2013:HPP

[LKG^{+}13] Eliana P. Lucas, Ichha Khanal, Pedro Gaspar, Georgina C. Fletcher, Cedric Polesello, Nicolas Tapon, and Barry J. Thompson. The Hippo pathway polarizes the actin cytoskeleton during collective migration of Drosophila border cells. Journal of Cell Biology, 201(6):875–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/6/875.

Lee:2012:NBN

[LKLA12] Sang Bae Lee, Chung Kwon Kim, Kyung-Hoon Lee, and Jee-Yin Ahn. S-nitrosylation of B23/nucleophosmin by GAPDH protects cells from the SIAH1–GAPDH death cascade. Journal of Cell Biology, 199(1):65–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/1/65.

Li:2013:SGC

[LKSG13] Yun R. Li, Oliver D. King, James Shorter, and Aaron D. Gitler. Stress granules as crucibles of ALS pathogenesis. Journal of Cell Biology, 201(3):361–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/201/3/361.

Luijsterburg:2012:DPC

[LLA^{+}12] Martijn S. Luijsterburg, Michael Lindh, Klara Acs, Mischa G. Vrouwe, Alex Pines, Haico van Attikum, Leon H. Mullenders, and Nico P. Dantuma. DDB2 promotes chromatin decondensation at UV-induced DNA damage. Journal of Cell Biology, 197(2):267–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/2/267.

Lin:2011:PRU

[LLcK^{+}11] Mei-Yao Lin, Yu-Min Lin, Te chan Kao, Hsiang-Hao Chuang, and Ruey-Hwa Chen. PDZ-RhoGEF ubiquitination by Cullin3–KLHL20 controls neurotrophin-induced neurite outgrowth. Journal of Cell Biology, 193(6):985–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/193/6/
985.

Lefkimmiatis:2013:IOC

[LLH13] Konstantinos Lefkimmiatis, Daniela Leronni, and Aldebaran M. Hofer. The inner and outer compartments of mitochondria are sites of distinct cAMP/PKA signaling dynamics. Journal of Cell Biology, 202(3):453–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/3/453.

Lu:2011:FPN

[LLK11] Lei Lu, Mark S. Ladinsky, and Tomas Kirchhausen. Formation of the postmitotic nuclear envelope from extended ER cisternae precedes nuclear pore assembly. Journal of Cell Biology, 194(3):425–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/3/425.

Lu:2010:GPC

[LLM^{+}10] Junjie Lu, Feng Li, Christopher S. Murphy, Michael W. Davidson, and David M. Gilbert. G2 phase chromatin lacks determinants of replication timing. Journal of Cell Biology, 189 (6):967–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/6/967.

Longatti:2012:TRA

[LLR^{+}12] Andrea Longatti, Christopher A. Lamb, Minoo Razi, Shin ichiro Yoshimura, Francis A. Barr, and Sharon A. Tooze. TBC1D14 regulates autophagosome formation via Rab11- and ULK1-positive recycling endosomes. Journal of Cell Biology, 197(5):659–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/5/659.

Laporte:2011:MSR

[LLS^{+}11] Damien Laporte, Anne Lebaudy, Annelise Sahin, Benoît Pinson, Johanna Ceschin, Bertrand Daignan-Fornier, and Isabelle Sagot. Metabolic status rather than cell cycle signals control quiescence entry and exit. Journal of Cell Biology, 192 (6):949–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/6/949.

Lemieux:2012:TCD

[LLT^{+}12] Mado Lemieux, Simon Labrecque, Christian Tardif, Étienne Labrie-Dion, Éric LeBel, and Paul De Koninck. Translocation of CaMKII to dendritic microtubules supports the plasticity of local synapses. Journal of Cell Biology, 198(6):1055– ??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/6/1055.

Lawson:2012:FPRa

[LLU^{+}12a] Christine Lawson, Ssang-Taek Lim, Sean Uryu, Xiao Lei Chen, David A. Calderwood, and David D. Schlaepfer. FAK promotes recruitment of talin to nascent adhesions to control cell motility. Journal of Cell Biology, 196(2):223–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
2/223.

Lawson:2012:FPRb

[LLU^{+}12b] Christine Lawson, Ssang-Taek Lim, Sean Uryu, Xiao Lei Chen, David A. Calderwood, and David D. Schlaepfer. FAK promotes recruitment of talin to nascent adhesions to control cell motility. Journal of Cell Biology, 196(3):387–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
3/387.

Levayer:2013:MCC

[LM13] Romain Levayer and Eduardo Moreno. Mechanisms of cell competition: Themes and variations. Journal of Cell Biology, 200(6):689–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/6/689.

Lampert:2013:MRF

[LMA^{+}13] Fabienne Lampert, Christine Mieck, Gregory M. Alushin, Eva Nogales, and Stefan Westermann. Molecular requirements for the formation of a kinetochore–microtubule interface by Dam1 and Ndc80 complexes. Journal of Cell Biology, 200 (1):21–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/200/1/21.

Lim:2012:NLF

[LMC^{+}12] Ssang-Taek Lim, Nichol L. G. Miller, Xiao Lei Chen, Isabelle Tancioni, Colin T. Walsh, Christine Lawson, Sean Uryu, Sara M. Weis, David A. Cheresh, and David D. Schlaepfer. Nuclear-localized focal adhesion kinase regulates inflammatory VCAM-1 expression. Journal of Cell Biology, 197(7):907–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
197/7/907.

Lumpkin:2010:CBT

[LMN10] Ellen A. Lumpkin, Kara L. Marshall, and Aislyn M. Nelson. The cell biology of touch. Journal of Cell Biology, 191(2):

237–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/2/237.

Lee:2010:MIM

[LMS10a] Sang Hyun Lee, Frank McCormick, and Hideyuki Saya. Mad2 inhibits the mitotic kinesin MKlp2. Journal of Cell Biology, 191(6):1069–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/6/1069.

Liu:2010:FAF

[LMS^{+}10b] Fei Liu, Justin D. Mih, Barry S. Shea, Alvin T. Kho, Asma S. Sharif, Andrew M. Tager, and Daniel J. Tschumperlin. Feedback amplification of fibrosis through matrix stiffening and COX-2 suppression. Journal of Cell Biology, 190(4):693–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
190/4/693.

Low:2010:PKR

[LMS^{+}10c] Pei Ching Low, Ryo Misaki, Kate Schroder, Amanda C. Stanley, Matthew J. Sweet, Rohan D. Teasdale, Bart Vanhaesebroeck, Frédéric A. Meunier, Tomohiko Taguchi, and Jennifer L. Stow. Phosphoinositide 3-kinase δ regulates membrane fission of Golgi carriers for selective cytokine secretion. Journal of Cell Biology, 190(6):1053–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/6/1053.

Leikina:2013:EAD

[LMS^{+}13] Evgenia Leikina, Kamran Melikov, Sarmistha Sanyal, Santosh K. Verma, Bokkee Eun, Claudia Gebert, Karl Pfeifer, Vladimir A. Lizunov, Michael M. Kozlov, and Leonid V. Chernomordik. Extracellular annexins and dynamin are important for sequential steps in myoblast fusion. Journal of Cell Biology, 200(1):109–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/1/109.

Lauffer:2010:SMP

[LMT^{+}10] Benjamin E. L. Lauffer, Cristina Melero, Paul Temkin, Cai Lei, Wanjin Hong, Tanja Kortemme, and Mark von Zastrow.

SNX27 mediates PDZ-directed sorting from endosomes to the plasma membrane. Journal of Cell Biology, 190(4):565–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
190/4/565.

Li:2012:RDM

[LMT^{+}12] Wenjing Li, Takashi Moriwaki, Tomomi Tani, Takashi Watanabe, Kozo Kaibuchi, and Gohta Goshima. Reconstitution of dynamic microtubules with Drosophila XMAP215, EB1, and Sentin. Journal of Cell Biology, 199(5):849–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
5/849.

Li:2011:EPM

[LMW^{+}11] Wenjing Li, Tomohiro Miki, Takashi Watanabe, Mai Kakeno, Ikuko Sugiyama, Kozo Kaibuchi, and Gohta Goshima. EB1 promotes microtubule dynamics by recruiting Sentin in Drosophila cells. Journal of Cell Biology, 193(6):973–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
6/973.

Laplante:2011:ADE

[LN11] Caroline Laplante and Laura A. Nilson. Asymmetric distribution of Echinoid defines the epidermal leading edge during Drosophila dorsal closure. Journal of Cell Biology, 192(2): 335–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/2/335.

Liu:2014:BCS

[LN14] Dongmei Liu and Peter Novick. Bem1p contributes to secretory pathway polarization through a direct interaction with Exo70p. Journal of Cell Biology, 207(1):59–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/207/1/59.

Lazarou:2013:PDP

[LNJ^{+}13] Michael Lazarou, Derek P. Narendra, Seok Min Jin, Ephrem Tekle, Soojay Banerjee, and Richard J. Youle. PINK1 drives Parkin self-association and HECT-like E3 activity upstream of mitochondrial binding. Journal of Cell Biology, 200(2): 163–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/200/2/163.

Lander:2011:FBP

[LNL11] Rachel Lander, Kara Nordin, and Carole LaBonne. The F-box protein Ppa is a common regulator of core EMT factors Twist, Snail, Slug, and Sip1. Journal of Cell Biology, 194(1):17–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
194/1/17.

Lewellis:2013:PSM

[LNS^{+}13] Stephen W. Lewellis, Danielle Nagelberg, Abhi Subedi, Alison Staton, Michelle LeBlanc, Antonio Giraldez, and Holger Knaut. Precise SDF1-mediated cell guidance is achieved through ligand clearance and microRNA-mediated decay. Journal of Cell Biology, 200(3):337–??, February 2013. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/200/3/
337.

Lee:2010:DCM

[LNT^{+}10] Joo-Yong Lee, Yoshito Nagano, J. Paul Taylor, Kah Leong Lim, and Tso-Pang Yao. Disease-causing mutations in Parkin impair mitochondrial ubiquitination, aggregation, and HDAC6-dependent mitophagy. Journal of Cell Biology, 189 (4):671–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/4/671.

Lesman:2014:CFR

[LNTR14] Ayelet Lesman, Jacob Notbohm, David A. Tirrell, and Guruswami Ravichandran. Contractile forces regulate cell division in three-dimensional environments. Journal of Cell Biology, 205(2):155–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/2/155.

Ludwig:2010:FMI

[LOR^{+}10] Alexander Ludwig, Grant P. Otto, Kirsi Riento, Emily Hams, Padraic G. Fallon, and Ben J. Nichols. Flotillin microdomains interact with the cortical cytoskeleton to control uropod formation and neutrophil recruitment. Journal of Cell Biology, 191(4):771–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/4/771.

Liu:2011:TKC

[LP11] Lunhua Liu and Carole A. Parent. TOR kinase complexes and cell migration. Journal of Cell Biology, 194(6):815–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
194/6/815.

Lu:2013:GAR

[LP13] Albert Lu and Suzanne R. Pfeffer. Golgi-associated RhoBTB3 targets Cyclin E for ubiquitylation and promotes cell cycle progression. Journal of Cell Biology, 203(2):233–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
2/233.

Larsen:2010:CRF

[LPG^{+}10] Dorthe Helena Larsen, Catherine Poinsignon, Thorkell Gudjonsson, Christoffel Dinant, Mark R. Payne, Flurina J. Hari, Jannie M. Rendtlew Danielsen, Patrice Menard, Jette Christensen Sand, Manuel Stucki, Claudia Lukas, Jiri Bartek, Jens S. Andersen, and Jiri Lukas. The chromatin-remodeling factor CHD4 coordinates signaling and repair after DNA damage. Journal of Cell Biology, 190(5):731–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
5/731.

Lee:2011:PPP

[LR11a] Kwanwoo Lee and Kunsoo Rhee. PLK1 phosphorylation of pericentrin initiates centrosome maturation at the onset of mitosis. Journal of Cell Biology, 195(7):1093–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
7/1093.

Liu:2011:MMS

[LR11b] Ling Liu and Thomas A. Rando. Manifestations and mechanisms of stem cell aging. Journal of Cell Biology, 193(2):

257–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/2/257.

Lerit:2013:PIA

[LR13] Dorothy A. Lerit and Nasser M. Rusan. PLP inhibits the activity of interphase centrosomes to ensure their proper segregation in stem cells. Journal of Cell Biology, 202(7):1013– ??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/7/1013.

Linkert:2010:MMA

[LRA^{+}10] Melissa Linkert, Curtis T. Rueden, Chris Allan, Jean-Marie Burel, Will Moore, Andrew Patterson, Brian Loranger, Josh Moore, Carlos Neves, Donald MacDonald, Aleksandra Tarkowska, Caitlin Sticco, Emma Hill, Mike Rossner, Kevin W. Eliceiri, and Jason R. Swedlow. Metadata matters: access to image data in the real world. Journal of Cell Biology, 189(5):777–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/5/777.

Lucic:2013:CET

[LRB13] Vladan Lučić, Alexander Rigort, and Wolfgang Baumeister. Cryo-electron tomography: The challenge of doing structural biology in situ. Journal of Cell Biology, 202(3):407–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
3/407.

Lafkas:2013:NMC

[LRH^{+}13] Daniel Lafkas, Veronica Rodilla, Mathilde Huyghe, Larissa Mourao, Hippokratis Kiaris, and Silvia Fre. Notch3 marks clonogenic mammary luminal progenitor cells in vivo. Journal of Cell Biology, 203(1):47–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/1/47.

Lawson:2013:PAK

[LS13a] Christine Lawson and David D. Schlaepfer. pHocal adhesion kinase regulation is on a FERM foundation. Journal of Cell Biology, 202(6):833–??, September 2013. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/202/6/833.

Listovsky:2013:SCM

[LS13b] Tamar Listovsky and Julian E. Sale. Sequestration of CDH1 by MAD2L2 prevents premature APC/C activation prior to anaphase onset. Journal of Cell Biology, 203(1):87–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
1/87.

Li:2011:GSR

[LSCF11] Fuchuan Li, Wen Shi, Mariana Capurro, and Jorge Filmus. Glypican-5 stimulates rhabdomyosarcoma cell proliferation by activating Hedgehog signaling. Journal of Cell Biology, 192(4): 691–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/4/691.

Liu:2010:TDS

[LSE^{+}10] Yuanyuan Liu, Claudia Schirra, Ludwig Edelmann, Ulf Matti, JeongSeop Rhee, Detlef Hof, Dieter Bruns, Nils Brose, Heiko Rieger, David R. Stevens, and Jens Rettig. Two distinct secretory vesicle–priming steps in adrenal chromaffin cells. Journal of Cell Biology, 190(6):1067–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/6/1067.

LeDreau:2014:SSA

[LSGVM14] Gwenvael Le Dréau, Murielle Saade, Irene Gutiérrez-Vallejo, and Elisa Martí. The strength of SMAD1/5 activity determines the mode of stem cell division in the developing spinal cord. Journal of Cell Biology, 204(4):591–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
4/591.

Lippi:2011:TAA

[LSM^{+}11] Giordano Lippi, Joern R. Steinert, Emma L. Marczylo, Sabina D’Oro, Roberto Fiore, Ian D. Forsythe, Gerhard Schratt, Michele Zoli, Pierluigi Nicotera, and Kenneth W. Young. Targeting of the Arpc3 actin nucleation factor by miR-29a/b regulates dendritic spine morphology. Journal of Cell Biology, 194(6):889–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/6/889.

Li:2010:NOL

[LSOT10] Gang Li, Christopher Scull, Lale Ozcan, and Ira Tabas. NADPH oxidase links endoplasmic reticulum stress, oxidative stress, and PKR activation to induce apoptosis. Journal of Cell Biology, 191(6):1113–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/6/1113.

Lo:2012:LRO

[LSS^{+}12] Bryan Lo, Geraldine Strasser, Meredith Sagolla, Cary D. Austin, Melissa Junttila, and Ira Mellman. Lkb1 regulates organogenesis and early oncogenesis along AMPK-dependent and -independent pathways. Journal of Cell Biology, 199(7): 1117–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/7/1117.

Lieu:2014:PIF

[LSW^{+}14] Kim G. Lieu, Eun-Hee Shim, Jinling Wang, Ravi K. Lokareddy, Tao Tao, Gino Cingolani, Gerard P. Zambetti, and David A. Jans. The p53-induced factor Ei24 inhibits nuclear import through an importin β-binding-like domain. Journal of Cell Biology, 205(3):301–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/205/3/301.

Lin:2011:RRP

[LT11] Jung-Chun Lin and Woan-Yuh Tarn. RBM4 down-regulates PTB and antagonizes its activity in muscle cell–specific alternative splicing. Journal of Cell Biology, 193(3):509–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
3/509.

Lee:2012:SSF

[LTJN^{+}12] Pei-Chih Lee, Kimberly M. Taylor-Jaffe, Kara M. Nordin, Maneeshi S. Prasad, Rachel M. Lander, and Carole LaBonne. SUMOylated SoxE factors recruit Grg4 and function as transcriptional repressors in the neural crest. Journal of Cell Biology, 198(5):799–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/5/799.

Liu:2010:RTP

[LVB^{+}10] Dan Liu, Mathijs Vleugel, Chelsea B. Backer, Tetsuya Hori, Tatsuo Fukagawa, Iain M. Cheeseman, and Michael A. Lampson. Regulated targeting of protein phosphatase 1 to the outer kinetochore by KNL1 opposes Aurora B kinase. Journal of Cell Biology, 188(6):809–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/6/809.

Luijsterburg:2010:SRA

[LvBG^{+}10] Martijn S. Luijsterburg, Gesa von Bornstaedt, Audrey M. Gourdin, Antonio Z. Politi, Martijn J. Moné, Daniël O. Warmerdam, Joachim Goedhart, Wim Vermeulen, Roel van Driel, and Thomas Höfer. Stochastic and reversible assembly of a multiprotein DNA repair complex ensures accurate target site recognition and efficient repair. Journal of Cell Biology, 189(3):445–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/3/445.

Lacroix:2010:TPS

[LvDG^{+}10] Benjamin Lacroix, Juliette van Dijk, Nicholas D. Gold, Julien Guizetti, Gudrun Aldrian-Herrada, Krzysztof Rogowski, Daniel W. Gerlich, and Carsten Janke. Tubulin polyglutamylation stimulates spastin-mediated microtubule severing. Journal of Cell Biology, 189(6):945–??, June 2010. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/189/6/
945.

Law:2013:LSW

[LVK^{+}13] Ah-Lai Law, Anne Vehlow, Maria Kotini, Lauren Dodgson, Daniel Soong, Eric Theveneau, Cristian Bodo, Eleanor Taylor, Christel Navarro, Upamali Perera, Magdalene Michael, Graham A. Dunn, Daimark Bennett, Roberto Mayor, and Matthias Krause. Lamellipodin and the Scar/WAVE complex cooperate to promote cell migration in vivo. Journal of Cell Biology, 203(4):673–??, November 2013. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/4/673.

Lochte:2014:LCM

[LWB^{+}14] Sara Löchte, Sharon Waichman, Oliver Beutel, Changjiang You, and Jacob Piehler. Live cell micropatterning reveals the dynamics of signaling complexes at the plasma membrane. Journal of Cell Biology, 207(3):407–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
3/407.

Lapointe:2012:SDU

[LWBH12] Jérôme Lapointe, Ying Wang, Eve Bigras, and Siegfried Hekimi. The submitochondrial distribution of ubiquinone affects respiration in long-lived Mclk1^{+/}− mice. Journal of Cell Biology, 199(2):215–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/2/215.

Lianga:2013:WCI

[LWK^{+}13] Noel Lianga, Elizabeth C. Williams, Erin K. Kennedy, Carole Doré, Sophie Pilon, Stéphanie L. Girard, Jean-Sebastien Deneault, and Adam D. Rudner. A Wee1 checkpoint inhibits anaphase onset. Journal of Cell Biology, 201(6):843–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
6/843.

Lin:2013:CIC

[LWL^{+}13] Yi-Nan Lin, Chien-Ting Wu, Yu-Chih Lin, Wen-Bin Hsu, Chieh-Ju C. Tang, Ching-Wen Chang, and Tang K. Tang. CEP120 interacts with CPAP and positively regulates centriole elongation. Journal of Cell Biology, 202(2):211–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
2/211.

Lu:2012:EMD

[LWW12] Pengfei Lu, Valerie M. Weaver, and Zena Werb. The extracellular matrix: a dynamic niche in cancer progression. Journal of Cell Biology, 196(4):395–??, February 2012. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/4/395.

Li:2010:SGA

[LWZ^{+}10] Yujie Li, Qing Wei, Yuxia Zhang, Kun Ling, and Jinghua Hu. The small GTPases ARL-13 and ARL-3 coordinate intraflagellar transport and ciliogenesis. Journal of Cell Biology, 189 (6):1039–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/6/1039.

Lutter:2012:SME

[LXTM12] Sophie Lutter, Sherry Xie, Florence Tatin, and Taija Makinen. Smooth muscle–endothelial cell communication activates Reelin signaling and regulates lymphatic vessel formation. Journal of Cell Biology, 197(6):837–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/6/837.

Long:2010:SAR

[LYB^{+}10] Kimberly R. Long, Yasunori Yamamoto, Adam L. Baker, Simon C. Watkins, Carolyn B. Coyne, James F. Conway, and Meir Aridor. Sar1 assembly regulates membrane constriction and ER export. Journal of Cell Biology, 190(1):115–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
1/115.

Lin:2013:SPE

[LYH^{+}13] Long Lin, Peiguo Yang, Xinxin Huang, Hui Zhang, Qun Lu, and Hong Zhang. The scaffold protein EPG-7 links cargo– receptor complexes with the autophagic assembly machinery. Journal of Cell Biology, 201(1):113–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/1/113.

Luo:2013:MPE

[LZLG13] Guangzuo Luo, Jian Zhang, Francis C. Luca, and Wei Guo. Mitotic phosphorylation of Exo84 disrupts exocyst assembly and arrests cell growth. Journal of Cell Biology, 202(1):97–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
202/1/97.

Li:2011:SCE

[LZR^{+}11] Xiao-Yan Li, Xiaoming Zhou, R. Grant Rowe, Yuexian Hu, David D. Schlaepfer, Dusko Ilić, Gregory Dressler, Ann Park, Jun-Lin Guan, and Stephen J. Weiss. Snail1 controls epithelial–mesenchymal lineage commitment in focal adhesion kinase–null embryonic cells. Journal of Cell Biology, 195(5): 729–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/5/729.

Li:2010:IBM

[LZW^{+}10] Dong Li, Jiaxi Zhou, Lu Wang, Myung Eun Shin, Pei Su, Xiaohua Lei, Haibin Kuang, Weixiang Guo, Hong Yang, Linzhao Cheng, Tetsuya S. Tanaka, Deborah E. Leckband, Albert B. Reynolds, Enkui Duan, and Fei Wang. Integrated biochemical and mechanical signals regulate multifaceted human embryonic stem cell functions. Journal of Cell Biology, 191(3): 631–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/3/631.

Li:2012:SSG

[LZW^{+}12] Yujie Li, Qing Zhang, Qing Wei, Yuxia Zhang, Kun Ling, and Jinghua Hu. SUMOylation of the small GTPase ARL-13 promotes ciliary targeting of sensory receptors. Journal of Cell Biology, 199(4):589–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/4/589.

Li:2013:SSG

[LZW^{+}13] Yujie Li, Qing Zhang, Qing Wei, Yuxia Zhang, Kun Ling, and Jinghua Hu. SUMOylation of the small GTPase ARL-13 promotes ciliary targeting of sensory receptors. Journal of Cell Biology, 200(3):357–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/3/357.

Li:2012:AGF

[LZY^{+}12] Wei Li, Wei Zou, Yihong Yang, Yongping Chai, Baohui Chen, Shiya Cheng, Dong Tian, Xiaochen Wang, Ronald D. Vale, and Guangshuo Ou. Autophagy genes function sequentially to promote apoptotic cell corpse degradation in the engulfing cell.

Journal of Cell Biology, 197(1):27–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/1/27.

Mukhopadhyay:2010:SPS

[MAD10] Debaditya Mukhopadhyay, Alexei Arnaoutov, and Mary Dasso. The SUMO protease SENP6 is essential for inner kinetochore assembly. Journal of Cell Biology, 188(5):681–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
188/5/681.

Myers:2011:DEM

[MAD^{+}11] Kenneth A. Myers, Kathryn T. Applegate, Gaudenz Danuser, Robert S. Fischer, and Clare M. Waterman. Distinct ECM mechanosensing pathways regulate microtubule dynamics to control endothelial cell branching morphogenesis. Journal of Cell Biology, 192(2):321–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/2/321.

Mitsushima:2010:RMD

[MAE^{+}10] Masaru Mitsushima, Kazuhiro Aoki, Miki Ebisuya, Shigeru Matsumura, Takuya Yamamoto, Michiyuki Matsuda, Fumiko Toyoshima, and Eisuke Nishida. Revolving movement of a dynamic cluster of actin filaments during mitosis. Journal of Cell Biology, 191(3):453–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/3/453.

Manjithaya:2010:USP

[MALS10] Ravi Manjithaya, Christophe Anjard, William F. Loomis, and Suresh Subramani. Unconventional secretion of Pichia pastoris Acb1 is dependent on GRASP protein, peroxisomal functions, and autophagosome formation. Journal of Cell Biology, 188 (4):537–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/4/537.

Ma:2011:PAM

[MAS11] Changle Ma, Gaurav Agrawal, and Suresh Subramani. Peroxisome assembly: matrix and membrane protein biogenesis. Journal of Cell Biology, 193(1):7–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/1/7.

Maldonado-Baez:2013:MDE

[MBCKD13] Lymarie Maldonado-Báez, Nelson B. Cole, Helmut Krämer, and Julie G. Donaldson. Microtubule-dependent endosomal sorting of clathrin-independent cargo by Hook1. Journal of Cell Biology, 201(2):233–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/201/2/233.

Mittal:2010:TFS

[MBK^{+}10] Ashwani Mittal, Shephali Bhatnagar, Akhilesh Kumar, Estelle Lach-Trifilieff, Sandrine Wauters, Hong Li, Denys Y. Makonchuk, David J. Glass, and Ashok Kumar. The TWEAK–Fn14 system is a critical regulator of denervationinduced skeletal muscle atrophy in mice. Journal of Cell Biology, 188(6):833–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/6/833.

Mittelmeier:2011:APC

[MBLD11] Telsa M. Mittelmeier, Joseph S. Boyd, Mary Rose Lamb, and Carol L. Dieckmann. Asymmetric properties of the Chlamydomonas reinhardtii cytoskeleton direct rhodopsin photoreceptor localization. Journal of Cell Biology, 193(4):741–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
4/741.

Mehlitz:2010:TRE

[MBM^{+}10] Adrian Mehlitz, Sebastian Banhart, André P. Mäurer, Alexis Kaushansky, Andrew G. Gordus, Julia Zielecki, Gavin MacBeath, and Thomas F. Meyer. Tarp regulates early Chlamydia-induced host cell survival through interactions with the human adaptor protein SHC1. Journal of Cell Biology, 190(1):143–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/1/143.

Morioka:2014:TKS

[MBO^{+}14] Sho Morioka, Peter Broglie, Emily Omori, Yuka Ikeda, Giichi Takaesu, Kunihiro Matsumoto, and Jun Ninomiya-Tsuji.

TAK1 kinase switches cell fate from apoptosis to necrosis following TNF stimulation. Journal of Cell Biology, 204(4): 607–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/4/607.

Ma:2011:HDS

[MBR^{+}11] Peisong Ma, Shannon L. Beck, Ronald W. Raab, Robert L. McKown, George L. Coffman, Atsushi Utani, William J. Chirico, Alan C. Rapraeger, and Gordon W. Laurie. Heparanase deglycanation of syndecan-1 is required for binding of the epithelial-restricted prosecretory mitogen lacritin. Journal of Cell Biology, 192(2):365–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/2/365.

Montresor:2013:JTK

[MBVT^{+}13] Alessio Montresor, Matteo Bolomini-Vittori, Lara Toffali, Barbara Rossi, Gabriela Constantin, and Carlo Laudanna. JAK tyrosine kinases promote hierarchical activation of Rho and Rap modules of integrin activation. Journal of Cell Biology, 203(6):1003–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/6/1003.

Mocciaro:2010:VCG

[MBZ^{+}10] Annamaria Mocciaro, Eli Berdougo, Kang Zeng, Elizabeth Black, Paola Vagnarelli, William Earnshaw, David Gillespie, Prasad Jallepalli, and Elmar Schiebel. Vertebrate cells genetically deficient for Cdc14A or Cdc14B retain DNA damage checkpoint proficiency but are impaired in DNA repair. Journal of Cell Biology, 189(4):631–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/4/631.

Mesnard:2010:IPC

[MC10] Daniel Mesnard and Daniel B. Constam. Imaging proprotein convertase activities and their regulation in the implanting mouse blastocyst. Journal of Cell Biology, 191(1):129–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
191/1/129.

Moore:2010:SPC

[MCHCC10] Jeffrey K. Moore, Prakash Chudalayandi, Richard A. Heil-Chapdelaine, and John A. Cooper. The spindle position checkpoint is coordinated by the Elm1 kinase. Journal of Cell Biology, 191(3):493–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/3/493.

McNally:2013:MSP

[McN13] Francis J. McNally. Mechanisms of spindle positioning. Journal of Cell Biology, 200(2):131–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/2/131.

Munoz:2013:ECW

[MCS^{+}13] Javier Muñoz, Juan Carlos G. Cortés, Matthias Sipiczki, Mariona Ramos, José Angel Clemente-Ramos, M. Belén Moreno, Ivone M. Martins, Pilar Pérez, and Juan Carlos Ribas. Extracellular cell wall β (1,3)glucan is required to couple septation to actomyosin ring contraction. Journal of Cell Biology, 203 (2):265–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/2/265.

Machicoane:2014:SDA

[MdFF^{+}14] Mickael Machicoane, Cristina A. de Frutos, Jenny Fink, Murielle Rocancourt, Yannis Lombardi, Sonia Garel, Matthieu Piel, and Arnaud Echard. SLK-dependent activation of ERMs controls LGN–NuMA localization and spindle orientation. Journal of Cell Biology, 205(6):791–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/205/6/791.

Mendez:2010:CMP

[MDP^{+}10] Pablo Mendez, Mathias De Roo, Lorenzo Poglia, Paul Klauser, and Dominique Muller. N-cadherin mediates plasticityinduced long-term spine stabilization. Journal of Cell Biology, 189(3):589–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/3/589.

Morgan:2013:BCS

[MDW^{+}13] Anthony J. Morgan, Lianne C. Davis, Siegfried K. T. Y. Wagner, Alexander M. Lewis, John Parrington, Grant C. Churchill, and Antony Galione. Bidirectional Ca^{2+} signaling occurs between the endoplasmic reticulum and acidic organelles. Journal of Cell Biology, 200(6):789–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/6/789.

Mellman:2013:NPM

[ME13] Ira Mellman and Scott D. Emr. A Nobel Prize for membrane traffic: Vesicles find their journey’s end. Journal of Cell Biology, 203(4):559–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/4/559.

Mejlvang:2014:NHS

[MFA^{+}14] Jakob Mejlvang, Yunpeng Feng, Constance Alabert, Kai J. Neelsen, Zuzana Jasencakova, Xiaobei Zhao, Michael Lees, Albin Sandelin, Philippe Pasero, Massimo Lopes, and Anja Groth. New histone supply regulates replication fork speed and PCNA unloading. Journal of Cell Biology, 204(1):29–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
204/1/29.

Martinez:2012:DCT

[MFB12] Paula Martínez, Juana M. Flores, and Maria A. Blasco. 53BP1 deficiency combined with telomere dysfunction activates ATRdependent DNA damage response. Journal of Cell Biology, 197 (2):283–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/2/283.

Minina:2013:AMD

[MFF^{+}13] Elena A. Minina, Lada H. Filonova, Kazutake Fukada, Eugene I. Savenkov, Vladimir Gogvadze, David Clapham, Victoria Sanchez-Vera, Maria F. Suarez, Boris Zhivotovsky, Geoffrey Daniel, Andrei Smertenko, and Peter V. Bozhkov. Autophagy and metacaspase determine the mode of cell death in plants. Journal of Cell Biology, 203(6):917–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
6/917.

Mandal:2010:MCG

[MFGB10] Sudip Mandal, William A. Freije, Preeta Guptan, and Utpal Banerjee. Metabolic control of G1–S transition: cyclin E degradation by p53-induced activation of the ubiquitin– proteasome system. Journal of Cell Biology, 188(4):473– ??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/4/473.

Fu:2013:JRD

[mFH13] Meng meng Fu and Erika L. F. Holzbaur. JIP1 regulates the directionality of APP axonal transport by coordinating kinesin and dynein motors. Journal of Cell Biology, 202(3):495–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
202/3/495.

Murphy:2014:PRR

[MFR^{+}14] Anar K. Murphy, Michael Fitzgerald, Teresa Ro, Jee Hyun Kim, Ariana I. Rabinowitsch, Dipanjan Chowdhury, Carl L. Schildkraut, and James A. Borowiec. Phosphorylated RPA recruits PALB2 to stalled DNA replication forks to facilitate fork recovery. Journal of Cell Biology, 206(4):493–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
4/493.

Martin:2010:ICF

[MGFG^{+}10] Adam C. Martin, Michael Gelbart, Rodrigo Fernandez-Gonzalez, Matthias Kaschube, and Eric F. Wieschaus. Integration of contractile forces during tissue invagination. Journal of Cell Biology, 188(5):735–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/5/735.

Mahuzier:2012:DSC

[MGG^{+}12] Alexia Mahuzier, Helori-Mael Gaudé, Valentina Grampa, Isabelle Anselme, Flora Silbermann, Margot Leroux-Berger, Delphine Delacour, Jerome Ezan, Mireille Montcouquiol, Sophie Saunier, Sylvie Schneider-Maunoury, and Christine Vesque. Dishevelled stabilization by the ciliopathy protein Rpgrip1l is essential for planar cell polarity. Journal of Cell Biology, 198(5):927–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/5/927.

Maia:2012:CPM

[MGK^{+}12] Ana R. R. Maia, Zaira Garcia, Lilian Kabeche, Marin Barisic, Stefano Maffini, Sandra Macedo-Ribeiro, Iain M. Cheeseman, Duane A. Compton, Irina Kaverina, and Helder Maiato. Cdk1 and Plk1 mediate a CLASP2 phospho-switch that stabilizes kinetochore–microtubule attachments. Journal of Cell Biology, 199(2):285–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/2/285.

Mari:2010:ACC

[MGR^{+}10] Muriel Mari, Janice Griffith, Ester Rieter, Lakshmi Krishnappa, Daniel J. Klionsky, and Fulvio Reggiori. An Atg9containing compartment that functions in the early steps of autophagosome biogenesis. Journal of Cell Biology, 190(6): 1005–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/6/1005.

Magiera:2014:DRS

[MGS14] Maria M. Magiera, Elisabeth Gueydon, and Etienne Schwob. DNA replication and spindle checkpoints cooperate during S phase to delay mitosis and preserve genome integrity. Journal of Cell Biology, 204(2):165–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/2/165.

Maciejowski:2010:MDA

[MGT^{+}10] John Maciejowski, Kelly A. George, Marie-Emilie Terret, Chao Zhang, Kevan M. Shokat, and Prasad V. Jallepalli. Mps1 directs the assembly of Cdc20 inhibitory complexes during interphase and mitosis to control M phase timing and spindle checkpoint signaling. Journal of Cell Biology, 190(1):89–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
190/1/89.

Maiden:2011:SLC

[MH11] Stephanie L. Maiden and Jeff Hardin. The secret life of αcatenin: Moonlighting in morphogenesis. Journal of Cell Biology, 195(4):543–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/4/543.

Morita:2014:AKR

[MH14] Tsuyoshi Morita and Ken’ichiro Hayashi. Arp5 is a key regulator of myocardin in smooth muscle cells. Journal of Cell Biology, 204(5):683–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/5/683.

Meyer:2012:PMP

[MHAK^{+}12] Aaron S. Meyer, Shannon K. Hughes-Alford, Jennifer E. Kay, Amalchi Castillo, Alan Wells, Frank B. Gertler, and Douglas A. Lauffenburger. 2d protrusion but not motility predicts growth factor–induced cancer cell migration in 3D collagen. Journal of Cell Biology, 197(6):721–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/6/721.

Marshall:2012:SDA

[MHC^{+}12] Jamie L. Marshall, Johan Holmberg, Eric Chou, Amber C. Ocampo, Jennifer Oh, Joy Lee, Angela K. Peter, Paul T. Martin, and Rachelle H. Crosbie-Watson. Sarcospan-dependent Akt activation is required for utrophin expression and muscle regeneration. Journal of Cell Biology, 197(7):1009–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
7/1009.

Morawe:2011:LEU

[MHCvSW11] Tobias Morawe, Mona Honemann-Capito, Walter von Stein, and Andreas Wodarz. Loss of the extraproteasomal ubiquitin receptor Rings lost impairs ring canal growth in Drosophila oogenesis. Journal of Cell Biology, 193(1):71–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/193/1/71.

Minami:2010:BES

[MHK^{+}10] Ryosuke Minami, Atsuko Hayakawa, Hiroki Kagawa, Yuko Yanagi, Hideyoshi Yokosawa, and Hiroyuki Kawahara. BAG-6 is essential for selective elimination of defective proteasomal substrates. Journal of Cell Biology, 190(4):637–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
4/637.

Matsumoto:2011:MPC

[MHKM11] Seiji Matsumoto, Motoshi Hayano, Yutaka Kanoh, and Hisao Masai. Multiple pathways can bypass the essential role of fission yeast Hsk1 kinase in DNA replication initiation. Journal of Cell Biology, 195(3):387–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/3/387.

Maskell:2010:MAA

[MHS10] Daniel P. Maskell, Xiao-Wen Hu, and Martin R. Singleton. Molecular architecture and assembly of the yeast kinetochore MIND complex. Journal of Cell Biology, 190(5):823– ??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/5/823.

Menoni:2012:NER

[MHV12] Hervé Menoni, Jan H. J. Hoeijmakers, and Wim Vermeulen. Nucleotide excision repair–initiating proteins bind to oxidative DNA lesions in vivo. Journal of Cell Biology, 199(7): 1037–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/7/1037.

Machlus:2013:IJM

[MI13a] Kellie R. Machlus and Joseph E. Italiano. The incredible journey: From megakaryocyte development to platelet formation. Journal of Cell Biology, 201(6):785–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/6/785.

Mohammadi:2013:CIE

[MI13b] Sina Mohammadi and Ralph R. Isberg. Cdc42 interacts with the exocyst complex to promote phagocytosis. Journal of Cell Biology, 200(1):81–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/1/81.

Misteli:2010:NEB

[Mis10] Tom Misteli. New editorial board members 2010. Journal of Cell Biology, 188(5):619–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/5/619.

Misteli:2013:EII

[Mis13] Tom Misteli. Eliminating the impact of the Impact Factor. Journal of Cell Biology, 201(5):651–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/5/651.

Mitch:2012:DKI

[Mit12a] Leslie Mitch. DGK-α keeps integrins close to the edge. Journal of Cell Biology, 196(2):??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/2/180.3.

Mitch:2012:LFL

[Mit12b] Leslie Mitch. Licensing factors lose their credentials. Journal of Cell Biology, 196(2):??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/2/180.1.

Mitch:2012:SMD

[Mit12c] Leslie Mitch. n-syb makes a dangerous delivery. Journal of Cell Biology, 196(2):??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/2/180.2.

Manneville:2010:DBG

[MJEM10] Jean-Baptiste Manneville, Muguette Jehanno, and Sandrine Etienne-Manneville. Dlg1 binds GKAP to control dynein association with microtubules, centrosome positioning, and cell polarity. Journal of Cell Biology, 191(3):585–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
3/585.

Manjithaya:2010:YMC

[MJFS10] Ravi Manjithaya, Shveta Jain, Jean-Claude Farré, and Suresh Subramani. A yeast MAPK cascade regulates pexophagy but not other autophagy pathways. Journal of Cell Biology, 189 (2):303–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/2/303.

Malureanu:2010:CHM

[MJJ^{+}10] Liviu Malureanu, Karthik B. Jeganathan, Fang Jin, Darren J. Baker, Janine H. van Ree, Oliver Gullon, Zheyan Chen, John R. Henley, and Jan M. van Deursen. Cdc20 hypomorphic mice fail to counteract de novo synthesis of cyclin B1 in mitosis. Journal of Cell Biology, 191(2):313–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
2/313.

Moyle:2014:BMI

[MKH^{+}14] Mark W. Moyle, Taekyung Kim, Neil Hattersley, Julien Espeut, Dhanya K. Cheerambathur, Karen Oegema, and Arshad Desai. A Bub1–Mad1 interaction targets the Mad1– Mad2 complex to unattached kinetochores to initiate the spindle checkpoint. Journal of Cell Biology, 204(5):647–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
5/647.

Mosammaparast:2013:HDL

[MKL^{+}13] Nima Mosammaparast, Haeyoung Kim, Benoit Laurent, Yu Zhao, Hui Jun Lim, Mona C. Majid, Sebastian Dango, Yuying Luo, Kristina Hempel, Mathew E. Sowa, Steven P. Gygi, Hanno Steen, J. Wade Harper, Bruce Yankner, and Yang Shi. The histone demethylase LSD1/KDM1A promotes the DNA damage response. Journal of Cell Biology, 203(3): 457–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/3/457.

Martinelli:2013:RCT

[MKS^{+}13] Roberta Martinelli, Masataka Kamei, Peter T. Sage, Ramiro Massol, Laya Varghese, Tracey Sciuto, Mourad Toporsian, Ann M. Dvorak, Tomas Kirchhausen, Timothy A. Springer, and Christopher V. Carman. Release of cellular tension signals self-restorative ventral lamellipodia to heal barrier micro-wounds. Journal of Cell Biology, 201(3):449–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
3/449.

Molli:2010:ACP

[MLBY^{+}10] Poonam R. Molli, Da-Qiang Li, Rozita Bagheri-Yarmand, Suresh B. Pakala, Hiroshi Katayama, Subrata Sen, Jyoti Iyer, Jonathan Chernoff, Ming-Ying Tsai, Sujit S. Nair, and Rakesh Kumar. Arpc1b, a centrosomal protein, is both an activator and substrate of Aurora A. Journal of Cell Biology, 190 (1):101–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/1/101.

Marchiando:2010:CDO

[MLG^{+}10] Amanda M. Marchiando, Le Shen, W. Vallen Graham, Christopher R. Weber, Brad T. Schwarz, Jotham R. Austin, David R. Raleigh, Yanfang Guan, Alastair J. M. Watson, Marshall H. Montrose, and Jerrold R. Turner. Caveolin-1-dependent occludin endocytosis is required for TNF-induced tight junction regulation in vivo. Journal of Cell Biology, 189 (1):111–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/1/111.

Magudia:2012:KRB

[MLH12] Kirti Magudia, Aurelia Lahoz, and Alan Hall. K-Ras and B– Raf oncogenes inhibit colon epithelial polarity establishment through up-regulation of c-myc. Journal of Cell Biology, 198 (2):185–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/2/185.

Magalhaes:2011:CPR

[MLM^{+}11] Marco A. O. Magalhaes, Daniel R. Larson, Christopher C. Mader, Jose Javier Bravo-Cordero, Hava Gil-Henn, Matthew Oser, Xiaoming Chen, Anthony J. Koleske, and John Condeelis. Cortactin phosphorylation regulates cell invasion through a pH-dependent pathway. Journal of Cell Biology, 195(5):903–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/5/903.

Madsen:2013:MLM

[MLM^{+}13] Daniel H. Madsen, Daniel Leonard, Andrius Masedunskas, Amanda Moyer, Henrik Jessen Jürgensen, Diane E. Peters, Panomwat Amornphimoltham, Arul Selvaraj, Susan S. Yamada, David A. Brenner, Sven Burgdorf, Lars H. Engelholm, Niels Behrendt, Kenn Holmbeck, Roberto Weigert, and Thomas H. Bugge. M2-like macrophages are responsible for collagen degradation through a mannose receptor-mediated pathway. Journal of Cell Biology, 202(6):951–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
6/951.

Morin-Leisk:2011:ISB

[MLSM^{+}11] Jeanne Morin-Leisk, Simran G. Saini, Xin Meng, Alexander M. Makhov, Peijun Zhang, and Tina H. Lee. An intramolecular salt bridge drives the soluble domain of GTP-bound atlastin into the postfusion conformation. Journal of Cell Biology, 195(4):605–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/4/605.

Makio:2013:IYN

[MLW13] Tadashi Makio, Diego L. Lapetina, and Richard W. Wozniak. Inheritance of yeast nuclear pore complexes requires the Nsp1p subcomplex. Journal of Cell Biology, 203(2):187–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
2/187.

Mettlen:2010:CAS

[MLY^{+}10] Marcel Mettlen, Dinah Loerke, Defne Yarar, Gaudenz Danuser, and Sandra L. Schmid. Cargo- and adaptor-specific mechanisms regulate clathrin-mediated endocytosis. Journal of Cell Biology, 188(6):919–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/6/919.

Morselli:2011:SRI

[MMB^{+}11] Eugenia Morselli, Guillermo Mariño, Martin V. Bennetzen, Tobias Eisenberg, Evgenia Megalou, Sabrina Schroeder, Sandra Cabrera, Paule Bénit, Pierre Rustin, Alfredo Criollo, Oliver Kepp, Lorenzo Galluzzi, Shensi Shen, Shoaib Ahmad Malik, Maria Chiara Maiuri, Yoshiyuki Horio, Carlos López-Otín, Jens S. Andersen, Nektarios Tavernarakis, Frank Madeo, and Guido Kroemer. Spermidine and resveratrol induce autophagy by distinct pathways converging on the acetylproteome. Journal of Cell Biology, 192(4):615–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
4/615.

Mitchell:2010:PLT

[MMC^{+}10] Jana M. Mitchell, Jörg Mansfeld, Juliana Capitanio, Ulrike Kutay, and Richard W. Wozniak. Pom121 links two essential subcomplexes of the nuclear pore complex core to the membrane. Journal of Cell Biology, 191(3):505–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
3/505.

Martinez-Morales:2011:FRA

[MMdCOM^{+}11] Patricia L. Martínez-Morales, Ruth Diez del Corral, Isabel Olivera-Martínez, Alejandra C. Quiroga, Raman M. Das, Julio A. Barbas, Kate G. Storey, and Aixa V. Morales. FGF and retinoic acid activity gradients control the timing of neural crest cell emigration in the trunk. Journal of Cell Biology, 194 (3):489–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/3/489.

Moree:2011:CCR

[MMFS11] Ben Moree, Corey B. Meyer, Colin J. Fuller, and Aaron F. Straight. CENP-c recruits M18BP1 to centromeres to promote CENP–A chromatin assembly. Journal of Cell Biology, 194(6): 855–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/6/855.

Mikawa:2014:SIS

[MMO^{+}14] Takumi Mikawa, Takeshi Maruyama, Koji Okamoto, Hitoshi Nakagama, Matilde E. Lleonart, Takeshi Tsusaka, Kousuke Hori, Itsuo Murakami, Taisuke Izumi, Akifumi Takaori-Kondo, Masayuki Yokode, Gordon Peters, David Beach, and Hiroshi Kondoh. Senescence-inducing stress promotes proteolysis of phosphoglycerate mutase via ubiquitin ligase Mdm2. Journal of Cell Biology, 204(5):729–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/5/729.

Matsunaga:2010:ARE

[MMS^{+}10] Kohichi Matsunaga, Eiji Morita, Tatsuya Saitoh, Shizuo Akira, Nicholas T. Ktistakis, Tetsuro Izumi, Takeshi Noda, and Tamotsu Yoshimori. Autophagy requires endoplasmic reticulum targeting of the PI3-kinase complex via Atg14L. Journal of Cell Biology, 190(4):511–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/4/511.

Mackay:2010:DNP

[MMU10a] Douglas R. Mackay, Masaki Makise, and Katharine S. Ullman. Defects in nuclear pore assembly lead to activation of an Aurora B-mediated abscission checkpoint. Journal of Cell Biology, 191(5):923–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/5/923.

Misaki:2010:PRP

[MMU^{+}10b] Ryo Misaki, Miki Morimatsu, Takefumi Uemura, Satoshi Waguri, Eiji Miyoshi, Naoyuki Taniguchi, Michiyuki Matsuda, and Tomohiko Taguchi. Palmitoylated Ras proteins traffic through recycling endosomes to the plasma membrane during exocytosis. Journal of Cell Biology, 191(1):23–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
1/23.

Manzoni:2010:OCR

[MMV^{+}10] Romilde Manzoni, Francesca Montani, Clara Visintin, Fabrice Caudron, Andrea Ciliberto, and Rosella Visintin. Oscillations in Cdc14 release and sequestration reveal a circuit underlying mitotic exit. Journal of Cell Biology, 190(2):209–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
2/209.

Mi-Mi:2012:ZLF

[MMVK^{+}12] Lei Mi-Mi, SarahBeth Votra, Kenneth Kemphues, Anthony Bretscher, and David Pruyne. Z-line formins promote contractile lattice growth and maintenance in striated muscles of C. elegans. Journal of Cell Biology, 198(1):87–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/198/1/87.

McIntosh:2013:CDF

[MOZ^{+}13] J. Richard McIntosh, Eileen O’Toole, Kirill Zhudenkov, Mary Morphew, Cindi Schwartz, Fazly I. Ataullakhanov, and Ekaterina L. Grishchuk. Conserved and divergent features of kinetochores and spindle microtubule ends from five species. Journal of Cell Biology, 200(4):459–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/4/459.

Martina:2013:RGM

[MP13] Jose A. Martina and Rosa Puertollano. Rag GTPases mediate amino acid-dependent recruitment of TFEB and MITF to lysosomes. Journal of Cell Biology, 200(4):475–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/200/
4/475.

Marzi:2012:DAM

[MPD^{+}12] Matteo J. Marzi, Eleonora M. R. Puggioni, Valentina Dall’Olio, Gabriele Bucci, Loris Bernard, Fabrizio Bianchi, Marco Crescenzi, Pier Paolo Di Fiore, and Francesco Nicassio. Differentiation-associated microRNAs antagonize the Rb–E2F pathway to restrict proliferation. Journal of Cell Biology, 199 (1):77–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/1/77.

Monette:2011:HRN

[MPM11] Anne Monette, Nelly Panté, and Andrew J. Mouland. HIV-1 remodels the nuclear pore complex. Journal of Cell Biology, 193(4):619–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/4/619.

Muslimov:2011:SCR

[MPRT11] Ilham A. Muslimov, Mihir V. Patel, Arthur Rose, and Henri Tiedge. Spatial code recognition in neuronal RNA targeting: Role of RNA–hnRNP A2 interactions. Journal of Cell Biology, 194(3):441–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/3/441.

Mast:2014:SCB

[MRA14] Fred D. Mast, Alexander V. Ratushny, and John D. Aitchison. Systems cell biology. Journal of Cell Biology, 206(6): 695–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/6/695.

Monteiro:2013:EWE

[MRCC^{+}13] Pedro Monteiro, Carine Rossé, Antonio Castro-Castro, Marie Irondelle, Emilie Lagoutte, Perrine Paul-Gilloteaux, Claire Desnos, Etienne Formstecher, François Darchen, David Perrais, Alexis Gautreau, Maud Hertzog, and Philippe Chavrier. Endosomal WASH and exocyst complexes control exocytosis of MT1-MMP at invadopodia. Journal of Cell Biology, 203(6): 1063–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/6/1063.

Mitra:2012:DDM

[MRLLS12] Kasturi Mitra, Richa Rikhy, Mary Lilly, and Jennifer Lippincott-Schwartz. DRP1-dependent mitochondrial fission initiates follicle cell differentiation during Drosophila oogenesis. Journal of Cell Biology, 197(4):487–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/4/487.

Moreau:2012:APA

[MRPR12] Kevin Moreau, Brinda Ravikumar, Claudia Puri, and David C. Rubinsztein. Arf6 promotes autophagosome formation via effects on phosphatidylinositol 4,5-bisphosphate and phospholipase D. Journal of Cell Biology, 196(4):483–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
4/483.

Moutin:2012:DRS

[MRR^{+}12] Enora Moutin, Fabrice Raynaud, Jonathan Roger, Emilie Pellegrino, Vincent Homburger, Federica Bertaso, Vincent Ollendorff, Joël Bockaert, Laurent Fagni, and Julie Perroy. Dynamic remodeling of scaffold interactions in dendritic spines controls synaptic excitability. Journal of Cell Biology, 198(2):251–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
198/2/251.

Mardin:2012:BTB

[MS12] Balca R. Mardin and Elmar Schiebel. Breaking the ties that bind: New advances in centrosome biology. Journal of Cell Biology, 197(1):11–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/1/11.

Matson:2014:CAB

[MS14] Daniel R. Matson and P. Todd Stukenberg. CENP-I and Aurora B act as a molecular switch that ties RZZ/Mad1 recruitment to kinetochore attachment status. Journal of Cell Biology, 205(4):541–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/4/541.

Masszi:2010:FDM

[MSC^{+}10] András Masszi, Pam Speight, Emmanuel Charbonney, Monika Lodyga, Hiroyasu Nakano, Katalin Szászi, and András Kapus. Fate-determining mechanisms in epithelial–myofibroblast transition: major inhibitory role for Smad3. Journal of Cell Biology, 188(3):383–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/188/3/383.

Matkovic:2013:BCD

[MSK^{+}13a] Tanja Matkovic, Matthias Siebert, Elena Knoche, Harald Depner, Sara Mertel, David Owald, Manuela Schmidt, Ulrich Thomas, Albert Sickmann, Dirk Kamin, Stefan W. Hell, Jörg Bürger, Christina Hollmann, Thorsten Mielke, Carolin Wichmann, and Stephan J. Sigrist. The Bruchpilot cytomatrix determines the size of the readily releasable pool of synaptic vesicles. Journal of Cell Biology, 202(4):667–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
4/667.

Mirouse:2013:LAM

[MSK^{+}13b] Vincent Mirouse, Lance L. Swick, Nevzat Kazgan, Daniel St Johnston, and Jay E. Brenman. LKB1 and AMPK maintain epithelial cell polarity under energetic stress. Journal of Cell Biology, 203(2):373–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/2/373.

Mangmool:2010:ADA

[MSR10] Supachoke Mangmool, Arun K. Shukla, and Howard A. Rockman. β-arrestin-dependent activation of Ca^{2+} /calmodulin kinase II after β_{1} –adrenergic receptor stimulation. Journal of Cell Biology, 189(3):573–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/3/573.

Matsuda:2010:PSM

[MSS^{+}10] Noriyuki Matsuda, Shigeto Sato, Kahori Shiba, Kei Okatsu, Keiko Saisho, Clement A. Gautier, Yu shin Sou, Shinji Saiki, Sumihiro Kawajiri, Fumiaki Sato, Mayumi Kimura, Masaaki Komatsu, Nobutaka Hattori, and Keiji Tanaka. PINK1 stabilized by mitochondrial depolarization recruits Parkin to damaged mitochondria and activates latent Parkin for mitophagy. Journal of Cell Biology, 189(2):211–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/2/211.

Moreira:2012:SCE

[MSS^{+}12] Karen E. Moreira, Sebastian Schuck, Bianca Schrul, Florian Fröhlich, James B. Moseley, Tobias C. Walther, and Peter Walter. Seg1 controls eisosome assembly and shape. Journal of Cell Biology, 198(3):405–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/3/405.

Morimoto:2012:CKD

[MSZ^{+}12] Akihiro Morimoto, Hiroki Shibuya, Xiaoqiang Zhu, Jihye Kim, Kei ichiro Ishiguro, Min Han, and Yoshinori Watanabe. A conserved KASH domain protein associates with telomeres, SUN1, and dynactin during mammalian meiosis. Journal of Cell Biology, 198(2):165–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/198/2/165.

Ma:2011:TRL

[MTG^{+}11] Nan Ma, Janel Titus, Alyssa Gable, Jennifer L. Ross, and Patricia Wadsworth. TPX2 regulates the localization and activity of Eg5 in the mammalian mitotic spindle. Journal of Cell Biology, 195(1):87–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/1/87.

McNees:2010:AST

[MTM^{+}10] Carolyn J. McNees, Agueda M. Tejera, Paula Martínez, Matilde Murga, Francisca Mulero, Oscar Fernandez-Capetillo, and Maria A. Blasco. ATR suppresses telomere fragility and recombination but is dispensable for elongation of short telomeres by telomerase. Journal of Cell Biology, 188(5):639–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
188/5/639.

Muslimov:2014:INM

[MTT^{+}14] Ilham A. Muslimov, Aliya Tuzhilin, Thean Hock Tang, Robert K. S. Wong, Riccardo Bianchi, and Henri Tiedge. Interactions of noncanonical motifs with hnRNP A2 promote activitydependent RNA transport in neurons. Journal of Cell Biology, 205(4):493–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/4/493.

Massone:2011:RPI

[MVC^{+}11] Sara Massone, Irene Vassallo, Manuele Castelnuovo, Gloria Fiorino, Elena Gatta, Mauro Robello, Roberta Borghi, Massimo Tabaton, Claudio Russo, Giorgio Dieci, Ranieri Cancedda, and Aldo Pagano. RNA polymerase III drives alternative splicing of the potassium channel–interacting protein contributing to brain complexity and neurodegeneration. Journal of Cell Biology, 193(5):851–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/5/851.

Muller:2011:PMT

[MVN11] Patricia A. J. Muller, Karen H. Vousden, and Jim C. Norman. p53 and its mutants in tumor cell migration and invasion. Journal of Cell Biology, 192(2):209–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/2/209.

Mick:2010:CCE

[MVP^{+}10] David U. Mick, Milena Vukotic, Heike Piechura, Helmut E. Meyer, Bettina Warscheid, Markus Deckers, and Peter Rehling. Coa3 and Cox14 are essential for negative feedback regulation of COX1 translation in mitochondria. Journal of Cell Biology, 191(1):141–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/1/141.

Mai:2011:CBR

[MVP^{+}11] Anja Mai, Stefan Veltel, Teijo Pellinen, Artur Padzik, Eleanor Coffey, Varpu Marjomäki, and Johanna Ivaska. Competitive binding of Rab21 and p120RasGAP to integrins regulates receptor traffic and migration. Journal of Cell Biology, 194 (2):291–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/2/291.

Maghzal:2010:TAE

[MVR^{+}10] Nadim Maghzal, Emily Vogt, Wolfgang Reintsch, James S. Fraser, and François Fagotto. The tumor-associated EpCAM regulates morphogenetic movements through intracellular signaling. Journal of Cell Biology, 191(3):645–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
3/645.

Mall:2012:MLD

[MWG^{+}12] Moritz Mall, Thomas Walter, Mátyás Gorjánácz, Iain F. Davidson, Thi Bach Nga Ly-Hartig, Jan Ellenberg, and Iain W. Mattaj. Mitotic lamin disassembly is triggered by lipid-mediated signaling. Journal of Cell Biology, 198(6):981– ??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/6/981.

Maday:2012:AID

[MWH12] Sandra Maday, Karen E. Wallace, and Erika L. F. Holzbaur. Autophagosomes initiate distally and mature during transport toward the cell soma in primary neurons. Journal of Cell Biology, 196(4):407–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/4/407.

Makhnevych:2012:HRS

[MWP^{+}12] Taras Makhnevych, Philip Wong, Oxana Pogoutse, Franco J. Vizeacoumar, Jack F. Greenblatt, Andrew Emili, and Walid A. Houry. Hsp110 is required for spindle length control. Journal of Cell Biology, 198(4):623–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/4/623.

Mao:2011:TMS

[MWZ^{+}11] Kai Mao, Ke Wang, Mantong Zhao, Tao Xu, and Daniel J. Klionsky. Two MAPK-signaling pathways are required for mitophagy in Saccharomyces cerevisiae. Journal of Cell Biology, 193(4):755–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/4/755.

Mahjoub:2010:CAL

[MXS10] Moe R. Mahjoub, Zhigang Xie, and Tim Stearns. Cep120 is asymmetrically localized to the daughter centriole and is essential for centriole assembly. Journal of Cell Biology, 191 (2):331–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/2/331.

Montembault:2010:NDN

[MZP^{+}10] Emilie Montembault, Wei Zhang, Marcin R. Przewloka, Vincent Archambault, Emeric W. Sevin, Ernest D. Laue, David M. Glover, and Pier Paolo D’Avino. Nessun Dorma, a novel centralspindlin partner, is required for cytokinesis in Drosophila spermatocytes. Journal of Cell Biology, 191(7):1351–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
191/7/1351.

Nance:2014:GKY

[Nan14] Jeremy Nance. Getting to know your neighbor: Cell polarization in early embryos. Journal of Cell Biology, 206(7): 823–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/7/823.

Nekrasova:2011:DCU

[NAS^{+}11] Oxana E. Nekrasova, Evangeline V. Amargo, William O. Smith, Jing Chen, Geri E. Kreitzer, and Kathleen J. Green. Desmosomal cadherins utilize distinct kinesins for assembly into desmosomes. Journal of Cell Biology, 195(7):1185–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
195/7/1185.

Nekrasova:2012:DCU

[NAS^{+}12] Oxana E. Nekrasova, Evangeline V. Amargo, William O. Smith, Jing Chen, Geri E. Kreitzer, and Kathleen J. Green. Desmosomal cadherins utilize distinct kinesins for assembly into desmosomes. Journal of Cell Biology, 196(2):297–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
196/2/297.

Nekrasova:2013:DCU

[NAS^{+}13] Oxana E. Nekrasova, Evangeline V. Amargo, William O. Smith, Jing Chen, Geri E. Kreitzer, and Kathleen J. Green. Desmosomal cadherins utilize distinct kinesins for assembly into desmosomes. Journal of Cell Biology, 201(7):1085–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
201/7/1085.

Nguyen:2010:HEP

[NB10] Hai Dang Nguyen and Anja-Katrin Bielinsky. HDM2 ERKs PCNA. Journal of Cell Biology, 190(4):487–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
4/487.

Nechipurenko:2012:FLM

[NB12] Inna V. Nechipurenko and Heather T. Broihier. FoxO limits microtubule stability and is itself negatively regulated by microtubule disruption. Journal of Cell Biology, 196(3): 345–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/3/345.

Nakatsu:2012:PSP

[NBC^{+}12] Fubito Nakatsu, Jeremy M. Baskin, Jeeyun Chung, Lukas B. Tanner, Guanghou Shui, Sang Yoon Lee, Michelle Pirruccello, Mingming Hao, Nicholas T. Ingolia, Markus R. Wenk, and Pietro De Camilli. PtdIns4P synthesis by PI4KIIIα at the plasma membrane and its impact on plasma membrane identity. Journal of Cell Biology, 199(6):1003–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
6/1003.

Ng:2012:SSR

[NBDB12] Mei Rosa Ng, Achim Besser, Gaudenz Danuser, and Joan S. Brugge. Substrate stiffness regulates cadherin-dependent collective migration through myosin–II contractility. Journal of Cell Biology, 199(3):545–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/3/545.

Novak:2011:WRM

[NBS^{+}11] Nurit Novak, Vered Bar, Helena Sabanay, Shahar Frechter, Martine Jaegle, Scott B. Snapper, Dies Meijer, and Elior Peles. N-WASP is required for membrane wrapping and myelination by Schwann cells. Journal of Cell Biology, 192(2):243–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
192/2/243.

Nalepa:2013:TSCa

[NBSE^{+}13a] Grzegorz Nalepa, Jill Barnholtz-Sloan, Rikki Enzor, Dilip Dey, Ying He, Jeff R. Gehlhausen, Amalia S. Lehmann, Su-Jung Park, Yanzhu Yang, Xianlin Yang, Shi Chen, Xiaowei Guan, Yanwen Chen, Jamie Renbarger, Feng-Chun Yang, Luis F. Parada, and Wade Clapp. The tumor suppressor CDKN3 controls mitosis. Journal of Cell Biology, 201(7):997–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
7/997.

Nalepa:2013:TSCb

[NBSE^{+}13b] Grzegorz Nalepa, Jill Barnholtz-Sloan, Rikki Enzor, Dilip Dey, Ying He, Jeff R. Gehlhausen, Amalia S. Lehmann, Su-Jung Park, Yanzhu Yang, Xianlin Yang, Shi Chen, Xiaowei Guan, Yanwen Chen, Jamie Renbarger, Feng-Chun Yang, Luis F. Parada, and Wade Clapp. The tumor suppressor CDKN3 controls mitosis. Journal of Cell Biology, 202(4):717–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
4/717.

Nanes:2012:PCB

[NCML^{+}12] Benjamin A. Nanes, Christine Chiasson-MacKenzie, Anthony M. Lowery, Noboru Ishiyama, Victor Faundez, Mitsuhiko Ikura, Peter A. Vincent, and Andrew P. Kowalczyk. p120-catenin binding masks an endocytic signal conserved in classical cadherins. Journal of Cell Biology, 199(2):365–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
199/2/365.

Nakajima:2011:IAD

[NCT^{+}11] Yuko Nakajima, Anthony Cormier, Randall G. Tyers, Adrianne Pigula, Yutian Peng, David G. Drubin, and Georjana Barnes. Ipl1/ Aurora-dependent phosphorylation of Sli15/ INCENP regulates CPC–spindle interaction to ensure proper microtubule dynamics. Journal of Cell Biology, 194(1):137–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
194/1/137.

Nola:2011:ARR

[NDS^{+}11] Sébastien Nola, Reiko Daigaku, Kasia Smolarczyk, Maryke Carstens, Belen Martin-Martin, Gregory Longmore, Maryse Bailly, and Vania M. M. Braga. Ajuba is required for Rac activation and maintenance of E-cadherin adhesion. Journal of Cell Biology, 195(5):855–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/5/855.

Nayak:2010:TRA

[NEMH^{+}10] Tania Nayak, Heather Edgerton-Morgan, Tetsuya Horio, Yi Xiong, Colin P. De Souza, Stephen A. Osmani, and Berl R. Oakley. γ-tubulin regulates the anaphase-promoting complex/ cyclosome during interphase. Journal of Cell Biology, 190(3): 317–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/3/317.

Ngok:2012:VAE

[NGL^{+}12] Siu P. Ngok, Rory Geyer, Miaoliang Liu, Antonis Kourtidis, Sudesh Agrawal, Chuanshen Wu, Himabindu Reddy Seerapu, Laura J. Lewis-Tuffin, Karen L. Moodie, Deborah Huveldt, Ruth Marx, Jay M. Baraban, Peter Storz, Arie Horowitz, and Panos Z. Anastasiadis. VEGF and Angiopoietin-1 exert opposing effects on cell junctions by regulating the Rho GEF Syx. Journal of Cell Biology, 199(7):1103–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
7/1103.

Noatynska:2012:MSD

[NGM12] Anna Noatynska, Monica Gotta, and Patrick Meraldi. Mitotic spindle (DIS)orientation and DISease: Cause or consequence? Journal of Cell Biology, 199(7):1025–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
7/1025.

Nievergall:2010:PRE

[NJS^{+}10] Eva Nievergall, Peter W. Janes, Carolin Stegmayer, Mary E. Vail, Fawaz G. Haj, Shyh Wei Teng, Benjamin G. Neel, Philippe I. Bastiaens, and Martin Lackmann. PTP1B regulates Eph receptor function and trafficking. Journal of Cell Biology, 191(6):1189–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/6/1189.

Ninomiya:2011:SRM

[NKH11] Kensuke Ninomiya, Naoyuki Kataoka, and Masatoshi Hagiwara. Stress-responsive maturation of Clk1/4 pre-mRNAs promotes phosphorylation of SR splicing factor. Journal of Cell Biology, 195(1):27–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/1/27.

Nistala:2010:FDM

[NLAS^{+}10] Harikiran Nistala, Sui Lee-Arteaga, Silvia Smaldone, Gabriella Siciliano, Luca Carta, Robert N. Ono, Gerhard Sengle, Emilio Arteaga-Solis, Regis Levasseur, Patricia Ducy, Lynn Y. Sakai, Gerard Karsenty, and Francesco Ramirez. Fibrillin-1 and -2 differentially modulate endogenous TGF-β and BMP bioavailability during bone formation. Journal of Cell Biology, 190(6): 1107–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/6/1107.

Nair:2013:NRN

[NLJ^{+}13] Ramya Nair, Juliane Lauks, SangYong Jung, Nancy E. Cooke, Heidi de Wit, Nils Brose, Manfred W. Kilimann, Matthijs Verhage, and JeongSeop Rhee. Neurobeachin regulates neurotransmitter receptor trafficking to synapses. Journal of Cell Biology, 200(1):61–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/1/61.

Nahm:2010:CSG

[NLP^{+}10] Minyeop Nahm, A. Ashleigh Long, Sang Kyoo Paik, Sungdae Kim, Yong Chul Bae, Kendal Broadie, and Seungbok Lee. The Cdc42-selective GAP Rich regulates postsynaptic development and retrograde BMP transsynaptic signaling. Journal of Cell Biology, 191(3):661–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/3/661.

Nabti:2014:DMR

[NMB^{+}14] Ibtissem Nabti, Petros Marangos, Jenny Bormann, Nobuaki R. Kudo, and John Carroll. Dual-mode regulation of the APC/C by CDK1 and MAPK controls meiosis I progression and fidelity. Journal of Cell Biology, 204(6):891–??, March 2014.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/204/6/
891.

Nakata:2011:PBK

[NNO^{+}11] Takao Nakata, Shinsuke Niwa, Yasushi Okada, Franck Perez, and Nobutaka Hirokawa. Preferential binding of a kinesin-1 motor to GTP-tubulin–rich microtubules underlies polarized vesicle transport. Journal of Cell Biology, 194(2):245–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
2/245.

Nakagawa:2011:PSS

[NNSH11] Shinichi Nakagawa, Takao Naganuma, Go Shioi, and Tetsuro Hirose. Paraspeckles are subpopulation-specific nuclear bodies that are not essential in mice. Journal of Cell Biology, 193 (1):31–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/1/31.

Ninagawa:2014:EIM

[NOS^{+}14] Satoshi Ninagawa, Tetsuya Okada, Yoshiki Sumitomo, Yukiko Kamiya, Koichi Kato, Satoshi Horimoto, Tokiro Ishikawa, Shunichi Takeda, Tetsushi Sakuma, Takashi Yamamoto, and Kazutoshi Mori. EDEM2 initiates mammalian glycoprotein ERAD by catalyzing the first mannose trimming step. Journal of Cell Biology, 206(3):347–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/3/347.

Nazarko:2014:PAB

[NOT^{+}14] Taras Y. Nazarko, Katharine Ozeki, Andreas Till, Geetha Ramakrishnan, Pouya Lotfi, Mingda Yan, and Suresh Subramani. Peroxisomal Atg37 binds Atg30 or palmitoyl–CoA to regulate phagophore formation during pexophagy. Journal of Cell Biology, 204(4):541–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/4/541.

Nakatsu:2010:IPS

[NPL^{+}10] Fubito Nakatsu, Rushika M. Perera, Louise Lucast, Roberto Zoncu, Jan Domin, Frank B. Gertler, Derek Toomre, and Pietro De Camilli. The inositol 5-phosphatase SHIP2 regulates endocytic clathrin-coated pit dynamics. Journal of Cell Biology, 190(3):307–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/3/307.

Nishimura:2013:BFS

[NRK^{+}13] Hirohito Nishimura, Ken Ritchie, Rinshi S. Kasai, Miki Goto, Nobuhiro Morone, Hiroyuki Sugimura, Koichiro Tanaka, Ichiro Sase, Akihiko Yoshimura, Yoshitaro Nakano, Takahiro K. Fujiwara, and Akihiro Kusumi. Biocompatible fluorescent silicon nanocrystals for single-molecule tracking and fluorescence imaging. Journal of Cell Biology, 202(6):967–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
6/967.

Nagae:2012:CSI

[NRM^{+}12] Masamichi Nagae, Suyong Re, Emiko Mihara, Terukazu Nogi, Yuji Sugita, and Junichi Takagi. Crystal structure of α5β1 integrin ectodomain: Atomic details of the fibronectin receptor. Journal of Cell Biology, 197(1):131–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/1/131.

Nalesso:2011:WMA

[NSB^{+}11] Giovanna Nalesso, Joanna Sherwood, Jessica Bertrand, Thomas Pap, Manoj Ramachandran, Cosimo De Bari, Costantino Pitzalis, and Francesco Dell’Accio. WNT-3A modulates articular chondrocyte phenotype by activating both canonical and noncanonical pathways. Journal of Cell Biology, 193(3):551–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/3/551.

Noree:2010:INF

[NSBW10] Chalongrat Noree, Brian K. Sato, Risa M. Broyer, and James E. Wilhelm. Identification of novel filament-forming proteins in Saccharomyces cerevisiae and Drosophila melanogaster. Journal of Cell Biology, 190(4):541–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/4/541.

Norris:2014:MMA

[NSD^{+}14] Stephen R. Norris, Virupakshi Soppina, Aslan S. Dizaji, Kristin I. Schimert, David Sept, Dawen Cai, Sivaraj Sivaramakrishnan, and Kristen J. Verhey. A method for multiprotein assembly in cells reveals independent action of kinesins in complex. Journal of Cell Biology, 207(3):393–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
3/393.

Nezis:2010:ADD

[NSS^{+}10] Ioannis P. Nezis, Bhupendra V. Shravage, Antonia P. Sagona, Trond Lamark, Geir Bjørkøy, Terje Johansen, Tor Erik Rusten, Andreas Brech, Eric H. Baehrecke, and Harald Stenmark. Autophagic degradation of dBruce controls DNA fragmentation in nurse cells during late Drosophila melanogaster oogenesis. Journal of Cell Biology, 190(4):523–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
4/523.

Nakaya:2013:EIR

[NSS13] Yukiko Nakaya, Erike W. Sukowati, and Guojun Sheng. Epiblast integrity requires CLASP and Dystroglycan-mediated microtubule anchoring to the basal cortex. Journal of Cell Biology, 202(4):637–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/4/637.

Neisch:2010:RRA

[NSSF10] Amanda L. Neisch, Olga Speck, Beth Stronach, and Richard G. Fehon. Rho1 regulates apoptosis via activation of the JNK signaling pathway at the plasma membrane. Journal of Cell Biology, 189(2):311–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/2/311.

Niessen:2013:ACE

[NSZ^{+}13] Michaela T. Niessen, Jeanie Scott, Julia G. Zielinski, Susanne Vorhagen, Panagiota A. Sotiropoulou, Cédric Blanpain, Michael Leitges, and Carien M. Niessen. aPKCλ controls epidermal homeostasis and stem cell fate through regulation of division orientation. Journal of Cell Biology, 202(6):887– ??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/6/887.

Nakajima:2011:LRC

[NT11] Hiroyuki Nakajima and Takuji Tanoue. Lulu2 regulates the circumferential actomyosin tensile system in epithelial cells through p114RhoGEF. Journal of Cell Biology, 195(2):245– ??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/2/245.

Nandi:2014:AIA

[NTSK14] Nilay Nandi, Lauren K. Tyra, Drew Stenesen, and Helmut Krämer. Acinus integrates AKT1 and subapoptotic caspase activities to regulate basal autophagy. Journal of Cell Biology, 207(2):253–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/2/253.

Nijenhuis:2013:TDC

[NvCL^{+}13] Wilco Nijenhuis, Eleonore von Castelmur, Dene Littler, Valeria De Marco, Eelco Tromer, Mathijs Vleugel, Maria H. J. van Osch, Berend Snel, Anastassis Perrakis, and Geert J. P. L. Kops. A TPR domain–containing N-terminal module of MPS1 is required for its kinetochore localization by Aurora B. Journal of Cell Biology, 201(2):217–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/2/217.

Nightingale:2011:AIC

[NWD^{+}11] Thomas D. Nightingale, Ian J. White, Emily L. Doyle, Mark Turmaine, Kimberly J. Harrison-Lavoie, Kathleen F. Webb, Louise P. Cramer, and Daniel F. Cutler. Actomyosin II contractility expels von Willebrand factor from Weibel–Palade bodies during exocytosis. Journal of Cell Biology, 194(4): 613–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/4/613.

Neelsen:2013:OIG

[NZHL13] Kai J. Neelsen, Isabella M. Y. Zanini, Raquel Herrador, and Massimo Lopes. Oncogenes induce genotoxic stress by mitotic processing of unusual replication intermediates. Journal of Cell Biology, 200(6):699–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/6/699.

Ouellet:2012:OSD

[OB12] Jimmy Ouellet and Yves Barral. Organelle segregation during mitosis: Lessons from asymmetrically dividing cells. Journal of Cell Biology, 196(3):305–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/3/305.

Oas:2014:ARF

[OBC14] Sandy T. Oas, Anton L. Bryantsev, and Richard M. Cripps. Arrest is a regulator of fiber-specific alternative splicing in the indirect flight muscles of Drosophila. Journal of Cell Biology, 206(7):895–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/7/895.

Okada:2010:APC

[OBD^{+}10] Kyoko Okada, Francesca Bartolini, Alexandra M. Deaconescu, James B. Moseley, Zvonimir Dogic, Nikolaus Grigorieff, Gregg G. Gundersen, and Bruce L. Goode. Adenomatous polyposis coli protein nucleates actin assembly and synergizes with the formin mDia1. Journal of Cell Biology, 189(7):1087–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
189/7/1087.

Ortiz:2010:SHR

[OBM^{+}10] Julio O. Ortiz, Florian Brandt, Valério R. F. Matias, Lau Sennels, Juri Rappsilber, Sjors H. W. Scheres, Matthias Eibauer, F. Ulrich Hartl, and Wolfgang Baumeister. Structure of hibernating ribosomes studied by cryoelectron tomography in vitro and in situ. Journal of Cell Biology, 190(4):613–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
4/613.

Oakes:2012:TRS

[OBSG12] Patrick W. Oakes, Yvonne Beckham, Jonathan Stricker, and Margaret L. Gardel. Tension is required but not sufficient for focal adhesion maturation without a stress fiber template. Journal of Cell Biology, 196(3):363–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
3/363.

ODonohue:2010:FDR

[OCF^{+}10] Marie-Françoise O’Donohue, Valérie Choesmel, Marlène Faubladier, Gwennaële Fichant, and Pierre-Emmanuel Gleizes. Functional dichotomy of ribosomal proteins during the synthesis of mammalian 40S ribosomal subunits. Journal of Cell Biology, 190(5):853–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/190/5/853.

Okreglak:2010:LAR

[OD10] Voytek Okreglak and David G. Drubin. Loss of Aip1 reveals a role in maintaining the actin monomer pool and an in vivo oligomer assembly pathway. Journal of Cell Biology, 188(6): 769–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/6/769.

Oeffinger:2010:JIS

[Oef10] Marlene Oeffinger. Joining the interface: a site for Nmd3 association on 60S ribosome subunits. Journal of Cell Biology, 189 (7):1071–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/7/1071.

Owald:2010:SHR

[OFS^{+}10] David Owald, Wernher Fouquet, Manuela Schmidt, Carolin Wichmann, Sara Mertel, Harald Depner, Frauke Christiansen, Christina Zube, Christine Quentin, Jorg Körner, Henning Urlaub, Karl Mechtler, and Stephan J. Sigrist. A Syd-1 homologue regulates pre- and postsynaptic maturation in Drosophila. Journal of Cell Biology, 188(4):565–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
4/565.

Olson:2012:HAE

[OGD^{+}12] Sara K. Olson, Garrett Greenan, Arshad Desai, Thomas Müller-Reichert, and Karen Oegema. Hierarchical assembly of the eggshell and permeability barrier in C. elegans. Journal of Cell Biology, 198(4):731–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/4/731.

Oh:2010:WMC

[OHC10] Jeong Su Oh, Seung Jin Han, and Marco Conti. Wee1B, Myt1, and Cdc25 function in distinct compartments of the mouse oocyte to control meiotic resumption. Journal of Cell Biology, 188(2):199–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/2/199.

Han:2012:MPE

[oHXK^{+}12] Sang oh Han, Kunhong Xiao, Jihee Kim, Jiao-Hui Wu, James W. Wisler, Nobuhiro Nakamura, Neil J. Freedman, and Sudha K. Shenoy. MARCH2 promotes endocytosis and lysosomal sorting of carvedilol-bound β_{2}-adrenergic receptors. Journal of Cell Biology, 199(5):817–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/5/817.

Okamoto:2014:OEC

[Oka14] Koji Okamoto. Organellophagy: Eliminating cellular building blocks via selective autophagy. Journal of Cell Biology, 205 (4):435–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/4/435.

Onishi:2013:DRR

[OKNP13] Masayuki Onishi, Nolan Ko, Ryuichi Nishihama, and John R. Pringle. Distinct roles of Rho1, Cdc42, and Cyk3 in septum formation and abscission during yeast cytokinesis. Journal of Cell Biology, 202(2):311–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/202/2/311.

Okiyoneda:2012:FCF

[OL12] Tsukasa Okiyoneda and Gergely L. Lukacs. Fixing cystic fibrosis by correcting CFTR domain assembly. Journal of Cell Biology, 199(2):199–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/2/199.

Ophir:2013:TSE

[OLB13] Michael J. Ophir, Beiyun C. Liu, and Stephen C. Bunnell. The N terminus of SKAP55 enables T cell adhesion to TCR and integrin ligands via distinct mechanisms. Journal of Cell Biology, 203(6):1021–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/6/1021.

Osorio:2011:RMA

[OLT11] Karen M. Osorio, Karin C. Lilja, and Tudorita Tumbar. Runx1 modulates adult hair follicle stem cell emergence and maintenance from distinct embryonic skin compartments. Journal of Cell Biology, 193(1):235–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/1/235.

Omran:2010:NPG

[Omr10] Heymut Omran. NPHP proteins: gatekeepers of the ciliary compartment. Journal of Cell Biology, 190(5):715–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
190/5/715.

Overmeer:2011:RPS

[OMV^{+}11] René M. Overmeer, Jill Moser, Marcel Volker, Hanneke Kool, Alan E. Tomkinson, Albert A. van Zeeland, Leon H. F. Mullenders, and Maria Fousteri. Replication protein A safeguards genome integrity by controlling NER incision events. Journal of Cell Biology, 192(3):401–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/3/401.

Okumura:2014:CBC

[OMW^{+}14] Eiichi Okumura, Atsushi Morita, Mizuho Wakai, Satoru Mochida, Masatoshi Hara, and Takeo Kishimoto. Cyclin B– Cdk1 inhibits protein phosphatase PP2A–B55 via a Greatwall kinase–independent mechanism. Journal of Cell Biology, 204 (6):881–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/6/881.

Obri:2014:HIP

[OMZK14] Arnaud Obri, Munevver Parla Makinistoglu, Hong Zhang, and Gerard Karsenty. HDAC4 integrates PTH and sympathetic signaling in osteoblasts. Journal of Cell Biology, 205 (6):771–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/6/771.

Onodera:2012:RPA

[ONH^{+}12] Yasuhito Onodera, Jin-Min Nam, Ari Hashimoto, Jim C. Norman, Hiroki Shirato, Shigeru Hashimoto, and Hisataka Sabe. Rab5c promotes AMAP1–PRKD2 complex formation to enhance β1 integrin recycling in EGF-induced cancer invasion. Journal of Cell Biology, 197(7):983–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/7/983.

Ossareh-Nazari:2014:ULE

[ONNB^{+}14] Batool Ossareh-Nazari, Carlos A. Niño, Mario H. Bengtson, Joong-Won Lee, Claudio A. P. Joazeiro, and Catherine Dargemont. Ubiquitylation by the Ltn1 E3 ligase protects 60S ribosomes from starvation-induced selective autophagy. Journal of Cell Biology, 204(6):909–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/6/909.

Oikawa:2012:TDF

[OOKH^{+}12] Tsukasa Oikawa, Masaaki Oyama, Hiroko Kozuka-Hata, Shunsuke Uehara, Nobuyuki Udagawa, Hideyuki Saya, and Koichi Matsuo. Tks5-dependent formation of circumferential podosomes/invadopodia mediates cell–cell fusion. Journal of Cell Biology, 197(4):553–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/4/553.

Osmani:2010:CLC

[OPCEM10] Naël Osmani, Florent Peglion, Philippe Chavrier, and Sandrine Etienne-Manneville. Cdc42 localization and cell polarity depend on membrane traffic. Journal of Cell Biology, 191(7): 1261–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/7/1261.

Orlandi:2012:GRG

[OPM^{+}12] Cesare Orlandi, Ekaterina Posokhova, Ikuo Masuho, Thomas A. Ray, Nazarul Hasan, Ronald G. Gregg, and Kirill A. Martemyanov. GPR158/ 179 regulate G protein signaling by controlling localization and activity of the RGS7 complexes. Journal of Cell Biology, 197(6):711–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/197/6/711.

Orr:2012:CBS

[Orr12] Harry T. Orr. Cell biology of spinocerebellar ataxia. Journal of Cell Biology, 197(2):167–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/197/2/167.

Ochocki:2013:NSP

[OS13] Joshua D. Ochocki and M. Celeste Simon. Nutrient-sensing pathways and metabolic regulation in stem cells. Journal of Cell Biology, 203(1):23–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/1/23.

Olsen:2014:GUB

[OSD^{+}14] Jessica M. Olsen, Masaaki Sato, Olof S. Dallner, Anna L. Sandström, Didier F. Pisani, Jean-Claude Chambard, Ez-Zoubir Amri, Dana S. Hutchinson, and Tore Bengtsson. Glucose uptake in brown fat cells is dependent on mTOR complex 2–promoted GLUT1 translocation. Journal of Cell Biology, 207(3):365–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/3/365.

Oda:2011:SFD

[OT11] Hiroki Oda and Masatoshi Takeichi. Structural and functional diversity of cadherin at the adherens junction. Journal of Cell Biology, 193(7):1137–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/7/1137.

Ong:2010:PEP

[OTLH10] Yan Shan Ong, Bor Luen Tang, Li Shen Loo, and Wanjin Hong. p125a exists as part of the mammalian Sec13/Sec31 COPII subcomplex to facilitate ER–Golgi transport. Journal of Cell Biology, 190(3):331–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/3/331.

Osman:2011:MHT

[OVL11] Christof Osman, Dennis R. Voelker, and Thomas Langer. Making heads or tails of phospholipids in mitochondria. Journal of Cell Biology, 192(1):7–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/1/7.

Ohkawa:2010:ABM

[OVW10] Taro Ohkawa, Loy E. Volkman, and Matthew D. Welch. Actinbased motility drives baculovirus transit to the nucleus and cell surface. Journal of Cell Biology, 190(2):187–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
2/187.

Otera:2010:MEF

[OWC^{+}10] Hidenori Otera, Chunxin Wang, Megan M. Cleland, Kiyoko Setoguchi, Sadaki Yokota, Richard J. Youle, and Katsuyoshi Mihara. Mff is an essential factor for mitochondrial recruitment of Drp1 during mitochondrial fission in mammalian cells. Journal of Cell Biology, 191(6):1141–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
6/1141.

Ono:2013:CII

[OYH13] Takao Ono, Daisuke Yamashita, and Tatsuya Hirano. Condensin II initiates sister chromatid resolution during S phase. Journal of Cell Biology, 200(4):429–??, February 2013. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/200/4/

Oda:2014:MBC

[OYYK14] Toshiyuki Oda, Haruaki Yanagisawa, Toshiki Yagi, and Masahide Kikkawa. Mechanosignaling between central apparatus and radial spokes controls axonemal dynein activity. Journal of Cell Biology, 204(5):807–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/5/807.

Ogun:2014:CAM

[OZ14] Oluwatobi Ogun and Marisa Zallocchi. Clarin-1 acts as a modulator of mechanotransduction activity and presynaptic ribbon assembly. Journal of Cell Biology, 207(3):375–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
3/375.

Oddoux:2013:MFS

[OZT^{+}13] Sarah Oddoux, Kristien J. Zaal, Victoria Tate, Aster Kenea, Shuktika A. Nandkeolyar, Ericka Reid, Wenhua Liu, and Evelyn Ralston. Microtubules that form the stationary lattice of muscle fibers are dynamic and nucleated at Golgi elements. Journal of Cell Biology, 203(2):205–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/2/205.

Pankiv:2010:FRE

[PAB^{+}10] Serhiy Pankiv, Endalkachew A. Alemu, Andreas Brech, Jack-Ansgar Bruun, Trond Lamark, Aud Øvervatn, Geir Bjørkøy, and Terje Johansen. FYCO1 is a Rab7 effector that binds to LC3 and PI3P to mediate microtubule plus end–directed vesicle transport. Journal of Cell Biology, 188(2):253–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
2/253.

Petsalaki:2011:PSR

[PAB^{+}11] Eleni Petsalaki, Tonia Akoumianaki, Elizabeth J. Black, David A. F. Gillespie, and George Zachos. Phosphorylation at serine 331 is required for Aurora B activation. Journal of Cell Biology, 195(3):449–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/3/449.

Pallanck:2010:CSM

[Pal10] Leo J. Pallanck. Culling sick mitochondria from the herd. Journal of Cell Biology, 191(7):1225–??, December 2010. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/191/7/
1225.

Palacios:2014:HHP

[Pal14] Isabel M. Palacios. Hop-on hop-off: Polysomes take a tour of the cell on endosomes. Journal of Cell Biology, 204(3): 287–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/3/287.

Papior:2012:OCS

[PASG^{+}12] Peer Papior, José M. Arteaga-Salas, Thomas Günther, Adam Grundhoff, and Aloys Schepers. Open chromatin structures regulate the efficiencies of pre–RC formation and replication initiation in Epstein–Barr virus. Journal of Cell Biology, 198 (4):509–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/4/509.

Pisconti:2013:FIN

[PBD^{+}13] Addolarata Pisconti, Silvia Brunelli, Monica Di Padova, Clara De Palma, Daniela Deponti, Silvia Baesso, Vittorio Sartorelli, Giulio Cossu, and Emilio Clementi. Follistatin induction by nitric oxide through cyclic GMP: a tightly regulated signaling pathway that controls myoblast fusion. Journal of Cell Biology, 200(3):359–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/3/359.

Prashar:2013:FMB

[PBG^{+}13] Akriti Prashar, Sonam Bhatia, Darren Gigliozzi, Tonya Martin, Carla Duncan, Cyril Guyard, and Mauricio R. Terebiznik. Filamentous morphology of bacteria delays the timing of phagosome morphogenesis in macrophages. Journal of Cell Biology, 203(6):1081–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/6/1081.

Pigino:2011:CTR

[PBM^{+}11] Gaia Pigino, Khanh Huy Bui, Aditi Maheshwari, Pietro Lupetti, Dennis Diener, and Takashi Ishikawa. Cryoelectron tomography of radial spokes in cilia and flagella. Journal of Cell Biology, 195(4):673–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/4/673.

Pichlo:2014:HDLa

[PBPW^{+}14a] Magdalena Pichlo, Stefanie Bungert-Plümke, Ingo Weyand, Reinhard Seifert, Wolfgang Bönigk, Timo Strünker, Nachiket Dilip Kashikar, Normann Goodwin, Astrid Müller, Heinz G. Körschen, Ursel Collienne, Patric Pelzer, Qui Van, Jörg Enderlein, Clementine Klemm, Eberhard Krause, Christian Trötschel, Ansgar Poetsch, Elisabeth Kremmer, and U. Benjamin Kaupp. High density and ligand affinity confer ultrasensitive signal detection by a guanylyl cyclase chemoreceptor. Journal of Cell Biology, 206(4):541–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/4/541.

Pichlo:2014:HDLb

[PBPW^{+}14b] Magdalena Pichlo, Stefanie Bungert-Plümke, Ingo Weyand, Reinhard Seifert, Wolfgang Bönigk, Timo Strünker, Nachiket Dilip Kashikar, Normann Goodwin, Astrid Müller, Heinz G. Körschen, Ursel Collienne, Patric Pelzer, Qui Van, Jörg Enderlein, Clementine Klemm, Eberhard Krause, Christian Trötschel, Ansgar Poetsch, Elisabeth Kremmer, and U. Benjamin Kaupp. High density and ligand affinity confer ultrasensitive signal detection by a guanylyl cyclase chemoreceptor. Journal of Cell Biology, 207(5):675–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
5/675.

Platt:2012:LSD

[PBvdS12] Frances M. Platt, Barry Boland, and Aarnoud C. van der Spoel. Lysosomal storage disorders: The cellular impact of lysosomal dysfunction. Journal of Cell Biology, 199(5):723– ??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/5/723.

Pan:2011:FRA

[PCC11] Yi-Ru Pan, Chien-Lin Chen, and Hong-Chen Chen. FAK is required for the assembly of podosome rosettes. Journal of Cell Biology, 195(1):113–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/1/113.

Poukkula:2011:CBR

[PCCR11] Minna Poukkula, Adam Cliffe, Rishita Changede, and Pernille Rørth. Cell behaviors regulated by guidance cues in collective migration of border cells. Journal of Cell Biology, 192(3): 513–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/3/513.

Pisconti:2010:SNC

[PCO^{+}10] Addolorata Pisconti, D. D. W. Cornelison, Hugo C. Olguín, Tiffany L. Antwine, and Bradley B. Olwin. Syndecan-3 and Notch cooperate in regulating adult myogenesis. Journal of Cell Biology, 190(3):427–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/190/3/427.

Patel:2014:DMD

[PDKG14] Dipal M. Patel, Adi D. Dubash, Geri Kreitzer, and Kathleen J. Green. Disease mutations in desmoplakin inhibit Cx43 membrane targeting mediated by desmoplakin–EB1 interactions. Journal of Cell Biology, 206(6):779–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
6/779.

Prosser:2011:ENC

[PDMBW11] Derek C. Prosser, Theodore G. Drivas, Lymarie Maldonado-Báez, and Beverly Wendland. Existence of a novel clathrinindependent endocytic pathway in yeast that depends on Rho1 and formin. Journal of Cell Biology, 195(4):657–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
4/657.

Pfeffer:2010:USA

[Pfe10] Suzanne R. Pfeffer. Unconventional secretion by autophagosome exocytosis. Journal of Cell Biology, 188(4):451–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/188/4/451.

Poea-Guyon:2013:VAM

[PGAE^{+}13] Sandrine Poëa-Guyon, Mohamed Raafet Ammar, Marie Erard, Muriel Amar, Alexandre W. Moreau, Philippe Fossier, Vincent Gleize, Nicolas Vitale, and Nicolas Morel. The V– ATPase membrane domain is a sensor of granular pH that controls the exocytotic machinery. Journal of Cell Biology, 203 (2):283–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/2/283.

Paul:2010:TAT

[PGB^{+}10] Pradyut K. Paul, Sanjay K. Gupta, Shephali Bhatnagar, Siva K. Panguluri, Bryant G. Darnay, Yongwon Choi, and Ashok Kumar. Targeted ablation of TRAF6 inhibits skeletal muscle wasting in mice. Journal of Cell Biology, 191(7): 1395–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/7/1395.

Petrie:2012:NSR

[PGCY12] Ryan J. Petrie, Núria Gavara, Richard S. Chadwick, and Kenneth M. Yamada. Nonpolarized signaling reveals two distinct modes of 3D cell migration. Journal of Cell Biology, 197 (3):439–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/3/439.

Pol:2014:BML

[PGP14] Albert Pol, Steven P. Gross, and Robert G. Parton. Biogenesis of the multifunctional lipid droplet: Lipids, proteins, and sites. Journal of Cell Biology, 204(5):635–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/5/635.

Parton:2010:RCT

[PH10] Robert G. Parton and Mark T. Howes. Revisiting caveolin trafficking: the end of the caveosome. Journal of Cell Biology, 191(3):439–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/3/439.

Parton:2011:PDO

[PHB^{+}11] Richard M. Parton, Russell S. Hamilton, Graeme Ball, Lei Yang, C. Fiona Cullen, Weiping Lu, Hiroyuki Ohkura, and Ilan Davis. A PAR-1-dependent orientation gradient of dynamic microtubules directs posterior cargo transport in the Drosophila oocyte. Journal of Cell Biology, 194(1):121–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
194/1/121.

Peters:2013:FDP

[PHB^{+}13] Lee Zeev Peters, Rotem Hazan, Michal Breker, Maya Schuldiner, and Shay Ben-Aroya. Formation and dissociation of proteasome storage granules are regulated by cytosolic pH. Journal of Cell Biology, 201(5):663–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/5/663.

Palacios:2010:SCT

[PHD^{+}10] Jose A. Palacios, Daniel Herranz, Maria Luigia De Bonis, Susana Velasco, Manuel Serrano, and Maria A. Blasco. SIRT1 contributes to telomere maintenance and augments global homologous recombination. Journal of Cell Biology, 191(7): 1299–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/7/1299.

Poulsen:2013:RAS

[PHW^{+}13] Sara L. Poulsen, Rebecca K. Hansen, Sebastian A. Wagner, Loes van Cuijk, Gijsbert J. van Belle, Werner Streicher, Mats Wikström, Chunaram Choudhary, Adriaan B. Houtsmuller, Jurgen A. Marteijn, Simon Bekker-Jensen, and Niels Mailand. RNF111/Arkadia is a SUMO-targeted ubiquitin ligase that facilitates the DNA damage response. Journal of Cell Biology, 201(6):797–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/6/797.

Peyre:2011:LBC

[PJS^{+}11] Elise Peyre, Florence Jaouen, Mehdi Saadaoui, Laurence Haren, Andreas Merdes, Pascale Durbec, and Xavier Morin. A lateral belt of cortical LGN and NuMA guides mitotic spindle movements and planar division in neuroepithelial cells. Journal of Cell Biology, 193(1):141–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/1/141.

Papic:2011:MMO

[PKD^{+}11] Dražen Papić, Katrin Krumpe, Jovana Dukanovic, Kai S. Dimmer, and Doron Rapaport. Multispan mitochondrial outer membrane protein Ugo1 follows a unique Mim1-dependent import pathway. Journal of Cell Biology, 194(3):397–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
3/397.

Pappas:2010:NRA

[PKG10] Christopher T. Pappas, Paul A. Krieg, and Carol C. Gregorio. Nebulin regulates actin filament lengths by a stabilization mechanism. Journal of Cell Biology, 189(5):859–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
5/859.

Posch:2010:SRA

[PKS^{+}10] Markus Posch, Guennadi A. Khoudoli, Sam Swift, Emma M. King, Jennifer G. DeLuca, and Jason R. Swedlow. Sds22 regulates aurora B activity and microtubule–kinetochore interactions at mitosis. Journal of Cell Biology, 191(1):61–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
1/61.

Poulson:2010:RCM

[PL10] Nicholas D. Poulson and Terry Lechler. Robust control of mitotic spindle orientation in the developing epidermis. Journal of Cell Biology, 191(5):915–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/5/915.

Park:2011:MDM

[PL11] Chan Bae Park and Nils-Göran Larsson. Mitochondrial DNA mutations in disease and aging. Journal of Cell Biology, 193 (5):809–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/5/809.

Peterson:2011:CUC

[PLC^{+}11] Shaun E. Peterson, Yinyin Li, Brian T. Chait, Max E. Gottesman, Richard Baer, and Jean Gautier. Cdk1 uncouples CtIPdependent resection and Rad51 filament formation during Mphase double-strand break repair. Journal of Cell Biology, 194(5):705–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/5/705.

Poulsen:2012:HRN

[PLL^{+}12] Maria Poulsen, Claudia Lukas, Jiri Lukas, Simon Bekker-Jensen, and Niels Mailand. Human RNF169 is a negative regulator of the ubiquitin-dependent response to DNA doublestrand breaks. Journal of Cell Biology, 197(2):189–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
2/189.

Peter:2013:BCP

[PLR^{+}13] Arun T. John Peter, Jens Lachmann, Meenakshi Rana, Madeleine Bunge, Margarita Cabrera, and Christian Ungermann. The BLOC-1 complex promotes endosomal maturation by recruiting the Rab5 GTPase-activating protein Msb3. Journal of Cell Biology, 201(1):97–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/1/97.

Primorac:2013:PRA

[PM13] Ivana Primorac and Andrea Musacchio. Panta rhei: The APC/C at steady state. Journal of Cell Biology, 201(2): 177–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/2/177.

Pranke:2011:SAM

[PMB^{+}11] Iwona M. Pranke, Vincent Morello, Joëlle Bigay, Kimberley Gibson, Jean-Marc Verbavatz, Bruno Antonny, and Catherine L. Jackson. α-synuclein and ALPS motifs are membrane curvature sensors whose contrasting chemistry mediates selective vesicle binding. Journal of Cell Biology, 194(1):89–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
1/89.

Perez-Moreno:2012:GAW

[PMF12] Mirna Perez-Moreno and Elaine Fuchs. Guilt by association: What p120-catenin has to hide. Journal of Cell Biology, 199 (2):211–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/2/211.

Papayannopoulos:2010:NEM

[PMHZ10] Venizelos Papayannopoulos, Kathleen D. Metzler, Abdul Hakkim, and Arturo Zychlinsky. Neutrophil elastase and myeloperoxidase regulate the formation of neutrophil extracellular traps. Journal of Cell Biology, 191(3):677–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
3/677.

Prota:2013:SBT

[PMK^{+}13] Andrea E. Prota, Maria M. Magiera, Marijn Kuijpers, Katja Bargsten, Daniel Frey, Mara Wieser, Rolf Jaussi, Casper C. Hoogenraad, Richard A. Kammerer, Carsten Janke, and Michel O. Steinmetz. Structural basis of tubulin tyrosination by tubulin tyrosine ligase. Journal of Cell Biology, 200(3): 259–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/200/3/259.

Prada:2011:RNGa

[PMP^{+}11a] Ilaria Prada, Julie Marchaland, Paola Podini, Lorenzo Magrassi, Rosalba D’Alessandro, Paola Bezzi, and Jacopo Meldolesi. REST/ NRSF governs the expression of dense-core vesicle gliosecretion in astrocytes. Journal of Cell Biology, 193 (3):537–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/3/537.

Prada:2011:RNGb

[PMP^{+}11b] Ilaria Prada, Julie Marchaland, Paola Podini, Lorenzo Magrassi, Rosalba D’Alessandro, Paola Bezzi, and Jacopo Meldolesi. REST/ NRSF governs the expression of dense-core vesicle gliosecretion in astrocytes. Journal of Cell Biology, 194 (3):505–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/3/505.

Park:2013:LBA

[PoLC^{+}13] Inai Park, Hae ock Lee, Eunhee Choi, Yoo-Kyung Lee, Mi-Sun Kwon, Jaewon Min, Pil-Gu Park, Seonju Lee, Young-Yun Kong, Gyungyub Gong, and Hyunsook Lee. Loss of BubR1 acetylation causes defects in spindle assembly checkpoint signaling and promotes tumor formation. Journal of Cell Biology, 202(2):295–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/2/295.

Powell:2010:PPF

[Pow10] Kendall Powell. Pier Paolo Di Fiore: Plumbing the depths of endocytosis. Journal of Cell Biology, 188(1):4–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
1/4.

Powell:2014:DFR

[Pow14a] Kendall Powell. Dan Fletcher: a recipe for cooking up cellular machines. Journal of Cell Biology, 207(3):320–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
3/320.

Powell:2014:DBP

[Pow14b] Kendall Powell. Dominique Bergmann: Passionate about plant polarity. Journal of Cell Biology, 207(6):680–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
6/680.

Petrovic:2010:MCP

[PPD^{+}10] Arsen Petrovic, Sebastiano Pasqualato, Prakash Dube, Veronica Krenn, Stefano Santaguida, Davide Cittaro, Silvia Monzani, Lucia Massimiliano, Jenny Keller, Aldo Tarricone, Alessio Maiolica, Holger Stark, and Andrea Musacchio. The MIS12 complex is a protein interaction hub for outer kinetochore assembly. Journal of Cell Biology, 190(5):835– ??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/5/835.

Plotnikova:2011:AKAa

[PPG11a] Olga V. Plotnikova, Elena N. Pugacheva, and Erica A. Golemis. Aurora A kinase activity influences calcium signaling in kidney cells. Journal of Cell Biology, 193(6):1021–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
6/1021.

Plotnikova:2011:AKAb

[PPG11b] Olga V. Plotnikova, Elena N. Pugacheva, and Erica A. Golemis. Aurora A kinase activity influences calcium signaling in kidney cells. Journal of Cell Biology, 194(1):157–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
1/157.

Pinon:2014:TBN

[PPV^{+}14] Perrine Pinon, Jenita Pärssinen, Patricia Vazquez, Michael Bachmann, Rolle Rahikainen, Marie-Claude Jacquier, Latifeh Azizi, Juha A. Määttä, Martin Bastmeyer, Vesa P. Hytönen, and Bernhard Wehrle-Haller. Talin-bound NPLY motif recruits integrin-signaling adapters to regulate cell spreading and mechanosensing. Journal of Cell Biology, 205(2):265–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/205/
2/265.

Park:2012:BPT

[PR12] Eunyong Park and Tom A. Rapoport. Bacterial protein translocation requires only one copy of the SecY complex in vivo. Journal of Cell Biology, 198(5):881–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
5/881.

Petit:2013:RFL

[PRFF13] Constance S. Petit, Agnes Roczniak-Ferguson, and Shawn M. Ferguson. Recruitment of folliculin to lysosomes supports the amino acid-dependent activation of Rag GTPases. Journal of Cell Biology, 202(7):1107–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/7/1107.

Prinz:2014:BGM

[Pri14] William A. Prinz. Bridging the gap: Membrane contact sites in signaling, metabolism, and organelle dynamics. Journal of Cell Biology, 205(6):759–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/205/6/759.

Puustinen:2014:COC

[PRM^{+}14] Pietri Puustinen, Anna Rytter, Monika Mortensen, Pekka Kohonen, José M. Moreira, and Marja Jäättelä. CIP2A oncoprotein controls cell growth and autophagy through mTORC1 activation. Journal of Cell Biology, 204(5):713–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
5/713.

Puttagunta:2011:RRC

[PSF^{+}11] Radhika Puttagunta, André Schmandke, Elisa Floriddia, Perrine Gaub, Natalie Fomin, Norbert B. Ghyselinck, and Simone Di Giovanni. RA-RAR-β counteracts myelin-dependent inhibition of neurite outgrowth via Lingo-1 repression. Journal of Cell Biology, 193(7):1147–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/7/1147.

Pocha:2011:CRR

[PSK11] Shirin Meher Pocha, Anna Shevchenko, and Elisabeth Knust. Crumbs regulates rhodopsin transport by interacting with and stabilizing myosin V. Journal of Cell Biology, 195(5):827– ??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/5/827.

Pasapera:2010:MIA

[PSR^{+}10] Ana M. Pasapera, Ian C. Schneider, Erin Rericha, David D. Schlaepfer, and Clare M. Waterman. Myosin II activity regulates vinculin recruitment to focal adhesions through FAKmediated paxillin phosphorylation. Journal of Cell Biology, 188(6):877–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/6/877.

Perdiguero:2011:PMR

[PSVRB^{+}11] Eusebio Perdiguero, Pedro Sousa-Victor, Vanessa Ruiz-Bonilla, Mercè Jardí, Carme Caelles, Antonio L. Serrano, and Pura Muñoz-Cánoves. p38/MKP-1-regulated AKT coordinates macrophage transitions and resolution of inflammation during tissue repair. Journal of Cell Biology, 195(2): 307–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/2/307.

Pitaval:2010:CSC

[PTBT10] Amandine Pitaval, Qingzong Tseng, Michel Bornens, and Manuel Théry. Cell shape and contractility regulate ciliogenesis in cell cycle–arrested cells. Journal of Cell Biology, 191 (2):303–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/2/303.

Pearse:2010:RUG

[PTS^{+}10] Bradley R. Pearse, Taku Tamura, Johan C. Sunryd, Gregory A. Grabowski, Randal J. Kaufman, and Daniel N. Hebert. The role of UDP–Glc: glycoprotein glucosyltransferase 1 in the maturation of an obligate substrate prosaposin. Journal of Cell Biology, 189(5):829–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/5/829.

Pignatelli:2012:HPI

[PTST12] Jeanine Pignatelli, David A. Tumbarello, Ronald P. Schmidt, and Christopher E. Turner. Hic-5 promotes invadopodia formation and invasion during TGF-β-induced epithelial– mesenchymal transition. Journal of Cell Biology, 197(3): 421–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/3/421.

Pfanner:2014:UNM

[PvdLA^{+}14] Nikolaus Pfanner, Martin van der Laan, Paolo Amati, Roderick A. Capaldi, Amy A. Caudy, Agnieszka Chacinska, Manjula Darshi, Markus Deckers, Suzanne Hoppins, Tateo Icho, Stefan Jakobs, Jianguo Ji, Vera Kozjak-Pavlovic, Chris Meisinger, Paul R. Odgren, Sang Ki Park, Peter Rehling, Andreas S. Reichert, M. Saeed Sheikh, Susan S. Taylor, Nobuo Tsuchida, Alexander M. van der Bliek, Ida J. van der Klei, Jonathan S. Weissman, Benedikt Westermann, Jiping Zha, Walter Neupert, and Jodi Nunnari. Uniform nomenclature for the mitochondrial contact site and cristae organizing system. Journal of Cell Biology, 204(7):1083–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/7/1083.

Pines:2012:PPN

[PVM^{+}12] Alex Pines, Mischa G. Vrouwe, Jurgen A. Marteijn, Dimitris Typas, Martijn S. Luijsterburg, Medine Cansoy, Paul Hensbergen, André Deelder, Anton de Groot, Syota Matsumoto, Kaoru Sugasawa, Nicolas Thoma, Wim Vermeulen, Harry Vrieling, and Leon Mullenders. PARP1 promotes nucleotide excision repair through DDB2 stabilization and recruitment of ALC1. Journal of Cell Biology, 199(2):235–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/2/235.

Pincus:2012:FLD

[PW12] David Pincus and Peter Walter. A first line of defense against ER stress. Journal of Cell Biology, 198(3):277–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
3/277.

Pietras:2011:CCR

[PWP11] Eric M. Pietras, Matthew R. Warr, and Emmanuelle Passegué. Cell cycle regulation in hematopoietic stem cells. Journal of Cell Biology, 195(5):709–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/5/709.

Ponugoti:2013:FPW

[PXZ^{+}13] Bhaskar Ponugoti, Fanxing Xu, Chenying Zhang, Chen Tian, Sandra Pacios, and Dana T. Graves. FOXO1 promotes wound healing through the up-regulation of TGF-β1 and prevention of oxidative stress. Journal of Cell Biology, 203(2):327–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
203/2/327.

Pettem:2013:IBA

[PYT^{+}13] Katherine L. Pettem, Daisaku Yokomaku, Hideto Takahashi, Yuan Ge, and Ann Marie Craig. Interaction between autismlinked MDGAs and neuroligins suppresses inhibitory synapse development. Journal of Cell Biology, 200(3):321–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/200/
3/321.

Petsalaki:2014:CPM

[PZ14] Eleni Petsalaki and George Zachos. Chk2 prevents mitotic exit when the majority of kinetochores are unattached. Journal of Cell Biology, 205(3):339–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/205/3/339.

Qu:2012:NMS

[QB12] Jian Qu and J. Michael Bishop. Nucleostemin maintains selfrenewal of embryonic stem cells and promotes reprogramming of somatic cells to pluripotency. Journal of Cell Biology, 197 (6):731–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/6/731.

Quintavalle:2010:MCP

[QECC10] Manuela Quintavalle, Leonardo Elia, Gianluigi Condorelli, and Sara A. Courtneidge. MicroRNA control of podosome formation in vascular smooth muscle cells in vivo and in vitro. Journal of Cell Biology, 189(1):13–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/1/13.

Quinones:2010:BPA

[QJO10] Gabriel A. Quinones, Janet Jin, and Anthony E. Oro. I-BAR protein antagonism of endocytosis mediates directional sensing during guided cell migration. Journal of Cell Biology, 189 (2):353–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/2/353.

Qin:2010:TCG

[QMHM10] Yi Qin, Walter H. Meisen, Yi Hao, and Ian G. Macara. Tuba, a Cdc42 GEF, is required for polarized spindle orientation during epithelial cyst formation. Journal of Cell Biology, 189 (4):661–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/4/661.

Qi:2012:BAC

[QTL^{+}12] Yanmei Qi, Xiaoxiang Tian, Jie Liu, Yaling Han, Alan M. Graham, M. Celeste Simon, Josef M. Penninger, Peter Carmeliet, and Shaohua Li. Bnip3 and AIF cooperate to induce apoptosis and cavitation during epithelial morphogenesis. Journal of Cell Biology, 198(1):103–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/1/103.

Qi:2013:BAC

[QTL^{+}13] Yanmei Qi, Xiaoxiang Tian, Jie Liu, Yaling Han, Alan M. Graham, M. Celeste Simon, Josef M. Penninger, Peter Carmeliet, and Shaohua Li. Bnip3 and AIF cooperate to induce apoptosis and cavitation during epithelial morphogenesis. Journal of Cell Biology, 201(1):165–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/201/1/165.

Qian:2011:TNA

[QWL^{+}11] Li Qian, Joshua D. Wythe, Jiandong Liu, Jerome Cartry, Georg Vogler, Bhagyalaxmi Mohapatra, Robyn T. Otway, Yu Huang, Isabelle N. King, Marjorie Maillet, Yi Zheng, Timothy Crawley, Ouarda Taghli-Lamallem, Christopher Semsarian, Sally Dunwoodie, David Winlaw, Richard P. Harvey, Diane Fatkin, Jeffrey A. Towbin, Jeffery D. Molkentin, Deepak Srivastava, Karen Ocorr, Benoit G. Bruneau, and Rolf Bodmer. Tinman/ nkx2-5 acts via miR-1 and upstream of Cdc42 to regulate heart function across species. Journal of Cell Biology, 193(7):1181–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/7/1181.

Rivera:2011:SPP

[RB11] Lee B. Rivera and Rolf A. Brekken. SPARC promotes pericyte recruitment via inhibition of endoglin-dependent TGFβ1 activity. Journal of Cell Biology, 193(7):1305–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
7/1305.

Rodal:2011:PET

[RBA^{+}11] Avital A. Rodal, Aline D. Blunk, Yulia Akbergenova, Ramon A. Jorquera, Lauren K. Buhl, and J. Troy Littleton. A presynaptic endosomal trafficking pathway controls synaptic growth signaling. Journal of Cell Biology, 193(1):201–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
1/201.

Rincon:2014:PRA

[RBB^{+}14] Sergio A. Rincon, Payal Bhatia, Claudia Bicho, Mercè Guzman-Vendrell, Vincent Fraisier, Weronika E. Borek, Flavia de Lima Alves, Florent Dingli, Damarys Loew, Juri Rappsilber, Kenneth E. Sawin, Sophie G. Martin, and Anne Paoletti. Pom1 regulates the assembly of Cdr2–Mid1 cortical nodes for robust spatial control of cytokinesis. Journal of Cell Biology, 206(1):61–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/1/61.

Ryan:2012:MSG

[RBF^{+}12] Scott D. Ryan, Kunal Bhanot, Andrew Ferrier, Yves De Repentigny, Alphonse Chu, Alexandre Blais, and Rashmi Kothary. Microtubule stability, Golgi organization, and transport flux require dystonin-a2–MAP1B interaction. Journal of Cell Biology, 196(6):727–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/6/727.

Rullo:2012:APS

[RBH^{+}12] Jacob Rullo, Henry Becker, Sharon J. Hyduk, Janice C. Wong, Genevieve Digby, Pamma D. Arora, Adrianet Puig Cano, John Hartwig, Christopher A. McCulloch, and Myron I. Cybulsky. Actin polymerization stabilizes α4β1 integrin anchors that mediate monocyte adhesion. Journal of Cell Biology, 197 (1):115–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/1/115.

Rzeczkowski:2011:CJT

[RBM^{+}11] Katharina Rzeczkowski, Knut Beuerlein, Helmut Müller, Oliver Dittrich-Breiholz, Heike Schneider, Daniela Kettner-Buhrow, Helmut Holtmann, and Michael Kracht. c-jun Nterminal kinase phosphorylates DCP1a to control formation of P bodies. Journal of Cell Biology, 194(4):581–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
4/581.

Rangrez:2013:DPI

[RBP^{+}13] Ashraf Yusuf Rangrez, Alexander Bernt, Reza Poyanmehr, Violetta Harazin, Inka Boomgaarden, Christian Kuhn, Astrid Rohrbeck, Derk Frank, and Norbert Frey. Dysbindin is a potent inducer of RhoA–SRF-mediated cardiomyocyte hypertrophy. Journal of Cell Biology, 203(4):643–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
4/643.

Ruoslahti:2010:TDN

[RBS10] Erkki Ruoslahti, Sangeeta N. Bhatia, and Michael J. Sailor. Targeting of drugs and nanoparticles to tumors. Journal of Cell Biology, 188(6):759–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/6/759.

Raleigh:2011:OSP

[RBY^{+}11] David R. Raleigh, Devin M. Boe, Dan Yu, Christopher R. Weber, Amanda M. Marchiando, Emily M. Bradford, Yingmin Wang, Licheng Wu, Eveline E. Schneeberger, Le Shen, and Jerrold R. Turner. Occludin S408 phosphorylation regulates tight junction protein interactions and barrier function.

Journal of Cell Biology, 193(3):565–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/3/565.

Rodier:2011:FFC

[RC11] Francis Rodier and Judith Campisi. Four faces of cellular senescence. Journal of Cell Biology, 192(4):547–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
4/547.

Rotty:2012:WIK

[RC12] Jeremy D. Rotty and Pierre A. Coulombe. A wound-induced keratin inhibits Src activity during keratinocyte migration and tissue repair. Journal of Cell Biology, 197(3):381–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
3/381.

Rago:2013:FCF

[RC13] Florencia Rago and Iain M. Cheeseman. The functions and consequences of force at kinetochores. Journal of Cell Biology, 200(5):557–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/5/557.

Ravel-Chapuis:2012:RBP

[RCBY^{+}12] Aymeric Ravel-Chapuis, Guy Bélanger, Ramesh S. Yadava, Mani S. Mahadevan, Luc DesGroseillers, Jocelyn Côté, and Bernard J. Jasmin. The RNA-binding protein Staufen1 is increased in DM1 skeletal muscle and promotes alternative premRNA splicing. Journal of Cell Biology, 196(6):699–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
6/699.

Rank:2012:KMK

[RCC^{+}12] Katherine C. Rank, Chun Ju Chen, Julia Cope, Ken Porche, Andreas Hoenger, Susan P. Gilbert, and Ivan Rayment. Kar3Vik1, a member of the Kinesin-14 superfamily, shows a novel kinesin microtubule binding pattern. Journal of Cell Biology, 197(7):957–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/7/957.

Rane:2010:SSI

[RCFH10] Neena S. Rane, Oishee Chakrabarti, Lionel Feigenbaum, and Ramanujan S. Hegde. Signal sequence insufficiency contributes to neurodegeneration caused by transmembrane prion protein. Journal of Cell Biology, 188(4):515–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
4/515.

Rodriguez:2010:EMK

[RCG^{+}10] Javier Rodríguez, Fernando Calvo, José M. González, Berta Casar, Vicente Andrés, and Piero Crespo. ERK1/ 2 MAP kinases promote cell cycle entry by rapid, kinase-independent disruption of retinoblastoma–lamin A complexes. Journal of Cell Biology, 191(5):967–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/5/967.

Rodriguez:2011:EMK

[RCG^{+}11] Javier Rodríguez, Fernando Calvo, José M. González, Berta Casar, Vicente Andrés, and Piero Crespo. ERK1/ 2 MAP kinases promote cell cycle entry by rapid, kinase-independent disruption of retinoblastoma–lamin A complexes. Journal of Cell Biology, 192(1):201–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/1/201.

Rainero:2012:DKC

[RCM^{+}12] Elena Rainero, Patrick T. Caswell, Patricia A. J. Muller, Joan Grindlay, Mary W. McCaffrey, Qifeng Zhang, Michael J. O. Wakelam, Karen H. Vousden, Andrea Graziani, and Jim C. Norman. Diacylglycerol kinase α controls RCP-dependent integrin trafficking to promote invasive migration. Journal of Cell Biology, 196(2):277–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/2/277.

Roscioli:2012:INR

[RDB^{+}12] Emanuele Roscioli, Laura Di Francesco, Alessio Bolognesi, Maria Giubettini, Serena Orlando, Amnon Harel, Maria Eugenia Schininà, and Patrizia Lavia. Importin-β negatively regulates multiple aspects of mitosis including RANGAP1 recruitment to kinetochores. Journal of Cell Biology, 196(4):

435–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/4/435.

Roubinet:2011:MNL

[RDC^{+}11] Chantal Roubinet, Barbara Decelle, Gaëtan Chicanne, Jonas F. Dorn, Bernard Payrastre, François Payre, and Sébastien Carreno. Molecular networks linked by Moesin drive remodeling of the cell cortex during mitosis. Journal of Cell Biology, 195(1):99–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/1/99.

Ruault:2011:CHS

[RDLT11] Myriam Ruault, Arnaud De Meyer, Isabelle Loïodice, and Angela Taddei. Clustering heterochromatin: Sir3 promotes telomere clustering independently of silencing in yeast. Journal of Cell Biology, 192(3):417–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/3/417.

Rizk:2014:KKS

[RDPG14] Rania S. Rizk, Katherine A. DiScipio, Kathleen G. Proudfoot, and Mohan L. Gupta. The kinesin-8 Kip3 scales anaphase spindle length by suppression of midzone microtubule polymerization. Journal of Cell Biology, 204(6):965–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
6/965.

Rodriguez-Fraticelli:2012:CCC

[RFAA^{+}12] Alejo E. Rodríguez-Fraticelli, Muriel Auzan, Miguel A. Alonso, Michel Bornens, and Fernando Martín-Belmonte. Cell confinement controls centrosome positioning and lumen initiation during epithelial morphogenesis. Journal of Cell Biology, 198(6): 1011–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/6/1011.

Ruggiano:2014:ADP

[RFC14] Annamaria Ruggiano, Ombretta Foresti, and Pedro Carvalho. ER-associated degradation: Protein quality control and beyond. Journal of Cell Biology, 204(6):869–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
6/869.

Ryu:2010:PDS

[RFK^{+}10] Hyunju Ryu, Maiko Furuta, Donald Kirkpatrick, Steven P. Gygi, and Yoshiaki Azuma. PIASy-dependent SUMOylation regulates DNA topoisomerase IIα activity. Journal of Cell Biology, 191(4):783–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/4/783.

Ray:2013:CPE

[RFL13] Samriddha Ray, Henry P. Foote, and Terry Lechler. β-catenin protects the epidermis from mechanical stresses. Journal of Cell Biology, 202(1):45–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/1/45.

Rojas:2012:RPSa

[RFRV12a] Ana Maria Rojas, Gloria Fuentes, Antonio Rausell, and Alfonso Valencia. The Ras protein superfamily: Evolutionary tree and role of conserved amino acids. Journal of Cell Biology, 196(2):189–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/2/189.

Rojas:2012:RPSb

[RFRV12b] Ana Maria Rojas, Gloria Fuentes, Antonio Rausell, and Alfonso Valencia. The Ras protein superfamily: Evolutionary tree and role of conserved amino acids. Journal of Cell Biology, 196(4):545–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/4/545.

Rodriguez-Fraticelli:2010:CGI

[RFVE^{+}10] Alejo E. Rodriguez-Fraticelli, Silvia Vergarajauregui, Dennis J. Eastburn, Anirban Datta, Miguel A. Alonso, Keith Mostov, and Fernando Martín-Belmonte. The Cdc42 GEF Intersectin 2 controls mitotic spindle orientation to form the lumen during epithelial morphogenesis. Journal of Cell Biology, 189 (4):725–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/4/725.

Rajapakse:2011:ENO

[RG11] Indika Rajapakse and Mark Groudine. On emerging nuclear order. Journal of Cell Biology, 192(5):711–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
5/711.

Ragkousi:2014:CDM

[RG14] Katerina Ragkousi and Matthew C. Gibson. Cell division and the maintenance of epithelial order. Journal of Cell Biology, 207(2):181–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/2/181.

Romao:2013:APS

[RGB^{+}13] Susana Romao, Nathalie Gasser, Andrea C. Becker, Bruno Guhl, Milica Bajagic, Danusia Vanoaica, Urs Ziegler, Joachim Roesler, Jörn Dengjel, Janine Reichenbach, and Christian Münz. Autophagy proteins stabilize pathogen-containing phagosomes for prolonged MHC II antigen processing. Journal of Cell Biology, 203(5):757–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/5/757.

Rossio:2010:RCR

[RGF^{+}10] Valentina Rossio, Elena Galati, Matteo Ferrari, Achille Pellicioli, Takashi Sutani, Katsuhiko Shirahige, Giovanna Lucchini, and Simonetta Piatti. The RSC chromatin-remodeling complex influences mitotic exit and adaptation to the spindle assembly checkpoint by controlling the Cdc14 phosphatase. Journal of Cell Biology, 191(5):981–??, November 2010. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/191/5/
981.

Randazzo:2013:ORA

[RGL^{+}13] Davide Randazzo, Emiliana Giacomello, Stefania Lorenzini, Daniela Rossi, Enrico Pierantozzi, Bert Blaauw, Carlo Reggiani, Stephan Lange, Angela K. Peter, Ju Chen, and Vincenzo Sorrentino. Obscurin is required for ankyrinB-dependent dystrophin localization and sarcolemma integrity. Journal of Cell Biology, 200(4):523–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/4/523.

Rutkowski:2010:RBC

[RH10] D. Thomas Rutkowski and Ramanujan S. Hegde. Regulation of basal cellular physiology by the homeostatic unfolded protein response. Journal of Cell Biology, 189(5):783–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
5/783.

Rozelle:2011:CPC

[RHK11] Daniel K. Rozelle, Scott D. Hansen, and Kenneth B. Kaplan. Chromosome passenger complexes control anaphase duration and spindle elongation via a kinesin-5 brake. Journal of Cell Biology, 193(2):285–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/2/285.

Rigo:2012:ABT

[RHKB12] Frank Rigo, Yimin Hua, Adrian R. Krainer, and C. Frank Bennett. Antisense-based therapy for the treatment of spinal muscular atrophy. Journal of Cell Biology, 199(1):21–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
1/21.

Reymond:2012:CPT

[RIG^{+}12] Nicolas Reymond, Jae Hong Im, Ritu Garg, Francisco M. Vega, Barbara Borda d’Agua, Philippe Riou, Susan Cox, Ferran Valderrama, Ruth J. Muschel, and Anne J. Ridley. Cdc42 promotes transendothelial migration of cancer cells through β1 integrin. Journal of Cell Biology, 199(4):653–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
4/653.

Ricke:2012:BKA

[RJM^{+}12] Robin M. Ricke, Karthik B. Jeganathan, Liviu Malureanu, Andrew M. Harrison, and Jan M. van Deursen. Bub1 kinase activity drives error correction and mitotic checkpoint control but not tumor suppression. Journal of Cell Biology, 199(6): 931–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/6/931.

Ricke:2011:BOI

[RJvD11] Robin M. Ricke, Karthik B. Jeganathan, and Jan M. van Deursen. Bub1 overexpression induces aneuploidy and tumor formation through Aurora B kinase hyperactivation. Journal of Cell Biology, 193(6):1049–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/6/1049.

Rahimov:2013:CMM

[RK13] Fedik Rahimov and Louis M. Kunkel. Cellular and molecular mechanisms underlying muscular dystrophy. Journal of Cell Biology, 201(4):499–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/4/499.

Rosic:2014:RCSa

[RKE14a] Silvana Rošić, Florian Köhler, and Sylvia Erhardt. Repetitive centromeric satellite RNA is essential for kinetochore formation and cell division. Journal of Cell Biology, 207(3): 335–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/3/335.

Rosic:2014:RCSb

[RKE14b] Silvana Rošić, Florian Köhler, and Sylvia Erhardt. Repetitive centromeric satellite RNA is essential for kinetochore formation and cell division. Journal of Cell Biology, 207(5): 673–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/5/673.

Ray:2010:MIT

[RKG^{+}10] Samriddha Ray, Kazunori Kume, Sneha Gupta, Wanzhong Ge, Mohan Balasubramanian, Dai Hirata, and Dannel McCollum. The mitosis-to-interphase transition is coordinated by cross talk between the SIN and MOR pathways in Schizosaccharomyces pombe. Journal of Cell Biology, 190(5):793– ??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/5/793.

Rubenstein:2012:ASE

[RKG^{+}12] Eric M. Rubenstein, Stefan G. Kreft, Wesley Greenblatt, Robert Swanson, and Mark Hochstrasser. Aberrant substrate engagement of the ER translocon triggers degradation by the Hrd1 ubiquitin ligase. Journal of Cell Biology, 197(6):761–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
197/6/761.

Rouvinski:2014:LIP

[RKK^{+}14] Alexander Rouvinski, Sharon Karniely, Maria Kounin, Sanaa Moussa, Miri D. Goldberg, Gabriela Warburg, Roman Lyakhovetsky, Dulce Papy-Garcia, Janine Kutzsche, Carsten Korth, George A. Carlson, Susan F. Godsave, Peter J. Peters, Katarina Luhr, Krister Kristensson, and Albert Taraboulos. Live imaging of prions reveals nascent PrP^{Sc} in cell-surface, raft-associated amyloid strings and webs. Journal of Cell Biology, 204(3):423–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/3/423.

Roux:2012:PBL

[RKRB12] Kyle J. Roux, Dae In Kim, Manfred Raida, and Brian Burke. A promiscuous biotin ligase fusion protein identifies proximal and interacting proteins in mammalian cells. Journal of Cell Biology, 196(6):801–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/6/801.

Riley:2010:UAA

[RKS^{+}10] Brigit E. Riley, Stephen E. Kaiser, Thomas A. Shaler, Aylwin C. Y. Ng, Taichi Hara, Mark S. Hipp, Kasper Lage, Ramnik J. Xavier, Kwon-Yul Ryu, Keiko Taguchi, Masayuki Yamamoto, Keiji Tanaka, Noboru Mizushima, Masaaki Komatsu, and Ron R. Kopito. Ubiquitin accumulation in autophagydeficient mice is dependent on the Nrf2-mediated stress response pathway: a potential role for protein aggregation in autophagic substrate selection. Journal of Cell Biology, 191(3): 537–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/3/537.

Revelo:2014:NPS

[RKT^{+}14] Natalia H. Revelo, Dirk Kamin, Sven Truckenbrodt, Aaron B. Wong, Kirsten Reuter-Jessen, Ellen Reisinger, Tobias Moser, and Silvio O. Rizzoli. A new probe for super-resolution imaging of membranes elucidates trafficking pathways. Journal of Cell Biology, 205(4):591–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/4/591.

Rohner:2013:PRP

[RKW^{+}13] Sabine Rohner, Veronique Kalck, Xuefei Wang, Kohta Ikegami, Jason D. Lieb, Susan M. Gasser, and Peter Meister. Promoter- and RNA polymerase II-dependent hsp-16 gene association with nuclear pores in Caenorhabditis elegans. Journal of Cell Biology, 200(5):589–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/5/589.

Reversi:2014:PMP

[RLS^{+}14] Alessandra Reversi, Eva Loeser, Devaraj Subramanian, Carsten Schultz, and Stefano De Renzis. Plasma membrane phosphoinositide balance regulates cell shape during Drosophila embryo morphogenesis. Journal of Cell Biology, 205(3):395–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/3/395.

Rome:2010:ACP

[RMF^{+}10] Pierre Romé, Emilie Montembault, Nathalie Franck, Aude Pascal, David M. Glover, and Régis Giet. Aurora A contributes to p150^{glued} phosphorylation and function during mitosis. Journal of Cell Biology, 189(4):651–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/4/651.

Reed:2012:HGC

[RMG^{+}12] Jonathan C. Reed, Britta Molter, Clair D. Geary, John Mc-Nevin, Julie McElrath, Samina Giri, Kevin C. Klein, and Jaisri R. Lingappa. HIV-1 Gag co-opts a cellular complex containing DDX6, a helicase that facilitates capsid assembly.

Journal of Cell Biology, 198(3):439–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/3/439.

Raffaello:2010:JTF

[RMM^{+}10] Anna Raffaello, Giulia Milan, Eva Masiero, Silvia Carnio, Donghoon Lee, Gerolamo Lanfranchi, Alfred Lewis Goldberg, and Marco Sandri. JunB transcription factor maintains skeletal muscle mass and promotes hypertrophy. Journal of Cell Biology, 191(1):101–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/1/101.

Rajgor:2014:MMP

[RMS^{+}14] Dipen Rajgor, Jason A. Mellad, Daniel Soong, Jerome B. Rattner, Marvin J. Fritzler, and Catherine M. Shanahan. Mammalian microtubule P-body dynamics are mediated by nesprin-1. Journal of Cell Biology, 205(4):457–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/205/4/457.

Rivera-Molina:2013:LCI

[RMT13] Felix Rivera-Molina and Derek Toomre. Live-cell imaging of exocyst links its spatiotemporal dynamics to various stages of vesicle fusion. Journal of Cell Biology, 201(5):673–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
5/673.

Reid:2012:EEN

[RN12] David W. Reid and Christopher V. Nicchitta. The enduring enigma of nuclear translation. Journal of Cell Biology, 197(1):7–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/1/7.

Rahman:2014:DSM

[RNS^{+}14] Motiur Rahman, Niraj K. Nirala, Alka Singh, Lihua Julie Zhu, Kaori Taguchi, Takeshi Bamba, Eiichiro Fukusaki, Leslie M. Shaw, David G. Lambright, Jairaj K. Acharya, and Usha R. Acharya. Drosophila Sirt2/mammalian SIRT3 deacetylates ATP synthase β and regulates complex V activity. Journal of Cell Biology, 206(2):289–??, July 2014. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/206/2/289.

Rossner:2010:NSS

[Ros10a] Mike Rossner. New style, same substance. Journal of Cell Biology, 189(2):195–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/2/195.

Rossner:2010:URA

[Ros10b] Mike Rossner. Updating realistic access. Journal of Cell Biology, 189(3):393–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/3/393.

Rossner:2013:NMN

[Ros13] Mike Rossner. New mandates? No problem for The Rockefeller University Press. Journal of Cell Biology, 201(1):7–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
1/7.

Rubio:2011:HAE

[RPK^{+}11] Claudia Rubio, David Pincus, Alexei Korennykh, Sebastian Schuck, Hana El-Samad, and Peter Walter. Homeostatic adaptation to endoplasmic reticulum stress depends on Ire1 kinase activity. Journal of Cell Biology, 193(1):171–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
1/171.

Rizzo:2013:DER

[RPM^{+}13] Riccardo Rizzo, Seetharaman Parashuraman, Peppino Mirabelli, Claudia Puri, John Lucocq, and Alberto Luini. The dynamics of engineered resident proteins in the mammalian Golgi complex relies on cisternal maturation. Journal of Cell Biology, 201(7):1027–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/7/1027.

Richter:2014:SFA

[RPO^{+}14] Viviane Richter, Catherine S. Palmer, Laura D. Osellame, Abeer P. Singh, Kirstin Elgass, David A. Stroud, Hiromi Sesaki, Marc Kvansakul, and Michael T. Ryan. Structural and functional analysis of MiD51, a dynamin receptor required for mitochondrial fission. Journal of Cell Biology, 204(4): 477–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/4/477.

Raposo:2013:EVE

[RS13] Graça Raposo and Willem Stoorvogel. Extracellular vesicles: Exosomes, microvesicles, and friends. Journal of Cell Biology, 200(4):373–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/4/373.

Rogers:2013:SMC

[RSB13] Crystal D. Rogers, Ankur Saxena, and Marianne E. Bronner. Sip1 mediates an E–cadherin–to–N–cadherin switch during cranial neural crest EMT. Journal of Cell Biology, 203(5): 835–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/5/835.

Raab:2012:CSS

[RSD^{+}12] Matthew Raab, Joe Swift, P. C. Dave P. Dingal, Palak Shah, Jae-Won Shin, and Dennis E. Discher. Crawling from soft to stiff matrix polarizes the cytoskeleton and phosphoregulates myosin–II heavy chain. Journal of Cell Biology, 199(4): 669–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/4/669.

Rohn:2011:CRS

[RSL^{+}11] Jennifer L. Rohn, David Sims, Tao Liu, Marina Fedorova, Frieder Schöck, Joseph Dopie, Maria K. Vartiainen, Amy A. Kiger, Norbert Perrimon, and Buzz Baum. Comparative RNAi screening identifies a conserved core metazoan actinome by phenotype. Journal of Cell Biology, 194(5):789–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
5/789.

Raghuram:2013:PPH

[RSM^{+}13] Nikhil Raghuram, Hilmar Strickfaden, Darin McDonald, Kylie Williams, He Fang, Craig Mizzen, Jeffrey J. Hayes, John Th’ng, and Michael J. Hendzel. Pin1 promotes histone H1 dephosphorylation and stabilizes its binding to chromatin. Journal of Cell Biology, 203(1):57–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/1/57.

Romer:2013:SMF

[RSRK13] Anthony I. Romer, Jagmohan Singh, Satish Rattan, and Robert S. Krauss. Smooth muscle fascicular reorientation is required for esophageal morphogenesis and dependent on Cdo. Journal of Cell Biology, 201(2):309–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/2/309.

Reczek:2013:IBC

[RSS^{+}13] Colleen R. Reczek, Matthias Szabolcs, Jeremy M. Stark, Thomas Ludwig, and Richard Baer. The interaction between CtIP and BRCA1 is not essential for resection-mediated DNA repair or tumor suppression. Journal of Cell Biology, 201 (5):693–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/5/693.

Rifkin:2010:BMG

[RT10] Daniel B. Rifkin and Vesna Todorovic. Bone matrix to growth factors: location, location, location. Journal of Cell Biology, 190(6):949–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/6/949.

Reggiori:2012:ART

[RT12] Fulvio Reggiori and Sharon A. Tooze. Autophagy regulation through Atg9 traffic. Journal of Cell Biology, 198(2):151–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
198/2/151.

Reboutier:2012:NBA

[RTC^{+}12] David Reboutier, Marie-Bérengère Troadec, Jean-Yves Cremet, Kenji Fukasawa, and Claude Prigent. Nucleophosmin/ b23 activates Aurora A at the centrosome through phosphorylation of serine 89. Journal of Cell Biology, 197(1):19–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/197/1/19.

Reboutier:2013:AICa

[RTC^{+}13a] David Reboutier, Marie-Bérengère Troadec, Jean-Yves Cremet, Lucie Chauvin, Vincent Guen, Patrick Salaun, and Claude Prigent. Aurora A is involved in central spindle assembly through phosphorylation of Ser 19 in P150Glued. Journal of Cell Biology, 201(1):65–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/1/65.

Reboutier:2013:AICb

[RTC^{+}13b] David Reboutier, Marie-Bérengère Troadec, Jean-Yves Cremet, Lucie Chauvin, Vincent Guen, Patrick Salaun, and Claude Prigent. Aurora A is involved in central spindle assembly through phosphorylation of Ser 19 in P150Glued. Journal of Cell Biology, 201(3):487–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/3/487.

Raaijmakers:2013:SDD

[RTM13] Jonne A. Raaijmakers, Marvin E. Tanenbaum, and René H. Medema. Systematic dissection of dynein regulators in mitosis. Journal of Cell Biology, 201(2):201–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/2/201.

Ricke:2013:AHD

[RvD13] Robin M. Ricke and Jan M. van Deursen. Aneuploidy in health, disease, and aging. Journal of Cell Biology, 201 (1):11–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/1/11.

Rankin:2010:LAM

[RW10] Kathleen E. Rankin and Linda Wordeman. Long astral microtubules uncouple mitotic spindles from the cytokinetic furrow. Journal of Cell Biology, 190(1):35–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/1/35.

Rossio:2011:SRC

[RY11] Valentina Rossio and Satoshi Yoshida. Spatial regulation of Cdc55–PP2A by Zds1/Zds2 controls mitotic entry and mitotic exit in budding yeast. Journal of Cell Biology, 193(3):445–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
193/3/445.

Riche:2013:ECR

[RZA^{+}13] Soizic Riche, Melissa Zouak, Françoise Argoul, Alain Arneodo, Jacques Pecreaux, and Marie Delattre. Evolutionary comparisons reveal a positional switch for spindle pole oscillations in Caenorhabditis embryos. Journal of Cell Biology, 201 (5):653–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/5/653.

Regairaz:2011:MMD

[RZF^{+}11] Marie Regairaz, Yong-Wei Zhang, Haiqing Fu, Keli K. Agama, Nalini Tata, Surbhi Agrawal, Mirit I. Aladjem, and Yves Pommier. Mus81-mediated DNA cleavage resolves replication forks stalled by topoisomerase I–DNA complexes. Journal of Cell Biology, 195(5):739–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/5/739.

Reuter:2014:BDO

[RZS^{+}14] Marcel Reuter, Alex Zelensky, Ihor Smal, Erik Meijering, Wiggert A. van Cappellen, H. Martijn de Gruiter, Gijsbert J. van Belle, Martin E. van Royen, Adriaan B. Houtsmuller, Jeroen Essers, Roland Kanaar, and Claire Wyman. BRCA2 diffuses as oligomeric clusters with RAD51 and changes mobility after DNA damage in live cells. Journal of Cell Biology, 207(5): 599–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/5/599.

Saitoh:2010:RII

[SA10a] Tatsuya Saitoh and Shizuo Akira. Regulation of innate immune responses by autophagy-related proteins. Journal of Cell Biology, 189(6):925–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/6/925.

Saleem:2010:SCB

[SA10b] Ramsey A. Saleem and John D. Aitchison. Systems cell biology of the mitotic spindle. Journal of Cell Biology, 188 (1):7–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/1/7.

Sabatini:2010:PSB

[Sab10] David D. Sabatini. Philip Siekevitz: Bridging biochemistry and cell biology. Journal of Cell Biology, 189(1):3–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
1/3.

Schmidt:2011:NER

[SAG^{+}11] Nadine Schmidt, Mohammed Akaaboune, Nadesan Gajendran, Isabel Martinez-Peña y Valenzuela, Sarah Wakefield, Raphael Thurnheer, and Hans Rudolf Brenner. Neuregulin/ ErbB regulate neuromuscular junction development by phosphorylation of α-dystrobrevin. Journal of Cell Biology, 195(7): 1171–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/7/1171.

Saraogi:2014:RCR

[SAoS14] Ishu Saraogi, David Akopian, and Shu ou Shan. Regulation of cargo recognition, commitment, and unloading drives cotranslational protein targeting. Journal of Cell Biology, 205 (5):693–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/5/693.

Schwarz:2014:CDR

[SB14] Dianne S. Schwarz and Michael D. Blower. The calciumdependent ribonuclease XendoU promotes ER network formation through local RNA degradation. Journal of Cell Biology, 207(1):41–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/1/41.

Sakamoto:2013:ABI

[SBEM13] Yasuhisa Sakamoto, Batiste Boëda, and Sandrine Etienne-Manneville. APC binds intermediate filaments and is required for their reorganization during cell migration. Journal of Cell Biology, 200(3):249–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/200/3/249.

Sengupta:2010:CNE

[SBP^{+}10a] Jayati Sengupta, Cyril Bussiere, Jesper Pallesen, Matthew West, Arlen W. Johnson, and Joachim Frank. Characterization of the nuclear export adaptor protein Nmd3 in association with the 60S ribosomal subunit. Journal of Cell Biology, 189 (7):1079–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/7/1079.

Sinclair:2010:DPL

[SBP^{+}10b] Paul Sinclair, Qian Bian, Matt Plutz, Edith Heard, and Andrew S. Belmont. Dynamic plasticity of large-scale chromatin structure revealed by self-assembly of engineered chromosome regions. Journal of Cell Biology, 190(5):761–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
5/761.

Schmidt:2011:RTI

[SBR^{+}11] Ute Schmidt, Eugenia Basyuk, Marie-Cécile Robert, Minoru Yoshida, Jean-Philippe Villemin, Didier Auboeuf, Stuart Aitken, and Edouard Bertrand. Real-time imaging of cotranscriptional splicing reveals a kinetic model that reduces noise: implications for alternative splicing regulation. Journal of Cell Biology, 193(5):819–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/5/819.

Schmidt:2012:ARC

[SBS^{+}12] Nadine Schmidt, Sreya Basu, Stefan Sladecek, Sabrina Gatti, Jeffrey van Haren, Susan Treves, Jan Pielage, Niels Galjart, and Hans Rudolf Brenner. Agrin regulates CLASP2-mediated capture of microtubules at the neuromuscular junction synaptic membrane. Journal of Cell Biology, 198(3):421–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
3/421.

Shah:2013:ADM

[SBTF13] Mehul Shah, Oscar Y. Baterina, Vanessa Taupin, and Marilyn G. Farquhar. ARH directs megalin to the endocytic recycling compartment to regulate its proteolysis and gene expression. Journal of Cell Biology, 202(1):113–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
1/113.

Sung:2010:CBV

[SC10a] Ching-Hwa Sung and Jen-Zen Chuang. The cell biology of vision. Journal of Cell Biology, 190(6):953–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
6/953.

Sutterlin:2010:GCB

[SC10b] Christine Sütterlin and Antonino Colanzi. The Golgi and the centrosome: building a functional partnership. Journal of Cell Biology, 188(5):621–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/5/621.

Sumigray:2011:LEC

[SCL11] Kaelyn D. Sumigray, Hsin Chen, and Terry Lechler. Lis1 is essential for cortical microtubule organization and desmosome stability in the epidermis. Journal of Cell Biology, 194(4): 631–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/4/631.

Sau:2014:SDE

[SCL^{+}14] Soumitra Sau, Michael N. Conrad, Chih-Ying Lee, David B. Kaback, Michael E. Dresser, and Makkuni Jayaram. A selfish DNA element engages a meiosis-specific motor and telomeres for germ-line propagation. Journal of Cell Biology, 205 (5):643–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/5/643.

Shin:2014:DER

[SCN^{+}14] Nah-Young Shin, Hyewon Choi, Lynn Neff, Yumei Wu, Hiroaki Saito, Shawn M. Ferguson, Pietro De Camilli, and Roland Baron. Dynamin and endocytosis are required for the fusion of osteoclasts and myoblasts. Journal of Cell Biology, 207 (1):73–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/1/73.

Stavoe:2012:NIS

[SCR12] Andrea K. H. Stavoe and Daniel A. Colón-Ramos. Netrin instructs synaptic vesicle clustering through Rac GTPase, MIG-10, and the actin cytoskeleton. Journal of Cell Biology, 197 (1):75–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/1/75.

Stevens:2010:DAC

[SDB^{+}10] Naomi R. Stevens, Jeroen Dobbelaere, Kathrin Brunk, Anna Franz, and Jordan W. Raff. Drosophila Ana2 is a conserved centriole duplication factor. Journal of Cell Biology, 188(3): 313–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/3/313.

So:2010:ASS

[SDC10] Sairei So, Anthony J. Davis, and David J. Chen. Autophosphorylation at serine 1981 stabilizes ATM at DNA damage sites. Journal of Cell Biology, 188(3):443–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
3/443.

Snow:2013:DNI

[SDD^{+}13] Chelsi J. Snow, Ashraf Dar, Anindya Dutta, Ralph H. Kehlenbach, and Bryce M. Paschal. Defective nuclear import of Tpr in Progeria reflects the Ran sensitivity of large cargo transport. Journal of Cell Biology, 201(4):541–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/4/541.

Sainski:2014:CGHa

[SDN^{+}14a] Amy M. Sainski, Haiming Dai, Sekar Natesampillai, Yuan-Ping Pang, Gary D. Bren, Nathan W. Cummins, Cristina Correia, X. Wei Meng, James E. Tarara, Marina Ramirez-Alvarado, David J. Katzmann, Christina Ochsenbauer, John C. Kappes, Scott H. Kaufmann, and Andrew D. Badley. Casp8p41 generated by HIV protease kills CD4 T cells through direct Bak activation. Journal of Cell Biology, 206(7):867– ??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/7/867.

Sainski:2014:CGHb

[SDN^{+}14b] Amy M. Sainski, Haiming Dai, Sekar Natesampillai, Yuan-Ping Pang, Gary D. Bren, Nathan W. Cummins, Cristina Correia, X. Wei Meng, James E. Tarara, Marina Ramirez-Alvarado, David J. Katzmann, Christina Ochsenbauer, John C. Kappes, Scott H. Kaufmann, and Andrew D. Badley. Casp8p41 generated by HIV protease kills CD4 T cells through direct Bak activation. Journal of Cell Biology, 207(1):159– ??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/1/159.

Sahly:2012:LUP

[SDS^{+}12a] Iman Sahly, Eric Dufour, Cataldo Schietroma, Vincent Michel, Amel Bahloul, Isabelle Perfettini, Elise Pepermans, Amrit Estivalet, Diane Carette, Asadollah Aghaie, Inga Ebermann, Andrea Lelli, Maria Iribarne, Jean-Pierre Hardelin, Dominique Weil, José-Alain Sahel, Aziz El-Amraoui, and Christine Petit. Localization of Usher 1 proteins to the photoreceptor calyceal processes, which are absent from mice. Journal of Cell Biology, 199(2):381–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/2/381.

Sivadas:2012:FKA

[SDS^{+}12b] Priyanka Sivadas, Jennifer M. Dienes, Martin St. Maurice, William D. Meek, and Pinfen Yang. A flagellar A-kinase anchoring protein with two amphipathic helices forms a structural scaffold in the radial spoke complex. Journal of Cell Biology, 199(4):639–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/4/639.

Sedwick:2010:ADU

[Sed10a] Caitlin Sedwick. Andy Dillin: Using aging research to probe biology. Journal of Cell Biology, 189(4):616–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
4/616.

Sedwick:2010:BWN

[Sed10b] Caitlin Sedwick. Bettina Winckler: Neuronal polarity on her mind. Journal of Cell Biology, 191(1):4–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/191/1/4.

Sedwick:2010:CTM

[Sed10c] Caitlin Sedwick. Craig Thompson: The method to cancer’s madness. Journal of Cell Biology, 191(4):696–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
4/696.

Sedwick:2010:DPG

[Sed10d] Caitlin Sedwick. David Pellman: Grasping the geometry of cancer. Journal of Cell Biology, 190(1):4–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/190/1/4.

Sedwick:2010:DHH

[Sed10e] Caitlin Sedwick. Doug Hilton: At home with blood cell biology. Journal of Cell Biology, 188(6):754–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/188/6/754.

Sedwick:2010:GWG

[Sed10f] Caitlin Sedwick. Graham Warren: Gaining ground on the Golgi. Journal of Cell Biology, 188(4):448–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
4/448.

Sedwick:2010:INA

[Sed10g] Caitlin Sedwick. Inke Näthke: The ABCs of APC. Journal of Cell Biology, 189(5):774–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/5/774.

Sedwick:2010:IVM

[Sed10h] Caitlin Sedwick. Isabelle Vernos: Motoring around the mitotic spindle. Journal of Cell Biology, 188(5):616–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
5/616.

Sedwick:2010:JHA

[Sed10i] Caitlin Sedwick. Jiahuai Han: Aflame on inflammation and p38. Journal of Cell Biology, 191(2):228–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
2/228.

Sedwick:2010:JBR

[Sed10j] Caitlin Sedwick. Joan Brugge: Running rings around cancer. Journal of Cell Biology, 189(6):922–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/6/922.

Sedwick:2010:LGM

[Sed10k] Caitlin Sedwick. Laurie Glimcher: Merging cell biology and immune function. Journal of Cell Biology, 189(2):192–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
2/192.

Sedwick:2010:LRS

[Sed10l] Caitlin Sedwick. Liz Rhoades: Single molecules tell many stories. Journal of Cell Biology, 191(6):1046–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
6/1046.

Sedwick:2010:NMA

[Sed10m] Caitlin Sedwick. Noboru Mizushima: All about autophagy. Journal of Cell Biology, 190(6):946–??, September 2010. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/190/6/

Sedwick:2010:RHP

[Sed10n] Caitlin Sedwick. Ramanujan Hegde: The prion puzzle and protein translocation. Journal of Cell Biology, 191(7):1222– ??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/7/1222.

Sedwick:2010:SSC

[Sed10o] Caitlin Sedwick. Sandra Schmid: Collaring endocytosis. Journal of Cell Biology, 190(2):162–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/2/162.

Sedwick:2010:TSK

[Sed10p] Caitlin Sedwick. Ted Salmon: Kinetochores at the core of it all. Journal of Cell Biology, 191(5):896–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
5/896.

Sedwick:2011:ASA

[Sed11a] Caitlin Sedwick. Alejandro Sánchez Alvarado: Bootstrapping flatworms into the molecular age. Journal of Cell Biology, 194 (4):510–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/4/510.

Sedwick:2011:AAG

[Sed11b] Caitlin Sedwick. Anna Akhmanova: Great tips on microtubules. Journal of Cell Biology, 195(2):168–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
2/168.

Sedwick:2011:BGF

[Sed11c] Caitlin Sedwick. Benny Geiger: a force in the study of focal adhesions. Journal of Cell Biology, 195(3):346–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
3/346.

Sedwick:2011:CDE

[Sed11d] Caitlin Sedwick. Carol Dieckmann: an eye on organellar biology. Journal of Cell Biology, 194(1):4–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/1/4.

Sedwick:2011:DBG

[Sed11e] Caitlin Sedwick. David Bilder: Getting to know epithelia inside and out. Journal of Cell Biology, 193(6):956–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
6/956.

Sedwick:2011:DML

[Sed11f] Caitlin Sedwick. Denise Montell: Lighting the way in border cell migration. Journal of Cell Biology, 195(4):540–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
195/4/540.

Sedwick:2011:ESG

[Sed11g] Caitlin Sedwick. Erik Sahai: Getting the whole picture of metastasis. Journal of Cell Biology, 193(3):428–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
3/428.

Sedwick:2011:HSH

[Sed11h] Caitlin Sedwick. Harald Stenmark: Hands on FYVE-fingers. Journal of Cell Biology, 192(4):544–??, February 2011. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/192/4/
544.

Sedwick:2011:JSQ

[Sed11i] Caitlin Sedwick. Jagesh Shah: a quantitative approach to cell biology. Journal of Cell Biology, 193(7):1134–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
7/1134.

Sedwick:2011:JSR

[Sed11j] Caitlin Sedwick. Joan Steitz: RNA is a many-splendored thing. Journal of Cell Biology, 192(5):708–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/5/708.

Sedwick:2011:MBK

[Sed11k] Caitlin Sedwick. María Blasco: Keeping a cap on cancer and aging. Journal of Cell Biology, 192(3):370–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
3/370.

Sedwick:2011:MOA

[Sed11l] Caitlin Sedwick. Michael Overholtzer: Answering existential questions on entosis. Journal of Cell Biology, 195(6): 924–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/6/924.

Sedwick:2011:PCI

[Sed11m] Caitlin Sedwick. Pascale Cossart: The ins and outs of Listeria. Journal of Cell Biology, 192(6):904–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/6/904.

Sedwick:2011:PWI

[Sed11n] Caitlin Sedwick. Peter Walter: Investigating how the ER handles secretory proteins. Journal of Cell Biology, 193(1):4–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
193/1/4.

Sedwick:2011:RLT

[Sed11o] Caitlin Sedwick. Rong Li: Tipping the balance in cell biology. Journal of Cell Biology, 195(7):1068–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
7/1068.

Sedwick:2011:RKR

[Sed11p] Caitlin Sedwick. Rüdiger Klein: Reading the guideposts for axon guidance. Journal of Cell Biology, 194(2):162–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
2/162.

Sedwick:2011:RLG

[Sed11q] Caitlin Sedwick. Ruth Lehmann: Germ cells do things differently. Journal of Cell Biology, 194(5):660–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
5/660.

Sedwick:2011:SIL

[Sed11r] Caitlin Sedwick. Shinya Inoué: Lighting the way in microscopy. Journal of Cell Biology, 194(6):810–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
6/810.

Sedwick:2011:UEB

[Sed11s] Caitlin Sedwick. Ulrike Eggert: Big things from small molecules. Journal of Cell Biology, 194(3):352–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
3/352.

Sedwick:2011:VWO

[Sed11t] Caitlin Sedwick. Valerie Weaver: Overcoming cancer’s stiff resistance. Journal of Cell Biology, 193(5):802–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
5/802.

Sedwick:2011:VBH

[Sed11u] Caitlin Sedwick. Vann Bennett: How ankyrin holds it all together. Journal of Cell Biology, 195(5):706–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
5/706.

Sedwick:2012:AKD

[Sed12a] Caitlin Sedwick. Alexey Khodjakov: Diving into the mitotic apparatus. Journal of Cell Biology, 197(6):694–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
6/694.

Sedwick:2012:BTD

[Sed12b] Caitlin Sedwick. Barry Thompson: The delicate choreography in growing epithelia. Journal of Cell Biology, 198(3):268–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
198/3/268.

Sedwick:2012:BNP

[Sed12c] Caitlin Sedwick. Bruce Nicklas: Pioneering studies on spindle forces. Journal of Cell Biology, 198(4):474–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
4/474.

Sedwick:2012:CBS

[Sed12d] Caitlin Sedwick. Cédric Blanpain: The stories stem cells tell. Journal of Cell Biology, 199(1):4–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/1/4.

Sedwick:2012:CMH

[Sed12e] Caitlin Sedwick. Coleen Murphy: How to stay young at heart, body, and mind. Journal of Cell Biology, 197(3):342–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
3/342.

Sedwick:2012:EWO

[Sed12f] Caitlin Sedwick. Ewald Weibel: an organelle of his very own, and more. Journal of Cell Biology, 197(1):4–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/197/1/4.

Sedwick:2012:FCN

[Sed12g] Caitlin Sedwick. Fernando Camargo: No limits to learning about stem cells. Journal of Cell Biology, 197(4):462–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
4/462.

Sedwick:2012:FIU

[Sed12h] Caitlin Sedwick. Fumiyo Ikeda: Ubiquitin lines up for action. Journal of Cell Biology, 199(3):405–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/3/405.

Sedwick:2012:GDS

[Sed12i] Caitlin Sedwick. Gaudenz Danuser: The symphony of the cell. Journal of Cell Biology, 199(4):568–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/199/4/
568.

Sedwick:2012:GRM

[Sed12j] Caitlin Sedwick. Graça Raposo: Melanosomes, more than skin deep. Journal of Cell Biology, 197(5):572–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
5/572.

Sedwick:2012:JLV

[Sed12k] Caitlin Sedwick. Jiri Lukas: Visualizing genome integrity maintenance. Journal of Cell Biology, 198(1):4–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/198/1/4.

Sedwick:2012:JHC

[Sed12l] Caitlin Sedwick. John Heuser: Capture the moment. Journal of Cell Biology, 199(6):868–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/6/868.

Sedwick:2012:KVG

[Sed12m] Caitlin Sedwick. Karen Vousden: Getting the big picture on p53. Journal of Cell Biology, 198(2):148–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
2/148.

Sedwick:2012:KBC

[Sed12n] Caitlin Sedwick. Keith Burridge: Cultivating knowledge on Rho. Journal of Cell Biology, 196(1):4–??, January 2012.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/196/1/4.

Sedwick:2012:LBS

[Sed12o] Caitlin Sedwick. Laurie Boyer: Stem cell circuitry for commitment. Journal of Cell Biology, 199(2):190–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
2/190.

Sedwick:2012:LHT

[Sed12p] Caitlin Sedwick. Li-huei Tsai: I well remember. Journal of Cell Biology, 199(7):1020–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/7/1020.

Sedwick:2012:MGB

[Sed12q] Caitlin Sedwick. Markus Grebe: Bringing plant polarity to light. Journal of Cell Biology, 197(7):854–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
7/854.

Sedwick:2012:OCF

[Sed12r] Caitlin Sedwick. Orna Cohen–Fix: Playing with nuclear morphology. Journal of Cell Biology, 196(5):550–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
5/550.

Sedwick:2012:PSS

[Sed12s] Caitlin Sedwick. Pamela Silver: Synthesizing a new biology. Journal of Cell Biology, 196(3):302–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
3/302.

Sedwick:2012:SGP

[Sed12t] Caitlin Sedwick. Sergio Grinstein: Phagocytosis step by step. Journal of Cell Biology, 196(4):392–??, February 2012. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/196/4/

Sedwick:2012:SBS

[Sed12u] Caitlin Sedwick. Stefano Bertuzzi: Science first. Journal of Cell Biology, 199(5):716–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/5/716.

Sedwick:2012:SBH

[Sed12v] Caitlin Sedwick. Sue Biggins: How kinetochores keep control of mitosis. Journal of Cell Biology, 196(6):668–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
6/668.

Sedwick:2012:TRC

[Sed12w] Caitlin Sedwick. Tannishtha Reya: Classic pathways, new views on cancer. Journal of Cell Biology, 198(5):766– ??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/5/766.

Sedwick:2012:TWF

[Sed12x] Caitlin Sedwick. Tobias Walther: Floating ideas on lipids. Journal of Cell Biology, 196(2):182–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/196/2/182.

Sedwick:2012:WLE

[Sed12y] Caitlin Sedwick. Wendell Lim: Exploring the path not chosen. Journal of Cell Biology, 198(6):956–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
6/956.

Sedwick:2012:YOA

[Sed12z] Caitlin Sedwick. Yoshinori Ohsumi: Autophagy from beginning to end. Journal of Cell Biology, 197(2):164–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
2/164.

Sedwick:2013:AYC

[Sed13a] Caitlin Sedwick. Alpha Yap: Cadherins hold it all together. Journal of Cell Biology, 201(5):648–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/5/648.

Sedwick:2013:AEA

[Sed13b] Caitlin Sedwick. Arnaud Echard: Adieu, ma soeur. Journal of Cell Biology, 203(6):868–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/6/868.

Sedwick:2013:BGC

[Sed13c] Caitlin Sedwick. Bob Goldstein: Cell biology by way of development. Journal of Cell Biology, 202(3):400–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
3/400.

Sedwick:2013:CPH

[Sed13d] Caitlin Sedwick. Carl-Philipp Heisenberg: Early embryos make a big move. Journal of Cell Biology, 200(3):238– ??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/200/3/238.

Sedwick:2013:CKS

[Sed13e] Caitlin Sedwick. Carla Koehler: Small TIMs are a big deal. Journal of Cell Biology, 201(3):358–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/3/358.

Sedwick:2013:CBL

[Sed13f] Caitlin Sedwick. Chris Bakal: Look and learn. Journal of Cell Biology, 203(3):378–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/3/378.

Sedwick:2013:DCR

[Sed13g] Caitlin Sedwick. Daniel Colón–Ramos: Observing and making connections. Journal of Cell Biology, 203(2):168–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
2/168.

Sedwick:2013:DSH

[Sed13h] Caitlin Sedwick. Didier Stainier: How function follows form. Journal of Cell Biology, 202(1):4–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/1/4.

Sedwick:2013:FBY

[Sed13i] Caitlin Sedwick. Frances Brodsky: So you think you know all about clathrin? Journal of Cell Biology, 202(6):830– ??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/6/830.

Sedwick:2013:GGH

[Sed13j] Caitlin Sedwick. Gillian Griffiths: How T cells get on target. Journal of Cell Biology, 200(1):4–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/1/4.

Sedwick:2013:HMH

[Sed13k] Caitlin Sedwick. Helder Maiato: Hot (+)TIPs on mitosis. Journal of Cell Biology, 202(5):722–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
5/722.

Sedwick:2013:JRE

[Sed13l] Caitlin Sedwick. Jody Rosenblatt: To extrude apically or basally, that is the question. Journal of Cell Biology, 202 (4):602–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/4/602.

Sedwick:2013:JBC

[Sed13m] Caitlin Sedwick. John Briggs: a closer look at HIV and coated vesicles. Journal of Cell Biology, 201(7):966–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
7/966.

Sedwick:2013:JRI

[Sed13n] Caitlin Sedwick. Jürgen Roth: Immunogold master. Journal of Cell Biology, 200(6):686–??, March 2013. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/6/686.

Sedwick:2013:MGI

[Sed13o] Caitlin Sedwick. Margaret Gardel: At the interface of physics and biology. Journal of Cell Biology, 201(2):170–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
2/170.

Sedwick:2013:MFB

[Sed13p] Caitlin Sedwick. Marilyn Farquhar: From the beginning. Journal of Cell Biology, 203(4):554–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/4/554.

Sedwick:2013:MHA

[Sed13q] Caitlin Sedwick. Martin Humphries: Attached to adhesion. Journal of Cell Biology, 200(5):554–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/5/554.

Sedwick:2013:MSL

[Sed13r] Caitlin Sedwick. Michael Sixt: Love the way they move. Journal of Cell Biology, 202(7):988–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/7/988.

Sedwick:2013:SRP

[Sed13s] Caitlin Sedwick. Samara Reck–Peterson: Dynein defies expectations. Journal of Cell Biology, 200(2):128–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/200/
2/128.

Sedwick:2013:SKB

[Sed13t] Caitlin Sedwick. Sandhya Koushika: Building new models and communities. Journal of Cell Biology, 201(1):4–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/201/1/4.

Sedwick:2013:SDP

[Sed13u] Caitlin Sedwick. Stephen Doxsey: At the (peri)center of things. Journal of Cell Biology, 200(4):364–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/200/
4/364.

Sedwick:2013:SEB

[Sed13v] Caitlin Sedwick. Suzanne Eaton: The beautiful logic of development. Journal of Cell Biology, 202(2):184–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
2/184.

Sedwick:2013:TSW

[Sed13w] Caitlin Sedwick. Trina Schroer: What’s cooking on dynactin. Journal of Cell Biology, 203(5):714–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
5/714.

Sedwick:2013:VAB

[Sed13x] Caitlin Sedwick. Victor Ambros: The broad scope of microRNAs. Journal of Cell Biology, 201(4):492–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/201/4/
492.

Sedwick:2013:YBM

[Sed13y] Caitlin Sedwick. Yohanns Bellaïche: Mastering multiscale morphology. Journal of Cell Biology, 203(1):4–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
1/4.

Sedwick:2013:YYC

[Sed13z] Caitlin Sedwick. Yukiko Yamashita: The centrosomes get there first. Journal of Cell Biology, 201(6):782–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
6/782.

Sedwick:2014:ACV

[Sed14a] Caitlin Sedwick. Adam Cohen: Visualizing cellular voltage. Journal of Cell Biology, 205(5):610–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/205/5/610.

Sedwick:2014:AMM

[Sed14b] Caitlin Sedwick. Alex Mogilner: Math illuminates biology. Journal of Cell Biology, 205(2):130–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/205/2/130.

Sedwick:2014:AGF

[Sed14c] Caitlin Sedwick. Amy Gladfelter: Fungi with a streak of individuality. Journal of Cell Biology, 204(4):464–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
4/464.

Sedwick:2014:BBA

[Sed14d] Caitlin Sedwick. Buzz Baum: The art of cell shape. Journal of Cell Biology, 206(3):332–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/206/3/332.

Sedwick:2014:CPM

[Sed14e] Caitlin Sedwick. Carole Parent: Migrating cells relay the message. Journal of Cell Biology, 205(3):286–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/205/3/
286.

Sedwick:2014:ECF

[Sed14f] Caitlin Sedwick. Elizabeth Chen: Fusing cells press closer. Journal of Cell Biology, 206(5):576–??, September 2014. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/206/5/
576.

Sedwick:2014:FMB

[Sed14g] Caitlin Sedwick. Fernando Martín–Belmonte: Epithelia embrace the space. Journal of Cell Biology, 205(6):756–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/205/
6/756.

Sedwick:2014:FPS

[Sed14h] Caitlin Sedwick. Feroz Papa: Saving cells from themselves. Journal of Cell Biology, 204(5):628–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/5/628.

Sedwick:2014:GGQ

[Sed14i] Caitlin Sedwick. Gohta Goshima: Questing for answers on the mitotic spindle. Journal of Cell Biology, 206(2):148–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
2/148.

Sedwick:2014:GON

[Sed14j] Caitlin Sedwick. Guangshuo Ou: New perspectives on Q cell biology. Journal of Cell Biology, 207(2):164–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
2/164.

Sedwick:2014:HMM

[Sed14k] Caitlin Sedwick. Heidi McBride: Mitochondria are well connected. Journal of Cell Biology, 206(4):454–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
4/454.

Sedwick:2014:JHM

[Sed14l] Caitlin Sedwick. Jonathon Howard: Motor proteins go walkabout. Journal of Cell Biology, 204(2):150–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
2/150.

Sedwick:2014:LMA

[Sed14m] Caitlin Sedwick. Laura Machesky: Actin opens the way. Journal of Cell Biology, 206(7):816–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/7/816.

Sedwick:2014:MRV

[Sed14n] Caitlin Sedwick. Margaret Robinson: Vesicles wear fancy coats. Journal of Cell Biology, 206(6):692–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
6/692.

Sedwick:2014:MFS

[Sed14o] Caitlin Sedwick. Matthew Freeman: The sharp end of rhomboids. Journal of Cell Biology, 205(1):4–??, April 2014. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/205/1/4.

Sedwick:2014:MSF

[Sed14p] Caitlin Sedwick. Melina Schuh: First comes the egg. Journal of Cell Biology, 204(7):1080–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/7/1080.

Sedwick:2014:MKT

[Sed14q] Caitlin Sedwick. Michael Kozlov: a twist in membrane physics. Journal of Cell Biology, 204(1):4–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/1/4.

Sedwick:2014:MMD

[Sed14r] Caitlin Sedwick. Miklós Müller: The deep history of eukaryotic metabolism. Journal of Cell Biology, 207(5):574–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
5/574.

Sedwick:2014:OLB

[Sed14s] Caitlin Sedwick. Ottoline Leyser: The beauty of plant genetics. Journal of Cell Biology, 204(3):284–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
3/284.

Sedwick:2014:PCN

[Sed14t] Caitlin Sedwick. Peter Cullen: Nexins have it sorted. Journal of Cell Biology, 205(4):432–??, May 2014. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/205/4/432.

Sedwick:2014:RVH

[Sed14u] Caitlin Sedwick. Raphael Valdivia: How Chlamydia settles in. Journal of Cell Biology, 207(1):4–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/207/1/4.

Sedwick:2014:SKN

[Sed14v] Caitlin Sedwick. Sally Kornbluth: Nature’s incredible contraptions. Journal of Cell Biology, 206(1):4–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/206/1/4.

Sedwick:2014:VGG

[Sed14w] Caitlin Sedwick. Valentina Greco: Got hair? Journal of Cell Biology, 207(4):436–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/4/436.

Sedwick:2014:YBL

[Sed14x] Caitlin Sedwick. Yves Barral: Lessons from yeast on growth, renewal, and old age. Journal of Cell Biology, 204(6):860–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
204/6/860.

Sevier:2010:NIO

[Sev10] Carolyn S. Sevier. New insights into oxidative folding. Journal of Cell Biology, 188(6):757–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/6/757.

Steed:2014:MCT

[SEV^{+}14] Emily Steed, Ahmed Elbediwy, Barbara Vacca, Sébastien Dupasquier, Sandra A. Hemkemeyer, Tesha Suddason, Ana C. Costa, Jean-Bernard Beaudry, Ceniz Zihni, Ewen Gallagher, Christophe E. Pierreux, Maria S. Balda, and Karl Matter. MarvelD3 couples tight junctions to the MEKK1–JNK pathway to regulate cell behavior and survival. Journal of Cell Biology, 204(5):821–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/5/821.

Serrels:2012:FTW

[SF12] Bryan Serrels and Margaret C. Frame. FAK and talin: Who is taking whom to the integrin engagement party? Journal of Cell Biology, 196(2):185–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/2/185.

Selvaraj:2012:LAF

[SFB^{+}12] Bhuvaneish Thangaraj Selvaraj, Nicolas Frank, Florian L. P. Bender, Esther Asan, and Michael Sendtner. Local axonal function of STAT3 rescues axon degeneration in the pmn model of motoneuron disease. Journal of Cell Biology, 199 (3):437–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/3/437.

Sahin:2014:OSI

[SFJ^{+}14] Umut Sahin, Omar Ferhi, Marion Jeanne, Shirine Benhenda, Caroline Berthier, Florence Jollivet, Michiko Niwa-Kawakita, Orestis Faklaris, Niclas Setterblad, Hugues de Thé, and Valérie Lallemand-Breitenbach. Oxidative stress–induced assembly of PML nuclear bodies controls sumoylation of partner proteins. Journal of Cell Biology, 204(6):931–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/6/931.

Schweizer:2013:SAC

[SFK^{+}13] Nina Schweizer, Cristina Ferrás, David M. Kern, Elsa Logarinho, Iain M. Cheeseman, and Helder Maiato. Spindle assembly checkpoint robustness requires Tpr-mediated regulation of Mad1/Mad2 proteostasis. Journal of Cell Biology, 203(6): 883–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/6/883.

Sumigray:2012:NMT

[SFL12] Kaelyn D. Sumigray, Henry P. Foote, and Terry Lechler. Noncentrosomal microtubules and type II myosins potentiate epidermal cell adhesion and barrier formation. Journal of Cell Biology, 199(3):513–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/3/513.

Salaun:2010:IDP

[SGC10] Christine Salaun, Jennifer Greaves, and Luke H. Chamberlain. The intracellular dynamic of protein palmitoylation. Journal of Cell Biology, 191(7):1229–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/7/1229.

Sun:2010:MTR

[SGD^{+}10] Yuting Sun, Yejing Ge, Jenny Drnevich, Yong Zhao, Mark Band, and Jie Chen. Mammalian target of rapamycin regulates miRNA-1 and follistatin in skeletal myogenesis. Journal of Cell Biology, 189(7):1157–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/7/1157.

Schoumacher:2010:AMV

[SGLV10] Marie Schoumacher, Robert D. Goldman, Daniel Louvard, and Danijela M. Vignjevic. Actin, microtubules, and vimentin intermediate filaments cooperate for elongation of invadopodia. Journal of Cell Biology, 189(3):541–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/3/541.

Sugiyama:2013:ECM

[SGT^{+}13] Nami Sugiyama, Erika Gucciardo, Olga Tatti, Markku Varjosalo, Marko Hyytiäinen, Matthias Gstaiger, and Kaisa Lehti. EphA2 cleavage by MT1-MMP triggers single cancer cell invasion via homotypic cell repulsion. Journal of Cell Biology, 201 (3):467–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/3/467.

Shah:2010:CTS

[Sha10] Jagesh V. Shah. Cells in tight spaces: the role of cell shape in cell function. Journal of Cell Biology, 191(2):233–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
2/233.

Steinberg:2010:CCS

[SHB^{+}10] Benjamin E. Steinberg, Kassidy K. Huynh, Alexandre Brodovitch, Sabrina Jabs, Tobias Stauber, Thomas J. Jentsch, and Sergio Grinstein. A cation counterflux supports lysosomal acidification. Journal of Cell Biology, 189(7):1171–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
7/1171.

Steinberg:2012:SPI

[SHBC12] Florian Steinberg, Kate J. Heesom, Mark D. Bass, and Peter J. Cullen. SNX17 protects integrins from degradation by sorting between lysosomal and recycling pathways. Journal of Cell Biology, 197(2):219–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/2/219.

Scott:2010:LKR

[SHC^{+}10] Rebecca W. Scott, Steven Hooper, Diane Crighton, Ang Li, Ireen König, June Munro, Elisabeth Trivier, Grant Wickman, Pierre Morin, Daniel R. Croft, John Dawson, Laura Machesky, Kurt I. Anderson, Erik A. Sahai, and Michael F. Olson. LIM kinases are required for invasive path generation by tumor and tumor-associated stromal cells. Journal of Cell Biology, 191 (1):169–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/1/169.

Sosanya:2013:DHA

[SHC^{+}13] Natasha M. Sosanya, Peggy P. C. Huang, Luisa P. Cacheaux, Chun Jung Chen, Kathleen Nguyen, Nora I. Perrone-Bizzozero, and Kimberly F. Raab-Graham. Degradation of high affinity HuD targets releases Kv1.1 mRNA from miR-129 repression by mTORC1. Journal of Cell Biology, 202(1):53–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
202/1/53.

Sheng:2014:MTA

[She14] Zu-Hang Sheng. Mitochondrial trafficking and anchoring in neurons: New insight and implications. Journal of Cell Biology, 204(7):1087–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/7/1087.

Schermelleh:2010:GSR

[SHL10] Lothar Schermelleh, Rainer Heintzmann, and Heinrich Leonhardt. A guide to super-resolution fluorescence microscopy. Journal of Cell Biology, 190(2):165–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/2/165.

Suzuki:2011:SMG

[SHN^{+}11] Aussie Suzuki, Tetsuya Hori, Tatsuya Nishino, Jiro Usukura, Atsushi Miyagi, Kosuke Morikawa, and Tatsuo Fukagawa. Spindle microtubules generate tension-dependent changes in the distribution of inner kinetochore proteins. Journal of Cell Biology, 193(1):125–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/1/125.

Short:2010:SLA

[Sho10a] Ben Short. α-synuclein leaves autophagy feeling compromised. Journal of Cell Biology, 190(6):??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/6/944.3.

Short:2010:CGH

[Sho10b] Ben Short. β-catenin gets an honorable discharge. Journal of Cell Biology, 190(6):945–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/6/945.

Short:2010:TKC

[Sho10c] Ben Short. γ-tubulin keeps the cell cycling. Journal of Cell Biology, 190(3):??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/3/280.1.

Short:2010:AML

[Sho10d] Ben Short. Acetylated microtubules let the ER slide. Journal of Cell Biology, 190(3):??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/190/3/280.2.

Short:2010:ASD

[Sho10e] Ben Short. Acid sphingomyelinase deals the seal. Journal of Cell Biology, 189(6):??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/6/920.3.

Short:2010:ATV

[Sho10f] Ben Short. The acid test of v–ATPase function. Journal of Cell Biology, 189(5):773–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/5/773.

Short:2010:AOH

[Sho10g] Ben Short. Actin oligomers hit the assembly line. Journal of Cell Biology, 188(6):753–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/6/753.

Short:2010:AMC

[Sho10h] Ben Short. Arf and Miz1 cause cells to lose their grip. Journal of Cell Biology, 188(6):??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/6/752.3.

Short:2010:ATC

[Sho10i] Ben Short. Augmin-ting the central spindle. Journal of Cell Biology, 191(2):??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/2/226.1.

Short:2010:AMH

[Sho10j] Ben Short. Autophagy membranes held in reserve. Journal of Cell Biology, 190(6):??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/190/6/944.2.

Short:2010:BKH

[Sho10k] Ben Short. Boi keeps Hedgehog close to home. Journal of Cell Biology, 191(5):??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/5/894.1.

Short:2010:CAN

[Sho10l] Ben Short. Casting the autophagy net. Journal of Cell Biology, 190(4):483–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/4/483.

Short:2010:CGR

[Sho10m] Ben Short. Cdc14 in good repair. Journal of Cell Biology, 189(4):??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/4/614.1.

Short:2010:CFA

[Sho10n] Ben Short. Cells find an alternative exit. Journal of Cell Biology, 191(4):??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/4/694.2.

Short:2010:CPF

[Sho10o] Ben Short. Cep152 and Plk4 form a double act. Journal of Cell Biology, 191(4):695–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/4/695.

Short:2010:CJW

[Sho10p] Ben Short. Christine Jacobs–Wagner: Drawing the bacterial organizational chart. Journal of Cell Biology, 189(3):390–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
189/3/390.

Short:2010:CGS

[Sho10q] Ben Short. The cytoplasm gets some new threads. Journal of Cell Biology, 190(4):??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/4/482.1.

Short:2010:DKF

[Sho10r] Ben Short. Daniel Klionsky: a full plate for autophagy. Journal of Cell Biology, 189(1):8–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/1/8.

Short:2010:DWE

[Sho10s] Ben Short. Deciding what to eat. Journal of Cell Biology, 189 (2):191–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/2/191.

Short:2010:DFF

[Sho10t] Ben Short. Deconstructing FAK function. Journal of Cell Biology, 189(6):??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/6/920.2.

Short:2010:DTM

[Sho10u] Ben Short. Disuse TWEAKs muscle loss. Journal of Cell Biology, 188(6):??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/6/752.2.

Short:2010:ETD

[Sho10v] Ben Short. EGF takes dual control of mRNA. Journal of Cell Biology, 188(3):??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/3/300.1.

Short:2010:ECD

[Sho10w] Ben Short. Epithelial cells on death Rho. Journal of Cell Biology, 189(2):??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/2/190.2.

Short:2010:FKS

[Sho10x] Ben Short. FLIPping the kill switch. Journal of Cell Biology, 190(3):281–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/3/281.

Short:2010:IPH

[Sho10y] Ben Short. Gli inhibitor pricked by Hedgehog. Journal of Cell Biology, 191(2):??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/2/226.3.

Short:2010:GKS

[Sho10z] Ben Short. Glycosphingolipids keep signaling in top–Notch condition. Journal of Cell Biology, 188(4):447–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
4/447.

Short:2010:HVH

[Sho10-27] Ben Short. How Vps41 HOPS between tethering functions. Journal of Cell Biology, 191(4):??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/4/694.3.

Short:2010:IPS

[Sho10-28] Ben Short. Identifying the prime suspects in vesicle release. Journal of Cell Biology, 188(3):??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/188/3/300.3.

Short:2010:SLT

[Sho10-29] Ben Short. In search of lost timing. Journal of Cell Biology, 189(6):921–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/6/921.

Short:2010:IRA

[Sho10-30] Ben Short. The integrin reconstruction act. Journal of Cell Biology, 188(1):3–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/1/3.

Short:2010:KYE

[Sho10-31] Ben Short. Kenneth Yamada: Exploring the paths of cell migration. Journal of Cell Biology, 188(2):178–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
2/178.

Short:2010:KHR

[Sho10-32] Ben Short. Kinetochores hang on for the ride. Journal of Cell Biology, 189(4):615–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/4/615.

Short:2010:LCG

[Sho10-33] Ben Short. A little Cdc20 goes a long way. Journal of Cell Biology, 191(2):227–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/2/227.

Short:2010:LAR

[Sho10-34] Ben Short. Long asters rock the spindle. Journal of Cell Biology, 190(1):??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/1/2.1.

Short:2010:MHT

[Sho10-35] Ben Short. Martin Hetzer: Taking the nuclear membrane beyond the barrier. Journal of Cell Biology, 190(4):484–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
190/4/484.

Short:2010:MNT

[Sho10-36] Ben Short. Maxence Nachury: a transporting view of the primary cilium. Journal of Cell Biology, 191(3):436–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
3/436.

Short:2010:MCC

[Sho10-37] Ben Short. Migrating cells CLIC into gear. Journal of Cell Biology, 190(4):??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/4/482.3.

Short:2010:MGC

[Sho10-38] Ben Short. MIM gives cells a sense of direction. Journal of Cell Biology, 189(2):??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/2/190.3.

Short:2010:MBD

[Sho10-39] Ben Short. Mónica Bettencourt–Dias: Centered on centrioles. Journal of Cell Biology, 190(5):710–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
5/710.

Short:2010:MPR

[Sho10-40] Ben Short. Myo1 provides a ring in the tail. Journal of Cell Biology, 191(7):??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/7/1220.3.

Short:2010:MIC

[Sho10-41] Ben Short. Myosin II contracts out GTPase activation. Journal of Cell Biology, 190(4):??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/190/4/482.2.

Short:2010:NDM

[Sho10-42] Ben Short. Nebulin doesn’t measure up. Journal of Cell Biology, 189(5):??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/5/772.3.

Short:2010:NFO

[Sho10-43] Ben Short. New function for an old master. Journal of Cell Biology, 188(1):??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/1/2.1.

Short:2010:PHM

[Sho10-44] Ben Short. p53 hides a multitude of CIN. Journal of Cell Biology, 188(3):301–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/3/301.

Short:2010:PSC

[Sho10-45] Ben Short. p53 strands cell cycle on archipelago. Journal of Cell Biology, 188(4):??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/4/446.2.

Short:2010:PRD

[Sho10-46] Ben Short. Parkin restrictions for damaged mitochondria. Journal of Cell Biology, 189(4):??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/4/614.2.

Short:2010:PPG

[Sho10-47] Ben Short. PIH proteins give dynein arms a hand. Journal of Cell Biology, 190(1):??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/1/2.3.

Short:2010:PPP

[Sho10-48] Ben Short. PIPs propel phagosomes. Journal of Cell Biology, 191(5):??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/5/894.3.

Short:2010:PFT

[Sho10-49] Ben Short. Plakophilin 1 is found in translation. Journal of Cell Biology, 188(4):??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/4/446.1.

Short:2010:PMC

[Sho10-50] Ben Short. Polyglutamylation makes the cut. Journal of Cell Biology, 189(6):??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/6/920.1.

Short:2010:PPB

[Sho10-51] Ben Short. PP6 puts the brakes on spindle assembly. Journal of Cell Biology, 191(7):1221–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/7/1221.

Short:2010:RGEa

[Sho10-52] Ben Short. Rab GEFs emerge from their DENN. Journal of Cell Biology, 191(2):??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/2/226.2.

Short:2010:RDE

[Sho10-53] Ben Short. Rab35 drives exosome secretion. Journal of Cell Biology, 189(2):??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/2/190.1.

Short:2010:RGEb

[Sho10-54] Ben Short. Rb gets ERK’d by the competition. Journal of Cell Biology, 191(5):??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/5/894.2.

Short:2010:SBM

[Sho10-55] Ben Short. Sar1 bends membranes into shape. Journal of Cell Biology, 190(1):3–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/1/3.

Short:2010:SMR

[Sho10-56] Ben Short. Sean Morrison: a root and branch approach to stem cells. Journal of Cell Biology, 189(7):1056–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
7/1056.

Short:2010:SUR

[Sho10-57] Ben Short. Segregating out UbcH10’s role in tumorigenesis. Journal of Cell Biology, 188(1):??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/188/1/2.2.

Short:2010:SSC

[Sho10-58] Ben Short. Setting the spindle’s compass. Journal of Cell Biology, 191(5):895–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/5/895.

Short:2010:SPM

[Sho10-59] Ben Short. A SOLO performance in meiotic cohesion. Journal of Cell Biology, 188(3):??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/3/300.2.

Short:2010:SSN

[Sho10-60] Ben Short. SPSB2 sets NO limits. Journal of Cell Biology, 190(1):??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/1/2.2.

Short:2010:SDC

[Sho10-61] Ben Short. SRGP-1 drives cell junctions round the bend. Journal of Cell Biology, 191(4):??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/4/694.1.

Short:2010:SG

[Sho10-62] Ben Short. The story of O-glycosylation. Journal of Cell Biology, 189(5):??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/5/772.2.

Short:2010:SMM

[Sho10-63] Ben Short. Stress management for mRNAs. Journal of Cell Biology, 189(5):??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/5/772.1.

Short:2010:SVW

[Sho10-64] Ben Short. Synaptic vesicles are well connected. Journal of Cell Biology, 188(1):??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/1/2.3.

Short:2010:SAN

[Sho10-65] Ben Short. Syndecan-3 adds a Notch to its belt. Journal of Cell Biology, 190(3):??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/3/280.3.

Short:2010:TRK

[Sho10-66] Ben Short. A team of rivals at the kinetochore. Journal of Cell Biology, 188(6):??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/6/752.1.

Short:2010:TGS

[Sho10-67] Ben Short. Telomeres get SIRT-ified. Journal of Cell Biology, 191(7):??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/7/1220.2.

Short:2010:TPM

[Sho10-68] Ben Short. A turning point for macrophages. Journal of Cell Biology, 189(4):??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/4/614.3.

Short:2010:UCF

[Sho10-69] Ben Short. Unpacking condensins’ function in ES cells. Journal of Cell Biology, 188(4):??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/4/446.3.

Short:2010:YGM

[Sho10-70] Ben Short. Yukiko Goda: Memories are made of this. Journal of Cell Biology, 190(3):282–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/3/282.

Short:2010:ZWZ

[Sho10-71] Ben Short. Zfp521 works its zinc fingers to the bone. Journal of Cell Biology, 191(7):??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/7/1220.1.

Short:2011:AAH

[Sho11a] Ben Short. AMPK amplifies Huntington’s disease. Journal of Cell Biology, 194(2):??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/2/160.1.

Short:2011:AAC

[Sho11b] Ben Short. Angelika Amon: Conquering the divide. Journal of Cell Biology, 193(2):254–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/2/254.

Short:2011:ADR

[Sho11c] Ben Short. Arf1 doubles up to release vesicles. Journal of Cell Biology, 194(5):??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/5/658.3.

Short:2011:ALC

[Sho11d] Ben Short. Aurora A limits calcium release. Journal of Cell Biology, 193(6):??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/6/954.3.

Short:2011:BIB

[Sho11e] Ben Short. Bigger isn’t better. Journal of Cell Biology, 193(4):??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/4/596.3.

Short:2011:BM

[Sho11f] Ben Short. BRCA1 on the move. Journal of Cell Biology, 192 (3):369–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/3/369.

Short:2011:BGB

[Sho11g] Ben Short. Bridging the gap between atlastin conformations. Journal of Cell Biology, 195(4):??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/4/538.1.

Short:2011:BYS

[Sho11h] Ben Short. Budding yeast star in their own biofilm. Journal of Cell Biology, 194(5):659–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/5/659.

Short:2011:CTB

[Sho11i] Ben Short. Cadherins trade bonds. Journal of Cell Biology, 192 (6):903–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/6/903.

Short:2011:CTM

[Sho11j] Ben Short. Cdc42 and Tinman march to the same beat. Journal of Cell Biology, 193(7):1133–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/7/1133.

Short:2011:CPT

[Sho11k] Ben Short. Cdks put Treslin on TopBP1. Journal of Cell Biology, 193(6):??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/6/954.1.

Short:2011:CDH

[Sho11l] Ben Short. Cell death helps give closure. Journal of Cell Biology, 195(6):??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/6/922.3.

Short:2011:CRT

[Sho11m] Ben Short. A COG in the retrograde transport machinery. Journal of Cell Biology, 194(3):??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/3/350.3.

Short:2011:CCS

[Sho11n] Ben Short. Cohesin and condensin spring into action. Journal of Cell Biology, 193(7):??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/7/1132.2.

Short:2011:CCD

[Sho11o] Ben Short. Connecting the cytokinesis dots. Journal of Cell Biology, 192(6):??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/6/902.2.

Short:2011:CPH

[Sho11p] Ben Short. CUPS provide a handle on Acb1 secretion. Journal of Cell Biology, 195(6):??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/6/922.2.

Short:2011:DMN

[Sho11q] Ben Short. DLK makes neuronal cutbacks. Journal of Cell Biology, 194(5):??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/5/658.2.

Short:2011:ECN

[Sho11r] Ben Short. Epac1 cAMPs out at nuclear pores. Journal of Cell Biology, 193(6):??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/6/954.2.

Short:2011:FAI

[Sho11s] Ben Short. A fifth amendment to the intestine’s constitution. Journal of Cell Biology, 192(5):??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/5/706.2.

Short:2011:MTH

[Sho11t] Ben Short. For Myo4p, two heads are better than one. Journal of Cell Biology, 195(4):??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/4/538.2.

Short:2011:GHM

[Sho11u] Ben Short. Good housekeeping maintains a healthy liver. Journal of Cell Biology, 195(2):??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/2/166.1.

Short:2011:HMA

[Sho11v] Ben Short. Harvey McMahon: Ahead of the curve on membrane dynamics. Journal of Cell Biology, 193(4):598–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
4/598.

Short:2011:HMP

[Sho11w] Ben Short. HIV-1 makes a pore adjustment. Journal of Cell Biology, 193(4):??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/4/596.1.

Short:2011:HPC

[Sho11x] Ben Short. HJURP puts the centromere in place. Journal of Cell Biology, 194(2):??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/2/160.2.

Short:2011:HSS

[Sho11y] Ben Short. How the ER shapes up and ships out. Journal of Cell Biology, 193(2):??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/2/252.2.

Short:2011:HCG

[Sho11z] Ben Short. Howard Chang: Gene regulation in time and space. Journal of Cell Biology, 195(1):4–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/1/4.

Short:2011:IAP

[Sho11-27] Ben Short. Imaging β amyloid’s pore performance. Journal of Cell Biology, 195(3):345–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/3/345.

Short:2011:IKE

[Sho11-28] Ben Short. IRSp53 keeps ECM signaling up to Par. Journal of Cell Biology, 192(3):??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/3/368.3.

Short:2011:LDF

[Sho11-29] Ben Short. Lipid droplets fatten up with Fsp27. Journal of Cell Biology, 195(6):??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/6/922.1.

Short:2011:LMS

[Sho11-30] Ben Short. Locating a mitochondrial scaffold on the map. Journal of Cell Biology, 195(2):167–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/2/167.

Short:2011:LE

[Sho11-31] Ben Short. Location is everything. Journal of Cell Biology, 192(6):??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/6/902.1.

Short:2011:MWM

[Sho11-32] Ben Short. Matthew Welch: The many branches of actin regulation. Journal of Cell Biology, 192(2):206–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
2/206.

Short:2011:MFS

[Sho11-33] Ben Short. Microtubules follow septins’ guidelines. Journal of Cell Biology, 194(2):161–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/2/161.

Short:2011:MPC

[Sho11-34] Ben Short. miRNAs prevent a change of heart. Journal of Cell Biology, 193(7):??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/7/1132.3.

Short:2011:MPB

[Sho11-35] Ben Short. Mutant phagosomes bite off more than they can chew. Journal of Cell Biology, 192(3):??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
3/368.1.

Short:2011:MIG

[Sho11-36] Ben Short. Myosin II gets polarity back to front. Journal of Cell Biology, 193(2):??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/2/252.3.

Short:2011:NSR

[Sho11-37] Ben Short. A new spin on radial spokes. Journal of Cell Biology, 195(4):??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/4/538.3.

Short:2011:NTS

[Sho11-38] Ben Short. New TIPs for successfully growing microtubules. Journal of Cell Biology, 193(6):955–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/6/955.

Short:2011:NES

[Sho11-39] Ben Short. Nuclear envelope starts with a clean sheet. Journal of Cell Biology, 194(3):351–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/3/351.

Short:2011:OBI

[Sho11-40] Ben Short. Osteoblasts are bone idle without Frizzled-9. Journal of Cell Biology, 192(6):??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/192/6/902.3.

Short:2011:PRK

[Sho11-41] Ben Short. PARP retains Ku at double-strand breaks. Journal of Cell Biology, 194(3):??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/3/350.1.

Short:2011:POB

[Sho11-42] Ben Short. Passengers operate the brakes. Journal of Cell Biology, 193(2):??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/2/252.1.

Short:2011:PCH

[Sho11-43] Ben Short. PKA cuts down on HDAC4 signaling. Journal of Cell Biology, 195(3):??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/3/344.2.

Short:2011:PGN

[Sho11-44] Ben Short. PLD gives nutritional advice to mTORC1. Journal of Cell Biology, 195(3):??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/3/344.3.

Short:2011:RMS

[Sho11-45] Ben Short. Rad54 mutant stays rooted to the spot. Journal of Cell Biology, 192(5):??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/5/706.1.

Short:2011:RDS

[Sho11-46] Ben Short. Ran drags spindles through the Mud. Journal of Cell Biology, 195(3):??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/3/344.1.

Short:2011:RDP

[Sho11-47] Ben Short. Repair defects put stem cells in a fix. Journal of Cell Biology, 193(2):253–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/2/253.

Short:2011:RAK

[Sho11-48] Ben Short. Retinoic acid keeps Lingo-1 quiet. Journal of Cell Biology, 193(7):??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/7/1132.1.

Short:2011:RTG

[Sho11-49] Ben Short. RNA targeting gets competitive. Journal of Cell Biology, 194(3):??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/3/350.2.

Short:2011:SCC

[Sho11-50] Ben Short. Schwann cells cover their territory. Journal of Cell Biology, 195(2):??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/2/166.3.

Short:2011:SMR

[Sho11-51] Ben Short. Sean Munro: Revealing the Golgi’s true identity. Journal of Cell Biology, 192(1):4–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/1/4.

Short:2011:SNS

[Sho11-52] Ben Short. Setting a new standard for kinetochores. Journal of Cell Biology, 195(4):539–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/4/539.

Short:2011:SRS

[Sho11-53] Ben Short. Shs1 rounds out the septin repertoire. Journal of Cell Biology, 195(6):923–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/6/923.

Short:2011:SHT

[Sho11-54] Ben Short. Sir3 helps telomeres stick together. Journal of Cell Biology, 192(3):??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/3/368.2.

Short:2011:SPS

[Sho11-55] Ben Short. Sphingolipid puts the squeeze on apoptotic cells. Journal of Cell Biology, 193(4):??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/4/596.2.

Short:2011:SHB

[Sho11-56] Ben Short. Sphingomyelinase helps bones get their minerals. Journal of Cell Biology, 194(2):??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/2/160.3.

Short:2011:TIS

[Sho11-57] Ben Short. Titin isn’t a sleeping giant. Journal of Cell Biology, 193(4):597–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/4/597.

Short:2011:TFP

[Sho11-58] Ben Short. The Tudor family produces heirs. Journal of Cell Biology, 192(5):??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/5/706.3.

Short:2011:TAF

[Sho11-59] Ben Short. Twins activate! Form of a new centriole! Journal of Cell Biology, 195(2):??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/2/166.2.

Short:2011:VLC

[Sho11-60] Ben Short. Vpr loosens chromatid ties. Journal of Cell Biology, 194(5):??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/5/658.1.

Short:2011:WS

[Sho11-61] Ben Short. The wrong suspect? Journal of Cell Biology, 192 (5):707–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/5/707.

Short:2012:TFG

[Sho12a] Ben Short. γ-tubulin flips the G1/S switch. Journal of Cell Biology, 198(5):??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/5/764.1.

Short:2012:TSC

[Sho12b] Ben Short. γ-tubulin stands up to be counted. Journal of Cell Biology, 197(1):??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/1/2.2.

Short:2012:ACG

[Sho12c] Ben Short. The APC/C gets Trim’d down for apoptosis. Journal of Cell Biology, 197(3):??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/197/3/340.1.

Short:2012:ARD

[Sho12d] Ben Short. Apoptotic regulators dispose of the midbody. Journal of Cell Biology, 199(7):??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/7/1018.1.

Short:2012:ASW

[Sho12e] Ben Short. ARL-13 SUMO wrestles receptors into cilia. Journal of Cell Biology, 199(4):??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/4/566.1.

Short:2012:BET

[Sho12f] Ben Short. Big enough for two. Journal of Cell Biology, 197(1):3–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/1/3.

Short:2012:BSS

[Sho12g] Ben Short. BubR1 shows self-control. Journal of Cell Biology, 198(2):??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/2/146.3.

Short:2012:BAA

[Sho12h] Ben Short. Building up actin at adherens junctions. Journal of Cell Biology, 196(1):3–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/1/3.

Short:2012:CHC

[Sho12i] Ben Short. Cdc42 helps cancer cells make their exit. Journal of Cell Biology, 199(4):567–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/4/567.

Short:2012:CPS

[Sho12j] Ben Short. Centrosomes are PKA’s sensitive spot. Journal of Cell Biology, 198(4):??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/4/472.3.

Short:2012:CHM

[Sho12k] Ben Short. Clathrin helps maintain the centrosome’s integrity. Journal of Cell Biology, 198(4):473–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/4/473.

Short:2012:DC

[Sho12l] Ben Short. Directing a chaperone. Journal of Cell Biology, 196 (6):667–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/6/667.

Short:2012:DRS

[Sho12m] Ben Short. Dnm1’s recruitment strategy. Journal of Cell Biology, 199(4):??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/4/566.2.

Short:2012:ECH

[Sho12n] Ben Short. An enzyme complex helps fatten up lipid droplets. Journal of Cell Biology, 198(5):??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/5/764.3.

Short:2012:SPM

[Sho12o] Ben Short. ER stress PERKs up an miRNA. Journal of Cell Biology, 196(6):??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/6/666.1.

Short:2012:FGB

[Sho12p] Ben Short. Farp1 gives both sides of the story. Journal of Cell Biology, 199(6):867–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/6/867.

Short:2012:FZR

[Sho12q] Ben Short. The fellowship of the Z ring. Journal of Cell Biology, 199(4):??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/4/566.3.

Short:2012:FUM

[Sho12r] Ben Short. A firmer understanding of muscle fibrosis. Journal of Cell Biology, 196(1):??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/1/2.3.

Short:2012:FAD

[Sho12s] Ben Short. Focal adhesions degrade the ECM. Journal of Cell Biology, 196(3):??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/3/300.3.

Short:2012:FMS

[Sho12t] Ben Short. Formin’ muscle sarcomeres. Journal of Cell Biology, 198(1):??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/1/2.3.

Short:2012:FLI

[Sho12u] Ben Short. The Fuzzy logistics of intraflagellar transport. Journal of Cell Biology, 198(1):??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/1/2.1.

Short:2012:FR

[Sho12v] Ben Short. Fyn regeneration. Journal of Cell Biology, 199 (2):189–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/2/189.

Short:2012:GSR

[Sho12w] Ben Short. Global sourcing for ring construction. Journal of Cell Biology, 199(5):??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/5/714.3.

Short:2012:GSF

[Sho12x] Ben Short. GW220 seals the fate of mRNAs. Journal of Cell Biology, 198(4):??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/4/472.2.

Short:2012:HEJ

[Sho12y] Ben Short. How endothelial junctions get deep–Syx’d. Journal of Cell Biology, 199(7):??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/7/1018.3.

Short:2012:HMC

[Sho12z] Ben Short. How meiotic chromosomes meet their match. Journal of Cell Biology, 196(1):??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/196/1/2.2.

Short:2012:IGB

[Sho12-27] Ben Short. IFT-A goes both ways. Journal of Cell Biology, 197(6):??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/6/692.3.

Short:2012:KHM

[Sho12-28] Ben Short. KIF4 helps mitotic chromosomes get in shape. Journal of Cell Biology, 199(5):715–??, November 2012. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/199/5/
715.

Short:2012:LDA

[Sho12-29] Ben Short. Leukocytes drop anchor. Journal of Cell Biology, 197(1):??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/1/2.3.

Short:2012:LTD

[Sho12-30] Ben Short. Lkb1 takes different paths to morphogenesis. Journal of Cell Biology, 199(7):1019–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/7/1019.

Short:2012:MN

[Sho12-31] Ben Short. Meet the neighbors. Journal of Cell Biology, 196(6):??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/6/666.3.

Short:2012:MSO

[Sho12-32] Ben Short. Microvillar scaffold is only temporary. Journal of Cell Biology, 198(2):??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/2/146.2.

Short:2012:MKO

[Sho12-33] Ben Short. miR-34s keep osteoblasts bone idle. Journal of Cell Biology, 197(4):461–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/4/461.

Short:2012:MMS

[Sho12-34] Ben Short. Mitochondrial morphology shapes differentiation. Journal of Cell Biology, 197(4):??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/4/460.1.

Short:2012:MAN

[Sho12-35] Ben Short. Myelinated axons need a 4.1G connection. Journal of Cell Biology, 196(3):??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/3/300.2.

Short:2012:MMC

[Sho12-36] Ben Short. The mystery of the (not) missing coenzyme. Journal of Cell Biology, 199(2):??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/2/188.1.

Short:2012:OEG

[Sho12-37] Ben Short. Opening up the ER’s gatekeeper. Journal of Cell Biology, 199(6):??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/6/866.2.

Short:2012:OCT

[Sho12-38] Ben Short. Organelles compete for their inheritance. Journal of Cell Biology, 198(1):??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/1/2.2.

Short:2012:OCS

[Sho12-39] Ben Short. Ovastacin cuts off sperm binding. Journal of Cell Biology, 197(1):??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/1/2.1.

Short:2012:PPS

[Sho12-40] Ben Short. PI(3,5)P_{2} sows the seeds of plant growth. Journal of Cell Biology, 198(2):147–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/2/147.

Short:2012:PPY

[Sho12-41] Ben Short. Prion puts yeast cells under arrest. Journal of Cell Biology, 197(3):??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/3/340.2.

Short:2012:PDO

[Sho12-42] Ben Short. Protrusion drives osteoclast fusion. Journal of Cell Biology, 197(4):??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/4/460.3.

Short:2012:PPP

[Sho12-43] Ben Short. Protrusion provides a prognosis. Journal of Cell Biology, 197(6):??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/6/692.1.

Short:2012:PCC

[Sho12-44] Ben Short. Putting a “cap” on clathrin-mediated endocytosis. Journal of Cell Biology, 197(4):??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/4/460.2.

Short:2012:PSS

[Sho12-45] Ben Short. Putting SUMO and SoxE into context. Journal of Cell Biology, 198(5):765–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/5/765.

Short:2012:PSR

[Sho12-46] Ben Short. A pyrimidine scheme to restore mRNA export. Journal of Cell Biology, 196(3):??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/196/3/300.1.

Short:2012:RKW

[Sho12-47] Ben Short. Rad53 keeps watch over mitochondrial DNA. Journal of Cell Biology, 198(5):??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/5/764.2.

Short:2012:RND

[Sho12-48] Ben Short. Release notes for dense-core vesicles. Journal of Cell Biology, 199(6):??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/6/866.1.

Short:2012:RRS

[Sho12-49] Ben Short. The ribosome runs a systems test. Journal of Cell Biology, 197(6):??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/6/692.2.

Short:2012:SCA

[Sho12-50] Ben Short. Schwann cells aid regeneration project. Journal of Cell Biology, 198(1):3–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/1/3.

Short:2012:SRM

[Sho12-51] Ben Short. Sds22 and Repo–Man keep anaphase chromosomes moving. Journal of Cell Biology, 198(2):??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/2/146.1.

Short:2012:STS

[Sho12-52] Ben Short. Sec and Tat share the workload. Journal of Cell Biology, 199(2):??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/2/188.2.

Short:2012:SPH

[Sho12-53] Ben Short. Secretory proteins hail a cab at the TGN. Journal of Cell Biology, 199(7):??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/7/1018.2.

Short:2012:SBM

[Sho12-54] Ben Short. Segregating Bub1’s mitotic functions. Journal of Cell Biology, 199(6):??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/6/866.3.

Short:2012:SEC

[Sho12-55] Ben Short. A SIMPLE explanation for Charcot–Marie– Tooth disease. Journal of Cell Biology, 199(5):??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
5/714.2.

Short:2012:SS

[Sho12-56] Ben Short. Simulating segregation. Journal of Cell Biology, 196(6):??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/6/666.2.

Short:2012:SFG

[Sho12-57] Ben Short. Stress fibers guide focal adhesions to maturity. Journal of Cell Biology, 196(3):301–??, February 2012. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/196/3/
301.

Short:2012:TCI

[Sho12-58] Ben Short. Talin is cut out for intercellular adhesion. Journal of Cell Biology, 197(6):693–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/197/6/693.

Short:2012:TIT

[Sho12-59] Ben Short. Tiam1 increases turnover. Journal of Cell Biology, 199(2):??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/2/188.3.

Short:2012:TKP

[Sho12-60] Ben Short. Trichoplein keeps primary cilia silent. Journal of Cell Biology, 197(3):341–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/197/3/341.

Short:2012:UIU

[Sho12-61] Ben Short. Ubiquitin isn’t ubiquitous in receptor trafficking. Journal of Cell Biology, 197(3):??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/3/340.3.

Short:2012:UKN

[Sho12-62] Ben Short. Ubiquitylation keeps the nucleus moving. Journal of Cell Biology, 196(1):??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/1/2.1.

Short:2012:WTS

[Sho12-63] Ben Short. WASP takes the sting out of SCAR mutants. Journal of Cell Biology, 198(4):??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/4/472.1.

Short:2012:WMA

[Sho12-64] Ben Short. Why a microRNA’s absence makes the heart grow larger. Journal of Cell Biology, 199(5):??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
5/714.1.

Short:2013:AST

[Sho13a] Ben Short. AKT shuts down the TOR network. Journal of Cell Biology, 203(4):??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/4/552.2.

Short:2013:ADC

[Sho13b] Ben Short. Arp2/ 3-deficient cells share their problems. Journal of Cell Biology, 203(6):867–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/6/867.

Short:2013:ABH

[Sho13c] Ben Short. Aurora B helps the central spindle measure up. Journal of Cell Biology, 202(4):601–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/4/601.

Short:2013:BED

[Sho13d] Ben Short. The BBSome’s export duty. Journal of Cell Biology, 201(2):??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/2/168.2.

Short:2013:BAF

[Sho13e] Ben Short. Breaking down Arp2/3 function. Journal of Cell Biology, 200(5):??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/5/552.2.

Short:2013:BRS

[Sho13f] Ben Short. Bruchpilot readies synaptic vesicles for release. Journal of Cell Biology, 202(4):??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/4/600.2.

Short:2013:BFD

[Sho13g] Ben Short. Bub3’s function doesn’t stop at the checkpoint. Journal of Cell Biology, 201(3):??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/3/356.2.

Short:2013:CO

[Sho13h] Ben Short. A call for oxygen in the ER. Journal of Cell Biology, 203(4):??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/4/552.3.

Short:2013:CCP

[Sho13i] Ben Short. The CCM complex puts a cap on integrin activity. Journal of Cell Biology, 202(3):399–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/3/399.

Short:2013:CMM

[Sho13j] Ben Short. CDK1 makes the meiotic spindle wait. Journal of Cell Biology, 202(2):??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/2/182.1.

Short:2013:CTK

[Sho13k] Ben Short. Chemokine turnover keeps neurons on track. Journal of Cell Biology, 200(3):??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/3/236.3.

Short:2013:CKR

[Sho13l] Ben Short. CUL4B keeps replication firing on all cylinders. Journal of Cell Biology, 200(6):??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/6/684.3.

Short:2013:DRF

[Sho13m] Ben Short. Determining the replication factory settings. Journal of Cell Biology, 202(7):??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/7/986.1.

Short:2013:EHD

[Sho13n] Ben Short. EB3 helps daughter cells spread their wings. Journal of Cell Biology, 201(5):??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/201/5/646.2.

Short:2013:ECH

[Sho13o] Ben Short. EphA2 cleavage helps tumor cells go it alone. Journal of Cell Biology, 201(3):??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/201/3/356.3.

Short:2013:ECC

[Sho13p] Ben Short. Epiblast cells CLASP onto the basement membrane. Journal of Cell Biology, 202(4):??, August 2013. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/202/4/
600.1.

Short:2013:ERH

[Sho13q] Ben Short. Epithelia restored by healing waves. Journal of Cell Biology, 202(2):??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/2/182.3.

Short:2013:FFM

[Sho13r] Ben Short. Fewer forks make faster progress. Journal of Cell Biology, 201(3):??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/3/356.1.

Short:2013:FCR

[Sho13s] Ben Short. Formin’ the cytokinetic ring. Journal of Cell Biology, 203(1):3–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/1/3.

Short:2013:FKA

[Sho13t] Ben Short. FRETting over kinesin activity. Journal of Cell Biology, 203(3):??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/3/376.2.

Short:2013:FOE

[Sho13u] Ben Short. Frog oocytes evade the checkpoint. Journal of Cell Biology, 201(2):??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/2/168.1.

Short:2013:GHR

[Sho13v] Ben Short. GEFs help Rabs find their target. Journal of Cell Biology, 200(3):??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/3/236.2.

Short:2013:HSG

[Sho13w] Ben Short. Heat shock gene resides in a pore neighborhood. Journal of Cell Biology, 200(5):??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/5/552.1.

Short:2013:HPT

[Sho13x] Ben Short. Hif-1α provides a Twist to neural crest migration. Journal of Cell Biology, 201(5):??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/5/646.3.

Short:2013:HHG

[Sho13y] Ben Short. Histone H1 gets Pin’d onto chromatin. Journal of Cell Biology, 203(1):??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/1/2.1.

Short:2013:HTI

[Sho13z] Ben Short. Hypoxia takes invadopodia up a Notch. Journal of Cell Biology, 201(2):??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/2/168.3.

Short:2013:KHM

[Sho13-27] Ben Short. KASH5 helps meiotic chromosomes LINC up. Journal of Cell Biology, 202(7):??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/7/986.2.

Short:2013:LB

[Sho13-28] Ben Short. Lamellipodin branches out. Journal of Cell Biology, 203(4):553–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/4/553.

Short:2013:LPM

[Sho13-29] Ben Short. Lis1 puts mRNAs in their place. Journal of Cell Biology, 202(3):398–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/3/398.

Short:2013:LPP

[Sho13-30] Ben Short. Low pH puts proteasomes in storage. Journal of Cell Biology, 201(5):??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/5/646.1.

Short:2013:MHM

[Sho13-31] Ben Short. MAD2L2 helps mitotic cells take it slow. Journal of Cell Biology, 203(1):??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/1/2.2.

Short:2013:MVG

[Sho13-32] Ben Short. A mature view of the Golgi. Journal of Cell Biology, 201(7):??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/7/964.2.

Short:2013:MHB

[Sho13-33] Ben Short. MDGAs help the brain lose its inhibitions. Journal of Cell Biology, 200(3):237–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/3/237.

Short:2013:MSN

[Sho13-34] Ben Short. Microtubules support nuclear nonproliferation arrangement. Journal of Cell Biology, 203(4):??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
4/552.1.

Short:2013:MOE

[Sho13-35] Ben Short. Misshapen orders an egg roll. Journal of Cell Biology, 200(6):??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/6/684.2.

Short:2013:MG

[Sho13-36] Ben Short. Moving the Greatwall. Journal of Cell Biology, 202(2):??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/2/182.2.

Short:2013:NVM

[Sho13-37] Ben Short. Notochord vacuoles make a rod for the vertebrate back. Journal of Cell Biology, 200(5):553–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/200/
5/553.

Short:2013:OIR

[Sho13-38] Ben Short. Oncogenes induce a reversal of replication’s fortunes. Journal of Cell Biology, 200(6):??, March 2013. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/200/6/
684.1.

Short:2013:PCM

[Sho13-39] Ben Short. p120 CLASPs microtubules to junctions. Journal of Cell Biology, 203(6):??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/6/866.3.

Short:2013:PTD

[Sho13-40] Ben Short. PEX26 takes the direct route. Journal of Cell Biology, 200(5):??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/5/552.3.

Short:2013:PHA

[Sho13-41] Ben Short. Phosphorylation helps Atg18 get the vacuole in shape. Journal of Cell Biology, 202(4):??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/202/4/
600.3.

Short:2013:PP

[Sho13-42] Ben Short. The plastic proteome. Journal of Cell Biology, 200 (6):685–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/200/6/685.

Short:2013:PHM

[Sho13-43] Ben Short. PLP helps the mother centrosome stay mum. Journal of Cell Biology, 202(7):987–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/7/987.

Short:2013:PVM

[Sho13-44] Ben Short. A potato virus makes its move. Journal of Cell Biology, 201(7):??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/7/964.1.

Short:2013:RGA

[Sho13-45] Ben Short. RNA granules act as egg timers. Journal of Cell Biology, 202(7):??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/7/986.3.

Short:2013:SSPa

[Sho13-46] Ben Short. Sanpodo seals precursors’ fate. Journal of Cell Biology, 201(3):357–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/3/357.

Short:2013:SCC

[Sho13-47] Ben Short. Senescent cells have a case of the SADS. Journal of Cell Biology, 203(6):??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/6/866.1.

Short:2013:SPD

[Sho13-48] Ben Short. Sequestration puts Drp1 on furlough. Journal of Cell Biology, 201(7):??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/7/964.3.

Short:2013:SIT

[Sho13-49] Ben Short. Setting an immuno–TRAP for PML nuclear bodies. Journal of Cell Biology, 201(2):169–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/2/169.

Short:2013:SGN

[Sho13-50] Ben Short. Sortilins get a new delivery route. Journal of Cell Biology, 203(3):377–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/3/377.

Short:2013:SCR

[Sho13-51] Ben Short. Spastin coordinates the recycling program. Journal of Cell Biology, 202(3):??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/202/3/398.3.

Short:2013:SCG

[Sho13-52] Ben Short. Stem cells get a cholesterol test. Journal of Cell Biology, 201(5):647–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/5/647.

Short:2013:SSPb

[Sho13-53] Ben Short. Switching on synaptic PP1. Journal of Cell Biology, 203(3):??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/3/376.3.

Short:2013:SPB

[Sho13-54] Ben Short. Syntaphilin puts the brakes on axonal mitochondria. Journal of Cell Biology, 202(2):183–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
2/183.

Short:2013:CJI

[Sho13-55] Ben Short. T cells JAK up integrin activity. Journal of Cell Biology, 203(6):??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/6/866.2.

Short:2013:TLC

[Sho13-56] Ben Short. Testing the limits of cell migration. Journal of Cell Biology, 201(7):965–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/7/965.

Short:2013:UIB

[Sho13-57] Ben Short. Ubiquitin isolates bacterial invaders. Journal of Cell Biology, 203(1):??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/203/1/2.3.

Short:2013:WTK

[Sho13-58] Ben Short. When two kinases are better than one. Journal of Cell Biology, 202(3):??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/3/398.1.

Short:2013:WTE

[Sho13-59] Ben Short. Working at the tail end of tubulin. Journal of Cell Biology, 200(3):??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/3/236.1.

Short:2013:YCC

[Sho13-60] Ben Short. Yeast centromeres coordinate their movements. Journal of Cell Biology, 203(3):??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/3/376.1.

Short:2014:ACC

[Sho14a] Ben Short. Acinus cleavage cuts down on autophagy. Journal of Cell Biology, 207(2):??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/2/162.3.

Short:2014:AAA

[Sho14b] Ben Short. Ankyrin aweigh for axonal transport. Journal of Cell Biology, 207(6):??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/6/678.3.

Short:2014:BDV

[Sho14c] Ben Short. Bem1p directs vesicular traffic. Journal of Cell Biology, 207(1):??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/1/2.2.

Short:2014:BMK

[Sho14d] Ben Short. BUB-1 makes kinetochores MAD. Journal of Cell Biology, 204(5):??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/5/626.1.

Short:2014:BYD

[Sho14e] Ben Short. Budding yeast do the Cdc42 two-step. Journal of Cell Biology, 206(1):??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/1/2.1.

Short:2014:BMM

[Sho14f] Ben Short. Building modular motor complexes. Journal of Cell Biology, 207(3):??, November 2014. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/3/318.3.

Short:2014:CBC

[Sho14g] Ben Short. CAL1 builds centromeres on the fly. Journal of Cell Biology, 204(3):??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/3/282.1.

Short:2014:COW

[Sho14h] Ben Short. CDK5 opens the way for DLC1. Journal of Cell Biology, 207(5):??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/5/572.2.

Short:2014:CFD

[Sho14i] Ben Short. Chibby’s function in the dock. Journal of Cell Biology, 207(1):??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/1/2.3.

Short:2014:CG

[Sho14j] Ben Short. Circumventing the Greatwall. Journal of Cell Biology, 204(6):??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/6/858.1.

Short:2014:CMB

[Sho14k] Ben Short. Clathrin’s muscle-building regimen. Journal of Cell Biology, 205(3):285–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/3/285.

Short:2014:CCA

[Sho14l] Ben Short. CNTD1’s crossover act. Journal of Cell Biology, 205(5):609–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/5/609.

Short:2014:DKR

[Sho14m] Ben Short. Defining the kinetochore’s rules of engagement. Journal of Cell Biology, 206(1):3–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/1/3.

Short:2014:DHR

[Sho14n] Ben Short. Deubiquitination helps Rad18 grow more tolerant. Journal of Cell Biology, 206(2):??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/2/146.3.

Short:2014:DLB

[Sho14o] Ben Short. Dynamin loss bulks up endocytosis. Journal of Cell Biology, 204(7):??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/7/1078.2.

Short:2014:ENS

[Sho14p] Ben Short. Endocytosis in its natural state. Journal of Cell Biology, 205(5):??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/5/608.3.

Short:2014:ELN

[Sho14q] Ben Short. Exploring the LINC to nuclear envelope spacing. Journal of Cell Biology, 206(2):??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/2/146.1.

Short:2014:FR

[Sho14r] Ben Short. The fellowship of the Rng. Journal of Cell Biology, 205(3):??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/3/284.2.

Short:2014:FRT

[Sho14s] Ben Short. First remodel, then recycle. Journal of Cell Biology, 205(1):3–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/1/3.

Short:2014:FDC

[Sho14t] Ben Short. The fluid dynamics of collective cell migration. Journal of Cell Biology, 206(1):??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/1/2.2.

Short:2014:GPL

[Sho14u] Ben Short. GAAC pathway limits autophagy to a light snack. Journal of Cell Biology, 206(2):??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/2/146.2.

Short:2014:HNK

[Sho14v] Ben Short. Harvesting a new KASH crop. Journal of Cell Biology, 205(5):??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/5/608.2.

Short:2014:HMK

[Sho14w] Ben Short. How mitosis keeps itself in order. Journal of Cell Biology, 207(1):3–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/1/3.

Short:2014:IIA

[Sho14x] Ben Short. If it ain’t broke, don’t fix it. Journal of Cell Biology, 206(5):??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/5/574.1.

Short:2014:KDR

[Sho14y] Ben Short. Keeping DNA replication below PAR. Journal of Cell Biology, 205(4):??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/4/430.1.

Short:2014:KEE

[Sho14z] Ben Short. KIF13B erects an endocytic scaffold. Journal of Cell Biology, 204(3):283–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/3/283.

Short:2014:KKO

[Sho14-27] Ben Short. KIF4A knows its own limits. Journal of Cell Biology, 207(6):??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/6/678.1.

Short:2014:LPS

[Sho14-28] Ben Short. Lamin-A provides stiff resistance to cell migration. Journal of Cell Biology, 204(5):??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/5/626.2.

Short:2014:LBG

[Sho14-29] Ben Short. Lasp brings a giant down to size. Journal of Cell Biology, 206(4):??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/4/452.3.

Short:2014:LEU

[Sho14-30] Ben Short. Late endosomes uproot focal adhesions. Journal of Cell Biology, 205(4):??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/4/430.2.

Short:2014:LHE

[Sho14-31] Ben Short. Lipids help epithelia stand tall. Journal of Cell Biology, 206(2):147–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/2/147.

Short:2014:MSN

[Sho14-32] Ben Short. Macrophages show neutrophils the exit. Journal of Cell Biology, 207(5):??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/5/572.1.

Short:2014:MHG

[Sho14-33] Ben Short. MagT1 helps a glycosylase gain acceptance. Journal of Cell Biology, 206(4):??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/206/4/452.1.

Short:2014:MPP

[Sho14-34] Ben Short. Mdm2 pulls the plug on glycolysis. Journal of Cell Biology, 204(5):627–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/5/627.

Short:2014:MCK

[Sho14-35] Ben Short. Mitochondrial caspase keeps autophagy in flux. Journal of Cell Biology, 205(4):431–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/205/4/431.

Short:2014:MPH

[Sho14-36] Ben Short. Motor proteins Hook on to early endosomes. Journal of Cell Biology, 204(6):859–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/6/859.

Short:2014:MHB

[Sho14-37] Ben Short. mTORC2 helps brown adipose tissue fuel up. Journal of Cell Biology, 207(3):??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/207/3/318.2.

Short:2014:MIP

[Sho14-38] Ben Short. A mutual interest in prostaglandin signaling. Journal of Cell Biology, 204(3):??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/3/282.2.

Short:2014:NPE

[Sho14-39] Ben Short. Netrin puts an end to the anchor cell’s vacillations. Journal of Cell Biology, 206(5):575–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
5/575.

Short:2014:NCM

[Sho14-40] Ben Short. A new CLUH to mitochondrial biogenesis. Journal of Cell Biology, 207(2):??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/2/162.1.

Short:2014:OSM

[Sho14-41] Ben Short. OPA1’s shortcut to mitochondrial fission. Journal of Cell Biology, 204(6):??, March 2014. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/6/858.2.

Short:2014:OGJ

[Sho14-42] Ben Short. Osmotic gradient is just the tonic for wounded epithelia. Journal of Cell Biology, 207(6):679–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
6/679.

Short:2014:PCI

[Sho14-43] Ben Short. p53 cuts off invading cells. Journal of Cell Biology, 204(7):1079–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/7/1079.

Short:2014:PHD

[Sho14-44] Ben Short. p73 helps developing sperm stick to the right path. Journal of Cell Biology, 204(7):??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/7/1078.3.

Short:2014:PDV

[Sho14-45] Ben Short. PIP_{2} directs vinculin assembly. Journal of Cell Biology, 207(5):??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/5/572.3.

Short:2014:PIC

[Sho14-46] Ben Short. Prion infections come with strings attached. Journal of Cell Biology, 204(3):??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/3/282.3.

Short:2014:PPP

[Sho14-47] Ben Short. Probing PtdIns4 P ’s localization. Journal of Cell Biology, 205(1):??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/1/2.3.

Short:2014:PCK

[Sho14-48] Ben Short. Protein coat keeps axons buttoned up. Journal of Cell Biology, 205(1):??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/1/2.2.

Short:2014:RGT

[Sho14-49] Ben Short. Red and green traffic signals. Journal of Cell Biology, 207(3):319–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/3/319.

Short:2014:ROG

[Sho14-50] Ben Short. A replication origin gets SNP’d out in fragile X syndrome. Journal of Cell Biology, 206(5):??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
5/574.2.

Short:2014:RHG

[Sho14-51] Ben Short. A ribonuclease helps the ER get in shape. Journal of Cell Biology, 207(1):??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/1/2.1.

Short:2014:SRG

[Sho14-52] Ben Short. Satellite RNA guides kinetochore assembly. Journal of Cell Biology, 207(3):??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/207/3/318.1.

Short:2014:SPL

[Sho14-53] Ben Short. Septins provide a link to epithelial migration. Journal of Cell Biology, 207(2):??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/207/2/162.2.

Short:2014:SAO

[Sho14-54] Ben Short. A short Anillin opens the way for germline development. Journal of Cell Biology, 206(1):??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/1/2.3.

Short:2014:SSS

[Sho14-55] Ben Short. Sperm’s sensitive steering machinery. Journal of Cell Biology, 206(4):??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/4/452.2.

Short:2014:SMD

[Sho14-56] Ben Short. Sphingomyelin is the master of its domain. Journal of Cell Biology, 206(5):??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/206/5/574.3.

Short:2014:SCS

[Sho14-57] Ben Short. The spindle checkpoint’s on-off switch. Journal of Cell Biology, 205(4):??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/4/430.3.

Short:2014:SMS

[Sho14-58] Ben Short. Spindle microtubules sustain the tension. Journal of Cell Biology, 205(3):??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/3/284.1.

Short:2014:SPM

[Sho14-59] Ben Short. SUN proteins melt the nuclear envelope. Journal of Cell Biology, 204(7):??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/7/1078.1.

Short:2014:STC

[Sho14-60] Ben Short. A surprising Twist to cell dissemination. Journal of Cell Biology, 204(5):??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/5/626.3.

Short:2014:TIS

[Sho14-61] Ben Short. Talin’s invasive side. Journal of Cell Biology, 205(5):??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/5/608.1.

Short:2014:TSO

[Sho14-62] Ben Short. Tat substrates open the portal. Journal of Cell Biology, 205(1):??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/1/2.1.

Short:2014:TXP

[Sho14-63] Ben Short. Tenascin-x and the pick of destiny. Journal of Cell Biology, 205(3):??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/3/284.3.

Short:2014:TGP

[Sho14-64] Ben Short. TGF-β gets primed for action. Journal of Cell Biology, 207(2):163–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/2/163.

Short:2014:TTA

[Sho14-65] Ben Short. Tracking traction in amoeboid cells. Journal of Cell Biology, 204(6):??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/6/858.3.

Short:2014:UDP

[Sho14-66] Ben Short. Understanding the DisAp-pearance of cilia orientation. Journal of Cell Biology, 207(6):??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/207/6/
678.2.

Short:2014:USC

[Sho14-67] Ben Short. Unlocking the secrets of chitinase secretion. Journal of Cell Biology, 207(5):573–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/207/5/573.

Short:2014:VGP

[Sho14-68] Ben Short. Vps74 gives phosphatase directions. Journal of Cell Biology, 206(4):453–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/206/4/453.

Shen:2012:SIW

[SHS^{+}12] Qing-Tao Shen, Peter P. Hsiue, Charles V. Sindelar, Matthew D. Welch, Kenneth G. Campellone, and Hong-Wei Wang. Structural insights into WHAMM-mediated cytoskeletal coordination during membrane remodeling. Journal of Cell Biology, 199(1):111–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/1/111.

Soares:2013:RVF

[SHS^{+}13] Helena Soares, Ricardo Henriques, Martin Sachse, Leandro Ventimiglia, Miguel A. Alonso, Christophe Zimmer, Maria-Isabel Thoulouze, and Andres Alcover. Regulated vesicle fusion generates signaling nanoterritories that control T-cell activation at the immunological synapse. Journal of Cell Biology, 203(1):??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/1/2031OIA112.

Stephens:2011:CCI

[SHV^{+}11] Andrew D. Stephens, Julian Haase, Leandra Vicci, Russell M. Taylor, and Kerry Bloom. Cohesin, condensin, and the intramolecular centromere loop together generate the mitotic chromatin spring. Journal of Cell Biology, 193(7):1167–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
193/7/1167.

Stephens:2013:PCL

[SHV^{+}13] Andrew D. Stephens, Rachel A. Haggerty, Paula A. Vasquez, Leandra Vicci, Chloe E. Snider, Fu Shi, Cory Quammen, Christopher Mullins, Julian Haase, Russell M. Taylor, Jolien S. Verdaasdonk, Michael R. Falvo, Yuan Jin, M. Gregory Forest, and Kerry Bloom. Pericentric chromatin loops function as a nonlinear spring in mitotic force balance. Journal of Cell Biology, 200(6):757–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/6/757.

Suzuki:2010:CBS

[SIO10] Tatsuya Suzuki, Hiroto Izumi, and Mutsuhito Ohno. Cajal body surveillance of U snRNA export complex assembly. Journal of Cell Biology, 190(4):603–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/4/603.

Sixt:2012:CMF

[Six12] Michael Sixt. Cell migration: Fibroblasts find a new way to get ahead. Journal of Cell Biology, 197(3):347–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
3/347.

Stankovic:2013:RVK

[SJ13] Ana Stankovic and Lars E. T. Jansen. Reductionism at the vertebrate kinetochore. Journal of Cell Biology, 200(1):7–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
200/1/7.

Sirois:2013:RNA

[SJM^{+}13] Cherilyn M. Sirois, Tengchuan Jin, Allison L. Miller, Damien Bertheloot, Hirotaka Nakamura, Gabor L. Horvath, Abubakar Mian, Jiansheng Jiang, Jacob Schrum, Lukas Bossaller, Karin Pelka, Natalio Garbi, Yambasu Brewah, Jane Tian, Chew-Shun Chang, Partha Chowdhury, Gary Sims, Roland Kolbeck, Anthony Coyle, Alison Humbles, T. Sam Xiao, and Eicke Latz. RAGE is a nucleic acid receptor that promotes inflammatory responses to DNA. Journal of Cell Biology, 203(1):??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
1/2031OIA111.

Smith:2014:ASD

[SJRV14] Alex J. Smith, Byung-Ju Jin, Julien Ratelade, and Alan S. Verkman. Aggregation state determines the localization and function of M1- and M23–aquaporin-4 in astrocytes. Journal of Cell Biology, 204(4):559–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/4/559.

Sohaskey:2010:ORO

[SJZ^{+}10] Michael L. Sohaskey, Yebin Jiang, Jenny J. Zhao, Andreas Mohr, Frank Roemer, and Richard M. Harland. Osteopotentia regulates osteoblast maturation, bone formation, and skeletal integrity in mice. Journal of Cell Biology, 189(3):511–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
3/511.

Stelter:2011:PNF

[SKFH11] Philipp Stelter, Ruth Kunze, Jessica Fischer, and Ed Hurt. Probing the nucleoporin FG repeat network defines structural and functional features of the nuclear pore complex. Journal of Cell Biology, 195(2):183–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/2/183.

Schmidt:2010:ABK

[SKH^{+}10] Jens C. Schmidt, Tomomi Kiyomitsu, Tetsuya Hori, Chelsea B. Backer, Tatsuo Fukagawa, and Iain M. Cheeseman. Aurora B kinase controls the targeting of the Astrin–SKAP complex to bioriented kinetochores. Journal of Cell Biology, 191(2): 269–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/2/269.

Sarkar:2010:ICR

[SKM10a] Sovan Sarkar, Rhian Kiely, and Peter J. McHugh. The Ino80 chromatin-remodeling complex restores chromatin structure during UV DNA damage repair. Journal of Cell Biology, 191(6):1061–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/6/1061.

Schwander:2010:CBH

[SKM10b] Martin Schwander, Bechara Kachar, and Ulrich Müller. The cell biology of hearing. Journal of Cell Biology, 190(1):9–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
190/1/9.

Schmidt:2012:CMV

[SKN^{+}12] Kerstin N. Schmidt, Stefanie Kuhns, Annett Neuner, Birgit Hub, Hanswalter Zentgraf, and Gislene Pereira. Cep164 mediates vesicular docking to the mother centriole during early steps of ciliogenesis. Journal of Cell Biology, 199(7):1083– ??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/7/1083.

Saner:2013:SAN

[SKN^{+}13] Nazan Saner, Jens Karschau, Toyoaki Natsume, Marek Gierliński, Renata Retkute, Michelle Hawkins, Conrad A. Nieduszynski, J. Julian Blow, Alessandro P. S. de Moura, and Tomoyuki U. Tanaka. Stochastic association of neighboring replicons creates replication factories in budding yeast. Journal of Cell Biology, 202(7):1001–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/7/1001.

Siddiqui:2011:CES

[SKV^{+}11] M. Rizwan Siddiqui, Yulia A. Komarova, Stephen M. Vogel, Xiaopei Gao, Marcelo G. Bonini, Johnson Rajasingh, You-Yang Zhao, Viktor Brovkovych, and Asrar B. Malik. Caveolin-1-eNOS signaling promotes p190RhoGAP–A nitration and endothelial permeability. Journal of Cell Biology, 193(5):841–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
193/5/841.

Saraya:2011:PRH

[SKVvdK11] Ruchi Saraya, Arjen M. Krikken, Marten Veenhuis, and Ida J. van der Klei. Peroxisome reintroduction in Hansenula polymorpha requires Pex25 and Rho1. Journal of Cell Biology, 193(5):885–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/5/885.

Slaidina:2014:TCG

[SL14] Maija Slaidina and Ruth Lehmann. Translational control in germline stem cell development. Journal of Cell Biology, 207 (1):13–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/1/13.

Srikumar:2013:GAS

[SLC^{+}13] Tharan Srikumar, Megan C. Lewicki, Michael Costanzo, Johnny M. Tkach, Harm van Bakel, Kyle Tsui, Erica S. Johnson, Grant W. Brown, Brenda J. Andrews, Charles Boone, Guri Giaever, Corey Nislow, and Brian Raught. Global analysis of SUMO chain function reveals multiple roles in chromatin regulation. Journal of Cell Biology, 201(1):145–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
1/145.

Shigeoka:2013:RBM

[SLH13] Toshiaki Shigeoka, Bo Lu, and Christine E. Holt. RNA-based mechanisms underlying axon guidance. Journal of Cell Biology, 202(7):991–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/7/991.

Sun:2014:SOM

[SLH^{+}14] Jianwei Sun, Fujian Lu, Huifang He, Junling Shen, Jane Messina, Rahel Mathew, Dapeng Wang, Amod A. Sarnaik, Wei-Chiao Chang, Minjung Kim, Heping Cheng, and Shengyu Yang. STIM1- and Orai1-mediated Ca^{2+} oscillation orchestrates invadopodium formation and melanoma invasion. Journal of Cell Biology, 207(4):535–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/207/4/535.

Snider:2013:GSR

[SLK^{+}13] Natasha T. Snider, Jessica M. Leonard, Raymond Kwan, Nicholas W. Griggs, Liangyou Rui, and M. Bishr Omary. Glucose and SIRT2 reciprocally mediate the regulation of keratin 8 by lysine acetylation. Journal of Cell Biology, 200(3): 241–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/200/3/241.

Schulz:2011:TPR

[SLM^{+}11] Christian Schulz, Oleksandr Lytovchenko, Jonathan Melin, Agnieszka Chacinska, Bernard Guiard, Piotr Neumann, Ralf Ficner, Olaf Jahn, Bernhard Schmidt, and Peter Rehling. Tim50’s presequence receptor domain is essential for signal driven transport across the TIM23 complex. Journal of Cell Biology, 195(4):643–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/195/4/643.

Shao:2013:XOM

[SLM^{+}13] Hua Shao, Ruizhen Li, Chunqi Ma, Eric Chen, and X. Johné Liu. Xenopus oocyte meiosis lacks spindle assembly checkpoint control. Journal of Cell Biology, 201(2):191–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
2/191.

Szulwach:2010:CTB

[SLS^{+}10] Keith E. Szulwach, Xuekun Li, Richard D. Smrt, Yujing Li, Yuping Luo, Li Lin, Nicholas J. Santistevan, Wendi Li, Xinyu Zhao, and Peng Jin. Cross talk between microRNA and epigenetic regulation in adult neurogenesis. Journal of Cell Biology, 189(1):127–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/1/127.

Strobl-Mazzulla:2012:PSR

[SMB12] Pablo H. Strobl-Mazzulla and Marianne E. Bronner. A PHD12–Snail2 repressive complex epigenetically mediates neural crest epithelial-to-mesenchymal transition. Journal of Cell Biology, 198(6):999–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/6/999.

Saadaoui:2014:DCP

[SMdP^{+}14] Mehdi Saadaoui, Mickaël Machicoane, Florencia di Pietro, Fred Etoc, Arnaud Echard, and Xavier Morin. Dlg1 controls planar spindle orientation in the neuroepithelium through direct interaction with LGN. Journal of Cell Biology, 206(6): 707–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/6/707.

Shahbazi:2013:CIP

[SME^{+}13] Marta N. Shahbazi, Diego Megias, Carolina Epifano, Anna Akhmanova, Gregg G. Gundersen, Elaine Fuchs, and Mirna Perez-Moreno. CLASP2 interacts with p120-catenin and governs microtubule dynamics at adherens junctions. Journal of Cell Biology, 203(6):1043–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/6/1043.

Simons:2014:UCB

[SMK14] Mikael Simons, Thomas Misgeld, and Martin Kerschensteiner. A unified cell biological perspective on axon–myelin injury.

Journal of Cell Biology, 206(3):335–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/3/335.

Sloan:2013:BEE

[SML^{+}13] Katherine E. Sloan, Sandy Mattijssen, Simon Lebaron, David Tollervey, Ger J. M. Pruijn, and Nicholas J. Watkins. Both endonucleolytic and exonucleolytic cleavage mediate ITS1 removal during human ribosomal RNA processing. Journal of Cell Biology, 200(5):577–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/5/577.

Stramer:2010:CMM

[SMM^{+}10] Brian Stramer, Severina Moreira, Tom Millard, Iwan Evans, Chieh-Yin Huang, Ola Sabet, Martin Milner, Graham Dunn, Paul Martin, and Will Wood. Clasp-mediated microtubule bundling regulates persistent motility and contact repulsion in Drosophila macrophages in vivo. Journal of Cell Biology, 189(4):681–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/4/681.

Specht:2011:HRS

[SMMB11] Sebastian Specht, Stephanie B. M. Miller, Axel Mogk, and Bernd Bukau. Hsp42 is required for sequestration of protein aggregates into deposition sites in Saccharomyces cerevisiae. Journal of Cell Biology, 195(4):617–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
4/617.

Solinet:2013:ABE

[SMS^{+}13] Sara Solinet, Kazi Mahmud, Shannon F. Stewman, Khaled Ben El Kadhi, Barbara Decelle, Lama Talje, Ao Ma, Benjamin H. Kwok, and Sébastien Carreno. The actin-binding ERM protein Moesin binds to and stabilizes microtubules at the cell cortex. Journal of Cell Biology, 202(2):251–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/2/251.

Sikirzhytski:2014:DKS

[SMS^{+}14] Vitali Sikirzhytski, Valentin Magidson, Jonathan B. Steinman, Jie He, Maël Le Berre, Irina Tikhonenko, Jeffrey G. Ault, Bruce F. McEwen, James K. Chen, Haixin Sui, Matthieu Piel, Tarun M. Kapoor, and Alexey Khodjakov. Direct kinetochore– spindle pole connections are not required for chromosome segregation. Journal of Cell Biology, 206(2):231–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
2/231.

Schulte-Merker:2011:LVM

[SMSP11] Stefan Schulte-Merker, Amélie Sabine, and Tatiana V. Petrova. Lymphatic vascular morphogenesis in development, physiology, and disease. Journal of Cell Biology, 193(4): 607–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/4/607.

Sugawara:2010:BHD

[SMT^{+}10] Yo Sugawara, Takuhiro Matsumura, Yuki Takegahara, Yingji Jin, Yoshikazu Tsukasaki, Masatoshi Takeichi, and Yukako Fujinaga. Botulinum hemagglutinin disrupts the intercellular epithelial barrier by directly binding E-cadherin. Journal of Cell Biology, 189(4):691–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/4/691.

Swanson:2013:HOU

[SMZL13] Eric C. Swanson, Benjamin Manning, Hong Zhang, and Jeanne B. Lawrence. Higher-order unfolding of satellite heterochromatin is a consistent and early event in cell senescence. Journal of Cell Biology, 203(6):929–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
6/929.

Schmidt:2011:KMS

[SNR^{+}11] Sarah Schmidt, Inaam Nakchbandi, Raphael Ruppert, Nina Kawelke, Michael W. Hess, Kristian Pfaller, Pierre Jurdic, Reinhard Fässler, and Markus Moser. Kindlin-3-mediated signaling from multiple integrin classes is required for osteoclastmediated bone resorption. Journal of Cell Biology, 192(5):

883–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/5/883.

Schwab:2013:FCR

[SNSyN13] Rebekka A. Schwab, Jadwiga Nieminuszczy, Kazuo Shin-ya, and Wojciech Niedzwiedz. FANCJ couples replication past natural fork barriers with maintenance of chromatin structure. Journal of Cell Biology, 201(1):33–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/1/33.

Shevchuk:2012:AMC

[SNT^{+}12] Andrew I. Shevchuk, Pavel Novak, Marcus Taylor, Ivan A. Diakonov, Azza Ziyadeh-Isleem, Marc Bitoun, Pascale Guicheney, Max J. Lab, Julia Gorelik, Christien J. Merrifield, David Klenerman, and Yuri E. Korchev. An alternative mechanism of clathrin-coated pit closure revealed by ion conductance microscopy. Journal of Cell Biology, 197(4):499–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
4/499.

Sen:2012:DPS

[SNZVK12] Arnab Sen, Zsanett Nagy-Zsvér-Vadas, and Michael P. Krahn. Drosophila PATJ supports adherens junction stability by modulating Myosin light chain activity. Journal of Cell Biology, 199(4):685–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/4/685.

Sen:2013:DPS

[SNZVK13] Arnab Sen, Zsanett Nagy-Zsvér-Vadas, and Michael P. Krahn. Drosophila PATJ supports adherens junction stability by modulating Myosin light chain activity. Journal of Cell Biology, 200(6):853–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/6/853.

Schatz:2011:TCD

[SOW^{+}11] Jonathan H. Schatz, Elisa Oricchio, Andrew L. Wolfe, Man Jiang, Irina Linkov, Jocelyn Maragulia, Weiji Shi, Zhigang Zhang, Rajasekhar K. Vinagolu, Nen C. Pagano, John A.

Porco, Julie Teruya-Feldstein, Neal Rosen, Andrew D. Zelenetz, Jerry Pelletier, and Hans-Guido Wendel. Targeting cap-dependent translation blocks converging survival signals by AKT and PIM kinases in lymphoma. Journal of Cell Biology, 194(5):??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/5/i9.

Stringer:2011:SUS

[SP11] Daniel K. Stringer and Robert C. Piper. A single ubiquitin is sufficient for cargo protein entry into MVBs in the absence of ESCRT ubiquitination. Journal of Cell Biology, 192(2): 229–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/2/229.

Saraiva:2013:HPC

[SPC^{+}13] Nuno Saraiva, David L. Prole, Guia Carrara, Benjamin F. Johnson, Colin W. Taylor, Maddy Parsons, and Geoffrey L. Smith. hGAAP promotes cell adhesion and migration via the stimulation of store-operated Ca^{2+} entry and calpain 2. Journal of Cell Biology, 202(4):699–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/4/699.

Sir:2013:LCC

[SPD^{+}13] Joo-Hee Sir, Monika Pütz, Owen Daly, Ciaran G. Morrison, Mark Dunning, John V. Kilmartin, and Fanni Gergely. Loss of centrioles causes chromosomal instability in vertebrate somatic cells. Journal of Cell Biology, 203(5):747–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
5/747.

Shteyn:2011:ARA

[SPF11] Elina Shteyn, Lucy Pigati, and Heike Fölsch. Arf6 regulates AP-1B-dependent sorting in polarized epithelial cells. Journal of Cell Biology, 194(6):873–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/6/873.

Shamir:2014:TID

[SPJ^{+}14] Eliah R. Shamir, Elisa Pappalardo, Danielle M. Jorgens, Kester Coutinho, Wen-Ting Tsai, Khaled Aziz, Manfred Auer, Phuoc T. Tran, Joel S. Bader, and Andrew J. Ewald. Twist1induced dissemination preserves epithelial identity and requires E-cadherin. Journal of Cell Biology, 204(5):839–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
204/5/839.

Sonneville:2012:DRL

[SQC^{+}12] Remi Sonneville, Matthieu Querenet, Ashley Craig, Anton Gartner, and J. Julian Blow. The dynamics of replication licensing in live Caenorhabditis elegans embryos. Journal of Cell Biology, 196(2):233–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/2/233.

Schoborg:2013:CIB

[SRBL13] Todd Schoborg, Ryan Rickels, Josh Barrios, and Mariano Labrador. Chromatin insulator bodies are nuclear structures that form in response to osmotic stress and cell death. Journal of Cell Biology, 202(2):261–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/202/2/261.

Shen:2014:PPS

[SRCP14] Yi Shen, Morgane Rosendale, Robert E. Campbell, and David Perrais. pHuji, a pH-sensitive red fluorescent protein for imaging of exo- and endocytosis. Journal of Cell Biology, 207(3): 419–??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/3/419.

Shen:2010:SBS

[SRKR10] Jingshi Shen, Shailendra S. Rathore, Lavan Khandan, and James E. Rothman. SNARE bundle and syntaxin N-peptide constitute a minimal complement for Munc18-1 activation of membrane fusion. Journal of Cell Biology, 190(1):55–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
1/55.

Smith:2013:ISB

[SRP^{+}13] Sarah E. Smith, Boris Rubinstein, Inês Mendes Pinto, Brian D. Slaughter, Jay R. Unruh, and Rong Li. Independence of symmetry breaking on Bem1-mediated autocatalytic activation of Cdc42. Journal of Cell Biology, 202(7):1091–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
7/1091.

Sandu:2010:DIA

[SRS10] Cristinel Sandu, Hyung Don Ryoo, and Hermann Steller. Drosophila IAP antagonists form multimeric complexes to promote cell death. Journal of Cell Biology, 190(6):1039– ??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/6/1039.

Suraneni:2012:ACR

[SRU^{+}12] Praveen Suraneni, Boris Rubinstein, Jay R. Unruh, Michael Durnin, Dorit Hanein, and Rong Li. The Arp2/3 complex is required for lamellipodia extension and directional fibroblast cell migration. Journal of Cell Biology, 197(2):239–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/197/
2/239.

Straub:2011:CHD

[SRZ^{+}11] Beate K. Straub, Steffen Rickelt, Ralf Zimbelmann, Christine Grund, Caecilia Kuhn, Marcus Iken, Michael Ott, Peter Schirmacher, and Werner W. Franke. E–n-cadherin heterodimers define novel adherens junctions connecting endoderm-derived cells. Journal of Cell Biology, 195(5):873–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
5/873.

Sicheri:2011:PBU

[SS11] Frank Sicheri and Robert H. Silverman. Putting the brakes on the unfolded protein response. Journal of Cell Biology, 193(1):17–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/1/17.

Scime:2010:OSM

[SSB^{+}10] Anthony Scimè, Vahab D. Soleimani, C. Florian Bentzinger, Mark A. Gillespie, Fabien Le Grand, Guillaume Grenier, Lisa Bevilacqua, Mary-Ellen Harper, and Michael A. Rudnicki. Oxidative status of muscle is determined by p107 regulation of PGC-1α. Journal of Cell Biology, 190(4):651–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
4/651.

Schaupp:2014:CEC

[SSD^{+}14] Andreas Schaupp, Ola Sabet, Irina Dudanova, Marion Ponserre, Philippe Bastiaens, and Rüdiger Klein. The composition of EphB2 clusters determines the strength in the cellular repulsion response. Journal of Cell Biology, 204(3):409–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
204/3/409.

Schiefermeier:2014:LEP

[SSdA^{+}14] Natalia Schiefermeier, Julia M. Scheffler, Mariana E. G. de Araujo, Taras Stasyk, Teodor Yordanov, Hannes L. Ebner, Martin Offterdinger, Sebastian Munck, Michael W. Hess, Sara A. Wickström, Anika Lange, Winfried Wunderlich, Reinhard Fässler, David Teis, and Lukas A. Huber. The late endosomal p14–MP1 (LAMTOR2/3) complex regulates focal adhesion dynamics during cell migration. Journal of Cell Biology, 205(4):525–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/4/525.

Stephens:2013:IPD

[SSH^{+}13] Andrew D. Stephens, Chloe E. Snider, Julian Haase, Rachel A. Haggerty, Paula A. Vasquez, M. Gregory Forest, and Kerry Bloom. Individual pericentromeres display coordinated motion and stretching in the yeast spindle. Journal of Cell Biology, 203(3):407–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/3/407.

Spangler:2013:LPP

[SSK^{+}13] Samantha A. Spangler, Sabine K. Schmitz, Josta T. Kevenaar, Esther de Graaff, Heidi de Wit, Jeroen Demmers, Ruud F.

Toonen, and Casper C. Hoogenraad. Liprin-α2 promotes the presynaptic recruitment and turnover of RIM1/CASK to facilitate synaptic transmission. Journal of Cell Biology, 201 (6):915–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/6/915.

Snider:2014:DTG

[SSK^{+}14] Chloe E. Snider, Andrew D. Stephens, Jacob G. Kirkland, Omar Hamdani, Rohinton T. Kamakaka, and Kerry Bloom. Dyskerin, tRNA genes, and condensin tether pericentric chromatin to the spindle axis in mitosis. Journal of Cell Biology, 207(2):189–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/2/189.

Schneider:2014:PMN

[SSL^{+}14] Katharina Schneider, Eric Seemann, Lutz Liebmann, Rashmi Ahuja, Dennis Koch, Martin Westermann, Christian A. Hübner, Michael M. Kessels, and Britta Qualmann. ProSAP1 and membrane nanodomain-associated syndapin I promote postsynapse formation and function. Journal of Cell Biology, 205(2):197–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/2/197.

Steinberg:2012:MDM

[SST^{+}12] Gero Steinberg, Martin Schuster, Ulrike Theisen, Sreedhar Kilaru, Andrew Forge, and Magdalena Martin-Urdiroz. Motordriven motility of fungal nuclear pores organizes chromosomes and fosters nucleocytoplasmic transport. Journal of Cell Biology, 198(3):343–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/3/343.

Samejima:2012:MCC

[SSV^{+}12] Kumiko Samejima, Itaru Samejima, Paola Vagnarelli, Hiromi Ogawa, Giulia Vargiu, David A. Kelly, Flavia de Lima Alves, Alastair Kerr, Lydia C. Green, Damien F. Hudson, Shinya Ohta, Carol A. Cooke, Christine J. Farr, Juri Rappsilber, and William C. Earnshaw. Mitotic chromosomes are compacted laterally by KIF4 and condensin and axially by topoisomerase IIα. Journal of Cell Biology, 199(5):755–??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
5/755.

Safferling:2013:WHR

[SSW^{+}13] Kai Safferling, Thomas Sütterlin, Kathi Westphal, Claudia Ernst, Kai Breuhahn, Merlin James, Dirk Jäger, Niels Halama, and Niels Grabe. Wound healing revised: a novel reepithelialization mechanism revealed by in vitro and in silico models. Journal of Cell Biology, 203(4):691–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
4/691.

Shen:2014:SAM

[SSZ^{+}14] Qing-Tao Shen, Amber L. Schuh, Yuqing Zheng, Kyle Quinney, Lei Wang, Michael Hanna, Julie C. Mitchell, Marisa S. Otegui, Paul Ahlquist, Qiang Cui, and Anjon Audhya. Structural analysis and modeling reveals new mechanisms governing ESCRT–III spiral filament assembly. Journal of Cell Biology, 206(6):763–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/6/763.

Santaguida:2010:DRM

[STD^{+}10] Stefano Santaguida, Anthony Tighe, Anna Morena D’Alise, Stephen S. Taylor, and Andrea Musacchio. Dissecting the role of MPS1 in chromosome biorientation and the spindle checkpoint through the small molecule inhibitor reversine. Journal of Cell Biology, 190(1):73–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/190/1/73.

Steinman:2012:DNP

[Ste12] Lawrence Steinman. The discovery of natalizumab, a potent therapeutic for multiple sclerosis. Journal of Cell Biology, 199 (3):413–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/3/413.

Shrimal:2013:ECT

[STG13] Shiteshu Shrimal, Steven F. Trueman, and Reid Gilmore. Extreme C-terminal sites are posttranslocationally glycosylated by the STT3B isoform of the OST. Journal of Cell Biology, 201(1):81–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/1/81.

Shimura:2011:EDH

[STI^{+}11] Mari Shimura, Yusuke Toyoda, Kenta Iijima, Masanobu Kinomoto, Kenzo Tokunaga, Kinya Yoda, Mitsuhiro Yanagida, Tetsutaro Sata, and Yukihito Ishizaka. Epigenetic displacement of HP1 from heterochromatin by HIV-1 Vpr causes premature sister chromatid separation. Journal of Cell Biology, 194(5):721–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/5/721.

Schwertz:2010:PPD

[SW10a] Hansjörg Schwertz and Andrew S. Weyrich. Platelet precursors display bipolar behavior. Journal of Cell Biology, 191(4): 699–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/4/699.

Sedmak:2010:ITM

[SW10b] Tina Sedmak and Uwe Wolfrum. Intraflagellar transport molecules in ciliary and nonciliary cells of the retina. Journal of Cell Biology, 189(1):171–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/1/171.

Stehbens:2012:TTH

[SW12] Samantha Stehbens and Torsten Wittmann. Targeting and transport: How microtubules control focal adhesion dynamics. Journal of Cell Biology, 198(4):481–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/4/481.

Swanson:2013:ND

[Swa13] Joel A. Swanson. The noodle defense. Journal of Cell Biology, 203(6):871–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/6/871.

Strack:2013:CDK

[SWC13] Stefan Strack, Theodore J. Wilson, and J. Thomas Cribbs. Cyclin-dependent kinases regulate splice-specific targeting of dynamin-related protein 1 to microtubules. Journal of Cell Biology, 201(7):1037–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/7/1037.

Suozzi:2012:SMO

[SWF12] Kathleen C. Suozzi, Xiaoyang Wu, and Elaine Fuchs. Spectraplakins: Master orchestrators of cytoskeletal dynamics. Journal of Cell Biology, 197(4):465–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/4/465.

Snider:2011:EDG

[SWS^{+}11] Natasha T. Snider, Sujith V. W. Weerasinghe, Amika Singla, Jessica M. Leonard, Shinichiro Hanada, Philip C. Andrews, Anna S. Lok, and M. Bishr Omary. Energy determinants GAPDH and NDPK act as genetic modifiers for hepatocyte inclusion formation. Journal of Cell Biology, 195(2):217–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
195/2/217.

Schulze:2013:LDB

[SWS^{+}13] Ryan J. Schulze, Shaun G. Weller, Barbara Schroeder, Eugene W. Krueger, Susan Chi, Carol A. Casey, and Mark A. McNiven. Lipid droplet breakdown requires Dynamin 2 for vesiculation of autolysosomal tubules in hepatocytes. Journal of Cell Biology, 203(2):315–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/2/315.

Smeenk:2010:NCR

[SWV^{+}10] Godelieve Smeenk, Wouter W. Wiegant, Hans Vrolijk, Aldo P. Solari, Albert Pastink, and Haico van Attikum. The NuRD chromatin–remodeling complex regulates signaling and repair of DNA damage. Journal of Cell Biology, 190(5):741– ??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/5/741.

Smutny:2010:NRC

[SY10] Michael Smutny and Alpha S. Yap. Neighborly relations: cadherins and mechanotransduction. Journal of Cell Biology, 189 (7):1075–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/7/1075.

Shen:2013:SHS

[SYH^{+}13] David Shen, Hua Yuan, Alex Hutagalung, Avani Verma, Daniel Kümmel, Xudong Wu, Karin Reinisch, James A. Mc-New, and Peter Novick. The synaptobrevin homologue Snc2p recruits the exocyst to secretory vesicles by binding to Sec6p. Journal of Cell Biology, 202(3):509–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/202/3/509.

Sen:2011:MSM

[SYK^{+}11] Anindya Sen, Takakazu Yokokura, Mark W. Kankel, Douglas N. Dimlich, Jan Manent, Subhabrata Sanyal, and Spyros Artavanis-Tsakonas. Modeling spinal muscular atrophy in Drosophila links Smn to FGF signaling. Journal of Cell Biology, 192(3):481–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/3/481.

Sen:2013:ILB

[SYS13] Mehmet Sen, Koichi Yuki, and Timothy A. Springer. An internal ligand-bound, metastable state of a leukocyte integrin, α_{X} β_{2}. Journal of Cell Biology, 203(4):629–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
4/629.

Saito:2014:CSE

[SYS^{+}14] Kota Saito, Koh Yamashiro, Noriko Shimazu, Tomoya Tanabe, Kenji Kontani, and Toshiaki Katada. Concentration of Sec12 at ER exit sites via interaction with cTAGE5 is required for collagen export. Journal of Cell Biology, 206(6): 751–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/6/751.

Scherer:2014:PDD

[SYV14] Julian Scherer, Julie Yi, and Richard B. Vallee. PKAdependent dynein switching from lysosomes to adenovirus: a novel form of host–virus competition. Journal of Cell Biology, 205(2):163–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/2/163.

Sarkar:2012:HRI

[SZ12a] Anjali A. Sarkar and Irene E. Zohn. Hectd1 regulates intracellular localization and secretion of Hsp90 to control cellular behavior of the cranial mesenchyme. Journal of Cell Biology, 196 (6):789–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/6/789.

Shiloh:2012:API

[SZ12b] Yosef Shiloh and Yael Ziv. The ATM protein: The importance of being active. Journal of Cell Biology, 198(3):273–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
3/273.

Sideridou:2011:CER

[SZE^{+}11] Maria Sideridou, Roubini Zakopoulou, Konstantinos Evangelou, Michalis Liontos, Athanassios Kotsinas, Emmanouil Rampakakis, Sarantis Gagos, Kaoru Kahata, Kristina Grabusic, Kalliopi Gkouskou, Ioannis P. Trougakos, Evangelos Kolettas, Alexandros G. Georgakilas, Sinisa Volarevic, Aristides G. Eliopoulos, Maria Zannis-Hadjopoulos, Aristidis Moustakas, and Vassilis G. Gorgoulis. Cdc6 expression represses Ecadherin transcription and activates adjacent replication origins. Journal of Cell Biology, 195(7):1123–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
7/1123.

Sens:2010:IPL

[SZJ^{+}10] Kristin L. Sens, Shiliang Zhang, Peng Jin, Rui Duan, Guofeng Zhang, Fengbao Luo, Lauren Parachini, and Elizabeth H. Chen. An invasive podosome-like structure promotes fusion pore formation during myoblast fusion. Journal of Cell Biology, 191(5):1013–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/5/1013.

Sdek:2011:RPC

[SZW^{+}11] Patima Sdek, Peng Zhao, Yaping Wang, Chang jiang Huang, Christopher Y. Ko, Peter C. Butler, James N. Weiss, and W. Robb MacLellan. Rb and p130 control cell cycle gene silencing to maintain the postmitotic phenotype in cardiac myocytes. Journal of Cell Biology, 194(3):407–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
3/407.

Thomanetz:2013:AMC

[TAC^{+}13] Venus Thomanetz, Nico Angliker, Dimitri Cloëtta, Regula M. Lustenberger, Manuel Schweighauser, Filippo Oliveri, Noboru Suzuki, and Markus A. Rüegg. Ablation of the mTORC2 component rictor in brain or Purkinje cells affects size and neuron morphology. Journal of Cell Biology, 201(2):293–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
2/293.

Takabe:2011:DFP

[TAGJ11] Wakako Takabe, Noah Alberts-Grill, and Hanjoong Jo. Disturbed flow: p53 SUMOylation in the turnover of endothelial cells. Journal of Cell Biology, 193(5):805–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/193/5/
805.

Thorpe:2011:BRF

[TALR11] Peter H. Thorpe, David Alvaro, Michael Lisby, and Rodney Rothstein. Bringing Rad52 foci into focus. Journal of Cell Biology, 194(5):665–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/5/665.

Tang:2012:AFR

[TB12] Vivian W. Tang and William M. Brieher. α-Actinin-4/ FSGS1 is required for Arp2/3-dependent actin assembly at the adherens junction. Journal of Cell Biology, 196(1):115–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
196/1/115.

Tang:2013:FCB

[TB13] Vivian W. Tang and William M. Brieher. FSGS3/ CD2AP is a barbed-end capping protein that stabilizes actin and strengthens adherens junctions. Journal of Cell Biology, 203(5):815– ??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/5/815.

Tsygankov:2014:CCP

[TBV^{+}14] Denis Tsygankov, Colleen G. Bilancia, Eric A. Vitriol, Klaus M. Hahn, Mark Peifer, and Timothy C. Elston. CellGeo: a computational platform for the analysis of shape changes in cells with complex geometries. Journal of Cell Biology, 204(3): 443–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/3/443.

Thompson:2010:PAH

[TC10] Sarah L. Thompson and Duane A. Compton. Proliferation of aneuploid human cells is limited by a p53-dependent mechanism. Journal of Cell Biology, 188(3):369–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
3/369.

Turgay:2014:SPF

[TCB^{+}14] Yagmur Turgay, Lysie Champion, Csaba Balazs, Michael Held, Alberto Toso, Daniel W. Gerlich, Patrick Meraldi, and Ulrike Kutay. SUN proteins facilitate the removal of membranes from chromatin during nuclear envelope breakdown. Journal of Cell Biology, 204(7):1099–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/7/1099.

Toret:2014:EDC

[TCN14] Christopher P. Toret, Caitlin Collins, and W. James Nelson. An Elmo–Dock complex locally controls Rho GTPases and actin remodeling during cadherin-mediated adhesion. Journal of Cell Biology, 207(5):577–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/207/5/577.

Tanaka:2010:PPM

[TCX^{+}10] Atsushi Tanaka, Megan M. Cleland, Shan Xu, Derek P. Narendra, Der-Fen Suen, Mariusz Karbowski, and Richard J. Youle. Proteasome and p97 mediate mitophagy and degradation of mitofusins induced by Parkin. Journal of Cell Biology, 191(7): 1367–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/7/1367.

Toret:2014:GWS

[TDV^{+}14] Christopher P. Toret, Michael V. D’Ambrosio, Ronald D. Vale, Michael A. Simon, and W. James Nelson. A genome-wide screen identifies conserved protein hubs required for cadherinmediated cell–cell adhesion. Journal of Cell Biology, 204(2): 265–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/2/265.

Theerthagiri:2010:NNC

[TESA10] Gandhi Theerthagiri, Nathalie Eisenhardt, Heinz Schwarz, and Wolfram Antonin. The nucleoporin Nup188 controls passage of membrane proteins across the nuclear pore complex. Journal of Cell Biology, 189(7):1129–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/7/1129.

Thomson:2010:RFA

[TGB10] Alexander M. Thomson, Peter J. Gillespie, and J. Julian Blow. Replication factory activation can be decoupled from the replication timing program by modulating Cdk levels. Journal of Cell Biology, 188(2):209–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/2/209.

Telley:2012:AMD

[TGES12] Ivo A. Telley, Imre Gáspár, Anne Ephrussi, and Thomas Surrey. Aster migration determines the length scale of nuclear separation in the Drosophila syncytial embryo. Journal of Cell Biology, 197(7):887–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/7/887.

Tamm:2011:BDS

[TGG^{+}11] Tiina Tamm, Agnes Grallert, Emily P. S. Grossman, Isabel Alvarez-Tabares, Frances E. Stevens, and Iain M. Hagan. Brr6 drives the Schizosaccharomyces pombe spindle pole body nuclear envelope insertion/extrusion cycle. Journal of Cell Biology, 195(3):467–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/3/467.

Talamas:2011:PSP

[TH11] Jessica A. Talamas and Martin W. Hetzer. POM121 and Sun1 play a role in early steps of interphase NPC assembly. Journal of Cell Biology, 194(1):27–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/1/27.

Tam:2010:EAS

[TID^{+}10] Christina Tam, Vincent Idone, Cecilia Devlin, Maria Cecilia Fernandes, Andrew Flannery, Xingxuan He, Edward Schuchman, Ira Tabas, and Norma W. Andrews. Exocytosis of acid sphingomyelinase by wounded cells promotes endocytosis and plasma membrane repair. Journal of Cell Biology, 189 (6):1027–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/6/1027.

Titos:2014:CLP

[TIM14] Iris Titos, Tsvetomira Ivanova, and Manuel Mendoza. Chromosome length and perinuclear attachment constrain resolution of DNA intertwines. Journal of Cell Biology, 206(6): 719–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/6/719.

Taguchi:2011:MED

[TIT11] Katsutoshi Taguchi, Takashi Ishiuchi, and Masatoshi Takeichi. Mechanosensitive EPLIN-dependent remodeling of adherens junctions regulates epithelial reshaping. Journal of Cell Biology, 194(4):643–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/4/643.

Taschner:2014:CSI

[TKB^{+}14] Michael Taschner, Fruzsina Kotsis, Philipp Braeuer, E. Wolfgang Kuehn, and Esben Lorentzen. Crystal structures of IFT70/52 and IFT52/46 provide insight into intraflagellar transport B core complex assembly. Journal of Cell Biology, 207(2):269–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/207/2/269.

Tang:2010:EHR

[TKL^{+}10] Daolin Tang, Rui Kang, Kristen M. Livesey, Chun-Wei Cheh, Adam Farkas, Patricia Loughran, George Hoppe, Marco E. Bianchi, Kevin J. Tracey, Herbert J. Zeh, and Michael T. Lotze. Endogenous HMGB1 regulates autophagy. Journal of Cell Biology, 190(5):881–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/5/881.

Tamada:2010:ARS

[TKMK10] Atsushi Tamada, Satoshi Kawase, Fujio Murakami, and Hiroyuki Kamiguchi. Autonomous right-screw rotation of growth cone filopodia drives neurite turning. Journal of Cell Biology, 188(3):429–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/3/429.

Takefuji:2013:RCC

[TKS^{+}13] Mikito Takefuji, Marcus Krüger, Kishor K. Sivaraj, Kozo Kaibuchi, Stefan Offermanns, and Nina Wettschureck. RhoGEF12 controls cardiac remodeling by integrating G protein– and integrin-dependent signaling cascades. Journal of Cell Biology, 201(1):??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/1/i1.

Treusch:2012:IDY

[TL12] Sebastian Treusch and Susan Lindquist. An intrinsically disordered yeast prion arrests the cell cycle by sequestering a spindle pole body component. Journal of Cell Biology, 197 (3):369–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/3/369.

Tilsner:2013:RTP

[TLL^{+}13] Jens Tilsner, Olga Linnik, Marion Louveaux, Ian M. Roberts, Sean N. Chapman, and Karl J. Oparka. Replication and trafficking of a plant virus are coupled at the entrances of plasmodesmata. Journal of Cell Biology, 201(7):981–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
7/981.

Tukachinsky:2010:MVH

[TLS10] Hanna Tukachinsky, Lyle V. Lopez, and Adrian Salic. A mechanism for vertebrate Hedgehog signaling: recruitment to cilia and dissociation of SuFu–Gli protein complexes. Journal of Cell Biology, 191(2):415–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/2/415.

Treuner-Lange:2014:RCP

[TLSA14] Anke Treuner-Lange and Lotte Søgaard-Andersen. Regulation of cell polarity in bacteria. Journal of Cell Biology, 206(1):7–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/1/7.

Tsai:2010:DAE

[TLTW10] Nien-Pei Tsai, Ya-Lun Lin, Yao-Chen Tsui, and Li-Na Wei. Dual action of epidermal growth factor: extracellular signalstimulated nuclear–cytoplasmic export and coordinated translation of selected messenger RNA. Journal of Cell Biology, 188 (3):325–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/3/325.

Thoma:2010:QIA

[TMG^{+}10] Claudio R. Thoma, Alexandre Matov, Katrin L. Gutbrodt, Christian R. Hoerner, Zlatko Smole, Wilhelm Krek, and Gaudenz Danuser. Quantitative image analysis identifies pVHL as a key regulator of microtubule dynamic instability. Journal of Cell Biology, 190(6):991–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/190/6/991.

Trueman:2012:GMT

[TMG12] Steven F. Trueman, Elisabet C. Mandon, and Reid Gilmore. A gating motif in the translocation channel sets the hydrophobicity threshold for signal sequence function. Journal of Cell Biology, 199(6):907–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/199/6/907.

Thon:2010:CMP

[TMPH^{+}10] Jonathan N. Thon, Alejandro Montalvo, Sunita Patel-Hett, Matthew T. Devine, Jennifer L. Richardson, Allen Ehrlicher, Mark K. Larson, Karin Hoffmeister, John H. Hartwig, and Joseph E. Italiano. Cytoskeletal mechanics of proplatelet maturation and platelet release. Journal of Cell Biology, 191(4): 861–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/4/861.

Terrin:2012:PPA

[TMS^{+}12] Anna Terrin, Stefania Monterisi, Alessandra Stangherlin, Anna Zoccarato, Andreas Koschinski, Nicoletta C. Surdo, Marco Mongillo, Akira Sawa, Niove E. Jordanides, Joanne C. Mountford, and Manuela Zaccolo. PKA and PDE4D3 anchoring to AKAP9 provides distinct regulation of cAMP signals at the centrosome. Journal of Cell Biology, 198(4):607–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
198/4/607.

Tao:2011:HMS

[TNH^{+}11] Yazhong Tao, Ronald L. Neppl, Zhan-Peng Huang, Jianfu Chen, Ru-Hang Tang, Ru Cao, Yi Zhang, Suk-Won Jin, and Da-Zhi Wang. The histone methyltransferase Set7/9 promotes myoblast differentiation and myofibril assembly. Journal of Cell Biology, 194(4):551–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/4/551.

Takats:2013:ASD

[TNV^{+}13] Szabolcs Takáts, Péter Nagy, Ágnes Varga, Karolina Pircs, Manuéla Kárpáti, Kata Varga, Attila L. Kovács, Krisztina Hegedűs, and Gábor Juhász. Autophagosomal Syntaxin17dependent lysosomal degradation maintains neuronal function in Drosophila. Journal of Cell Biology, 201(4):531–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
4/531.

TerBush:2012:DFC

[TO12] Allan D. TerBush and Katherine W. Osteryoung. Distinct functions of chloroplast FtsZ1 and FtsZ2 in Z-ring structure and remodeling. Journal of Cell Biology, 199(4):623– ??, November 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/4/623.

Tamura:2013:APC

[TOI^{+}13] Naoki Tamura, Masahide Oku, Moemi Ito, Nobuo N. Noda, Fuyuhiko Inagaki, and Yasuyoshi Sakai. Atg18 phosphoregulation controls organellar dynamics by modulating its phosphoinositide-binding activity. Journal of Cell Biology, 202 (4):685–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/4/685.

Toulmay:2013:DIR

[TP13] Alexandre Toulmay and William A. Prinz. Direct imaging reveals stable, micrometer-scale lipid domains that segregate proteins in live cells. Journal of Cell Biology, 202(1):35–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
202/1/35.

Thon:2012:GHP

[TPM^{+}12] Jonathan N. Thon, Christopher G. Peters, Kellie R. Machlus, Rukhsana Aslam, Jesse Rowley, Hannah Macleod, Matthew T. Devine, Tobias A. Fuchs, Andrew S. Weyrich, John W. Semple, Robert Flaumenhaft, and Joseph E. Italiano. T granules in human platelets function in TLR9 organization and signaling. Journal of Cell Biology, 198(4):561–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/4/561.

Tomatis:2013:MVS

[TPM^{+}13] Vanesa M. Tomatis, Andreas Papadopulos, Nancy T. Malintan, Sally Martin, Tristan Wallis, Rachel S. Gormal, John Kendrick-Jones, Folma Buss, and Frédéric A. Meunier. Myosin VI small insert isoform maintains exocytosis by tethering secretory granules to the cortical actin. Journal of Cell Biology, 200(3):301–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/3/301.

Teekakirikul:2012:HCT

[TPSS12] Polakit Teekakirikul, Robert F. Padera, J. G. Seidman, and Christine E. Seidman. Hypertrophic cardiomyopathy: Translating cellular cross talk into therapeutics. Journal of Cell Biology, 199(3):417–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/3/417.

Taylor:2014:CSR

[TPZ^{+}14] Jackson Taylor, Andrea Pereyra, Tan Zhang, Maria Laura Messi, Zhong-Min Wang, Claudia Hereñú, Pei-Fen Kuan, and Osvaldo Delbono. The Ca_{v} β_{1a} subunit regulates gene expression and suppresses myogenin in muscle progenitor cells. Journal of Cell Biology, 205(6):829–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/205/6/829.

Tripathi:2014:CMR

[TQM^{+}14] Brajendra K. Tripathi, Xiaolan Qian, Philipp Mertins, Dunrui Wang, Alex G. Papageorge, Steven A. Carr, and Douglas R. Lowy. CDK5 is a major regulator of the tumor suppressor DLC1. Journal of Cell Biology, 207(5):627–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/207/
5/627.

Tsun:2011:CDI

[TQS^{+}11] Andy Tsun, Ihjaaz Qureshi, Jane C. Stinchcombe, Misty R. Jenkins, Maike de la Roche, Joanna Kleczkowska, Rose Zamoyska, and Gillian M. Griffiths. Centrosome docking at the immunological synapse is controlled by Lck signaling. Journal of Cell Biology, 192(4):663–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/4/663.

Tauzin:2014:RSF

[TSB^{+}14] Sebastien Tauzin, Taylor W. Starnes, Francisco Barros Becker, Pui ying Lam, and Anna Huttenlocher. Redox and Src family kinase signaling control leukocyte wound attraction and neutrophil reverse migration. Journal of Cell Biology, 207(5): 589–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/5/589.

Tata:2014:IBR

[TSH^{+}14] Aleksandra Tata, David C. Stoppel, Shangyu Hong, Ayal Ben-Zvi, Tiao Xie, and Chenghua Gu. An image-based RNAi screen identifies SH3BP1 as a key effector of Semaphorin 3E– PlexinD1 signaling. Journal of Cell Biology, 205(4):573–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
205/4/573.

Tang:2012:NPA

[TSL12] Xianying Tang, Bryan St. Germain, and Wei-Lih Lee. A novel patch assembly domain in Num1 mediates dynein anchoring at the cortex during spindle positioning. Journal of Cell Biology, 196(6):743–??, March 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/6/743.

Thievessen:2013:VAI

[TTB^{+}13] Ingo Thievessen, Peter M. Thompson, Sylvain Berlemont, Karen M. Plevock, Sergey V. Plotnikov, Alice Zemljic-Harpf, Robert S. Ross, Michael W. Davidson, Gaudenz Danuser, Sharon L. Campbell, and Clare M. Waterman. Vinculin–actin interaction couples actin retrograde flow to focal adhesions, but is dispensable for focal adhesion growth. Journal of Cell Biology, 202(1):163–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/1/163.

Tarantino:2014:TED

[TTC^{+}14] Nadine Tarantino, Jean-Yves Tinevez, Elizabeth Faris Crowell, Bertrand Boisson, Ricardo Henriques, Musa Mhlanga, Fabrice Agou, Alain Israël, and Emmanuel Laplantine. TNF and IL-1 exhibit distinct ubiquitin requirements for inducing NEMO– IKK supramolecular structures. Journal of Cell Biology, 204 (2):231–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/2/231.

Tanaka:2014:HTS

[TTM^{+}14] Chikara Tanaka, Li-Jing Tan, Keisuke Mochida, Hiromi Kirisako, Michiko Koizumi, Eri Asai, Machiko Sakoh-Nakatogawa, Yoshinori Ohsumi, and Hitoshi Nakatogawa. Hrr25 triggers selective autophagy–related pathways by phosphorylating receptor proteins. Journal of Cell Biology, 207 (1):91–??, October 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/207/1/91.

Tien:2010:CDN

[TUG^{+}10] Jerry F. Tien, Neil T. Umbreit, Daniel R. Gestaut, Andrew D. Franck, Jeremy Cooper, Linda Wordeman, Tamir Gonen, Charles L. Asbury, and Trisha N. Davis. Cooperation of the Dam1 and Ndc80 kinetochore complexes enhances microtubule coupling and is regulated by aurora B. Journal of Cell Biology, 189(4):713–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/189/4/713.

Toretsky:2014:AFU

[TW14] Jeffrey A. Toretsky and Peter E. Wright. Assemblages: Functional units formed by cellular phase separation. Journal of Cell Biology, 206(5):579–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/5/579.

Tateishi:2013:TAH

[TYN^{+}13] Kazuhiro Tateishi, Yuji Yamazaki, Tomoki Nishida, Shin Watanabe, Koshi Kunimoto, Hiroaki Ishikawa, and Sachiko Tsukita. Two appendages homologous between basal bodies and centrioles are formed using distinct Odf2 domains. Journal of Cell Biology, 203(3):417–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/3/417.

Ulbricht:2012:PPM

[UAH^{+}12] Tobias Ulbricht, Mohammad Alzrigat, Almut Horch, Nina Reuter, Anna von Mikecz, Viktor Steimle, Eberhard Schmitt, Oliver H. Krämer, Thomas Stamminger, and Peter Hemmerich. PML promotes MHC class II gene expression by stabilizing the class II transactivator. Journal of Cell Biology, 199 (1):49–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/1/49.

Uehara:2010:FCS

[UG10] Ryota Uehara and Gohta Goshima. Functional central spindle assembly requires de novo microtubule generation in the interchromosomal region during anaphase. Journal of Cell Biology, 191(2):259–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/2/259.

Updike:2011:PGE

[UHKS11] Dustin L. Updike, Stephanie J. Hachey, Jeremy Kreher, and Susan Strome. P granules extend the nuclear pore complex environment in the C. elegans germ line. Journal of Cell Biology, 192(6):939–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/6/939.

Upadhyay:2013:SCS

[UKZ^{+}13] Alok Upadhyay, Vasundhara Kandachar, Diana Zitserman, Xin Tong, and Fabrice Roegiers. Sanpodo controls sensory organ precursor fate by directing Notch trafficking and binding γ-secretase. Journal of Cell Biology, 201(3):439–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
3/439.

Uehara:2013:ABK

[UTK^{+}13] Ryota Uehara, Yuki Tsukada, Tomoko Kamasaki, Ina Poser, Kinya Yoda, Daniel W. Gerlich, and Gohta Goshima. Aurora B and Kif2A control microtubule length for assembly of a functional central spindle during anaphase. Journal of Cell Biology, 202(4):623–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/4/623.

VanKeymeulen:2012:TES

[VB12] Alexandra Van Keymeulen and Cédric Blanpain. Tracing epithelial stem cells during development, homeostasis, and repair. Journal of Cell Biology, 197(5):575–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/197/5/
575.

vonBlume:2012:CRC

[vBAK^{+}12] Julia von Blume, Anne-Marie Alleaume, Christine Kienzle, Amado Carreras-Sureda, Miguel Valverde, and Vivek Malhotra. Cab45 is required for Ca^{2+}-dependent secretory cargo sorting at the trans-Golgi network. Journal of Cell Biology, 199(7):1057–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/7/1057.

Valacca:2010:SRE

[VBB^{+}10] Cristina Valacca, Serena Bonomi, Emanuele Buratti, Simona Pedrotti, Francisco Ernesto Baralle, Claudio Sette, Claudia Ghigna, and Giuseppe Biamonti. Sam68 regulates EMT through alternative splicing–activated nonsensemediated mRNA decay of the SF2/ASF proto-oncogene. Journal of Cell Biology, 191(1):87–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/1/87.

VandenAbeele:2013:FIC

[VBG^{+}13] Fabien Vanden Abeele, Gabriel Bidaux, Dmitri Gordienko, Benjamin Beck, Yuri V. Panchin, Ancha V. Baranova, Dmitry V. Ivanov, Roman Skryma, and Natalia Prevarskaya. Functional implications of calcium permeability of the channel formed by pannexin 1. Journal of Cell Biology, 201(5):777–??, May 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
201/5/777.

Vernieri:2013:ASC

[VCF^{+}13] Claudio Vernieri, Elena Chiroli, Valentina Francia, Fridolin Gross, and Andrea Ciliberto. Adaptation to the spindle checkpoint is regulated by the interplay between Cdc28/Clbs and PP2A^{Cdc55}. Journal of Cell Biology, 202(5):765–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
5/765.

vandeBospoort:2012:MCL

[vdBFS^{+}12] Rhea van de Bospoort, Margherita Farina, Sabine K. Schmitz, Arthur de Jong, Heidi de Wit, Matthijs Verhage, and Ruud F. Toonen. Munc13 controls the location and efficiency of densecore vesicle release in neurons. Journal of Cell Biology, 199(6): 883–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/6/883.

vanderVaart:2011:SLM

[vdVMG^{+}11] Babet van der Vaart, Cristina Manatschal, Ilya Grigoriev, Vincent Olieric, Susana Montenegro Gouveia, Saša Bjelić, Jeroen Demmers, Ivan Vorobjev, Casper C. Hoogenraad, Michel O. Steinmetz, and Anna Akhmanova. SLAIN2 links microtubule plus end–tracking proteins and controls microtubule growth in interphase. Journal of Cell Biology, 193(6):1083–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
6/1083.

Vidal-Eychenie:2013:DSS

[VEDBC13] Sophie Vidal-Eychenié, Chantal Décaillet, Jihane Basbous, and Angelos Constantinou. DNA structure-specific priming of ATR activation by DNA–PKcs. Journal of Cell Biology, 202 (3):421–??, August 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/3/421.

Viatour:2011:NSI

[VES^{+}11] Patrick Viatour, Ursula Ehmer, Louis A. Saddic, Craig Dorrell, Jesper B. Andersen, Chenwei Lin, Anne-Flore Zmoos, Pawel K. Mazur, Bethany E. Schaffer, Austin Ostermeier, Hannes Vogel, Karl G. Sylvester, Snorri S. Thorgeirsson, Markus Grompe, and Julien Sage. Notch signaling inhibits hepatocellular carcinoma following inactivation of the RB pathway. Journal of Cell Biology, 194(5):??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
5/i11.

Vega:2011:RRD

[VFNR11] Francisco M. Vega, Gilbert Fruhwirth, Tony Ng, and Anne J. Ridley. RhoA and RhoC have distinct roles in migration and invasion by acting through different targets. Journal of Cell Biology, 193(4):655–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/4/655.

Vitureira:2013:IBH

[VG13] Nathalia Vitureira and Yukiko Goda. The interplay between Hebbian and homeostatic synaptic plasticity. Journal of Cell Biology, 203(2):175–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/203/2/175.

vanGalen:2014:SHR

[vGCMA^{+}14] Josse van Galen, Felix Campelo, Emma Martínez-Alonso, Margherita Scarpa, José Ángel Martínez-Menárguez, and Vivek Malhotra. Sphingomyelin homeostasis is required to form functional enzymatic domains at the trans-Golgi network. Journal of Cell Biology, 206(5):609–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
5/609.

Vassilopoulos:2014:ASC

[VGL^{+}14] Stéphane Vassilopoulos, Christel Gentil, Jeanne Lainé, Pierre-Olivier Buclez, Agathe Franck, Arnaud Ferry, Guillaume Précigout, Robyn Roth, John E. Heuser, Frances M. Brodsky, Luis Garcia, Gisèle Bonne, Thomas Voit, France Piétri-Rouxel, and Marc Bitoun. Actin scaffolding by clathrin heavy chain is required for skeletal muscle sarcomere organization. Journal of Cell Biology, 205(3):377–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/205/3/377.

vanGisbergen:2012:CIF

[vGLWB12] Peter A. C. van Gisbergen, Ming Li, Shu-Zon Wu, and Magdalena Bezanilla. Class II formin targeting to the cell cortex by binding PI(3,5)P_{2} is essential for polarized growth. Journal of Cell Biology, 198(2):235–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/2/235.

Velichkova:2010:DMCa

[VJK^{+}10a] Michaella Velichkova, Joe Juan, Pavan Kadandale, Steve Jean, Inês Ribeiro, Vignesh Raman, Chris Stefan, and Amy A. Kiger. Drosophila Mtm and class II PI3K coregulate a PI(3)P pool with cortical and endolysosomal functions. Journal of Cell Biology, 190(3):407–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/3/407.

Velichkova:2010:DMCb

[VJK^{+}10b] Michaella Velichkova, Joe Juan, Pavan Kadandale, Steve Jean, Inês Ribeiro, Vignesh Raman, Chris Stefan, and Amy A. Kiger. Drosophila Mtm and class II PI3K coregulate a PI(3)P pool with cortical and endolysosomal functions. Journal of Cell Biology, 190(5):941–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/190/5/941.

Veltman:2012:SKD

[VKMI12] Douwe M. Veltman, Jason S. King, Laura M. Machesky, and Robert H. Insall. SCAR knockouts in Dictyostelium: WASP assumes SCAR’s position and upstream regulators in pseudopods. Journal of Cell Biology, 198(4):501–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
4/501.

Vermillion:2014:CPM

[VLG14] Katie L. Vermillion, Kevin A. Lidberg, and Laura S. Gammill. Cytoplasmic protein methylation is essential for neural crest migration. Journal of Cell Biology, 204(1):95–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
1/95.

Vogler:2014:CFA

[VLI^{+}14] Georg Vogler, Jiandong Liu, Timothy W. Iafe, Ede Migh, József Mihály, and Rolf Bodmer. Cdc42 and formin activity control non-muscle myosin dynamics during Drosophila heart morphogenesis. Journal of Cell Biology, 206(7):909– ??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/7/909.

Veltman:2014:PDM

[VLKI14] Douwe M. Veltman, Michael G. Lemieux, David A. Knecht, and Robert H. Insall. PIP_{3}-dependent macropinocytosis is incompatible with chemotaxis. Journal of Cell Biology, 204(4): 497–??, February 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/204/4/497.

Vavassori:2014:NLO

[VM14] Stefano Vavassori and Andreas Mayer. A new life for an old pump: V–ATPase and neurotransmitter release. Journal of Cell Biology, 205(1):7–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/205/1/7.

Vicente-Manzanares:2011:MII

[VMNLB^{+}11] Miguel Vicente-Manzanares, Karen Newell-Litwa, Alexia I. Bachir, Leanna A. Whitmore, and Alan Rick Horwitz. Myosin IIA/IIB restrict adhesive and protrusive signaling to generate front–back polarity in migrating cells. Journal of Cell Biology, 193(2):381–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/2/381.

Viswanatha:2012:LPM

[VOSB12] Raghuvir Viswanatha, Patrice Y. Ohouo, Marcus B. Smolka, and Anthony Bretscher. Local phosphocycling mediated by LOK/SLK restricts ezrin function to the apical aspect of epithelial cells. Journal of Cell Biology, 199(6):969–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/199/
6/969.

VanHook:2014:CER

[VPC^{+}14] Matthew J. Van Hook, Caitlyn M. Parmelee, Minghui Chen, Karlene M. Cork, Carina Curto, and Wallace B. Thoreson. Calmodulin enhances ribbon replenishment and shapes filtering of synaptic transmission by cone photoreceptors. Journal of Cell Biology, 207(3):??, November 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/207/3/2073OIA192.

vanRee:2010:OEU

[vRJMvD10] Janine H. van Ree, Karthik B. Jeganathan, Liviu Malureanu, and Jan M. van Deursen. Overexpression of the E2 ubiquitin– conjugating enzyme UbcH10 causes chromosome missegregation and tumor formation. Journal of Cell Biology, 188(1): 83–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/1/83.

Vernay:2012:SPB

[VSG^{+}12] Aurélia Vernay, Sébastien Schaub, Isabelle Guillas, Martine Bassilana, and Robert A. Arkowitz. A steep phosphoinositide bis-phosphate gradient forms during fungal filamentous growth. Journal of Cell Biology, 198(4):711–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
4/711.

Vachova:2011:FDE

[VŠH^{+}11] Libuše Váchová, Vratislav Šťovíček, Otakar Hlaváček, Oleksandr Chernyavskiy, Luděk Štěpánek, Lucie Kubínová, and Zdena Palková. Flo11p, drug efflux pumps, and the extracellular matrix cooperate to form biofilm yeast colonies. Journal of Cell Biology, 194(5):679–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/194/5/679.

Valerio-Santiago:2011:TLS

[VSMC11] Mauricio Valerio-Santiago and Fernando Monje-Casas. Tem1 localization to the spindle pole bodies is essential for mitotic exit and impairs spindle checkpoint function. Journal of Cell Biology, 192(4):599–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/192/4/599.

Vasquez:2014:DMP

[VTM14] Claudia G. Vasquez, Mike Tworoger, and Adam C. Martin. Dynamic myosin phosphorylation regulates contractile pulses and tissue integrity during epithelial morphogenesis. Journal of Cell Biology, 206(3):435–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/3/435.

Vleugel:2013:ABR

[VTO^{+}13] Mathijs Vleugel, Eelco Tromer, Manja Omerzu, Vincent Groenewold, Wilco Nijenhuis, Berend Snel, and Geert J. P. L. Kops. Arrayed BUB recruitment modules in the kinetochore scaffold KNL1 promote accurate chromosome segregation. Journal of Cell Biology, 203(6):943–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/6/943.

Vizeacoumar:2010:IHT

[VvDV^{+}10] Franco J. Vizeacoumar, Nydia van Dyk, Frederick S. Vizeacoumar, Vincent Cheung, Jingjing Li, Yaroslav Sydorskyy, Nicolle Case, Zhijian Li, Alessandro Datti, Corey Nislow, Brian Raught, Zhaolei Zhang, Brendan Frey, Kerry Bloom, Charles Boone, and Brenda J. Andrews. Integrating highthroughput genetic interaction mapping and high-content screening to explore yeast spindle morphogenesis. Journal of Cell Biology, 188(1):69–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/1/69.

Varma:2013:SAC

[VWC^{+}13] Dileep Varma, Xiaohu Wan, Dhanya Cheerambathur, Reto Gassmann, Aussie Suzuki, Josh Lawrimore, Arshad Desai, and E. D. Salmon. Spindle assembly checkpoint proteins are positioned close to core microtubule attachment sites at kinetochores. Journal of Cell Biology, 202(5):735–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
5/735.

VandeWalle:2013:SNR

[VWD^{+}13] Inge Van de Walle, Els Waegemans, Jelle De Medts, Greet De Smet, Magda De Smedt, Sylvia Snauwaert, Bart Vandekerckhove, Tessa Kerre, Georges Leclercq, Jean Plum, Thomas Gridley, Tao Wang, Ute Koch, Freddy Radtke, and Tom Taghon. Specific Notch receptor–ligand interactions control human TCR-ab/gd development by inducing differential Notch signal strength. Journal of Cell Biology, 201(1):??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/201/1/i2.

Vacher:2011:CMP

[VYC^{+}11] Hélène Vacher, Jae-Won Yang, Oscar Cerda, Amapola Autillo-Touati, Bénédicte Dargent, and James S. Trimmer. Cdkmediated phosphorylation of the Kvβ2 auxiliary subunit regulates Kv1 channel axonal targeting. Journal of Cell Biology, 192(5):813–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/5/813.

Van:2010:CPS

[VYM^{+}10] Christopher Van, Shan Yan, W. Matthew Michael, Shou Waga, and Karlene A. Cimprich. Continued primer synthesis at stalled replication forks contributes to checkpoint activation. Journal of Cell Biology, 189(2):233–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/2/233.

vanZon:2010:ACR

[vZOtR^{+}10] Wouter van Zon, Janneke Ogink, Bas ter Riet, René H. Medema, Hein te Riele, and Rob M. F. Wolthuis. The APC/ C recruits cyclin B1–Cdk1–Cks in prometaphase before D box recognition to control mitotic exit. Journal of Cell Biology, 190 (4):587–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/4/587.

Wirt:2010:GAD

[WAG^{+}10] Stacey E. Wirt, Adam S. Adler, Véronique Gebala, James M. Weimann, Bethany E. Schaffer, Louis A. Saddic, Patrick Viatour, Hannes Vogel, Howard Y. Chang, Alex Meissner, and Julien Sage. G1 arrest and differentiation can occur independently of Rb family function. Journal of Cell Biology, 191(4): 809–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/4/809.

Welf:2012:MFR

[WAJ^{+}12] Erik S. Welf, Shoeb Ahmed, Heath E. Johnson, Adam T. Melvin, and Jason M. Haugh. Migrating fibroblasts reorient directionality by a metastable, PI3K-dependent mechanism. Journal of Cell Biology, 197(1):105–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/197/1/105.

Wemmer:2011:BBS

[WAW^{+}11] Megan Wemmer, Ishara Azmi, Matthew West, Brian Davies, David Katzmann, and Greg Odorizzi. Bro1 binding to Snf7 regulates ESCRT–III membrane scission activity in yeast. Journal of Cell Biology, 192(2):295–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/192/2/295.

Weissbein:2014:GMP

[WBBD14] Uri Weissbein, Nissim Benvenisty, and Uri Ben-David. Genome maintenance in pluripotent stem cells. Journal of Cell Biology, 204(2):153–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/2/153.

White:2011:DHL

[WBcY^{+}11] Anne E. White, Brandon D. Burch, Xiao cui Yang, Pamela Y. Gasdaska, Zbigniew Dominski, William F. Marzluff, and Robert J. Duronio. Drosophila histone locus bodies form by hierarchical recruitment of components. Journal of Cell Biology, 193(4):677–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/4/677.

Wang:2011:ABD

[WBL11] Enxiu Wang, Edward R. Ballister, and Michael A. Lampson. Aurora B dynamics at centromeres create a diffusion-based phosphorylation gradient. Journal of Cell Biology, 194(4): 539–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/194/4/539.

Wang:2013:PMS

[WBMCSS13] Haifeng Wang, Ingrid Brust-Mascher, Gul Civelekoglu-Scholey, and Jonathan M. Scholey. Patronin mediates a switch from kinesin-13-dependent poleward flux to anaphase B spindle elongation. Journal of Cell Biology, 203(1):35–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
1/35.

Windoffer:2011:CMD

[WBML11] Reinhard Windoffer, Michael Beil, Thomas M. Magin, and Rudolf E. Leube. Cytoskeleton in motion: the dynamics of keratin intermediate filaments in epithelia. Journal of Cell Biology, 194(5):669–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/194/5/669.

Wang:2011:CTT

[WBS11] Pengbo Wang, Christoph Ballestrem, and Charles H. Streuli. The C terminus of talin links integrins to cell cycle progression. Journal of Cell Biology, 195(3):499–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/3/499.

Wandke:2012:HCP

[WBS^{+}12] Cornelia Wandke, Marin Barisic, Reinhard Sigl, Veronika Rauch, Frank Wolf, Ana C. Amaro, Chia H. Tan, Antonio J. Pereira, Ulrike Kutay, Helder Maiato, Patrick Meraldi, and Stephan Geley. Human chromokinesins promote chromosome congression and spindle microtubule dynamics during mitosis. Journal of Cell Biology, 198(5):847–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
5/847.

Winslow:2010:SIM

[WCC^{+}10] Ashley R. Winslow, Chien-Wen Chen, Silvia Corrochano, Abraham Acevedo-Arozena, David E. Gordon, Andrew A. Peden, Maike Lichtenberg, Fiona M. Menzies, Brinda Ravikumar, Sara Imarisio, Steve Brown, Cahir J. O’Kane, and David C. Rubinsztein. α-synuclein impairs macroautophagy: implications for Parkinson’s disease. Journal of Cell Biology, 190(6):1023–??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/6/1023.

Williams:2012:JDS

[WCM12a] Elizabeth H. Williams, Pamela Carpentier, and Tom Misteli. The JCB DataViewer scales up. Journal of Cell Biology, 198 (3):271–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/3/271.

Williams:2012:MRR

[WCM12b] Elizabeth H. Williams, Pamela A. Carpentier, and Tom Misteli. Minimizing the “Re” in Review. Journal of Cell Biology, 197(3):345–??, April 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/3/345.

Winbanks:2013:BMP

[WCQ^{+}13] Catherine E. Winbanks, Justin L. Chen, Hongwei Qian, Yingying Liu, Bianca C. Bernardo, Claudia Beyer, Kevin I. Watt, Rachel E. Thomson, Timothy Connor, Bradley J. Turner, Julie R. McMullen, Lars Larsson, Sean L. McGee, Craig A. Harrison, and Paul Gregorevic. The bone morphogenetic protein axis is a positive regulator of skeletal muscle mass. Journal of Cell Biology, 203(2):345–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/2/345.

Wilson:2011:FEN

[WD11] Katherine L. Wilson and Scott C. Dawson. Functional evolution of nuclear structure. Journal of Cell Biology, 195(2): 171–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/2/171.

Woodruff:2010:MSD

[WDB10] Jeffrey B. Woodruff, David G. Drubin, and Georjana Barnes. Mitotic spindle disassembly occurs via distinct subprocesses driven by the anaphase-promoting complex, Aurora B kinase, and kinesin-8. Journal of Cell Biology, 191(4):795–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
191/4/795.

Worth:2013:DCC

[WDG^{+}13] Daniel C. Worth, Catherine N. Daly, Sara Geraldo, Fazal Oozeer, and Phillip R. Gordon-Weeks. Drebrin contains a cryptic F-actin–bundling activity regulated by Cdk5 phosphorylation. Journal of Cell Biology, 202(5):793–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
5/793.

Wang:2014:CCR

[WEK^{+}14] Dong Wang, Daniel Epstein, Ossama Khalaf, Sankaranarayanan Srinivasan, W. Ryan Williamson, Amir Fayyazuddin, Florante A. Quiocho, and P. Robin Hiesinger. Ca^{2+}- calmodulin regulates SNARE assembly and spontaneous neurotransmitter release via v-ATPase subunit V0a1. Journal of Cell Biology, 205(1):21–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/205/1/21.

Fong:2013:WGS

[wFLW^{+}13] Ka wing Fong, Yujing Li, Wenqi Wang, Wenbin Ma, Kunpeng Li, Robert Z. Qi, Dan Liu, Zhou Songyang, and Junjie Chen. Whole-genome screening identifies proteins localized to distinct nuclear bodies. Journal of Cell Biology, 203(1): 149–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/1/149.

Wickstead:2011:EC

[WG11] Bill Wickstead and Keith Gull. The evolution of the cytoskeleton. Journal of Cell Biology, 194(4):513–??, August 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
4/513.

Wang:2011:MCT

[WGC11] Jiadong Wang, Zihua Gong, and Junjie Chen. MDC1 collaborates with TopBP1 in DNA replication checkpoint control. Journal of Cell Biology, 193(2):267–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/2/267.

Wu:2012:ATC

[WGM^{+}12] Yao Wu, Yiwen Gu, Mary K. Morphew, Jun Yao, Felix L. Yeh, Min Dong, and Edwin R. Chapman. All three components of the neuronal SNARE complex contribute to secretory vesicle docking. Journal of Cell Biology, 198(3):323–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
3/323.

Wang:2013:CCR

[WGN^{+}13] Peng Wang, Jacob A. Galan, Karine Normandin, Éric Bonneil, Gilles R. Hickson, Philippe P. Roux, Pierre Thibault, and Vincent Archambault. Cell cycle regulation of Greatwall kinase nuclear localization facilitates mitotic progression. Journal of Cell Biology, 202(2):277–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/202/2/277.

Whitehouse:2012:MSH

[WGR^{+}12] Sarah Whitehouse, Vicki A. M. Gold, Alice Robson, William J. Allen, Richard B. Sessions, and Ian Collinson. Mobility of the SecA 2-helix-finger is not essential for polypeptide translocation via the SecYEG complex. Journal of Cell Biology, 199(6): 919–??, December 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/199/6/919.

Wang:2013:VAM

[WH13] Dong Wang and P. Robin Hiesinger. The vesicular ATPase: a missing link between acidification and exocytosis. Journal of Cell Biology, 203(2):171–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/203/2/171.

Wu:2013:LAI

[WHA^{+}13] Congying Wu, Elizabeth M. Haynes, Sreeja B. Asokan, Jeremy M. Simon, Norman E. Sharpless, Albert S. Baldwin, Ian J. Davis, Gary L. Johnson, and James E. Bear. Loss of Arp2/3 induces an NF-κB-dependent, nonautonomous effect on chemotactic signaling. Journal of Cell Biology, 203(6):

907–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/6/907.

Worth:2010:ISR

[WHDR^{+}10] Daniel C. Worth, Kairbaan Hodivala-Dilke, Stephen D. Robinson, Samantha J. King, Penny E. Morton, Frank B. Gertler, Martin J. Humphries, and Maddy Parsons. α vβ3 integrin spatially regulates VASP and RIAM to control adhesion dynamics and migration. Journal of Cell Biology, 189(2):369–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
2/369.

Wang:2011:DEN

[WHF^{+}11] Zhimin Wang, Ceba Humphrey, Nicole Frilot, Gaofeng Wang, Zhongzhen Nie, Nader H. Moniri, and Yehia Daaka. Dynamin2- and endothelial nitric oxide synthase–regulated invasion of bladder epithelial cells by uropathogenic Escherichia coli. Journal of Cell Biology, 192(1):101–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/192/1/
101.

Werner:2011:AMD

[WHH^{+}11] Michael E. Werner, Peter Hwang, Fawn Huisman, Peter Taborek, Clare C. Yu, and Brian J. Mitchell. Actin and microtubules drive differential aspects of planar cell polarity in multiciliated cells. Journal of Cell Biology, 195(1):19–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
1/19.

Wurzenberger:2012:SRM

[WHL^{+}12] Claudia Wurzenberger, Michael Held, Michael A. Lampson, Ina Poser, Anthony A. Hyman, and Daniel W. Gerlich. Sds22 and Repo–Man stabilize chromosome segregation by counteracting Aurora B on anaphase kinetochores. Journal of Cell Biology, 198(2):173–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/2/173.

Woo:2012:NSR

[WHWS12] Stephanie Woo, Michael P. Housley, Orion D. Weiner, and Didier Y. R. Stainier. Nodal signaling regulates endodermal cell motility and actin dynamics via Rac1 and Prex1. Journal of Cell Biology, 198(5):941–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/198/5/941.

Wee:2011:CBR

[WJPD11] Brett Wee, Christopher A. Johnston, Kenneth E. Prehoda, and Chris Q. Doe. Canoe binds RanGTP to promote Pins^{TPR}/ Mud-mediated spindle orientation. Journal of Cell Biology, 195(3):369–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/195/3/369.

Willenborg:2011:IBF

[WJW^{+}11] Carly Willenborg, Jian Jing, Christine Wu, Hugo Matern, Jerome Schaack, Jemima Burden, and Rytis Prekeris. Interaction between FIP5 and SNX18 regulates epithelial lumen formation. Journal of Cell Biology, 195(1):71–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/195/1/71.

Wang:2012:IUP

[WK12] Shiyu Wang and Randal J. Kaufman. The impact of the unfolded protein response on human disease. Journal of Cell Biology, 197(7):857–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/7/857.

Wolf:2010:PST

[WKGB^{+}10] Annika Wolf, Malgorzata Krause-Gruszczynska, Olaf Birkenmeier, Antje Ostareck-Lederer, Stefan Hüttelmaier, and Mechthild Hatzfeld. Plakophilin 1 stimulates translation by promoting eIF4A1 activity. Journal of Cell Biology, 188(4): 463–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/4/463.

Woo:2013:API

[WKN^{+}13] Jooyeon Woo, Seok-Kyu Kwon, Jungyong Nam, Seungwon Choi, Hideto Takahashi, Dilja Krueger, Joohyun Park, Yeunkum Lee, Jin Young Bae, Dongmin Lee, Jaewon Ko, Hyun Kim, Myoung-Hwan Kim, Yong Chul Bae, Sunghoe Chang, Ann Marie Craig, and Eunjoon Kim. The adhesion protein IgSF9b is coupled to neuroligin 2 via S–SCAM to promote inhibitory synapse development. Journal of Cell Biology, 201 (6):929–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/6/929.

Wang:2011:RSM

[WLGC11] Zhao Wang, Huisheng Liu, Yiwen Gu, and Edwin R. Chapman. Reconstituted synaptotagmin I mediates vesicle docking, priming, and fusion. Journal of Cell Biology, 195(7): 1159–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/195/7/1159.

Williams:2011:MNM

[WLK^{+}11] Corey L. Williams, Chunmei Li, Katarzyna Kida, Peter N. Inglis, Swetha Mohan, Lucie Semenec, Nathan J. Bialas, Rachel M. Stupay, Nansheng Chen, Oliver E. Blacque, Bradley K. Yoder, and Michel R. Leroux. MKS and NPHP modules cooperate to establish basal body/transition zone membrane associations and ciliary gate function during ciliogenesis. Journal of Cell Biology, 192(6):1023–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
6/1023.

Wang:2014:UNS

[WLN^{+}14] Zheng Wang, Lara M. Linden, Kaleb M. Naegeli, Joshua W. Ziel, Qiuyi Chi, Elliott J. Hagedorn, Natasha S. Savage, and David R. Sherwood. UNC-6 (netrin) stabilizes oscillatory clustering of the UNC-40 (DCC) receptor to orient polarity. Journal of Cell Biology, 206(5):619–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/206/5/619.

Wu:2011:VPT

[WLW11] Chih-Hang Wu, Shu-Chuan Lee, and Chao-Wen Wang. Viral protein targeting to the cortical endoplasmic reticulum is required for cell–cell spreading in plants. Journal of Cell Biology, 193(3):521–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/3/521.

Wang:2014:NPR

[WLZ^{+}14] Ning Wang, Libera Lo Presti, Yi-Hua Zhu, Minhee Kang, Zhengrong Wu, Sophie G. Martin, and Jian-Qiu Wu. The novel proteins Rng8 and Rng9 regulate the myosin–V Myo51 during fission yeast cytokinesis. Journal of Cell Biology, 205 (3):357–??, May 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/205/3/357.

Walther:2010:MSB

[WM10] Tobias C. Walther and Matthias Mann. Mass spectrometrybased proteomics in cell biology. Journal of Cell Biology, 190 (4):491–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/4/491.

Williams:2011:NTJ

[WM11] Elizabeth H. Williams and Tom Misteli. New Tools for JCB. Journal of Cell Biology, 194(5):663–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
5/663.

Wang:2012:IMD

[WM12] Yu Wang and Mark A. McNiven. Invasive matrix degradation at focal adhesions occurs via protease recruitment by a FAK–p130Cas complex. Journal of Cell Biology, 196(3): 375–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/196/3/375.

Wiedemann:2010:ICT

[WMB^{+}10] Sonja M. Wiedemann, Silke N. Mildner, Clemens Bönisch, Lars Israel, Andreas Maiser, Sarah Matheisl, Tobias Straub, Rainer Merkl, Heinrich Leonhardt, Elisabeth Kremmer, Lothar Schermelleh, and Sandra B. Hake. Identification and characterization of two novel primate-specific histone H3 variants, H3.X and H3.Y. Journal of Cell Biology, 190(5):777– ??, September 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/190/5/777.

Wang:2012:ADM

[WMB12] Jack T. Wang, Zachary A. Medress, and Ben A. Barres. Axon degeneration: Molecular mechanisms of a self-destruction pathway. Journal of Cell Biology, 196(1):7–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/196/1/7.

Wang:2014:SEV

[WMC14] Chao-Wen Wang, Yu-Hsuan Miao, and Yi-Shun Chang. A sterol-enriched vacuolar microdomain mediates stationary phase lipophagy in budding yeast. Journal of Cell Biology, 206 (3):357–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/3/357.

Webster:2010:VTM

[WMCF10] Micah T. Webster, J. Michael McCaffery, and Orna Cohen-Fix. Vesicle trafficking maintains nuclear shape in Saccharomyces cerevisiae during membrane proliferation. Journal of Cell Biology, 191(6):1079–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/191/6/1079.

Werner:2014:RIR

[WMP^{+}14] Michael E. Werner, Jennifer W. Mitchell, William Putzbach, Elizabeth Bacon, Sun K. Kim, and Brian J. Mitchell. Radial intercalation is regulated by the Par complex and the microtubule-stabilizing protein CLAMP/Spef1. Journal of Cell Biology, 206(3):367–??, August 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/206/3/367.

Winkle:2014:NNS

[WMV^{+}14] Cortney C. Winkle, Leslie M. McClain, Juli G. Valtschanoff, Charles S. Park, Christopher Maglione, and Stephanie L. Gupton. A novel Netrin-1–sensitive mechanism promotes local SNARE-mediated exocytosis during axon branching. Journal of Cell Biology, 205(2):217–??, April 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/205/2/217.

Wang:2013:GPN

[WOG13] Dong Wang, Damien O’Halloran, and Miriam B. Goodman. GCY-8, PDE-2, and NCS-1 are critical elements of the cGMPdependent thermotransduction cascade in the AFD neurons responsible for C. elegans thermotaxis. Journal of Cell Biology, 203(1):??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/1/2031OIA114.

Walters:2014:TQC

[WP14] Robert Walters and Roy Parker. Is there quality control of localized mRNAs? Journal of Cell Biology, 204(6):863–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
204/6/863.

Wilson:2011:GDCa

[WPL^{+}11a] Deanna G. Wilson, Khanhky Phamluong, Li Li, Mei Sun, Tim C. Cao, Peter S. Liu, Zora Modrusan, Wendy N. Sandoval, Linda Rangell, Richard A. D. Carano, Andrew S. Peterson, and Mark J. Solloway. Global defects in collagen secretion in a Mia3/TANGO1 knockout mouse. Journal of Cell Biology, 193(5):935–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/193/5/935.

Wilson:2011:GDCb

[WPL^{+}11b] Deanna G. Wilson, Khanhky Phamluong, Li Li, Mei Sun, Tim C. Cao, Peter S. Liu, Zora Modrusan, Wendy N. Sandoval, Linda Rangell, Richard A. D. Carano, Andrew S. Peterson, and Mark J. Solloway. Global defects in collagen secretion in a Mia3/TANGO1 knockout mouse. Journal of Cell Biology, 194(2):347–??, July 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/194/2/347.

Wang:2014:MMM

[WPM14] Xiaoxi Wang and Andrea Page-McCaw. A matrix metalloproteinase mediates long-distance attenuation of stem cell proliferation. Journal of Cell Biology, 206(7):923–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
7/923.

Weigert:2013:ICB

[WPSA13] Roberto Weigert, Natalie Porat-Shliom, and Panomwat Amornphimoltham. Imaging cell biology in live animals: Ready for prime time. Journal of Cell Biology, 201(7):969–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
201/7/969.

Wynne:2012:DDP

[WRCD12] David J. Wynne, Ofer Rog, Peter M. Carlton, and Abby F. Dernburg. Dynein-dependent processive chromosome motions promote homologous pairing in C. elegans meiosis. Journal of Cell Biology, 196(1):47–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/1/47.

Wang:2013:MMD

[WRF^{+}13] Songyu Wang, Fabian B. Romano, Christine M. Field, Tim J. Mitchison, and Tom A. Rapoport. Multiple mechanisms determine ER network morphology during the cell cycle in Xenopus egg extracts. Journal of Cell Biology, 203(5):801–??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
5/801.

Whittle:2010:SSS

[WS10] James R. R. Whittle and Thomas U. Schwartz. Structure of the Sec13–Sec16 edge element, a template for assembly of the COPII vesicle coat. Journal of Cell Biology, 190(3):347–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
190/3/347.

Wang:2011:CCC

[WSUT11] Won-Jing Wang, Rajesh Kumar Soni, Kunihiro Uryu, and Meng-Fu Bryan Tsou. The conversion of centrioles to centrosomes: essential coupling of duplication with segregation. Journal of Cell Biology, 193(4):727–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/4/727.

Wei:2012:MIO

[WSZ^{+}12] Jianwen Wei, Yu Shi, Lihua Zheng, Bin Zhou, Hiroyuki Inose, Ji Wang, X. Edward Guo, Rudolf Grosschedl, and Gerard Karsenty. miR-34s inhibit osteoblast proliferation and differentiation in the mouse by targeting SATB2. Journal of Cell Biology, 197(4):509–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/4/509.

Winkler:2012:HTH

[WTBM12] Juliane Winkler, Jens Tyedmers, Bernd Bukau, and Axel Mogk. Hsp70 targets Hsp100 chaperones to substrates for protein disaggregation and prion fragmentation. Journal of Cell Biology, 198(3):387–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/3/387.

Wen:2011:VHI

[WTH^{+}11] Lei Wen, Fu-Lei Tang, Yan Hong, Shi-Wen Luo, Chun-Lei Wang, Wanxia He, Chengyong Shen, Ji-Ung Jung, Fei Xiong, Dae hoon Lee, Quan-Guang Zhang, Darrell Brann, Tae-Wan Kim, Riqiang Yan, Lin Mei, and Wen-Cheng Xiong. VPS35 haploinsufficiency increases Alzheimer’s disease neuropathology. Journal of Cell Biology, 195(5):765–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
5/765.

Wolf:2013:PLC

[WtLK^{+}13] Katarina Wolf, Mariska te Lindert, Marina Krause, Stephanie Alexander, Joost te Riet, Amanda L. Willis, Robert M. Hoffman, Carl G. Figdor, Stephen J. Weiss, and Peter Friedl. Physical limits of cell migration: Control by ECM space and nuclear deformation and tuning by proteolysis and traction force.

Journal of Cell Biology, 201(7):1069–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/7/1069.

Wang:2012:HIR

[WUD^{+}12] Fangwei Wang, Natalia P. Ulyanova, John R. Daum, Debasis Patnaik, Anna V. Kateneva, Gary J. Gorbsky, and Jonathan M. G. Higgins. Haspin inhibitors reveal centromeric functions of Aurora B in chromosome segregation. Journal of Cell Biology, 199(2):251–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/2/251.

Wloka:2013:IMI

[WVT^{+}13] Carsten Wloka, Elizabeth A. Vallen, Lydia Thé, Xiaodong Fang, Younghoon Oh, and Erfei Bi. Immobile myosin–II plays a scaffolding role during cytokinesis in budding yeast. Journal of Cell Biology, 200(3):271–??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/3/271.

Wakana:2013:KEI

[WVvG^{+}13] Yuichi Wakana, Julien Villeneuve, Josse van Galen, David Cruz-Garcia, Mitsuo Tagaya, and Vivek Malhotra. Kinesin-5/Eg5 is important for transport of CARTS from the trans-Golgi network to the cell surface. Journal of Cell Biology, 202 (2):241–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/2/241.

Walter:2010:STB

[WWB^{+}10] Alexander M. Walter, Katrin Wiederhold, Dieter Bruns, Dirk Fasshauer, and Jakob B. Sørensen. Synaptobrevin Nterminally bound to syntaxin–SNAP–25 defines the primed vesicle state in regulated exocytosis. Journal of Cell Biology, 188(3):401–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/188/3/401.

Williamson:2010:DFV

[WWHH10] W. Ryan Williamson, Dong Wang, Adam S. Haberman, and P. Robin Hiesinger. A dual function of V0-ATPase a1 provides an endolysosomal degradation mechanism in Drosophila melanogaster photoreceptors. Journal of Cell Biology, 189 (5):885–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/5/885.

Wang:2012:TIP

[WWM^{+}12] Shujie Wang, Takashi Watanabe, Kenji Matsuzawa, Akira Katsumi, Mai Kakeno, Toshinori Matsui, Feng Ye, Kazuhide Sato, Kiyoko Murase, Ikuko Sugiyama, Kazushi Kimura, Akira Mizoguchi, Mark H. Ginsberg, John G. Collard, and Kozo Kaibuchi. Tiam1 interaction with the PAR complex promotes talin-mediated Rac1 activation during polarized cell migration. Journal of Cell Biology, 199(2):331–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/2/331.

Wynne:2012:RIA

[WWS^{+}12] Joseph P. Wynne, Jinhua Wu, Wenjuan Su, Adam Mor, Nikolaos Patsoukis, Vassiliki A. Boussiotis, Stevan R. Hubbard, and Mark R. Philips. Rap1-interacting adapter molecule (RIAM) associates with the plasma membrane via a proximity detector. Journal of Cell Biology, 199(2):317–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/199/2/317.

Winbanks:2012:FMS

[WWT^{+}12] Catherine E. Winbanks, Kate L. Weeks, Rachel E. Thomson, Patricio V. Sepulveda, Claudia Beyer, Hongwei Qian, Justin L. Chen, James M. Allen, Graeme I. Lancaster, Mark A. Febbraio, Craig A. Harrison, Julie R. McMullen, Jeffrey S. Chamberlain, and Paul Gregorevic. Follistatin-mediated skeletal muscle hypertrophy is regulated by Smad3 and mTOR independently of myostatin. Journal of Cell Biology, 197 (7):997–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/7/997.

West:2011:AYS

[WZHV11] Matt West, Nesia Zurek, Andreas Hoenger, and Gia K. Voeltz. A 3D analysis of yeast ER structure reveals how ER domains are organized by membrane curvature. Journal of Cell Biology, 193(2):333–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/2/333.

Xu:2013:PFE

[XBC^{+}13] Peng Xu, Ryan D. Baldridge, Richard J. Chi, Christopher G. Burd, and Todd R. Graham. Phosphatidylserine flipping enhances membrane curvature and negative charge required for vesicular transport. Journal of Cell Biology, 202(6):875– ??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/6/875.

Xiong:2012:MIR

[XG12] Yue Xiong and Kun-Liang Guan. Mechanistic insights into the regulation of metabolic enzymes by acetylation. Journal of Cell Biology, 198(2):155–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/198/2/155.

Xu:2010:APS

[XHB^{+}10] Naihan Xu, Nadia Hegarat, Elizabeth J. Black, Mary T. Scott, Helfrid Hochegger, and David A. Gillespie. Akt/ PKB suppresses DNA damage processing and checkpoint activation in late G2. Journal of Cell Biology, 190(3):297–??, August 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/190/
3/297.

Xu:2013:SMH

[XHS^{+}13] Qiaoqiao Xu, Shanshan Huang, Mingke Song, Chuan-En Wang, Sen Yan, Xudong Liu, Marta A. Gaertig, Shan Ping Yu, He Li, Shihua Li, and Xiao-Jiang Li. Synaptic mutant huntingtin inhibits synapsin-1 phosphorylation and causes neurological symptoms. Journal of Cell Biology, 202(7):1123– ??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/7/1123.

Xu:2010:IAB

[XOV^{+}10] Zhenjie Xu, Hiromi Ogawa, Paola Vagnarelli, Jan H. Bergmann, Damien F. Hudson, Sandrine Ruchaud, Tatsuo Fukagawa, William C. Earnshaw, and Kumiko Samejima. INCENP-aurora B interactions modulate kinase activity and chromosome passenger complex localization. Journal of Cell Biology, 188(4):611–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/4/611.

Xu:2011:DMI

[XRO^{+}11] Yingke Xu, Bradley R. Rubin, Charisse M. Orme, Alexander Karpikov, Chenfei Yu, Jonathan S. Bogan, and Derek K. Toomre. Dual-mode of insulin action controls GLUT4 vesicle exocytosis. Journal of Cell Biology, 193(4):643–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
4/643.

Xu:2010:PAR

[XSJ^{+}10] Ye Xu, Yingli Sun, Xiaofeng Jiang, Marina K. Ayrapetov, Patryk Moskwa, Shenghong Yang, David M. Weinstock, and Brendan D. Price. The p400 ATPase regulates nucleosome stability and chromatin ubiquitination during DNA repair. Journal of Cell Biology, 191(1):31–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/1/31.

Xu:2011:MRC

[XTH^{+}11] Dan Xu, Fumitaka Takeshita, Yumiko Hino, Saori Fukunaga, Yasusei Kudo, Aya Tamaki, Junko Matsunaga, Ryou u Takahashi, Takashi Takata, Akira Shimamoto, Takahiro Ochiya, and Hidetoshi Tahara. miR-22 represses cancer progression by inducing cellular senescence. Journal of Cell Biology, 193 (2):409–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/193/2/409.

Xia:2013:VLP

[XTX^{+}13] Wen-Fang Xia, Fu-Lei Tang, Lei Xiong, Shan Xiong, Ji-Ung Jung, Dae-Hoon Lee, Xing-Sheng Li, Xu Feng, Lin Mei, and Wen-Cheng Xiong. Vps35 loss promotes hyperresorptive osteoclastogenesis and osteoporosis via sustained RANKL signaling. Journal of Cell Biology, 200(6):821–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/200/6/821.

Xiang:2010:GGP

[XW10] Yi Xiang and Yanzhuang Wang. GRASP55 and GRASP65 play complementary and essential roles in Golgi cisternal stacking. Journal of Cell Biology, 188(2):237–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/188/
2/237.

Xiong:2010:PTW

[XWE^{+}10] Xin Xiong, Xin Wang, Ronny Ewanek, Pavan Bhat, Aaron DiAntonio, and Catherine A. Collins. Protein turnover of the Wallenda/DLK kinase regulates a retrograde response to axonal injury. Journal of Cell Biology, 191(1):211–??, October 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
1/211.

Xu:2010:NMV

[XYM^{+}10] Yunling Xu, Li Yuan, Judy Mak, Luc Pardanaud, Maresa Caunt, Ian Kasman, Bruno Larrivée, Raquel del Toro, Steven Suchting, Alexander Medvinsky, Jillian Silva, Jian Yang, Jean-Léon Thomas, Alexander W. Koch, Kari Alitalo, Anne Eichmann, and Anil Bagri. Neuropilin-2 mediates VEGF–C– induced lymphatic sprouting together with VEGFR3. Journal of Cell Biology, 188(1):115–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/188/1/115.

Xu:2012:FDC

[XZC^{+}12] Ningyi Xu, Shaobing O. Zhang, Ronald A. Cole, Sean A. McKinney, Fengli Guo, Joel T. Haas, Sudheer Bobba, Robert V. Farese, and Ho Yi Mak. The FATP1–DGAT2 complex facilitates lipid droplet expansion at the ER–lipid droplet interface. Journal of Cell Biology, 198(5):895–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
5/895.

Yamashita:2013:NTS

[Yam13] Yukiko M. Yamashita. Nonrandom template segregation: a way to break the symmetry of stem cells. Journal of Cell Biology, 203(1):7–??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/203/1/7.

Yonamine:2011:SKT

[YBN^{+}11] Ikuko Yonamine, Takeshi Bamba, Niraj K. Nirala, Nahid Jesmin, Teresa Kosakowska-Cholody, Kunio Nagashima, Eiichiro Fukusaki, Jairaj K. Acharya, and Usha Acharya. Sphingosine kinases and their metabolites modulate endolysosomal trafficking in photoreceptors. Journal of Cell Biology, 192(4): 557–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/4/557.

Yoon:2010:DPS

[YCP10] Je-Hyun Yoon, Eui-Ju Choi, and Roy Parker. Dcp2 phosphorylation by Ste20 modulates stress granule assembly and mRNA decay in Saccharomyces cerevisiae. Journal of Cell Biology, 189(5):813–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/5/813.

Yoon:2011:CIP

[YDB^{+}11] Mee-Sup Yoon, Guangwei Du, Jonathan M. Backer, Michael A. Frohman, and Jie Chen. Class III PI-3-kinase activates phospholipase D in an amino acid–sensing mTORC1 pathway. Journal of Cell Biology, 195(3):435–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/3/435.

Yoo:2012:ERS

[YFLH12] Sa Kan Yoo, Christina M. Freisinger, Danny C. LeBert, and Anna Huttenlocher. Early redox, Src family kinase, and calcium signaling integrate wound responses and tissue regeneration in zebrafish. Journal of Cell Biology, 199(2):225–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
199/2/225.

Yuan:2012:DCT

[YFO12] Kai Yuan, Jeffrey A. Farrell, and Patrick H. O’Farrell. Different cyclin types collaborate to reverse the S-phase checkpoint and permit prompt mitosis. Journal of Cell Biology, 198(6): 973–??, September 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/6/973.

Yagita:2013:TAP

[YHF13] Yuichi Yagita, Takahide Hiromasa, and Yukio Fujiki. Tailanchored PEX26 targets peroxisomes via a PEX19-dependent and TRC40-independent class I pathway. Journal of Cell Biology, 200(5):651–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/5/651.

Yamauchi:2014:PMA

[YHG^{+}14] Shota Yamauchi, Yan Yan Hou, Alvin Kunyao Guo, Hiroaki Hirata, Wataru Nakajima, Ai Kia Yip, Cheng han Yu, Ichiro Harada, Keng-Hwee Chiam, Yasuhiro Sawada, Nobuyuki Tanaka, and Keiko Kawauchi. p53-mediated activation of the mitochondrial protease HtrA2/Omi prevents cell invasion. Journal of Cell Biology, 204(7):1191–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/204/
7/1191.

Yamamoto:2010:DPP

[YHK10] Ryosuke Yamamoto, Masafumi Hirono, and Ritsu Kamiya. Discrete PIH proteins function in the cytoplasmic preassembly of different subsets of axonemal dyneins. Journal of Cell Biology, 190(1):65–??, July 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/190/1/65.

Ye:2010:RTE

[YHT^{+}10] Feng Ye, Guiqing Hu, Dianne Taylor, Boris Ratnikov, Andrey A. Bobkov, Mark A. McLean, Stephen G. Sligar, Kenneth A. Taylor, and Mark H. Ginsberg. Recreation of the terminal events in physiological integrin activation. Journal of Cell Biology, 188(1):157–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/1/157.

Yoshida:2013:MOC

[YKT^{+}13] Masashi Yoshida, Satoshi Katsuyama, Kazuki Tateho, Hiroto Nakamura, Junpei Miyoshi, Tatsunori Ohba, Hirotada Matsuhara, Futaba Miki, Koei Okazaki, Tokuko Haraguchi, Osami Niwa, Yasushi Hiraoka, and Ayumu Yamamoto. Microtubuleorganizing center formation at telomeres induces meiotic telomere clustering. Journal of Cell Biology, 200(4):385– ??, February 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/200/4/385.

Yamamoto:2012:AVI

[YKW^{+}12] Hayashi Yamamoto, Soichiro Kakuta, Tomonobu M. Watanabe, Akira Kitamura, Takayuki Sekito, Chika Kondo-Kakuta, Rie Ichikawa, Masataka Kinjo, and Yoshinori Ohsumi. Atg9 vesicles are an important membrane source during early steps of autophagosome formation. Journal of Cell Biology, 198 (2):219–??, July 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/198/2/219.

Lee:2013:ARL

[yLFAM13] Kyoo young Lee, Haiqing Fu, Mirit I. Aladjem, and Kyungjae Myung. ATAD5 regulates the lifespan of DNA replication factories by modulating PCNA level on the chromatin. Journal of Cell Biology, 200(1):31–??, January 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/200/1/31.

Yang:2010:FGF

[YMM^{+}10] Jingxuan Yang, Michael Meyer, Anna-Katharina Müller, Friederike Böhm, Richard Grose, Tina Dauwalder, Francois Verrey, Manfred Kopf, Juha Partanen, Wilhelm Bloch, David M. Ornitz, and Sabine Werner. Fibroblast growth factor receptors 1 and 2 in keratinocytes control the epidermal barrier and cutaneous homeostasis. Journal of Cell Biology, 188 (6):935–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/6/935.

Yano:2013:AMT

[YMT^{+}13] Tomoki Yano, Takeshi Matsui, Atsushi Tamura, Masami Uji, and Sachiko Tsukita. The association of microtubules with tight junctions is promoted by cingulin phosphorylation by AMPK. Journal of Cell Biology, 203(4):605–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/203/
4/605.

Yamamizu:2010:CNC

[YMU^{+}10] Kohei Yamamizu, Taichi Matsunaga, Hideki Uosaki, Hiroyuki Fukushima, Shiori Katayama, Mina Hiraoka-Kanie, Kohnosuke Mitani, and Jun K. Yamashita. Convergence of Notch and β-catenin signaling induces arterial fate in vascular progenitors. Journal of Cell Biology, 189(2):325–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/189/
2/325.

Yamamizu:2013:CNC

[YMU^{+}13] Kohei Yamamizu, Taichi Matsunaga, Hideki Uosaki, Hiroyuki Fukushima, Shiori Katayama, Mina Hiraoka-Kanie, Kohnosuke Mitani, and Jun K. Yamashita. Convergence of Notch and β-catenin signaling induces arterial fate in vascular progenitors. Journal of Cell Biology, 202(1):179–??, July 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/202/
1/179.

Yabuta:2011:TRR

[YOA^{+}11] Yukihiro Yabuta, Hiroshi Ohta, Takaya Abe, Kazuki Kurimoto, Shinichiro Chuma, and Mitinori Saitou. TDRD5 is required for retrotransposon silencing, chromatoid body assembly, and spermiogenesis in mice. Journal of Cell Biology, 192 (5):781–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/5/781.

Yi:2011:HRI

[YOMM^{+}11] Julie Y. Yi, Kassandra M. Ori-McKenney, Richard J. McKenney, Michael Vershinin, Steven P. Gross, and Richard B. Vallee. High-resolution imaging reveals indirect coordination of opposite motors and a role for LIS1 in high-load axonal transport. Journal of Cell Biology, 195(2):193–??, October 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
2/193.

Yajima:2012:CCT

[YON^{+}12] Hiroaki Yajima, Toshihiko Ogura, Ryo Nitta, Yasushi Okada, Chikara Sato, and Nobutaka Hirokawa. Conformational changes in tubulin in GMPCPP and GDP-taxol microtubules observed by cryoelectron microscopy. Journal of Cell Biology, 198(3):315–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/198/3/315.

Yi:2013:SAB

[YRU^{+}13] Kexi Yi, Boris Rubinstein, Jay R. Unruh, Fengli Guo, Brian D. Slaughter, and Rong Li. Sequential actin-based pushing forces drive meiosis I chromosome migration and symmetry breaking in oocytes. Journal of Cell Biology, 200(5):567–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/200/
5/567.

Yamamoto:2013:CCN

[YSaY^{+}13] Ryosuke Yamamoto, Kangkang Song, Haru aki Yanagisawa, Laura Fox, Toshiki Yagi, Maureen Wirschell, Masafumi Hirono, Ritsu Kamiya, Daniela Nicastro, and Winfield S. Sale. The MIA complex is a conserved and novel dynein regulator essential for normal ciliary motility. Journal of Cell Biology, 201 (2):263–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/201/2/263.

Yu:2010:CRS

[YSM10] Fang Yu, Lu Sun, and Khaled Machaca. Constitutive recycling of the store-operated Ca^{2+} channel Orai1 and its internalization during meiosis. Journal of Cell Biology, 191(3): 523–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/191/3/523.

Yen:2010:COG

[YSN^{+}10] Wei-Lien Yen, Takahiro Shintani, Usha Nair, Yang Cao, Brian C. Richardson, Zhijian Li, Frederick M. Hughson, Misuzu Baba, and Daniel J. Klionsky. The conserved oligomeric Golgi complex is involved in double-membrane vesicle formation during autophagy. Journal of Cell Biology, 188(1):

101–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/1/101.

Yamaguchi:2011:LIA

[YSN^{+}11] Yoshifumi Yamaguchi, Naomi Shinotsuka, Keiko Nonomura, Kiwamu Takemoto, Keisuke Kuida, Hiroki Yosida, and Masayuki Miura. Live imaging of apoptosis in a novel transgenic mouse highlights its role in neural tube closure. Journal of Cell Biology, 195(6):1047–??, December 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/195/6/1047.

Yamashita:2011:MRC

[YSO^{+}11] Daisuke Yamashita, Keishi Shintomi, Takao Ono, Ioannis Gavvovidis, Detlev Schindler, Heidemarie Neitzel, Marc Trimborn, and Tatsuya Hirano. MCPH1 regulates chromosome condensation and shaping as a composite modulator of condensin II. Journal of Cell Biology, 194(6):841–??, September 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/194/
6/841.

Yu:2011:CTC

[YST^{+}11] Min Yu, Shannon Stott, Mehmet Toner, Shyamala Maheswaran, and Daniel A. Haber. Circulating tumor cells: approaches to isolation and characterization. Journal of Cell Biology, 192(3):373–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/3/373.

Yan:2011:RPP

[YTM^{+}11] Hong Yan, Thomas Toczylowski, Jill McCane, Chinyi Chen, and Shuren Liao. Replication protein A promotes 5′ → 3′ end processing during homology-dependent DNA double-strand break repair. Journal of Cell Biology, 192(2):251–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/192/
2/251.

Yan:2010:SMP

[YTT^{+}10] Rihui Yan, Sharon E. Thomas, Jui-He Tsai, Yukihiro Yamada, and Bruce D. McKee. SOLO: a meiotic protein required for centromere cohesion, coorientation, and SMC1 localization in Drosophila melanogaster. Journal of Cell Biology, 188(3): 335–??, February 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/3/335.

Yi:2013:CRC

[YWC^{+}13] Jason Yi, Xufeng Wu, Andrew H. Chung, James K. Chen, Tarun M. Kapoor, and John A. Hammer. Centrosome repositioning in T cells is biphasic and driven by microtubule end-on capture-shrinkage. Journal of Cell Biology, 202(5): 779–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/202/5/779.

Yamamoto:2012:KDA

[YWJ^{+}12] Kenta Yamamoto, Yunyue Wang, Wenxia Jiang, Xiangyu Liu, Richard L. Dubois, Chyuan-Sheng Lin, Thomas Ludwig, Christopher J. Bakkenist, and Shan Zha. Kinase-dead ATM protein causes genomic instability and early embryonic lethality in mice. Journal of Cell Biology, 198(3):305–??, August 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/198/
3/305.

Yano:2011:TRC

[YYA^{+}11] Tomoki Yano, Yuji Yamazaki, Makoto Adachi, Katsuya Okawa, Philippe Fort, Masami Uji, Shoichiro Tsukita, and Sachiko Tsukita. Tara up-regulates E-cadherin transcription by binding to the Trio RhoGEF and inhibiting Rac signaling. Journal of Cell Biology, 193(2):319–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/2/319.

Yamaguchi:2011:PKS

[YYM^{+}11] Hideki Yamaguchi, Shuhei Yoshida, Emi Muroi, Nachi Yoshida, Masahiro Kawamura, Zen Kouchi, Yoshikazu Nakamura, Ryuichi Sakai, and Kiyoko Fukami. Phosphoinositide 3-kinase signaling pathway mediated by p110α regulates invadopodia formation. Journal of Cell Biology, 193(7):1275–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
193/7/1275.

Yasuda:2013:RBP

[YZL^{+}13] Kyota Yasuda, Huaye Zhang, David Loiselle, Timothy Haystead, Ian G. Macara, and Stavroula Mili. The RNAbinding protein Fus directs translation of localized mRNAs in APC–RNP granules. Journal of Cell Biology, 203(5):737– ??, December 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/5/737.

Yu:2012:WCD

[YZM^{+}12a] Xinzi Yu, Tobias Zech, Laura McDonald, Esther Garcia Gonzalez, Ang Li, Iain Macpherson, Juliane P. Schwarz, Heather Spence, Kinga Futó, Paul Timpson, Colin Nixon, Yafeng Ma, Ines M. Anton, Balázs Visegrády, Robert H. Insall, Karin Oien, Karen Blyth, Jim C. Norman, and Laura M. Machesky. N-WASP coordinates the delivery and F-actin-mediated capture of MT1-MMP at invasive pseudopods. Journal of Cell Biology, 199(3):527–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/199/3/527.

Yu:2012:SSI

[YZM^{+}12b] Yamei Yu, Jianghai Zhu, Li-Zhi Mi, Thomas Walz, Hao Sun, JianFeng Chen, and Timothy A. Springer. Structural specializations of α_{4}β_{7}, an integrin that mediates rolling adhesion. Journal of Cell Biology, 196(1):131–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/196/1/131.

Yang:2012:ACD

[YZPF12] Qing Yang, Xiao-Feng Zhang, Thomas D. Pollard, and Paul Forscher. Arp2/ 3 complex-dependent actin networks constrain myosin II function in driving retrograde actin flow. Journal of Cell Biology, 197(7):939–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/197/7/939.

Zeng:2010:PPR

[ZBBG10] Kang Zeng, Ricardo Nunes Bastos, Francis A. Barr, and Ulrike Gruneberg. Protein phosphatase 6 regulates mitotic spindle formation by controlling the T-loop phosphorylation state of Aurora A bound to its activator TPX2. Journal of Cell Biology, 191(7):1315–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/191/7/1315.

Zaidel-Bar:2010:FBD

[ZBJL^{+}10] Ronen Zaidel-Bar, Michael J. Joyce, Allison M. Lynch, Kristen Witte, Anjon Audhya, and Jeff Hardin. The F–BAR domain of SRGP-1 facilitates cell–cell adhesion during C. elegans morphogenesis. Journal of Cell Biology, 191(4):761–??, November 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/191/
4/761.

Zhou:2011:NIM

[ZC11] Shan Zhou and Lihsia Chen. Neural integrity is maintained by dystrophin in C. elegans. Journal of Cell Biology, 192(2): 349–??, January 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/2/349.

Zito:2010:EPSa

[ZCB^{+}10a] Ester Zito, King-Tung Chin, Jaime Blais, Heather P. Harding, and David Ron. ERO1-β, a pancreas-specific disulfide oxidase, promotes insulin biogenesis and glucose homeostasis. Journal of Cell Biology, 188(6):821–??, March 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/188/6/821.

Zito:2010:EPSb

[ZCB^{+}10b] Ester Zito, King-Tung Chin, Jaime Blais, Heather P. Harding, and David Ron. ERO1-β, a pancreas-specific disulfide oxidase, promotes insulin biogenesis and glucose homeostasis. Journal of Cell Biology, 189(4):769–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/189/4/769.

Zapata:2014:PMR

[ZDM^{+}14] Jessica Zapata, Noah Dephoure, Tracy MacDonough, Yaxin Yu, Emily J. Parnell, Meghan Mooring, Steven P. Gygi, David J. Stillman, and Douglas R. Kellogg. PP2A^{Rts1} is a master regulator of pathways that control cell size. Journal of Cell Biology, 204(3):359–??, February 2014. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/3/359.

Zhang:2012:IPS

[ZDS^{+}12] Liang Zhang, Priyabrata Das, Mirco Schmolke, Balaji Manicassamy, Yaming Wang, Xiaoyi Deng, Ling Cai, Benjamin P. Tu, Christian V. Forst, Michael G. Roth, David E. Levy, Adolfo García-Sastre, Jef de Brabander, Margaret A. Phillips, and Beatriz M. A. Fontoura. Inhibition of pyrimidine synthesis reverses viral virulence factor-mediated block of mRNA nuclear export. Journal of Cell Biology, 196(3):315–??, February 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/196/
3/315.

Zyss:2011:CKD

[ZEG11] Deborah Zyss, Hani Ebrahimi, and Fanni Gergely. Casein kinase I delta controls centrosome positioning during T cell activation. Journal of Cell Biology, 195(5):781–??, November 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/195/
5/781.

Zimmer:2011:PCO

[ZF11] Christophe Zimmer and Emmanuelle Fabre. Principles of chromosomal organization: lessons from yeast. Journal of Cell Biology, 192(5):723–??, March 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/192/5/723.

Zhou:2013:RGC

[ZFA^{+}13] Xuan Zhou, Maria Carolina Florian, Paritha Arumugam, Xiaoyi Chen, Jose A. Cancelas, Richard Lang, Punam Malik, Hartmut Geiger, and Yi Zheng. RhoA GTPase controls cytokinesis and programmed necrosis of hematopoietic progenitors. Journal of Cell Biology, 203(1):??, October 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/1/2031OIA113.

Zhang:2013:ARC

[ZFP^{+}13] Zhao Zhang, Jing Feng, Chenyu Pan, Xiangdong Lv, Wenqing Wu, Zhaocai Zhou, Feng Liu, Lei Zhang, and Yun Zhao.

Atrophin–Rpd3 complex represses Hedgehog signaling by acting as a corepressor of Ci^{R}. Journal of Cell Biology, 203(4): 575–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/203/4/575.

Zambrano:2014:THR

[ZGCG^{+}14] Alberto Zambrano, Verónica García-Carpizo, María Esther Gallardo, Raquel Villamuera, Maria Ana Gómez-Ferrería, Angel Pascual, Nicolas Buisine, Laurent M. Sachs, Rafael Garesse, and Ana Aranda. The thyroid hormone receptor β induces DNA damage and premature senescence. Journal of Cell Biology, 204(1):129–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/1/129.

Zhou:2012:NPS

[ZGEM12] Xiao Zhou, Katja Graumann, David E. Evans, and Iris Meier. Novel plant SUN–KASH bridges are involved in RanGAP anchoring and nuclear shape determination. Journal of Cell Biology, 196(2):203–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/196/2/203.

Zhou:2014:IUS

[ZGW^{+}14] Xiao Zhou, Katja Graumann, Lennart Wirthmueller, Jonathan D. G. Jones, and Iris Meier. Identification of unique SUNinteracting nuclear envelope proteins with diverse functions in plants. Journal of Cell Biology, 205(5):677–??, June 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/205/5/
677.

Zhang:2012:RRR

[ZIG^{+}12] Donglei Zhang, Nora R. Isack, Doreen R. Glodowski, Jie Liu, Carlos Chih-Hsiung Chen, X. Z. Shawn Xu, Barth D. Grant, and Christopher Rongo. RAB-6.2 and the retromer regulate glutamate receptor recycling through a retrograde pathway. Journal of Cell Biology, 196(1):85–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/196/1/85.

Zheng:2010:HHP

[ZJP^{+}10] Yupeng Zheng, Sam John, James J. Pesavento, Jennifer R. Schultz-Norton, R. Louis Schiltz, Sonjoon Baek, Ann M. Nardulli, Gordon L. Hager, Neil L. Kelleher, and Craig A. Mizzen. Histone H1 phosphorylation is associated with transcription by RNA polymerases I and II. Journal of Cell Biology, 189 (3):407–??, May 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/3/407.

Zanet:2012:FPF

[ZJP^{+}12] Jennifer Zanet, Asier Jayo, Serge Plaza, Tom Millard, Maddy Parsons, and Brian Stramer. Fascin promotes filopodia formation independent of its role in actin bundling. Journal of Cell Biology, 197(4):477–??, May 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/197/4/477.

Zylkiewicz:2011:TCC

[ŻKC^{+}11] Eliza Żyłkiewicz, Monika Kijańska, Won-Chan Choi, Urszula Derewenda, Zygmunt S. Derewenda, and P. Todd Stukenberg. The N-terminal coiled-coil of Ndel1 is a regulated scaffold that recruits LIS1 to dynein. Journal of Cell Biology, 192(3): 433–??, February 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/192/3/433.

Zuleger:2011:SAS

[ZKR^{+}11] Nikolaj Zuleger, David A. Kelly, A. Christine Richardson, Alastair R. W. Kerr, Martin W. Goldberg, Andrew B. Goryachev, and Eric C. Schirmer. System analysis shows distinct mechanisms and common principles of nuclear envelope protein dynamics. Journal of Cell Biology, 193(1):109–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
1/109.

Zhu:2013:SFA

[ZKW^{+}13] Qingwei Zhu, Yong Hwan Kim, Douglas Wang, S. Paul Oh, and Kunxin Luo. SnoN facilitates ALK1-Smad1/5 signaling during embryonic angiogenesis. Journal of Cell Biology, 202(6):937–??, September 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/202/6/937.

Zeman:2014:DDS

[ZLFC14] Michelle K. Zeman, Jia-Ren Lin, Raimundo Freire, and Karlene A. Cimprich. DNA damage-specific deubiquitination regulates Rad18 functions to suppress mutagenesis. Journal of Cell Biology, 206(2):183–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/206/2/183.

Zhu:2014:ARR

[ZLH^{+}14] Ying-Ting Zhu, Fu Li, Bo Han, Sean Tighe, Suzhen Zhang, Szu-Yu Chen, Xin Liu, and Scheffer C. G. Tseng. Activation of RhoA–ROCK–BMP signaling reprograms adult human corneal endothelial cells. Journal of Cell Biology, 206(6): 799–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/6/799.

Zhang:2013:CHS

[ZLJ^{+}13] Ke Zhang, Zhihong Li, Manish Jaiswal, Vafa Bayat, Bo Xiong, Hector Sandoval, Wu-Lin Charng, Gabriela David, Claire Haueter, Shinya Yamamoto, Brett H. Graham, and Hugo J. Bellen. The C8ORF38 homologue Sicily is a cytosolic chaperone for a mitochondrial complex I subunit. Journal of Cell Biology, 200(6):807–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/6/807.

Zeng:2013:HPP

[ZLW^{+}13] An Zeng, Yong-Qin Li, Chen Wang, Xiao-Shuai Han, Ge Li, Jian-Yong Wang, Dang-Sheng Li, Yong-Wen Qin, Yufang Shi, Gary Brewer, and Qing Jing. Heterochromatin protein 1 promotes self-renewal and triggers regenerative proliferation in adult stem cells. Journal of Cell Biology, 201(3):409–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/201/
3/409.

Zihni:2014:DDC

[ZME^{+}14] Ceniz Zihni, Peter M. G. Munro, Ahmed Elbediwy, Nicholas H. Keep, Stephen J. Terry, John Harris, Maria S. Balda, and Karl Matter. Dbl3 drives Cdc42 signaling at the apical margin to regulate junction position and apical differentiation. Journal of Cell Biology, 204(1):111–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/1/111.

Zou:2013:CPR

[ZMW^{+}13] Yongxin Zou, Jun Mi, Wenxing Wang, Juanjuan Lu, Wei Zhao, Zhaojian Liu, Huili Hu, Yang Yang, Xiaoxing Gao, Baichun Jiang, Changshun Shao, and Yaoqin Gong. CUL4B promotes replication licensing by up-regulating the CDK2– CDC6 cascade. Journal of Cell Biology, 200(6):743–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/200/
6/743.

Ziani:2014:SOA

[ZNA^{+}14] Salim Ziani, Zita Nagy, Sergey Alekseev, Evi Soutoglou, Jean-Marc Egly, and Frédéric Coin. Sequential and ordered assembly of a large DNA repair complex on undamaged chromatin. Journal of Cell Biology, 206(5):589–??, September 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/206/
5/589.

Zhang:2011:IIL

[ZNH^{+}11] He Zhang, Beibei Niu, Ji-Fan Hu, Shengfang Ge, Haibo Wang, Tao Li, Jianqun Ling, Brandon N. Steelman, Guanxiang Qian, and Andrew R. Hoffman. Interruption of intrachromosomal looping by CCCTC binding factor decoy proteins abrogates genomic imprinting of human insulin-like growth factor II. Journal of Cell Biology, 193(3):475–??, May 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/193/3/475.

Zhong:2013:LOF

[ZNP^{+}13] Yuan Zhong, Tittu Nellimoottil, Jared M. Peace, Simon R. V. Knott, Sandra K. Villwock, Janis M. Yee, Jeffrey M. Jancuska, Sanket Rege, Marianne Tecklenburg, Robert A. Sclafani, Simon Tavaré, and Oscar M. Aparicio. The level of origin firing inversely affects the rate of replication fork progression. Journal of Cell Biology, 201(3):373–??, April 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/201/3/373.

Zylbersztejn:2012:VSS

[ZPB^{+}12] Kathleen Zylbersztejn, Maja Petkovic, Andrea Burgo, Marie Deck, Sonia Garel, Séverine Marcos, Evelyne Bloch-Gallego, Fatiha Nothias, Guido Serini, Dominique Bagnard, Thomas Binz, and Thierry Galli. The vesicular SNARE Synaptobrevin is required for Semaphorin 3A axonal repulsion. Journal of Cell Biology, 196(1):37–??, January 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/196/1/37.

Zhao:2010:RKI

[ZPS^{+}10] Xiaofeng Zhao, Xu Peng, Shaogang Sun, Ann Y. J. Park, and Jun-Lin Guan. Role of kinase-independent and -dependent functions of FAK in endothelial cell survival and barrier function during embryonic development. Journal of Cell Biology, 189(6):955–??, June 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/189/6/955.

Zhang:2014:HND

[ZQA^{+}14] Jun Zhang, Rongde Qiu, Herbert N. Arst, Miguel A. Peñalva, and Xin Xiang. HookA is a novel dynein–early endosome linker critical for cargo movement in vivo. Journal of Cell Biology, 204(6):1009–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/204/6/1009.

Zaytsev:2014:APK

[ZSD^{+}14] Anatoly V. Zaytsev, Lynsie J. R. Sundin, Keith F. DeLuca, Ekaterina L. Grishchuk, and Jennifer G. DeLuca. Accurate phosphoregulation of kinetochore–microtubule affinity requires unconstrained molecular interactions. Journal of Cell Biology, 206(1):45–??, July 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/206/1/45.

Zhang:2010:ACM

[ZSH10] Keman Zhang, Jingfeng Sha, and Marian L. Harter. Activation of Cdc6 by MyoD is associated with the expansion of quiescent myogenic satellite cells. Journal of Cell Biology, 188 (1):39–??, January 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/188/1/39.

Zhang:2012:ADR

[ZSK12] Fangliang Zhang, Sougata Saha, and Anna Kashina. Arginylationdependent regulation of a proteolytic product of talin is essential for cell–cell adhesion. Journal of Cell Biology, 197 (6):819–??, June 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/197/6/819.

Zhu:2013:MNP

[ZSK^{+}13] Mei Zhu, Florian Settele, Sachin Kotak, Luis Sanchez-Pulido, Lena Ehret, Chris P. Ponting, Pierre Gönczy, and Ingrid Hoffmann. MISP is a novel Plk1 substrate required for proper spindle orientation and mitotic progression. Journal of Cell Biology, 200(6):773–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/6/773.

Zhang:2013:MDO

[ZSZ^{+}13] Chuansheng Zhang, Keiichiro Susuki, Daniel R. Zollinger, Jeffrey L. Dupree, and Matthew N. Rasband. Membrane domain organization of myelinated axons requires βII spectrin. Journal of Cell Biology, 203(3):437–??, November 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/203/3/437.

Zatulovskiy:2014:BDC

[ZTBK14] Evgeny Zatulovskiy, Richard Tyson, Till Bretschneider, and Robert R. Kay. Bleb-driven chemotaxis of Dictyostelium cells. Journal of Cell Biology, 204(6):1027–??, March 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/204/6/1027.

Zhang:2014:THR

[ZWL^{+}14] Duo Zhang, Xiaoyun Wang, Yuying Li, Lei Zhao, Minghua Lu, Xuan Yao, Hongfeng Xia, Yu cheng Wang, Mo-Fang Liu, Jingjing Jiang, Xihua Li, and Hao Ying. Thyroid hormone regulates muscle fiber type conversion via miR-133a1. Journal of Cell Biology, 207(6):753–??, December 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/207/6/753.

Zhang:2011:PSD

[ZYF^{+}11] Jun Zhang, Xuanli Yao, Lauren Fischer, Juan F. Abenza, Miguel A. Peñalva, and Xin Xiang. The p25 subunit of the dynactin complex is required for dynein–early endosome interaction. Journal of Cell Biology, 193(7):1245–??, June 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/193/
7/1245.

Zhang:2011:CBM

[ZYH^{+}11] Shichuan Zhang, Hirohiko Yajima, HoangDinh Huynh, Junke Zheng, Elsa Callen, Hua-Tang Chen, Nancy Wong, Samuel Bunting, Yu-Fen Lin, Mengxia Li, Kyung-Jone Lee, Michael Story, Eric Gapud, Barry P. Sleckman, André Nussenzweig, Cheng Cheng Zhang, David J. Chen, and Benjamin P. C. Chen. Congenital bone marrow failure in DNA–PKcs mutant mice associated with deficiencies in DNA repair. Journal of Cell Biology, 193(2):295–??, April 2011. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/193/2/295.

Zhu:2013:CIH

[ZZS13] Jieqing Zhu, Jianghai Zhu, and Timothy A. Springer. Complete integrin headpiece opening in eight steps. Journal of Cell Biology, 201(7):1053–??, June 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/201/7/1053.

Zheng:2010:LRM

[ZZW^{+}10] Zhen Zheng, Huabin Zhu, Qingwen Wan, Jing Liu, Zhuoni Xiao, David P. Siderovski, and Quansheng Du. LGN regulates mitotic spindle orientation during epithelial morphogenesis. Journal of Cell Biology, 189(2):275–??, April 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/189/2/275.

Zhao:2013:MDE

[ZZW^{+}13] Hongyu Zhao, Yan G. Zhao, Xingwei Wang, Lanjun Xu, Lin Miao, Du Feng, Quan Chen, Attila L. Kovács, Dongsheng Fan, and Hong Zhang. Mice deficient in Epg5 exhibit selective neuronal vulnerability to degeneration. Journal of Cell Biology, 200(6):731–??, March 2013. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/200/6/731.

Zhang:2014:LNR

[ZZW^{+}14] He Zhang, Michael J. Zeitz, Hong Wang, Beibei Niu, Shengfang Ge, Wei Li, Jiuwei Cui, Guanjun Wang, Guanxiang Qian, Michael J. Higgins, Xianqun Fan, Andrew R. Hoffman, and Ji-Fan Hu. Long noncoding RNA-mediated intrachromosomal interactions promote imprinting at the Kcnq1 locus. Journal of Cell Biology, 204(1):61–??, January 2014. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/204/1/61.